A Complete Bibliography of Publications in Future Generation Computer Systems: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

11 December 2023 Version 3.02

Title word cross-reference

(L, α) [SLW11]. 2 [AKB18b, AM19a]. 3 [AKB18b, Bro19, CPD^{+}15, DJH^{+}19, GWW^{+}19, SHN10, VF18, XJY^{+}18]. 5 [AT19a, DPK^{+}19, NLS19, YYS^{+}19]. 80/20 [MRH17]. ^{+} [PYH^{+}18]. ^{2} [WWT^{+}16]. ^{3} [LHM14]. _{3} [TMDZ15]. B [WCWC19, WCWC20]. c [Bu18, LSZ^{+}16]. ϵ [GPS13]. K [LZY^{+}19b, PLL^{+}18, ZLT^{+}19, ZZC14, CCJ16, MLW^{+}18a, PSLZ18, STA17a, TDBR18]. M [YH18]. N [RW18, ZXM^{+}19, CFL^{+}18]. Q [CLL^{+}18a]. R^{2} [TY11]. T^{2}K^{2} [TDBR18]. T^{2}K^{2}D^{2} [TDBR18]. τ [RGDML16]. u [HZM14].

* [Vin16]. *AIDA [WM14].

-anonymity [ZLT^{+}19]. -aware [LWS^{+}12]. -based [YZ12, TMDZ15]. -constraint [GPS13]. -core [PSLZ18, PLL^{+}18]. -D [AM19a, VF18]. -dimensional [CFL^{+}18]. -diversity [SLW11]. -driven [HB19]. -fuzzifying [YH18]. -gram [ZXM^{+}19]. -Learning [CLL^{+}18a]. -Lop [RGDML16]. -means [Bu18, LSZ^{+}16, LZY^{+}19b]. -pillbox [HZM14].

0th [RCMT18].

1 [PZY16]. 100

[KSW^{+}13, SME^{+}21, WLRL18]. 108

[DP21a].

2 [MK16b, PZY17]. 2.0 [GHLW18, LLMP13, MDA^{+}19]. 2.5 [XTL^{+}19]. 2009 [MBC^{+}11, RW13]. 2010 [CC11]. 2012 [TCG14]. 2019 [Ano19q, Ano19n, Ano19p, Ano19o, Ano19r, Ano19m, Ano19k, Ano19l]. 234Compositor [NOF18]. 2K [HSP^{+}13, SKF^{+}11].

3 [LC17]. 3.0 [VEET18]. 37C [Cha14b].

4.0 [ZWJ^{+}19a]. 4.0-based [FG18]. 4CaaSt [MVG^{+}14]. 4K [HKU^{+}11, HSP^{+}13, KSK^{+}11, MBC^{+}11, SKF^{+}11]. 4K/2K [SKF^{+}11].

5G [ELAEAVAM19, HHZ19].

6 [BBvdB^{+}11]. 60p [KSK^{+}11]. 61850 [YS16]. 68 [BFS^{+}17a].

72 [HYS18].

802.11af [AAQ^{+}19]. 86 [LBJ^{+}24]. 88 [GHEB^{+}23, NDZ^{+}18a, NDZ^{+}19]. 89 [ABMMC22]. 8K [KSK^{+}11].

90 [AB19a]. 91 [Bo20, DP20, DP21a, DP21b, ZMZ^{+}20]. 98 [HZX^{+}20, JLC^{+}20, WWP20, WCWC20, YWG^{+}20, YTQ20, wZcZN^{+}20].

AAA [MLM16]. AAA-based [MLM16]. AAL [AMR^{+}19]. aaService [GMP^{+}17]. abbreviation [DZLA19]. ABC [JFZL17, XLW^{+}17]. ABE [HQZH14, ZSW^{+}18b]. Abnormal [LSL^{+}15, CXZC18, LWR^{+}19, RJN^{+}19, WLZ^{+}14, ZCZ^{+}18]. abnormality [GRS^{+}19]. absolute [HKP10]. Abstract [GGC17, FTD17, YJL^{+}19]. abstracted [AAD^{+}13]. abstraction [AFS16]. abuse [JSMG18, QRW^{+}18]. AC [HMW14]. accelerate [RGCCL18]. accelerated [FBM19, HSP^{+}13, LZX16]. Accelerating [CLDC19, FRB^{+}14, GXW^{+}19, XFJ^{+}19, KKL11]. acceleration [ABF^{+}15a, HZDS19, XZL^{+}19]. Acceptance [SAPA17]. Access [AMSPL19, BL13, DCMB15, HRR^{+}14, MDA^{+}19, ASAA18, AMHJ10, BCN^{+}19, BR18, CWL^{+}19, CJK^{+}18, CCM^{+}14, CdRRdCB19, DW11, DLDTGMMP16, FXG^{+}19, FNA11, FS18, FSM^{+}18b, GHMX10, HAJ^{+}19, HZL18a, HLCL16, KRD^{+}19, KIS11, LLMP13, LZXW13, LHO17, LLW^{+}18a, LLW^{+}19a, LZLL18b, MAC14, MYHZ18, MYBMM18, MLM16, Mer13, NRV^{+}17, NJ16, NA19, PFRC16, QGT^{+}18, QCX18, RPH19, SAK19, SMSF18, SCL14, SYK^{+}17, SZR18, VPP^{+}19, Wan18a, WG13, XYML19, YXZG18a, YWJ^{+}19, YAX^{+}18, ZZ15, ZCL^{+}18, ZDW^{+}16, ZFH^{+}18, ZSL^{+}19b]. access-right [RPH19]. accessibility [RMSPP17]. Accessing [CLH10, YSC^{+}15, YCY10]. accident [PWP^{+}18]. account [WNR19]. accountability [HCL^{+}17]. accountable [Wan18b, XX14]. accounting [BBC^{+}12]. accounts [LZP^{+}18]. accuracy [DSM^{+}19]. Accurate [TMB^{+}19, CSL19, FZT^{+}18, FWB13a, FWB13b, IDM^{+}16, MAPA19, NS17b, WN10]. Accurately [CPSD18]. ACEIS [GACM17]. Achieving [DW11, GLD^{+}19b, JLC18, KHG13, TSWL17, VPP^{+}19, XX14, ZZX^{+}19, NJHT11, WCL^{+}17a, WHS^{+}18]. ACID [KJI11]. Acknowledgement [Ano10, Ano11a, Ano12a]. ACM [KZ17]. ACO [GPJC17]. ACO-based [GPJC17]. Acoustic [KWK^{+}18, CJG^{+}18, FZW^{+}18, HAAWH^{+}18, HST^{+}18, TSD18, WCB^{+}18, WTP^{+}13]. acquaintance [RQN^{+}19]. acquisition [AK19, BDZ13, LLS^{+}19, XYLZ18]. across [AM19b, CTVB12, LZP^{+}18, LLM^{+}16, LSMVML13, SG17, SFR15, TMMVL12, UNM^{+}16, WTR^{+}13, ZZH^{+}16, dlFVPSHL^{+}14, SMSF18]. Action [UMUB19]. actions [ABG18, SSC^{+}19].

Activation [TWG^{+}19, SSZ^{+}17, ZBL^{+}14]. Active [KMI11, SFR15, STC15, BTP19, FGG13, LQK^{+}16, SYT^{+}19, XFLL16]. ActiveSort [LQK^{+}16]. activities [TMB^{+}19, dlFVPSHL^{+}14]. Activity [BAKB19, LTC^{+}19, TCCW19, BDE17, CHWW13, CZXL18, GMP^{+}17, HUMA18, HLT^{+}18, HMMW19, IFD^{+}19, KSS19, KKP19, MSS^{+}13, OCW14, RM19, dSGD13, GMP^{+}17]. Activity-aaService [GMP^{+}17]. Activity-based [BAKB19]. activity-travel [BAKB19]. ActOn [XCGD10]. ActOn-based [XCGD10]. actor [LYW^{+}16]. Actuation [SST18]. actuator [ARSMY19]. Ad [CNP^{+}19, KKN18, KIAD17, LLYW19, SVK19, YFY^{+}13, AAS^{+}19, BLMU19, HHK18, LAQ^{+}19, LLJ^{+}11, SGGCR^{+}16, VCD^{+}18, ZF16]. Ad-hoc [CNP^{+}19, LLYW19, SVK19, YFY^{+}13, AAS^{+}19, BLMU19, HHK18, LAQ^{+}19]. AdaBoost [LLS^{+}14]. ADAMAS [RM16]. adaptability [HRVW18]. adaptable [MSE19, PMLVLS^{+}13]. adaptation [AKM18, CLNR18, FTD17, FA11b, PWB^{+}13, PSBB15, SJ18]. adaptations [Kyr19]. adapter [LZL^{+}16]. Adapting [SPR^{+}10, SJV12, EKSDN19, JLRS18]. Adaptive [AS14, BML18, CJG^{+}18, DBP19, DP17, EP12, HCMJ19, Ima19, IDCJ11, JGB19, KPJ19, LZY^{+}19a, LJ17a, MCJ19, MBS13, MDB^{+}18b, MYK16, PIP18a, QPTGG^{+}12, RT16, SME^{+}21, SRdlPG19, SVK19, TPBS14, UGBM^{+}17, WLA18a, YMD^{+}13, AEME^{+}18, CZT^{+}15, CQW^{+}19, CPD^{+}15, CFMC19, CLL18b, CXC^{+}18, CB10, DST10, DFG^{+}19, DDD18, DFRW17, DHL18, DLS^{+}12, ESW^{+}17, EKGS14, FTH16, HGG^{+}14, HAP15, HXC^{+}18, IKLL12, JNR12, JSZ^{+}19, KKB14, KJI11, KK10b, KS18c, LLpC12, LBD^{+}19, LHM14, LSL^{+}18, LW18b, LKJ17, LSG^{+}19, MWCK19, MdFTGM19, MJRM16, MAA^{+}19, NDZ^{+}18a, NDZ^{+}18b, NDZ^{+}19, NQQL13, PdAF12, PNZ14, PPB16, RWO^{+}19, RWZ^{+}19, SME^{+}19, SMG18, SEPV19, SYL18, TJZ^{+}15, WCF^{+}15, WXZ^{+}18b, XLL^{+}14, XWRZ19, XJZ^{+}19, YP12, YDD^{+}18, ZCK^{+}15, ZTKX19, ZBL^{+}14, ZAC^{+}18, vWMBS14]. Adaptively [YPCK12, CLH10, JDW^{+}14]. address [ABF^{+}15a, BCdV^{+}19, GFD14]. Addressing [HNCJ13, ZL12, BFS^{+}17a, BFS^{+}17b]. adequacy [CMS^{+}18]. adjusting [Lea13]. adjustment [TJ18]. ADL [Bae14]. administration [SB19a]. admissible [QPTGG^{+}12]. admission [Che13a, IS18, MWPVB12]. adoption [CKR16, NZOCJ^{+}19]. ADSP [SVK19]. adult [WNR19]. adults [CAS^{+}18]. Advance [CJK^{+}18, DVB14, CKP^{+}19, TCCC11]. Advanced [BRB19a, CSYY18, DO15, GCCPGBGS10, PPS^{+}19, PPB16, PZY16, PZY17, VLAC^{+}13, ZZLR18, ALL^{+}18, Cuz14, DFRW17, GHP^{+}18, MLC^{+}11, MVG^{+}14, MFL18, RAdARP19, RRH16, WdL16, YCH19, VOV17]. Advances [CLCY18, DPDS14, FBS18, KGVW14, Pal16, TCG14, WQ14, ADLW12, BB12, CSC18, HMS15, KA13, MGL^{+}18, YHA^{+}19, ZCYZ18]. advancing [GDP^{+}18]. Advantages [SW17]. adversary [NAM^{+}19]. advertisement [AMHJ10]. Advertisements [BCMM18, NK18]. advisor [CRTN17]. advisory [BDE17]. aerial [MND^{+}19]. aerosol [XAW^{+}10]. AES [FAA^{+}18]. affairs [LLW^{+}18b]. affect [PKA19]. affect-related [PKA19]. affecting [DSM^{+}19]. Affective [AW19, NPH19, XFM16, FGW^{+}19, LCC19, LJ19a, NKB19, QLM^{+}18, SA19]. affinities [HAF^{+}16]. AFIRM [MDB^{+}18b]. AFT [PIP18a]. after [ZXW19]. Against [SCH^{+}17, IDKD19, JL14, JCL^{+}15, JSMG18, JCL^{+}19, KIAD17, KdGP^{+}19, NNC^{+}19, QRW^{+}18, WLYL11, XWRZ19]. Age [TBS^{+}18, LZT^{+}19]. Age-related [TBS^{+}18]. Agent [MND^{+}19, BDP11a, Bar11, BGMLS17, COC10, DRGC^{+}19, FCY18, GJKP18, GGS13, HKG^{+}16, KMB^{+}17, KVK^{+}18, Kim14, KB16, LBD18, LJ17b, LFY^{+}19, MFSV19, SSG17, Sko19, WWX^{+}17, WXZL11, WWZ18, WHCW19, ZMN19]. Agent-based [MND^{+}19, BDP11a, BGMLS17, DRGC^{+}19, FCY18, GGS13, HKG^{+}16, KVK^{+}18, LBD18, ZMN19]. Agents [KSS11, AB19c, ADH^{+}16, KFBKD14, LCZR12, PBV^{+}13, SB19a, SMS14a, VRGR16, YP12, PVHTP19]. Aggregate [CFP^{+}19, GLB^{+}18, JSC^{+}15, LLL^{+}18, Wan19, YFY^{+}13]. Aggregated [IHK^{+}18, KYB^{+}19, RGC^{+}10, WWG19a]. aggregated-task [WWG19a]. Aggregating [SB17b, HQ10]. aggregation [BMZ10, CLR16, CIK10, CXC^{+}18, DLS^{+}12, ECPF17b, FRZ19, KLP19, KV12, LCBF13, LLQS14, PvSS17, RBLvM14, TAS^{+}18, WWDF18, ZGL19, ZZH^{+}18]. aggregation-capable [PvSS17]. aggregator [LLQS14]. agile [LG18]. agility [FLR^{+}16]. agility-oriented [FLR^{+}16]. Aging [FPR18, DLZ^{+}14]. Aging-related [FPR18]. agnostic [KDG^{+}19]. AGRA [CES^{+}19]. agree [CAC^{+}15]. Agreement [LYW^{+}16, PB18, AQRH^{+}18, APK^{+}18, IOV^{+}18, KLW^{+}16, MLC^{+}18a, MJRM16, ODK^{+}17, OSANAM19, PRS12, RZ16]. Agreement-based [PB18]. Agricultural [ZL18, CZH^{+}18, HPZL18, LBJ^{+}18, LBJ^{+}24]. agriculture [APK^{+}18, BNJ16, BW19]. ahead [Eng14, WYBS11]. AHP [ABMMC22, ABMMC18]. AI [CES^{+}19, LLT^{+}19, QLM^{+}18]. AI-augmented [CES^{+}19]. AI-defined [LLT^{+}19]. AI-enabled [QLM^{+}18]. aid [ABM19]. aided [GSC^{+}19, LYL15, MGA^{+}18, Sko19]. AIEM [QLM^{+}18]. AIM [TWG^{+}19]. air [LLN^{+}18, ZXW^{+}18]. air-land [LLN^{+}18]. Airavata [GRZ^{+}19]. aircraft [TYWZ18]. al. [RLM18]. alert [NJKH13, SSB13]. alerting [MVL^{+}18a]. alerts [CPP16]. algebra [JLRS18, MKM11, WZ16]. algebraic [Che13b, MdOO^{+}17]. Algorithm [CT19b, GAA19, HMF^{+}19, JGB19, MPI^{+}18, SE19, TA18, TKA18b, ABMESM18, AHEM17, AA18, ATF11, AIA^{+}18b, AS18a, AY16, Bag11, Bag19, BZMY10, BMT12, CT19a, CPDJ13, CLRL17, CZY^{+}18, CZL^{+}18a, CLZ18, CJ14, Che18, CWJ^{+}18a, CGSJ18, CCC19, CHY^{+}18, CMZ^{+}18, CZL^{+}18b, DYY^{+}19, DQXW19, DC19, EAA16, EA17, FLT17, sGbKS19, GPS^{+}17, HZW19, HPZL18, HLL18, HHM^{+}19, HEES19, HXWW18, HLL^{+}17, HBN^{+}13, HZLH19, HLZ18, IS18, ISS^{+}15, JGFB18, KHMB13, KB18, Kim18, KLP19, KKS18b, LC17, LLpC12, LCL^{+}16, LSZ^{+}16, LGY^{+}16, LJJ18, LZY^{+}19b, LZZ19, LZL^{+}19a, LLL^{+}19, LC15, LWH^{+}18, LJC^{+}19, LAH10, LZXG12, LWZ18, LLSL18, LW18b, LYH^{+}19, LZWF19, MWL18a, MNV12, MHC14, MLC18b, MZYA19, MAÇ17, MGA^{+}18, MAA^{+}19, MFT^{+}17, MCSA18, MRN19, NPP12, NWL17, OB19, PC18a, PKF14, PRC^{+}14, QCY^{+}19, QMSG12, RGAT18, RS17a, RWZ^{+}19, SR12, SGRT19, SOD18, SMS^{+}19]. algorithm [SPD^{+}19, SYAL13, SLB^{+}17, SJL^{+}18, TDFZ18, TLL^{+}11, TJZ^{+}15, TZLL18, TM19, TWW^{+}18, TdPF^{+}17, VAdlP12, VPT^{+}15, VMN^{+}18, WYBS11, WCL^{+}17b, WJZ^{+}17, WDD18, WWZZ18, WNR19, WZ18, WLA17a, WLA18a, WCC14, WWZ18, WHCW19, XY15, XRPT18, XHL^{+}19, YWCC18, YHL^{+}19, YZWG18, YSZW18, YPJ19, YDD^{+}18, ZRZL18, ZWL13, ZYCZ19, ZBL^{+}14, ZRZ^{+}14, ZL18, Zin18, ST11]. Algorithmic [GTEL^{+}18, AB18a, WM14]. Algorithms [CCMGF18, MJDN15, RGCCL18, ANE13, ACWJ19, AMÇ19, AGKZ18, AB16, ACD^{+}19, CLCMG^{+}18, CZY^{+}19, CGM^{+}18, CSQL17, Cuz14, ESPN17, EP13, FK12, FGM11, FM10a, GLJ19, HH19, HKS18, HZ10, JOSD19, Kha12, KGT15, KVHT10, LLC11, LCW^{+}18, MCA^{+}18, Mér17, MVCC10, NAD^{+}18, SF19, SC16, SG13, VSDD13, WLQ10, WWT^{+}16, XWJ^{+}16, XLW^{+}17, YFY^{+}13, ZGV19, ZLG^{+}14, ZAC^{+}18]. alias [MWYC12]. ALICE [JCA^{+}19, PcFP^{+}17]. Aligning [SBLW14]. alignment [KKvdB^{+}17]. all-to-all [ZTKF17]. alleviate [DV13]. Alloc [QMSG12]. allocate [LLF^{+}18a]. Allocating [MDD15]. Allocation [DXL^{+}18, AOIS10, AJR^{+}19, AAQ^{+}19, BAA^{+}19, BAB12, BCDP12, dCCDFdO15, DVB14, DVD12, DEG^{+}17, EZTL19, EMJ^{+}13, FDPR17, FEPC18, GCZ^{+}19, HMH17, HH19, HDLW13, HLL^{+}17, IAL10, JSS^{+}12, JMAG19, KOT18, KPJ19, KKW^{+}14, KA19, LN13, LPK17, LPK18, LWD^{+}14, LLZ^{+}18a, LTTL19, LZCX19, LSL^{+}18, NCS12, PSPP16, PdASM18, PPL^{+}15, QMSG12, QPTGG^{+}12, RCTY19, SPD^{+}19, SCMS12, TTH15, TSBH11, WCW18, WY17, YCY10, YSC^{+}15, YLJL18, YC13, YW12, YMY^{+}17, ZGL^{+}18, ZXZL18, ZB19]. allocations [SMS14b]. allowing [VRGR16]. along [DFLO17]. alternating [SMC18]. Alzheimer [NDZ^{+}18a, NDZ^{+}19, NDZ^{+}18b]. Ambient [VKT^{+}19, YGS16, CRC^{+}19, FKT14, MLGGB^{+}17, NPH19, OdI14, SJSA19, TF17, VRGR16]. ambulance [KKS^{+}18a]. AMD [DRC^{+}19]. America [PKP19]. among [GSY^{+}17, HS19, LLF^{+}18a, SAGGB17, SCZ^{+}19, VGD^{+}19, dFPFG19]. amphibious [LLN^{+}18]. analgesic [WWA19]. Analyse [ACMM19, Bel16]. analyses [ATdC^{+}16, GCBM17, JHC10, OdOD^{+}13, SBA^{+}17]. Analysing [BDP11a, WTP^{+}13]. Analysis [AM17, BBC^{+}17, BHE^{+}19, BGC19b, CMEA^{+}19, KVR15, KCH^{+}13, KKA18, Li15, LRYJ17, MCA^{+}18, MG16, MCRB19, MKS^{+}19, SHL^{+}19a, SB11, WTG^{+}14, WCWC20, AAYL19, ASV^{+}13, AB18a, AMR^{+}19, ABG17, ASYF18, AMÇ19, AGKZ18, ATX13, AAS^{+}19, AAI^{+}19, AEME^{+}18, Bag19, BSRR18, BBC^{+}13, Bo19, Bo20, CSV^{+}12, CA15a, CCRV13, CFM17, CGIP14, CPE^{+}17, Cha15, CHC^{+}17, CLL18b, CRYG18, CWUS19, CDMR19, DMC^{+}19, DZH18, DCBF19, DLS14, DMPS19, DDMPG17, DMM^{+}18, DNW^{+}19, DMM14, DZLA19, ETR^{+}13, FAA^{+}18, Fer13, FNA12, GHMX10, GAB^{+}14, GHP^{+}18, GLNT13, GBY16, GRMSOG18, GPJC17, GRX19, sGbKS19, GRS^{+}19, HHL11, HPP^{+}18, HAM18, HMW14, HZX^{+}19, HZX^{+}20, HLL12, IPCA^{+}16, JBC16, JLS19, JYZ^{+}19, KI19, KPS18, KCV11, LCGPC19, LSB^{+}18, LOR^{+}18, LWW^{+}18, LCH^{+}18, LQF19, LWR^{+}19, MLL15, MZH^{+}17, MFC^{+}19, MZP^{+}19, MT17, MSA^{+}19, MK19a, MHdlS19, MEBA12, MMC^{+}18]. analysis [MSM^{+}18a, MBL^{+}19, NKP16, NJ17, NAM^{+}19, OKF10, OMPSPL^{+}19, OCCK14, PVN^{+}12, PBL^{+}18, PTM^{+}18, PPS^{+}18, RJN^{+}19, RHH^{+}19, RS16, RAA^{+}18, RLM18, RGCCL18, SP18a, SOA17, SHN10, SVB^{+}19, SG17, SWW^{+}13, SAG19, TSD18, TBK^{+}10, TLSC17, TBR^{+}19, TM19, TKTG19, TSGVRGS19, VCL^{+}19, VR12, WJS^{+}18, WDJC18, WXZ^{+}18a, WYJ^{+}19, WZH^{+}19, WZY^{+}19, WMA^{+}19, WZL18, WCWC19, WLA18b, WZ16, WWA19, WLZ^{+}19, XFTZ16, XL19, YPCK12, YPJ19, YN18, ZZDM^{+}18, ZSX^{+}15, ZWZ18, ZWWL18, ZMZ^{+}19, ZRZR19, ZMZ^{+}20, ZCZ^{+}18, ZWJ19b, ZXL14, dACAM13, DDJ^{+}13]. analysts [SOA17]. analytic [LJC^{+}19, MMVP13]. analytical [CPGBC16, GS13, JS13, PBA18]. Analytics [DP20, DP21a, DP21b, MGL^{+}18, PN13, ABD^{+}19, AAA^{+}19, ATM^{+}19, ADLM18, BA17, BAJ^{+}19, BMP^{+}16, CZXL18, CMP^{+}17, DP19, DGCGH^{+}17, FSV^{+}19, FFGP^{+}19, FRM^{+}18, GQLX18, HHZ16, IHA18, KPB18, KMST19, LJ19a, LHPC^{+}19, LIC18, MID16, ML19, MDT^{+}18, Osm19, SSA^{+}19, SG19, SJSA19, WSQ^{+}18, WHYZ18, WLHH18, XWZ^{+}19, YSHM19, ZZH^{+}16, ZWQ^{+}19, uRYS^{+}19]. analyze [RWV^{+}13]. Analyzer [VOV17, SHLJ13]. Analyzing [CRB^{+}16, CBLS13, HSB^{+}18, PLL^{+}18, HHH^{+}19, HHD^{+}12, UZ11, WWSL19, WMLS14]. AndrODet [MdFTGM19]. Android [GMCM18, AAI^{+}19, DMM^{+}18, GMCM16, JCA^{+}19, LLWZ18, MH19, MHdlS19, MdFTGM19, MRL14, PZC19, SGS^{+}18, WWH^{+}17, WLW^{+}18, WMJW18, XWRZ19, ZRZR19, ZCQ^{+}16]. AndroKit [AAI^{+}19]. Aneka [CVKB12, TSB18, VCKB12]. anesthetic [WWA19]. ANFIS [BBMG10]. angiography [DGA18, RDSA18]. angle [XZ16]. angle-based [XZ16]. Angular [DJH^{+}19]. animal [KS18d]. ankle [Bo19, Bo20]. annealing [Cha11]. annotating [AAF18]. annotation [GMP^{+}16]. anomalies [FPR18, LYS^{+}19, RKB18]. Anomaly [HZW^{+}16, ZGV19, AMI16, BS17, CYW^{+}19, FJ18, GNGG17, VSBN19, WWZZ18, WXZ^{+}18b]. anomaly-based [VSBN19]. anonymity [AMQS^{+}19, TVV13, WHS^{+}18, YHL16, ZLT^{+}19]. anonymization [KC19a, NK17]. Anonymizing [ZLL17a]. Anonymous [FHZW18, WWDF18, AIB^{+}18, ECE^{+}19, LHM14, LYL15, MLC^{+}18a, VCD^{+}18, YZL^{+}18, ZLL^{+}19, ZMH^{+}18]. answer [YLG^{+}16]. Answering [ZLXZ18, LLYW19]. Ant [TA18, FZHH14, SH19, TV16, WZ13, ZHHQ18, PT16]. anti [BBH18, CMZ^{+}18]. anti-collision [CMZ^{+}18]. anti-unification [BBH18]. AntiCheetah [DLMS15]. Anticipating [Lea15]. anxiety [TCH19]. AODV [ASAA18]. AOFAS [Bo20, Bo19]. APA [JNR12]. Apache [CSG^{+}18, FPR18, GRZ^{+}19, HSV^{+}17]. Apart [GMM18]. APHID [SGKC10]. API [BC17, XLL^{+}18a]. APIs [BBC^{+}17, RMJ^{+}18]. Apnea [DDD^{+}19, HJA^{+}19]. App [PS13, CMVA18, WWH^{+}17]. Appearance [RSY^{+}18]. appliance [FTK17, KTKN11, SH19]. Application [BB17, BR18, CBC^{+}19, CT19c, GEAR13, GRS^{+}19, GZZ^{+}18, HMH17, KMZJ16, KKK^{+}19, KKJJ10, LOR^{+}18, LRMS19, MED16, MCRB19, PdASM18, RMDB18, SCL18, SWW^{+}13, SCL14, WMBV17, YYW^{+}19, ASW11, ANG^{+}19, AHP16, AS14, BGC^{+}19a, BWR12, BKKM11, CPDJ13, CHJS^{+}10, CPE^{+}17, Ciu10a, CRB^{+}16, CZL^{+}18b, CRWZ19, DFGR14, DdM10, GVDT16, GMCM16, GMCM18, HLNM11, IG12, IFD^{+}19, JNHL18, JEB18, KANS18, KSS19, KOT18, KIJ^{+}19, KDG^{+}19, KMK^{+}14, KKA18, LLES19, LN18, LZL^{+}12, LYYW19, LPD^{+}13, LBB^{+}19, MSM^{+}18b, PAL^{+}19, PGTBC18, RBN13, RVC16a, RPMG10, SRZD15, SEMJ11, SI18, SSL13, SCJ^{+}19a, TDFZ18, TOD17, TMW^{+}17, VTTK17, VCL^{+}19, WLYL11, WWD^{+}14, WSL^{+}19, WLH^{+}19, XLL^{+}14, YSC^{+}19, YZWG18, YY11, ZJW^{+}14, dCRL^{+}19, ABS11, FCD^{+}14]. Application-aware [RMDB18]. application-based [WWD^{+}14]. application-layer [ZJW^{+}14]. Application-level [KKK^{+}19, LBB^{+}19, RPMG10, SRZD15, SEMJ11, SSL13, WLYL11]. application-specific [KANS18]. Applications [MG18, PN13, RC18, RC19, SZW^{+}19, WHW17, ZZLR18, ZYA^{+}18, AHS^{+}18, APAZ17, AOIS10, ALTG19, AUSA19, ACH^{+}11, AKCY^{+}17, ABS^{+}18, AASI17, AHU^{+}19, AHL11, AJY12, AS19b, ACCM19, AB16, BGI14, BKS^{+}18, BC15, BC17, BDNP13, CGN18, CGCB^{+}12, CVKB12, CTVB12, CVT19, CLCMG^{+}18, CPGBC16, CSdCM^{+}17, CRVZ15, CA15b, CCL11, CMNK19, CES^{+}19, CXL^{+}17, CW13a, CGL15, CGSV17, DQC^{+}19, DST10, DST14, DRGC^{+}19, DMM^{+}18, DPL14, EAS^{+}18, EAED18, EET18, EMJ^{+}13, ECA^{+}18, EKGS14, ETR^{+}13, FG18, FTP14, FRM^{+}18, FFC12, Frî14, FM10b, FSP^{+}18, GVURIVBV14, GBS10, GAW^{+}18, GHD19, GHEB^{+}18, GHEB^{+}23, GAJP18, GMD19, GKW^{+}12, GIM16, GIK18, GPS^{+}17, HGG^{+}14, HC17, HMF^{+}19, HPGMM18, HKPT10, HXA^{+}17, HHW11, HSC15, IDCJ11, JOPW14, JP18, JLQ18, JMAG19, JBP^{+}18, KMB16, KWK^{+}18, KK19, KZ14, KC14]. applications [KSC^{+}19, Kol18, KMU19, KAS^{+}18, KFK19, KKKM17, KKKM18, LJLW13, LMZ^{+}14, LRYJ17, LTTL19, LWTL19b, LCL14, LZYC13, LZX16, LBB^{+}19, LZHY19, MBB10, MFN13, MAJD18, MZC10, MZL^{+}19, MAD^{+}16, MSM^{+}18a, MDD15, MRL14, MND^{+}19, MMPF19, MMRL17, MVC^{+}13, NF13, NAGD18, NS10, NLV^{+}19, OWX19, PKY^{+}17, PC17, PRL^{+}19, PMPC13, PVHTP19, QCYJ17, RRS10, RBJ^{+}13, RMCN^{+}10, RVC16b, dRRRR^{+}18, dRRdQGR^{+}18, RSJ^{+}14, RBC^{+}15, SB19b, SHN10, SBAD^{+}18, SPR^{+}10, SJ18, SNC18, STB^{+}19, SB17b, Sip12, SLSS19, SSKK13, Śle14, SG14, SK19, ŠCJ^{+}19b, SVN10b, TZQ18, TKA^{+}18a, TKTG19, TVB18, TAHS14, TBB^{+}17, TCBC18, TSB18, TGM^{+}19b, TSAER18, URKM19, VCKB12, VETT16, VL19, VSKS19, VRS^{+}19, WYBS11, WTM^{+}17, WJS^{+}18, WY17, WLW^{+}19, WCVL12, WBKL16, WOPW13, WG13, WWT^{+}16, WHYZ17, WPS^{+}18, XYLZ18, XLZ18, XLL^{+}19b, YZC^{+}19, YNSM12, YFY^{+}13, YWCC18, YJHZ14]. applications [YL16, ZME^{+}15, ZZ15, ZCK^{+}15, ZCS^{+}16, ZZZC19, ZHW19, ZSMS18, ZBF14, dSFP^{+}17, dFPFG19]. applied [DDJ^{+}13, MGMT18, SSFFR19, ZPPE17]. applying [KNI^{+}18, MSE19]. Approach [SAH19, ABZK15, ABDH19, ARP^{+}19, AR18, AIP^{+}19, ACSdRR17, AVPV17, AL18, Ano12r, ASY^{+}18, AB17, AB18c, ABN19, ATX13, AMR18, AMBC19, Bae14, BBD^{+}19, Bal16, BBWB^{+}18, BRXdS11, BFS^{+}17a, BFS^{+}17b, Bel16, BRH18, BMZ10, BRB19a, Bu18, CQW^{+}19, CGCB^{+}12, CWJD19, CCRV13, CA13, CFG^{+}19, Cha11, CES^{+}19, CAL^{+}18, CMP^{+}17, CTU19, CGM^{+}19, DEL19, DHW^{+}17, DA16, DV13, DLDTGMMP16, DMMM11, DC18a, DdM10, DLXR14, DdSdN^{+}19, DPL14, EG18, ECA^{+}18, FTK^{+}14, FJL^{+}16, FEPC18, FPPD14, GHEB^{+}18, GHEB^{+}23, GAJP18, GLD^{+}19b, GNGG17, HO17, HDKC18, HHL11, Ham17, Ham19, HZC10, HLYW17, HSBE19, HHK^{+}16, HFL^{+}19, HZ19, Hua10, HQ10, HHZ16, JLCC12, JS13, KSS11, KZA^{+}18, KTKN11, KKB14, KJI11, KK14, LBD18, LA19, LY17, LJLW13, LXL^{+}17, LPY^{+}18, LYW^{+}18a, LDJL19, LRMS19, LWS^{+}12, LSL^{+}15, LHW^{+}18, LO19, LCCM18, LKJ17]. approach [dSMAdR^{+}17, ML11, MdOO^{+}17, MdMMNS^{+}19, MEBA12, MSBA16, MHZK18, MBL^{+}19, MSS^{+}16, MAB^{+}15, MSM^{+}13, MMRL17, NSR^{+}19, NS17a, NJH^{+}18, NGB18, NJKH13, OMKM^{+}19, ODC19, PZC19, Pal13, PC17, PNZ14, PSY^{+}19, PVHTP19, PS19, Pon19, PDW^{+}11, PPB16, QZD^{+}18, RAKJ18, RGN^{+}18, RDSA18, RLP12, RW18, RWO^{+}19, SSHC19, SBCF16, SSFFR19, SD18, SAM^{+}19, SZ12, SMS14a, SSJ19, SCZ^{+}19, SEPV19, SSI19, SGJ18, SCMS12, SBA^{+}17, SMM^{+}14, SRKS18, SL19, SK19, ŠCJ^{+}19b, SLL^{+}17, SAC11, TM19, TV16, TOS18, TMS^{+}17, TCCW19, UZ11, UPP17, VV16, WN10, WPY19, WXZL11, YAO14, YP12, YZ12, YNLY19, ZBCT17, ZYB^{+}18, ZZJY16, ZWJ^{+}18]. Approaches [GTEL^{+}18, ASAAM^{+}19, ALK15, BMU18, CMS^{+}18, CDMR19, EDH^{+}13, HHS^{+}18, LIH^{+}19, MOFGP18, SSG19, SB11, SNXB17, TAHS14]. Approaching [CAC^{+}10]. approximate [DK14, FZT^{+}18, GPS13, sGbKS19, SK19, WLW^{+}19, YDQC19]. approximation [BTG19, SC16]. apps [AMRM18, DC18b, PZC19, WLW^{+}18, WTTH19, HXC^{+}18]. AppTCP [WWD^{+}14].

APT [DC18b, ZCW19]. APTs [LA19]. AR-RRNS [CBT^{+}19]. arbitrary [kHsZwJW18, JLL17, VMSRM12]. archaeological [LKK^{+}16]. ArchaeoSTOR [GML^{+}13]. archeological [GML^{+}13]. Archimedean [TM19]. Architectural [KR14, GBMP13, LLES19, LTZ15, QCD16, TZL^{+}18, TSOB15]. Architecture [ACC^{+}19a, CBS17, GLM^{+}12, MBMTJR18, PGCML^{+}19, SCL18, ZDL^{+}13, ZLG^{+}14, AdI14, BBC^{+}17, BLO^{+}18, BCC^{+}17, BRH18, BDH14, CJ14, CWW^{+}13, CWW^{+}16, DVD12, DGCGH^{+}17, DJ13, EDH^{+}13, FNA11, FRM^{+}18, GDJ^{+}13, GD10, GIK18, GSY^{+}17, HIA18a, HDO16, JYY^{+}17, KR19, KS11, KSS19, KB18, LBJ^{+}18, LBJ^{+}24, LLW^{+}12a, LKN^{+}13, LFP^{+}17, LRC^{+}18, LML^{+}19, LMA^{+}19, LMCSE19, LHBC16, LYH^{+}19, MVL^{+}18a, MPCAF15, MPF^{+}16, MRH17, MRS^{+}18a, OBG^{+}18, PSW^{+}14, PBC^{+}11, PPMM^{+}18, PMLVLS^{+}13, PSBB15, RBGA18, RHPV17, RCOP^{+}11, SGKC10, SP18b, TSTD16, UKK^{+}19, VETT16, VDK12, VFHB14, WZW19b, WSH^{+}16, WZ16, XWW19, YÁJG^{+}15, YS16, ZMP10]. Architecture-based [ZLG^{+}14]. Architectures [DO15, HYZS16, TKRA14, VDTK12, BFP18, BPC^{+}14, CA15a, CDG^{+}14, CGIP14, CCL11, CPSRG14, DLH^{+}17, DDB14, DM12, EBCP18, FFC12, GRCP^{+}17, GLJ19, IS18, IGB^{+}14, JPB17, LX13, MAC14, MCA^{+}18, PSPP16, Pal13, PMCC18, PPPS18, RC18, RC19, SB19c, SCAC^{+}19, SMBMT^{+}18, dLB10]. Area [AS18a, PP10, AL14, GG10, HLYW17, LCBF13, NAGD18, SCS^{+}18, ZGV19]. Argumentation [DKFKF18, OAMS18]. ARM [OBG^{+}18, RRP^{+}14, SGN^{+}17]. ARM-based [OBG^{+}18, RRP^{+}14]. ARMCO [PPB16]. ARMOR [MML^{+}18]. arranging [GVdBdL15]. array [WHZL10, CSV^{+}19]. arrhythmia [MdMMNS^{+}19, SD18]. arrhythmias [AFO^{+}18]. arrival [MRN19, WMLS14]. art [CsZzG^{+}13, LCC19, SJV^{+}15, dCTVC18]. arthritis [YTQ19, YTQ20]. arthroscopic [Bo19, Bo20]. artifact [KYZ19]. artifact-centric [KYZ19]. Artificial [LFY^{+}19, SLS^{+}19, BZMY10, Che18, CZL^{+}18b, GJKP18, HXWW18, HLZ18, KS18c, MGA^{+}18, SMS^{+}19, XLW^{+}17, YZWG18, GKTK15]. artistic [UNM^{+}16]. as-a-Service [ACC^{+}19c]. ASA [SCH^{+}17]. ASIP [ABF^{+}15a]. Aspect [MWQ^{+}19, MZP^{+}19]. Aspect-based [MWQ^{+}19]. aspect-level [MZP^{+}19]. aspects [AR18, BBvdB^{+}11, CSL18, MZP^{+}19, MZC10, QCD16, WM14]. assemblies [RSRA18]. assess [FFGP^{+}19, Kim18]. Assessing [PRL^{+}19, SG15, ZGS^{+}13]. Assessment [FW19, vdPGZ^{+}16, AB19a, ABGMC19, AB21, AJ19, ACSV18, BBC^{+}17, CCL19, FJA^{+}18, GAMC19, KHG^{+}18, LCGPC19, LZT^{+}19, PRSR14, PB18, PKA19, RHH^{+}19, YYS^{+}19]. Asset [PLLP19]. assets [GTM19]. assigning [HCL^{+}17]. assignment [ADBO18, AAM^{+}16, DLXR14, FGW^{+}19, HYG^{+}19, KMT14, KRZ^{+}19, LTC12, LXD17, RS17a, TdPF^{+}17, VAdlP12, VTTK17, WZM^{+}18]. assistance [SZD^{+}17]. Assistant [VKT^{+}19, FGW^{+}19, HIA^{+}18b, LCMX16, SNC18]. assisted [BDS^{+}10, CCC19, FKT14, FFL^{+}19, HMA^{+}18a, HDLW13, JYZ^{+}19, LZY^{+}19a, LLC14a, LNY^{+}18, LYL^{+}19, NDA^{+}19, RSK16, SJSA19, TMB^{+}19, VRS^{+}19, WXYL15, WXZ^{+}18a, ZYZ^{+}18, ZWQ^{+}19, HLL^{+}19, WLML17]. assisting [SGRT19]. Assistive [KPG19]. association [AVPV17, EBOY14, LHW^{+}18, LWZ^{+}19b, PGCML^{+}19, RAKJ18, XL19, ZDW^{+}18]. associations [GJ18, KZCW13]. Associative [HMMW19, SSZ^{+}17, CMP^{+}17]. assortment [SSHC19]. assurance [GSC11, LLW^{+}12a, WWRS16, WLW^{+}19]. astronomy [LBM18]. astrophysics [LKA^{+}19]. asymmetric [CRYG18]. asynchronous [CCL11, PvSS17, WHCZ18]. asynchronously [SLG^{+}17]. ATLAS [JLRS18, KVR15]. atmospheric [LN18, PPZ12, WJZ^{+}17, ZDL^{+}13]. atom [ZWZ19]. atomic [Bag11]. attack [AQAR^{+}18, AAC^{+}19, BZHV19, CM17, DYY^{+}19, DC18a, DCC13, JNHL18, KAW12, LSL^{+}15, LMM19, NZL^{+}15, NAM^{+}19, SPT^{+}18, SCLC19, WLYL11, YWJ^{+}18, ZCW19]. attacker [PLGMCdF18]. Attacks [JL14, LW19, IDKD19, JCL^{+}19, KIAD17, KPS18, KdGP^{+}19, NNC^{+}19, OA17, QKC19, SCLC19, SCH^{+}17, SSB13, TA18, VS13, ZJW^{+}14]. attention [FJA^{+}18, LLT^{+}19]. attention-scoring [FJA^{+}18]. attestation [ECE^{+}19, GZW18]. attitudes [RWY^{+}18]. attractor [TZD^{+}19]. Attribute [CDL18, LAL^{+}15, LHL15, LWL^{+}18, Rao17, SYK^{+}17, WLXZ18, CIK10, FLT17, FRZ19, GB10, HZL18a, HYS17, HYS18, JSMG18, LYL15, LDZW19, MWQ^{+}14, QRW^{+}18, QGT^{+}18, SMSF18, SCL14, SCZ^{+}14, VPP^{+}19, XZP^{+}19, XTZ^{+}19, YCT15, ZCL^{+}18]. Attribute-Based [LAL^{+}15, LHL15, Rao17, CDL18, LWL^{+}18, SYK^{+}17, WLXZ18, HZL18a, HYS17, HYS18, JSMG18, LYL15, LDZW19, MWQ^{+}14, QRW^{+}18, QGT^{+}18, SMSF18, SCZ^{+}14, VPP^{+}19, XZP^{+}19, XTZ^{+}19, YCT15]. attributes [BFS^{+}17a, BFS^{+}17b]. attribution [Mil11, NAAC19]. Auction [ZBL^{+}14, IAL10, YLJL18, ZXZL18]. Auction-based [ZBL^{+}14]. audio [BDMO11, LYXT14]. Auditable [ED16]. auditing [LYY^{+}18, YNY^{+}14, YXA^{+}16]. auditory [CPP^{+}18, YXY18]. augmentation [AGKZ18, GBRM18]. augmented [CES^{+}19, KMV^{+}15, LWZ^{+}19a, RMSPP17]. Augmenting [HMZ18]. August [Ano19q]. Aura [HHK18]. AusPlots [TSTL16]. authenticated [ODK^{+}17, OSANAM19, PPG19, YZL^{+}18, ZXWA18]. Authentication [AMSPL19, ASO14, ABB19b, JLX^{+}19, AMN18, AMKC19, AHM^{+}18, APK^{+}18, AIM^{+}19, Alp18, AKB^{+}18a, BK19, BDL^{+}19, BDM^{+}19, BLAN^{+}16, CJXX19, CHS11, DEL19, ED19, FHZW18, GHD19, GAI^{+}18, GLB^{+}18, Ham19, HZW19, HLC16, HPL^{+}19, IOV^{+}18, JKAU19, KVvE18, KLW^{+}16, KLW^{+}17, LC17, LNK^{+}18, LWK^{+}18, LW19, LHM14, LH13b, LYL15, LNY^{+}18, LWF^{+}17, MCN^{+}18, MZL^{+}19, PYH^{+}18, PLGMCdF18, PCK19, SGGCR^{+}16, SCS^{+}18, SYW17, VCD^{+}18, Wan18b, WDKV19, WLS^{+}18, YHL16, ZZY^{+}19, ZLY^{+}19]. authenticators [SYY^{+}17]. authoritative [WMA18]. Authorization [KWR^{+}13, ABTF16, BPC^{+}14, CHC^{+}17, CZZ^{+}18, CBPP18, DLLZ17, GVdBdL15, HHD^{+}12, LXZ^{+}18, PBC^{+}11, SBL18]. authorization-security [ABTF16]. authors [AK18a]. authorship [AsRA^{+}19]. Auto [WLH16, DWS12, EPB18, KTTK17, LCL^{+}18, LMA^{+}19, RSY^{+}18, SSI19, TV16, dACNC16, vW19]. auto-encoder [LCL^{+}18]. auto-scaling [DWS12, EPB18, KTTK17, LMA^{+}19, SSI19, dACNC16]. auto-tuner [vW19]. autocorrelation [DH16]. autoencoder [BR19, UMUB19]. Automata [TA18, CBBdL16, PRN14, RGAT18]. Automata-based [TA18, CBBdL16, RGAT18]. Automated [AD18, AFO^{+}18, Asu13, BGNI19, CPP16, FTD17, GACM17, GGTRRC16, GMM18, KHO^{+}19, LJ17a, LJW^{+}19a, NUPA19, PKI^{+}18, RBGA18, RHKC15, AAF18, CsZzG^{+}13, COC10, GBRM18, GRS^{+}19, Hua10, JBR^{+}16, LZT^{+}19, RSRA18, dSSCdL19, YPHZ14]. Automatic [CMVA18, HWWT12, HAAR^{+}19, LLN^{+}18, MCdA16, XZ14a, ZLZ13, ZSL^{+}19b, ZH17, AdI14, AMM16, BBH18, DZLA19, GMMM18, HLZ^{+}19, MDM^{+}19, TMDZ15]. Automatically [CHSA18, BCF16]. Automating [XLL^{+}18a]. automation [ACPI19, DVJ^{+}15, JBC16, WBKL16, WTP^{+}13]. automaton [LPY^{+}18]. automotive [LXT^{+}19]. Autonomic [AHU^{+}19, Erd13, NLV^{+}19, YVCB10, ASAB^{+}18, ENC^{+}12, GAJP18, HCZW17, IGB^{+}14, MSBA16, MTD18, MAB^{+}15, PVN^{+}12, SMPC12, TTB^{+}13, TCR^{+}12, Vin16, ZYW^{+}18]. Autonomous [RYH^{+}19, DP19, DP20, DP21a, DP21b, DBS14, FGW^{+}19, PBV^{+}13, SB19a, TCCC11, YP12]. AUTOSAR [KR19]. autoscaling [ZTL^{+}19]. AuTrA [KJI11]. AUV [KKN18]. AUV-based [KKN18]. auxiliary [Wan19]. Availability [NKP16, ASD12, KKAS19, LCH^{+}18, RLP12, TSWL17, VVC^{+}12, WLML17]. available [Frî14, GVI13, JSS^{+}12, SB14]. average [AMÇ19, NJKF18, YK17]. average-case [AMÇ19]. average-utility [YK17]. avoid [OWX19]. Avoidance [SSK^{+}19, Ciu10b, DPK^{+}19, NLS19, ZGL19]. Aware [ABTA18, BLO^{+}18, BTP19, JGB19, KIMR15, KV17, RMA^{+}18, SV16, TKA18b, WLL^{+}19b, WDW^{+}19, AHEM17, AGR19, ADAAD12, ARSMY19, ADA^{+}19, AS19b, ALK15, AMR^{+}19, ACCD17, AMT^{+}12, AGA16, AB17, AB18c, ABN19, ACSV18, AC18, AEME^{+}18, BAA^{+}19, BAB12, BPC^{+}14, BCDP12, BRB^{+}19b, BMK^{+}14b, BCR^{+}12, CXZ^{+}19, CCIP18, CBC^{+}19, CCCT14, CHY^{+}18, DWJM18, DPK^{+}19, DKV14, DFG^{+}19, DDD18, DSBC19, DA18, DCF19, DCMW17, EMHE18, FG18, FKT14, GMMM18, GJGB19, GHD19, GDJ^{+}13, GJ15, GD10, GS16b, GKTK15, GBKJ18, HGG^{+}14, HCMJ19, HNKÖ18, HCJ14, HH19, HSC15, HZ19, HBN^{+}13, JGFB18, JLQ^{+}17, JNR12, JLD^{+}19, JEB18, KMB16, KS18a, KANS18, KSF^{+}13, KRD^{+}19, KCS14, KKW^{+}14, KBB^{+}16, KIC12, LTC12, LLC14a, LPK17, LPK18, LGY^{+}16, LYYY18, LJGW18, LAQ^{+}19, LWS^{+}12, Lok12, MCL^{+}16, MHC14, MBM18, MHZK18, MQL^{+}19, MPR^{+}16]. aware [MBA19, MMPF19, NKB19, NK15, NSSA^{+}14, NJ16, PP10, PC18a, PLZX19, PFPJ18, PNZ14, PAB^{+}14, PRC^{+}14, PSW^{+}19, Qur19, RMRSA19, RMDB18, SSG19, SC16, SHRE16, SDTA19, SAK19, SCY^{+}18, SMG18, SSSJ19a, SRN^{+}18, SK19, SSSJ19b, SCCS11, SCH^{+}17, SSP17, SLY^{+}19, SGL^{+}19, TKR^{+}15, TCCC11, TCR^{+}12, TuIS^{+}19, UGBM^{+}17, UDvdW^{+}18, VNAMM19, VDTK12, VPT^{+}15, VGC^{+}13, WKC^{+}13, WWC14, WCC^{+}16, WLZ^{+}16, WZZ16, WWQ^{+}18, WCH^{+}18, WSZC18, WZM^{+}18, WPY19, WWZ^{+}19, WS10, WZ13, WYH^{+}17, WZS^{+}18, XDH^{+}17, XCZ^{+}19, YIA17, YCX18, ZMTT16, ZL13, ZME^{+}15, ZHHC17, ZGL^{+}18, ZWS^{+}12, ZSL^{+}19a, ZSGJ19, ZQB^{+}18, ZZLH18, ZAC^{+}18, ZFH^{+}18, ZTD^{+}18, ZWYH19, dSGD19, dACAM13, KC14]. awareness [AL14, ABF^{+}15b, GRTV10, LCL14, RR18, SHBP10, WSC^{+}19, XFTZ16, ZRZR19]. AWS [GAMC19].

B [WWT^{+}16]. B.A.T.M.A.N. [DCBF19]. B.A.T.M.A.N.-based [DCBF19]. Back [SKS^{+}18, CWSW14]. back-end [CWSW14]. backbone [ZJW^{+}14]. Background [LXT^{+}19, TWZP18, HCC^{+}14, YPHZ14]. backhaul [WCW18]. backup [PJDO13, WZZ16]. backward [NJB19]. Bacteria [VR12]. Bacterial [RC13]. bad [TWG^{+}19]. bag [APAZ17, CLRL17, CA15b, GGS13, MVC^{+}13, TVB18]. bag-of-task [CLRL17, MVC^{+}13, TVB18]. bag-of-tasks [APAZ17, CA15b, GGS13]. bagging [ArMS19]. balance [CWW^{+}16, DXL^{+}18]. balanced [AS18a, CTR^{+}17, CFH^{+}19, DC19, SDTA19, WWT^{+}16, YWCC18]. Balancing [GXL^{+}18, LCZR12, MD12, AA18, HLW12, IS18, KRZ12, LZXW13, LN18, LYH^{+}19, MKM11, PZA18, PRC^{+}14, SMG18, SJTN18, SLH^{+}19, TZLL18, WCF^{+}15, ZGB^{+}17, ZFC17, ZFC18]. band [SYCH18, VAdlP12, WYJ^{+}19, WGX^{+}19]. bands [SLH^{+}19]. Bandwidth [XDH^{+}17, AHP16, AHP^{+}18, KWK16, WLP10, WQG15]. Bandwidth-aware [XDH^{+}17]. Bank [WY17, NUPA19, YW12]. banking [KVvE18]. BANs [BLL^{+}19]. Barrier [WPJ16]. barter [SM10]. bartering [GBRM18, ZA13]. Base [BBCN18, Cha15]. Based [GBY16, GGC18, HXA^{+}17, LAL^{+}15, LHL15, ML17, RRKA19, Rao17, SME^{+}21, Sar18a, ST11, SCH^{+}17, YG18, AD18, ABZK15, AMN18, ABMESM18, ABMMC18, ABM19, ABMMC22, ANA16, AJY15a, AA18, ARP^{+}19, AJR^{+}19, APRC16, AKP^{+}18, ANG^{+}19, AJ19, AAAQJ^{+}18, ADAAD12, AASI17, ABB^{+}19a, AAQ^{+}19, ARP14, ASO14, AIA^{+}18a, AIM^{+}19, ACHP19, ATdC^{+}16, ADH^{+}16, AK14, ACMM19, AAM^{+}19, AMKM18, ABF^{+}15a, ASAA18, AMT^{+}12, ACCM19, AKG^{+}17, AB16, AMPZ16, AZO^{+}19, AM19a, ACPI19, BK19, BAJ^{+}19, BDE17, BBC^{+}17, BG12, Bae14, BBD^{+}19, BRL19, BKB11, BDP11a, BAKB19, BDL^{+}19, BARMB14, BFP18, BDM^{+}19, BBT19, BCdV^{+}19, BRH18, BBC^{+}12, BMU18, BW13, BT17, BGMLS17, Bu18, BW19, CT19a, CSV^{+}12, CMZ^{+}12, CEP19a, CLRL17, CQW^{+}19, CAC^{+}10, CLZ18, CLL^{+}18a, CPP^{+}18, CPD^{+}15, CPE^{+}17, CFL^{+}15, CLL18b, CXZC18, CMI^{+}19, CES^{+}19, CCS^{+}10, CLL^{+}14]. based [CYZK15, CWJ16, CJG^{+}18, CLH^{+}18, CLR18, CRYG18, CZH^{+}18, CZZ^{+}18, CLC^{+}19, CT19c, CLS19a, CYJ19, CZ19, CGZL19, CXWT19, CLS^{+}19b, CJS19, CMP^{+}17, CWL^{+}19, CS12, CB10, CGSJ18, CGSV17, CLK11, CFF14, CDL18, CMZ^{+}18, CZL^{+}18b, CWLZ19, CRWZ19, CRTN17, CBBdL16, DQC^{+}19, DZZ^{+}15, DJZ^{+}15, DCBF19, DDD^{+}19, DRGC^{+}19, DMPP16, DNJG17, DLDTGMMP16, DA18, DMMM11, DdM10, DLS^{+}12, DNW^{+}19, DCC13, DKFKF18, DdSdN^{+}19, DZLA19, DK14, DNP14, ELAEAVAM19, EAA16, EBOY14, EMJ^{+}13, EPB18, EA17, FTK^{+}14, FG18, FH13, FHYH15, FLT17, FCY18, FSY^{+}19, FJA^{+}18, FSM^{+}18a, FJJ^{+}18, FMV14, FSV^{+}19, FHZW18, FY19, FW19, FAL^{+}19, Fer13, FEPC18, FTD17, FFC12, FKOC11, FWB13b, FA11b, FTH16, FPP^{+}18, FM10b, FSP^{+}18, GAFFOG12, GHY^{+}18, GAW^{+}18, GSC11, GDJ^{+}13, GA13, GSLI12, GAJP18, GS15, GCK18, GJF^{+}12, GCCL18, GCTLA^{+}19, GPJC17, GB10, GPV^{+}14]. based [GP11, GMP^{+}17, GIK18, sGbKS19, GSR^{+}19, GPVN19, GGMS18, GGS13, HAJ^{+}19, HIA18a, HO17, HDKC18, HHL11, HKA^{+}18, Ham19, HZW19, HSM13, HPZL18, HZL18a, HLL18, HLH^{+}18, HZW^{+}18, HLW12, HUY^{+}19, HJA^{+}19, HAP15, HDO16, HAT19, HHM^{+}19, HLYW17, HNKÖ18, HSBE19, HCZW17, HHW^{+}19, HHK^{+}16, HIA^{+}18b, HMC19, HZL18b, HLL^{+}11, HMW14, HLC16, HDH^{+}18, HZZ^{+}18, HMW^{+}19, HZ10, HZM14, HYS17, HCW^{+}18, HYS18, HXC^{+}18, HYF18, HAA^{+}16, HSS17, HKG^{+}16, HHS^{+}18, IASK14, IS18, IOV^{+}18, IFD^{+}19, ISS^{+}15, IdAP19, JKAU19, JCMPPC^{+}18, JGFB18, JGB19, JXC^{+}19, JH16, JHC18, JL14, JLS19, JLQZ18, JFZL17, JSMG18, JLY^{+}18, JTL^{+}19, JLX^{+}19, JXZ^{+}19, JCL^{+}19, JEB18, Jun18, KZA^{+}18, KNI^{+}18, KMB^{+}17, KVK^{+}18, KKN18, KRD^{+}19, KKB14, KSS19, KH19, KHMB13, KOT18, KZA11, KKB^{+}19, KRZ12, KID^{+}16, Kim18, KPJ19, KDG^{+}19, KSC^{+}19, KGS^{+}19, KX11, KXS^{+}16]. based [KLJS19, KARP14, KMK^{+}14, KS17b, KCH^{+}13, KB16, KLV^{+}18, KLW^{+}17, KP18, LBD18, LCH^{+}11, LLYW19, LA19, LLpC12, LJ17a, LKCS18, LY18a, LCL^{+}19, LJ17b, LBJ^{+}18, LBJ^{+}24, LYJ10, Li10, LK12, LZXW13, LCL^{+}16, LLC^{+}16, LSZ^{+}16, LYYY17, LZL^{+}17, LFL^{+}17, LXL^{+}17, LCW^{+}18, LYYY18, LOR^{+}18, LZP^{+}18, LYW^{+}18a, LJJ18, LNK^{+}18, LWK^{+}18, LZY^{+}19b, LZZ19, LTMW19, LSCL19, LXRS19, LLZ^{+}19, LZH^{+}18, LZT^{+}19, LZL^{+}19a, LLL^{+}19, LLW^{+}19b, LSYC18, LN18, LXMW15, LWH^{+}18, LNLA19, LRMS19, LSH^{+}11, LLJ^{+}11, LWX13, LC13, LLS^{+}14, LvW14, LYL15, LCMX16, LWL^{+}18, LDS^{+}18, LSL^{+}18, LXM^{+}18, LWYS18, LCH^{+}18, LHW^{+}18, LLGY18, LW18b, LNY^{+}18, LGZY18, LXT^{+}19, LMM19, LWXY19, LZJL19, LDZW19, LMCSE19, LBU^{+}10, LKJ17, LWW^{+}13, LL16, LWF^{+}17, LQF19, LWR^{+}19, LZY^{+}16, LYH^{+}19, LWZ^{+}19b, LCY^{+}19b, LHY^{+}19, LZWF19, MWQ^{+}14, MML^{+}18, MWL18a, MYHZ18, MWQ^{+}19, MFC^{+}19, MSS^{+}13, MJM^{+}16, MCN^{+}18, MLC^{+}18a, MK17, MBB10, MHC14, MLC18b, MCT^{+}15, MBM18, MGV^{+}18, MHdlS19]. based [MS19, MDB^{+}18b, MSBA16, MPI^{+}18, MLM16, MHZK18, MZL^{+}19, MQL^{+}19, Mér17, MBL^{+}19, MZYA19, MLW^{+}18b, MJRM16, MGA^{+}18, MCdA16, MSE19, MAPA19, MDM^{+}19, MYK16, MND^{+}19, MW12, MWMA10, MCG^{+}15, MRS18b, NSR^{+}19, NV11, NNRA19, NRV^{+}17, NCS12, NS10, NK18, NGB18, NLM^{+}16, NUPA19, NZL^{+}15, NJKH13, NLS19, NAAC19, NAM^{+}19, NLV^{+}19, OdOD^{+}13, OMPSPL^{+}19, OMD^{+}18, OBG^{+}18, OB19, PZC19, PYH^{+}18, PNGFJ13, PBV^{+}13, PFRC16, PC18a, PVGD^{+}19, PS10, PKY^{+}17, PC17, PWP^{+}18, PLW^{+}19, PBC^{+}11, PMCC18, PGCML^{+}19, PLGMCdF18, PCK19, PAB^{+}14, PQBP17, PSY^{+}19, PVHTP19, PS19, PTT12, PPS^{+}18, PB18, PSBB15, PSK^{+}10, PKA19, QZD^{+}18, QRW^{+}18, QGT^{+}18, QZM^{+}18, QCY^{+}19, Qur19, RCM17, RQN^{+}19, RBA17, RGAT18, RYH^{+}19, RDSA18, RC13, RRP^{+}14, RGGH18, RT15, RS17a, RD14, RJS^{+}19, RSY^{+}18, RGM^{+}19, RGSL18, RCMT18, RW18, RM19]. based [RCTY19, SR12, SSHC19, SMF^{+}19, SME^{+}19, SCL18, SPT^{+}18, SC19, SLTK19, SS13, SB19a, SAGL10, SGGCR^{+}16, SGRT19, SYJ^{+}19a, SYJ^{+}19b, SHS^{+}19, SMC18, SD18, SAM^{+}19, SJR13, SSST17, SZ12, STA17a, SP18b, SRP19, SBLW14, SYT^{+}19, SHL^{+}19a, SJ19, SSL^{+}19, SCZ^{+}19, SMSF18, SH19, SS17, SB17a, SGB^{+}18, SB18, SGBK19, SSKK13, SCN^{+}14, SIL^{+}13, SAPA17, SCL14, SCMS12, SMM^{+}14, SLS^{+}19, SKS17, SZD^{+}17, SM18, SYK^{+}17, ŠCJ^{+}19b, SSSJ19b, SCZ^{+}14, SBK18, SMZ^{+}16, SSZ^{+}17, SLL^{+}17, SLD^{+}18, SZR18, SJL^{+}18, SCH^{+}19, TTC^{+}14, TX14, TF18, TJWS10, TOD17, TZL^{+}18, TMB^{+}19, TM19, TNY17, TMM^{+}13, TKA^{+}18a, TA18, TY11, TMDZ15, TZD^{+}19, TCR^{+}12, TYWZ18, TCCW19, TIHT14, Tur18, URC19, UGBM^{+}17, UPP17, VVB11, VTTK17, VPP^{+}19, VSKS19, VCL^{+}19, VSBN19, VOCHC17, VKK14, VKT^{+}19, WWX^{+}17, WCHL10, WW11, WLYL11, WWCN13, WWD^{+}14, WLZ^{+}14, WZZ16, WMX^{+}17]. based [WWVJ17, WRCC17, WCL^{+}17b, WWDF18, WZH^{+}18, WCB^{+}18, Wan18a, WJS^{+}18, WHCZ18, WWZZ18, WCW18, WZWC18, WWSL19, WLL^{+}19a, WGC19, WDZ19, WY17, WZL18, WMA18, WZ18, WFQ^{+}10, WW13, WZ13, WYH^{+}17, WZCH17, WLXZ18, WPS^{+}18, WZML18, WWZ18, WWA19, WLZ^{+}19, WHCW19, WSY^{+}19, WBR19, XTL^{+}19, XFJ^{+}19, XHW19, XWJ^{+}16, XKBA18, XZW^{+}19, XKJ^{+}18, XCGD10, XL19, XZZ^{+}19, XZP^{+}19, XLL^{+}14, XZ14b, XY15, XWL^{+}15, XZ16, XFM16, XWX^{+}17, XRPT18, XTT18, XZZ^{+}18, XJY^{+}18, XTZ^{+}19, XLL^{+}19b, XHL^{+}19, XXX^{+}19, XLL^{+}19c, XJZ^{+}19, XCZ^{+}19, YMLT13, YAO14, YMW^{+}18, YWJ^{+}19, YFY^{+}13, YZLQ14, YWL^{+}17, YLWW18, YJS18, YZW^{+}18, YWZ^{+}18, YXA^{+}18, YWJ^{+}18, YJY^{+}18, YZL^{+}19, YHL^{+}19, YLL^{+}19, YCT15, YSZW18, YARH18, YPJ19, YW12, YP12, YN18, YXY18, YS16, YZ12, YSL19, YPHZ14, YXA^{+}16, YWY^{+}17, YL18, YGY^{+}19, YJL^{+}19, YWF^{+}10, YMY^{+}17, YDQC19, YL16, ZF16, ZRZL18, ZSJ19, ZXW19, ZZF^{+}19, ZA13, ZWL13, ZSZ14, ZZH^{+}16, ZWHC17, ZCH^{+}17, ZCX^{+}18, ZCLW18, ZHL^{+}18, ZXL^{+}18]. based [ZXZL18, ZZXL18, ZWWL18, ZZBP19, ZXM^{+}19, ZLT^{+}19, ZZZC19, ZTL^{+}19, ZZJY16, ZYC^{+}19, ZDL^{+}19, ZBL^{+}14, ZLG^{+}14, ZLL^{+}16, ZCZ^{+}18, ZFC17, ZFC18, ZWX^{+}19, ZZC14, ZDW^{+}16, ZCDV19, ZLY^{+}19, ZZS^{+}19, ZWJ19b, ZRZ^{+}14, ZaTZ^{+}17, ZL18, Zhu18, ZWGC19, ZMN19, ZCW19, ZDW^{+}18, ZWMC19, dSBN19, KK19, ASAAM^{+}19, KKK^{+}19, PMDS18]. based-wireless [HKA^{+}18]. bases [ABTF16, GZS14]. basis [TSBH11]. Bat [Zin18]. Batch [SS13, WSQ^{+}16, AQB15, GZQ^{+}19, JLL17, SVN10b, TF18, VCD^{+}18, WWZ^{+}19]. batch-oriented [AQB15]. batching [LZCX19]. battery [CFMC19]. battlefield [WSZC18]. Bayesian [CFL^{+}15, CAPG18, KIMR15, LYYY17, TAHS14, WWA19, ZSGJ19, ZCW19]. BC [DKJ19]. BCN [ESPN17]. BDEv [VEET18]. BDP [PPA18]. BDWatchdog [EET18]. be [SOA17, WWG^{+}19b]. Beaconing [YLWW18]. bearing [CRWZ19, RCW^{+}19]. BECloud [Bel16]. Bee [HLZ18, CZL^{+}18b, HXWW18, SMS^{+}19, XLW^{+}17]. BeeCup [XZZ^{+}14]. before [MAJD18]. Behavior [CJHH13, BDWM17, CMEA^{+}19, CYW^{+}19, FFGP^{+}19, FM10b, LRJG19, Lee12, LNB14, LJW^{+}19b, NLLC19, RWV^{+}13, SHRE16, SZ12, TIHT14, WSU^{+}10, WWH^{+}17, YHH^{+}19, ZCZ^{+}18, ZWJ19b]. behavioral [CPP^{+}18, CWJ16, CT19c, KYZ19, KMK^{+}14, NJ17, SBCF16, SSA^{+}19, WNR19]. behavioral-based [KMK^{+}14]. behaviors [ABD^{+}19, WTG^{+}14]. behaviour [HIA^{+}18c, PRW14]. Behavioural [RT16, HAAR^{+}19, PLLA18, PLA18]. Behind [DMZ12]. Belief [AZO^{+}19, KLJS19]. belt [TTC^{+}14]. benchmark [ABG18, ZYB^{+}18]. Benchmarking [DFGR14, GVDT16, MMLO18, PPPS18, SZV19, TDBR18, Wri19, CCD^{+}10, ECA^{+}18, LSB^{+}18, TZQ18]. benchmarks [GMB19]. Benders [FBM19]. benefit [MEBA12, SC16]. benefit-aware [SC16]. benefits [KK14, PNGFJ13, SC19, VVB11]. benign [WLW^{+}18]. Best [KZ17, KS19]. beta [DJH^{+}19]. better [FTD17, Mér17]. between [BFP18, BDA19, BCF16, DGCGH^{+}17, HA18, LZCX19, LZL^{+}16, MFL18, NJKF18, PWP^{+}18, XLZ^{+}14, YAO14, YZW^{+}18, YTQ19, YTQ20]. betweenness [MBM18, MCSA18]. Beyond [KSK^{+}11, LWW^{+}13, PN13, Eng14, KDHP16, RR18]. BG [ABG18]. BGP [IGB^{+}14]. BGP-inspired [IGB^{+}14]. BGQoS [AJY15b]. BHR [ST11]. Bi [ABS^{+}18, CGSJ18, WWC14, ZZH^{+}16]. Bi-directional [ABS^{+}18]. bi-level [WWC14, ZZH^{+}16]. bi-objective [CGSJ18]. biased [PdASM18, ZSFZ19]. Bicycle [ZMN19, ZWQ^{+}19]. bidding [WCHL10]. bidding-based [WCHL10]. Bidirectional [MAÇ17, WHBC19]. Big [AAA^{+}19, ASYF18, BA17, BAJ^{+}19, CMEA^{+}19, CCRL18, DP20, DP21a, DP21b, EET18, GBB18, JGFB18, Jun17, KS17a, KPB18, KLH^{+}18, LIC18, MGL^{+}18, NNLH18, PN13, PPA18, RPA^{+}18, SNXB17, SCG^{+}18, VEET18, WSQ^{+}18, WRK^{+}15, WJS^{+}18, WXGM18, YDD^{+}18, YZI18, ZCYZ18, ABD^{+}19, ATH^{+}19, ATM^{+}19, ACSV18, BGI14, Bha18, Bu18, CQW^{+}19, CGIP14, CCMP18, CZXL18, CRYG18, CZZ^{+}18, CLS19a, CDB^{+}19, CSQL17, DJPM18, DP19, DQXW19, DXL^{+}18, DZLA19, EAS^{+}18, FRM^{+}18, FS18, GQLX18, GLC19, GBY16, GSY^{+}17, HPGMM18, HZDS19, Hsu14, kHsZwJW18, HZ19, HHZ16, HXL^{+}18, IAM^{+}18, JH16, JHC18, JLC18, KOP^{+}17, KP18, LFP^{+}17, LLZ^{+}18a, LZH^{+}18, LJ19b, LWH^{+}18, LYZC15, LSHW17, LHW^{+}18, LWR^{+}19, MWW^{+}15, MLC18b, MVL^{+}18a, MPCAF15, MGMT18, NJ17, OFD17, Osm19, PFRC16, PvSS17, PYM18, PPPS18, QGX18, QCX18, RTHB17, SZV19, SB17a, SBA^{+}17, SBD^{+}18, SZD^{+}17, SSZ^{+}17, SJSA19, TZQ18, TF18]. big [TBR^{+}19, TWZP18, WXPL17, WMY^{+}18, WZW19b, WLHH18, XB14, XFM16, XLL^{+}19a, XTF^{+}19, YWCC18, YZG^{+}18, YWLL19, YSHM19, ZYB^{+}18, ZFS^{+}18, ZHL^{+}18, ZWZ18, ZQB^{+}18, Zhu18, uRYS^{+}19, PPS^{+}18, AKP^{+}18, BBD^{+}19, CRW^{+}16, CW16, DX14, GS16a, KKAS19, MCR^{+}16, Sha16, ZS16]. big-data [BGI14, ZYB^{+}18, ZQB^{+}18]. Bigdata [WLZ^{+}19]. BigFCM [GGN17]. BigFlow [VSBN19]. BIGSEA [AAA^{+}19]. bike [TCH19]. bile [HZX^{+}19, HZX^{+}20]. bilinear [ZZ15]. bill [MD12]. billing [MK17]. billion [CCZ^{+}19]. billion-scale [CCZ^{+}19]. Bin [BB17, LZCX19, MAA^{+}19, SK12]. bin-packing [LZCX19]. Binary [GK18, SBCF16, HXWW18, JLS19, SMS^{+}19]. binding [MLL15, MMVS19, MLBS11]. Bio [FM10a, KR19, Kim18, PMBS14, XZZ^{+}14, ZYW^{+}18]. Bio-inspired [FM10a, XZZ^{+}14, ZYW^{+}18]. bio-medical [PMBS14]. bio-security [KR19]. bio-signal-processing [Kim18]. biochemical [HZX^{+}19, HZX^{+}20]. BioClimate [FEB^{+}19]. biocollections [MMF16]. Biodiversity [FEB^{+}19]. biofeedback [BDE17, KMU19]. biogeography [ZLL^{+}16]. biogeography-based [ZLL^{+}16]. Bioinformatics [DP17, CRSdS10, EDH^{+}13, KTTK17, MMC^{+}18, SBAD^{+}18]. Bioinspired [CCMGF18, CLCMG^{+}18]. biological [PB17, SWW^{+}13]. biology [CSdCM^{+}17]. biomarker [WZH^{+}19]. Biomedical [CRSdS10, AIA^{+}18b, CMD^{+}14, DMPP16, DZLA19, KTTK17]. Biometric [Alp18, BCN^{+}19, BDM^{+}19, KHMB13, AHM^{+}18, BK19, BWR12, CCL19, DEL19, GEAR13, Ham19, KZA11, LTJK12, LK12, NWMG17, Sar18a]. Biometric-based [BDM^{+}19]. Biometrics [BW13, FFC12, AGBR19, FHZW18, Kha12, KS18d, KLW^{+}17, LNK^{+}18, TA19]. Biometrics-based [FFC12, FHZW18, KLW^{+}17]. biopsy [BDS^{+}10]. bipartite [QGX18, TJ18]. birth [PSI19]. bit [KHJ10, YFY^{+}13]. bit-rate [KHJ10]. bit-store [YFY^{+}13]. Bitcoin [HCW^{+}18, HS19, VSM^{+}19, WLGL19, ZMH^{+}18]. Bitcoin-based [HCW^{+}18]. bitmap [CH10]. bitonic [GDAS18]. BitTorrent [AY16, WSU^{+}10, WLQ10]. Black [VVB13b, CCL19]. blackbox [CLL^{+}14]. blackbox-oriented [CLL^{+}14]. blacklisting [OWX19]. Blanc [OBG^{+}18]. blank [FZT^{+}18]. blind [CJS19, LLW^{+}19a, ZaTZ^{+}17]. Block [HS19, CLR18, MHY^{+}18, RP18, YLL^{+}19, ZSW^{+}18a]. Blockchain [CLC^{+}19, MJGW18, SP18b, SSSJ19b, ABB^{+}19a, Che18, CLS19a, CLS^{+}19b, DKJ19, FSY^{+}19, FY19, IFD^{+}19, JLX^{+}19, KS18b, KGS^{+}19, LBJ^{+}18, LBJ^{+}24, LML^{+}19, LXL^{+}19, MKS^{+}19, MQN19, QNM^{+}19, RMC^{+}18, SI19, SSL^{+}19, Sko19, URC19, XLL^{+}19b, YZL^{+}19, ZWX^{+}19, ZWGC19]. blockchain-aided [Sko19]. Blockchain-based [SSSJ19b, FSY^{+}19, IFD^{+}19, JLX^{+}19, XLL^{+}19b, YZL^{+}19, ZWGC19]. blockchained [LTMW19]. blockchains [HZLH19]. Blog [CC19, ZWWL18]. blood [QXZ^{+}19, RAA^{+}19]. blood-cooling [QXZ^{+}19]. bloom [ZL13, ML17, SGB^{+}18]. Blue [KBVH14]. blueprint [JYY^{+}17]. Board [Ano19b, Ano19c, Ano19d, JBP^{+}18, Ano11c, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano14e, Ano14f, Ano14g, Ano14h, Ano15k, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano16v, Ano16w, Ano16x, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano19a, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j]. Body [AS18a, FPPD14, ASO14, AKM18, CJ14, FP14, HLYW17, IASK14, SCS^{+}18, THA^{+}17]. body-sensor [ASO14]. BodyCloud [FPPD14]. BOINC [FKBG10, VKK14]. bone [LZT^{+}19, YTQ19, YTQ20]. BonFIRE [GCV^{+}14]. Boosting [CPW19, GCBM17, KMV^{+}15, MWYC12, LMCSE19]. booting [SPSP17]. bootstrap [DDJ^{+}13]. bootstrapping [PGCML^{+}19]. borne [SM18]. Bot [KMST19]. Bot-IoT [KMST19]. botnet [KMST19]. Bound [MGMT18, BMT12, VD16]. Boundary [AS18b]. bounding [PYH^{+}18]. bounds [LSG18]. box [CCL19, CSG^{+}18, VVB13b, ZSW^{+}18a]. BPM [SJV^{+}15]. BPMN [GHLW18]. BPSO [KP18]. brain [ABS^{+}18, ASYF18, DSM^{+}19, GRS^{+}19]. brain-computer [ABS^{+}18]. Branch [MGMT18, PGCC^{+}10, BMT12, WZF^{+}19, VD16]. Branch-and-Bound [MGMT18, VD16]. breaches [NAM^{+}19]. Breast [SYT^{+}19, KIJ^{+}19, WZH^{+}19]. BRelax [SSG17]. bridge [SB16]. Bridging [Bha18, DHC^{+}17, NJKF18, SG17, KH19, dCRL^{+}19]. brighter [SM10]. Bro [KBdLG18]. broadcast [HZ10, WSQ^{+}18, WDZ19, XZP^{+}19]. broadcasting [LRZ^{+}18, YY11]. Broker [AWN^{+}13, BS11, FA11b, MSBA16, MGG^{+}17, PKY^{+}17, RGC^{+}10, SAK^{+}10]. brokerage [CCIP18, OG18]. Brokered [BB17]. Brokering [ACCD17, CS12, TMMVL12]. Brownian [GK18]. browsers [AAI^{+}19]. browsing [FD12]. BTEM [AZO^{+}19]. budget [ABP16, CXL^{+}17, WWT^{+}16]. budget-deadline [ABP16]. buffer [SCH^{+}19]. Buffers [YZI18, dSCD^{+}19]. build [BRXdS11, RAdARP19]. Building [BH13, FKOC11, HLT^{+}18, KTTK17, PWB^{+}13, RR18, dVXB^{+}11, ABMM18, HSS17, LRL^{+}14, LCY19a, PVHTP19]. buildings [PPS^{+}18]. built [GZS14]. Bulk [GGLD10, dRRdQGR^{+}18]. Bulk-Synchronous [dRRdQGR^{+}18]. Bumper [DR15]. Burst [YZI18, KK10b, ZCS^{+}16, dSCD^{+}19]. bursty [GHMX10, TCBPR16]. Bus [HHXL13, CCC19]. bus-assisted [CCC19]. bushfire [GAW^{+}18]. Business [AJY15b, ACC^{+}16, BGRBA19, Cha14a, DDMPG17, FG14, JC15, QCD16, SJV^{+}15, BBT19, Cha14b, CKR16, CYJ19, KYZ19, LCCM18, MCL^{+}16, MVG^{+}14, MAB^{+}15, MCWP16, WBR19, YNSM12, YJL^{+}19, ZCX^{+}18]. Business-driven [FG14]. business-oriented [YNSM12].

c [DNW^{+}19, BEWZ10, MDT^{+}18]. cache [CXC^{+}18, DZZ^{+}15, HCZW17, LTTL19, MVL18b, PRS^{+}13, Tao10, ZWL^{+}16, ZCQ^{+}16]. caches [YDT19]. caching [CXC^{+}18, CdRRdCB19, GSP^{+}17, HMA18b, LW18a, LSG^{+}19, RHMGC14, TGM^{+}19a, XWX^{+}17, ZCLW18, ZLT^{+}19]. CAD [LHY^{+}19]. Calculating [SSRQ19, YAA^{+}19]. calculations [ML11, RMHMG17]. calculus [KAP19]. calendar [YSL19]. calendar-based [YSL19]. calendaring [AHP^{+}18]. Calibers [AHP^{+}18]. calibration [HLZ18, JCA^{+}19]. calls [PZC19]. calm [SCH^{+}19]. CalmWPC [SCH^{+}19]. camera [GWW^{+}19, YJS18]. CAMP [HRR^{+}14]. campus [GLD^{+}19a, MFG^{+}14]. CAN [KR19]. cancer [ABM19, JLC^{+}20, KMK^{+}19, KIJ^{+}19, LYYW19, WZH^{+}19, WWA19, YWG^{+}19, YWG^{+}20, wZcZN^{+}19, wZcZN^{+}20]. cancerous [LJW^{+}19a]. CANF [FJ18]. canonical [Bu18, GLC19]. Canopy [LZY^{+}19b]. Canopy- [LZY^{+}19b]. can’t [GMCM16, GMCM18]. capabilities [BBWB^{+}18, HMZ18]. capability [ZCL^{+}18]. capable [PvSS17]. capacities [JLL17, Li15]. Capacity [BB17, CMB17, HMH17, KIC12, RZ16, Bha18, CWL^{+}18, GBRM18, KCM19, KMJ18, dACNC16]. Capacity-driven [RZ16]. capillary [APRC16]. capital [NZOCJ^{+}19]. capsule [KHO^{+}19]. capsulorhexis [LSW^{+}19]. CAPTCHA [GPV^{+}14, RM16]. caption [DQXW19]. capture [Asu13]. Capturing [BCG^{+}19, NHH^{+}19]. car [KXS^{+}16, LLN^{+}18]. carbon [VVC^{+}12]. carcinoma [MGA^{+}18]. cardiac [CAC^{+}10, DNW^{+}19, MdMMNS^{+}19, SWW^{+}13]. cardiological [KKS^{+}18a]. cardiomyopathy [DNW^{+}19]. cardiotocographic [LQF19]. cardiovascular [NDA^{+}19]. cards [CHS11, NWMG17, SGGCR^{+}16]. care [ABC^{+}18, CAS^{+}18, KIJ^{+}19, MRS^{+}18a, NDZ^{+}18a, NDZ^{+}18b, NDZ^{+}19, SPS18, BS11, SAK^{+}10, BGMLS17]. Carlo [CPGdS^{+}13, CPGBC16, DDJ^{+}13, JH16]. carpooling [HKG^{+}16]. carrier [LML^{+}19]. cars [MHW^{+}16]. CartoonPlus [JHC10]. cascade [GJY18]. cascaded [HHH^{+}19]. cascading [HA19]. Case [LLES19, MW12, SGN^{+}17, AR18, AMÇ19, Asu13, APR^{+}19, CGN18, CLAL19, CFF14, DDMPG17, FGM11, HMA18b, MNC^{+}18, MKS18, MOFGP18, RRP^{+}14, SHS^{+}19, SB11, TuIS^{+}19, WLH16, WRCC17, XLL^{+}19b, ZDL^{+}13, ZXL14]. CASQ [LZY^{+}19a]. Catalan [SAM^{+}19]. catalogue [MRT^{+}19]. cataract [LSW^{+}19]. Catcher [BSE^{+}13]. categorizing [WLW^{+}18]. Category [LZY^{+}19b, GGLW18, ZH17]. Causal [HKS18, MFT^{+}17, TSWL17]. cause [CPP16, LWR^{+}19]. caused [MHW^{+}16]. cavities [DNJG17]. cavity [HLZ^{+}19]. CCA [ZSW^{+}18b]. CCA-secure [ZSW^{+}18b].

CCBKE [LZYC13]. CCGrid [KZ17]. CCTV [LHJC18]. cDBN [Jun18]. CDN [LLT^{+}19]. celiac [KHO^{+}19]. cell [DPK^{+}19, EAA16, JLC^{+}20, YWG^{+}19, YWG^{+}20]. CellML [CAC^{+}10]. CellML-based [CAC^{+}10]. cells [JLC^{+}20, YWG^{+}19, YWG^{+}20]. Cellular [TA18, TZD^{+}19]. CenLocShare [XCS^{+}18]. Center [BMK^{+}14a, SE19, GDS18, HDB18, KAEC^{+}18, LN13, LWD^{+}14, LGL^{+}17, LGZY18, LWW^{+}13, Man15, MP17, NKP16, SPT^{+}18, VOS12, XDH^{+}17, YPLZ17, YGYW16, ZGL19, ZWJ^{+}18]. centered [ABZK15, AB19b, JNS^{+}19, NRR^{+}15]. centers [AD19, AAAQJ^{+}18, AK18b, BAB12, DKK^{+}13, FNCR11, GS16b, GBF^{+}12, GFW^{+}18, KKBK19, LSYC18, LWW^{+}18, ODC19, Pon19, QMSG12, Qur19, RT15, RMRSA19, SLD^{+}15, SLB^{+}17, SLZ^{+}18, TDC^{+}14, VTTK17, WTR^{+}13, WWQ^{+}18, ZZH^{+}16, ZXD^{+}19, ZAC^{+}18]. central [SZG^{+}19]. Centrality [LG18, AdVAGF18, GJY18, LXM^{+}18, MBM18, MCSA18]. centralized [XCS^{+}18]. centred [LASL16]. Centres [CAB^{+}18, WCH^{+}18]. Centric [WSQ^{+}18, APBdI17, ACL^{+}18, AT18a, AHM^{+}18, BS17, DGR^{+}15, DMM^{+}18, DGdL15, GZLZ16, GMCM16, GMCM18, JTS13, KYZ19, LKN^{+}13, PIP18a, QGT^{+}18, RLL^{+}17, SST18, SZK18, SRKS18, TGM^{+}19a, TGM^{+}19b, XLL^{+}14, ZW10]. CEPSim [HCB16]. CERN [NNLH18]. Certain [YH18]. certificate [LL16]. certificate-based [LL16]. certificateless [LHO17]. certification [CB10]. certified [LH13b]. CFD [LZX16, OBG^{+}18]. CG [KLP19]. CG-E2S2 [KLP19]. CGLXTouch [PDW^{+}11]. chaff [KHMB13]. chain [ABMM18, AB19a, ABGMC19, AB21, BDA19, BDNP13, CLL18b, Che18, HHXL13, JLQ18, LBJ^{+}18, LBJ^{+}24, SLL^{+}18, YLL^{+}19, ZSL^{+}19a, FY19]. chain-generated [HHXL13]. chaining [SLY^{+}19]. Challenge [KZA11, STMV18, SGM11]. Challenge-response-based [KZA11]. Challenges [CCRL18, CDFW18, GS16a, GRL11, JC15, Jun17, KMU19, MWW^{+}15, PWA^{+}19, YS16, AUSA19, AR17, ALK15, AC10, ALL^{+}18, CAC^{+}10, CRSdS10, CCIP18, CSC18, CBBC^{+}17, EGVT18, Eng14, FFC^{+}18, FJKK17, GFD14, KS18b, KAW12, KLH^{+}18, KARP14, NS10, OFMZ18, PYM18, PDH18, RLM18, SP18a, SJV^{+}15, SWY^{+}18, URC19, UKK^{+}19, VCE^{+}19, WWVJ17, WGX^{+}19, YHA^{+}19, RMC^{+}18]. Chances [ALL^{+}18]. Change [FEB^{+}19, HKT^{+}19, PWB^{+}13, ZXW19]. changes [BNJ16]. channel [ABS^{+}18, AAQ^{+}19, CWW^{+}16, DJH^{+}19, GQXL18, SWW^{+}13, SHL^{+}19a, YZC^{+}19, YN18]. channels [GHMX10, KKB^{+}19]. Chaos [AIA^{+}18a, XRPT18]. Chaos-based [AIA^{+}18a]. chaotic [Bro19, KLW^{+}16, LWK^{+}18]. character [LJ19a, TJ18]. Characteristics [WCWC19, WCWC20, YXY18, MPP13, SHL^{+}19a]. Characterization [AHD^{+}19, BARMB14, BZHV19, GPVN19, KRD^{+}19, PLGMCdF18, VEET18, dSFP^{+}17]. Characterizing [GdVC10, JTB13, JCD^{+}13, MSS^{+}13, WWH^{+}17, MMAA19]. Chargeback [BKS^{+}14]. chargers [CZM^{+}18]. charging [HZLH19, RYH^{+}19, YHL^{+}19]. chatbot [CPMG19]. cheating [DLMS15]. check [Che13b, LZL^{+}17]. checking [LXJD18, PSS13, RWV^{+}13, SYY^{+}17, YZN^{+}15, YXA^{+}16]. checkpointing [CPGdS^{+}13, CRVZ15, FMSSM12, FRB^{+}14, KKJJ10, LBB^{+}19, LM12, RPMG10]. chemical [CHWW13, YWLL19]. chemistry [ZDL^{+}13]. Cherenkov [CSV^{+}19]. children [XL19]. China [HPZL18, PWP^{+}18]. Chinese [TJ18, ZWWL18]. Chip [MBM18, BFP18, CPSD18]. chips [SSG19]. CHIS [CCMP18]. choice [SC16]. choking [AY16]. cholangiocarcinoma [ZMZ^{+}19, ZMZ^{+}20]. cholangiopancreatography [HZX^{+}19, HZX^{+}20]. chromosome [hKBB11]. chronic [HEES19, VFHB14, WCWC19, WCWC20]. chronological [MPP13]. Chunk [ZZJ17]. chunking [WLA17a, XFJ^{+}19]. CineGrid [DGdL15, GHO^{+}11, KGdL11, LSH^{+}11, WdL16]. CIoT [GTEL^{+}18, ZZX^{+}19, ZZLZ18]. cipher [ZSW^{+}18a]. ciphers [MHY^{+}18]. Ciphertext [JSMG18, Rao17, LDZW19, QRW^{+}18, LAL^{+}15, LHL15]. Ciphertext-Policy [Rao17, JSMG18, LDZW19, QRW^{+}18, LAL^{+}15, LHL15]. ciphertexts [WXLY16]. circuit [AEK^{+}18, BDG^{+}19]. circulating [JLC^{+}20, YWG^{+}19, YWG^{+}20]. circulation [WJZ^{+}17]. circumstance [ZLY^{+}19]. circumvent [DC17]. Circumventing [DC18b]. citation [HHL11, HQ10]. citations [CZ14]. citie [AT18b]. Cities [HSS17, RMSPP17, Sta17b, APBdI17, AR18, AK19, BAKB19, DGR^{+}19, HCZW17, HMA^{+}18a, HMA18b, KAS^{+}18, KS18d, NWL17, Osm19, RGSL18, SRdlPG19, SB19c, TLL^{+}19, UGBM^{+}17, GMLGB^{+}17]. Citizen [APBdI17, AB19b, ALFR16, NZOCJ^{+}19]. citizen-centered [AB19b]. Citizen-centric [APBdI17]. City [CMNK19, MPI^{+}18, RPA^{+}18, BBC^{+}17, CCC19, DDMPG17, EAED18, GAI^{+}18, KPA17, KFK19, LNK^{+}18, LSV^{+}18, LCY19a, LCY^{+}19b, LLW^{+}18b, PC17, SP18b, TWZP18, TCB^{+}17, WHBC19, XYLZ18, YJS18, dSK^{+}19, CGSV17, FAMA^{+}17, hKRM17, UPP17]. civil [AB19b]. civilian [MND^{+}19]. CL [SVK19]. CL-ADSP [SVK19]. Clairvoyance [BOP^{+}14]. CLAPP [GNGG17]. class [DKK^{+}13, GWW^{+}19, GS13, GD10, RHH^{+}19]. class-statistic [GWW^{+}19]. classes [CMB17, JOPW14, PSLZ18]. classic [MCA^{+}18]. classical [MAÇ17]. Classification [BD18, MWQ^{+}19, ZXM^{+}19, AHMS18, AAM^{+}19, AGA18, CMI^{+}19, CLY14, CTU19, DDD^{+}19, DH16, EP13, FTK^{+}14, FMV14, GLVC18, GOLL19, HZW19, HMA18b, JOSD19, KMK^{+}19, KIJ^{+}19, hKBB11, LXL^{+}17, LXT^{+}19, MLG13, MG14, MdMMNS^{+}19, MHdlS19, MCRB19, MZYA19, MRS^{+}18a, NUPA19, PRW14, RRKA19, RGGH18, RD14, RHH^{+}19, SD18, TSRG17, XZZ^{+}18, YARH18]. classifications [Bag16, CWUS19]. classifier [KLV^{+}18, ZSQ^{+}19]. classifiers [IdAP19, KI19, WLW^{+}18]. Classifying [BCMM18, DSM^{+}19]. classroom [Kim18]. clean [MRL14]. CLEM [CJN^{+}15]. Clickstream [FFGP^{+}19]. Client [ZW10, BK19, CSL18, DSD^{+}11, JLX^{+}19, KuRAk^{+}18, MG14]. Client-centric [ZW10]. client-server [BK19]. clients [LH13b]. cliff [SCH^{+}19]. Climate [FEB^{+}19, BNJ16, FQBCF15, MLC18b, PWB^{+}13, FNA12]. Climate-G [FNA12]. Clinical [WWP19, WWP20, CPE^{+}17, JNS^{+}19, LDY^{+}18]. cloaking [NZL^{+}15]. clock [FSY^{+}19]. clone [SMRM13]. CloneSpot [MH19]. CloRExPa [DLS14]. closed [HXY13]. closed-form [HXY13]. CLOSER [CCRL18]. cloth [HX19]. clothing [HYC^{+}18]. Cloud [AJR^{+}19, ASAB^{+}18, AASI17, ABTA18, AAJ17, AM17, AKB^{+}18a, ACCD17, AMPZ16, BB13, BDL^{+}19, BdDPP16, BCJT13, CLRL18, CHWW13, CRM^{+}16, CDL^{+}16, CPD^{+}15, Cha14a, CKR16, CDFZ16, DMC^{+}19, DXA14, DLS14, FMN^{+}17, FSV^{+}19, FPP^{+}18, GAW^{+}18, GJ15, GBRM18, GTCZG^{+}18, GMP^{+}17, HMH17, HMM18, HSBE19, Hel16, HM19, HXA^{+}17, HMA^{+}18a, IASK14, JC15, JY15, JGB19, JTL^{+}19, JTBS15, KKK^{+}19, KIMR15, KLV^{+}18, KJ18, LSB^{+}18, LSD^{+}17, LOR^{+}18, LJGW18, MID16, MP17, MAB^{+}15, MM18, Pal16, PN13, PLLA18, PDH18, RAA^{+}18, SV16, SB19a, SST18, SGN^{+}17, SS17, SE19, SZK18, SAPA17, SPKG18, TMW^{+}17, TMMVL12, VLAC^{+}13, WBKL16, XDWL15, YG18, YHL16, YWZ^{+}18, YARH18, YAP16, YXA^{+}16, ZAC^{+}18, ABZK15, ABMC18, ANA16, ABDH19, AA18, ALTG19, AD19, ANG^{+}19, AAAQJ^{+}18, AFSH^{+}18, AFSH^{+}19, AMQS^{+}19, ABB^{+}19a, AEK^{+}18, ADA^{+}19]. cloud [ASO14, ALK15, AL18, AMGCC18, ABP18, AMMC18, AMKM18, ATM^{+}19, Ano12r, AB16, AAD^{+}13, ABN17, ADBO18, AMS19, ADDV16, ALL^{+}18, AK18b, BKS^{+}14, BDE17, BFN18, BMU16, BGC^{+}19a, BCN^{+}19, BKS^{+}18, BFS^{+}17a, BFS^{+}17b, BC15, BBCN18, Bel16, BBT19, BPC^{+}14, BT17, BGRBA19, BB12, BDH14, BR10, CMX^{+}16, CA15a, CVT19, CDG^{+}14, CLZ18, CLAL19, CMB17, CCIP18, CTR^{+}17, CXDM18, CFG^{+}19, Cha14b, CFL^{+}15, CRW^{+}16, CW16, CLL18b, CM17, Che13b, CWSW14, CLL^{+}14, CAC^{+}15, CXL^{+}17, CLR18, CCY^{+}18, CXWT19, CW13a, CJS19, CGSJ18, CGL15, CPSRG14, CR14, CBLS13, CSQL17, CDL18, DC18b, DZZ^{+}15, DJZ^{+}15, DRC^{+}19, DL19, DST14, DRGC^{+}19, DLLZ17, DMPP16, DEG^{+}17, DQLW15, DHL18, DLL^{+}19, DR18, DCC13, DWS12, DSPA18, DHC^{+}17, DCMW17, DJ13, EKSDN19, EBOY14, EAS^{+}18, Erd13, EP12, EMJ^{+}13, EPB18, ECA^{+}18, ETR^{+}13]. cloud [EA17, FH13, FHYH15, FCY18, FLT^{+}19, FLL^{+}19, FLR^{+}16, FSM^{+}18a, FJL^{+}16, FW19, FQBCF15, FLR13, FDPR17, FEPC18, FPGK18, FKT14, FM17, FP13, Frî14, FCD^{+}14, FSP^{+}18, GGTRRC16, GPJA^{+}14, GVBG17, GEG14, GdCP19, GVB13, GJGB19, GGJ13, GSLI12, GAJP18, GS16b, GJF^{+}12, GCCL18, GB10, GTSAR^{+}14, GPS13, GMP^{+}18, GJKP18, GFW^{+}18, GSR^{+}19, GLB^{+}18, GZQ^{+}19, GGC18, GGS13, HIA18a, HFM19, HSM13, HGG^{+}14, HQH16, HLL18, HCMJ19, HJA^{+}19, HHK18, HLYW17, HCNT14, HZL^{+}19, HB19, HHK^{+}16, HCB16, HNQ^{+}18, HZ19, HZW^{+}16, HYS17, HYS18, HHS^{+}18, IDCJ11, IS18, IKLL12, IGB^{+}14, JAAD^{+}16, JTB13, JTS13, JBR^{+}16, JCL^{+}15, JZJ^{+}18, JMAG19, JEB18, JOSD19, KS18a, KZ17, KC14, KRD^{+}19, KYB^{+}19, KMT14, KKKM13, KAW12, KCS14, KMJ18, KADJ14, KDG^{+}19, KSC^{+}19, KLH^{+}18, KACN16, KARP14, KMV^{+}15, KTTK17, KKKM17, KKKM18]. cloud [KGLY18, KS18c, KLW^{+}17, LBD18, LCH^{+}11, LZY^{+}19a, LCL^{+}19, LXD17, LLW^{+}12a, LJLW13, LCHW14, LXK^{+}14, LLH^{+}17, LPY^{+}18, LLF^{+}18a, LLWW18, LYY^{+}18, Li18, LWTL19b, LSCL19, LWTL19a, LAL^{+}15, LSYC18, LCL14, LXMW15, LHX^{+}18, LWW^{+}18, LZYC13, LYZC15, LYL15, LHL15, LCMX16, LPL^{+}16, LSL^{+}18, LDY^{+}18, LCH^{+}18, LNY^{+}18, LZLL18b, LYL^{+}19, LPBB^{+}18, LHBC16, LSJ^{+}14, MZH^{+}17, MG14, MG16, MK17, MGT18, MFN13, MHC14, Man15, MCT^{+}15, MYW^{+}19, MAJD18, MKK13, MCAS19, MS19, MBMTJR18, MDB^{+}18a, MKH13, MK19b, MFG^{+}14, MNY^{+}19, MSBA16, MLM16, MHZK18, MVG^{+}14, MRH17, MMVP13, MAA^{+}19, MCdA16, MSE19, MSM^{+}13, MYK16, MWL^{+}18b, MCG^{+}15, MBV^{+}15, MCWP16, Nag16, NRV^{+}17, NCS12, NNC^{+}19, NLV^{+}19, OdOD^{+}13, ODK^{+}17, PZA18, PBV^{+}13, PC18a, PVN^{+}12, PLA18, PVGD^{+}19, PFPJ18, PKY^{+}17, PPG19, PPPS18, PBA18, POJ^{+}16, PTM^{+}18, Pon19, PPB16, PKI^{+}18, QWCW19, QCYJ17, QCD16, RGAT18]. cloud [RT16, RRKA19, RML^{+}19, RZ16, RT15, Rao17, RS17a, RP18, RHPV17, RMJ^{+}18, RTS^{+}16, RMRSA19, RGVGGSSZ14, RMHMG17, SRZD15, SF19, SB14, SPD^{+}19, SJV^{+}15, SHRE16, SBAD^{+}18, SBK^{+}16, SJTN18, Sha16, SSSJ19a, SLD^{+}15, SYY^{+}17, SSI19, SAR18b, SV15, SG14, SKS17, SRKS18, SM18, SYK^{+}17, SB16, SBP^{+}17, SMS13, SK19, SCJ^{+}19a, ŠCJ^{+}19b, SBL18, SMZ^{+}16, SLZ^{+}18, SDH^{+}19, TWdLZ19, TCC18, TLC^{+}15, TSWL17, TDLC17, TOD17, TZL^{+}18, TKA^{+}18a, TKTG19, TBB^{+}17, TCBPR16, TCBC18, TYWZ18, TMS^{+}17, VB18, VPP^{+}19, VSDD13, VCDK18, VGD^{+}19, VSP^{+}14, WHMO13, WWCN13, WWC14, WXYL15, WCF^{+}15, WMK16, WZZ16, WMX^{+}17, WCL^{+}17a, WWDF18, WJS^{+}18, WMY^{+}18, WXZ^{+}18a, WXZ^{+}18b, Wan19, WLA17a, WLA17b, WLA18a, WCC14, WYH^{+}17, WZCH17, WBR19, XX14, XKBA18, XZW^{+}19, XWM18, XTT18, XWL^{+}18, XLL^{+}19a, XTZ^{+}19, XYML19, XWjZyF19, YMW^{+}18, YFY^{+}13, YLHJ14, YSC^{+}15, YLN15, YWL^{+}17]. cloud [YJS18, YCD^{+}19, YSHM19, YH19, YN18, YKK13, YQZ^{+}19, YAGG18, YNY^{+}14, YZN^{+}15, YGYW16, YYLC10, YDQC19, ZCM19, ZRZL18, ZYZ^{+}18, ZLZ13, ZCL^{+}14, ZME^{+}15, ZZH^{+}16, ZCS^{+}16, ZWHC17, ZHHC17, ZXZL18, ZZXL18, ZTL^{+}19, ZSL^{+}19a, ZCZ^{+}18, ZDW^{+}16, ZZC18, ZFH^{+}18, ZLY^{+}19, ZZS^{+}19, ZaTZ^{+}17, ZZZ17, Zhu18, ZWGC19, ZCW19, ZL12, ZZQ^{+}13, ZBF14, dSBN19, dACAM13, dACNC16, BAB12, CFVP12, CMG^{+}19, CJN^{+}15, CFF14, CPP16, DRZ^{+}19, DDJ^{+}13, DHC^{+}17, ENC^{+}12, FG14, FP14, FMRS18, IG12, JS12, JBR^{+}16, KKB14, KMK^{+}14, KS17b, LLAH13, MGR11, MEBA12, MBS13, NJH^{+}18, NHH^{+}19, PMPC13, PMDS18, SMBMT^{+}18, SG15, SMZ^{+}16, SYL18, SYQ^{+}19, TCN^{+}14, WLML17, YNSM12, ZLR^{+}15]. cloud-agnostic [KDG^{+}19]. Cloud-assisted [HMA^{+}18a, LZY^{+}19a, LYL^{+}19, WXYL15, ZYZ^{+}18]. Cloud-aware [GJ15]. Cloud-Based [HXA^{+}17, AMPZ16, CPD^{+}15, FPP^{+}18, GMP^{+}17, LOR^{+}18, YARH18, ASO14, BDE17, CFL^{+}15, DMPP16, EPB18, FW19, GSR^{+}19, GGC18, HJA^{+}19, HLYW17, HHK^{+}16, KRD^{+}19, KSC^{+}19, LCH^{+}11, LSYC18, MK17, MYK16, MCG^{+}15, PC18a, PKY^{+}17, TKA^{+}18a, TYWZ18, WZZ16, WJS^{+}18, XKBA18, XTT18, YMW^{+}18, YN18, ZZXL18, PMDS18]. cloud-computing [ZLZ13]. cloud-connected [RHPV17]. cloud-distributed [AB16]. cloud-edge [CFG^{+}19, CXWT19, DLL^{+}19, XZW^{+}19, XLL^{+}19a]. Cloud-Edges [BDL^{+}19]. cloud-enabled [ATM^{+}19, TOD17]. cloud-end [DJZ^{+}15]. Cloud-Fog [PDH18, MWL^{+}18b, ZSL^{+}19a]. Cloud-FuSeR [SMZ^{+}16]. cloud-hosted [YKK13]. cloud-integrated [FCD^{+}14]. cloud-native [SCJ^{+}19a, ŠCJ^{+}19b, TBB^{+}17]. Cloud-of-Clouds [WLML17]. Cloud-of-Things [CMG^{+}19]. cloud-oriented [FKT14]. Cloud-P2P [SYL18]. CloudFlow [ZME^{+}15]. CloudGrid [CCRV13]. CloudIntell [MAY18]. CloudLaunch [ALTG19]. cloudlet [ZGL^{+}18]. cloudlets [Lok12]. CLOUDRB [SG14]. Clouds [GLNT13, GGA^{+}17, KSF^{+}13, LX13, MG18, MAD^{+}16, PMLVLS^{+}13, SPdSR^{+}17, WRK^{+}15, APAZ17, AMQS^{+}19, AS19b, AMGCC18, AK14, ABN19, ASB18, AEME^{+}18, BL15, CLRL17, CTVB12, CGBAP18, CKR16, CBN16, dCCDFdO15, CRTN17, DT16, FWB13a, GCBM17, GZWQ13, HZZ^{+}14, HBN^{+}13, HFT16, IHK^{+}18, KuRAk^{+}18, KH18a, KFK19, LZ10, LLC^{+}14b, LMZ^{+}14, LGY^{+}16, LLZ^{+}18a, LLZ^{+}19, LLW^{+}12b, LPV^{+}16, LSHW17, LWZ18, LPD^{+}13, LYW^{+}16, LL16, LSMVML13, MJM^{+}16, MJDN15, MGV^{+}18, MGG^{+}17, MDD15, MG10, OG18, PFRC16, PS19, QGT^{+}18, RMVG^{+}10, RMCMD12, RCTY19, STMV18, SEMJ11, SSL13, SBA^{+}17, SJV12, SNP19, TZST14, TLSC17, TdPF^{+}17, TVB18, TSB18, TCN^{+}16, UDST19, VVB13a, VHML11, VCKB12, WHS^{+}18, WWZ^{+}19, WG13, WLH^{+}19, XSMS15, XJWW15, YLJL18, ZYCZ19, ZQB^{+}18, ZT19, ZB19, dCTVC18, dOOO^{+}13, ANE13, ATX13, CVKB12, WLML17, YMY^{+}17]. CloudSME [TKA^{+}18a]. CloudSNAP [MGLPPJ13]. CloudStore [LSB^{+}18]. CloudWave [SNP19]. Cloudy [BFN18]. ClowdFlows [KOP^{+}17]. Cluster [BB13, CWJ16, KV17, BARMB14, BYL^{+}18, CDG^{+}14, EMHE18, GCV^{+}14, HLL^{+}17, KZ17, LZXW13, LJL12, LHPC^{+}19, LKTC14, MNV12, MFG^{+}14, RBJ^{+}13, STC15, SMS14b, SBP^{+}17, TM19, WKC^{+}13, WHW16, WJZ^{+}17, YLHJ14, YYW^{+}19, dACAM13]. cluster-based [BARMB14, LZXW13]. Cluster-group [CWJ16]. cluster/datacenter [LKTC14]. clustered [PRC^{+}14, SCY^{+}18, GK18]. ClusterGrids [KKJJ10]. Clustering [FJ18, GBY16, HO17, LJ17a, LHJC18, PC18a, SAH19, TF18, ZWHC17, BK16, BCF16, Bu18, CJ14, CdSDS15, CSP13, DNW^{+}19, EP13, FGM11, GPJC17, GNGG17, hKBB11, KP18, LZZ19, LYS^{+}19, LXF19, LRMS19, LZXG12, LSV^{+}18, NK17, RCM17, SV16, TTC^{+}14, WCL^{+}17b, WTG^{+}19, XZZ^{+}14]. Clustering-based [LJ17a, LRMS19]. Clusters [MG10, AHEM17, ADAAD12, ALM^{+}10, BC15, BL13, BTM10, CRM^{+}16, CP17, CSdCM^{+}17, ÇBCA15, FQBCF15, GCBM17, JBP^{+}18, KSS11, LLZ^{+}18a, MJM^{+}16, PK11, RGDML16, SK12, VVB11, XDHL12, YIA17, ZGB^{+}17, ZBCT17, ZTD^{+}18]. cluttering [Gra15]. CMPs [MVL18b]. CMS [RAA^{+}18]. CMSS [CEP19a]. CNN [RBGA18, UMUB19, ZDM^{+}19]. CNN-LSTM [ZDM^{+}19]. CNNs [AAM^{+}19, LYL^{+}19]. Co [DXL^{+}18, DVB14, DBD^{+}14, GS15, VL19, XZL^{+}19, YCY10, YSC^{+}15, LLS^{+}14]. Co-AdaBoost [LLS^{+}14]. Co-Allocation [DXL^{+}18, DVB14, YCY10, YSC^{+}15]. co-design [DBD^{+}14, XZL^{+}19]. co-processor [GS15]. co-simulation [VL19]. Coalition [KIMR15]. CoAP [GCK18]. CoAP-based [GCK18]. coarse [SJ18, TKK^{+}14]. coarse-grained [TKK^{+}14]. CoCaMAAL [FKT14]. CoCoA [RJS^{+}19]. Code [AsRA^{+}19, BGS^{+}19, HDO16, HCL^{+}17, JLS19, LLC14a, OBG^{+}18, PSK^{+}10, SJ14, TLL^{+}19, WMJW18, vW19]. codecs [EBCP18]. coded [PWMX17, XWM18]. codes [GIM16]. coding [CJXX19, LDX19, LSD11, LYS12]. coefficient [QMCX19, SLD^{+}18, ZWZ19]. coevolution [LD17]. coevolutionary [CZY^{+}18]. Coflow [LZZ19]. Cognitive [AT19b, GWW^{+}19, MK19a, ASAA18, AS18b, BKS^{+}18, BMK^{+}14b, CLH^{+}18, Che18, CGM^{+}19, CWLZ19, GOLL19, HMW^{+}19, LLW^{+}19c, LLW^{+}19d, NDZ^{+}18a, NDZ^{+}18b, NDZ^{+}19, STB^{+}19, TSD18, WLL^{+}19a, WLZ^{+}19, XZZ^{+}19, ZYZ^{+}18, GTEL^{+}18]. Cognitive-inspired [GWW^{+}19, CWLZ19, XZZ^{+}19]. coherence [MVL18b]. cohesive [QC18]. Collaboration [ASW11, DPDS14, GFD14, AkBAL^{+}19, BJM^{+}17, BAPS14, BP10, HKU^{+}11, HLL^{+}11, HYS17, HYS18, KKBK19, LZH^{+}18, LLS^{+}19, LSW^{+}19, OPT^{+}17, PPS^{+}19, POJ^{+}16, RMA^{+}16, SKF^{+}11, UNM^{+}16, WTP^{+}13, YLJ^{+}17]. Collaborations [GDZ^{+}19, TAB^{+}18, EYY19]. Collaborative [BDF^{+}16, BDG^{+}19, FGW^{+}19, HGM15, HAM18, LTTL19, LZY^{+}19b, OMKM^{+}19, PLLP19, PVGD^{+}19, WWTF18, WYL^{+}18, AM19b, BRXdS11, BDCC19, BDMO11, CFM19, CPE^{+}17, CSL17, CSL19, CMP^{+}17, CNP^{+}19, CPLH19, DMPP16, DFRW17, DV13, DPS16, DLL^{+}19, DKFKF18, FLR^{+}16, FdAGdAFV19, GMD19, GMP^{+}16, IGB^{+}14, KZA^{+}18, KKB18, KIAD17, KWK16, LCL^{+}16, LCW^{+}18, LTMW19, LQLX10, LHCC18, LHY^{+}19, MZH^{+}17, ML11, MVG18, MOBD18, MQL^{+}19, MED16, NRR^{+}15, PS10, PDW^{+}11, RW18, RRH16, SHBP10, Sip12, SCL14, TQL^{+}19, WCVL12, WLB11, ZL13, ZCZ^{+}18, ZWYH19, ZRZ^{+}14, ZBF14, dlFVPSHL^{+}14, FGG13]. collaboratively [GZS14]. collaboratively-built [GZS14]. collapsed [KWB19]. collapses [KMC18]. collating [AB18b]. collection [DGCGH^{+}17, HMM18, HYC^{+}18, KC19a, LLW^{+}19d, PPS^{+}18, TSTL16]. collection-processing [PPS^{+}18]. collections [YDK11]. Collective [Jun17, CFGM16, SSST17]. collision [CMZ^{+}18, NLS19]. collusion [LXJD18]. collusive [ZZH^{+}18]. Colony [PT16, TA18, CZL^{+}18b, FZHH14, HXWW18, HLZ18, SMS^{+}19, SH19, TV16, WZ13, XLW^{+}17, ZHHQ18]. Colony-based [TA18]. color [Bro19, DNW^{+}19, GPV^{+}14, HX19, MSM^{+}18b, RM16, XJZ^{+}19]. Colorectal [KMB^{+}17]. colored [DC19]. coloured [AEM10]. Column [GVA^{+}16]. Combination [CWJD19, LLW^{+}18a, SYT^{+}19, WQZ19, XRPT18]. combinational [CWW^{+}16]. Combinatorial [HMH17]. combine [SDST18]. combined [SMBMT^{+}18]. combiner [MLC18b]. Combining [AS19a, AGBR19, CPGBC16, CFGM16, Che13a, DMC^{+}19, MK19b, MWPVB12, OCW14, FZHH14, GAA19, SK12, WWZC19]. CoMD [PAL^{+}19]. CoMe4ACloud [ASAB^{+}18]. COMFIT [dFBP^{+}17]. comfort [FHHM19, Mat18]. comfortability [WPS^{+}18]. comments [CXZC18]. commerce [ZL18]. commercial [ABN19, CBN16, DQC^{+}19, SMM^{+}14, TKA^{+}18a]. Commodity [JBP^{+}18, KDHP16, LKTC14]. Common [GAB^{+}14, CCL19, CZ14, HZX^{+}19, HZX^{+}20]. commons [BFN18]. communicating [MZD^{+}16]. Communication [ADALZ14, DKV14, AK19, AAQ^{+}19, AKG^{+}17, Bag11, BWR12, BTM10, CDDR17, DSBC19, DdM10, EAA16, GQXL18, GXL^{+}12, GZLZ16, GVBG17, GBY16, HAAWH^{+}18, KH19, KMU19, LXF19, LWW^{+}16, MCN^{+}18, MWCK19, MWL^{+}18b, MCG^{+}15, PSK^{+}10, QMCX19, RPA^{+}18, RHPV17, SRZD15, SAH19, SLS^{+}19, SCEC18, SAVS19, URKM19, UPP17, VSKS19, VCDK18, VV16, WPGN^{+}18]. communication-layer [WPGN^{+}18]. communication/networking [GXL^{+}12]. Communications [EYY19, AMN18, CFVP12, GZL^{+}18, GSC^{+}19, HSC15, HCHH19, MHY^{+}18, QGX18, RGDML16, SAGGB17, WQ14, WSZC18, WGX^{+}19, XHW19, YXZG18a, YYS^{+}19, ZZY^{+}19, dSBN19]. communities [BOHCC17, DMPP16, FK11, HSB^{+}18, PRSR14]. Community [àCKPM19, GDZ^{+}19, RMA^{+}18, BFN18, BR19, CWL^{+}18, CWJD19, CGL^{+}10, FPPD14, FLN^{+}18, GHJ^{+}19, GG10, HZ19, HBN^{+}13, LFL^{+}17, LJJ18, LLZ^{+}18a, SOD18, SEHS19, SGBK19, SZG^{+}19, SJL^{+}18, YMLT13, ZDL^{+}13, ZCL^{+}19, HRR^{+}14]. Community-Aware [RMA^{+}18, HZ19, HBN^{+}13]. community-based [YMLT13]. companies [STMV18]. Comparative [CA15a, TAHS14, BMU18, CDMR19, DNW^{+}19, GRS^{+}19, OdOD^{+}13, RHH^{+}19, ZN12]. compare [CW16]. Comparing [HHS^{+}18, PFS^{+}13]. Comparison [CHS^{+}18, MFG^{+}14, MOFGP18, CRWZ19, GLJ19, JS13, KBVH14, KS17b, SI18, ZTKF17]. compatible [GA13, HDLW13, SOM^{+}19]. compensation [LZLL18a, LZWF19, YXY18]. competency [KRZ12]. Competition [GPVN19, JLQ18, LLS^{+}19, SIL^{+}13]. Competitive [AGKZ18, PS10, SSL12, WPJ16]. compiler [GMB19, LJS17]. complementary [PRS12]. complementing [RS16]. complete [BP13, KJ18, MSM^{+}13]. Completely [GMM18]. completion [CND^{+}19, LZ10]. Complex [BCD^{+}18, CCMGF18, CCRL18, HCB16, LJ17a, BKB11, Bal16, BWR12, BW13, CSC18, Fer13, FPP^{+}18, GRZ^{+}19, GJY18, HSB^{+}18, JSZ^{+}19, KAP19, MSO18, PKI^{+}18, RWY^{+}18, SW17, Sun10, SJL^{+}18, WSZH18, WMC19, ZYTC15, ZSP17]. complexes [CGV10]. COMPLEXIS [CCRL18]. Complexity [GMMM18, AMÇ19, ABP16, GVURIVBV14, TWW^{+}18, ZZF^{+}19]. Compliance [SS17, HHL11, LCCM18]. Compliance-based [SS17]. compliant [LLCF11]. Component [BR18, SAPA17, ACPI19, HIA18a, LASL16, LTZ15, MZYA19, PSBB15, SVN10b]. Component-based [SAPA17, ACPI19, PSBB15]. Component-oriented [BR18]. components [KSS11]. Composable [LFP^{+}17, ÖE13]. Composing [AMM^{+}19b, RBC^{+}15]. Composite [AM17, Lin18, ATF11, MBA19, OCW14]. Compositing [MWQ^{+}19, NOF18]. Composition [JGB19, SZW^{+}19, AdI14, ACCM19, BML18, FTD17, GMMM18, GJGB19, JGFB18, LPMY18, LJ17b, LWS^{+}12, LO19, PKB19, SJ18, SDC11, TSOB15, TSTD16, UGBM^{+}17, WPY19, WZ13, XSMS15, XWW19, XLW^{+}17, XCZ^{+}19, YWLL19, ZZLH18, LKA^{+}19]. compositional [dSMAdR^{+}17]. compositions [Lok12]. Comprehensive [GHJ^{+}19, Tao10, ZGL19, HHK^{+}16, LDJL19, NJKH13]. compress [NS17a]. Compressed [JL14, ML17, BD18]. compressible [WWP19, WWP20]. compressing [JDW^{+}14]. compression [AMBC19, CCD^{+}19, CPD^{+}15, DHW^{+}17, HSP^{+}13, IASK14, OSC14, SMS16, YZWG18]. compressive [GSY^{+}17]. Compromise [CLK11, NAAC19]. Compromised [SSA^{+}19]. CompSci07 [AC10]. Compulsory [QRW^{+}18]. Comput [AB19a, AB21, ABMMC22, BFS^{+}17a, Bo20, Cha14b, DP20, DP21a, DP21b, GHEB^{+}23, HZX^{+}20, HYS18, JLC^{+}20, NDZ^{+}18a, NDZ^{+}19, SME^{+}21, WWP20, WCWC20, YWG^{+}20, YTQ20, ZMZ^{+}20, wZcZN^{+}20]. computation [AHU^{+}19, Bal16, BWG19, CAPG18, CWJ^{+}18b, DEL19, EZTL19, EHT10, GLC19, GEAR13, HHK18, HLT^{+}19, KGLY18, LJC^{+}19, LLJ^{+}11, LLM^{+}16, LSAM13, PB18, STC15, VV16, WMK16, WLXZ18, XLL^{+}19a, XXQ^{+}19, YDT19]. computation-driven [LJC^{+}19]. Computational [AC10, BDS^{+}10, Jun17, SA19, XA10, BMT12, BL15, CBBC^{+}17, CGJ^{+}10, DVB14, DLS^{+}12, GKTK15, GSN^{+}18, GPS^{+}17, Hua10, IAL10, KHG13, KX11, KKW^{+}14, LAH10, MFN13, RM11, SPdSR^{+}17, SSST17, SSL12, SKS17, SDF^{+}19, SZR18, TBK^{+}10, WSC^{+}19, WCKW10, WS10, ZZDM^{+}18]. Computationally [VCD^{+}18, DST10]. computations [AQB15, HCW^{+}18, SK12, SK19]. Computer [DPDS14, JO11, LBJ^{+}24, ADAAD12, ABS^{+}18, BDS^{+}10, CAC^{+}10, CRSdS10, GMLGB^{+}17, JBP^{+}18, KdGP^{+}19, LWW^{+}13, MGA^{+}18, NSSA^{+}14, PPS^{+}19, WY19]. computer-assisted [BDS^{+}10]. Computers [GMM18]. Computing [AR17, ADALZ14, APS^{+}19, ABP18, AKB^{+}18a, BB13, BCG^{+}19, CMA11, CFP^{+}19, DXA14, FSV^{+}19, FBS18, Hel16, JJH19, KFBKD14, LWW^{+}13, MGR11, OFMZ18, Pal16, PN13, PZY16, SAPA17, SG14, SPKG18, WRK^{+}15, WZM^{+}18, WSB^{+}15, Wri19, WXGM18, XDWL15, YHL16, YWZ^{+}18, AZH18, AHS^{+}18, ABMC18, ANA16, AJY15a, ABS11, ASW11, ANG^{+}19, ADLW12, AkBAL^{+}19, ADA^{+}19, ALK15, AL18, AMGCC18, ARB12, ADBO18, AMS19, BJ12, BBWB^{+}18, BAV16, BMU16, BKS^{+}18, BFS^{+}17a, BFS^{+}17b, BYL^{+}18, BAB12, BOP^{+}14, BDZ13, BKY18, BT17, BdDPP16, BCP18, BKB18b, BB12, BDMO11, BR10, BRHH18, CMZ^{+}12, CPDJ13, CLRL18, CHWW13, CGCB^{+}12, CA15a, CLCMG^{+}18, CPGdS^{+}13, CLZ18, CTR^{+}17, CDDR17, CMT16, CC11, CKR16, CLL18b, CWJ16, CXL^{+}17, CLH^{+}18, CCY^{+}18, CXWT19, CW13b, CLCY18, CGSJ18, CGM^{+}19, CNP^{+}19, CCCT14, CBLS13, CSP13, CWLZ19]. computing [Cuz14, DMC^{+}19, DQC^{+}19, DJZ^{+}15, DST14, DDD18, DSD^{+}11, DRGC^{+}19, DLMS15, DDB14, DWS12, DHC^{+}17, DCMW17, EZTL19, EAS^{+}18, EMJ^{+}13, EA17, FLL^{+}19, FLR13, FRB^{+}14, FP14, FP13, GHY^{+}18, GVBG17, GVB13, GAW^{+}18, GGJ13, GDJ^{+}13, GJ15, GCCL18, GB10, GMP^{+}18, GRX19, GCZ^{+}19, GZWQ13, GTMZ17, HFM19, HSM13, HAP11, HCZW17, HB19, HKPT10, HMZ18, HMA18b, HZDS19, HNQ^{+}18, HMW^{+}19, HZ10, HYS17, HYS18, HXC^{+}18, HLT^{+}19, HKP10, IFD^{+}19, JCA^{+}19, JSMG18, JEB18, KNI^{+}18, KMB^{+}17, KANS18, KZ17, KKKM13, KB18, KAH^{+}19, KMI11, KAW12, KMJ18, KPJ19, KLH^{+}18, KACN16, KGS^{+}19, KJ12, KH18b, KMV^{+}15, KGLY18, KJ18, LSB^{+}18, Len16, LPK17, LPK18, LXD17, LXJD18, LS10, LLW^{+}12a, LCHW14, LXK^{+}14, LFP^{+}17, LLH^{+}17, LZS18, LW18a, Li18, LTTL19, LLW^{+}19c, LCL14, LC15, LXMW15, LHX^{+}18, LZYC13, LYL15, LHL15, LPL^{+}16, LDY^{+}18, LZLL18b, LLZ^{+}18b]. computing [LPBB^{+}18, LSJ^{+}14, LYH^{+}19, LCY^{+}19b, MWW^{+}15, MLC^{+}18a, MDA^{+}19, MS19, MKH13, MEBA12, MK19b, MGMT18, MMC^{+}18, MRH17, MPR^{+}16, MSS^{+}16, MLSF16, MSE19, MGA^{+}19, NPH19, NRV^{+}17, NSSA^{+}14, OdI14, ODK^{+}17, OB17, PZA18, Pal13, PdASM18, PRL^{+}19, PMMAM13, PGTBC18, PMCC18, PSY^{+}19, PPB16, PZY17, PPLL17, PS13, QRW^{+}18, RGN^{+}18, RGAT18, RC13, RML^{+}19, RBN13, Rao17, RR18, RLM18, SB14, SPJ17, SBB^{+}10, SSSJ19a, SA19, SCAC^{+}19, SG13, SSZ13, SMM^{+}14, SBD^{+}18, SYK^{+}17, SJV12, SK19, SG15, SSP17, SLL^{+}18, SGL^{+}19, TLC^{+}15, TZLL18, TGM^{+}19a, TKTG19, TuIS^{+}19, TGM^{+}19b, Tur18, VNAMM19, VVC^{+}12, VAR14, VB18, VPT^{+}15, VCL^{+}19, VSDD13, WTR^{+}13, WWCN13, WQ14, WWC14, WWRS16, WCL^{+}17a, WWDF18, Wan18a, WMY^{+}18, WXZ^{+}18a, WLL^{+}19a, WDKV19, WXZL11, WZCH17, WHYZ17, WMJW18, WLHH18, WPS^{+}18, WHZ19, XZ11, XX14, XZP^{+}19, XLL^{+}19a, XTZ^{+}19]. computing [XXQ^{+}19, YFY^{+}13, YJS18, YSHM19, YW12, YVCB10, YCH19, YY11, YGYW16, YCX18, YAX^{+}18, ZAA^{+}14, ZGZ^{+}10, ZLZ13, ZZH^{+}16, ZWHC17, ZBCT17, ZTKF17, ZCX^{+}18, ZCL^{+}18, ZXZL18, ZZF18, ZZ19, ZYC^{+}19, ZSL^{+}19a, ZLY^{+}19, ZSH12, ZZZ17, ZCW19, ZL12, ZSW^{+}18b, dSGD13, dACAM13, vKvWD^{+}13, AHL11, CC11, LLAH13, MGR11, SGN^{+}17, YXA^{+}18]. computing-enabled [LW18a, XXQ^{+}19]. concentration [KAS^{+}18]. Concept [HLV^{+}16, WZL18, HXC^{+}18, IdAP19, KI19, KWB19, TSRG17, ZDW^{+}18]. Concepts [OFMZ18, TAB^{+}18]. conceptual [CIK10, UZ11, dSGD19]. Concurrency [Vin16]. Concurrent [CHS^{+}18, SZW^{+}19, AB17, AB18c, GGS13, RBJ^{+}13, RGDML16, RM11]. Conditional [MAÇ17, BDCC19, CLM^{+}14a, HYF18, IOV^{+}18, LY18b, TA19]. conditions [MLGGB^{+}17, MSE19, YZC^{+}19]. Conference [CC11, KZ17]. confidence [DCC13]. confidence-based [DCC13]. confidentiality [GSR^{+}19]. Configurable [CBT^{+}19, XWRZ19]. configuration [CLAL19, CFMC19, DDD^{+}19, GGTRRC16, LMZ^{+}14, MBS13, PKI^{+}18]. configurations [CMVA18]. configure [KPM^{+}18]. Configuring [BYL^{+}18]. conflict [WCH^{+}18]. conflict-aware [WCH^{+}18]. conflicts [DR15]. confluence [SISGS18]. Congestion [Ciu10b, HDB18, MBM18, NLLC19, SSK^{+}19, KXS^{+}16, LTN10, LGP^{+}19, LAQ^{+}19, MWPVB12, RJS^{+}19]. Congestion-aware [MBM18]. congestion-free [LGP^{+}19]. conjunctive [HZL^{+}19, XTZ^{+}19, YQZ^{+}19]. Connected [BRH18, HAJ^{+}19, XXQ^{+}19, KMU19, LYC18, NLLC19, PECA19, RHPV17]. connectivity [AC18, LY18b, VOS12]. Connex [JCA^{+}19]. conquer [KTB18]. Conqueror [KTB18]. conscious [ZA14]. consensus [CFGM16, CFM19, LTC^{+}19, MMF16]. consequence [OA17]. conservation [Mat18, PPMM^{+}18]. conserving [NQQL13]. consideration [FW19, GGLW18]. Considerations [CJPC19]. Considering [KMJ18, BAKB19, CT19b, WMBV17, ZZH^{+}16]. Consistency [KLP19, GNVST14, HKS18, LWTL19a, Sip12, TSWL17, ZW10]. Consistency-guaranteed [KLP19]. consistent [LGP^{+}19, ZSP17]. consolidated [SS13, ZFW14, ZLL^{+}16]. consolidation [ADBO18, CFF14, FNCR11, HMH17, HZZ^{+}14, HH19, IDM^{+}16, JFZL17, KCV11, LYYY17, LYYY18, LLWW18, Man15, MP17, RT15, SSSJ19a, TDSH16, WCC^{+}16, WLA18a]. consolidation-aware [WCC^{+}16]. consortium [HZLH19]. constant [KHJ10]. CONStanza [PDDS10]. constrain [GWW^{+}19]. constrained [APAZ17, ANE13, ABP16, AB17, ABN17, AB18c, CXL^{+}17, CLM14b, DSCJ18, KCK16, MJDN15, MRN19, NK15, PCK19, RPH19, SSL12, SGJ18, TWdLZ19, TLL^{+}11, VVB13a, ZYB^{+}18, ZT19]. Constraint [LPMY18, SSG17, DQLW15, DKFKF18, FCY18, GPS13, LYYY18, TSTD16, WLP10, XJY^{+}18]. constraint-based [DKFKF18, LYYY18]. constraints [CSL18, GQLX18, Li18, LLZ^{+}19, LD17, LPL^{+}16, LTZ15, SZK16, TSB18, WTM^{+}17, WZWC18]. Construct [HPZL18]. Constructing [CFL^{+}18, STB^{+}19, ZZL^{+}10, GNGG17, HQ10, NJ19]. Construction [KMZJ16, CPLH19, DPS16, LW19, LLS^{+}14, LWXY19, PRN14, SSRQ19, TWZP18, YPHZ14]. Consumer [AAC^{+}19, GMCM16, GMCM18, FFC12, HMA18b, ZL18]. Consumer-centric [GMCM16, GMCM18].

Consumer-facing [AAC^{+}19]. consumers [CAC^{+}15]. consumption [BdM11, CRB^{+}16, DPBK16, HCHH19, IPCA^{+}16, LCZR12, LIH^{+}19, MPF^{+}16, RMRSA19, SRP19, SYQ^{+}19, VGC^{+}13, WOPW13, ZAC^{+}18]. contact [CTU19]. Container [LYH^{+}19, CMG^{+}19, KS17b, MG19, PMCC18, ZTL^{+}19]. Container-based [LYH^{+}19, KS17b, PMCC18, ZTL^{+}19]. containerized [AS19b, GCTLA^{+}19]. containers [BGRBA19, DSS19, DL19, MK19b, MEW^{+}19, SF19]. contend [BA17]. Content [ARP14, SNP19, WSQ^{+}18, AAF18, AAS^{+}19, BCR^{+}12, CdRRdCB19, DFLO17, FLT17, GVI13, GGH^{+}19, HCC^{+}14, HDLW13, HPL^{+}19, LLMP13, LMZ^{+}14, LLF^{+}18a, LZP^{+}18, LLWW18, LSYC18, LWZ^{+}19b, RLL^{+}17, WLA17a, XFJ^{+}19, XWJ^{+}16, YZZC19]. content-based [FLT17, XWJ^{+}16]. content-defined [WLA17a]. content-defined-chunking [XFJ^{+}19]. Contention [KIMR15, CNR19, DLXR14]. Contention-Aware [KIMR15]. Contents [Ano11b, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, BTP19, CWJD19, MSM^{+}18b, SKS^{+}18]. contests [WXYL15]. Context [ABTA18, ACSV18, BTP19, CBC^{+}19, KRD^{+}19, MBA19, VNAMM19, ZGZ^{+}10, AGA16, BN17, CHY^{+}18, FG18, FNA12, FKT14, GA13, HSB^{+}18, JOSD19, LRJG19, Lok12, NKB19, POJ^{+}16, RMSPP17, TCB^{+}17, UGBM^{+}17, WWG19a, WYH^{+}17, WYL^{+}18, ZWYH19, dSGD19]. Context-Aware [ABTA18, BTP19, ACSV18, CBC^{+}19, KRD^{+}19, MBA19, VNAMM19, CHY^{+}18, FKT14, Lok12, NKB19, UGBM^{+}17, WYH^{+}17, ZWYH19, dSGD19]. context-cloudlets [Lok12]. context-sensitive [JOSD19, WYL^{+}18]. contexts [XYLZ18]. Contextual [MLL15, dSMAdR^{+}17, PNZ14]. Contiki [ZAI^{+}18]. continental [MBC^{+}11]. continents [UNM^{+}16]. contingency [DCC^{+}14]. continual [ZLXZ18]. continued [WTS14]. continuity [MBMTJR18]. continuous [KC19a, PLGMCdF18, ZLT^{+}19]. contour [WZF^{+}19]. contours [HWWT12, HZM14]. contract [LML^{+}19, VVB15]. contracts [WGC19]. contracture [AMAY19, MSA^{+}19]. contrasting [MRL14]. Contributing [FTK17]. contribution [MCWP16]. contributions [ALFR16]. Control [AMSPL19, CLL^{+}18a, CDL^{+}16, KESL17, YLWW18, AHP16, AMHJ10, BR18, Che13a, CYZK15, CZM^{+}18, CZZ^{+}18, CJK^{+}18, CNR19, DCMB15, DLDTGMMP16, DHL18, FXG^{+}19, FJL^{+}16, FNCR11, FS18, FM10b, FSM^{+}18b, GHMX10, GZLZ16, GHD19, HAJ^{+}19, HLW12, HYG^{+}19, Hua10, HDB18, IS18, KRD^{+}19, KKS^{+}18a, LHO17, LLW^{+}19a, LLW^{+}19c, LAQ^{+}19, LZLL18b, LLZ^{+}18b, LFY^{+}19, MYHZ18, MM10, MLGGB^{+}17, Mer13, MLW^{+}18b, MWPVB12, NLLC19, NRV^{+}17, NWL17, NJ16, NA19, PWWD18, QCX18, RJS^{+}19, RGSL18, SRdlPG19, SAK19, SMSF18, SCL14, SYK^{+}17, SZR18, TDC^{+}14, TMJH19, TV16, VPP^{+}19, VS19, VFHB14, Wan18a, WSY^{+}19, XSMS15, XZZ^{+}19, XYML19, YXZG18a, YWJ^{+}19, YAX^{+}18, ZYK17, ZCL^{+}18, ZFH^{+}18, ZSL^{+}19b, dSBN19, vdLLE19, BL13]. control-Application [BR18]. control-theoretic [FJL^{+}16, Hua10]. Controllable [FH13, YD18, ZWGC19]. controller [FJL^{+}16, HYG^{+}19, NLV^{+}19, ZXD^{+}19]. controllers [dSFD^{+}19]. Controlling [GCD^{+}18, PECA19]. controls [QGT^{+}18]. convenient [WHYZ17]. Convergence [Kim14, SPS18, WMA^{+}19, LJJ18]. Convergent [MFT^{+}17, TSWL17, XZW^{+}19]. converter [VS19]. convolution [HZW19, LYW^{+}18b, vWMBS14]. Convolutional [FS19, AAYL19, AsRA^{+}19, AFO^{+}18, AMBB18, ASYF18, AMM19a, CZH^{+}18, GLC19, HUY^{+}19, LZT^{+}19, LLU^{+}18, MMAA19, TBS^{+}18]. cooling [Len16, QXZ^{+}19]. cooperation [BDP11a, FMN^{+}17, XDHL12]. cooperation-oriented [XDHL12]. Cooperative [EA13, HDH^{+}18, JO11, PMFH11, RRB10, SYJA19, ABDH19, CZY^{+}18, DPL14, FAL^{+}19, FTH16, LJ17b, LD17, LWZ^{+}19b, Mer13, PPS^{+}19, PLP^{+}19, PNZ14, YC13, Zhu18]. coordinated [CRVZ15, ZXD^{+}19]. coordinates [HHZ16]. Coordinating [GDR^{+}14]. Coordination [YZZC19, FTP14, MSBA16]. coordinator [CTVB12]. Coordinators [WLL^{+}19b]. cope [ZCS^{+}16]. Copernicus [PPLL17]. copilots [PWP^{+}18]. copy [LXRS19, ZTKX19]. copy-on-write [LXRS19]. coral [FEPC18]. coral-reefs [FEPC18]. Core [TAB^{+}18, EDH^{+}13, HAF^{+}16, HYZS16, JLY^{+}18, LC14, LLG^{+}16, LG16b, MWYC12, MAC14, MCA^{+}18, MBM18, MGMT18, MCF^{+}11, MMRL17, PSPP16, PLL^{+}18, PSLZ18, PK11, PRC^{+}14, QZM^{+}18, RS17b, SK18, VD16, WJZ^{+}17, WWZC19, WSH^{+}16, YMW13, YLJ^{+}17, YZ12, YDT19, ZAB15, ZLG^{+}14, ZHHQ18, VK17]. core-shutdown [WWZC19]. CoreFlow [KBdLG18]. cores [Eng14, HLZ18]. corner [HTL^{+}18]. correct [JMAG19]. correction [KDHP16, LZL19b]. Correlation [FAA^{+}18, GJY18, HZX^{+}19, HZX^{+}20, GHP^{+}18, LA19, YL18, YGY^{+}19]. correlations [LD17, dFPFG19]. Corrigendum [AB19a, AB21, BFS^{+}17a, Cha14b, HYS18, NDZ^{+}18a, NDZ^{+}19, SME^{+}21]. corrugation [LSZ^{+}18]. CoSHE [PMDS18]. COSHH [RD14]. Cost [APAZ17, ALK15, AK14, BKKM11, KV17, MFN13, MEBA12, WMX^{+}17, ZQB^{+}18, ADA^{+}19, AL18, ABN19, BSE^{+}13, BBB^{+}19, CRM^{+}16, CHSA18, CWJ^{+}18b, CSQL17, DMC^{+}19, DR18, FdAGdAFV19, GBS10, GD10, GS16b, HGG^{+}14, KK16, LGY^{+}16, LGL^{+}17, LSCL19, LEW19, LML^{+}19, LYL^{+}19, MJDN15, MFG^{+}14, OCCK14, SGJ18, SBA^{+}17, SK19, SYAL13, TDSH16, VVB13a, WY17, WWZ^{+}19, ZG19, ZXZL18, ZYCZ19, ZSQ^{+}19, ZZS^{+}19]. cost- [MJDN15]. Cost-aware [ALK15, ABN19, GS16b, HGG^{+}14]. Cost-benefit [MEBA12]. cost-constrained [KCK16]. Cost-effective [WMX^{+}17, CWJ^{+}18b, CSQL17, DMC^{+}19, KK16, LML^{+}19, MFG^{+}14, SBA^{+}17, SK19, WY17]. cost-efficient [CRM^{+}16, GD10, SGJ18, VVB13a]. cost-sensitive [ZSQ^{+}19]. costs [KACN16, MD12]. count [XWX^{+}17]. counter [DLDTGMMP16, VOCHC17]. counter-forensic [VOCHC17]. counter-measures [DLDTGMMP16]. Countering [MWCK19]. countermeasures [FAA^{+}18]. counters [CLP^{+}14]. couple [BC17]. coupled [BC15, LLS^{+}14]. Coupling [BC15, sGbKS19]. couplings [PM14]. Courier [AT18b, ZFW14]. coverage [KWB19, LWH^{+}18, LZXG12, LLW^{+}19d, MLL15, WYJ^{+}19]. covert [MWCK19]. COWB [DMPP16]. CPFL [AHEM17]. CPRS [LCH^{+}11]. CPS [FSV^{+}19, LLW^{+}19b]. CPU [ASB18, BBMG10, CDG^{+}14, CNR19, DLS^{+}12, GJF^{+}12, IPCA^{+}16, JCA^{+}19, KDHP16, KANS18, KAEC^{+}18, MDB^{+}18a, PS19, RBJ^{+}13, TTH15, VD16, WMQ^{+}16]. CPU-based [GJF^{+}12]. CPU/GPU [CDG^{+}14]. CPUs [ZHHQ18]. CR [CLL^{+}18a]. Cracking [VSM^{+}19]. CRAFT [BDG^{+}19]. craniofacial [CPE^{+}17]. crashes [PWP^{+}18]. crawler [BLMU19]. Cray [KBVH14]. CREAM [AAB^{+}10]. Creating [BAB13, RGVGGSSZ14, SMS13, WMC19]. creation [ACM^{+}18, ACD^{+}19, GRZ^{+}19, HA16, WLB11]. credentials [SSA^{+}19]. credibility [PB18]. credit [KCS14]. creditable [HCL^{+}17]. CRFID [CWSW14]. crisis [SCL14, VDK12]. Criteria [ACD^{+}19, AFSH^{+}19, EG18, FJJ^{+}18, KHG^{+}18, KA19, MLSF16, PdASM18, SDH^{+}19, ZZH^{+}18]. criterion [FTK^{+}14]. Critical [AHEM17, AL14, AMR^{+}19, BMR15, BKB18b, DNJG17, FDPR17, KSC^{+}19, LZL19b, OWX19, RSRA18, RRU^{+}18, SOA17, STB^{+}19, ŠCJ^{+}19b, TWdLZ19, WTM^{+}17, WWT^{+}16, WHYZ18, XTF^{+}19]. criticality [CXZ^{+}19]. Cross [FSM^{+}18b, MFT^{+}17, SVK19, YZL^{+}18, ZZBP19, dSBN19, AKCY^{+}17, JBM^{+}18, LZH^{+}18, LXMW15, LSD11, LWW^{+}16, QZD^{+}18, SSG19, SSB13, XWM18, ZN12]. cross-datacenters [XWM18]. cross-disciplinary [LZH^{+}18]. Cross-domain [YZL^{+}18, SSB13]. Cross-Layer [SVK19, dSBN19, AKCY^{+}17, LXMW15, LWW^{+}16, ZN12]. cross-layered [SSG19]. Cross-Mapping [MFT^{+}17]. Cross-platform [ZZBP19, QZD^{+}18]. cross-stream [LSD11]. crossed [CFL^{+}18]. crowd [AMPZ16, CFGM16, CFM19, CWW^{+}13, KAS^{+}18, SOM^{+}19, VNAMM19]. crowdcomputing [WHS^{+}18]. crowdsensing [DWJM18, KRZ^{+}19, MAD^{+}16, YLWW18, YZL^{+}19]. crowdsourced [LLGY18, MMF16, MDM^{+}19]. crowdsourcing [àCKPM19, FY19, LDS^{+}18, SYJA19, WLA18b]. crowdsourcing-based [LDS^{+}18]. Crypto [ArMS19, XZL^{+}19]. Crypto-ransomware [ArMS19]. Cryptoanalysis [CJS19]. cryptocurrencies [ZLL^{+}19]. cryptographic [OMPSPL^{+}19, WMX^{+}17]. Cryptography [CDFZ16, DXA14, AMN18, MCN^{+}18, MK17, YXA^{+}18, ZXW^{+}18]. Cryptosystem [WSQ^{+}16, GLB^{+}18, YY11]. cryptosystems [Wan19]. CSCWD [Bar11]. CSTP [LSD11]. CT [HLZ^{+}19, KMK^{+}19]. CTPSO [XRPT18]. cube [CFL^{+}18]. Cuckoo [CWJ^{+}18a]. CUDA [KSC^{+}19]. CUIDATS [ABC^{+}18]. Cultural [CMP^{+}17, PC18b, CCMP18, MKS18, PPMM^{+}18, WDJC18]. CUPUS [AMPZ16]. curation [GZS14, GML^{+}13]. curative [Bo19, Bo20]. currency [PRS12, SI19]. Current [KARP14, LAQ^{+}19, KBVH14]. curvature [Jun18]. curve [AMN18, MCN^{+}18]. custom [ACCM19]. customizable [MDD15]. customization [SSW^{+}19]. customized [BAJ^{+}19, CTU19]. Cyber [CFH^{+}19, HLV^{+}16, LWW^{+}16, NLM^{+}16, OFD17, RVC16b, RVC16a, SYJ^{+}19a, SHS^{+}19, UNM^{+}16, ZZLR18, AHS^{+}18, APRC16, ABD^{+}19, BK16, CM17, DWJM18, EAED18, EG18, GVBG17, GAW^{+}18, GHD19, GGDM^{+}18, GMP^{+}17, HAAR^{+}19, KB16, LLES19, LSZ^{+}18, LLF^{+}18b, LLS^{+}14, LSL^{+}15, LZY^{+}16, MMPF19, NAAC19, OA17, PKF14, PTD^{+}18, SMS14a, SZK16, Sko19, SZD^{+}17, SM18, SDK19, WLZ^{+}16, WZH^{+}18, XLZ18, YPJ19, YS16, APRC16, Zhu14]. Cyber-Enabled [ZZLR18, LSZ^{+}18, LLF^{+}18b, YPJ19]. cyber-foraging [LLES19]. Cyber-Physical [CFH^{+}19, HLV^{+}16, LWW^{+}16, OFD17, RVC16a, ABD^{+}19, BK16, EG18, GVBG17, GHD19, GMP^{+}17, HAAR^{+}19, KB16, LLS^{+}14, LSL^{+}15, LZY^{+}16, MMPF19, OA17, PTD^{+}18, SZK16, Sko19, SZD^{+}17, SM18, WLZ^{+}16, XLZ18, YS16, APRC16, Zhu14]. Cyber-physical-social-thinking [NLM^{+}16]. CyberGuarder [LLW^{+}12a]. cyberinfrastructure [KS19]. cyberinfrastructures [CGBAP18, DGGH11]. CyberLiveApp [LJLW13]. Cybermatics [NLM^{+}16, ZZLR18]. cybersecurity [FPP^{+}18]. CyberShip [SME^{+}21, SME^{+}19]. CyberShip-IoT [SME^{+}21, SME^{+}19]. cycle [SFR15, SCJ^{+}19a]. cycles [JCA^{+}19]. CycloidGrid [GDJ^{+}13]. cyclone [RRKA19]. cytology [KIJ^{+}19].

D [AKB18b, AM19a, AKB18b, Bro19, CPD^{+}15, DJH^{+}19, GWW^{+}19, SHN10, VF18, XJY^{+}18, ZDL^{+}19]. D-S [ZDL^{+}19]. D/ [AKB18b]. D2D [YXZG18a, YZZC19]. D4Science [ACC^{+}19b]. DAC [WLML17]. DAD [KK19]. DAG [ZZZC19]. DAGs [SK12]. daisyworld [PM14]. Dalton [ASV^{+}13]. damage [QXZ^{+}19]. DAP [RAdARP19]. DAPSYS [NFK10]. DARGOS [PMLVLS^{+}13]. dark [YLJ^{+}17]. DARS [SYL18]. Data [ADBM19, AAA^{+}19, BA17, BSRR18, BMK^{+}14a, CMEA^{+}19, CLNR18, CCRL18, CMNK19, CAB^{+}18, CDH^{+}19, DYY^{+}19, DGR^{+}15, DP20, DP21a, DP21b, DLH^{+}17, EET18, FZT^{+}18, FA11a, GBB18, GRL11, HSBE19, HX19, HCX^{+}19, JL14, Jun17, KK19, KPS18, KPB18, KLH^{+}18, KT17, KIMR15, LZL^{+}16, LXF19, LWYS18, MGL^{+}18, MCR^{+}16, NNLH18, PN13, PPS^{+}18, PPA18, QC18, RPA^{+}18, SVK19, SMS16, SFR15, SE19, SNXB17, SCG^{+}18, VEET18, WRK^{+}15, WJS^{+}18, WXGM18, WZML18, WDW^{+}19, YZC^{+}19, YCY10, YDD^{+}18, YZI18, YCZJ18, ZLTY10, ZCX^{+}18, ZCYZ18, AOIS10, AHEM17, AD19, APBdI17, ABD^{+}19, AK19, AAAQJ^{+}18, AFSH^{+}18, AFSH^{+}19, ABB^{+}19a, ATH^{+}19, ARP14, AHYF19, ASO14, AIM^{+}19, ACHP19, ATdC^{+}16, AWN^{+}13, ASYF18, ASD12, AMT^{+}12, ATM^{+}19, ACK^{+}15, AGA18, ACSV18, dRADFG18, Asu13, ALL^{+}18, ACC^{+}16, AK18b, AMM^{+}19b, AMBC19, APR^{+}19, BAJ^{+}19, BLL^{+}19, BK16, BGI14, BCN^{+}19, BBD^{+}13]. data [BJM^{+}17, BAP17a, BAP17b, BAB12, BDWM17, BBC^{+}13, BDZ13, Bha18, BKY18, BDCC19, BN17, Bu18, BWG19, CZT^{+}15, CQW^{+}19, CMX^{+}16, CVT19, CD16, CLZ18, CZY^{+}19, CCD^{+}19, CTR^{+}17, CHJS^{+}10, CFM17, CGIP14, CPD^{+}15, CCMP18, CLH10, CCS^{+}10, Che13b, CPA14, CZXL18, CAL^{+}18, CRYG18, CZZ^{+}18, CLS19a, CSL19, CLDC19, CBT^{+}19, CDB^{+}19, CCM^{+}14, CPSRG14, CSP13, CSQL17, CDL18, CMZ^{+}18, CBBdL16, Cuz14, CLM14b, DW11, DHW^{+}17, DJPM18, DP19, DMMM11, DQXW19, DGD^{+}16, DGCGH^{+}17, DKK^{+}13, DLS^{+}12, DXL^{+}18, DZLA19, ELAEAVAM19, EZTL19, EAS^{+}18, EU19, EP13, ED16, FTA^{+}14, FG18, FLT^{+}19, FZW^{+}18, FD12, FSV^{+}19, FZHH14, FTK17, FNCR11, FNA11, FRM^{+}18, FS18, GQLX18, GACM17, GLC19, GVURIVBV14, GZS14, GHD19, GSC11, GJ15, GBY16, GML^{+}13, GS16b, GS15, GBF^{+}12, GCCL18, GLVC18, GCD^{+}18, GTCZG^{+}18, GDS18, GRX19, GHJ^{+}19, GG10, GFW^{+}18]. data [GLD^{+}19b, GPVN19, GSY^{+}17, GLB^{+}18, GZQ^{+}19, HKA^{+}18, HSM13, HAT19, HLYW17, HHH^{+}19, HPGMM18, HZDS19, Hsu14, HLCL16, HYS17, HDB18, HYS18, HXL^{+}18, HYF18, IAM^{+}18, IHA18, JTS13, JNS^{+}19, JH16, JHC18, JLS19, JSC^{+}15, JCL^{+}15, JFZL17, JLC18, JZJ^{+}18, JLD^{+}19, KS18a, KKN18, KRD^{+}19, KHJ10, KKBK19, KIS11, KP12, KR14, KC19a, KC19b, KSW^{+}13, KS17a, KLP19, KMC18, KCV11, KGVW14, KXS^{+}16, KLMB19, KBdLG18, KH18b, KGT15, KBB^{+}16, KAS^{+}18, KKS^{+}18a, KOP^{+}17, KK16, KAEC^{+}18, KP18, LLYW19, LLpC12, LY17, LY18a, LBM18, LN13, LXD17, LKN^{+}13, LCHW14, LXK^{+}14, LLC^{+}14b, LWD^{+}14, LLQS14, LNB14, LRYJ17, LFP^{+}17, LZL^{+}17, LGL^{+}17, LOR^{+}18, LYY^{+}18, LLY^{+}18, LLZ^{+}18a, LTTL19, LAL^{+}15, LZH^{+}18, LLL^{+}19, LJY12, LSYC18, LJ19b, LWH^{+}18, LHX^{+}18, LWW^{+}18, LSD11, LvW14, LYZC15, LSHW17, LWT18, LHW^{+}18, LRC^{+}18, LZW^{+}18, LGZY18, LKFB18, LLL^{+}18, LML^{+}19, LLW^{+}19d, LIH^{+}19, LSAM13, LKK^{+}16, LHPC^{+}19]. data [LWW^{+}13, LL16, LWR^{+}19, LZY^{+}16, LIC18, MWW^{+}15, MSS^{+}13, MNV12, MID16, MPP13, Man15, MLC18b, MVL^{+}18a, MPCAF15, Mat18, MP17, MZD^{+}16, MTD18, MMVP13, MFL18, MDT^{+}18, MFT^{+}17, MDM^{+}19, MWPVB12, MVCC10, NS17a, NNRA19, NJB19, NAD^{+}18, NK17, NWMG17, NKP16, NJ17, NSSA^{+}14, NAM^{+}19, OFD17, Osm19, OSANAM19, PPZ12, PFRC16, PJDO13, PvSS17, PVN^{+}12, PSW^{+}14, PKY^{+}17, PWMX17, PWP^{+}18, PYM18, PPG19, PECA19, PGCC^{+}10, PMBS14, PPPS18, PSY^{+}19, PSW^{+}19, PIP18a, Pon19, PRN14, QGX18, QGT^{+}18, QCX18, QCZH19, QMSG12, QC13, Qur19, RGGH18, RTHB17, RT15, RSK16, RMRSA19, RWO^{+}19, RHKC15, RGCCL18, RSJ^{+}14, SR12, SPT^{+}18, SZV19, SAGL10, SPMC10, SMPC12, SP18a, ST11, SOA17, SGKC10, SA14, SG17, SLD^{+}15, SYY^{+}17, SSJ19, SLC^{+}17, STC15, SB17b, SB17a, SGB^{+}18, SB18, SAR18b, SGBK19, Śle14, SBA^{+}17]. data [SBD^{+}18, SZD^{+}17, SYK^{+}17, SAG19, SG19, SRN^{+}18, Sta17b, SLW11, SSZ^{+}17, SLB^{+}17, SLZ^{+}18, SYQ^{+}19, SAC11, SSLF^{+}10, SJSA19, TZBK13, TZST14, THA^{+}17, TZQ18, TDC^{+}14, TF18, TZLL18, TKR^{+}15, TLSC17, TOD17, TBR^{+}19, TLL^{+}19, TSRG17, TSGVRGS19, TdPF^{+}17, TCN^{+}14, TWZP18, TSS^{+}19, TBdL16, TSTL16, TCBC18, TYWZ18, TAS^{+}18, TSB18, Tor13, TGM^{+}19b, TCCW19, TSAER18, UMUB19, USK16, VTTK17, VGD^{+}19, VOS12, VGC^{+}13, WSQ^{+}18, WZW^{+}19a, WTR^{+}13, WTG^{+}14, WLLF16, WLZ^{+}16, WZZ16, WXPL17, WWQ^{+}18, WCH^{+}18, WMY^{+}18, WZE19, WZW19b, WLL^{+}19a, Wan19, WDZ19, WLA17a, WTP^{+}13, WBJM14, WHW17, WHYZ17, WHYZ18, WLHH18, WZWW18, WHZ19, XB14, XFJ^{+}19, XWZ^{+}19, XFM16, XDH^{+}17, XLL^{+}19a, XTF^{+}19, XJZ^{+}19, XAW^{+}10, XXB19, YLVY15, YMW^{+}18, YXZG18a, YMW13, YLN15, YPLZ17, YWCC18, YZG^{+}18, YXZG18b, YWLL19, YSHM19, YKK13, YQZ^{+}19, YZN^{+}15, YXA^{+}16, YGYW16, YWY^{+}17, YYLC10, YWF^{+}10, YL16, ZZDM^{+}18, ZMP10, ZGL19]. data [ZCW11, ZME^{+}15, ZSX^{+}15, ZCK^{+}15, ZZH^{+}16, ZTKF17, ZYB^{+}18, ZFS^{+}18, ZHL^{+}18, ZWZ18, ZZZC19, ZYCZ19, ZWQ^{+}19, ZYC^{+}19, ZXD^{+}19, ZQB^{+}18, ZAC^{+}18, ZW10, ZWJ^{+}18, Zhu18, ZWGC19, ZSBB19, dSCD^{+}19, dFPFG19, uRYS^{+}19, vKvWD^{+}13, AKP^{+}18, BBD^{+}19, CAS^{+}16, CRW^{+}16, CW16, DX14, GS16a, KKAS19, MCR^{+}16, PVHTP19, Sha16, ZS16]. Data-as-a-Service [HSBE19]. data-aware [AHEM17, TKR^{+}15, ZME^{+}15]. data-based [CZZ^{+}18, GPVN19]. data-capacity [Bha18]. Data-centric [DGR^{+}15, JTS13]. data-check [LZL^{+}17]. Data-driven [FZT^{+}18, HX19, HCX^{+}19, LXF19, KKS^{+}18a, TSS^{+}19]. data-flow [GS15]. data-intensive [Cuz14, GVURIVBV14, JFZL17, KGVW14, LTTL19, RSJ^{+}14, SLC^{+}17, SBD^{+}18, TdPF^{+}17, TBdL16, TSB18, WTR^{+}13, WHYZ17, WHZ19, dSCD^{+}19]. data-locality-aware [JLD^{+}19]. data-matching [PRN14]. DataABC [JFZL17]. Database [LSG^{+}19, ASAAM^{+}19, Cha15, CWSW14, LZL^{+}16, LRMS19, PB17, WCL^{+}17a, YXD18, YWF^{+}10]. databases [ATS14, FSM^{+}18a, GAYTC18, KMB16, KYB^{+}19, LY18a, LLCF11, MQN19, NA19, PDDS10]. Datacenter [YMY^{+}17, ESPN17, KMJ18, LKTC14, SHRE16]. datacenters [LJGW18, LPBB^{+}18, MRH17, SSP17, WCC14, XWM18, ZG19]. datacentre [Len16]. datacentres [PLLA18]. dataflow [PPLL17]. dataset [KMST19, WZZ16]. datasets [AMBB18, Ans11, BRXdS11, CLY14, CYJ19, FGM11, HHZ16, MR19, PPZ12, RVST17, WMLS14, ZLXZ18].

Dataspace [CDH^{+}19]. Dataspaces [HLCL16]. date [HMA18b]. day [KLW^{+}18]. DB [PYH^{+}18]. DBaaS [BBB^{+}19]. DBMS [MGV^{+}18]. DBMS-based [MGV^{+}18]. DC [QMCX19]. DDoS [CWLZ19, DCC13, SPT^{+}18, TA18, VS13, VSP^{+}14, ZJW^{+}14]. Deadline [ANE13, PNGFJ13, VCKB12, ZT19, APAZ17, ABP16, ABN17, ABN19, DQLW15, FCY18, HLW12, KDG^{+}19, MJDN15, SGJ18, TWdLZ19, TSB18, VVB13a, WTM^{+}17, ZYB^{+}18, ZQB^{+}18]. deadline-aware [ZQB^{+}18]. deadline-based [KDG^{+}19]. Deadline-constrained [ANE13, ZT19, APAZ17, MJDN15, SGJ18, TWdLZ19, VVB13a, ZYB^{+}18]. Deadline-driven [VCKB12]. deadlines [BBB16]. deadlock [CWW^{+}16]. deal [MCG^{+}15, SHBP10]. deblocking [WC14]. decay [PKA19]. December [Ano19n]. decentralised [PWA^{+}19, STMV18]. Decentralized [CLNR18, FY19, HNKÖ18, JTS13, LJC^{+}19, MPC^{+}18, ÖEE13, RLP12, SHRE16, TGM11, XTZ^{+}19, ATF11, ABH18, CCL11, DA18, DCF19, FWB13a, HBN^{+}13, MML^{+}18, MT17, MOBD18, MGLPPJ13, NLV^{+}19, RRB10, SAK19, SCMS12, TY11, YLA18, ZMH^{+}18]. deceptive [CWZ^{+}17]. decide [MOBD18]. decision [AFSH^{+}19, APR^{+}19, BBWB^{+}18, BKB18b, CLY14, GBKJ18, JNS^{+}19, JXZ^{+}19, KMI11, KKS^{+}18a, KFBKD14, LPK17, LPK18, LRMS19, LYS12, LDY^{+}18, RT16, TSS^{+}19, VDK12, WY17, XYLZ18, YCH19]. decision-making [APR^{+}19, JXZ^{+}19]. decisions [ABMMC18, ABMMC22, LTZ15, TWdLZ19, WGM15]. declarative [ATdC^{+}16]. decoction [QXZ^{+}19]. decoding [PSPP16]. Decomposed [YSC^{+}19]. decomposition [AB18a, Bu18, FBM19, Ima19, LLW^{+}18a, LYW^{+}18b, MSBA16, NJ19, PLL^{+}18, PSLZ18, SMC18, WHZL10]. decomposition-coordination [MSBA16]. decoupled [LSJ^{+}14]. decrease [KIS11]. decreased [WCWC19, WCWC20]. Decreasing [VGC^{+}13]. decryption [XTZ^{+}19, XYML19, ZSW^{+}18b]. dedicated [BBB16]. deductive [RB12]. deduplicated [MKRD19]. deduplication [FLL^{+}19, KH18b, SAR18b, WLA17a, WLA17b, XFJ^{+}19, ZFH^{+}18, WLML17]. Deduplication-Assisted [WLML17]. DEED [YZC^{+}19]. Deep [ASYF18, AAM^{+}19, AMM19a, DDD^{+}19, EU19, GSC^{+}19, HDA^{+}19, LYYW19, SB19c, ZTC^{+}19, ZYC^{+}19, ZZLZ18, ABDH19, AQAR^{+}18, ASAAM^{+}19, BWG19, BW19, CLCMG^{+}18, CFMC19, CAL^{+}18, DFG^{+}19, DC18a, GLC19, HDKC18, HUMA18, HHH^{+}19, KMK^{+}19, KSS19, KLJS19, LBD^{+}19, LLH^{+}17, LZT^{+}19, LXT^{+}19, LLU^{+}18, MFSV19, RSY^{+}18, SD18, SYT^{+}19, SHL^{+}19b, TBS^{+}18, TYWZ18, UMUB19, WCB^{+}18, WZF^{+}19, YWLL19, ZZ19]. deep-learning [CAL^{+}18]. deep-regression [AQAR^{+}18]. DeepNet [CPW19]. defense [ASA19, CWLZ19, DCC13, GHYK18, NNC^{+}19, VSP^{+}14, ZJW^{+}14]. deferrable [CSQL17]. define [ABD^{+}19]. Defined [BRH18, CWL^{+}19, IDKD19, RGSL18, SMG18, ACHP19, CJK^{+}18, CKP^{+}19, CS19, GZLZ16, GHYK18, GTSP^{+}19, GXL^{+}18, HYG^{+}19, JAAD^{+}16, KJI11, LGP^{+}19, LLT^{+}19, LLW^{+}19c, LXM^{+}18, LRC^{+}18, MKRD19, NAGD18, QCY^{+}19, WLA17a, XFJ^{+}19, ZGL19, ZWJ^{+}19a, ZBCT17, ZZSZ18]. Defining [UDST19, MEBA12]. definition [GHO^{+}11]. deforestation [ALFR16]. deformable [RDSA18]. deformation [JHC10]. defragmentation [RT15]. Degeneration [TBS^{+}18]. Degree [ZSFZ19, LJJ18]. Degree-biased [ZSFZ19]. Delay [ACC^{+}19a, BARMB14, RJS^{+}19, CLRL17, LCL14, RMDB18, SZK16, VSKS19, ZTKX19]. delay-based [CLRL17]. delay-sensitive [LCL14]. delegated [SBL18]. delegation [AH11, JSMG18, RPH19, XZP^{+}19]. deletion [LZLL18a]. Delivering [ACC^{+}19c, PMDS18, QCD16]. Delivery [SST18, FAL^{+}19, HAT19, HDLW13, KKN18, LMZ^{+}14, RMVG^{+}10, Zhu18]. deluge [TGM^{+}19b]. Demand [CAB^{+}18, SLSS19, DEG^{+}17, Len16, LWH^{+}18, LWZ18, MHW^{+}16, PYH17, SH19, TDC^{+}14, SCN^{+}14]. demand-oriented [LWZ18]. demands [KPM^{+}18]. dementia [NDZ^{+}18a, NDZ^{+}18b, NDZ^{+}19]. demonstration [VRGR16]. demosaicking [HLC16]. dendritic [EAA16]. Dendron [PMK18]. Deniable [DA18]. Denial [IDKD19]. denoising [BTG19, GDAS18]. Dense [RCW^{+}19]. DenseNet [CTU19]. Density [GBY16, ZSZ14, LZXW13, LZZ19]. Density-based [ZSZ14]. deontic [SZR18]. dependable [BCC^{+}17, WY17]. dependence [XFM16]. dependencies [BBI13]. Dependent [BB17, GTMZ17, MCRB19]. deploy [ALTG19]. deployed [KSS19]. deploying [MVC^{+}13]. Deployment [TCB^{+}17, BARMB14, BGRBA19, CZY^{+}18, CGL15, KTKN11, LWH^{+}18, LLW^{+}12b, LSMVML13, MAJD18, MGLPPJ13, MCSA18, MW12, RCMT18, ŠCJ^{+}19b, VCL^{+}19, VSDD13, dOWdAS^{+}18, WG13, WLH^{+}19, ZLZ13]. deployments [LPD^{+}13]. depression [CFL^{+}15]. Depth [HZZ^{+}18, XTL^{+}19]. derive [LN13]. derived [YMW13]. Deriving [MdOO^{+}17, CHSA18]. descent [MWL18a]. describing [BBT19]. description [DJPM18, KGdL11, PLCGS11, Sun10]. descriptions [XLL^{+}18a]. descriptor [LZL^{+}12, PSS^{+}18]. Design [AMN18, AAB^{+}10, Bag19, CBS17, CLS19a, DCBF19, DVD12, HZDS19, JO11, JNR12, KO11, KANS18, KLW^{+}17, LJS17, MK17, MCWP16, PMT10, RPH19, SPR^{+}10, SYCH18, TMM^{+}13, TBB^{+}17, WDKV19, XKJ^{+}18, ZYA^{+}18, ZZQ^{+}13, dSK^{+}19, AAD^{+}13, BFP18, BSRR18, DSD^{+}11, DBD^{+}14, ECPF17b, FZT^{+}18, GdCP19, HIA^{+}18b, HX19, HLL^{+}19, HHZ19, JLQZ18, LRYJ17, MOBD18, MNY^{+}19, MHY^{+}18, MWMA10, NSR^{+}19, OSANAM19, RHH^{+}16, RCOP^{+}11, TWW^{+}18, UZ11, VPP^{+}19, WWD^{+}14, WXYL15, WSL^{+}19, WGM15, dOWdAS^{+}18, XZL^{+}19, ZDL^{+}13, ZLG^{+}14]. Designed [DGR^{+}19]. Designing [AB18a, LTMW19, OdOD^{+}13, XLL^{+}19b, CFPC17, DGCGH^{+}17, Mér17]. designs [BDG^{+}19]. desirable [PKB19]. desktop [CLL^{+}14, FK12, KWK16, KJ12, KKL11, LJLW13, RLP12, SWW^{+}18, TPBS14, TKTG19, VKK14, CB10]. DESRP [LWZ18]. desynchronized [Tur18]. details [SCLC19]. detect [JXC^{+}19, LYS^{+}19, PZC19, SAG19]. detected [BCF16]. Detecting [BOHCC17, RKB18, WLW^{+}18, AMM16, CZ14, LA19, SM18]. Detection [GMLGB^{+}17, GHP^{+}18, GLXF17, LMM19, SAPA17, ZJW^{+}14, AD18, ABDH19, AMI16, ASAAM^{+}19, ArMS19, AIP^{+}19, AHMS18, ASYF18, AKM18, AS18b, ALFR16, AM19b, BTG19, BBH18, BR19, CXZC18, CYW^{+}19, CTU19, CCDP19, CRC^{+}19, CWLZ19, DJPM18, DNJG17, DC18a, ENC^{+}12, FJ18, FZHH14, FS19, FCD^{+}14, GAFFOG12, GPV^{+}14, GRS^{+}19, GNGG17, GSY^{+}17, GSC^{+}19, GMCM16, GMCM18, HTL^{+}18, HJA^{+}19, HNCJ13, HZZ^{+}18, HZW^{+}16, HAA^{+}16, HIA^{+}18c, JNHL18, KHWZ18, KIAD17, KIJ^{+}19, KAW12, LRL^{+}14, LYJ10, LLN^{+}18, LJJ18, LTMW19, LYC^{+}19, LYXT14, LSL^{+}15, LJW^{+}19a, LXT^{+}19, LYYW19, LWR^{+}19, MH19, MAY18, MGA^{+}18, NK18, NGB18, NJ17, NO19, PMK18, PRW14, RBGA18, SPT^{+}18, SC19, SOD18, SD18, SMRM13, SI18, SSA^{+}19, SEHS19, SYT^{+}19, SCLC19, SB18, SGBK19, dSSCdL19, SGS^{+}18, SBK18, SJL^{+}18, TBS^{+}18, TMB^{+}19, VSBN19, WWH^{+}17, WJS^{+}18, WWZZ18, WXZ^{+}18b, WCM^{+}19, XFTZ16, XJZ^{+}19, YWJ^{+}18, YARH18]. detection [YCXW18, YJL^{+}19, ZZH^{+}18, ZGV19, ZRZR19]. detector [MdFTGM19]. deterioration [WWX^{+}17]. determinant [GSC11]. Determinants [CT19c]. determination [KK10b]. determine [FM10b]. Determining [SOA17, SS17]. determinism [WPGN^{+}18]. deterministic [GOLL19, WLGL19]. Deterrence [SMF^{+}19]. develop [Ham19, ZBF14]. Developing [GTM19, HZZ^{+}14, HCJ14, LLMP13, SR19, Ham17, Kim14, LJS17, LKJ17, SWW^{+}13, TKA^{+}18a, WCH^{+}18]. development [CGL15, FSP^{+}18, KMST19, MMPF19, SBLW14, Sip12, SHJS^{+}10, ŠCJ^{+}19b, TBK^{+}10, TF17, TCBC18, dFBP^{+}17]. developments [SSZ13, YAGG18]. Deviation [WYH^{+}17, HKP10]. Deviation-based [WYH^{+}17]. Device [HCHH19, WWVJ17, CLK11, DSBC19, FRM^{+}18, KPG19, LPL^{+}16, MKS18, SZG^{+}19]. Device-to-Device [HCHH19, WWVJ17, DSBC19]. Devices [BDM^{+}19, WLP18, APRC16, AT19a, AKB^{+}18a, ACD^{+}19, BSRR18, BOP^{+}14, CRRC18, CBPP18, CRC^{+}19, DC17, EYY19, GMM18, HMZ18, HCW^{+}18, JCA^{+}19, JKLK17, KMU19, KKA18, LKJ^{+}19, LJ19b, LCMX16, LNY^{+}18, LYL^{+}19, MMC^{+}18, PCK19, RSRA18, SAGGB17, SGS^{+}18, TLC^{+}15, VFHB14, Wan18b, WWG19a, WGX^{+}19, ZDM^{+}19]. DevOps [WBKL16]. DEXIN [FLT17]. dexmedetomidine [FHHM19]. DFINT [QC18]. DGC [TF18]. DHT [HNKÖ18, HZ10, PTT12, TJWS10]. DHT- [TJWS10]. DHT-based [HNKÖ18, HZ10, PTT12]. diabetes [GP11, VFHB14]. diabetic [NDA^{+}19]. diagnosis [AAN^{+}18, AASI17, CLZ18, CFL^{+}15, KHO^{+}19, KLV^{+}18, SB19c]. diagonal [PSS^{+}18]. diagram [WZW^{+}19a]. Dichotomous [RRKA19]. dictionary [DPS16, ZWWL18]. difference [BG12, LLL^{+}19]. different [BARMB14, DFGR14, JZWL17, RJN^{+}19, RWY^{+}18, SHH^{+}19, SCG^{+}18]. differential [HZL18b, KS18c, LWZ18, PSY^{+}19, TM19, WZE19]. differentially [WLL^{+}19a, ZLXZ18]. differentiation [TSBH11]. DiffServ [BLO^{+}18]. diffusion [DH16, TWG^{+}19, WRCC17, CSV^{+}12]. diffusive [BMZ10]. digest [BK19]. Digital [AFS16, LCHW14, MLC^{+}11, QC13, QC18, SR19, ACMM19, Bro19, CCMP18, DGA18, GTM19, LCH^{+}11, LQF19, MJGW18, PKP19, QCX18, SI19, SSFFR19, SSA^{+}19, Wei11, ZSMS18]. digitised [OPT^{+}17]. digitized [SYT^{+}19]. dimension [ZAB15]. dimensional [CFL^{+}18, DJPM18, EP13, FNA12, GDAS18, HZW19, PM14, SWG^{+}16, SS17, SJL^{+}17, WZW^{+}19a, WW11, YPCK12, ZFW14, ZZZ17]. dimensionality [MZYA19]. dimensions [EHT10, WMBV17]. Direct [EDH^{+}13, ECE^{+}19]. directed [DK14]. direction [DHW^{+}17, ZG18]. direction-preserving [DHW^{+}17]. Directional [CLL^{+}18a, ABS^{+}18, XZW^{+}19, YPCK12]. Directions [CDFZ16, CGBAP18, GVBG17, GBMP13, LLWZ18, MND^{+}19, QKC19, URC19, VB18, YS16, dCTVC18]. disabilities [RMSPP17]. disabled [HFT16]. disaster [AT19b, DCC^{+}14, NKP16, RCMT18, SGRT19, SA14]. disaster-inspired [AT19b]. disciplinary [LZH^{+}18]. disciplines [LKK^{+}16]. disclosure [LNY^{+}18]. discomfort [XJY^{+}18]. Discover [ALTG19, AVPV17]. Discovering [LLG^{+}16, WFQ^{+}10, Fer13, LC15]. Discovery [BBC^{+}13, BCF16, KK11, AAS^{+}19, BMH10, BCR^{+}12, CWJD19, CCT13, ÇÖ13, CH10, DMPS19, EK11, EBOY14, EA17, FFPS10, GACM17, GDJ^{+}13, GB10, HCC^{+}14, HZC10, HSV^{+}17, KKL11, KIC12, LCBF13, Li10, LZH^{+}18, MLL15, MCL^{+}16, MHC14, Pal13, Pip10, PTT12, PPL^{+}15, SB17a, YSC^{+}19, YCZJ18, YLA18, ZAB15, ZL13, ZZX^{+}19, ZCL^{+}19]. discrete [BMP^{+}16, LCW^{+}18]. discretization [LTJK12, RGGH18]. discriminant [CLY14, HLL12, PZC19, YPCK12]. discriminate [AHMS18]. discrimination [HLH^{+}18, XLL^{+}19c]. discriminative [WHYZ18]. discussion [BOHCC17]. disease [AAN^{+}18, ARP^{+}19, AASI17, DRC^{+}19, HUY^{+}19, KHO^{+}19, KLV^{+}18, NDA^{+}19, WCWC19, WCWC20, XL19, ZZXL18]. diseases [HEES19, SM18, VFHB14]. disk [DLXR14, JLI^{+}13, NQQL13]. disorders [AHMS18, CPE^{+}17]. dispatch [WLH16]. dispersed [ED19]. dispersion [ZWZ19]. Display [KID^{+}16, KWK16, NRR^{+}15, RRH16, WY19, XJY^{+}18, YDK11]. displays [BOP^{+}14, PDK10, RMA^{+}16]. dissemination [BK16, GRTV10, MZD^{+}16, ROK19, SJ14, TLL^{+}19, WLP10, ZWJ19b]. Dissipative [WZL18]. Distance [PYH^{+}18, TKA18b, ZL13, CZL^{+}18b, HKU^{+}11, LYW^{+}18a, OMKM^{+}19, SPT^{+}18, WWD^{+}14, ZSGJ19]. Distance-aware [ZL13, ZSGJ19]. distance-fitness-based [CZL^{+}18b]. distinguishing [SM18]. distress [MNC^{+}18]. Distributed [AWYJ16, Bag16, CGV10, CZY^{+}18, CMT16, CMNK19, DRC^{+}19, DC18a, EH10, GLVC18, HCZW17, KKS^{+}18a, LSG^{+}19, ML17, OdI14, PT16, PZY16, PIP18b, SSG17, SSL13, TCG14, VVC^{+}12, WMLS14, XHL^{+}19, ABS11, AdI14, AS19b, AB18a, AB19c, AKB^{+}18a, ABF^{+}15b, AB16, ADLM18, Bag11, Bag19, Bal16, BVFGWA15, BC15, BDZ13, BR19, BGMLS17, BRNR15, CMZ^{+}12, CPGdS^{+}13, CZL^{+}18a, CZY^{+}19, CRC13, CLR16, CDDR17, CCL11, CBT^{+}19, CCM^{+}14, CPLH19, CGM^{+}18, CBBdL16, CLM14b, DR15, DdM10, DDB14, DM12, EK11, EBOY14, ECPF17a, ECPF17b, FGG13, FFPS10, FM10a, FSP^{+}18, GA13, GAI^{+}18, GIM16, GM11, GKTK15, GSY^{+}17, HKPT10, HST^{+}18, HZDS19, HLNM11, IS18, JRJ^{+}11, JLD^{+}19, KMB16, KS11, KANS18, KYB^{+}19, KB18, KKBK19, KBB^{+}16, KOP^{+}17, KJ18, LCBF13, LZL^{+}17, LGW^{+}17, LLW^{+}19c, LC15, LSH^{+}11, LZXG12, LKK^{+}16, LMA^{+}19]. distributed [LBU^{+}10, LCCM18, LM12, MYHZ18, MDA^{+}19, MSS^{+}16, MLSF16, MAA^{+}19, MKS^{+}19, MGLPPJ13, MROD10, MW12, MFSV19, MQN19, NFK10, ODK^{+}17, PBV^{+}13, PdASM18, PSW^{+}14, PGTBC18, PMT10, PZY17, PSBB15, RHH^{+}16, RLRC13, RMHCMG15, RMHMG17, SB19a, SGRT19, SLG^{+}17, SMSF18, STB^{+}19, SSLF^{+}10, TBdL16, THT12, UDvdW^{+}18, VPT^{+}15, VOS12, WHZL10, WTR^{+}13, WWW^{+}16, WHS^{+}17, WWG^{+}19b, WHYZ17, WHYZ18, XWZ^{+}19, XXX^{+}19, XXB19, YJS18, YZW14, YWF^{+}10, ZWL13, ZWW^{+}13, ZSX^{+}15, ZWHC17, ZTKF17, ZW10, ZB19, ZSBB19, dSGD13, dSFD^{+}19, dLB10, BCF^{+}10, YMY^{+}17]. Distribution [KK19, KT17, AC18, BGP^{+}17, CRC13, CdRRdCB19, DJH^{+}19, FHZW18, GZL^{+}18, HFM19, HAAWH^{+}18, KTKN11, LBD18, LJY10, LSYC18, LJW^{+}19b, MOFGP18, NK17, SA14, Tor13, USK16, VCDK18, ZTKF17]. distributions [BARMB14]. Disturbances [VLAC^{+}13]. divergence [VS13]. DiVers [HDO16]. diverse [BCdV^{+}19, CCJ16, CFPC17, NK18]. diversified [LYW^{+}18a]. diversity [BCdV^{+}19, SLW11, ZSJ19]. divide [KTB18]. divide-and-conquer [KTB18]. divider [DJJ^{+}18]. divisible [AOIS10, GOBL16]. DMPO [YCX18]. DMZ [XLL18b]. DNA [NRV^{+}17, ZRZL18]. DNS [LKJ^{+}19]. do [ACBM15, AMRM18, LMZ^{+}14]. do-it-yourself [LMZ^{+}14]. DOC [SSG17]. Doc2vec [CXZC18]. Docker [DSS19, DL19]. docking [IPG^{+}18, TKTG19]. document [LLS^{+}14, TDBR18]. documentation [WGM15]. documents [BD18, CZ14, CZ19, SW17, YJL^{+}19]. Doddington [MCRB19]. DOE [RS16]. DOM [GGC18]. DOM-Based [GGC18]. domain [AMI16, FSM^{+}18b, GZLZ16, GDAS18, GSY^{+}17, KZA^{+}18, LKA^{+}19, SC19, SMC18, SCEC18, SLY^{+}19, SSB13, WMBV17, XWZ^{+}19, XLW^{+}17, YZL^{+}18]. domain-oriented [XLW^{+}17]. domain-specific [KZA^{+}18]. domains [AL14, KOT18, LLM^{+}16, SG17]. dome [KO11]. domestic [SH19]. dominance [ZZS^{+}19]. Dominated [BRH18]. Donating [WPJ16]. door [ZTC^{+}19]. Double [LEW19, DCBF19, LBJ^{+}18, LBJ^{+}24, LY18b]. double-orbit [LY18b]. down [SCH^{+}19]. download [Li15]. DPM [CGM^{+}18]. DPRank [LXM^{+}18]. dragon [CLR16]. Dragonfly [CT19a]. drainage [ZMZ^{+}19, ZMZ^{+}20]. DRAM [CSJ^{+}17]. drift [HXC^{+}18, IdAP19, TSRG17]. drilling [LZL^{+}19a]. drive [SYW17, VS19]. drive-thru [SYW17]. driven [AJY12, BFS^{+}17a, BFS^{+}17b, CVKB12, CSL19, DCC^{+}14, DWS12, FFC^{+}18, FZT^{+}18, FTP14, FG14, FFC12, GEG14, GBF^{+}12, GMP^{+}16, GHJ^{+}19, GKTK15, Ham17, HHK18, HB19, HX19, HCX^{+}19, KCK16, KADJ14, Kol18, KKS^{+}18a, LJC^{+}19, LXF19, LvW14, MFC^{+}19, NJH^{+}18, NJ16, Pal13, PMK18, PdAF12, RZ16, RBN13, SMS14b, TSS^{+}19, TCBC18, VCKB12, VETT16, VPP^{+}19, WLH^{+}19, XWW19, XJZ^{+}19, YMLT13, ZCM19, ZAA^{+}14, ZCX^{+}18, ZLT^{+}19]. driver [JCMPPC^{+}18]. drivers [PWP^{+}18]. drives [PLZX19]. driving [FPL^{+}19, OMKM^{+}19]. drone [ASAAM^{+}19, LLN^{+}18]. drones [PECA19]. Drop [CNP^{+}19]. droplets [QC13]. DRTHIS [HDA^{+}19]. DRTM [DJZ^{+}15]. drug [HSV^{+}17, KKL11, MMVS19]. drug-protein [MMVS19]. drugs [WWA19]. drying [Jun18]. DSM [CCS^{+}10].

DSM-based [CCS^{+}10]. DStore [XXB19]. DT [vKvWD^{+}13]. DTA [XWJ^{+}16]. DTN [WSZC18]. DTRM [LHX^{+}18]. Dual [HPL^{+}19, BWR12, BW13, ESPN17, QCZH19, VAdlP12, XZZ^{+}19]. dual-band [VAdlP12]. dual-port [ESPN17]. duct [HZX^{+}19, HZX^{+}20]. duplex [WGX^{+}19]. during [AKM18, BDM^{+}19, BWR12, RS17b, SK18, SLC^{+}17, SE19]. DUSKG [WWH^{+}19]. DVFS [BBC^{+}12, KAEC^{+}18, SK19, TuIS^{+}19, WCC14]. DVS [LLC14a]. DWARM [WZS^{+}18]. DWT [AM19a, KHO^{+}19]. DXT [HSP^{+}13]. Dyck [SAM^{+}19]. Dynamic [AMT^{+}12, ABN19, ABB19b, BK19, CBC^{+}19, CsZzG^{+}13, DLDTGMMP16, EG18, GGDM^{+}18, HYG^{+}19, JEB18, KKK^{+}19, KMK^{+}14, LWTL19b, LD17, LZL^{+}19a, LCL^{+}18, LWSY18, MNV12, MDO^{+}15, NV11, NSSA^{+}14, PSPP16, RSJ^{+}14, SME^{+}21, ST11, SE19, SCJ^{+}19a, TMW^{+}17, TMP15, TSAER18, WCHL10, XWJ^{+}16, YZC^{+}19, YCX18, ZAB15, AHL11, AL14, ASD12, BML18, BN17, BTM10, CLRL17, CDDR17, CJHH13, CWJ^{+}18a, CGSV17, CNP^{+}19, CS19, CBBdL16, DSK^{+}14, DBP19, DGR^{+}15, FLT17, FMN^{+}17, FAL^{+}19, FWB13a, GQLX18, GEG14, GS16b, GB10, GLB^{+}18, GZQ^{+}19, HZW^{+}18, HRVW18, HMW14, HBN^{+}13, JBR^{+}16, JSS^{+}12, JK17, KIS11, KCM19, KID^{+}16, KC19a, KH18b, KAS^{+}18, LY18a, LXJD18, LYYY18, LWTL19a, LHM14, LJL12, LN18, MPC^{+}18, MWCK19, MTD18, MDD15, MSE19, MW12, NPP12, OWX19, PAL^{+}19, PNZ14, POJ^{+}16, PF17, RML^{+}19, RB18, SR12, SME^{+}19, SHP^{+}16]. dynamic [SRZD15, SLD^{+}15, SLG^{+}17, SGJ18, SSC^{+}19, SYAL13, SYL18, SYQ^{+}19, UDST19, VAdlP12, WCM^{+}19, WWZC19, WDZ19, WLA18a, WZ13, XWjZyF19, YZL^{+}18, YZ12, YMY^{+}17, YDQC19, ZSMS18]. dynamical [BKB18a, XTT18]. dynamically [ABN17, LKJ17, QCYJ17].

DynamicCloudSim [BL15]. dynamics [BMZ10, GPS^{+}17, JTB13, KVK^{+}18, MCRB19, MR19, WJLW18].

e-commerce [ZL18]. E-Health [AMSPL19, WMX^{+}17, KIJ^{+}19, RGN^{+}18, VRS^{+}19, PKY^{+}17, YZL^{+}18, YZG^{+}18, ZAA^{+}14]. e-Healthcare [IFD^{+}19, ZZXL18, JNS^{+}19, LZLL18b]. e-infrastructure [MSS^{+}13]. E-learning [CJN^{+}15, CLL^{+}14, TNY17]. e-markets [VPT^{+}10]. e-payment [CLM^{+}14a]. e-Science [BDP11b, CBN16, SAGL10, AC10, BH13, KA13]. e-Social [LSAM13]. e-VLBI [WWD^{+}14]. E2S2 [KLP19]. eagle [GJGB19, LPY^{+}18]. EAMSuS [MPI^{+}18]. Early [AAJ17, ZPPE17, ArMS19, SPT^{+}18, UZ11, VMN^{+}18]. Earth [CCM^{+}14, GSGPP^{+}19]. earthquake [BSE^{+}13, ZPPE17]. Easy [WWG^{+}19b, ZDW^{+}16, vdPGZ^{+}16]. Easy-to-be [WWG^{+}19b]. easy-to-use [vdPGZ^{+}16]. EasyGateway [GRZ^{+}19]. EATDDS [ARSMY19]. Eavesdrop [FZW^{+}18]. EbH [GMdFPLC17]. EC2 [BC15, FEÁ19]. ECG [GMdFPLC17, HZW19, IASK14, PVN^{+}12, PLGMCdF18, SD18, WZY^{+}19, WPS^{+}18]. ECG-based [PLGMCdF18, SD18]. ecoinformatics [PWB^{+}13]. ecological [KSSG16, PSW^{+}14, TSTL16]. Economic [AB19a, ABGMC19, AB21, TBR^{+}19, YW12, ATM^{+}19, CC19, DQC^{+}19, HAP11, HAP15, LTN10, RMCMD12, WDZ19, dACAM13]. Economic-based [YW12]. Economical [CFH^{+}19, GFW^{+}18, LLZ^{+}18a]. Economics [AAC^{+}19, ABP18, BAV16, VAR14, ARB12, AR10, TL19]. economy [ABP18, FdAGdAFV19, KHG13]. ecosystem [AMPZ16, DDMPG17, GHGP19, GPJA^{+}14, MVL^{+}18a, Sha16]. ecosystems [CGM^{+}19, SRdlPG19, ZZDM^{+}18]. ECU [KR19]. EDAK [ABB19b]. EDAWS [WHYZ18]. Edge [AJR^{+}19, AR17, AHM^{+}18, CLH^{+}18, DMC^{+}19, GSP^{+}17, JJH19, KAH^{+}19, LOR^{+}18, LSCL19, MPLM18, OFMZ18, SST18, WZM^{+}18, ZZF^{+}19, AHS^{+}18, ANG^{+}19, AHU^{+}19, AK18b, AMBC19, BCN^{+}19, CFG^{+}19, CXWT19, CGSV17, CGM^{+}19, DLL^{+}19, EZTL19, EBCP18, GRX19, GCZ^{+}19, HMZ18, HMA18b, HMW^{+}19, HLT^{+}18, HLT^{+}19, IG12, KZ17, KA19, LW18a, LTTL19, LWTL19b, LWTL19a, LY18b, LWZ^{+}19b, LCY^{+}19b, MLC^{+}18a, MCAS19, MBMTJR18, QZD^{+}18, RGN^{+}18, RR18, RLM18, SPJ17, SYJA19, SCAC^{+}19, SLL^{+}18, TGM^{+}19a, WMJW18, WLH^{+}19, XZW^{+}19, XZP^{+}19, XLL^{+}19a, XXQ^{+}19, YCH19, YCX18, ZZ19, ZYC^{+}19, GQXL18]. Edge-based [ZZF^{+}19, CGSV17]. Edge-centric [AHM^{+}18]. edge-cloud [AK18b, LWTL19a]. edge-crowdsourcing [SYJA19]. edge-IoT [EBCP18]. Edge-of-Things [GQXL18]. edge-stream [GRX19]. edge/cloud [ANG^{+}19]. Edges [BDL^{+}19, ZZF^{+}19]. edible [Jun18]. editing [LHCC18]. Editorial [ADLW12, Ano19b, Ano19c, Ano19d, CRW^{+}16, GVTdL18, PC18b, RB13, SZV19, TKRA14, YAGG18, FBS18, SSZ13, Ano11c, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano14e, Ano14f, Ano14g, Ano14h, Ano15k, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano16v, Ano16w, Ano16x, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano19a, Ano19e, Ano19f]. Editorial [Ano19g, Ano19h, Ano19i, Ano19j]. Editors [BB13]. EDS [DLH^{+}17]. education [KACN16, MM18, SG19]. eduGAIN [TOS18]. EEG [AHD^{+}19, AAM^{+}19, CMI^{+}19, WZY^{+}19]. Effect [FHHM19, PLGMCdF18, Bo19, Bo20, JLC^{+}20, RWY^{+}18, WWP19, WWP20, YWG^{+}19, YWG^{+}20]. Effective [KV17, PKF14, SCBK^{+}16, CWJ^{+}18b, CSQL17, DMC^{+}19, DLS^{+}12, HHL11, HYC^{+}18, KS18a, KWB19, KK16, LML^{+}19, MFG^{+}14, PFRC16, RT16, SBA^{+}17, SK19, WWH^{+}17, WMX^{+}17, WWH^{+}19, WY17, ZBCT17, ZRZ^{+}14]. Effectiveness [YZI18, ECPF17a, Man15]. Effects [JK17, QXZ^{+}19, CN17, GPS^{+}17, LRYJ17, LLS^{+}19, PM14]. efficiency [DDB14, FTK17, GLJ19, IDM^{+}16, JLRS18, KSSG16, KHG^{+}18, KKvdB^{+}17, KV12, KdGP^{+}19, LSB^{+}18, LPL^{+}16, PLLA18, SLD^{+}15, SEPV19, Sta17b, SG15, TKA18b, THT12, VEET18, WXLY16, vKLA^{+}19]. Efficient [ABB19b, CXL^{+}17, CAPG18, CXC^{+}18, DLH^{+}17, ELAEAVAM19, FXG^{+}19, FGG13, HFM19, HZL18a, HDLW13, KCV11, LLYW19, LQK^{+}16, LAL^{+}14, LXK^{+}14, LHO17, LFHQ19, LLL^{+}19, LNLA19, LLM^{+}16, NWMG17, OBG^{+}18, PRS^{+}13, PWMX17, PPS^{+}18, PRN14, RM11, SJ14, SZW^{+}19, TLC^{+}15, VKK14, WCC^{+}16, WHS^{+}17, WCM^{+}19, WZCH17, XWX^{+}17, YZC^{+}19, YHL16, YLA18, YNLY19, ZGB^{+}17, ZXJ^{+}14, ZWL^{+}16, ZFS^{+}18, ZWJ^{+}18, ABMM18, AR15, AQRH^{+}18, ARP14, AHYF19, AS18a, AAM^{+}16, AK18b, AMBC19, BK16, BZ19, BAB12, BMH10, Bu18, CMX^{+}16, CRM^{+}16, CP17, CFMC19, CLM^{+}14a, CSJ^{+}17, CZXL18, DSBC19, DHW^{+}17, DVD12, DQLW15, DR18, DNP14, ECE^{+}19, GdCP19, GD10, GK18, GBF^{+}12, GCZ^{+}19, HMH17, HXY13, HYS17, HYS18, HLT^{+}19, IOV^{+}18, JJH19, JYY^{+}17, JLCC12, JFZL17, KKN18, KKB18, KMT14, KLP19, Kyr19, LBD18, LY17, LY18a, LBM18, LLQS14, LCL^{+}16, LLC^{+}16, LYY^{+}18, LLW^{+}19a, LAL^{+}15, LSYC18, LC15]. efficient [LZXG12, LG16b, LCMX16, LWZ18, LFY^{+}19, MNV12, MPP13, MMF16, MP17, MZD^{+}16, MLM16, MAÇ17, MMRL17, NRV^{+}17, NDA^{+}19, NWL17, PSPP16, PPZ12, PDH18, PLP^{+}19, PPMM^{+}18, PF17, PPLL17, QZM^{+}18, QMSG12, Qur19, Rao17, RGSL18, RCTY19, SB14, SC16, SAGL10, SJTN18, SSSJ19a, SGJ18, SRN^{+}18, SCBK^{+}16, SK19, SYAL13, SLA^{+}16, SLB^{+}17, SLY^{+}19, TDSH16, TPBS14, VVB13a, VTTK17, VS19, VCD^{+}18, VGD^{+}19, VGC^{+}13, WHMO13, WCC14, Wu16, WPJ16, WHYZ18, WLHH18, XZZ^{+}14, XZL^{+}19, XDH^{+}17, XLL18b, YPLZ17, YYD^{+}14, YZWG18, YDT19, ZL13, ZTKF17, ZCL^{+}18, ZZSZ18, ZZXL18, ZLL^{+}19, ZYCZ19, ZXD^{+}19, ZFC17, ZFC18, MPI^{+}18]. Efficiently [XTF^{+}19, ZZC18, BSE^{+}13, LFH^{+}15]. effort [MZC10]. efforts [GDZ^{+}19]. EFG [YGY^{+}19]. EGEE [FKBG10, VHML10]. EGI [DRZ^{+}19, SLSS19]. Ego [GPJC17]. eHealth [FFC^{+}18, PZA18, dSGD19]. eID [FRZ19, SGGCR^{+}16]. EKF [BRB19a]. EKF-MRPL [BRB19a]. Elastic [ISS^{+}15, LML^{+}19, BKKM11, CRM^{+}16, CVKB12, CTVB12, CLNR18, FEPC18, GGS13, LPBB^{+}18, LSJ^{+}14, MWQ^{+}14, SGL^{+}19, UDvdW^{+}18, MTD18, SJV^{+}15]. Elastic-PPQ [MTD18]. elasticity [Bel16, BBT19, CBLS13, FJL^{+}16, HGG^{+}14, KC14, LSB^{+}18, MAB^{+}15, MCdA16, NLV^{+}19, dRRRR^{+}18, ZTL^{+}19]. elderly [AMR^{+}19]. eLearning [FJA^{+}18]. electric [MHW^{+}16, SLTK19, XZZ^{+}19]. electricity [MHW^{+}16, MD12]. electrocardiogram [BLL^{+}19, DNW^{+}19, MdMMNS^{+}19, OMPSPL^{+}19]. electrodes [WPS^{+}18]. electromyogram [NUPA19, PSI19]. electronic [BAB13, CLC^{+}19, DLL^{+}19, HPP^{+}18, YKÖ17, ZDL^{+}13]. Elements [CBS17, GBMP13, SMM^{+}14]. elephant [AS18b]. elevated [WCWC19, WCWC20]. Eley [YZI18]. elimination [ZHL^{+}18]. elite [LZWF19]. elliptic [AMN18, MCN^{+}18]. EMA [BRB^{+}19b]. EMA-RPL [BRB^{+}19b]. Embedded [ADH^{+}16, HYZS16, LSAM13, AMR18, CXZ^{+}19, CRRC18, EYY19, FLR^{+}16, QZM^{+}18, STC15, VRGR16, ZDM^{+}19, LLS^{+}14]. embedding [CZ12, FBM19, LYS^{+}19, LWSY18, NS19, SYAL13, YHH^{+}19, ZSFZ19]. emergence [ZSS^{+}18]. emergency [BDM^{+}19, GAW^{+}18, ODC19, RGSL18, SSK^{+}19]. Emergent [MVCC10, SHRE16]. Emerging [AWYJ16, HM19, KLH^{+}18, RR18, YGS16, CDMR19, ZGL^{+}18]. EMIH [GSY^{+}19]. EMIH-FPMIPv6 [GSY^{+}19]. emission [JCMPPC^{+}18]. EML [TDFZ18]. emotion [CPP^{+}18, CC19, DYC^{+}18, LRL^{+}14, YXY18]. emotional [ADH^{+}16]. emotional-based [ADH^{+}16]. Emotions [CRC^{+}19]. empathetic [HZM14]. emphasis [TAHS14]. Empirical [IKLL12, CT19c, GAB^{+}14, RMJ^{+}18]. employing [LRC^{+}18, RSY^{+}18]. empowered [ML11]. Emulating [ACPI19]. emulation [FPP^{+}18, SGN^{+}17]. enable [MPP13]. Enabled [HFT16, ZZLR18, AT19a, ATA19, AKB^{+}18a, ABF^{+}15a, ATM^{+}19, AMPZ16, BS11, DWJM18, FXG^{+}19, IASK14, LSZ^{+}18, LLF^{+}18b, LW18a, MID16, MPC^{+}18, MGN^{+}16, QLM^{+}18, QCY^{+}19, STB^{+}19, TOD17, TAS^{+}18, UKK^{+}19, VOS12, XLL^{+}19a, XXQ^{+}19, XAW^{+}10, YPJ19, GRL11]. Enabled/disabled [HFT16]. enablers [Bel16]. Enabling [AUSA19, BMH10, BP10, CDH^{+}19, EKGS14, GSGPP^{+}19, HGG^{+}14, IJCR19, JRJ^{+}11, KDHP16, KPM^{+}18, LCHW14, MGA^{+}19, RMHCMG15, TKK^{+}14, TCN^{+}16, WHBC19, WCKW10, YIA17, ZL13, ZLR^{+}15, AZH18, DFRW17, KPS18, LLAH13, PGCML^{+}19, SSLF^{+}10, WDR^{+}19, YCH19].

Enacting [ACC^{+}19b, CAB^{+}18].

Enactment

[HMS15, Bal16, JTBS15, dlFVPSHL^{+}14]. encoder [LCL^{+}18]. encoders [RSY^{+}18]. encoding [JLCC12, KP12, TA19, XWM18]. encrypted [AHM^{+}18, BGP^{+}17, CDL18, JLC18, KH18b, LXK^{+}14, NJB19, YXD18, YQZ^{+}19, ZFH^{+}18]. encryption [AMHJ10, BLL^{+}19, BGP^{+}17, Bro19, CLC^{+}19, CXWT19, DLZ16, DLLZ17, FAA^{+}18, FH13, HQZH14, HZL18a, HFT16, HYS17, HYS18, HYF18, JSMG18, KHMB13, LLL^{+}18, LDZW19, LL16, QRW^{+}18, Wan18a, WDZ19, WZCH17, XZP^{+}19, XTZ^{+}19, YWJ^{+}19, YCT15, ZXJ^{+}14, ZDW^{+}16, GMdFPLC17, LAL^{+}15, Sar18a]. End [BGP^{+}17, LOR^{+}18, MGN^{+}16, RHPV17, AHP16, ASAB^{+}18, CWSW14, DJZ^{+}15, GRZ^{+}19, GGC17, LLMP13, LASL16, MWL^{+}18b, PvSS17, PWP^{+}18, RAA^{+}18, RMDB18, Wan18b]. end-devices [Wan18b]. End-to-end [BGP^{+}17, LOR^{+}18, MGN^{+}16, RHPV17, AHP16, ASAB^{+}18, MWL^{+}18b, PvSS17, RMDB18]. end-user [LASL16, RAA^{+}18]. endoscopy [KHO^{+}19]. ENEAGRID [MDA^{+}19]. Energy [ARSMY19, ADA^{+}19, AS18a, AC18, BAB12, BBC^{+}12, BCDP12, BRB^{+}19b, CP17, DQLW15, DDB14, DMM14, DCMW17, GBF^{+}12, GCZ^{+}19, HCHH19, KKvdB^{+}17, KCS14, KV17, LBM18, LGL^{+}17, LYYY18, LSYC18, LG16b, MGT18, MMPF19, NDA^{+}19, OCCK14, PLP^{+}19, RGM^{+}19, SYJ^{+}19a, SCY^{+}18, SZW^{+}19, SLY^{+}19, TuIS^{+}19, THT12, URKM19, VEET18, WKC^{+}13, WWZC19, WWZ^{+}19, Wu16, WLHH18, XZL^{+}19, YZC^{+}19, YYD^{+}14, ZMTT16, ZZLH18, AK18b, AMBC19, BDP11a, BdM11, BMK^{+}14b, CJ14, CLP^{+}14, DKV14, DPBK16, DSBC19, DYI^{+}16, DNP14, FTK17, GDS18, GFW^{+}18, GAA19, HCMJ19, HFL^{+}19, HLL^{+}17, IPCA^{+}16, JLRS18, JFZL17, JEB18, KANS18, KSF^{+}13, KHG^{+}18, KMT14, KLP19, KKW^{+}14, LTC12, Len16, LN13, LPK17, LPK18, LLQS14, LOR^{+}18, LJGW18, Li18, LCZR12, LZXG12, LCMX16, LPL^{+}16, LIH^{+}19, LFY^{+}19, LYH^{+}19, LCY^{+}19b, MNV12, MPF^{+}16, Mat18, MZD^{+}16, MFSV19, NSSA^{+}14, NQQL13, PSPP16, PLLA18, PTD^{+}18]. energy [PPMM^{+}18, QCY^{+}19, QMSG12, QCD16, Qur19, RYH^{+}19, RCTY19, SHP^{+}16, SB14, SMS14a, SSSJ19a, SLD^{+}15, SK19, SG15, TSD18, TDSH16, TPBS14, TSGVRGS19, TKA18b, VF18, VVC^{+}12, VTTK17, VGC^{+}13, WWC14, WLH16, WCC14, XZZ^{+}14, XDH^{+}17, XDHL12, YPLZ17, ZG19, ZAA^{+}14, ZXD^{+}19, ZAC^{+}18, ZZZ17, dACAM13, vKLA^{+}19, DYY^{+}19]. Energy-aware [ARSMY19, ADA^{+}19, BAB12, BCDP12, DCMW17, LYYY18, MMPF19, WKC^{+}13, ZMTT16, ZZLH18, BMK^{+}14b, HCMJ19, JEB18, KANS18, KSF^{+}13, KCS14, LTC12, LPK17, LPK18, NSSA^{+}14, ZAC^{+}18, dACAM13]. energy-balancing [LYH^{+}19]. energy-conserving [NQQL13]. Energy-credit [KCS14]. energy-driven [ZAA^{+}14]. Energy-efficiency [DDB14, SLD^{+}15, vKLA^{+}19]. Energy-Efficient [SZW^{+}19, YZC^{+}19, CP17, GBF^{+}12, LBM18, LSYC18, PLP^{+}19, SLY^{+}19, Wu16, WLHH18, XZL^{+}19, DNP14, JFZL17, KLP19, LLQS14, LZXG12, LCMX16, LFY^{+}19, PPMM^{+}18, Qur19, RCTY19, SB14, SSSJ19a, SK19, VTTK17, WCC14, XZZ^{+}14, YPLZ17, ZXD^{+}19]. energy-incentivized [BDP11a]. energy-saving [CJ14, QCD16]. Enforcement [SME^{+}21, EKSDN19, Hua10, MG14, MG16, SME^{+}19]. Enforcing [TTH15]. engagement [BDWM17, FTK17, Kim18]. Engine [BBCN18, Bal16, DLH^{+}17, FPL^{+}19, KGT15, RLRC13, SSSJ19b, XLL18b, PS13]. Engineering [BLO^{+}18, BZS18, GHD19, AC16, AAJ17, BAP17a, BAP17b, SK18, dSSCdL19, Zhu14, ZWMC19].

Engineering-DiffServ [BLO^{+}18]. Engines [HMS15, CMP^{+}17, GRCP^{+}17, XLZ^{+}14]. enhance [HCL^{+}17, LHX^{+}18, MK19b, NNRA19, RM16]. Enhanced [HLL12, SMS14a, AMN18, ArMS19, AM19a, BCC^{+}17, CT19a, CSP13, DMPS19, HLW12, KID^{+}16, KKKM17, KKKM18, NV11, RTS^{+}16, RGVGGSSZ14, SLW11, YZN^{+}15, ZZZC19]. enhancement [BMU18, CHS11, DZZ^{+}15, DGA18, LCL^{+}19, MYBMM18, SVN10b, WYJ^{+}19, ZLT^{+}19]. enhancements [PSJ^{+}12]. Enhancing [AMR18, BAKB19, BHE^{+}19, CWL^{+}18, IPG^{+}18, KX11, MZL^{+}19, TCH19, YMW13, ZCLW18, AGBR19, HYG^{+}19, SMPC12]. enough [BBD^{+}13]. enriched [LRJG19]. Enriching [KBdLG18, GTM19]. enrollment [DEL19]. Ensemble [SB18, BMP^{+}16, CMT16, CTU19, LPK17, LPK18, LYS^{+}19, PPLL17, RGAT18, WLW^{+}18]. ensembles [IdAP19, MJDN15]. ensure [CBT^{+}19]. Ensuring [MROD10, SYQ^{+}19]. Entangled [ADDV16]. Enterprise [ECA^{+}18, SSA^{+}19, dVXB^{+}11]. enterprises [YLL^{+}19]. entities [LWZ^{+}19a, NS19, XLZ^{+}14]. entity [BZHV19, LXL^{+}17, WLLF16]. entropy [RDSA18, RGGH18, WLZ^{+}14]. entropy-based [RGGH18]. envelopment [KPS18]. Environment [PMDS18, YG18, ANA16, ABB^{+}19a, AJY12, AHYF19, AKB^{+}18a, AAD^{+}13, AK18b, BAJ^{+}19, BJ12, Bar11, BKS^{+}18, BDZ13, BBMG10, CCL11, CMI^{+}19, DJZ^{+}15, DL19, DW11, DLMS15, DR18, DCC13, DGdL15, EK11, EHT10, EA17, FMSSM12, FHZW18, FS19, GCK18, GAI^{+}18, HKP10, JPB17, JLI^{+}13, JOSD19, KVK^{+}18, KPA17, KKP19, KTTK17, KCCL18, LLC11, LJLW13, Li18, LTTL19, LJY10, MRT^{+}19, MSBA16, MVG^{+}14, MRS18b, NRV^{+}17, NHH^{+}19, PP10, PVN^{+}12, POJ^{+}16, QZD^{+}18, RGAT18, RYH^{+}19, RPA^{+}18, RGVGGSSZ14, RM11, SV16, SLTK19, SDWS13, STA17a, SSI19, SGB^{+}18, SYQ^{+}19, TCC18, TZLL18, TMB^{+}19, VSP^{+}14, VKT^{+}19, WCHL10, WY19, XZW^{+}19, XWjZyF19, YLN15, ZZJY16, ZCZ^{+}18, ZZL^{+}10, dFBP^{+}17, BCF^{+}10, LWW^{+}13]. Environmental [YCD^{+}19, BBWB^{+}18, GRL11, PSW^{+}14, RHKC15, VDK12, WTP^{+}13]. Environments [ACC^{+}19c, BB17, CCMGF18, CDH^{+}19, TF17, ABC^{+}18, ABS11, ADA^{+}19, ALK15, AMT^{+}12, AFB^{+}10, ACK^{+}15, BML18, BOWD^{+}19, BdM11, BBC^{+}12, BN17, BBB16, BDP11b, BDH14, BCG^{+}19, CPDJ13, CRC13, CFVP12, CFG^{+}19, CBC^{+}19, CFPC17, CCCT14, DVD12, DEG^{+}17, EMM12, FDGR14, FdAGdAFV19, Frî14, FSM^{+}18b, GL19, GKW^{+}12, GCTLA^{+}19, GTSP^{+}19, GKTK15, HFM19, HB19, HCB16, HPP^{+}18, HHW11, JRJ^{+}11, JTB13, JTS13, KKB14, KWK16, KH18a, LTMW19, LLS^{+}19, LZCX19, LC15, MCT^{+}15, MOBD18, MROD10, MMPF19, MMRL17, NNRA19, NRR^{+}15, NAD^{+}18, NJ16, NOF18, PBV^{+}13, PLA18, PECA19, PGTBC18, RPH19, RBN13, RMHMG17, RRH16, SB14, SPdSR^{+}17, SBAD^{+}18, Sip12, SCN^{+}14, SL19, TDFZ18, TCBPR16, TMP15, TAKV12, UMUB19, VPP^{+}19, VCL^{+}19, VD16, WHW16, WSZC18, WHBC19, WSC^{+}19, XTT18, ZGZ^{+}10]. Envy [YLJL18]. Envy-free [YLJL18]. ePASS [SCZ^{+}14]. Ephemeral [CCMGF18, CLCMG^{+}18, MOBD18]. epidemic [GRTV10, XL19]. epidemics [OCCK14]. episode [AMKM18]. episodes [DDD^{+}19]. equal [DMM14, ZXW19]. equal-energy [DMM14]. equality [WZCH17]. equation [PCC18]. equilibrium [GPVN19]. Equivalent [WZE19]. era [QCX18, SDDG17]. erasable [LY18a, YL16]. erasure [GIM16, HDO16, PWMX17, XWM18]. erasure-coded [PWMX17, XWM18]. Erlang [Tur18]. Erlang-based [Tur18]. erosion [YTQ19, YTQ20]. Erratum [DP20, DP21a]. error [CCDP19, GHMX10, KDHP16, LRMS19, WWSL19]. errors [WZE19]. escalation [XWRZ19]. eScience [BCJT13, FA11a, HGM15, KZ14, KMZJ16, MK16a, MK16b]. eSciGrid [SAGL10]. ESLEA [SHJS^{+}10]. ESNET [ZWDP18]. ESPM [LJY10]. Essence [PKP19]. Establishing [BFN18, SZK18, TAB^{+}18, HPP^{+}18]. Establishment [AMAY19, CRRC18, PGCML^{+}19, XZ14b, ZXW^{+}18]. Estimating [SHH^{+}19]. Estimation [AdVAGF18, KSSG16, CBK^{+}17, HJA^{+}19, JOPW14, KS18a, KXS^{+}16, KMK^{+}14, LSL^{+}15, LGZY18, NS17b, PS10, PcFP^{+}17, RAKJ18, RMDB18, SHP^{+}16, TMDZ15, WOPW13, YZWG18, YSZW18, ZWZ19, WSZH18]. estimations [IDM^{+}16]. Ethernet [ZWJ^{+}19a]. ethnic [HX19]. EUBrazilCloudConnect [FEB^{+}19]. European [CPGBC16]. EV [HZLH19]. evacuation [LG18]. Evaluate [YTQ19, LCGPC19, YTQ20]. evaluated [BRXdS11]. Evaluating [ABTA18, HA18, HH19, Len16, LLS^{+}19, SC19, XWjZyF19, ABMC18, ALFR16, BBT19, OA17, RMJ^{+}18, VK17]. Evaluation [CGN18, CYJ19, CPP16, FK12, GLJ19, KVvE18, MG19, SS17, YDNV16, AB19a, ABGMC19, AB21, AMAY19, AZO^{+}19, AEM10, BGI14, BARMB14, BDZ13, BMU18, àCKPM19, CCRV13, CHS^{+}18, GS13, GEAR13, GOLL19, JAAD^{+}16, JWW14, JFZL17, LSD^{+}17, LW19, LZL^{+}19a, MZC10, RPMG10, SHH^{+}19, TDLC17, VVB11, VRS^{+}19, WWD^{+}14, ZMP10, ZWDP18, ZXZL18, dSK^{+}19, dOOO^{+}13]. evaluations [DQC^{+}19, SPR^{+}10]. evaporative [BMZ10]. Event [HCB16, LJ17a, MLW^{+}18b, ML17, NJH^{+}18, ZSP17, AHYF19, BKB11, BMP^{+}16, FLT17, JSZ^{+}19, KNI^{+}18, KAP19, KZCW13, LLG^{+}16, LWXY19, LCCM18, MWQ^{+}14, NJ17, NJ19, PQBP17, PKB19, TCN^{+}16, VETT16, WHMO13]. Event-Based [ML17, LWXY19, PQBP17]. Event-driven [NJH^{+}18, VETT16]. events [AMPS19, DdM10, KBdLG18, KAS^{+}18, LC13, NHH^{+}19, XWL^{+}15]. Evidence [AJ19, CPE^{+}17, WW11, WQZ19, ZGZ^{+}10, ZDL^{+}19]. evidence-based [WW11]. evidences [ACMM19]. eVLBI [SHJS^{+}10]. Evolution [CSV^{+}12, DDD^{+}19, JGB19, DFLO17, GVBG17, HZL18b, HCX^{+}19, KKS18b, KS18c, LWZ18, LO19, RGVGGSSZ14, TM19, WZL18]. Evolution-based [DDD^{+}19]. Evolutionary [KOT18, LFL^{+}17, RCMT18, dCTVC18, GKTK15, GLJ19, JGFB18, LWYS18, MDB^{+}18a, MCSA18, NF13, PKB19, SG13, TdPF^{+}17]. evolutive [DLDTGMMP16]. evolvement [Kyr19]. Evolving [DSCJ18, LWXY19, LRBW17, SISGS18, KZS^{+}19, NAD^{+}18, SLB^{+}17, THA^{+}17]. exact [CSdCM^{+}17]. examine [RDSA18]. Examining [NZOCJ^{+}19]. exams [CND^{+}19]. Exascale [DBD^{+}14, Eng14, FBS18, DYI^{+}16]. exceptions [GGLD10]. exchange [AHL11, OPT^{+}17, TLSC17, WWW^{+}16, ZA13, ZXWA18, KGdL11, LSH^{+}11]. exchanges [CKP^{+}19]. exclusion [Bag19]. exclusive [CXC^{+}18, FLT17, WDZ19]. executables [AD18]. executing [CCL11, KTTK17, TKA^{+}18a, TVB18]. Execution [CBK^{+}17, ANG^{+}19, ATF11, AEK^{+}18, BBI13, CLRL17, CLC11, CsZzG^{+}13, DJZ^{+}15, DLS14, DGdL15, EHT10, HMM18, KS18a, KHG13, KC14, KuRAk^{+}18, KN10, KDG^{+}19, KGT15, LAL^{+}14, LWS^{+}12, MMRL17, NF13, PdAF12, PAB^{+}14, PKB19, QCYJ17, RBC^{+}15, SV16, SHP^{+}16, SPdSR^{+}17, SLC^{+}17, SSSJ19b, TZST14, TGM11, dOWdAS^{+}18, ZZZC19, ZZL^{+}10]. executions [CRVZ15, CdSDS15, HHD^{+}12]. EXEHDA [dSMAdR^{+}17]. exhibits [VPA^{+}18]. exhiSTORY [VPA^{+}18]. existing [BCP18, DGY^{+}18, MFG^{+}14]. exome [CMX^{+}16]. Expanded [Sha16]. Expanding [CGCB^{+}12, SZG^{+}19]. expansion [DZLA19, LSCL19, MLL15]. Experience [HXL^{+}18, SHS^{+}19, AKB18b, GTCZG^{+}18, KKP19, QLM^{+}18, SSST17, SHH^{+}19, TBB^{+}17, XFM16, dSGD19]. experiment [BRNR15, JCA^{+}19, MdOO^{+}17]. Experimental [LWW^{+}18, CCT13, CPMG19, DCBF19, IAM^{+}18, MLC^{+}11]. experimentation [CPGBC16, GTSAR^{+}14]. experimentations [VRGR16]. experiments [GCV^{+}14, MMVS19, PBC^{+}16, PBC^{+}17, dOWdAS^{+}18]. Expert [CAS^{+}18, DBS14, WGM15]. expertise [DKFKF18, KZA^{+}18, SG17]. explain [NZOCJ^{+}19]. explicit [MWPVB12, TSOB15, XLZ^{+}14]. exploit [EDH^{+}13, GGC17]. exploitation [SK12]. Exploiting [AM17, BCF16, CGIP14, CSL17, CLP^{+}14, MPP13, PQBP17, RGN^{+}18, WC14, CCJ16, CPSRG14, HCC^{+}14, HDO16, HMZ18, KAS^{+}18, VVC^{+}12, XFJ^{+}19]. Exploration [BBD^{+}19, BMP^{+}16, DBD^{+}14, RTHB17, RLRC13, SGRT19]. Exploratory [CFM17]. explore [dSMAdR^{+}17]. Exploring [BFP18, DRZ^{+}19, HBN^{+}13, LWF^{+}17, MKS18, PAL^{+}19, SSW^{+}19, YZW^{+}18, LCY19a, SNC18]. exponential [WSL^{+}19]. exponentiation [WSQ^{+}16]. Exposing [CGL15]. exposure [WG13, XYML19]. expressive [SCZ^{+}14]. extend [MK19b]. extended [SCL14, ZGZ^{+}10, ZWWL18]. Extending [BDP11b, LWT18, RGDML16, TKTG19, KSC^{+}19]. Extensible [MEW^{+}19, SB19b, LKN^{+}13]. Extension [GSY^{+}19, VKK14, ZXW19, ZXJ^{+}14, ZWL^{+}16]. extensions [TMJH19]. External [LYZC15, BAKB19, CN17, LQK^{+}16, MFT^{+}17, NHH^{+}19]. extraction [ASY^{+}18, Cha11, GWC^{+}16, HZW19, HWWT12, HAAR^{+}19, MdMMNS^{+}19, PRW14]. Extractive [CZ19]. extrema [PSS^{+}18]. Extreme [AMPS19, FRB^{+}14, KNI^{+}18, LJW^{+}19b, WTG^{+}19, dSFP^{+}17, TKRA14]. extreme-scale [FRB^{+}14, dSFP^{+}17]. eye [KCCL18, MKS18]. Eyetracking [PKA19]. Eyetracking-based [PKA19].

F [KK19]. F-DAD [KK19]. F2C [SMBMT^{+}18]. Face [GPV^{+}14, CZ12, EU19, HWWT12, HNQ^{+}18, HLL12, LTJK12, LLSL18, PSS^{+}18, YPCK12]. FaceDCAPTCHA [GPV^{+}14]. faceted [XZ14a]. facial [EU19]. facilitate [GVdBdL15, XLL18b]. facilitating [NAD^{+}18]. facing [AAC^{+}19]. Factor [AMSPL19, ED19, HDH^{+}18, LWD^{+}14, LNK^{+}18, LW19, LZWF19, RSJ^{+}14, WLS^{+}18]. factorization [GGMS18, WTS14, WYL^{+}18, ZCH^{+}17]. factors [CN17, LCGPC19, WGM15, CsZW14]. factory [AJR^{+}19, HMW^{+}19, HMW^{+}19]. facts [CZ14]. failover [AS19b]. Failure [SSSJ19a, BMR15, DCF19, DNW^{+}19, LPMY18, LJGW18, MSI^{+}12, NLLC19, RRU^{+}18, YIA17]. Failure-aware [SSSJ19a, YIA17]. failures [BDNP13, GJY18, HA19, JS13, RAdARP19]. Fair [CA15b, HCW^{+}18, JJH19, WWW^{+}16, YZ12, GSGPP^{+}19, WXYL15]. Fair-Play [WXYL15]. fairshare [ÖEE13]. false [LY17]. families [ZXM^{+}19]. family [GGS13, MCG^{+}15, SLW11]. FAMOUS [KKA18]. farming [KK16]. farms [MD12]. farthest [FJ18]. Fast [BLAN^{+}16, DSK^{+}14, GGN17, GEAR13, LGP^{+}19, MH19, SSK^{+}19, WLP10, XKBA18, ZFY18, CHWW13, DST14, FLT17, HIA^{+}18c, JFZL17, KHMB13, LSZ^{+}16, LPY^{+}18, LLL^{+}19, LC15, LZYC13, NS17b, RRU^{+}18, WWD^{+}14, YIA17, ZA14, MDA^{+}19].

FastDesk [SWW^{+}18]. fat [WWQ^{+}18]. fat-tree [WWQ^{+}18]. Fault [DK14, GCV^{+}14, LYW^{+}16, SE19, SAPA17, ABF^{+}15b, AFB^{+}10, AGKZ18, AMR18, CLZ18, CXZ^{+}19, CY12, CCL11, DZZ^{+}15, GdVC10, LCBF13, PIP18a, RWZ^{+}19, SPR^{+}10, WWSL19, WDR^{+}19, WHCW19]. fault-prone [AGKZ18]. fault-tolerance [WDR^{+}19]. Fault-tolerant [DK14, GCV^{+}14, LYW^{+}16, AFB^{+}10, CXZ^{+}19, DZZ^{+}15, LCBF13, RWZ^{+}19, SPR^{+}10, WHCW19]. FBAC [YWJ^{+}19]. FCA [BCdV^{+}19]. FCA-based [BCdV^{+}19]. FCM [GGN17]. FDE [KKB^{+}19]. fear [MGA^{+}18]. Feasibility [OMPSPL^{+}19]. FEATHERS [BAKB19]. Feature [MWQ^{+}19, SLD^{+}18, AUSA19, AAM^{+}19, CZY^{+}19, Cha11, CMI^{+}19, CWUS19, DH16, FTK^{+}14, GLC19, GNGG17, HZW19, HLZ^{+}19, HX19, JLS19, JLQZ18, KP18, LZL^{+}12, LHY^{+}19, MBC^{+}11, MdMMNS^{+}19, MZYA19, MDD15, MGA^{+}18, MBA19, NK18, PCC18, PSS^{+}18, QCZH19, dSSCdL19]. Feature-based [MWQ^{+}19, LHY^{+}19]. Features [NS17b, AMBB18, AKP^{+}18, AGA18, BLMU19, CPP^{+}18, CAL^{+}18, FNA12, GAFFOG12, KHO^{+}19, LYC^{+}19, LYXT14, LZL19b, SHL^{+}19a, WLZ^{+}14, XPL19, YXY18, YGY^{+}19]. FEC [EBCP18]. Federate [BCF^{+}10]. Federated [BR10, DRZ^{+}19, APAZ17, AK14, CFVP12, dCCDFdO15, CGJ^{+}10, GCV^{+}14, HFM19, PP10, PBV^{+}13, PFRC16, PPL^{+}15, SLB^{+}17, SLZ^{+}18, ZZDM^{+}18, EH10, Lea13, Lea15]. Federation [FMN^{+}17, VHML10, COC10, KBB^{+}16, TOS18, YNSM12, CCM^{+}14]. federations [Erd13, FRZ19, MLM16, MPR^{+}16, PVGD^{+}19, RMHMG17]. Feedback [GSLI12, CWJ^{+}18a, PQBP17, ZXD^{+}19]. Feedback-based [GSLI12]. feeding [KO11]. FENNEL [SLG^{+}17]. Fermat [WTS14]. fetch [GTMZ17]. fetching [SR12]. FGCS [CRW^{+}16, GVTdL18, PC18b, SZV19]. FHE [MS19]. FHE-based [MS19]. Fi [AKM18, DCBF19, KWB19, SLS10]. fiber [WXZ^{+}18a]. fidelity [LBD^{+}19]. field [CGIP14, HPP^{+}18, KS17a, LZH^{+}18, LLU^{+}18, SMC18, TSTL16]. fight [WZH^{+}18]. File [AHEM17, DLXR14, GTMZ17, KO11, Li15, MLG13, MFL18, MM18, MMLO18, SYQ^{+}19, USK16, UDvdW^{+}18, WLP10, WZS^{+}18, YCY10, YSC^{+}15, ZSX^{+}15, MBC^{+}11]. files [LLF^{+}18a]. filesystem [ZYZ^{+}18]. filter [GCTLA^{+}19, HAA^{+}16, HIA^{+}18c, JYZ^{+}18, LLW^{+}19a, NUPA19, SGB^{+}18, WLL^{+}19a, WC14]. filter-bank [NUPA19]. filter-wrapper [HIA^{+}18c]. Filtering [KMC18, LZY^{+}19b, CLAL19, DCC13, EAA16, GCD^{+}18, GDAS18, KKB18, LCL^{+}16, LCW^{+}18, LQLX10, MG19, MVG18, PQBP17, RW18, TQL^{+}19, ZRZ^{+}14]. Filters [ML17, LLYW19, ZL13]. FIM [HDH^{+}18]. financial [AMI16, KK14, PKP19, QGT^{+}18, WGC19]. find [BGRBA19]. Finding [BBH18, CSdCM^{+}17, LXM^{+}18, LG18]. Fine [SJ18, AMBB18, CZZ^{+}18, JLY^{+}18, LCL^{+}18, LZLL18b, LZHY19, MLW^{+}18a, VVB11, WWH^{+}19, WC14, XTL^{+}19, XYML19, YAX^{+}18, ZMP10, ZSL^{+}19b]. fine-grain [WC14]. fine-grained [AMBB18, CZZ^{+}18, JLY^{+}18, LZLL18b, LZHY19, MLW^{+}18a, VVB11, WWH^{+}19, XTL^{+}19, XYML19, YAX^{+}18, ZMP10, ZSL^{+}19b]. fine-tuning [LCL^{+}18]. finer [ZFS^{+}18]. finer-grained [ZFS^{+}18]. finger [LWF^{+}17]. fingerprint [GAFFOG12, LZL^{+}12]. fingerprinting [KPB18, RS17b]. Finite [XWW19]. Finite-State-Machine [XWW19]. FinTech [LLW^{+}19a, MZL^{+}19, NAAC19]. FIPIP [JLY^{+}18]. firefly [LZWF19]. firewall [HCNT14]. first [Che13a, GTCZG^{+}18]. Fisher [CLY14, HLL12]. FitCNN [LYL^{+}19]. Fitness [JGB19, CZL^{+}18b]. fitting [CLC11, MK19a]. fixed [LZL^{+}12]. fixed-length [LZL^{+}12]. fixture [LSW^{+}19]. Flame [VETT16]. Flame-MR [VETT16]. flash [DLZ^{+}14, LPK17, LPK18, SCH^{+}19]. flash-based [SCH^{+}19]. flash-flood [LPK17, LPK18]. Flat [OCCK14]. flexibility [CAB^{+}18]. Flexible [CRC13, LCBF13, MWL^{+}18b, YXZG18a, AHP16, BZMY10, Bar11, DKJ19, FTD17, HLL^{+}11, IG12, LTC^{+}19, NOF18, RPH19, SGS^{+}18, TMP15, WCL^{+}17a, WDZ19, YCXW18, ZAB15, ZFH^{+}18]. Flight [JCA^{+}19]. Floating [MDM^{+}19, KXS^{+}16]. flood [LPK17, LPK18]. flow [ABMESM18, CT19a, CT19b, CAL^{+}18, FM10b, GS15, GXL^{+}18, KLJS19, LGP^{+}19, LPS19, LRC^{+}18, TMJH19, XSMS15, XDH^{+}17]. flower [GHEB^{+}18, GHEB^{+}23, GJ18]. flows [CEP19a, HAF^{+}16, KDHP16, KC19b, KGT15, WBJM14]. Flubber [JLI^{+}13]. fluid [GPS^{+}17, WJLW18]. Flux [DDJ^{+}13]. fly [BTM10, HPZL18]. flying [SNP19]. focal [AHD^{+}19]. focused [CY12]. Focusing [HPP^{+}18]. foetal [LQF19]. Fog [FSV^{+}19, MMC^{+}18, NWT19, PDH18, PGTBC18, SM18, VRS^{+}19, WZH^{+}18, AZH18, AkBAL^{+}19, CLCY18, CdRRdCB19, FXG^{+}19, FFC^{+}18, GLJ19, HDA^{+}19, HCW^{+}18, HXC^{+}18, IFD^{+}19, JSMG18, KGS^{+}19, KH18b, LLYW19, LDY^{+}18, LLZ^{+}18b, LYH^{+}19, MWL^{+}18b, MGA^{+}19, NDA^{+}19, NLS19, PRL^{+}19, PSY^{+}19, QRW^{+}18, RGN^{+}18, RR18, TuIS^{+}19, TGM^{+}19b, VCL^{+}19, VKT^{+}19, WWDF18, Wan18a, WDKV19, WPS^{+}18, WSY^{+}19, YSHM19, YAX^{+}18, ZCX^{+}18, ZCL^{+}18, ZZF18, ZSL^{+}19a, ZSW^{+}18b, AkBAL^{+}19, KK19, RLM18, SMBMT^{+}18, YXA^{+}18]. Fog-2-Fog [AkBAL^{+}19]. Fog-assisted [VRS^{+}19, NDA^{+}19]. Fog-based [WZH^{+}18, WWDF18, KK19]. Fog-cloud [SM18, LDY^{+}18]. fog-computing-based [PSY^{+}19]. fog-driven [FFC^{+}18]. fog-enabled [FXG^{+}19]. Fog-to-Cloud [SMBMT^{+}18]. fog/edge [RR18]. foggy [CLCY18]. fold [AHMS18]. folding [SSLF^{+}10]. followee [CCJ16]. following [MYW^{+}19]. foot [XL19]. footprint [BEWZ10, DYI^{+}16, LYW^{+}18b, VVC^{+}12, KK11]. footprints [WNR19]. foraging [LLES19, RC13, VR12]. Force [VLAC^{+}13, FZT^{+}18]. forecast [LCL^{+}18]. Forecasting [PLA18, CLRL18, Che14, HZL18b, JH16, KSSG16, LPK17, LPK18, LLW^{+}18a, Lin18, PGTBC18, SG15, YGY^{+}19]. Forensic [SK18, ACMM19, DC18b, KMST19, QC13, QC18, VOCHC17, XFTZ16, KKA18]. Forensics [JL14, AAI^{+}19, ASA19, CDFW18, LCHW14, PD11, YHA^{+}19]. forest [BYL^{+}18, IdAP19, JNHL18]. forest-based [IdAP19]. Fork [LTC^{+}19]. Fork-free [LTC^{+}19]. form [GTSAR^{+}14, HXY13]. Formal [PLCGS11, AMMC18, MKM11, VV16, XYLZ18]. formalism [BGI14]. formalisms [OCW14]. Formalization [WZ16, ZW10]. formally [HKA^{+}18]. format [AHL11, ZWL^{+}16]. formation [DBS14, FMRS18]. forming [FZT^{+}18]. forms [YCZJ18]. formula [RS16]. fortification [KPS18]. forward [KDHP16, KJ18, LAQ^{+}19, NJB19, KS17a]. Forwarding [KIMR15, FFL^{+}19, LGP^{+}19, MDB^{+}18b, MVL18b, MRS18b, RWO^{+}19, WZWW18, ZCW11, ZCDV19]. FOSS [JMAG19]. found [SSLF^{+}10]. foundation [XLL^{+}19c]. four [SYCH18, CsZW14]. four-band [SYCH18]. Four-factors [CsZW14]. Fourier [LZL19b]. foxtail [DJJ^{+}18]. FPGA [ABF^{+}15a, SAVS19]. FPGA-based [ABF^{+}15a]. FPGAs [CJPC19]. FPM [CsZW14]. FPMIPv6 [GSY^{+}19, GSY^{+}19]. fractal [IASK14]. fraction [WTS14]. Fractional [BWR12, BW13]. fractures [RBGA18, WWP19, WWP20]. fragmentation [MSE19]. Fragmented [ML17, CCS^{+}10, FdAGdAFV19]. Fragmented-Iterated [ML17]. frame [JLY^{+}18, LYS12]. Framework [ACC^{+}19a, ABTA18, GTEL^{+}18, KV17, MK19a, MPI^{+}18, MWMA10, SME^{+}21, SDK19, WWVJ17, YAA^{+}19, ABMM18, ABMC18, ABMMC18, AB19a, ABGMC19, ABM19, AB21, ABMMC22, AAF18, AK19, ASAB^{+}18, AT18a, AASI17, AB18b, ASA19, BAKB19, BBT19, BGS^{+}19, BOP^{+}14, BDH14, CCL19, CVT19, CPGdS^{+}13, CMEA^{+}19, ÇBCA15, CKR16, CCS^{+}10, CFPC17, CCDP19, CGM^{+}18, DMPP16, DPS16, DLL^{+}19, EAA16, FTA^{+}14, FHYH15, FJA^{+}18, FPGK18, GVB13, GGC17, GAMC19, GVdBdL15, GMP^{+}17, GIK18, GBKJ18, HAP15, HHK18, HEES19, HMA^{+}18a, HLL^{+}11, HXC^{+}18, HAA^{+}16, HIA^{+}18c, HPL^{+}19, IFD^{+}19, ISS^{+}15, JSZ^{+}19, KSF^{+}13, KC14, KSS19, KH19, KPA17, KuRAk^{+}18, KIJ^{+}19, KAS^{+}18, KGLY18, LQK^{+}16, LKJ^{+}19, LLF^{+}18b, LvW14, LLM^{+}16, LSL^{+}18, LYL^{+}19, LKK^{+}16, LSJ^{+}14, LQF19, LWZ^{+}19b, MML^{+}18, MNV12, MID16, ML11, MMLO18, NK15, NJ18, NJ16, NOF18, NAAC19]. framework [NAM^{+}19, Osm19, ÖE13, PBV^{+}13, PLCGS11, PSW^{+}19, PMT10, PcFP^{+}17, Qur19, RBA17, RSRA18, RAdARP19, SME^{+}19, SRZD15, SMPC12, SYJ^{+}19a, SA14, SBAD^{+}18, SCLC19, SMSF18, SAK^{+}10, SG14, SDC11, SYW17, SGS^{+}18, SBK18, SYAL13, SLB^{+}17, SDH^{+}19, SJSA19, TZL^{+}18, TY11, VK17, WLLF16, WMQ^{+}16, dOWdAS^{+}18, WLA18b, WG13, WHYZ17, WHYZ18, XFLL16, XWRZ19, YMLT13, YSC^{+}19, YPHZ14, YCZJ18, ZSX^{+}15, ZCK^{+}15, ZYTC15, ZMN19, ZZQ^{+}13]. frameworks [BDP11a, DJ13, EET18, HZZ^{+}14, IAM^{+}18, MGT18, VEET18]. fraud [AAC^{+}19, CLS^{+}19b, SC19, ZZH^{+}18]. FRDT [KK11]. Free [KTB18, CWL^{+}18, CWW^{+}16, ED19, FZW^{+}18, GWW^{+}19, LTC12, LZLL18a, LGP^{+}19, LTC^{+}19, LL16, PWA^{+}19, SMM^{+}14, TJ18, WSU^{+}10, YLJL18]. free-riding [WSU^{+}10]. frequency [AQB15, CJHH13, IPCA^{+}16, KANS18, KWB19]. frequent [CEP19b, ÇÖ13, CLM14b, IHA18, KZCW13, LY17, LC15, LHW^{+}18]. fresh [GWC^{+}16]. freshness [JZJ^{+}18]. friend [MML^{+}18]. friendly [ABB^{+}19a, KuRAk^{+}18, KLW^{+}16, WHZ19]. frog [LCW^{+}18]. front [GRZ^{+}19, LLMP13]. front-end [GRZ^{+}19, LLMP13]. Frontier [KPG19, GML^{+}13, TZST14]. FRP [RRU^{+}18]. fruit [HPZL18]. fruits [HMA18b]. FTL [PLZX19]. Full [JZJ^{+}18, BRL19, JK17, SCJ^{+}19a, WGX^{+}19, ZLL17a]. fully [CCL11, HKS18, TY11, ZXJ^{+}14]. Function [YWJ^{+}19, MCJ19, SLL^{+}18, SLY^{+}19, WLXZ18, ZSL^{+}19a, ZZBZ19]. Function-based [YWJ^{+}19]. functional [BBH18, BRHH18, GSV^{+}10, YAO14]. functionality [KID^{+}16, MK19b]. functions [BBH18, CNR19, FJKK17, HH19, KTB18, YCL^{+}19]. fundamental [AGKZ18, EHT10, YH18]. fungible [ABP18]. fungus [Jun18]. FuSeR [SMZ^{+}16]. Fusing [GJ18, LYXT14]. Fusion [SCG^{+}18, AAM^{+}19, CMZ^{+}18, DGA18, FTK^{+}14, GHEB^{+}18, GHEB^{+}23, GCCL18, sGbKS19, HLZ^{+}19, Ima19, JYZ^{+}18, KCCL18, LSL^{+}15, LWT18, LYW^{+}18b, LFY^{+}19, LZY^{+}16, PCC18, TCC18, YLVY15, YZG^{+}18, dFPFG19]. Future [AB19a, AB21, ABMMC22, ADALZ14, BFS^{+}17a, Bo20, ChK11, Cha14b, CDFZ16, DP20, DP21a, DP21b, DO15, GHEB^{+}23, HYZS16, HZX^{+}20, HYS18, JLC^{+}20, LBJ^{+}24, NDZ^{+}18a, NDZ^{+}19, SME^{+}21, WWP20, WCWC20, YWG^{+}20, YTQ20, ZMZ^{+}20, wZcZN^{+}20, AR18, AS19a, ACSdRR17, AGMT17, BS17, CLCY18, GBMP13, KARP14, LLMP13, MYBMM18, PTD^{+}18, PWA^{+}19, QKC19, TSS^{+}19, URC19, URKM19, Zhu14, dCTVC18]. fuzzifying [YH18]. fuzziness [FLR^{+}16]. fuzziness-embedded [FLR^{+}16]. Fuzzy [AS18b, HZL18b, IS18, KLV^{+}18, NDZ^{+}18a, NDZ^{+}18b, NDZ^{+}19, RRKA19, SMZ^{+}16, XL19, Alp18, Bag16, Bu18, CPE^{+}17, Che14, CAL^{+}18, Che18, DNW^{+}19, GBY16, KHMB13, LSZ^{+}16, LZL^{+}19a, LAH10, LWYS18, MLGGB^{+}17, OMD^{+}18, RAKJ18, SYJ^{+}19b, YAO14, ZZS^{+}19, ZDW^{+}18]. FvRS [XTF^{+}19].

G [AT19a, AVPV17, DPK^{+}19, FNA12, NLS19, WTR^{+}13, YYS^{+}19]. G-enabled [AT19a]. G-Hadoop [WTR^{+}13]. G-SPAMINE [AVPV17]. GA [HMH17]. gain [JCL^{+}19]. Gait [AAN^{+}18, LZL19b, WCB^{+}18]. Gait-based [WCB^{+}18]. gallstones [YWLL19]. Game [FEPC18, IS18, KIMR15, LKCS18, YLWW18, ZCDV19, AY16, DPL14, FAL^{+}19, HS19, HSBE19, JLQ^{+}17, JLCC12, JCL^{+}19, KMC18, KK14, LJ17b, LN13, LY19, LWYS18, SSHC19, SSJ19, SBK18, TLSC17, WWRS16, YMY^{+}17]. game-based [FAL^{+}19, HSBE19]. game-theoretic [JLQ^{+}17, JLCC12, KK14, TLSC17]. games [AW19, JLQ18, LCC19, YC13]. GAMESH [BCC^{+}17]. GAN [TA19]. Gang [PK11]. gap [Bha18, NJKF18]. gaps [KAW12]. Garbled [WMK16, AEK^{+}18]. GARCH [Lin18]. gas [HHM^{+}19]. gastroscope [DLL^{+}19]. Gate [RCW^{+}19]. Gateway [FEB^{+}19, MPLM18, ACCM19, DRZ^{+}19, GRZ^{+}19, KZS^{+}19, MDA^{+}19, SVB^{+}19, SDF^{+}19, SG19, ZZBZ19]. Gateways [GDZ^{+}19, BOWD^{+}19, GDP^{+}18, GLD^{+}19a, GRCP^{+}17, KS19, RGN^{+}18]. Gather [FD12]. Gathering [GDP^{+}18, FG18, SNP19]. gatherings [KVK^{+}18]. gating [LLC14a]. Gauss [Jun18]. Gaussian [GSC^{+}19, KKB^{+}19]. Gazetteer [CAS^{+}16]. GbE [WLRL18]. Gbps [KSW^{+}13]. GC [SEPV19]. GC-Wise [SEPV19]. gCube [ACC^{+}19c]. Gearing [WMJW18]. GenApp [SB19b]. gender [CPW19, RSY^{+}18]. gene [JLC^{+}20, MSS^{+}16, YWG^{+}19, YWG^{+}20, YPJ19, KBVH14]. Gene/Q [KBVH14]. Gener [AB19a, AB21, ABMMC22, BFS^{+}17a, Bo20, Cha14b, DP20, DP21a, DP21b, GHEB^{+}23, HZX^{+}20, HYS18, JLC^{+}20, NDZ^{+}18a, NDZ^{+}19, SME^{+}21, WWP20, WCWC20, YWG^{+}20, YTQ20, ZMZ^{+}20, wZcZN^{+}20]. General [CJXX19, BRNR15, FWB13a, GXW^{+}19, HQZH14, LLT^{+}19, PCC18, VVB11, WJZ^{+}17]. general-purpose [BRNR15]. generalizable [LvW14]. generalization [WWZZ18]. generalized [BCdV^{+}19, PSLZ18]. generated [AHL11, DGGH11, HHXL13, LZP^{+}18, LJ19b, RTHB17, SK18, ZHL^{+}18]. Generating [JWW14]. Generation [ChK11, DO15, HYZS16, LJ17a, LBJ^{+}24, BDP11a, BGC19b, CJXX19, CsZzG^{+}13, DCMB15, DQXW19, GLM^{+}12, GMdFPLC17, HCX^{+}19, KVR15, KHMB13, MYBMM18, MVL18b, MR19, NDZ^{+}18a, NDZ^{+}18b, NDZ^{+}19, OMPSPL^{+}19, PJDO13, SB19b, VB18, YCZJ18, ZCX^{+}18, ZNC^{+}18]. generators [RNR18]. generic [DSPA18, FK11, GD10, LKK^{+}16, MRT^{+}19, SMSF18, TKA^{+}18a, TCN^{+}14, WMC19]. Genetic [JGB19, PMK18, ATF11, ACCD17, BRL19, DC19, DSCJ18, GdVC10, ISS^{+}15, KX11, MGA^{+}18, MRN19, PC18a, SF19, WYBS11, WLQ10, ZRZL18, ZLG^{+}14]. genetic-based [KX11]. genomic [JHC10, LJW^{+}19a, SBA^{+}17]. genomics [CJPC19, OdOD^{+}13]. Genotypes [JGB19]. geo [KB18, KKBK19, MPC^{+}18, WLA18b, XWZ^{+}19, ZBF14, DWJM18, YMY^{+}17]. geo-collaborative [ZBF14]. geo-crowdsourcing [WLA18b]. geo-distributed [KB18, KKBK19, XWZ^{+}19, YMY^{+}17]. Geo-QTI [DWJM18]. geo-replication [MPC^{+}18]. geographic [CES^{+}19, DWJM18, LC17]. geographical [CSL17, MBV^{+}15]. geographically [ZB19]. geometric [SHL^{+}19a]. Geometrical [SNXB17]. geometry [WWA19]. geoparsing [CGZL19]. geoscience [FFPS10]. Geospatial [GFD14, CCM^{+}14, CPLH19, CGL^{+}10, DGGH11, KZS^{+}19, LJ19b, MPP13, SBD^{+}18]. geostatistics [SDST18]. geriatric [HZM14]. gestures [CRRC18]. getting [PECA19]. giga [PDK10, PDK10]. giga-pixel [PDK10]. Giga-stack [PDK10]. gigabit [KDHP16]. GIS [FCD^{+}14, WZW19b]. GISs [LZXW13]. Giving [GTSAR^{+}14, GTCZG^{+}18, IG12]. GJMF [ÖE13]. gland [WZF^{+}19]. gLite [AAB^{+}10, KMV^{+}15, VKK14]. Global [MWYC12, BOWD^{+}19, CLL18b, GJ18, LSAM13, LCdPMCT19, MLC^{+}11, MVC^{+}13, PSJ^{+}12, RRB10, SCAC^{+}19, LNK^{+}18]. Globus [CLM^{+}16]. glottal [YXY18]. GMA [WNR19]. GMBS [KK10a]. GMonE [MSM^{+}13]. Go [CDH^{+}19]. good [SOA17]. goods [MEBA12]. Google [PS13]. GoSharing [LWZ^{+}19b]. gossip [AMRM18]. Gossiping [Fer13, ECPF17a]. gouty [YTQ19, YTQ20]. governance [BDH14]. Governing [IPCA^{+}16]. Government [LLW^{+}18b, SDK19, WLA18b]. governmental [PSY^{+}19]. governors [KANS18]. GPaaS [HB19]. GPGPUs [OSC14, ZXL^{+}18]. GPOS [JK17]. GPU [ADH^{+}16, CDG^{+}14, CPSD18, CSdCM^{+}17, CSP13, DNJG17, HSP^{+}13, KDHP16, LZX16, MAPA19, RBJ^{+}13, VD16, WMQ^{+}16, vW19]. GPU-accelerated [HSP^{+}13, LZX16]. GPU-based [MAPA19]. GPUs [CJPC19, CCZ^{+}19, RNR18, vWMBS14]. GRADE [FK11]. gradient [LMCSE19, MWL18a, RJS^{+}19]. grain [WC14]. grained [AMBB18, CZZ^{+}18, JLY^{+}18, LZLL18b, LZHY19, MLW^{+}18a, SJ18, TKK^{+}14, VVB11, WWH^{+}19, XTL^{+}19, XYML19, YAX^{+}18, ZMP10, ZFS^{+}18, ZSL^{+}19b]. gram [ZXM^{+}19]. Granary [YZW14]. granularity [MVC^{+}13]. Graph [MFC^{+}19, VSDD13, AdI14, AB18a, ADLM18, BGC^{+}19a, CWJD19, CWJ16, CWJ^{+}18b, CGM^{+}18, HO17, HB19, HDH^{+}18, kHsZwJW18, HZ19, JHC18, JLD^{+}19, LWXY19, NS17a, NJ19, NS19, QCZH19, SLG^{+}17, SMM^{+}14, WWH^{+}19, HB19, KWR^{+}13]. Graph-based [MFC^{+}19, HO17, SMM^{+}14]. graph-regularization [QCZH19]. graphical [MZL^{+}19]. graphics [GXW^{+}19]. graphs [BR19, BKB18a, DCMB15, DKFKF18, GRMSOG18, JWW14, KKI14, LYW^{+}18a, LSG18, LL18, MG11, Nag16, WWG^{+}19b, vdLLE19]. GRASP [PFRC16]. GRASP-based [PFRC16]. Gravitational [CT19b]. gray [CSG^{+}18, GDAS18]. gray-box [CSG^{+}18]. Greedy [SOD18, SOM^{+}19]. greedy-proof [SOM^{+}19]. Green [GDS18, KJFS12, NNLH18, YZZC19, DWS12, GCZ^{+}19, HCHH19, LLW^{+}12a, LJL12, MMPF19, TuIS^{+}19, WCC14, YLHJ14, BMK^{+}14a, ZSH12]. greener [DKK^{+}13, VTTK17]. Greening [GFW^{+}18]. GRelC [FNA11]. grey [CT19a]. GREYC [AGBR19]. GREYC-Hashing [AGBR19]. Grid [AJY15a, AJY12, AJY15b, AC10, BB13, CMA11, CPGBC16, CC11, CCM^{+}14, JY15, KCH^{+}13, MJRM16, NJ16, PAC^{+}17, AMN18, AOIS10, ASW11, AT18b, ATA19, AAQ^{+}19, AMT^{+}12, AEM10, AH11, BCC^{+}17, BMH10, CSV^{+}12, CLL^{+}14, CPSRG14, CCCT14, DST10, DW11, DLS^{+}12, EK11, EMM12, FG18, FK12, FMSSM12, FQBCF15, FFPS10, GD10, GSV^{+}10, GFW^{+}18, HAP11, HHW11, ISS^{+}15, KVR15, KZ17, KIS11, KCK16, KCV11, KV12, KKL11, LLC11, LYJ10, LXJD18, MCN^{+}18, MLC^{+}18a, ML11, MGR11, MLBS11, MW12, MVCC10, NAGD18, OB17, ÖE13, PSS13, RRS10, RC13, RLP12, RWV^{+}13, RMCMD12, RM11, SMPC12, ST11, SPR^{+}10, SSKK13, TV16, TSBH11, VS13, VHML11, VR12, WY17, YHL^{+}19, YY11, ZCLW18, ZZY^{+}19, CS19, FGM11, MM10, AR10, AM10, BKB11, BP10, BR10, CWW^{+}13, COC10]. Grid [CS12, CCD^{+}10, CGL^{+}10, DSD^{+}11, GRL11, HKPT10, HJC10, JS12, KWR^{+}13, KMI11, KRZ12, KN10, KVHT10, Li10, LSL^{+}15, MBB10, Mer13, MWMA10, MG10, NF13, NJHT11, PP10, PBC^{+}11, PDDS10, RMCN^{+}10, RGC^{+}10, SAGL10, SBB^{+}10, STA17a, SIL^{+}13, SSLF^{+}10, TLL^{+}11, VPT^{+}10, WCHL10, WXZL11, XA10, XAW^{+}10, YW12, ZMP10, ZLTY10]. grid-based [CSV^{+}12, LYJ10, MW12, YHL^{+}19]. Grid-computing [YW12]. Grid-enabled [GRL11, XAW^{+}10]. Grid-enabling [SSLF^{+}10]. GridFTP [RSK16]. gridification [MZC10]. Grids [HAP15, YCY10, ZS16, ACH^{+}11, AR15, AWN^{+}13, ASD12, BMT12, CLH10, CH10, CCS^{+}10, CGJ^{+}10, DVB14, DPL14, GCBM17, Hua10, HBN^{+}13, IAL10, KK10a, KK11, KX11, KKW^{+}14, KJ12, KK10b, KIC12, LLpC12, LAH10, MLG13, MGV^{+}18, MWPVB12, MVC^{+}13, OM10, PS10, PT16, PGCC^{+}10, QPTGG^{+}12, RRB10, RAdARP19, SR12, SHBP10, SPMC10, SEMJ11, SSL12, SCCS11, SAC11, SSB13, TZBK13, TJWS10, TPBS14, Tor13, VKK14, XY15, ZZL^{+}10, AL14, CB10, EH10, GXL^{+}12, GBS10, Lea13, Lea15, RPMG10, SVN10b, VDTK12, WS10, XCGD10]. GridSim [HLW12]. GridUFO [MWMA10]. GridWay [RMHCMG15, TCR^{+}12]. GridWay-based [TCR^{+}12]. GriF [ML11]. groundwater [ZWZ19]. Group [AJY15a, CB10, WLL^{+}19b, Bag11, BCF16, CDDR17, CLM^{+}16, CWJ16, FMRS18, GCK18, GZL^{+}18, GNVST14, IOV^{+}18, LO19, MHC14, NWD^{+}18, RLP12, SCEC18, SOM^{+}19, VCDK18, YZL^{+}18, ZLL^{+}19, GVdBdL15]. Group-based [AJY15a, CB10]. group-key [IOV^{+}18]. grouped [PWMX17]. groups [BCF16, ZZ15]. growing [YJHZ14]. growth [KZCW13]. GSA [CGSJ18]. GSP [KSS11].

GSR [LC17]. GTSIM [ZCDV19]. GTSIM-POP [ZCDV19]. guarantee [CMG^{+}19, LGL^{+}17, SCZ^{+}14]. guaranteed [KLP19, XTT18]. guaranteeing [CM17, LLW^{+}12b, UDST19]. guarantees [GJF^{+}12, LHPC^{+}19, SBK^{+}16, ZFW14]. Guest [BB13]. guests [JK17]. GUI [SB19b]. guidance [JYZ^{+}18]. Guided [GZS14, ZSMS18, JGFB18]. guifi.net [BFN18]. GWpilot [RMHCMG15]. GXP [TMM^{+}13].

H.264 [WC14]. Haar [MGA^{+}18]. Hadoop [BAJ^{+}19, CRB^{+}16, GGN17, IPCA^{+}16, NK17, RD14, TF18, WTR^{+}13, WLHH18, YH19, YIA17, ZTD^{+}18]. Hamiltonian [SJR13]. hand [KPG19, NUPA19, XL19]. handheld [RSRA18]. handle [GGDM^{+}18, LZW^{+}18]. Handling [BMZ10, HHL11, SGB^{+}18, TWdLZ19, XWL^{+}18]. handoff [SLS10]. handover [GSY^{+}19, LLF^{+}18b, YHL16]. handshake [LWL^{+}18]. handshaking [CJG^{+}18]. handshaking-based [CJG^{+}18]. Hardware [MHY^{+}18, VL19, YLJ^{+}17, AMN18, FFC12, KDHP16, KKvdB^{+}17, LAL^{+}14, PAB^{+}14, RSK16, TLC^{+}15, YÁJG^{+}15, CLK11]. hardware-assisted [RSK16]. hardware-aware [PAB^{+}14]. Harmonizing [LTZ15]. harnessing [FQBCF15, ZZC18]. Harris [HTL^{+}18, HMF^{+}19]. Hartmann [MAPA19]. harvesting [DJJ^{+}18, TSD18, VF18]. Hash [JL14, CHS^{+}18, HLL18]. hashing [QZD^{+}18, ZZLZ18, AGBR19]. hashtag [ZZJY16]. Haskell [BBH18]. haul [RSK16]. hawks [HMF^{+}19]. hazard [LHJC18]. HB [PYH^{+}18]. HCN [ESPN17]. HD [HSP^{+}13]. HDFS [TMDZ15, WZML18, YSC^{+}15]. HDS [PIP18b]. head [FJJ^{+}18]. Healing [LHBC16, AFB^{+}10, GZL^{+}18, dSGD13]. Health [AMSPL19, BSRR18, DP20, DP21a, DP21b, KIJ^{+}19, LHL15, RGN^{+}18, Rao17, VRS^{+}19, ABZK15, ABC^{+}18, APR^{+}19, CBN16, CLC^{+}19, CRWZ19, DP19, EAS^{+}18, GFD14, Ham19, HIA^{+}18b, MID16, ML19, PKY^{+}17, SPS18, TCN^{+}14, WMX^{+}17, YZL^{+}18, YZG^{+}18, YKÖ17, ZAA^{+}14, LG16a]. health-shocks [MID16]. Healthcare [IFD^{+}19, WRK^{+}15, WLP18, ZZXL18, AASI17, ABB^{+}19a, ASO14, AHMS18, APR^{+}19, CPD^{+}15, CLH^{+}18, CZZ^{+}18, FGW^{+}19, FFC^{+}18, FRM^{+}18, HLYW17, HXA^{+}17, HZM14, JNS^{+}19, JP18, JTBS15, KKB18, Kim14, KLV^{+}18, LWL^{+}18, LZLL18b, MVL^{+}18a, MGN^{+}16, MGA^{+}19, PMDS18, PSW^{+}19, RBA17, RGN^{+}18, RYH^{+}19, SAVS19, SJSA19, UKK^{+}19, WLS^{+}18, XKBA18, ZCK^{+}15, ZCDV19]. heap [KP12]. Heart [GMdFPLC17, DH16, DNW^{+}19, HJA^{+}19, RJN^{+}19, WLZ^{+}14]. heartbeat [SD18]. HEFT [ZZS^{+}19]. height [CFL^{+}18]. HELIO [BBC^{+}13]. heliophysics [BBC^{+}13]. Help [SCN^{+}14]. Help-On-Demand [SCN^{+}14]. hemotherapy [ATdC^{+}16]. Henry [HHM^{+}19]. HEPart [YWCC18]. hepatitis [WCWC19, WCWC20]. heritage [CGN18, MKS18, PPMM^{+}18, PC18b, WDJC18]. heterogeneity [BL15, PLLA18]. Heterogeneous [KV17, PZY16, PIP18b, VF18, ADAAD12, ABTF16, AMGCC18, ACK^{+}15, ACCM19, ABP16, AB17, AB18c, ABB19b, BDNP13, BBB16, CPGdS^{+}13, CDG^{+}14, CSdCM^{+}17, CGBAP18, CXDM18, CDDR17, ÇBCA15, CXL^{+}17, CSL18, DPK^{+}19, DT16, DR18, DCMW17, EP12, EMHE18, ECPF17b, FMSSM12, GS16b, GVA^{+}16, GSY^{+}19, IPG^{+}18, JLY^{+}18, KANS18, KKB14, LYJ10, LZXW13, LGY^{+}16, LWZ18, MT17, MRT^{+}19, MRS18b, NNRA19, NK18, OG18, PSPP16, PNZ14, PZY17, PDDS10, RD14, SHP^{+}16, SB19a, SAGGB17, SJTN18, SZK16, SLS10, SOIS12, SFR15, SSL13, SB16, SK12, TTH15, TMP15, VPT^{+}15, VD16, WHW16, WHS^{+}17, WCW18, WZWC18, WWZC19, WSC^{+}19, YLJ^{+}17, YS16, YCZJ18, YHH^{+}19, ZG19, ZMP10]. Heuristic [AA18, AL18, CSL18, ESW^{+}17, GSR^{+}19, KMT14, RC13, RS17a, XA10, XY15, ZZZC19]. Heuristic-based [AA18]. Heuristics [NBB18, RT15, BAB12, àCKPM19, DST14, DNP14, GVA^{+}16, GGS13, Man15, WCHL10]. HIB [CZXL18]. HIB-tree [CZXL18]. hidden [WQG15, XZP^{+}19]. Hiding [DSM^{+}19, CCD^{+}19, HZL18a, Sha16]. Hierarchical [BMT12, LLW^{+}18a, ABF^{+}15b, CWW^{+}13, EK11, FK12, HYS17, HYS18, KX11, OCCK14, PRC^{+}14, QPTGG^{+}12, RKB18, TTC^{+}14, WLGL19, ZB19, ZLTY10]. hierarchical-deterministic [WLGL19]. hierarchies [ECPF17b]. hierarchy [XL19, ZH17]. High [AJY12, APS^{+}19, AQB15, FBS18, GGJ13, HGM15, KDHP16, Kha12, LYS12, MKH13, MDT^{+}18, PMMAM13, PIP18b, SG14, WHMO13, WQG15, WMY^{+}18, Wri19, WXGM18, YTQ19, YTQ20, ZSX^{+}15, AFS16, APRC16, dRADFG18, BC17, BGMLS17, BRHH18, CND^{+}19, CMS^{+}18, Che14, DJPM18, DCC^{+}14, DJJ^{+}18, EP13, FdAGdAFV19, GAFFOG12, GSC11, GHO^{+}11, GG10, HAF^{+}16, HKPT10, HDB18, JRJ^{+}11, JLC18, KHG13, KMB^{+}17, KPM^{+}18, KC19b, KMK^{+}14, LRYJ17, LHX^{+}18, LSH^{+}11, MWL18a, MAPA19, NWMG17, NAAC19, PDW^{+}11, RPA^{+}18, RS17b, SB14, SPMC10, SGKC10, SEMJ11, SK18, SSZ13, SSP17, SYL18, TSWL17, TBdL16, TCN^{+}16, Tur18, VSBN19, WXLY16, WdL16, Wei11, YDK11, YK17, YNLY19, ZYZ^{+}18, ZBCT17, ZCQ^{+}16, ZYTC15, AHL11, MLC^{+}11]. high-available [SB14]. high-dimensional [DJPM18]. high-level [dRADFG18, CMS^{+}18, KMK^{+}14, NAAC19]. high-order [Che14, DJPM18]. High-Performance [SG14, WXGM18, KDHP16, MDT^{+}18, HKPT10, LRYJ17, LHX^{+}18, SGKC10, SEMJ11, SSP17, TBdL16, Tur18, ZYZ^{+}18]. High-resolution [YTQ19, YTQ20, DCC^{+}14, JRJ^{+}11, YDK11]. High-speed [LYS12, DJJ^{+}18, HDB18, KC19b, LSH^{+}11, MWL18a, MAPA19, RPA^{+}18, VSBN19]. high-stakes [CND^{+}19]. High-throughput [ZSX^{+}15]. Highly [CD16, LN18, MRH17, CWJ^{+}18b, Frî14, GVI13, JSS^{+}12, MAC14, MCA^{+}18, PMLVLS^{+}13, Tor13, ZYCZ19]. highly-threaded [MAC14, MCA^{+}18]. historical [JH16]. history [YWJ^{+}18, YSL19]. hitting [CSdCM^{+}17]. HM [dSMAdR^{+}17]. HMD [SHH^{+}19]. hoc [AAS^{+}19, BLMU19, CNP^{+}19, HHK18, KKN18, KIAD17, LLYW19, LAQ^{+}19, LLJ^{+}11, SGGCR^{+}16, SVK19, VCD^{+}18, YFY^{+}13, ZF16]. HOG [LZL19b]. holder [FZT^{+}18]. holes [SK12]. Holistic [BBWB^{+}18, LJGW18, Ano12r]. holographic [CGN18]. home [ACPI19, AC18, CBPP18, JBC16, MVL18b, PMDS18, KADJ14]. home-forwarding [MVL18b]. homes [Bae14, LRJG19, Mat18, OCW14, TOD17, TZL^{+}18, YSHM19, GMLGB^{+}17, HSS17]. Homogeneous [TKA18b, PSPP16, SSP17]. homomorphic [CJXX19, MK17, RTS^{+}16, ZXJ^{+}14]. Honeypot [PD11]. Hop [WWTF18, KESL17, LLAW17, SAH19, TZD^{+}19]. HOPE [LM12]. Hopfield [TZBK13]. Horizons [Hel16]. horizontal [KAEC^{+}18]. Horse [CLK11]. host [MDB^{+}18a]. hosted [YKK13]. hostile [GL19]. hosting [PVN^{+}12]. hot [LZL^{+}17]. HPC [LPK18, SG14, AHEM17, ALM^{+}10, BC15, BC17, BL13, CXDM18, CLDC19, CLP^{+}14, CGL15, CRTN17, DYI^{+}16, ETR^{+}13, FDGR14, GDP^{+}18, JOPW14, KBVH14, KKBK19, LPK17, LZX16, MAJD18, MGR11, MDO^{+}15, MSM^{+}18a, RRP^{+}14, SNC18, TZQ18, TKR^{+}15, WMLS14, WOPW13, WG13, WDR^{+}19, YZI18, YDT19, ZGB^{+}17, ZME^{+}15]. HPC/cloud [BC15]. HPDMnet [MLC^{+}11]. HPS [BGMLS17, PIP18b]. HSim [LLAH13]. HTM [WHZ19]. HTML [LYC^{+}19]. HTTP [WLYL11, dCRL^{+}19]. HTTP-based [WLYL11]. HTTPS [XZL^{+}19]. HUBzero [SG19]. huge [CWW^{+}13]. Human [LY19, WCB^{+}18, ZTC^{+}19, ABD^{+}19, AS19a, BAPS14, CZXL18, CPMG19, GMP^{+}17, HZW19, HUMA18, HZZ^{+}18, HZM14, HMMW19, KSS19, KKP19, OPT^{+}17, PYH^{+}18, PB18, PKA19, UPP17, WY19, XLL^{+}19c, YXY18]. human-chatbot [CPMG19]. human-computer [WY19]. human-workpiece [OPT^{+}17]. humanistic [HZM14]. humanities [BH13]. Humans [GMM18, AS19a]. hunt [ZSS^{+}18]. Hunting [GGC18, HDKC18, HDA^{+}19]. HVS [BMU18]. HW [XZL^{+}19]. HW/SW [XZL^{+}19]. Hybrid [CLL^{+}18a, CWW^{+}13, HLZ^{+}19, HEES19, LDY^{+}18, MGR11, MKM11, OPO13, TCC18, VMN^{+}18, WM14, YJHZ14, ABMESM18, APAZ17, ABC^{+}18, AR18, AFSH^{+}19, ATH^{+}19, AB18a, AMGCC18, AK14, AM19a, BC15, BBB^{+}19, BCD^{+}18, CT19a, CRM^{+}16, CVKB12, CRC13, CGBAP18, CFG^{+}19, CAC^{+}15, CSJ^{+}17, CGSJ18, CRTN17, DBS14, EAS^{+}18, FJL^{+}16, FDPR17, FWB13a, GAJP18, HZL^{+}19, HZLH19, HIA^{+}18c, LBD18, LA19, LLC^{+}14b, LDX19, LQLX10, LZX16, LTC^{+}19, LSV^{+}18, dSMAdR^{+}17, MFN13, MPF^{+}16, MLSF16, Nag16, PPLL17, PRN14, RBN13, RS17a, ROK19, STMV18, SP18b, SLZ^{+}18, TZST14, TNY17, TdPF^{+}17, TSB18, VF18, VVB13a, VPT^{+}15, VCKB12, WXLY16, WMQ^{+}16, WXPL17, YLN15, ZCH^{+}17, ZYW^{+}18, ZWJ^{+}18, LM12]. hybrid-cloud [LBD18]. hybrid-indexed [WXLY16]. hybrid-multi [HIA^{+}18c]. HybridNN [FWB13b]. Hybrids [HAA^{+}16]. hydrological [HLZ18]. hylomorphisms [CHSA18]. hyper [AL18, RC13]. hyper-heuristic [AL18, RC13]. hypercubes [DK14]. HyperFlow [Bal16]. Hypergraph [RRKA19, SKS17, YWCC18]. hyperheuristics [MOBD18]. Hyperscale [RMRSA19]. hyperspace [NLM^{+}16]. hyperspectral [Cha11]. hypervisor [ASB18, RS16].

I-TASSER [ZZBZ19]. I/O

[CSL18, DLZ^{+}14, DLXR14, DYI^{+}16,

TDFZ18, ZXL14]. IaaS [AA18, AMMC18, AEME^{+}18, CMB17, GLNT13, LCH^{+}18, MJDN15, MGG^{+}17, NCS12, TTB^{+}13, TTH15, VVB15, YDQC19, ZT19]. iBike [ZWQ^{+}19]. IBRIDIA [ATH^{+}19]. iCAFE [SSK^{+}19]. ICAuth [SBL18]. ICPADS [TCG14]. ICs [VF18]. ID [Wan18a, WDZ19]. ID-based [Wan18a, WDZ19]. IDE [SCJ^{+}19a]. identifiable [SWW^{+}13]. Identification [WLZ^{+}14, AsRA^{+}19, AFO^{+}18, ARP^{+}19, AJ19, ASAAM^{+}19, BRL19, CPW19, CPP16, HNQ^{+}18, HZW^{+}16, KZA11, LLQS14, LWT18, VOCHC17, WLLF16, WCB^{+}18, WNR19, ZWJ19b]. identify [HIA^{+}18c, KAS^{+}18, MBL^{+}19]. Identifying [AHS^{+}18, BTP19, PSLZ18, WCM^{+}19, XTF^{+}19]. identities [TOS18]. Identity [HSM13, HYF18, ZDW^{+}16, CLM^{+}16, HCL^{+}17, LK12, MLM16, PLCGS11, WMX^{+}17, Wan18b, WZCH17, YWJ^{+}19, YXA^{+}16, Sar18a]. Identity-based [HSM13, HYF18, ZDW^{+}16, LK12, WZCH17, YWJ^{+}19, YXA^{+}16]. Idle [CLR18]. iDoctor [ZCH^{+}17]. IDS [HNCJ13]. IEC [YS16]. IEEE [AAQ^{+}19, KZ17]. IEEE/ACM [KZ17]. IgA [QXZ^{+}19]. IGRC [YHL^{+}19]. II [RVC16a]. III [YDD^{+}18]. illegitimate [WTTH19]. illumination [LLU^{+}18]. Image [MSM^{+}18b, YXD18, AMBB18, AM19a, Bro19, DGA18, FMV14, GHEB^{+}18, GHEB^{+}23, GPV^{+}14, GDAS18, sGbKS19, HLC16, Ima19, JYZ^{+}18, KO11, KZA11, hKBB11, KVHT10, LHJC18, LSZ^{+}16, LZL^{+}12, MZYA19, NOF18, OMKM^{+}19, RRKA19, VOCHC17, XWX^{+}17, XPL19, YDK11, YCY10, YSC^{+}15, YJS18, ZFY18, ZSMS18, CSV^{+}12]. image-retrieval [YJS18]. imagery [AAM^{+}19, CMI^{+}19, PDK10]. Images [YTQ19, AM19a, BCdV^{+}19, Cha11, CZ19, DSM^{+}19, DCC^{+}14, FCD^{+}14, GRS^{+}19, HAM18, HWWT12, JP18, KMK^{+}19, KIJ^{+}19, KHO^{+}19, LLU^{+}18, MAPA19, QMCX19, TA19, WWCN13, XJWW15, YTQ20]. imaging [HLZ^{+}19, OM10, SYCH18]. imbalance [CdSDS15, PAL^{+}19]. iMeter [YZLQ14]. Immersive [WY19]. immune [BZMY10, CZY^{+}18, LFL^{+}17]. Immunization [ZHL^{+}18]. Immunization-based [ZHL^{+}18]. Impact [HAF^{+}16, KESL17, RA12, dACNC16, ABMM18, AGKZ18, BOWD^{+}19, CRVZ15, CRB^{+}16, CBLS13, HHD^{+}12, dSCD^{+}19]. impacts [BNJ16, HH19, PLLA18, SCG^{+}18]. impaired [KPG19]. impelled [PRN14]. imperfect [EU19]. Implantable [BDM^{+}19, ABS^{+}18]. Implementation [YCY10, YCD^{+}19, YCL^{+}19, AMN18, AAB^{+}10, ACHP19, GVDT16, GHLW18, HZDS19, HLV^{+}16, HHZ19, JYY^{+}17, JNR12, KO11, KANS18, KDG^{+}19, KVHT10, LJS17, LGW^{+}17, MK17, OBG^{+}18, PMT10, RPH19, SPR^{+}10, SAVS19, TLC^{+}15, TMM^{+}13, TBB^{+}17, XKJ^{+}18, ZN12, ZZQ^{+}13]. implementations [DFGR14, ZSW^{+}18a]. implemented [BTM10]. Implementing [CMD^{+}14, KFK19, PK11]. Implications [PSL19, CHS^{+}18, IHK^{+}18, LPD^{+}13]. implicit [CWJ16, XLZ^{+}14]. importance [AMT^{+}12]. important [LXM^{+}18]. imported [XLL^{+}19b]. imprecise [KRD^{+}19, SK12]. IMPRECO [CGV10]. imprint [CWUS19]. improve [CLP^{+}14, GMLGB^{+}17, LFH^{+}15, OdOD^{+}13, PLZX19, RMSPP17]. Improved [ESPN17, SE19, TA18, TV16, WLL^{+}19a, AMÇ19, BBI13, HZL18b, HLZ18, JLS19, JLQZ18, KKAS19, KI19, LZXG12, TM19, WN10, YHL^{+}19, YPJ19, YDD^{+}18]. improvement [CHS11, MJM^{+}16, RJN^{+}19, WQZ19, WTS14]. improvements [BGC19b]. Improving [AkBAL^{+}19, AMMC18, BVFGWA15, BL13, ECPF17a, GIM16, HAF^{+}16, HCNT14, HMA18b, HXL^{+}18, LLC11, LYJ10, MYW^{+}19, PAB^{+}14, SG13, WSC^{+}19, WLML17, YZI18, ASD12, GVI13, GJKP18, SAH19, SBA^{+}17, XTF^{+}19]. improvisation [VSKS19]. imputation [CRYG18]. In-band [WGX^{+}19]. In-Mapper [MLC18b]. In-memory [GQLX18, BYL^{+}18, GXW^{+}19, HZDS19, USK16, UDvdW^{+}18, WZS^{+}18]. In-network [PJDO13, ZZDM^{+}18]. in-place [PWMX17]. in-transit [AHP16]. incentive [DWJM18, GA13, HHK18, HDLW13, JLQ^{+}17, LLW^{+}19d, LWZ^{+}19b, SOM^{+}19, XY15, ZA13, ZCDV19]. incentive-based [XY15]. incentive-compatible [HDLW13, SOM^{+}19]. incentive-driven [HHK18]. incentives [KRZ^{+}19]. incentivized [BDP11a]. incidence [XL19]. incident [CES^{+}19, FM17]. incident-supporting [CES^{+}19]. incidents [TMS^{+}17, dSGD13]. inclusive [FLT17]. incompatible [SL19]. inconsistency [YWJ^{+}18]. incorporate [YMW13]. Incorporating [LYS^{+}19]. Increasing [IDM^{+}16, KKAS19, KH18a, WWX^{+}17]. increment [TWG^{+}19]. Incremental [PBL^{+}18, XWM18, ArMS19, CAPG18, FRB^{+}14, KMB16, LKM14, LY18a, LPY^{+}18, YNLY19, ZFS^{+}18]. incrementally [YGYW16]. Independent [FMSSM12, CFL^{+}18, EG18, GVA^{+}16]. index [CZXL18, CPLH19, WXPL17, Lin18]. indexed [LSL^{+}15, WXLY16, YNLY19]. indexing [ARP14, FMV14, LJ19b, RTS^{+}16, SB17a, WZW^{+}19a, WHMO13]. indicative [ZNC^{+}18]. indicators [DPS16, HZX^{+}19, HZX^{+}20, HAAR^{+}19, NAAC19, SA19, KJFS12]. indirect [VS19].

INDIS [GVTdL18]. indistinguishability [YZW^{+}18]. indistinguishability-based [YZW^{+}18]. individual [CN17, HZM14, NZOCJ^{+}19, PS10]. Indoor [SLK17, YCD^{+}19, HHH^{+}19, HST^{+}18, HDH^{+}18, HZZ^{+}18, LvW14, Mat18, MLGGB^{+}17, MKS18, OMD^{+}18, PECA19, YYD^{+}14]. induced [LSV^{+}18]. induction [PMK18]. Industrial [ZWJ^{+}19a, DM12, GHD19, LHO17, LW18a, LW19, PTD^{+}18, PWA^{+}19, uRYS^{+}19, NSR^{+}19]. Industry [JC15, FG18, QGT^{+}18, VCL^{+}19, ZSMS18, ZWJ^{+}19a]. infants [CPE^{+}17]. Inference [CBC^{+}19, JLS19, LBD^{+}19, MFT^{+}17, NA19, SL19, TCC18, WMBV17]. inferred [YWY^{+}17]. inferring [DYC^{+}18, XTL^{+}19]. InfiniBand [KBVH14]. Influence [GGH^{+}19, NBB18, HMC19, MFT^{+}17, TL19, WJLW18, WDD18, WGM15]. influencing [KSC^{+}19]. Influential [ZSJ19, WCM^{+}19, ZZC14]. Informal [GPJA^{+}14]. Information [ABS11, AT18a, BBB^{+}11, ChK11, CCRL18, DP17, MSO18, TAB^{+}18, URKM19, ZZ15, AHL11, AAA^{+}19, ATdC^{+}16, AM19b, BG12, BT17, BNJ16, BGP^{+}17, BMH10, CPW19, CCJ16, CWZ^{+}17, CWJ^{+}18a, DSM^{+}19, DMPS19, DGGH11, ELAEAVAM19, FAL^{+}19, GZLZ16, GRTV10, GCCL18, GCTLA^{+}19, GFD14, GWC^{+}16, HZC10, HPP^{+}18, JTL^{+}19, KPB18, KV12, KAS^{+}18, KB16, LHJC18, LNB14, LWK^{+}18, dSMAdR^{+}17, MVG18, Mat18, MMF16, OB17, PARMF14, RLL^{+}17, RGC^{+}10, SMF^{+}19, SPT^{+}18, SSL^{+}19, SBD^{+}18, SLD^{+}18, TWG^{+}19, TGM^{+}19a, TJWS10, TSOB15, TQL^{+}19, TGM^{+}19b, TCH19, VS13, VPT^{+}10, WRCC17, WXZ^{+}18a, WZ18, WWZ18, XSMS15, XCGD10, XLL^{+}14, XPL19, XLL^{+}19c, YSL19, YHH^{+}19, ZWS^{+}12, ZWJ19b, vdHSL^{+}15]. Information-centric [AT18a, RLL^{+}17, TGM^{+}19b]. Information-Powered [TAB^{+}18].

Information-theoretical [ZZ15].

inframetric [FWB13b]. InfraPhenoGrid [PAC^{+}17]. infrared [wZcZN^{+}19, wZcZN^{+}20]. Infrastructure [HGM15, TSTL16, VDTK12, AB19b, ATM^{+}19, BBWB^{+}18, BMU16, BC15, BH13, BCD^{+}18, CCMP18, CCM^{+}14, DWS12, FG14, GGTRRC16, GAI^{+}18, GRX19, HMM18, HZW^{+}16, KZ14, KuRAk^{+}18, KCH^{+}13, KK14, MSS^{+}13, MLC^{+}18a, MRT^{+}19, MBS13, MGLPPJ13, PP10, PAC^{+}17, RAdARP19, RMVG^{+}10, XHW19, ANE13, CPGBC16, GRL11, LPD^{+}13]. infrastructure-as-a-service [KuRAk^{+}18, KK14, LPD^{+}13]. infrastructure-level [FG14]. Infrastructures [CMNK19, CMZ^{+}12, CRM^{+}16, CA15a, DQC^{+}19, ENC^{+}12, EMJ^{+}13, FQBCF15, FPGK18, GDP^{+}18, GFD14, HZ10, KdGP^{+}19, LTN10, LPK17, LPK18, LPBB^{+}18, MT17, MFN13, MPR^{+}16, MLSF16, PGCML^{+}19, PPL^{+}15, PWA^{+}19, RMHCMG15, SFR15, VHML10, WTM^{+}17, WHYZ18, dSGD13]. infused [ZSP17]. Inhibition [JLC^{+}20, YWG^{+}19, YWG^{+}20]. initialization [PCK19]. initiative [ASAAM^{+}19, MLC^{+}11, YCH19]. Injecting [GRTV10]. injuries [PWP^{+}18]. injury [Bo19, Bo20]. innovations [SSFFR19]. Innovative [Cuz14, TGM^{+}19b]. InOt [PLA18]. input [WHCZ18]. input/output [WHCZ18]. inputs [Wan19]. insecurity [BP13]. Insertion [GK18, MAÇ17, LZLL18a]. InsideNet [MMAA19]. insider [SMF^{+}19]. insight [BCD^{+}18]. insights [AC16, GTSAR^{+}14, WLHH18]. inspection [HLL^{+}19, LSZ^{+}18, RSRA18]. inspired [AT19b, BW13, CWLZ19, FM10a, GWW^{+}19, IGB^{+}14, KKS18b, LCW^{+}18, VV16, XZZ^{+}14, XZZ^{+}19, ZYW^{+}18, dSFD^{+}19]. instability [Bo19, Bo20]. installations [ACPI19]. installment [SOIS12]. instance [FCY18, WWZ^{+}19]. instance-intensive [WWZ^{+}19]. instances [CLRL18, DR18, FEÁ19, GdCP19]. Institute [GDZ^{+}19]. institutional [BH13]. institutions [PKP19]. instrumentation [BDZ13]. insurance [ABZK15]. integer [GGA^{+}17]. integrate [BOP^{+}14, GRCP^{+}17]. Integrated [AS18b, AIP^{+}19, CM17, CSV^{+}19, FHYH15, FCD^{+}14, HFL^{+}19, KAS^{+}18, MPF^{+}16, PDH18, SGKC10, SH19, WZW19b, WS10, WHCW19, YZLQ14, YH19]. Integrating [HC17, LCL14, LHY^{+}19, BDZ13, CT19c, PSW^{+}14, TQL^{+}19]. Integration [APRC16, BdDPP16, CCY^{+}18, DP17, FP14, HHZ19, RAA^{+}18, URC19, GACM17, HAAR^{+}19, JNS^{+}19, LLF^{+}18b, MGV^{+}18, RMC^{+}18, SMM^{+}14, SPKG18, SAC11, WLLF16, YWF^{+}10, ZMP10]. Integrity [DL19, FLT^{+}19, GSC11, HKA^{+}18, JZJ^{+}18, LSG18, LXRS19, LYZC15, PPG19, WWRS16, YXA^{+}16]. Intel [CDMR19]. Intellect [AS18b]. Intelligence [ACC^{+}16, Cha14a, Jun17, SLS^{+}19, AB18b, Cha14b, GJKP18, GSN^{+}18, HDA^{+}19, HLT^{+}18, JNR12, LSV^{+}18, LFY^{+}19, MCWP16, NPH19, NAM^{+}19, OdI14, OB19, QC18, SSST17, SJL^{+}18, WM14, YZWG18, YJHZ14]. intelligence-based [OB19]. Intelligent [ACWJ19, BDCC19, DX14, GCD^{+}18, Hsu14, Kim18, LRZ^{+}18, LWZ^{+}19a, NNC^{+}19, XKJ^{+}18, XXX^{+}19, YCD^{+}19, YSL19, ZWQ^{+}19, AHMS18, BKS^{+}18, CPDJ13, COC10, CRC^{+}19, HMW^{+}19, HCX^{+}19, Kol18, LLZ^{+}18b, LXT^{+}19, LLW^{+}19d, LQF19, LWZ^{+}19b, MAY18, NAD^{+}18, NWL17, OMD^{+}18, PTD^{+}18, PKY^{+}17, PLP^{+}19, SSI19, Śle14, SDC11, VFHB14, XB14, XTL^{+}19, XZZ^{+}19, XLZ18, ZDL^{+}19, SSK^{+}19]. intelligent/cognitive [BKS^{+}18]. intensity [GZZ^{+}18]. intensity-modulated [GZZ^{+}18]. Intensive [DO15, ATdC^{+}16, BMK^{+}14b, Cuz14, DST10, GVURIVBV14, GJ15, IDCJ11, JFZL17, KS18a, KGVW14, LTTL19, LFHQ19, NSSA^{+}14, RSJ^{+}14, SLC^{+}17, SBD^{+}18, TdPF^{+}17, TBdL16, TCBC18, TSB18, VVB13b, WTR^{+}13, WZZ16, WWZ^{+}19, WHYZ17, WHZ19, YFY^{+}13, dSCD^{+}19]. intent [KPM^{+}18]. intention [CT19c, JLQZ18]. intentions [FTD17]. Inter [OMPSPL^{+}19, CLL18b, DPK^{+}19, GGLW18, LD17, LYS12, SRZD15, SB16, SBP^{+}17, SBL18, VOS12]. inter-cell [DPK^{+}19]. inter-cloud [CLL18b, SRZD15, SB16, SBP^{+}17, SBL18]. inter-connectivity [VOS12]. Inter-Pulse [OMPSPL^{+}19]. inter-purchase [GGLW18]. inter-service [LD17]. inter-view [LYS12]. interaction [AKB18b, CGN18, CPMG19, FJA^{+}18, HYC^{+}18, JRJ^{+}11, KID^{+}16, Nag16, RRH16, XZZ^{+}19, ZMN19]. interactions [OPT^{+}17, SDH^{+}19, UZ11, ZGS^{+}13]. Interactive [CAS^{+}16, MG18, SHN10, ABG18, CGN18, GKW^{+}12, HHZ16, OPO13, PTM^{+}18, SSFFR19, SL19, WY19, XFM16]. interception [MGLPPJ13]. intercloud [Erd13]. interconnect [YGYW16]. interconnected [FCY18, FGW^{+}19, MMLO18]. interconnection [GS13, LL18]. intercontinental [CKP^{+}19]. interdiction [KPS18]. Interdisciplinary [AC10, NS10]. interest [AS19a, JXZ^{+}19, LWSY18, LW18b, MRS18b, XZ16]. Interface [WLB11, ABS^{+}18, àCKPM19, DCBF19, GHY^{+}18, MFL18, MYK16, PSBB15, SRdlPG19]. interface-adaptation [PSBB15]. interface-based [GHY^{+}18]. interfaces [ABF^{+}15a, dRADFG18]. Interference [vdLLE19, DPK^{+}19, SS13, SMM^{+}14, VVB13b, ZF16]. interferon [WCWC19, WCWC20]. intergenerational [MCG^{+}15]. intermediate [LZW^{+}18, TZLL18, WZZ16]. International [CC11, CGJ^{+}10, GDZ^{+}19, MLC^{+}11]. Internet [AMQS^{+}19, DP20, DP21a, DP21b, GTEL^{+}18, HZW^{+}18, HAT19, IJCR19, JBM^{+}18, LKCS18, MWQ^{+}19, NSR^{+}19, RC18, RC19, SYJA19, SSW^{+}19, VKT^{+}19, YJHZ14, ZAI^{+}18, LTC12, QCY^{+}19, ABMM18, AKP^{+}18, ACWJ19, AR18, AGR19, AT18a, AT19b, ACSdRR17, AHYF19, AVPV17, ASAA18, AMPZ16, BA17, BAJ^{+}19, BS17, BZ19, BRH18, BGS^{+}19, BdDPP16, BRB19a, BRB^{+}19b, BWG19, CT19a, CCRL18, CMI^{+}19, Che14, CBT^{+}19, CMP^{+}17, CBPP18, CDFW18, CMZ^{+}18, CDH^{+}19, DSK^{+}14, DPK^{+}19, DZH18, DYY^{+}19, DGR^{+}19, DP19, DC18a, ESW^{+}17, EAS^{+}18, FG18, FJJ^{+}18, FTK17, FRM^{+}18, FPL^{+}19, GLC19, GMLGB^{+}17, GBB18, GHYK18, GMD19, GCK18, GZW18, GLD^{+}19b, GBMP13, HAJ^{+}19, HDKC18, HKA^{+}18, Ham19, HZL18a, HHW^{+}19, HHK^{+}16, HPP^{+}18, HFL^{+}19, HIA^{+}18b, HCL^{+}17, HNQ^{+}18, HXL^{+}18, HSS17, JKAU19, JYY^{+}17, JYZ^{+}19, KWK^{+}18, KK19, KOT18, KSK^{+}19, KCM19, Kim18, KLH^{+}18, KLMB19, KLJS19, KMST19, LLMP13]. Internet [LKJ^{+}19, LBD^{+}19, LS10, LHO17, LGL^{+}17, LYC18, LWZ^{+}19a, LW19, LXF19, LDS^{+}18, LYYW19, LLW^{+}19d, LZJL19, LRBW17, LWW^{+}13, MK17, MVL^{+}18a, MGL^{+}18, MLGGB^{+}17, MMC^{+}18, MDM^{+}19, MGN^{+}16, MPLM18, NWT19, NJ18, NLS19, OFD17, PC18a, PTD^{+}18, PLGMCdF18, PPMM^{+}18, PC18b, PVHTP19, QZM^{+}18, QCZH19, RGN^{+}18, RMSPP17, RJS^{+}19, RACA18, RHPV17, RGM^{+}19, RMDB18, SAGGB17, SYJ^{+}19a, SHS^{+}19, SST18, SAH19, SWY^{+}18, SHL^{+}19a, SDDG17, SYW17, SCZ^{+}14, SLH^{+}19, SCG^{+}18, TLSC17, TLL^{+}19, TSS^{+}19, URKM19, WWX^{+}17, WZW^{+}19a, WYJ^{+}19, WZ18, WWZ18, WLZ^{+}19, WDW^{+}19, XWW19, XXQ^{+}19, YLVY15, YWZ^{+}18, YCT15, YWLL19, YHA^{+}19, YN18, YAP16, ZPPE17, ZYA^{+}18, dFBP^{+}17, uRYS^{+}19]. Internet-based [VKT^{+}19, LWW^{+}13]. Internet-of-Things [FRM^{+}18, GCK18, LKJ^{+}19, LW19, RGN^{+}18]. Interoperability [FKBG10, AFS16, ACWJ19, BA17, BNJ16, BGNI19, CXDM18, DGCGH^{+}17, EHT10, GTCZG^{+}18, HA18, MRS^{+}18a, NAD^{+}18, SEMJ11, SGM11, TKK^{+}14]. Interoperable [KH19, CMD^{+}14, DMMM11, KK10a, KKB14, NWT19, SCY^{+}18]. interpretability [BDA19]. intersection [NWL17]. interval [FZT^{+}18]. Intervals [OMPSPL^{+}19]. Interweaving [RM16]. InTime [PBA18]. intra [GZLZ16, HMW14, JLY^{+}18]. intra-AC [HMW14]. intra-domain [GZLZ16]. intra-frame [JLY^{+}18]. intrinsic [AIP^{+}19]. Introducing [ACSdRR17]. introduction [BB13]. Intrusion [GLXF17, ABDH19, ATX13, DYY^{+}19, FZHH14, HNCJ13, KHWZ18, LYJ10, LTMW19, PMK18, SI18, SBK18, VSBN19, XFTZ16]. intuitive [MKS18]. Invariant [WWSL19, GMdFPLC17]. Investigating [CWZ^{+}17, NAM^{+}19]. Investigation [PWP^{+}18, AAN^{+}18, KR14, PLLA18]. Investigations [AMM16, DC18b]. Investment [YAA^{+}19]. investments [HA18]. Invited [ECPF17a, RW13]. Invocation [MBB10]. IO [MYW^{+}19]. IoFClime [MLGGB^{+}17]. IOFollow [MYW^{+}19]. IoMT [HJA^{+}19, KIJ^{+}19, LQF19, RYH^{+}19, SJSA19, TSAER18]. IoMT-based [LQF19]. ionic [SWW^{+}13]. iOS [DC17, DC18b]. IoT [ABMM18, AT18a, AHYF19, ASAA18, BZ19, GBMP13, KK19, PPMM^{+}18, RMDB18, SME^{+}21, ZYA^{+}18, Zhu18, AZH18, AHS^{+}18, AUSA19, AMSPL19, AMKC19, AJ19, AT19a, ATA19, AHU^{+}19, AMR^{+}19, AKB^{+}18a, AGA18, AMBC19, APR^{+}19, BDL^{+}19, BKS^{+}18, BGNI19, Bu18, BW19, CFP^{+}19, CBC^{+}19, CES^{+}19, CGM^{+}19, DRC^{+}19, DSBC19, DdSdN^{+}19, DSPA18, EBCP18, FSY^{+}19, FXG^{+}19, FFC^{+}18, FJA^{+}18, FMN^{+}17, GAI^{+}18, GZL^{+}18, HC17, HLL18, HHK18, HSBE19, HST^{+}18, HLT^{+}18, IHA18, JP18, JSC^{+}15, KSS19, KH19, KS18b, KPS18, hKRM17, KPJ19, KLP19, Kol18, KMST19, KLV^{+}18, KCCL18, LJS17, LOR^{+}18, LTMW19, LJ19b, LYZC15, LMM19, LYL^{+}19, LMCSE19, LWF^{+}17, LYH^{+}19, Mat18, MDB^{+}18b, MPI^{+}18, MOFGP18, MWL^{+}18b, MRS18b, MGA^{+}19, NAD^{+}18, NGB18, NDA^{+}19, OSANAM19, PFPJ18, PDH18, PGCML^{+}19, PCK19, PWA^{+}19, PPS^{+}18, RTHB17, RMC^{+}18, SME^{+}19, SISGS18, SDST18, SCY^{+}18]. IoT [SSL^{+}19, SGB^{+}18, SB18, SPS18, SL19, SPKG18, SLD^{+}18, SZW^{+}19, TF18, TOD17, TZL^{+}18, TMB^{+}19, TSGVRGS19, TAS^{+}18, TGM^{+}19b, URC19, UGBM^{+}17, VF18, VRS^{+}19, Wan18b, WWZ^{+}19, WLP18, WYH^{+}17, WGX^{+}19, XLL^{+}19a, XYML19, YZC^{+}19, YWJ^{+}19, YH18, YZG^{+}18, YSHM19, YAGG18, ZYZ^{+}18, ZLY^{+}19, dCRL^{+}19]. IoT- [TOD17]. IoT-aware [AMR^{+}19]. IoT-based [CES^{+}19, FJA^{+}18, PPS^{+}18, TZL^{+}18, UGBM^{+}17, ZLY^{+}19]. IoT-Cloud [HSBE19]. IoT-enabled [ATA19, AKB^{+}18a, TAS^{+}18, XLL^{+}19a]. IoT-FBAC [YWJ^{+}19]. IoT-NDN [MRS18b]. IoT-wireless [SCY^{+}18]. IoTBD [CCRL18]. IoTs [MKS^{+}19, ZSW^{+}18a]. IoTSim [BGC^{+}19a]. IoTSim-Stream [BGC^{+}19a]. IoV [CWL^{+}19, SLS^{+}19]. IoVs [TZD^{+}19, WLH^{+}19]. IP [JL14, KHJ10, RS17b, SK18]. IPV6 [LLL^{+}19]. Iris [GTEL^{+}18, CTU19, NWMG17, MCAS19]. iRobot [HMW^{+}19]. iRobot-Factory [HMW^{+}19]. iron [YLL^{+}19]. irrelevant [KMC18]. irrigation [SDST18]. islands [LG16b]. isolation [CW13a, JK17, MK19b, ZGB^{+}17]. isometry [YSZW18]. Isomorphisms [TX14]. ISP [ASA19]. ISSM [CT19c]. Issue [ADALZ14, CSYY18, DPDS14, DO15, GMP^{+}18, HYZS16, JY15, TCG14, TKRA14, YGS16, YJHZ14, ADLW12, ARB12, BB13, BDF^{+}16, CRW^{+}16, CPSRG14, DFRW17, DDB14, KZ14, KJ12, LNB14, PC18b, SZV19]. Issues [GTSAR^{+}14, LLWZ18, LKA^{+}19, UKK^{+}19, HHZ19, KLH^{+}18, LNB14, MKS^{+}19, SWY^{+}18, SYK^{+}17, URC19, YKÖ17, ZZF18, ZL12]. item [LCL^{+}16]. item-based [LCL^{+}16]. ITEMa [CGM^{+}19]. Items [BB17, CEP19b, CCMP18, ÇÖ13]. itemsets [CLM14b, LHW^{+}18, YK17]. Iterated [ML17, SOD18, WH19]. iteration [DGR^{+}15]. iterative [CCL11, FM10b, GZWQ13]. ITö [WDD18]. IV [HHXL13].

JACEP2P [CCL11]. JACEP2P-V2 [CCL11]. jammer [WWTF18]. Java [CDMR19, PLL^{+}18, SEPV19]. JetStream [TCN^{+}16]. JGRIM [MZC10]. Job [BCMM18, CCD^{+}10, CRTN17, KSC^{+}19, AAB^{+}10, BZMY10, BCC^{+}17, FCY18, GD10, GPS13, KHG13, KPJ19, LZ10, LLC11, LGL^{+}17, MLBS11, ÖE13, QPTGG^{+}12, SWG^{+}16, TZBK13, TZST14, WWC14, WMLS14, WS10, WXZL11, XY15]. Job-resource [CCD^{+}10]. job-shop [BZMY10]. JobPruner [SNC18]. jobs [GD10, JLL17, LAH10, LBU^{+}10, RM11, RMHCMG15, WDR^{+}19, ZQB^{+}18, ZA14]. Joint [AAQ^{+}19, KK10b, AMAY19, ECA^{+}18, HLL^{+}19, HXY13, MSA^{+}19, PSW^{+}19, WZF^{+}19, YHL^{+}19, YZWG18]. jointly [HMZ18]. JPEG [HSP^{+}13, QMCX19]. July [Ano19o]. June [Ano19p]. JXTA [AMHJ10].

K-LZF [JJH19]. K-Tree [WTG^{+}19]. Kalman [WLL^{+}19a]. KEcruiser [TMJH19]. Keeping [KCCL18, MPC^{+}18, SSC^{+}19]. Kernel [CAPG18, JP18, LLSL18, LZY^{+}16, PRW14, TMJH19, vW19]. kernels [CRYG18]. Key [ABB19b, BBvdB^{+}11, GLB^{+}18, PDH18, VCDK18, WSQ^{+}16, ZZH^{+}16, AQRH^{+}18, APK^{+}18, CRRC18, FHZW18, FNA12, GAYTC18, GMdFPLC17, GZL^{+}18, HAAWH^{+}18, IOV^{+}18, JSMG18, KLW^{+}16, LLH^{+}17, LYY^{+}18, LZYC13, MLC^{+}18a, ODK^{+}17, OSANAM19, PSLZ18, Wan19, WDKV19, WZ18, WHZ19, XZ14b, XYML19, YZL^{+}18, ZXW^{+}18, ZXWA18]. Key-aggregate [GLB^{+}18, Wan19]. key-agreement [APK^{+}18]. key-delegation [JSMG18]. key-value [GAYTC18, WHZ19]. keyless [ZXW^{+}18]. keynote [CCRL18]. keys [LH13b, LLL^{+}18]. keystroke [FZW^{+}18, MCRB19, MR19]. keyword [CZZ^{+}18, DLLZ17, HZL^{+}19, LXK^{+}14, LWZ^{+}19a, TDBR18, WHMO13, WXLY16, XTZ^{+}19, YD18, YQZ^{+}19, ZSZ14]. KID [HMMW19]. kind [WTTH19]. kinds [WGM15]. Kinect [AKG^{+}17]. Kinect-based [AKG^{+}17]. kleptographically [WLGL19]. kleptographically-secure [WLGL19]. Knapsack [WH19, HXWW18, OB19]. knee [AMAY19]. Knowle [XWL^{+}15]. Knowledge [BBCN18, NS19, ZSP17, ZS16, ABTF16, BBC^{+}17, BCdV^{+}19, DMPP16, FFPS10, GZS14, GRMSOG18, LLS^{+}19, LvW14, LTZ15, MED16, NJ19, PMT10, SHS^{+}19, Śle14, TCC18, TNY17, TF17, WWH^{+}19, XLL^{+}19c, YPHZ14, NJ18]. knowledge-based [TNY17]. Knowledge-infused [ZSP17]. Korea [HPP^{+}18]. KP [HQZH14]. KP-ABE [HQZH14]. KRAS [JLC^{+}20, YWG^{+}20, YWG^{+}19]. Kronecker [BKB18a]. Kruskal [MSA^{+}19]. Kuaa [dOWdAS^{+}18].

L [TSAER18]. L-RNN [TSAER18]. L2P2 [SLL^{+}17]. label [GLVC18, SLL^{+}17, XHL^{+}19]. labeled [LWT18]. labeling [CFM19]. laboratories [BOWD^{+}19]. laboratory [BDZ13, CGL^{+}10, ZDL^{+}13, SVN^{+}10a]. LACO [AMSPL19]. LADRA [LWR^{+}19]. LAG [DW11]. lambda [GIK18, PT16, GAMC19]. land [LLN^{+}18]. landing [TYWZ18]. Landmark [RCM17]. Language [DZLA19, EHT10, EP12, JSZ^{+}19, SDWS13, Sun10, XLL^{+}18a]. LANs [HMW14]. Laplacian [DGA18]. Large [CGJ^{+}10, EGVT18, FQBCF15, GSV^{+}10, LKM14, PB17, AMBB18, AR15, ASAAM^{+}19, ARP14, AB19c, AQB15, BC15, BR19, BAPS14, BKB18a, BCD^{+}18, CEP19a, CZT^{+}15, CZY^{+}18, CRVZ15, CA13, CA15b, CLY14, CZXL18, CWJ^{+}18b, CGM^{+}18, CSP13, DEL19, DC19, DKJ19, DDB14, DPL14, FDPR17, FPGK18, FWB13a, GLNT13, GRMSOG18, GLVC18, GIM16, HB19, HST^{+}18, HLCL16, HZ10, IPG^{+}18, IS18, JRJ^{+}11, JHC10, JTBS15, KKI14, KAS^{+}18, LTN10, LSZ^{+}16, LYW^{+}18a, LM12, MPR^{+}16, MMVS19, NBB18, NS10, NJKH13, PPZ12, PLA18, PGCC^{+}10, PF17, PPLL17, PKP19, STMV18, SJL^{+}18, TDFZ18, TJWS10, TY11, THT12, WSZH18, WTG^{+}14, WLLF16, WWH^{+}17, WRCC17, WCH^{+}18, WDD18, WCW18, WY19, WWG^{+}19b, WBJM14, WS10, XWL^{+}15, YDK11, YDT19, ZG19, ZAB15, ZWW^{+}13, ZSFZ19, ZW10, ZYTC15, ZXL14, ZA14, dSK^{+}19]. large-dimension [ZAB15]. Large-Scale [EGVT18, FQBCF15, GSV^{+}10, LKM14, PB17, BAPS14, BCD^{+}18, CZT^{+}15, CZY^{+}18, CRVZ15, CA13, CA15b, CZXL18, CGM^{+}18, DEL19, DPL14, FDPR17, FWB13a, GLVC18, HB19, HST^{+}18, HLCL16, HZ10, IPG^{+}18, IS18, JTBS15, LTN10, LSZ^{+}16, MPR^{+}16, MMVS19, NS10, PPZ12, PLA18, PF17, PPLL17, TJWS10, TY11, THT12, WRCC17, WCH^{+}18, WDD18, WCW18, WY19, WWG^{+}19b, WBJM14, WS10, ZSFZ19, ZW10, ZYTC15, ZXL14, ZA14, dSK^{+}19]. large-scale-application [TDFZ18]. Lark [ZBCT17]. Late [MLBS11]. Latency [AS19b, MHC14, CW13a, FFL^{+}19, GSC^{+}19, HSP^{+}13, KIS11, SLL^{+}18, WPGN^{+}18, WWZZ18, ZZSZ18]. Latency-aware [AS19b]. latency-sensitive [CW13a]. latterly [AAYL19]. lattice [CJS19, ZaTZ^{+}17, ZDW^{+}18]. lattice-based [CJS19, ZaTZ^{+}17]. lattices [ECE^{+}19]. law [QCZH19, XLL^{+}19c]. Layer [SVK19, AKCY^{+}17, AAM^{+}19, BGNI19, FNA11, GZW18, HDA^{+}19, JBM^{+}18, KSK^{+}19, LXMW15, LWW^{+}16, MYK16, NWT19, SCS^{+}18, TZL^{+}18, TJ18, VL19, WPGN^{+}18, YJL^{+}19, ZN12, ZFY18, ZJW^{+}14, dCRL^{+}19, dSBN19]. layered [AMQS^{+}19, PDK10, PSBB15, SSG19, SKF^{+}11, SBK18, WJS^{+}18, KJFS12]. LayerMover [ZFY18]. layout [KR14]. lazy [TV16]. LBBA [SC16]. LBBSRT [ZFC18, ZFC17]. LBS [SLL^{+}17]. LDA [ZZJY16]. LDA-based [ZZJY16]. Leading [KZ17]. Leading-edge [KZ17]. Leakage [Wan18a, DLZ16, LZL^{+}19a, YAX^{+}18]. leakage-resilient [DLZ16, YAX^{+}18]. leaking [AMRM18]. leaping [LCW^{+}18]. learners [GAA19]. Learning [AAM^{+}19, BCMM18, CLL^{+}18a, IdAP19, LRJG19, MHdlS19, MGL^{+}18, RT16, ZTC^{+}19, ZWMC19, AD18, AAN^{+}18, ABDH19, ASAAM^{+}19, AMKM18, ASY^{+}18, AMBC19, BRL19, BDA19, BTP19, BW19, BMP^{+}16, CMEA^{+}19, CFMC19, CMT16, CJN^{+}15, CMI^{+}19, CLL^{+}14, CAL^{+}18, CRWZ19, DRC^{+}19, DFG^{+}19, DA16, DGR^{+}19, DC18a, DGCGH^{+}17, ECPF17a, FSV^{+}19, FFGP^{+}19, GLC19, GHGP19, GPJA^{+}14, GHP^{+}18, GLD^{+}19b, HSV^{+}17, HUMA18, HPGMM18, HX19, HXL^{+}18, HHS^{+}18, JSZ^{+}19, JOSD19, KMK^{+}19, KI19, KAP19, KMI11, KKS18b, LHJC18, LLH^{+}17, LLY^{+}18, LJW^{+}19a, LXT^{+}19, LYYW19, LJW^{+}19b, LZWF19, MFSV19, NK18, NGB18, NAAC19, NAM^{+}19, NLV^{+}19, PZC19, PBL^{+}18, PWWD18, QCY^{+}19, QCZH19, RGAT18, RSY^{+}18, RHH^{+}19, RGM^{+}19, SD18, SI18, SRP19, SYT^{+}19, SNC18, dSSCdL19, SZD^{+}17, SR19, SHL^{+}19b, TNY17, TYWZ18, VRGR16, WZF^{+}19, dOWdAS^{+}18, WLB11, WXZL11, XZZ^{+}14, XHW19, XKBA18, XLL^{+}18a, XLL^{+}14, YYW^{+}19, YWLL19, ZGV19]. learning [ZXM^{+}19, ZZ19, ZSQ^{+}19, ZYC^{+}19, ZSL^{+}19b, dlFVPSHL^{+}14, ZBF14]. Learning-based [ZWMC19, HHS^{+}18, NAAC19]. leave [KMJ18]. left [DNW^{+}19]. legacy [DW11]. legal [CCIP18]. legal-rule [CCIP18]. LEGIoT [MPLM18]. length [LZL^{+}12]. lens [CTU19]. Let [MOBD18, BKY18]. Level [LYW^{+}16, AD18, AFS16, AJY12, AMR18, dRADFG18, CMS^{+}18, CXC^{+}18, Ciu10a, CCD^{+}10, CHY^{+}18, EMHE18, FG14, FWB13a, GOBL16, GJF^{+}12, GLXF17, JLI^{+}13, KKK^{+}19, KMK^{+}14, LA19, LBB^{+}19, LSG^{+}19, MZP^{+}19, MTD18, MJRM16, MLBS11, NAAC19, PLLA18, PRS12, QPTGG^{+}12, RZ16, RPMG10, RS17b, RMHCMG15, SRZD15, SVN^{+}10a, SEMJ11, SK18, SSL13, TSBH11, WLYL11, WWC14, WZWC18, WDR^{+}19, YDT19, ZLZ13, ZCK^{+}15, ZZH^{+}16]. levels [Kim18, LYJ10, LWT18, TTB^{+}13, UDST19]. Leveraging [CFM19, CFMC19, LLF^{+}18a, DW11, DLZ^{+}14, HXL^{+}18]. LHC [RWV^{+}13]. libraries [DSS19]. library [AFS16, JLRS18]. licensed [SLH^{+}19]. licenses [CMZ^{+}12]. life [AFO^{+}18, CBBC^{+}17, CRWZ19, GCBM17, GPS^{+}17, JOPW14, LWZ^{+}19a, ML19, RCW^{+}19, SFR15, SCJ^{+}19a, WOPW13, ML19]. life-long [ML19]. life-threatening [AFO^{+}18]. lifecycle [KAS^{+}18, LYW^{+}16, SPS18]. Lifetime [CXZ^{+}19, PLZX19, KCM19]. Lifetime-aware [CXZ^{+}19, PLZX19]. light [AKB^{+}18a, Eng14, LLZ^{+}18b, LLU^{+}18, PCK19, SJR13, YSL19, ZWX^{+}19]. light-based [SJR13]. light-weight [Eng14]. lighten [ZDM^{+}19]. lightpaths [CGD10]. Lightweight [AMSPL19, GAI^{+}18, HKT^{+}19, MPLM18, ZLY^{+}19, AMN18, AMKC19, BLL^{+}19, CMS^{+}18, CRRC18, DL19, HZL^{+}19, HZ10, MCN^{+}18, MHY^{+}18, MLBS11, OSANAM19, dRRRR^{+}18, SCS^{+}18, WSL^{+}19, WLS^{+}18, YCT15, ZZX^{+}19, ZZY^{+}19, ZSW^{+}18a]. limb [KPG19]. limitations [DGY^{+}18, NS17b]. limited [JLCC12, YDQC19]. limits [LN13]. line [BMR15, CCZ^{+}19, CCDP19, LLW^{+}19c, TJ18]. linear [ADAAD12, BCdV^{+}19, BDNP13, CZT^{+}15, CLY14, GGA^{+}17, JLRS18, JLQ18, TGM11, WWX^{+}17, YPCK12, ZZC18]. lines [MdOO^{+}17]. lingual [CGZL19]. linguistic [CPE^{+}17, ECPF17b]. Link [GCCPGBGS10, HQ10, KP18, LTC12, Sun10, WFQ^{+}10, Zhu10, AK18a, BCF16, CGM^{+}18, GGC17, GXL^{+}18, MDB^{+}18b, NLLC19, RRU^{+}18, SMG18, SRP19, SGS^{+}18, XWL^{+}15, XHL^{+}19, YCZJ18, ZGL19, ZS10]. linkable [ZLL^{+}19]. linkage [TTC^{+}14]. Linked [CDH^{+}19, DMMM11, Ans11, BBD^{+}13, CPSRG14, Li10, CAS^{+}16, PVHTP19]. linking [GTM19]. links [ACD^{+}19, CWJD19, KCM19, WLRL18]. LiReK [CRRC18]. LISA [PCK19]. list [BBI13, DNP14, LY18a, YK17, YNLY19]. list-based [DNP14]. literature [ABP18, AMS19, MND^{+}19, MCWP16, SDWS13]. Live [DK17, FS19, JDW^{+}14, AS14, CFMC19, HZW^{+}16, JFZL17, KO11, KSK^{+}11, LJLW13, LRZ^{+}18, LFHQ19, MYW^{+}19, MCAS19, SLA^{+}16, TCN^{+}16, YMD^{+}13]. live-broadcasting [LRZ^{+}18]. live-streaming [MCAS19]. liveness [GAFFOG12]. liver [wZcZN^{+}19, wZcZN^{+}20]. living [FKT14, SJSA19, TMB^{+}19, VKT^{+}19]. load [AA18, AB19c, AS18a, BBMG10, CT19a, CRC13, DLS^{+}12, DXL^{+}18, GOBL16, HLW12, IS18, KRZ12, KIC12, LZXW13, LLW^{+}18a, LN18, MKM11, PAL^{+}19, PRC^{+}14, PBC^{+}16, PBC^{+}17, QCY^{+}19, SDTA19, SL11, SMG18, SJTN18, SYL18, TDC^{+}14, TZLL18, VVB15, ZGB^{+}17, ZGL^{+}18, ZFC17, ZFC18]. load-aware [KIC12, ZGL^{+}18]. load-balance [DXL^{+}18]. load-balanced [SDTA19]. load-balancing [AA18, LZXW13, SJTN18]. loading [HSC15]. loads [NJKF18]. loan [WGC19]. LoC [WGC19]. Local [LBB^{+}19, PSS^{+}18, ABMESM18, BCdV^{+}19, FMRS18, GJ18, Hua10, HLL12, KMT14, LZL^{+}12, PM14, RCMT18, SWW^{+}13, WCM^{+}19, WH19, XZW^{+}19]. Locality [HSC15, SDTA19, ZXL^{+}18, HNKÖ18, JLD^{+}19, NNRA19, QZD^{+}18, SCCS11, USK16, WWC14, WWQ^{+}18]. Locality-aware [SDTA19, SCCS11, WWQ^{+}18]. locality-sensitive [QZD^{+}18]. Localization [MWL18a, DdSdN^{+}19, GAI^{+}18, HHH^{+}19, HXY13, SYJ^{+}19b, SJL^{+}17, WWSL19]. locally [BRL19]. Location [AHEM17, CT19c, DA18, LLW^{+}19a, LNLA19, DGY^{+}18, FWB13b, GCCL18, GBKJ18, HZW^{+}18, HHW^{+}19, HMC19, MQL^{+}19, MVCC10, NWD^{+}18, NZL^{+}15, PLW^{+}19, SLL^{+}17, WZE19, XCS^{+}18, XCZ^{+}19, YXA^{+}18, YZL^{+}19, ZCLW18, ZLT^{+}19, ZWS^{+}12, SCH^{+}17]. Location-based [CT19c, DA18, LNLA19, HMC19, MQL^{+}19, NZL^{+}15, PLW^{+}19, XCZ^{+}19, ZCLW18, ZLT^{+}19]. location-label [SLL^{+}17]. location-sharing [XCS^{+}18]. locations [Alp18, dlFVPSHL^{+}14]. LOFAR [BBB^{+}11]. Log [LWR^{+}19]. Log-based [LWR^{+}19]. logging [EH10, LBB^{+}19, LM12]. Logic [KB16, GBY16, LLJ^{+}11, LvW14, MYHZ18, MLGGB^{+}17, OMD^{+}18, SYJ^{+}19b, SZR18, TWW^{+}18]. Logic-based [KB16]. Logistic [RRKA19]. logistical [Zhu18]. logistics [ATH^{+}19, YLL^{+}19]. logs [PKB19, WLZ^{+}19]. Long [RSK16, DQXW19, DLS^{+}12, HKU^{+}11, HKG^{+}16, JLS19, ML19, PM14, WWD^{+}14]. long-distance [HKU^{+}11]. Long-haul [RSK16]. long-range [PM14]. long-term [HKG^{+}16]. look [WYBS11, YH18]. look-ahead [WYBS11]. Loom [BCD^{+}18]. loop [EKGS14, LTC12]. loop-free [LTC12]. loops [ACD^{+}19]. Lop [RGDML16]. LoRa [DBP19]. loss [JCL^{+}19, KDHP16, RMDB18]. losses [DJJ^{+}18]. lossless [CCD^{+}19, HDB18]. lossy [ASO14]. Low [ABP16, GVURIVBV14, KWK16, LLU^{+}18, MMC^{+}18, RS17b, SLL^{+}18, BTG19, BSE^{+}13, DMC^{+}19, FFL^{+}19, GSC^{+}19, HSP^{+}13, JCMPPC^{+}18, LA19, LEW19, LYL^{+}19, LZHY19, SPT^{+}18, TLC^{+}15, TWW^{+}18, VVC^{+}12, WPGN^{+}18, WWG19a, vdLLE19]. low-cost [BSE^{+}13, LEW19, LYL^{+}19]. low-footprint [VVC^{+}12]. Low-latency [SLL^{+}18, FFL^{+}19, GSC^{+}19, HSP^{+}13, WPGN^{+}18]. low-level [LA19]. Low-power [MMC^{+}18, DMC^{+}19, WWG19a, vdLLE19]. low-resource [TLC^{+}15]. Low-time [ABP16]. LP [SISGS18]. LPCMsim [HKT^{+}19]. LPTD [ZZX^{+}19]. LR [GHMX10]. LR-WPANs [GHMX10]. LSTM [YGY^{+}19, ZDM^{+}19]. LSTM-EFG [YGY^{+}19]. LTE [GSP^{+}17, VL19]. lucrative [SH19]. lung [HLZ^{+}19, JLC^{+}20, KMK^{+}19, LYYW19, YWG^{+}19, YWG^{+}20]. LZF [JJH19]. LZSS [OSC14].

m [ZCDV19]. m-healthcare [ZCDV19]. MAC [CLL^{+}18a, CJG^{+}18]. Machine [BCMM18, DGR^{+}19, LY19, MHdlS19, MGL^{+}18, NK18, XHW19, XWW19, ZTC^{+}19, AD18, AAN^{+}18, AD19, ASY^{+}18, AGKZ18, AMBC19, BDA19, CFMC19, CFVP12, CMI^{+}19, CLY14, CGZL19, dCCDFdO15, DRC^{+}19, DPBK16, DA16, FSV^{+}19, GHP^{+}18, GLD^{+}19b, HSV^{+}17, HPGMM18, HSC15, HXL^{+}18, HAA^{+}16, HHS^{+}18, JNR12, JOSD19, KI19, KCS14, LJS17, LYYY18, LLWW18, LLY^{+}18, LFHQ19, LXRS19, LLZ^{+}19, LJW^{+}19a, LSW^{+}19, LPBB^{+}18, LZY^{+}16, LJW^{+}19b, NAAC19, NAM^{+}19, PZC19, PFPJ18, Pon19, RHH^{+}19, RGM^{+}19, RCTY19, SI18, SHLJ13, SNC18, dSSCdL19, SBP^{+}17, VVB13b, dOWdAS^{+}18, XJWW15, YLHJ14, YPLZ17, ZXW19, ZYZ^{+}18, ZHHC17, ZFY18, ZGV19, ZXM^{+}19, ZLL^{+}16, ZSL^{+}19b]. machine-based [LZY^{+}16]. Machine-Learning [MHdlS19, DA16, GHP^{+}18, JOSD19]. Machines [DK17, LYYY17, AS19a, ADA^{+}19, AAM^{+}16, AS14, CHS^{+}18, DEG^{+}17, DQLW15, DCMW17, GVI13, HMH17, HZZ^{+}14, JLRS18, JLL17, JDW^{+}14, KSSG16, LLF^{+}18a, LJL12, LC13, Man15, MK19b, PRC^{+}14, SS13, SDDG17, SLA^{+}16, TMMVL12, WQG15, WHCZ18, XWX^{+}17, ZWHC17]. Macular [TBS^{+}18]. magnetic [HZX^{+}19, HZX^{+}20, RDSA18]. Main [BLMU19, HX19, TTC^{+}14]. main-belt [TTC^{+}14]. maintain [Mat18]. maintenance [CPLH19, DPS16, LWTL19a, ZH17]. Major [HSS17]. make [SI19, TMM^{+}13]. Maker [JBR^{+}16]. makes [MMC^{+}18]. Makespan [JZWL17, JLL17, WYBS11, ZZS^{+}19]. Making [RRP^{+}14, ZDW^{+}16, APR^{+}19, JXZ^{+}19, KK10a, KKS^{+}18a]. malapps [WWH^{+}17]. Malicious [YJL^{+}19, BTP19, LLQS14, LMM19, MRL14, PZC19, WLW^{+}18]. malware [AD18, HDKC18, HLH^{+}18, HAA^{+}16, HIA^{+}18c, HAAR^{+}19, MHdlS19, MAY18, QKC19, ZRZR19]. mammals [WMBV17]. mammograms [SYT^{+}19]. mammographic [BMU18, FMV14]. man [JCL^{+}19, LSW^{+}19]. man-in-the-middle [JCL^{+}19]. man-machine [LSW^{+}19]. manage [CCMP18, CdRRdCB19, EAS^{+}18, MAB^{+}15, OAMS18, SFR15, TMS^{+}17]. managed [CRM^{+}16, LHPC^{+}19, NJHT11]. Management [AR10, ABB19b, CDH^{+}19, DP20, DP21a, DP21b, GS16a, SE19, WXYL15, AB19a, ABGMC19, AB21, ANG^{+}19, AAB^{+}10, ACHP19, AMPS19, AEME^{+}18, ACC^{+}16, AK18b, BAJ^{+}19, BLO^{+}18, BBWB^{+}18, BFP18, BCF^{+}10, BAB12, BN17, BDH14, BCD^{+}18, BRNR15, BR10, CMZ^{+}12, CMEA^{+}19, CXDM18, CPD^{+}15, CFG^{+}19, CLL18b, CLM^{+}16, CYZK15, CLS19a, CLDC19, CAB^{+}18, CS19, CR14, DVJ^{+}15, DMPP16, DP19, DYI^{+}16, DSPA18, EG18, ECPF17b, FA11a, FG14, FM17, FTH16, GCM^{+}11, GLM^{+}12, GEG14, GBS10, GKW^{+}12, GCK18, GBF^{+}12, GGDM^{+}18, GHJ^{+}19, HAP15, HMM18, HCJ14, HST^{+}18, HZP^{+}14, HHS^{+}18, JBM^{+}18, KSF^{+}13, KRZ^{+}19, KMI11, KADJ14, KS17a, KGS^{+}19, Kol18, KH18b, KARP14, KAS^{+}18, KVHT10, LKN^{+}13, LMZ^{+}14, LLW^{+}12b, LKK^{+}16, LYW^{+}16, MJGW18, MPCAF15, MBS13, MSBA16, MED16, MKRD19, NJH^{+}18, NHH^{+}19, NSSA^{+}14, NJKH13, OB17, ÖE13]. management [PVGD^{+}19, PSJ^{+}12, PRS^{+}13, PB17, PLCGS11, PMT10, PF17, PPB16, PPA18, QLM^{+}18, RGSL18, RMVG^{+}10, RM11, SMS14a, SH19, SCL14, SPS18, SAK^{+}10, SSW^{+}19, SCJ^{+}19a, SCH^{+}19, TOD17, TTB^{+}13, TF17, TSGVRGS19, TAHS14, TY11, TCBPR16, URKM19, VAdlP12, VDK12, VCDK18, WW11, WCF^{+}15, WMY^{+}18, WGC19, WDKV19, WLA18a, WZS^{+}18, XLL^{+}14, XLL18b, XAW^{+}10, YMLT13, YZL^{+}18, YCD^{+}19, YJL^{+}19, YMD^{+}13, ZCM19, ZWW^{+}13, ZCS^{+}16, ZYK17, ZLR^{+}15, ZWGC19, dSFP^{+}17, dSGD19, dSFD^{+}19, dCTVC18, SJV^{+}15]. manager [KMT14]. managers [GDR^{+}14]. Managing [BBCN18, DGGH11, HLNM11, HKP10, MBMTJR18, SDF^{+}19, ZG19, CFGM16, LRJG19, SG14, TBB^{+}17, YLHJ14]. Mandatory [BL13]. MANET [AdI14]. MANETs [ZYW^{+}18]. maneuvering [BRL19, LWX13]. Manifestations [PKP19]. manifold [XLL^{+}14]. manipulate [VOCHC17]. manner [kHsZwJW18]. manufacturing [HMW^{+}19, HCX^{+}19, SSST17, SSW^{+}19, XWjZyF19]. Many [HYZS16, CLH10, EDH^{+}13, JLY^{+}18, LC14, MAC14, MCA^{+}18, MGMT18, YLJ^{+}17, YDT19, ZAB15, ZCL^{+}14, ZLG^{+}14]. Many-core [HYZS16, EDH^{+}13, JLY^{+}18, LC14, MAC14, MCA^{+}18, MGMT18, YLJ^{+}17, YDT19, ZAB15, ZLG^{+}14]. many-task [MGMT18]. manycore [HTL^{+}18, JPB17, Li18]. map [Bro19, HX19, LLW^{+}18a, LWK^{+}18, MZL^{+}19, SCCS11, STA17a]. map-based [LWK^{+}18, MZL^{+}19]. Mapper [MLC18b, MSS^{+}16]. Mapping [MRT^{+}19, Mil11, MFT^{+}17, YG18, AUSA19, DSK^{+}14, DKV14, DST10, DST14, HSC15, LvW14, MBM18, MEBA12, YWZ^{+}18]. MApReduce [FDGR14, ACK^{+}15, DDJ^{+}13, DFGR14, GAMC19, GZWQ13, HSC15, JGFB18, JZWL17, JS13, LKM14, LQK^{+}16, LLC^{+}16, LSZ^{+}16, LC14, LLAH13, LSJ^{+}14, MNV12, MLC18b, NNRA19, SEMJ11, SDTA19, SB17a, SJV12, VETT16, WTR^{+}13, XX14, XTT18, YWF^{+}10, ZCK^{+}15, ZFS^{+}18, ZWZ18]. MapReduce-based [LLC^{+}16, LSZ^{+}16, SB17a, YWF^{+}10]. Maps [AS18b, Che18, DH16, DMM14, KLW^{+}16]. March [BFS^{+}17a]. margin [CZ12]. MARIANE [FDGR14]. Marine [YWZ^{+}18, FCD^{+}14, WMBV17]. marital [MNC^{+}18]. Market [JBR^{+}16, JSS^{+}12, Lin18, RT16, RA12, SSJ19, SCMS12, WBR19]. market-based [SCMS12]. marketplace [JBR^{+}16, MVG^{+}14, VMSRM12]. marketplaces [ABH18]. markets [BAB13, MG16, PRS12, VPT^{+}10]. Markov [CLL18b, LvW14, LW18b, WQG15]. Mashup [PC18a]. mashups [dVXB^{+}11]. MASi [GHJ^{+}19]. masking [BMU18]. mass [KVK^{+}18, SYT^{+}19]. mass-gatherings [KVK^{+}18]. masses [BR18, BMU18]. massive [DGD^{+}16, FGM11, LXL^{+}17, NS17a, SGB^{+}18, WMLS14, YYS^{+}19, ZRZR19]. massively [NOF18, PDK10]. Master [LC14, PCC18]. Master-worker [LC14]. Matching [LZP^{+}18, TOS18, AHYF19, FLT17, GWW^{+}19, LLC^{+}16, LLAW17, MWQ^{+}14, NWMG17, PRN14, QGX18, TJ18, WHCZ18, XWjZyF19, YZL^{+}18, YZWG18, YSZW18]. Matchmaking [DT16, CCD^{+}10, LCBF13, LDS^{+}18, LZJL19, YAO14]. MatchTree [LCBF13]. materials [HLCL16, MZD^{+}16]. maternal [APR^{+}19]. matrices [WHCZ18]. Matrix [XZ14b, GGMS18, Jun18, KR14, LXM^{+}18, LGZY18, SJL^{+}17, VS19, WYL^{+}18, ZWL^{+}16, ZCH^{+}17, ZZC18]. Matrix-based [XZ14b]. matrix-vector [ZWL^{+}16]. matroids [YH18]. matters [PD11]. maximal [LHW^{+}18, SLD^{+}18]. maximization [HMC19, KCM19, MG14, NBB18, TL19, WDD18]. maximize [YDQC19]. Maximizing [MFL18, SMS14b, WJLW18]. maximum [CZ12]. May [Ano19r]. MC1 [LPS19]. MCA [SAH19]. MCA-V2I [SAH19]. MCDM [HFL^{+}19, SMZ^{+}16]. MCMC [LvW14]. MCS [FY19]. MCS-Chain [FY19]. MDP [MRS18b]. MDP-IoT [MRS18b]. mdtmFTP [ZWDP18]. meager [DV13]. meal [MOFGP18]. mean [CGIP14, HKP10]. means [AMHJ10, Bu18, DNW^{+}19, LSZ^{+}16, LZY^{+}19b, PT16]. measure [RAKJ18]. measurement [BDZ13, CZT^{+}15, GSY^{+}17, LXRS19, LRC^{+}18, WLRL18, XLL^{+}19c]. Measurements [CSC18]. measures [DLDTGMMP16, GJY18, LG18]. Measuring [HRVW18, KdGP^{+}19, PARMF14, VLAC^{+}13, dSCD^{+}19, vKLA^{+}19]. MEC [SISGS18]. Mechanism [ABB19b, WDW^{+}19, AR15, ABF^{+}15b, AZO^{+}19, CSJ^{+}17, CWL^{+}19, CB10, CWLZ19, DWJM18, DLS^{+}12, EK11, ECPF17b, GA13, GZW18, HLW12, HAP15, HDLW13, KKN18, LZCX19, LJY12, LXMW15, LHX^{+}18, LQLX10, LLGY18, LLW^{+}19d, MVL18b, PWWD18, RYH^{+}19, SYJ^{+}19b, SMG18, SJTN18, SSL^{+}19, SOM^{+}19, TGM^{+}19a, TMP15, TAKV12, WZE19, XDHL12, YLJL18, YXA^{+}16, ZA13, ZXZL18, ZCDV19, ZA14]. mechanisms [BDZ13, CLAL19, CCCT14, DC18b, GHYK18, KRZ^{+}19, KKW^{+}14, PLGMCdF18, PMFH11, SGN^{+}17, SSB13, TMMVL12, WRCC17, YNY^{+}14]. Mechatronics [CJN^{+}15]. MECOM [JDW^{+}14]. Media [MLC^{+}11, MPI^{+}18, BTP19, DVD12, GHO^{+}11, GTM19, LLG^{+}16, MCG^{+}15, NO19, WSN18, WdL16, WLA18b, YXZG18b, ZG18]. Media-based [MPI^{+}18]. median [KPS18]. mediated [ZGS^{+}13]. mediation [RVST17]. mediator [ABTF16]. mediators [BGNI19]. Medical [BDM^{+}19, CT19a, CMI^{+}19, GLD^{+}19b, sGbKS19, JYZ^{+}19, MZYA19, OM10, AIA^{+}18a, AIB^{+}18, AM19a, DSM^{+}19, KVHT10, LYYW19, LHBC16, PMBS14, PSW^{+}19, SB17b, SPS18, SZG^{+}19, TIHT14, VSKS19, VFHB14, WLS^{+}18, WPS^{+}18, XPL19, YCY10, YSC^{+}15, YLN15, YWLL19, ZCH^{+}17, AMKC19]. medicine [BJM^{+}17, FFC^{+}18, LWK^{+}18]. medium [BAKB19, GHMX10]. medium-sized [BAKB19]. meet [BR18]. meets [CCT13, MGR11, PYH^{+}18]. member [SCZ^{+}19]. membership [GNVST14, PF17]. MemEFS [UDvdW^{+}18]. Memory [DLZ16, HKT^{+}19, MWQ^{+}19, BYL^{+}18, BEWZ10, CPSD18, DKK^{+}13, DKJ19, FJL^{+}16, GQLX18, GXW^{+}19, GVI13, HHXL13, HZDS19, HMMW19, JLS19, JDW^{+}14, KP12, LAL^{+}14, LLWW18, LFHQ19, LYW^{+}18b, LBU^{+}10, MAC14, MCdA16, RKB18, SEPV19, USK16, UDvdW^{+}18, WHZL10, WZS^{+}18]. memory-based [MCdA16]. memory-performance [SEPV19]. MEnSuS [AK18b]. merging [CPGdS^{+}13, TSOB15]. mesh [DCBF19, XZ14b]. meshes [ZGV19]. Mesos [LHPC^{+}19]. Message [WHZL10, CJXX19, FFL^{+}19, GSC^{+}19, LC17, LBB^{+}19, NWL17, ROK19, RMDB18, SL19, ZA14, LM12]. Meta [CBS17, ESW^{+}17, PT16, PB17, TCCC11, WZH^{+}19, XL19, ZSQ^{+}19]. meta-analysis [WZH^{+}19]. meta-database [PB17]. meta-heuristic [ESW^{+}17]. meta-learning [ZSQ^{+}19]. Meta-model [CBS17]. meta-scheduling [PT16, TCCC11]. Metabolic [DDJ^{+}13]. metacomputing [BBMG10]. metadata [GHJ^{+}19, LFH^{+}15, MRT^{+}19, Mil11, YMW13]. metadata-driven [GHJ^{+}19]. metagenomics [DMC^{+}19, MMC^{+}18]. metaheuristic [TF18]. Metal [SPSP17]. Metal-as-a-service [SPSP17]. metamodel [GHGP19, WBKL16]. metascheduler [TCR^{+}12, VHML10]. Meteor [ZHW19]. meteorological [HCMJ19, LHJC18]. metered [YVCB10]. metering [JLC18, RAdARP19, TAS^{+}18]. meters [MK17]. Method [KPG19, AFSH^{+}19, AIA^{+}18a, BLL^{+}19, CZXL18, CYW^{+}19, DHL18, DCC13, DXL^{+}18, DZLA19, EMM12, FJ18, FW19, FWB13a, GAFFOG12, GOBL16, GPS13, GLXF17, HCMJ19, HNKÖ18, IAL10, JLS19, KIS11, LZXW13, LLC^{+}16, LYYY17, LCW^{+}18, LYYY18, LLW^{+}18a, LXRS19, LJ19b, LZL^{+}12, LZX16, LZW^{+}18, LXT^{+}19, LZL19b, LSW^{+}19, MQL^{+}19, OA17, PdASM18, PNZ14, PVHTP19, PDK10, SMC18, SWW^{+}13, SOIS12, dSSCdL19, SAG19, SYQ^{+}19, TDC^{+}14, TMDZ15, TSS^{+}19, TYWZ18, VS19, VOCHC17, WCM^{+}19, XZZ^{+}19, XLL^{+}19a, XXQ^{+}19, XJZ^{+}19, YLHJ14, YJS18, YJY^{+}18, ZXW19, ZZBP19, ZDL^{+}19, ZCZ^{+}18, ZZZ17, Zhu18]. methodical [MS19]. methodological [CVT19, CGM^{+}19]. Methodology [Ham17, ACMM19, KKS18b, LCY19a, MSBA16, MED16, TF17]. Methods [ALFR16, AAC^{+}19, AMM16, BMP^{+}16, CXZC18, CLR18, Cuz14, FLT17, GK18, KVvE18, LY19, MLG13, MG19, MYBMM18, PRW14, SPR^{+}10, XA10, ZN12, dLB10]. metocean [KNI^{+}18]. metric [LYW^{+}18a, LWSY18, PSL19]. metric-based [LYW^{+}18a]. metrics [AdVAGF18, BBI13, CdSDS15, GJF^{+}12, LCGPC19, SPT^{+}18, WMA^{+}19]. Metropolitan [PP10]. MGPV [VGD^{+}19]. MiCADO [KKK^{+}19]. micro [ZWWL18]. micro-blog [ZWWL18]. microarchitectural [VEET18]. microarray [CZY^{+}19]. microblog [ZZJY16]. microblogging [CCJ16, DCF19, HO17, MNC^{+}18]. microgrids [RWZ^{+}19]. microscopic [Wei11]. microservice [MFC^{+}19, ŠCJ^{+}19b, KKK^{+}19]. microservice-based [ŠCJ^{+}19b, KKK^{+}19]. microservices [KFK19]. Microsoft [QC13]. MID [GLXF17]. middle [JCL^{+}19]. Middleware [BS17, GLXF17, KT17, PZY16, SCL18, ARSMY19, AMPZ16, BGNI19, FTP14, FKT14, GACM17, GVBG17, GTMZ17, KK10a, KKJJ10, LGW^{+}17, MKM11, MWL^{+}18b, NJHT11, PZY17, TDLC17]. middleware-layer [BGNI19]. Middleware-level [GLXF17]. Midgar [SAGGB17, GMLGB^{+}17]. MidHDC [PZY16, PZY17]. MidSHM [SCL18]. MIFaaS [FMN^{+}17]. MIFIM [BS17]. MIFIM-Middleware [BS17]. MigPF [dRRdQGR^{+}18]. migrating [ALM^{+}10, GVDT16]. Migration [DK17, AS14, CPDJ13, DKV14, DPBK16, FNCR11, GGTRRC16, HH19, HZW^{+}16, JDW^{+}14, LJLW13, LWTL19b, LFHQ19, LLZ^{+}19, MYW^{+}19, PFPJ18, RCTY19, SE19, SLA^{+}16, ZMTT16, ZZJ17, ZHHC17, ZFY18, ZZ19, ZSL^{+}19a]. migrations [PK11, ZG19]. MIH [GSY^{+}19]. Milky [SVB^{+}19]. millet [DJJ^{+}18]. million [Eng14]. Mimic [LZS18]. MIMO [WCW18, YYS^{+}19]. Miner [ZSQ^{+}19]. mini [TF18]. mini-batch [TF18]. minimal [TVV13]. minimization [APAZ17, HLL^{+}17, JZWL17, LGL^{+}17, MFN13, TWG^{+}19, ZF16]. Minimizing [JLL17, ZAC^{+}18, ZZS^{+}19, DLXR14, LSCL19]. minimum [LY17]. Mining [CEP19b, CLM14b, FGM11, FZHH14, IHA18, JL14, MNC^{+}18, MRL14, XLZ^{+}14, YXZG18b, YK17, ZDW^{+}18, AMKM18, CPA14, DGD^{+}16, HS19, HXC^{+}18, KOP^{+}17, LY17, LY18a, LC15, RAKJ18, SGKC10, SSLF^{+}10, THA^{+}17, TNY17, TIHT14, WTG^{+}14, WWZ18, YLVY15, YLG^{+}16, YL16, YNLY19, ZZC14]. MIoT [BLMU19]. mismatch [MCG^{+}15]. Missing [APR^{+}19, EBOY14, TSAER18]. mission [RRU^{+}18]. mission-critical [RRU^{+}18]. Misty [AMQS^{+}19]. misunderstood [CSL19]. mitigate [SMF^{+}19]. Mitigating [LYW^{+}18b, OWX19, ASB18, TA18, WSQ^{+}18]. mitigation [HA19]. mitigations [KdGP^{+}19]. mix [LMM19, QC18]. Mixed [GGA^{+}17, SSFFR19, SZK16, CXZ^{+}19, WHW16]. mixed-criticality [CXZ^{+}19]. mixed-parallel [WHW16]. mixing [ZMH^{+}18]. Mixture [BBMG10]. ML [GGLD10]. mmWave [NLLC19]. MObile [KKA18, AR17, AT18b, BDE17, BRB19a, BRB^{+}19b, CFL^{+}15, CT19c, FG18, FMN^{+}17, FLR13, KIAD17, KRZ^{+}19, KIMR15, MAD^{+}16, NKB19, QKC19, RLM18, SVK19, TMW^{+}17, VFHB14, WZM^{+}18, WZWW18, YHL16, ZSL^{+}19a, AHS^{+}18, ANG^{+}19, ADLW12, AHU^{+}19, AB18a, AB19c, ADH^{+}16, AMGCC18, AMPZ16, AMRM18, ALL^{+}18, BJ12, BDP11a, BOP^{+}14, BN17, BCP18, CGBAP18, CWJ16, CZM^{+}18, CZZ^{+}18, CMVA18, DSD^{+}11, DdSdN^{+}19, DYC^{+}18, EAA16, EZTL19, FHZW18, FY19, GD10, GSP^{+}17, GMM18, GSY^{+}17, GGC18, GNVST14, HHK18, HMZ18, HXC^{+}18, HLT^{+}19, JLQ^{+}17, JXC^{+}19, JKLK17, JOSD19, KKKM13, KB18, KGLY18, LW18a, LLL^{+}19, LH13b, LCZR12, LXMW15, LWH^{+}18, LLJ^{+}11, LCMX16, LPL^{+}16, LWL^{+}18, LWT18, LLGY18, LZLL18b, Lok12, LFY^{+}19, LLAW17, LWZ^{+}19b, MCAS19, MKS18, ODK^{+}17, PZA18, PRS^{+}13, PKY^{+}17, QZD^{+}18, QGT^{+}18, RS17a, RCOP^{+}11, SB19a, SPJ17, SLK17, STA17a, SYK^{+}17, SOM^{+}19, TLC^{+}15]. mobile [TGM^{+}19a, TLL^{+}19, VNAMM19, VL19, VSDD13, VCDK18, VGD^{+}19, WLZ^{+}16, WWVJ17, WCL^{+}17b, WWG19a, WWTF18, WLA17b, WMJW18, XZZ^{+}14, XCS^{+}18, XXX^{+}19, YLWW18, YCXW18, YWY^{+}17, YCX18, ZF16, ZZ19, ZYC^{+}19, ZDW^{+}16, ZLL^{+}17b, SGN^{+}17, ZHL^{+}18]. Mobile-aware [ZSL^{+}19a]. mobile-IoT [AHS^{+}18]. Mobile-IoT-Federation-as-a-Service [FMN^{+}17]. Mobility [GD10, LNK^{+}18, BRB19a, BRB^{+}19b, EAED18, FAMA^{+}17, GVBG17, GCK18, GSP^{+}17, GMP^{+}17, MDB^{+}18b, MGN^{+}16, PC17, PRL^{+}19, SBP^{+}17, WZM^{+}18, WHBC19, YCX18]. Mobility-aware [GD10, YCX18]. MOBT [WLGL19]. modal [CCL19, CZ19, XPL19]. mode [JKLK17, LYS12, dSSCdL19, WMQ^{+}16, YZZC19, ZZJ17]. Model [Ans11, DWS12, EHT10, FMN^{+}17, LCH^{+}18, PSS13, RBN13, SST18, TCBC18, WHCZ18, ZCM19, AOIS10, AMAY19, AFSH^{+}18, AQAR^{+}18, ArMS19, ABTF16, AJY12, AL14, AB18a, ATdC^{+}16, AMGCC18, AK14, AMKM18, AKM18, Bal16, BAKB19, BKS^{+}18, BBvdB^{+}11, BdM11, BR19, BDMO11, BWG19, CGIP14, CIK10, CW16, CSG^{+}18, CBS17, Che14, CAC^{+}15, CCY^{+}18, CYJ19, CPLH19, CdRRdCB19, CBBdL16, DDMPG17, DNW^{+}19, DZLA19, EAS^{+}18, EMJ^{+}13, FLR^{+}16, FKOC11, FWB13b, GLC19, GS13, GZZ^{+}18, HAJ^{+}19, Ham17, HPZL18, HAM18, HZM14, HMMW19, HLZ18, HKG^{+}16, IG12, JFZL17, JCL^{+}19, JSS^{+}12, Jun18, KMK^{+}19, KYZ19, Kim14, KKP19, LHJC18, LZL^{+}17, LZLL18a, LYC^{+}19, LZT^{+}19, LLL^{+}19, LC14, LWW^{+}18, LLJ^{+}11, LDS^{+}18, LWSY18, LW18b, LZJL19, LBU^{+}10, MAC14, MWL18a, MYHZ18, MG16, MJM^{+}16, MK17, MS19, MCR^{+}16, MBA19, NSR^{+}19, NV11, NRV^{+}17, NA19, PGCC^{+}10, RRKA19, RDSA18]. model [RZ16, RR18, RWV^{+}13, RGDML16, RMCMD12, SSHC19, SMF^{+}19, SHP^{+}16, SS13, SHBP10, SL11, SZ12, SGM11, SFR15, SCL14, SKS17, TLSC17, TZL^{+}18, TMDZ15, TQL^{+}19, TGM^{+}19b, WWC14, WY17, WYH^{+}17, WWA19, XWW19, XZ16, YNSM12, YMW13, YZLQ14, YWLL19, YCH19, YXY18, YWY^{+}17, YJL^{+}19, ZZDM^{+}18, ZRZL18, ZZF^{+}19, ZLTY10, ZME^{+}15, ZZJY16, ZNC^{+}18, ZDM^{+}19, ZL18, dSGD19, CsZW14, GM11, HLCL16, ML19, Mil11, MCF^{+}11, SB11]. Model-based [LCH^{+}18, DZLA19, MJM^{+}16, NSR^{+}19, SZ12]. Model-driven [DWS12, RBN13, TCBC18, ZCM19, Ham17]. Modeling [AY16, BDWM17, CC19, DLZ^{+}14, DCC^{+}14, DYC^{+}18, FX10, GL19, GLNT13, GS16a, HHD^{+}12, LJ19a, LZH^{+}18, MZP^{+}19, MG18, Nag16, RLL^{+}17, VCL^{+}19, WRCC17, WBJM14, Wri19, YKK13, AKP^{+}18, AT18a, ABS^{+}18, AMMC18, BRL19, BCF16, BGMLS17, CPGBC16, CAC^{+}10, CPSD18, CAPG18, CSL19, CRWZ19, DJH^{+}19, FSM^{+}18a, GIK18, HLT^{+}19, KVR15, KMJ18, KB16, LN18, LPD^{+}13, LZHY19, LHY^{+}19, MGT18, MR19, MSS^{+}16, MHY^{+}18, NKP16, NO19, RMRSA19, SRP19, SLD^{+}15, SHL^{+}19a, TZQ18, TSS^{+}19, WBR19, XFM16, ZYTC15, ZCW19]. Modelling [BGC^{+}19a, CFP^{+}19, DPBK16, HCB16, Kyr19, Lin18, MHW^{+}16, WCH^{+}18, AEM10, CWW^{+}13, EPB18, KMB^{+}17, KGVW14, LSD^{+}17, LASL16, OCW14, SWW^{+}13, SCN^{+}14]. Models [BAP17a, BAP17b, BS17, BGI14, BCdV^{+}19, BBC^{+}12, BNJ16, CMS^{+}18, CHSA18, FEÁ19, HAP11, HLV^{+}16, IKLL12, KI19, KRZ^{+}19, LOR^{+}18, Lin18, LKJ17, LTZ15, dSMAdR^{+}17, MHZK18, Mér17, NBB18, OCCK14, RB12, RHH^{+}19, RLRC13, SWW^{+}13, SLW11, VVB13b, WJZ^{+}17, XA10, YZW^{+}18, vdHSL^{+}15]. modern [LRYJ17, MDT^{+}18, TKR^{+}15, ZME^{+}15]. modes [FAA^{+}18]. modifiable [CLS^{+}19b]. Modified [BRL19, KSS19, SLTK19, ST11]. MODIS [HRR^{+}14]. Modular [DSPA18, WSL^{+}19, Bro19, FSP^{+}18, PSW^{+}14, YP12]. modulated [GZZ^{+}18]. module [WPGN^{+}18]. MOEA [WHCW19]. MOF [DKJ19]. MOF-BC [DKJ19]. moldable [HZZ^{+}14, MLBS11]. molecular [ASV^{+}13, TKTG19]. momentum [DHL18]. monitor [SHLJ13, ZYZ^{+}18, vdLLE19]. Monitoring [DP20, DP21a, DP21b, KS18d, LRZ^{+}18, RMCN^{+}10, VLAC^{+}13, YCD^{+}19, ABC^{+}18, AASI17, AB19b, ASO14, APK^{+}18, AB19c, AMR^{+}19, AIB^{+}18, BKB11, BBWB^{+}18, BCC^{+}17, CA15a, CJHH13, CZH^{+}18, DP19, EET18, FZW^{+}18, FGG13, FAMA^{+}17, FWB13a, FCD^{+}14, GMP^{+}17, GKTK15, GJKP18, HEES19, JNHL18, KSF^{+}13, KBdLG18, LCCM18, LQF19, MVL^{+}18a, MSM^{+}13, MOFGP18, MMLO18, NDZ^{+}18a, NDZ^{+}18b, NDZ^{+}19, NDA^{+}19, PMLVLS^{+}13, RRS10, RYH^{+}19, RAA^{+}19, SB19a, SSL13, TCN^{+}14, TSTL16, TSBH11, WHMO13, WXZ^{+}18b, WPS^{+}18, YCL^{+}19, ZZSZ18, ZZQ^{+}13]. Mono [CCL19]. Mont [OBG^{+}18]. Mont-Blanc [OBG^{+}18]. Monte [CPGdS^{+}13, CPGBC16, DDJ^{+}13, JH16]. Monte-Carlo [CPGBC16]. mood [CC19]. Morpho [LSJ^{+}14]. Morphological [YP12, GDAS18]. MOSIX [MKM11]. mosquito [SM18]. most [CLC11]. Moth [CT19b]. motifs [GAB^{+}14]. motion [KZCW13, SYW17]. motions [CRRC18]. motivators [NZOCJ^{+}19]. motor [AAM^{+}19, CMI^{+}19, RMSPP17, XZZ^{+}19]. motorized [ZMN19]. Motus [GK18]. mouth [XL19]. MOVE [BJ12]. Movement [BRB19a, KSW^{+}13, KPG19]. movements [AKM18, NUPA19, RCMT18]. MOVICAB [HNCJ13]. movie [MBC^{+}11]. movies [Wei11]. Moving [LSAM13, ATS14, ASA19, BTG19, TTC^{+}14]. MP [GLXF17]. MP-MID [GLXF17]. MPI [ACH^{+}11, IPG^{+}18, JS13, LBB^{+}19, LZHY19, PSK^{+}10, RGDML16, WDR^{+}19, ZA14]. MPLS [BLO^{+}18]. MQRP [FG18]. MQTT [RMDB18]. MR [DSM^{+}19, GRS^{+}19, VETT16, YWF^{+}10]. MRA [ACK^{+}15]. MRI [GSV^{+}10]. mrMoulder [CQW^{+}19]. MRPL [BRB19a]. MSNP [WWVJ17]. MUC [QCYJ17]. Multi [AJY15a, AJR^{+}19, AB18c, BMU16, BB17, CGZL19, DNJG17, DNP14, FJJ^{+}18, GTEL^{+}18, GOBL16, GHEB^{+}18, GHEB^{+}23, GCCL18, GLXF17, HMH17, JLQ18, JYZ^{+}18, KA19, LBD^{+}19, LLH^{+}17, LPL^{+}16, LPV^{+}16, LZWF19, MZH^{+}17, MGMT18, MLSF16, MFSV19, MMRL17, PdASM18, PWWD18, RCW^{+}19, SA14, SAH19, SKF^{+}11, SS17, TZL^{+}18, WZW^{+}19a, WW11, WWTF18, XPL19, YDT19, ZFW14, ZCL^{+}14, AD18, ABDH19, AFSH^{+}19, AAM^{+}19, ABN19, BGI14, Bar11, BPC^{+}14, BTP19, BWG19, CCL19, CZY^{+}18, CJXX19, CZ19, CXC^{+}18, CCCT14, DLLZ17, DEG^{+}17, FTK^{+}14, FLT17, FHZW18, FZT^{+}18, FNA12, GQXL18, GLVC18, GTSP^{+}19, GPS13, GHJ^{+}19, GJKP18, sGbKS19, GDR^{+}14, GGLW18, HGG^{+}14, HLZ^{+}19, HFL^{+}19, HIA^{+}18c, HLZ18, Ima19, IDCJ11, IGB^{+}14, JRJ^{+}11, JP18, JTBS15, KDHP16, KKB18, KMT14, KIS11, KESL17, KKKM17, KKKM18, KB16, KGLY18, KLW^{+}17, LJ17b, LXK^{+}14, LWD^{+}14, LYYY18, LWTL19b, LLZ^{+}19, LDJL19]. multi [LH13b, LYS12, LG16b, LW18b, LPD^{+}13, LSG^{+}19, LLAW17, MWYC12, MZP^{+}19, MDD15, MYK16, MRN19, PK11, PRC^{+}14, PDW^{+}11, PMLVLS^{+}13, QCYJ17, QZM^{+}18, QCZH19, RHH^{+}19, RMJ^{+}18, RRU^{+}18, RB18, RMHCMG15, RRH16, STMV18, SSG17, SBAD^{+}18, SEHS19, SWG^{+}16, SCS^{+}18, SCLC19, SOIS12, Sko19, SWW^{+}18, SBK18, SLY^{+}19, SVN10b, TKA^{+}18a, TZD^{+}19, TCBPR16, TCN^{+}16, VD16, WHMO13, WWC14, WHW16, WJZ^{+}17, WJS^{+}18, WSH^{+}16, WXZL11, WSQ^{+}16, WLXZ18, WHCW19, XWZ^{+}19, XYLZ18, YJS18, YQZ^{+}19, YZ12, YJL^{+}19, ZGB^{+}17, ZWL13, ZZH^{+}18, ZTKX19, ZLG^{+}14, ZLL^{+}16, ZHHQ18, ZT19, Zin18, dFPFG19, AB17, HLL^{+}19]. multi- [ZLG^{+}14]. Multi-agent [MFSV19, Bar11, GJKP18, KB16, LJ17b, SSG17, Sko19, WXZL11, WHCW19]. Multi-Algorithmic [GTEL^{+}18]. multi-aspects [MZP^{+}19]. multi-attribute [FLT17]. multi-camera [YJS18]. Multi-Capacity [BB17, HMH17]. multi-category [GGLW18]. multi-channel [GQXL18]. multi-class [RHH^{+}19]. multi-Cloud [JTBS15, ABDH19, DEG^{+}17, KKKM17, KKKM18, RMJ^{+}18, SBAD^{+}18, TKA^{+}18a]. multi-cloud-server [KLW^{+}17]. multi-cluster [WHW16]. multi-community [GHJ^{+}19]. multi-component [SVN10b]. multi-copy [ZTKX19]. Multi-core [MGMT18, MMRL17, LG16b, MWYC12, PK11, PRC^{+}14, QZM^{+}18, VD16, WJZ^{+}17, WSH^{+}16, YZ12, ZHHQ18]. multi-cores [HLZ18]. multi-CPU [VD16]. Multi-criteria [FJJ^{+}18, KA19, MLSF16, ZZH^{+}18]. multi-criterion [FTK^{+}14]. Multi-device [LPL^{+}16]. Multi-dimensional [SS17, WZW^{+}19a, WW11, ZFW14, FNA12, SWG^{+}16]. multi-domain [SLY^{+}19, XWZ^{+}19]. multi-exponentiation [WSQ^{+}16]. multi-factor [LWD^{+}14]. multi-feature [HLZ^{+}19]. Multi-fidelity [LBD^{+}19]. multi-formalism [BGI14]. multi-function [WLXZ18]. multi-generation [CJXX19]. multi-gigabit [KDHP16]. multi-GPU [VD16]. Multi-GPU-based [DNJG17]. multi-grid [CCCT14]. Multi-Hop [WWTF18, SAH19, KESL17, LLAW17, TZD^{+}19]. Multi-information [GCCL18]. multi-installment [SOIS12]. multi-job [WWC14]. multi-kernel [JP18]. Multi-key [LLH^{+}17]. multi-keyword [LXK^{+}14, WHMO13, YQZ^{+}19]. multi-label [GLVC18]. Multi-layer [TZL^{+}18, AAM^{+}19, SCS^{+}18, YJL^{+}19]. multi-layered [WJS^{+}18]. Multi-level [YDT19, AD18, CXC^{+}18, GOBL16, LSG^{+}19, RMHCMG15]. Multi-lingual [CGZL19]. multi-link-failure [RRU^{+}18]. multi-modal [CCL19, CZ19]. Multi-objective [AJR^{+}19, DNP14, GOBL16, JLQ18, LPV^{+}16, LZWF19, ZCL^{+}14, CZY^{+}18, FZT^{+}18, GPS13, HFL^{+}19, LWTL19b, LDJL19, STMV18, SEHS19, WWC14, ZLL^{+}16, Zin18]. multi-order [LW18b]. multi-party [KKB18]. Multi-point [SKF^{+}11]. multi-processor [MRN19]. multi-projection [BWG19]. Multi-Protocol [GLXF17]. Multi-provider [BMU16, IGB^{+}14]. Multi-QoS [AB18c, AB17]. multi-relay [ZWL13]. multi-resolution [sGbKS19]. multi-resource [LYYY18, LLZ^{+}19, ZT19]. Multi-scale [JYZ^{+}18, RCW^{+}19, Ima19, XYLZ18]. Multi-scheduler [AJY15a]. multi-sensor [dFPFG19]. multi-server [FHZW18, LH13b]. Multi-site [SA14, TCN^{+}16]. Multi-spectral [GHEB^{+}18, GHEB^{+}23]. multi-stage [SCLC19]. multi-start [KMT14]. multi-tenancy [BPC^{+}14, TCBPR16]. multi-tenant [GTSP^{+}19, MDD15, PMLVLS^{+}13, RB18, SWW^{+}18, ZGB^{+}17]. Multi-threaded [PWWD18]. multi-tier [GDR^{+}14, HGG^{+}14, IDCJ11, KIS11, LPD^{+}13]. multi-touch [PDW^{+}11]. multi-trusted [WW11]. Multi-UAV [HLL^{+}19]. multi-user [DLLZ17, JRJ^{+}11, KGLY18, PDW^{+}11, RRH16]. Multi-valued [MZH^{+}17]. multi-version [QCYJ17]. multi-view [BTP19, LYS12, QCZH19]. multi-wireless [MYK16]. multi-workflow [ABN19]. Multiagent [RML^{+}19]. multibiometrics [GEAR13]. multicast [ZWJ^{+}18]. multicasting [EBCP18]. Multicenter [BJM^{+}17]. multichannel [MYBMM18]. Multicore [LJC^{+}19, ADH^{+}16, BFP18, CHS^{+}18, HAF^{+}16, JLRS18, KR19, KR14, LRYJ17, PSPP16, RGDML16, WCC^{+}16, WTG^{+}19, WPJ16, WLRL18, HYZS16]. multicores [SHP^{+}16]. Multidimensional [CLR16, PLW^{+}19, JLC18, YC13]. Multifaceted [ABTA18, GBF^{+}12]. MultiGrain [MSS^{+}16]. MultiGrain/MAPPER [MSS^{+}16]. Multilayer [SEHS19]. Multilevel [SRP19]. Multimedia [ACM^{+}18, HXA^{+}17, KCM19, LBD^{+}19, WDW^{+}19, YWL^{+}17, AMPS19, AMPP19, CdRRdCB19, GSY^{+}17, HLYW17, HYX^{+}19, HXL^{+}18, HPL^{+}19, HCHH19, IG12, KLJS19, LXF19, PKY^{+}17, RHH^{+}16, SLS^{+}19, TGM^{+}19a, WLH^{+}19, vKvWD^{+}13]. Multimodal [Sar18a, AHM^{+}18, EAA16, HUY^{+}19]. Multimodel [KVK^{+}18]. multinode [WTG^{+}19]. Multiobjective [CZL^{+}18a, CZY^{+}19]. multipath [HYX^{+}19, KN10, VSKS19]. Multiple [CWW^{+}16, SKS^{+}18, YZZC19, Ans11, AM19b, BBB16, CTVB12, CMB17, CAPG18, CSL18, DPL14, EG18, GOLL19, GAA19, HQ10, KSK^{+}11, LLY^{+}18, Li18, LMM19, LSMVML13, MJM^{+}16, PC17, RM11, RMHMG17, SG17, SLA^{+}16, SVN10b, TMMVL12, WTM^{+}17, WLGL19, XDHL12, YXZG18b, YD18, YQZ^{+}19, YL18, ZYB^{+}18, ZYCZ19, dlFVPSHL^{+}14, dCTVC18, dFPFG19]. Multiple-combinational-channel [CWW^{+}16]. multiple-mix-attack [LMM19]. multiple-perspective [HQ10]. multiplexing [HFM19]. multiplication [ZWL^{+}16]. multiplicative [TWW^{+}18]. multipliers [WMLS14]. multiply [CSQL17]. Multipolarization [FLN^{+}18]. multiprocessor [DKV14, SC16, ZMTT16]. multiprocessors [HHXL13]. multiresolution [GRS^{+}19]. Multiscale [APS^{+}19, BC15, BC17, BEWZ10, MSS^{+}16, RBC^{+}15]. multiside [CRYG18]. multisite [DST10, LPV^{+}16]. multispectral [Ima19, SYCH18]. multistage [GS13, KYB^{+}19]. multitask [MSO18]. Multiuser [NRR^{+}15, dSBN19]. Multiuser-centered [NRR^{+}15]. multivariate [WSL^{+}19]. multiworkflow [AHEM17]. MUSCLE [BC17]. MUSCLE-HPC [BC17]. music [PMFH11]. mussel [PRW14]. Mutation [SH19, JGFB18, JLC^{+}20, TM19, YWG^{+}19, YWG^{+}20, YDD^{+}18]. mutual [Bag19, BDL^{+}19, JKAU19, KLW^{+}16, VCD^{+}18, XPL19]. MWM [WMLS14]. MyGeoHub [KZS^{+}19]. myocardial [WXZ^{+}18a].

name [ACSdRR17]. named [RWO^{+}19, ADBM19]. Naming [ACSdRR17, LKJ^{+}19]. NAND [SCH^{+}19]. nanoprobe [wZcZN^{+}19, wZcZN^{+}20]. Narrow [WYJ^{+}19]. Narrow-band [WYJ^{+}19]. Nash [GPVN19]. nasopharyngeal [MGA^{+}18]. National [SVN^{+}10a]. native [SB19b, SCJ^{+}19a, ŠCJ^{+}19b, TBB^{+}17]. natural [DCC^{+}14, FS19, RRH16, XLL^{+}18a]. Nature [VV16, MOBD18]. Nature-inspired [VV16]. navigation [LLS^{+}19, XZ14a]. NB [HEES19]. NB-WOA [HEES19]. NBA [JLQZ18]. NBA-based [JLQZ18]. NDGF [KVR15]. NDN [ADBM19, MDB^{+}18b, MRS18b]. NDN-over-ZigBee [ADBM19]. NDN/IoT [MDB^{+}18b]. near [PGTBC18, wZcZN^{+}19, wZcZN^{+}20]. nearest [FJ18, GLVC18]. necessary [XWL^{+}18]. need [SOA17]. needs [GFD14]. Negative [YLH^{+}19, DYC^{+}18, HZX^{+}19, HZX^{+}20, QCZH19]. negotiation [COC10, LZYC13, MG14, MG16, RT16, RZ16, SSL12]. neighbor [CZL^{+}18b, FJ18, GLVC18]. Neighborhood [ZF16, CZ12, CSL17, NAGD18, TM19, WYH^{+}17, ZSJ19, ZRZ^{+}14]. Neighborhood-based [ZF16, ZRZ^{+}14]. neighbors [CSL19]. Nephropathy [QXZ^{+}19]. nervous [UPP17]. net [YDQC19, dLB10]. NetCoM [RW13]. NetCoM-2009 [RW13]. Netflow [KC19b]. nets [AEM10]. Network [ACC^{+}19a, AS18b, BZHV19, BHE^{+}19, BCR^{+}12, CBN16, CW13a, CWL^{+}19, DDD^{+}19, FS19, HDKC18, JL14, MLC^{+}11, MBM18, TMW^{+}17, TCCC11, VDTK12, WWTF18, WLRL18, ZHHC17, AAYL19, AFO^{+}18, ABD^{+}19, AQAR^{+}18, AAQ^{+}19, ABF^{+}15a, AFB^{+}10, AR10, ASA19, BMU16, CEP19a, CWL^{+}18, CFL^{+}15, CJXX19, CWW^{+}16, CZH^{+}18, CS12, CFPC17, CDB^{+}19, CS19, CNR19, DPK^{+}19, DCBF19, DBP19, DPBK16, DVB14, DCF19, FBM19, FZHH14, Fer13, FJKK17, FWB13b, FFL^{+}19, GOBL16, GBY16, GVdBdL15, GSC^{+}19, GGC18, HZW19, HPZL18, HAF^{+}16, HUY^{+}19, HLYW17, HHH^{+}19, HCNT14, HH19, HSP^{+}13, HSB^{+}18, HZL18b, HDH^{+}18, KR19, KSS19, KLJS19, KBdLG18, KMST19, KK10b, KAS^{+}18, KS18c, LJ19a, LZXW13, LNB14, LMZ^{+}14, LYYY17, LXL^{+}17, LLT^{+}19, LLZ^{+}19, LZT^{+}19, LLL^{+}19, LLW^{+}19b, LSYC18, LCL^{+}18, LDX19, LSH^{+}11, LWT18, LCdPMCT19, LMCSE19, LKTC14, MYHZ18, MHC14, MVG18, MBC^{+}11, MWCK19]. network [MBL^{+}19, MGA^{+}18, NAGD18, PP10, PJDO13, PMK18, PBL^{+}18, PECA19, Pip10, PTT12, RR18, RGCCL18, SSG19, SGRT19, SB19c, SP18b, STC15, SKS^{+}18, SCMS12, SNP19, SYAL13, TZBK13, TBS^{+}18, TCR^{+}12, UZ11, UDvdW^{+}18, VL19, VCDK18, VOS12, WTG^{+}14, WRCC17, WCB^{+}18, WWZZ18, WCW18, WZF^{+}19, WMA^{+}19, WLA18a, XWL^{+}15, YMLT13, YCL^{+}19, YCXW18, YHH^{+}19, ZGB^{+}17, ZZDM^{+}18, ZZF^{+}19, ZGV19, ZSFZ19, ZWJ^{+}18, ZS10, ZCW19, ZWMC19, AS18a, BRH18, BSE^{+}13, GPJC17, GKTK15, HQ10, MCJ19, RGSL18, Sun10, WFQ^{+}10, Zhu10]. network-aided [GSC^{+}19]. Network-aware [BCR^{+}12, TCCC11, VDTK12, ZHHC17, PP10, TCR^{+}12, UDvdW^{+}18]. network-based [LYYY17, RGSL18]. network-driven [YMLT13]. network-failure [DCF19]. Network-on-Chip [MBM18]. network-on-chips [SSG19]. networked [PMFH11]. Networking [SMG18, ABG18, CJK^{+}18, GXL^{+}12, GZLZ16, LLF^{+}18b, LGP^{+}19, MCJ19, MT17, PYM18, RLL^{+}17, RWO^{+}19, SP18a, SCZ^{+}19, SL19, TGM^{+}19a, TBdL16, WWVJ17, WGX^{+}19, ZBCT17, ZWJ19b, ADBM19]. Networks [AZO^{+}19, EGVT18, FPPD14, HGM15, KIAD17, LY19, MWQ^{+}19, PLLP19, RQN^{+}19, RCW^{+}19, SYJ^{+}19b, SVK19, SRP19, SMS16, TKA18b, WZWW18, AsRA^{+}19, AJ19, AQRH^{+}18, ARSMY19, ASO14, APK^{+}18, ACM^{+}18, AMPS19, AMPP19, ASYF18, AIB^{+}18, AMM19a, AAS^{+}19, AC18, ASA19, BAA^{+}19, BDP11a, BLAN^{+}16, BMK^{+}14b, BCR^{+}12, CEP19b, CZY^{+}18, ÇÖ13, CJ14, CWJ16, CJG^{+}18, CZM^{+}18, Che18, CZZ^{+}18, CW13b, CSC18, CdRRdCB19, CGJ^{+}10, CHY^{+}18, DBP19, DKJ19, DYC^{+}18, ELAEAVAM19, ESPN17, EYY19, EA13, FG18, FJJ^{+}18, FZHH14, FW19, FAL^{+}19, FP14, FLN^{+}18, GRTV10, GS13, GJY18, GSP^{+}17, GHO^{+}11, GBKJ18, GWC^{+}16, GLXF17, GXL^{+}18, GZL^{+}18, GSN^{+}18, GGMS18, GNVST14, HKA^{+}18, HAAWH^{+}18, HA19, HCC^{+}14, HAM18, HMC19, HYG^{+}19, HDB18, HYF18, HAAR^{+}19, IASK14, JLQ^{+}17, JXC^{+}19, JWW14, KWK^{+}18, KRZ12, KPM^{+}18, KC19b, KCV11, KV12, KESL17, KLW^{+}16, LZY^{+}19a, LLYW19, LC17]. networks [LBD^{+}19, LWD^{+}14, LLQS14, Li15, LFL^{+}17, LCW^{+}18, LZLL18a, LZP^{+}18, LW18a, LL18, LLW^{+}19c, LAQ^{+}19, LCZR12, LY18b, LWH^{+}18, LQLX10, LLJ^{+}11, LZXG12, LWX13, LvW14, LWL^{+}18, LYW^{+}18b, LGZY18, LMM19, LWXY19, LO19, LHBC16, LLU^{+}18, LFY^{+}19, LZY^{+}16, LLAW17, LWZ^{+}19b, LJW^{+}19b, MML^{+}18, MG19, MNC^{+}18, MDB^{+}18a, MDB^{+}18b, MFL18, MSS^{+}16, MLW^{+}18b, MSO18, MYK16, MSM^{+}18b, MMAA19, NBB18, NLLC19, NS10, NJKF18, NSSA^{+}14, NJKH13, ODC19, OSANAM19, PSJ^{+}12, PRS^{+}13, PC17, PDH18, PMMAM13, PPPS18, PWWD18, PF17, RWY^{+}18, RHMGC14, RWZ^{+}19, SJ14, SPT^{+}18, SGGCR^{+}16, SISGS18, SOD18, SCY^{+}18, SCS^{+}18, SLS10, STC15, SGBK19, SCEC18, SJL^{+}18, SLZ^{+}18, SLY^{+}19, TSD18, TY11, TL19, VSBN19, VCD^{+}18, WSQ^{+}18, WSZH18, WLP10, WN10, WWD^{+}14, WJLW18, WDD18, WDJC18, WCM^{+}19, WMC19, WdL16, WLQ10, WLS^{+}18, WLZ^{+}19, WSY^{+}19, XCS^{+}18, XLL^{+}14, XZ14b, XDH^{+}17, XXX^{+}19, YPLZ17, YHL^{+}19, YL18, YLA18]. networks [YLH^{+}19, ZF16, ZSJ19, ZGL19, ZLL17a, ZZSZ18, ZZH^{+}18, ZTKX19, ZWS^{+}12, ZNC^{+}18, ZBL^{+}14, ZCL^{+}19, ZZC14, ZZLH18, ZSS^{+}18, ZCDV19, BMK^{+}14a, IDKD19, LNK^{+}18, ZHL^{+}18]. networks-based [FG18]. Networks-Generated [ZHL^{+}18]. Neural [DDD^{+}19, FS19, GKTK15, HDKC18, LMCSE19, SKS^{+}18, AAYL19, AsRA^{+}19, AFO^{+}18, ASYF18, CZH^{+}18, Che18, GSC^{+}19, HZW19, HPZL18, HUY^{+}19, HHH^{+}19, HZL18b, KLV^{+}18, KS18c, LBD^{+}19, LZT^{+}19, LLW^{+}19c, LCL^{+}18, LYW^{+}18b, LLU^{+}18, MDB^{+}18a, MGA^{+}18, MMAA19, PWWD18, SB19c, TZBK13, TBS^{+}18, WCB^{+}18, WZF^{+}19, WLA18a, dLB10]. neutrosophic [ABMMC18, ABM19, ABMMC22]. news [CC19, LRL^{+}14, XWL^{+}15]. Newton [Jun18]. Next [VB18, BDP11a, GLM^{+}12, JXZ^{+}19, MVL18b]. next-generation [MVL18b]. nextgen [VCE^{+}19]. Nexus [CLM^{+}16]. NFV [ASA19]. NIR [TA19]. NLP [ACMM19]. NMCDA [ABMC18]. NN [STA17a]. node [CWJD19, HHZ16, LC17, LWD^{+}14, LJJ18, LKFB18, QZM^{+}18, ROK19, SYJ^{+}19b, ZSJ19]. nodes [AC18, GZW18, JXC^{+}19, KS11, LLYW19, LMM19, RYH^{+}19, RBJ^{+}13, ZZC14]. noise [MFT^{+}17]. noisy [MFT^{+}17]. NOMA [KKB^{+}19, SCEC18]. nomadic [EA13]. Non [NGB18, RNR18, YC13, AFO^{+}18, AM19a, BBB16, CZT^{+}15, CW16, CLR18, GHYK18, GOLL19, JLC^{+}20, KKB^{+}19, QCZH19, UMUB19, WWX^{+}17, WCHL10, XZW^{+}19, YAO14, YWG^{+}19, YWG^{+}20]. non-cloud [CW16]. Non-cooperative [YC13]. non-dedicated [BBB16]. non-deterministic [GOLL19]. non-functional [YAO14]. non-Gaussian [KKB^{+}19]. non-increasing [WWX^{+}17]. non-linear [CZT^{+}15]. non-local-convergent [XZW^{+}19]. non-medical [AM19a]. non-negative [QCZH19]. Non-parametric [NGB18]. non-patchable [GHYK18]. non-preemptive [CLR18]. non-reserved [WCHL10]. non-shockable [AFO^{+}18]. non-small [JLC^{+}20, YWG^{+}19, YWG^{+}20]. Non-standard [RNR18]. non-stationary [CZT^{+}15, UMUB19]. nonlinear [KHO^{+}19]. Norm [SZR18]. Norm-based [SZR18]. normal [RJN^{+}19, WLZ^{+}14]. normative [BDP11a]. normed [AB19c]. Nornir [DDD18]. North [PKP19]. NoSQL [BGI14, Cha15, FSM^{+}18a, GAYTC18, LZL^{+}16]. note [RTHB17, TKR^{+}15]. notebook [ZDL^{+}13]. nothing [KMB16]. notice [AB21, ABMMC22, Bo20, DP21a, DP21b, GHEB^{+}23, HZX^{+}20, JLC^{+}20, LBJ^{+}24, WWP20, WCWC20, YWG^{+}20, YTQ20, ZMZ^{+}20, wZcZN^{+}20]. notions [DGY^{+}18]. nourishing [QXZ^{+}19]. NovaGenesis [ACSdRR17]. Novel [BLO^{+}18, CDB^{+}19, FGM11, ABM19, AS19a, AK19, ArMS19, ASY^{+}18, ATX13, CPDJ13, CGCB^{+}12, CMI^{+}19, CZ12, CAL^{+}18, CGM^{+}18, CHY^{+}18, DdM10, DC19, HZW19, HLL18, HHM^{+}19, HHH^{+}19, HXWW18, HHZ16, JLY^{+}18, JXZ^{+}19, KHWZ18, KPG19, LLF^{+}18b, LYW^{+}18a, LXRS19, LJY12, LH13b, LJ19b, LLJ^{+}11, LLSL18, LHW^{+}18, LZW^{+}18, ML19, MdMMNS^{+}19, NPP12, NZL^{+}15, Osm19, QCY^{+}19, RAKJ18, RRU^{+}18, SMS^{+}19, SAM^{+}19, SGJ18, SYW17, ŠCJ^{+}19b, TLL^{+}19, TMJH19, TYWZ18, TMS^{+}17, UPP17, VGD^{+}19, WLLF16, WDD18, WXZL11, XZW^{+}19, YAO14, YK17, ZAA^{+}14, ZA13, ZYW^{+}18, ZCQ^{+}16, ZWZ19, wZcZN^{+}19, wZcZN^{+}20]. November [Ano19m]. NOVI [vdHSL^{+}15]. NP [BP13]. NP-complete [BP13]. NSGA [YDD^{+}18]. NSST [GDAS18]. NUCA [BFP18]. NUCA-based [BFP18]. NUMA [HTL^{+}18, LRYJ17]. number [BGC19b, LSG18, PSS^{+}18, RNR18]. numbers [SAM^{+}19]. numerical [CCY^{+}18, ZRZL18, ZXL14]. Nuoxus [CdRRdCB19]. nutrition [CAS^{+}18]. NVM [CSJ^{+}17].

O [CSL18, DLZ^{+}14, DLXR14, DYI^{+}16, LLW^{+}12b, TDFZ18, ZXL14]. Oak [SVN^{+}10a]. Oasis [XFLL16]. obfuscation [MdFTGM19]. Object [AMBB18, GSY^{+}17, ATS14, BTG19, BBB^{+}19, GWW^{+}19, GTSP^{+}19, MGA^{+}18, PSLZ18, RACA18, SLK17, Sun10, XFLL16, YARH18, Sun10]. Object-oriented [AMBB18, PSLZ18, Sun10]. objective [AJR^{+}19, CZY^{+}18, CGSJ18, DNP14, FZT^{+}18, GOBL16, GPS13, HFL^{+}19, JLQ18, LWTL19b, LDJL19, LPV^{+}16, LZWF19, STMV18, SEHS19, WWC14, ZCL^{+}14, ZLL^{+}16, Zin18]. Objects [MDA^{+}19, SAGGB17, GSGPP^{+}19, HC17, TTC^{+}14]. oblivious [Tso19, ZWL^{+}16]. Observation [LCdPMCT19, PD11, RVST17]. Observer [KS17a]. obstacle [ZDL^{+}19]. Obstacles [KT17]. Obstructive [DDD^{+}19]. OCCIware [ZCM19]. occupancy [GXL^{+}18, PAB^{+}14]. oceanographic [CGL^{+}10]. October [Ano19k]. ocular [MAPA19]. ODROID [LPS19]. ODROID-MC1 [LPS19]. OFDM [KKB^{+}19]. OFDMA [ZN12]. off [BDA19, CPSD18, DGCGH^{+}17, GBS10, KCM19, PMBS14, SMM^{+}14]. off-chip [CPSD18]. off-the-shelf [SMM^{+}14]. offering [CMB17]. office [BdM11]. offline [CYJ19, SHLJ13, WLGL19, XTZ^{+}19]. Offloading [AZH18, WLZ^{+}16, AHU^{+}19, BGS^{+}19, EZTL19, HHK18, HLT^{+}19, JOSD19, KPJ19, KGLY18, LPL^{+}16, MCAS19, SGN^{+}17, WMJW18, XLL^{+}19a, XXQ^{+}19, YZC^{+}19, YCH19, YCX18, ZYC^{+}19]. offs [DMM14]. offset [HLL18, KK10b]. oil [FCD^{+}14]. OLAP [EDD^{+}10]. older [CAS^{+}18]. OLTP [KMB16]. OMAS [Bar11]. omnipotence [MBV^{+}15]. omnipresence [MBV^{+}15]. omniscience [MBV^{+}15]. OMPI [ACH^{+}11]. OmpSs [YÁJG^{+}15]. OMS [WLZ^{+}14]. OMS-WPD [WLZ^{+}14]. on-campus [GLD^{+}19a]. on-chip [CPSD18]. on-demand [DEG^{+}17, LWH^{+}18]. On-line [BMR15, LLW^{+}19c, CCDP19, TJ18]. On-siteDriverID [SGGCR^{+}16]. One [CCD^{+}19, FLT^{+}19, RM19, MZP^{+}19, SYCH18, Tso19]. One-pass [CCD^{+}19]. Online [AMPP19, HHW11, KOP^{+}17, MG18, RGGH18, SRP19, dSSCdL19, VVB13a, AMQS^{+}19, AQRH^{+}18, AMKM18, ADBO18, ABH18, BCMM18, CYJ19, CYW^{+}19, CSQL17, DHW^{+}17, DdSdN^{+}19, GWC^{+}16, GGC18, GSN^{+}18, HLT^{+}18, HYF18, JWW14, Jun18, KAP19, KVvE18, LRL^{+}14, LPY^{+}18, LBU^{+}10, MML^{+}18, MSM^{+}18b, NZOCJ^{+}19, NJ17, PQBP17, SSHC19, SC16, SKS^{+}18, SGBK19, SYW17, TL19, WXYL15, WHW16, WJLW18, WXZ^{+}18b, XCS^{+}18, XWL^{+}15, XTZ^{+}19, YLA18, ZLL17a, ZYB^{+}18, ZXZL18, ZCL^{+}19, dSFD^{+}19]. online/offline [XTZ^{+}19]. Ontological [SCN^{+}14, OCW14]. Ontologies [CBS17, KGdL11]. Ontology [BNJ16, GMP^{+}16, TOD17, TF17, AC16, Bae14, CXDM18, CFL^{+}15, GTM19, HCX^{+}19, LKN^{+}13, RGVGGSSZ14, SBLW14, SMZ^{+}16, SSZ^{+}17, TJWS10, TZL^{+}18, TNY17, VPP^{+}19, WMBV17]. Ontology-based [TOD17, Bae14, SBLW14, TJWS10, TZL^{+}18]. ontology-centric [LKN^{+}13]. Ontology-driven [GMP^{+}16, VPP^{+}19]. opcodes [ZXM^{+}19]. Open [CAS^{+}16, PVHTP19, QC18, ASAAM^{+}19, ACC^{+}19b, BFN18, CCM^{+}14, HKP10, KS11, KS18b, KTTK17, MG11, MKS^{+}19, MQN19, PWA^{+}19, SJV^{+}15, SWY^{+}18, SGM11, Sko19, SYK^{+}17, TSGVRGS19, TCB^{+}17, YHA^{+}19, GM11, Mil11, MCF^{+}11, SB11]. OpenCL [CJPC19]. OpenMOLE [RLRC13]. OpenNebula [KMT14]. OpenStack [CFF14]. Operating [WCWC19, WCWC20, ZYA^{+}18, BBI13, DDMPG17, LSW^{+}19, ZAI^{+}18, Zin18]. operation [ALL^{+}18, ODC19, Sun10, WC14, ZTC^{+}19]. operational [RB12]. operations [EG18, GDAS18, LHCC18, LHY^{+}19, MBB10, TMDZ15, ZZC18, vWMBS14]. operator [SH19, YDD^{+}18]. operators [ECPF17b, TSOB15]. opinion [ZNC^{+}18]. opinionated [MZP^{+}19]. OPM [FKOC11, LLCF11]. OPM-compliant [LLCF11]. Opportunistic [CFP^{+}19, OFMZ18, ZHL^{+}18, FFL^{+}19, JXC^{+}19, JCA^{+}19, MCAS19, PJDO13, WSZC18, WZWW18]. Opportunities [AR17, Jun17, PYM18, AZH18, ALK15, CLCY18, CBBC^{+}17, CDFW18, GS16a, GRL11, GFD14, JC15, MWW^{+}15, OFMZ18, QCD16, RMC^{+}18, WGX^{+}19]. Optical [MSM^{+}18a, CGJ^{+}10, GHO^{+}11, KK10b, LKTC14, YGYW16]. Optimal [AOIS10, CT19b, DEG^{+}17, JGB19, KMK^{+}19, KCM19, LXD17, BAA^{+}19, BGRBA19, CXWT19, CJS19, EBCP18, FTK^{+}14, HZLH19, IDKD19, KA19, Len16, LSMVML13, SRKS18, SMS13, XYLZ18, ZLTY10, ZWL13, ZYCZ19, ZaTZ^{+}17, ZB19]. OPTIMIS [Ano12r]. Optimisation [XRPT18, AL18, AR10, HFL^{+}19, PTD^{+}18, TVB18, VVB15, WHW17]. Optimising [KACN16]. Optimistic [LM12, Pon19, WWW^{+}16]. Optimization [LXJD18, SSG17, TA18, WPS^{+}18, ZCS^{+}16, ABMESM18, ANA16, AKP^{+}18, AT18a, AC18, BRL19, BMU16, CT19a, CPGBC16, CZL^{+}18a, CHC^{+}17, DDD^{+}19, ECA^{+}18, FZT^{+}18, GSLI12, GHEB^{+}18, GHEB^{+}23, GGA^{+}17, GLJ19, GZZ^{+}18, HPZL18, HAP15, HHM^{+}19, HMF^{+}19, HZP^{+}14, HZL18b, HLZ18, HCHH19, HKP10, ISS^{+}15, JCL^{+}19, KKS18b, LLWW18, LAH10, LLW^{+}12b, LJW^{+}19b, MGMT18, MFSV19, PFRC16, QWCW19, RD14, SCY^{+}18, SEHS19, SSJ19, SGN^{+}17, SH19, SSI19, SGB^{+}18, SLS^{+}19, SHL^{+}19b, TWZP18, TV16, VR12, WWX^{+}17, WCL^{+}17b, WWQ^{+}18, WZWC18, WWZC19, WYJ^{+}19, WZ13, XHW19, XZW^{+}19, XDHL12, YDD^{+}18, YKK13, ZRZL18, ZN12, ZZJ17, ZWZ18, ZYW^{+}18, ZWZ19, ZLG^{+}14, ZLL^{+}16, ZQB^{+}18, ZSMS18, ZHHQ18, ZXL14, MWMA10, PT16]. optimization-based [BRL19, HAP15]. optimize [CdSDS15, HPGMM18]. Optimized [CAB^{+}18, FDPR17, GTMZ17, KR19, TKA18b, AKCY^{+}17, AAD^{+}13, BKKM11, DR18, DKJ19, GS16b, GSY^{+}19, LXL^{+}17, LLF^{+}18b, LWTL19b, LJY10, TF18, TMMVL12, UMUB19, VS19, ZCDV19, dSBN19]. optimizer [DHL18, SJ19]. Optimizing [BBB^{+}19, dCCDFdO15, MNY^{+}19, NF13, OSC14, SSI19, WYBS11, WLQ10, ZHW19, Zin18, ZSBB19, vWMBS14, FEPC18, HSC15, MFL18, VSDD13, WBJM14, WLH^{+}19, ZSMS18, vW19]. optimum [IdAP19]. optimum-path [IdAP19]. oracles [KGS^{+}19]. orbit [LY18b]. Orchestrating [CKP^{+}19]. Orchestration [MT17, EKSDN19, JTS13, SLL^{+}18, SLY^{+}19, VCE^{+}19, XWZ^{+}19]. Orchestrator [KKK^{+}19]. order [Che14, DJPM18, LW18b, TJ18]. ordered [Bag11]. Ordering [HMH17, Bag19, BBI13]. OREA [TKA18b]. organisations [SMF^{+}19]. Organism [ANA16]. organisms [EA17]. Organization [MG10, ZZH^{+}18]. Organizational [HKG^{+}16, GCTLA^{+}19, KS19]. Organizational-based [HKG^{+}16]. Organizations [YAA^{+}19, Li10, SCZ^{+}19, SMSF18, SR19]. organized [RP18]. organizing [MPR^{+}16, dRRRR^{+}18, dRRdQGR^{+}18, SJL^{+}18, VRGR16, XWL^{+}15]. orientated [ZDL^{+}13]. Oriented [BZS18, GLXF17, SCL18, VDTK12, YAA^{+}19, AMBB18, AWN^{+}13, AB18b, AQB15, BAA^{+}19, BR18, CJPC19, CPDJ13, CLL^{+}14, DVD12, DM12, EAED18, EG18, FLR^{+}16, FKT14, GVURIVBV14, GA13, HYX^{+}19, HLT^{+}18, JKLK17, KK16, LX13, LWZ18, MLSF16, MRS^{+}18a, NK15, ÖE13, PSLZ18, PBA18, SBLW14, SSJ19, SZD^{+}17, SLL^{+}18, SAC11, TAKV12, VDK12, WDJC18, WZWW18, XLW^{+}17, XDHL12, YNSM12, YZW14, Sun10]. OSG [VHML10]. OSINT [QC18]. OSLN [Sun10]. OSNs [GSN^{+}18, TWG^{+}19]. OSPF [WMA^{+}19]. Osseous [DGA18]. osteochondral [Bo19, Bo20]. osteoporotic [WWP19, WWP20]. other [HXA^{+}17, ZZC18]. outdoor [MLGGB^{+}17]. outlier [DJPM18, NGB18]. outliers [SAG19]. outpatients [JYZ^{+}19]. Output [JKLK17, GOLL19, HPZL18, WHCZ18]. Output-oriented [JKLK17]. outputs [CCRL18]. outsource [XTZ^{+}19]. Outsourced [GZQ^{+}19, DLMS15, FS18, LYZC15, WWRS16, WCL^{+}17a, XXB19, ZSW^{+}18b]. outsourcing [BGRBA19, ED16, HCW^{+}18, JCL^{+}15, ZCL^{+}18, dACAM13]. Over-sampling [MK19a]. over-the-air [ZXW^{+}18]. overbooked [LPBB^{+}18]. Overcoming [USK16]. overcommitted [CJHH13, ZHHC17]. overhead [CXWT19, LZHY19, MVL18b, RS17b]. Overlap [KKB^{+}19, LJJ18]. overlapping [BR19, SEHS19, SJL^{+}18, SAC11]. Overlay [SCMS12, CW13b, EA13, MLBS11, Pip10, PTT12, PIP18a, TY11]. Overload [CNR19, LYYY18]. overloaded [SPJ17]. oversubscription [MCdA16]. Overview [ZG18, AT19a, ATA19, AB19b, BCP18, FBS18]. own [VPA^{+}18]. Owner [HLH^{+}18, SBL18]. owners [YQZ^{+}19]. Ownership [AMSPL19, KH18b, LCL^{+}19, RACA18, SK18].

P [JLL17, FK11, MHZK18, SSLF^{+}10]. P-batch [JLL17]. P-found [SSLF^{+}10]. P-GRADE [FK11]. P-Spec [MHZK18]. P2P [AR10, BCR^{+}12, CEP19b, CsZW14, CW13b, GDJ^{+}13, HCC^{+}14, HNKÖ18, HDLW13, LS10, LJY12, LQLX10, OPO13, OCCK14, PRS^{+}13, RLP12, RHMGC14, SAGL10, SYL18, SYQ^{+}19, WN10, WW13, YMD^{+}13, ZZL^{+}10]. P2P-assisted [HDLW13]. P2P-based [GDJ^{+}13, SAGL10]. P2P-content [BCR^{+}12]. PaaS [KKKM18, KKKM17]. paced [SYT^{+}19]. packet [KDHP16, LLL^{+}19, MSM^{+}18a, ZCDV19]. packets [HCL^{+}17]. Packing [BB17, LLZ^{+}19, LZCX19, MAA^{+}19, SK12, ZT19]. Page [LLZ^{+}19, BRXdS11, LXL^{+}17]. Page-sharing-based [LLZ^{+}19]. PageRank [MK19a, RCM17]. Pages [Ano19q, Ano19n, Ano19p, Ano19o, Ano19r, Ano19m, JDW^{+}14]. Pairing [MLC^{+}18a, LL16, WMC19]. pairing-free [LL16]. Pairwise [CRWZ19, XZ14b, YWJ^{+}18]. Palmprint [BG12, CPW19]. panchromatic [GHEB^{+}18, GHEB^{+}23, Ima19]. pansharpening [YJY^{+}18]. paper [ECPF17a]. Papers [TCG14, CC11, CCRL18, KZ17, RW13]. para [LC13]. para-virtualized [LC13]. paradigm [AHS^{+}18, AHP^{+}18, BLMU19, BDP11b, BRHH18, GBY16, LC14, WMC19, XLW^{+}17, ZAA^{+}14]. paradigms [GSN^{+}18, OB17, RR18, SISGS18]. Parallel [AJY15a, BTM10, CDG^{+}14, CLZ18, EDD^{+}10, KKI14, KAP19, dRRdQGR^{+}18, TCG14, TKRA14, ZHHQ18, dRADFG18, AS14, BBH18, BC17, BKB18a, BGMLS17, CGCB^{+}12, CVT19, CPGdS^{+}13, CZY^{+}18, CZL^{+}18a, CZY^{+}19, CRVZ15, CMS^{+}18, CHSA18, CCL11, CXL^{+}17, dCCDFdO15, CSP13, DST14, FGM11, FM10b, GVDT16, GBS10, GGLD10, GZWQ13, HZ19, HHZ16, HLZ18, ISS^{+}15, JPB17, JLY^{+}18, JEB18, LYW^{+}18a, Li18, LJ19b, LSD11, MFL18, NFK10, NQQL13, NOF18, OdOD^{+}13, PRN14, QPTGG^{+}12, dRRRR^{+}18, SHN10, SLC^{+}17, SJL^{+}18, SVN10b, TTC^{+}14, TF18, WKC^{+}13, WHW16, WJZ^{+}17, WWZC19, WPY19, WZF^{+}19, WHYZ17, YSC^{+}19, ZXW19, ZXL14, dLB10, dOOO^{+}13, vKvWD^{+}13, NS17b, VD16, WTG^{+}19]. parallelism [CDG^{+}14, HPGMM18, WSC^{+}19, WC14, XFJ^{+}19]. parallelizable [Tor13]. Parallelization [MMVS19, SMC18]. Parallelizing [GWC^{+}16]. parameter [CQW^{+}19, SNC18, SIL^{+}13, SVN10b]. parameters [HLZ18, KSC^{+}19, NBB18, SWW^{+}13]. parametric [NGB18]. Pareto [PS19, TZST14, WWX^{+}17]. Pareto-based [PS19]. Parkinson [AAN^{+}18, ARP^{+}19, AASI17]. ParSA [ZSX^{+}15]. Part [MK16b, PZY16, PZY17, RVC16a]. Partial [WWT^{+}16, PBC^{+}16, PBC^{+}17, YCX18]. partial-load [PBC^{+}16, PBC^{+}17]. Partially [XZP^{+}19, HKS18]. participant [MLW^{+}18b]. participating [TDC^{+}14]. participation [CAB^{+}18, EKGS14, NZOCJ^{+}19]. participatory [LLGY18]. Particle [WCL^{+}17b, XRPT18, LAH10, XZW^{+}19, ZSMS18]. particularization [CPSRG14]. partition [JLY^{+}18, LZW^{+}18, RSJ^{+}14, TF18]. partition-based [TF18]. partitioned [DLXR14]. Partitioner [LZW^{+}18]. Partitioning [ATF11, kHsZwJW18, SLG^{+}17, LXD17, VSDD13, WWG^{+}19b, YWCC18]. partitions [SAC11]. party [ED19, KKB18, OSANAM19]. PASCAL [LMA^{+}19]. pass [CCD^{+}19, LY18a]. passing [Ciu10a, GSC^{+}19]. passive [BGC19b, CsZW14]. password [IOV^{+}18, LZS18, MZL^{+}19, ZXWA18]. password-based [IOV^{+}18]. Past [AGMT17]. PAT [WHS^{+}18]. patchable [GHYK18]. Path [AHEM17, Che13a, DLS14, IdAP19, KN10, NJHT11, SJR13, WWT^{+}16, ZWMC19]. pathological [DSM^{+}19]. paths [TWdLZ19, WW11, YXZG18b, ZCW19]. patience [dACNC16]. patient [AIB^{+}18, JNS^{+}19, RRS10, WCWC20, ZCDV19]. patient-centered [JNS^{+}19]. patient-optimized [ZCDV19]. patients [ABM19, DNW^{+}19, FHHM19, HEES19, HZX^{+}19, HZX^{+}20, MSA^{+}19, NDA^{+}19, VFHB14, WCWC19, WWA19]. Pattern [JHC18, dRADFG18, KZCW13, KPJ19, LHJC18, LY18a, MRS^{+}18a, PSS^{+}18, RAKJ18, THA^{+}17, TNY17, TSOB15, YL16, YNLY19]. Patterns [APS^{+}19, AM17, AVPV17, ABG17, AW19, CDG^{+}14, IHA18, LRJG19, LY17, LC15, MRL14, NAM^{+}19, PSK^{+}10]. PAU [FW19]. Paving [dRADFG18]. Pay [CDH^{+}19]. Pay-As-You-Go [CDH^{+}19]. payer [ZLL^{+}19]. Payload [VOV17, JKAU19]. payload-based [JKAU19]. payment [CLM^{+}14a, SYW17, ZWX^{+}19]. payments [HCW^{+}18]. PCP [WWT^{+}16]. PCP-B [WWT^{+}16]. PCRLB [LWX13]. PCRLB-based [LWX13]. PD [PKI^{+}18]. PDDRA [SR12]. PDE [BEWZ10]. peak [NJKF18]. peaks [LZZ19]. pearls [BBH18]. pedestrian [KVK^{+}18, RSY^{+}18]. peer [CCT13, ÇÖ13, KIC12, Li15, MROD10, PSJ^{+}12, PMMAM13, PRS12, PIP18a, PPLL17, SM10, WLP10, WLQ10, XKBA18, ZA13, FX10, HJC10, TPBS14]. peer-to-peer [CCT13, ÇÖ13, KIC12, Li15, MROD10, PMMAM13, PRS12, PIP18a, PPLL17, SM10, WLP10, XKBA18, ZA13, FX10, HJC10, TPBS14]. PeerAppear [CPLH19]. Pegasus [DVJ^{+}15]. pegylated [WCWC19, WCWC20]. pegylated-interferon [WCWC19, WCWC20]. penalty [GZZ^{+}18]. people [GMLGB^{+}17, GGMS18, KPG19, RMSPP17]. people-to-people [GGMS18]. per-application [XLL^{+}14]. perceived [XJY^{+}18]. perceived-constraint [XJY^{+}18]. Perception [KSK^{+}19, GGH^{+}19, LWH^{+}18, ZYW^{+}18]. perceptron [LMM19]. perceptron-based [LMM19]. percolation [Pal13]. percolation-driven [Pal13]. percutaneous [WWP19, WWP20]. Performance [AFSH^{+}18, AHL11, APS^{+}19, AAS^{+}19, BGI14, BDZ13, CHC^{+}17, DQC^{+}19, EPB18, ETR^{+}13, FBS18, HMW14, JS13, KI19, KJFS12, KSW^{+}13, LPS19, LRC^{+}18, LPD^{+}13, LZHY19, MLC^{+}11, MYBMM18, PIP18b, RPMG10, SI18, SG14, VRS^{+}19, WDJC18, WYJ^{+}19, Wri19, WXGM18, dOOO^{+}13, AkBAL^{+}19, AKB18b, ASB18, AEME^{+}18, BARMB14, BDA19, BFS^{+}17a, BFS^{+}17b, BC17, BGMLS17, BRHH18, CN17, CCRV13, CSG^{+}18, CW13a, CLP^{+}14, CS12, DCBF19, DPS16, DGCGH^{+}17, EP13, FSM^{+}18a, FPR18, GVDT16, GAFFOG12, GS13, GGJ13, GVI13, GSC11, GLNT13, GAMC19, GEAR13, GRMSOG18, GIK18, GJKP18, GG10, HAF^{+}16, HKPT10, JLC18, KDHP16, KMB^{+}17, KBVH14, Kha12, KPM^{+}18, KSC^{+}19, KCV11, KKW^{+}14, KS17b, KAEC^{+}18, LRYJ17, LSD^{+}17, LLT^{+}19, LHX^{+}18, LFH^{+}15, MJM^{+}16, MYW^{+}19, MAJD18, MBMTJR18, MKH13, MD12, MSBA16, MSM^{+}18a, MFL18, MDT^{+}18, MROD10, MWMA10]. performance [NNRA19, OG18, OKF10, PLZX19, PS10, PSL19, PMBS14, PKA19, PDDS10, QWCW19, RMJ^{+}18, RS16, RLL^{+}17, RKB18, RCTY19, SPMC10, SHN10, SGKC10, SEMJ11, SEPV19, SSI19, SSZ13, SBA^{+}17, SSP17, SCH^{+}19, SVN10b, Tao10, TMDZ15, TBdL16, TuIS^{+}19, Tur18, WJS^{+}18, WMY^{+}18, WZWC18, WHW17, XTF^{+}19, XDHL12, YZLQ14, ZG19, ZMP10, ZN12, ZYZ^{+}18, ZLTY10, ZZJ17, ZCQ^{+}16, ZYTC15]. performance-aware [KKW^{+}14]. performance-critical [XTF^{+}19]. performance-to-power [RCTY19]. performances [CGIP14, GIM16, SAH19, UNM^{+}16]. performing [SOD18]. perimeter [RR18]. periocular [TA19]. periodic [IHA18, SHP^{+}16, SMS14b]. periodic-frequent [IHA18]. permission [LXZ^{+}18, MRL14]. permissions [KKA18]. permutation [ABMESM18]. perpetual [RA12]. persistence [MEW^{+}19]. Persistent [CSYY18, ALL^{+}18, GHP^{+}18, ZSP17]. Personal [GTCZG^{+}18, Rao17, WLP18, ALL^{+}18, HCC^{+}14, KZA11, LWF^{+}17, XCZ^{+}19, BSRR18, LHL15]. personalisation [SCN^{+}14]. personalization [APR^{+}19, FHYH15, MVG18]. Personalized [CT19c, NWD^{+}18, ZCH^{+}17, BJ12, GJ18, PARMF14, WWH^{+}19, WMA18, WLS^{+}18, ZZJY16, ZSGJ19]. personnel [CLS19a]. perspective [BDH14, GVBG17, HQ10, TBR^{+}19, WWRS16, WZ16]. Perspectives [SBB^{+}10, AC10, KKvdB^{+}17]. Pervasive [AT18b, CMA11, RAA^{+}19, CPDJ13, CC11, Kol18, NDZ^{+}18a, NDZ^{+}18b, NDZ^{+}19, OB17, Pal13, SCN^{+}14, SMM^{+}14, ZGZ^{+}10]. Pessimistic [LM12]. pet [KS18d]. peta [LSH^{+}11]. peta-scale [LSH^{+}11]. petabyte [KLW^{+}18]. Petri [AEM10]. Petroleum [CFH^{+}19]. PFRF [LLpC12]. Phase [HKT^{+}19, PM14, BG12, FM10b, HZP^{+}14, HZ19, WH19]. phase-based [FM10b]. phase-difference [BG12]. phased [AWN^{+}13]. phases [KRZ12]. phenomena [CC19]. phenomics [PAC^{+}17]. PHESS [SA19]. PHFS [KIS11]. phishing [LYC^{+}19, SZ12, WFQ^{+}10]. phone [DYC^{+}18, HHH^{+}19, YWY^{+}17]. photodynamic [ZMZ^{+}19, ZMZ^{+}20]. photoelectric [XKJ^{+}18]. photography [DJJ^{+}18]. photonic [WdL16]. Photoplethysmogram [RAA^{+}19]. phylogenetic [OdOD^{+}13]. phylogenomic [dOOO^{+}13]. Physical [CFH^{+}19, GMM18, HLV^{+}16, LWW^{+}16, SHS^{+}19, ABD^{+}19, ABF^{+}15a, AAM^{+}16, BDE17, BK16, CM17, DWJM18, EAED18, EG18, GVBG17, GAW^{+}18, GHD19, GMP^{+}17, HAAR^{+}19, JHC10, KB16, LLS^{+}14, LSL^{+}15, LZY^{+}16, MMPF19, NLM^{+}16, OFD17, OA17, PKF14, PTD^{+}18, RVC16b, RVC16a, SMS14a, SZK16, Sko19, SZD^{+}17, SM18, VL19, WLZ^{+}16, XLZ18, YS16, APRC16, Zhu14]. physics [HHM^{+}19]. physics-based [HHM^{+}19]. physiological [GGH^{+}19]. PICADOR [BGP^{+}17]. PiCo [MDT^{+}18]. Picos [YÁJG^{+}15]. picture [LYZC15, RM19]. pictures [WM14]. PID [dSFD^{+}19]. PID-inspired [dSFD^{+}19]. piecewise [GZZ^{+}18]. pillar [RHH^{+}19]. pillbox [HZM14]. pilot [RMHCMG15]. pinning [KAEC^{+}18]. Pipeline [PPG19, KN10, ZZJ17, HRR^{+}14]. Pipeline-integrity [PPG19]. pipelined [PWMX17]. pipelines [MDT^{+}18]. pipsCloud [WMY^{+}18]. Pixel [XTL^{+}19, PDK10]. Pixel-wise [XTL^{+}19]. PKI [JLX^{+}19]. place [PWMX17]. placement [AD19, AK14, ACK^{+}15, BBB^{+}19, CRTN17, EBCP18, GLJ19, LLWW18, LWTL19b, LWZ18, LPBB^{+}18, MLG13, MNV12, MDD15, Pon19, SMBMT^{+}18, TZLL18, TMW^{+}17, TMMVL12, YPLZ17, YYLC10, ZWHC17, ZLL^{+}16]. Placing [HZ19, KN10]. plagiarism [AIP^{+}19]. plan [ABZK15, LWS^{+}12]. plane [CBBdL16, GZLZ16, HYG^{+}19, LLW^{+}19c]. Planning [WTM^{+}17, BA17, CMB17, CYJ19, DCC^{+}14, SSHC19, SA14, XCZ^{+}19, ZWMC19]. plant [PAC^{+}17]. Platform [GPS^{+}17, PECA19, SV16, AMQS^{+}19, ABB^{+}19a, AAA^{+}19, AAI^{+}19, CQW^{+}19, CVKB12, CGSV17, CSQL17, FPP^{+}18, FCD^{+}14, FSP^{+}18, GTMZ17, JNS^{+}19, JCA^{+}19, KSS19, KSC^{+}19, KGS^{+}19, KKS^{+}18a, KKKM17, KKKM18, LYJ10, LC14, LSH^{+}11, LXL^{+}19, LLW^{+}18b, MAD^{+}16, NV11, NKB19, NDZ^{+}18a, NDZ^{+}18b, NDZ^{+}19, PWB^{+}13, PPLL17, QZD^{+}18, SAGGB17, SA19, SBD^{+}18, TKA^{+}18a, TSGVRGS19, TCN^{+}14, TCB^{+}17, VF18, WQG15, WCC^{+}16, WZW19b, XFLL16, YWL^{+}17, ZZBP19, dSK^{+}19, CLM^{+}16]. Platform-as-a-Service [CLM^{+}16]. platforms [AUSA19, àCKPM19, CCT13, CA15b, DXL^{+}18, HA16, HA18, KMK^{+}14, LCH^{+}11, LOR^{+}18, LLW^{+}18a, LRZ^{+}18, LZX16, MCdA16, NNC^{+}19, PPPS18, PPA18, RML^{+}19, RB18, SB16, SBP^{+}17, SG15, ZCL^{+}14, ZZQ^{+}13, dSBN19]. play [dRRRR^{+}18, WXYL15]. Playing [Mér17]. PLC [XKJ^{+}18]. plot [ARP^{+}19]. plot-based [ARP^{+}19]. plug [dRRRR^{+}18]. plug-and-play [dRRRR^{+}18]. plumes [Sha16]. PM [XTL^{+}19]. PMC [BBC^{+}12]. PMC-based [BBC^{+}12]. PMI [TLC^{+}15]. PMSM [VS19]. Point [KPG19, CWL^{+}19, JXZ^{+}19, LWSY18, LW18b, SKF^{+}11, XWL^{+}18]. point-of-interest [JXZ^{+}19, LWSY18, LW18b]. pointer [KP12, KPG19]. pointing [MKS18]. Points [YZZC19, DNJG17, LZL19b]. Poisonedwater [WN10]. Poisson [LZLL18a]. PoKeMon [FZW^{+}18]. policies [ADBO18, BFP18, CXC^{+}18, DT16, EPB18, KDG^{+}19, Lea13, LHM14, MG14, MVC^{+}13, PZA18, SMSF18]. Policy [NCS12, Rao17, SME^{+}21, SVK19, AEK^{+}18, BBI13, DLH^{+}17, FS18, GCTLA^{+}19, HZL18a, Hua10, JFZL17, JSMG18, LJY10, LDZW19, MHZK18, QRW^{+}18, SME^{+}19, XZP^{+}19, XLL18b, LAL^{+}15, LHL15]. policy-based [GCTLA^{+}19]. policy-hidden [XZP^{+}19]. pollination [GHEB^{+}18, GHEB^{+}23]. polling [WWG19a]. polyadic [Bu18, GLC19]. polymorphic [AMM16]. Polynomial [CLS^{+}19b]. Polynomial-based [CLS^{+}19b]. Polynomials [TX14]. polysomnographies [KCH^{+}13]. pools [HS19]. POP [ZCDV19]. Popularity [WZML18, ZLL17a]. Popularity-based [WZML18]. population [KX11]. pornographic [LYXT14]. port [ESPN17]. portability [HA18, PSL19]. portable [MMC^{+}18, PMPC13]. portal [FFPS10, RMA^{+}16, FK11, FNA12]. portfolio [HKP10]. ports [CSL18]. Position [FZW^{+}18, YXA^{+}18, BSE^{+}13, ZF16]. position-based [ZF16]. positioning [CCC19, HDH^{+}18, LWW^{+}16, OMD^{+}18]. positive [WJLW18]. positives [LY17]. possession [Che13b, GZQ^{+}19, SYY^{+}17, YZN^{+}15]. post [BDMO11]. post-production [BDMO11]. posts [HO17, NO19]. posture [TMB^{+}19].

Potential [WZH^{+}19, ABD^{+}19, AAC^{+}19, KPS18, LZH^{+}18, MKS18, WLYL11]. Power [ADAAD12, AAM^{+}16, AEME^{+}18, CLL^{+}18a, JLCC12, KESL17, KAEC^{+}18, LLC14a, MP17, PZA18, SCEC18, TKA18b, XLL^{+}19c, AAQ^{+}19, BBC^{+}12, CT19a, CT19b, CRB^{+}16, DMC^{+}19, DDD18, DHC^{+}17, FAA^{+}18, FPL^{+}19, GRTV10, GVdBdL15, GFW^{+}18, HH19, JOPW14, JNR12, JKLK17, KCM19, LWW^{+}18, MMC^{+}18, NSSA^{+}14, Qur19, RMRSA19, RCTY19, VGC^{+}13, WZWC18, WWG19a, WOPW13, YZLQ14, YLHJ14, YGY^{+}19, ZAC^{+}18, vdLLE19]. Power-Aware [TKA18b, ADAAD12, AEME^{+}18, LLC14a, DDD18, JNR12, Qur19]. power-awareness [GRTV10]. Power-efficient [AAM^{+}16, JLCC12]. Power-performance [KAEC^{+}18]. power/rate [KCM19]. Powered [TAB^{+}18, CFMC19, GDS18, GCZ^{+}19, PSW^{+}14, ZZBZ19]. powerful [ABM19, BSE^{+}13]. PP [WZML18]. PPDP [ZZXL18]. PPFSCADA [FTA^{+}14]. PPG [RAA^{+}19]. PPQ [MTD18]. ppXen [ASB18]. Practical [ÇÖ13, KGT15, WOPW13, BGS^{+}19, HCL^{+}17, MdOO^{+}17, QMSG12, WDR^{+}19, ZSW^{+}18a]. Practice [CDFZ16, XDWL15, PMPC13, WGM15]. practices [KS19, TIHT14]. pre [SR12, ZCQ^{+}16]. pre-cache [ZCQ^{+}16]. pre-fetching [SR12]. precaution [AQRH^{+}18]. precedence [MRN19, TLL^{+}11]. precise [AMÇ19, GGN17, WHS^{+}18, HYF18]. precoded [KKB^{+}19]. precoding [ZWL13]. precompiler [BEWZ10]. preconditioning [BTM10]. Predatory [XRPT18]. predicate [FH13, HFT16]. predication [RGM^{+}19]. predict [AS19a, ABG17, CLP^{+}14, ZLTY10]. predictability [BVFGWA15]. predictable [MWPVB12]. Predicting [BDA19, CND^{+}19, MAJD18, MDB^{+}18a, SPSP17, DRC^{+}19, LLT^{+}19, WQG15, YWLL19]. prediction [AFSH^{+}18, AQAR^{+}18, ADA^{+}19, AMKM18, AK18a, BMR15, BZHV19, BBMG10, BRB19a, CHWW13, CGV10, CSL17, CAL^{+}18, CSL19, CGM^{+}18, CRWZ19, DMZ12, DdM10, DLS^{+}12, FEÁ19, GAW^{+}18, GSP^{+}17, GP11, GAA19, HPZL18, HUY^{+}19, HSV^{+}17, HHS^{+}18, IKLL12, JHC18, JLY^{+}18, Jun18, KXS^{+}16, KLJS19, KLV^{+}18, KS18c, Kyr19, LWSY18, LMCSE19, LCY^{+}19b, LJW^{+}19b, MZH^{+}17, MID16, MCT^{+}15, MPF^{+}16, MQL^{+}19, NWD^{+}18, PSI19, PNGFJ13, PTD^{+}18, RGAT18, RRKA19, RLP12, RCW^{+}19, SMRM13, SL11, SRP19, TYWZ18, VVB15, WHW17, WYH^{+}17, WYL^{+}18, XHL^{+}19, YLH^{+}19, YHH^{+}19, ZZXL18, ZZBP19, ZZBZ19, ZWYH19, ZXL14, FGG13]. predictions [CRTN17, NF13]. Predictive [BMP^{+}16, JYZ^{+}19, LBD18, ML19, NJH^{+}18]. predictor [XJY^{+}18]. preemptible [GdCP19]. preemptive [CLR18]. Preface [ARB12, KJ12, LH13a, LSHW17, Śle14, XKC13]. Preference [ZZC14, MTD18, SRP19, ZL18]. Preference-based [ZZC14]. preferences [BCF16, LLGY18, MBL^{+}19, OAMS18]. Preferred [LZY^{+}19b]. prefetching [WZWC18]. prefix [DZLA19]. pregnancy [MRS^{+}18a]. premiere [MBC^{+}11]. preprocessing [MCSA18, RHKC15]. Preprocessor [VOV17]. prescription [HIA^{+}18b]. presence [KDHP16, MFT^{+}17, RBLvM14]. present [AGMT17]. preservation [KRZ^{+}19, PLW^{+}19, SYJA19, URC19, XXQ^{+}19]. preserved [CRYG18, ZCL^{+}19]. Preserving [TSOB15, YDNV16, ALL^{+}18, CZ12, DHW^{+}17, FTA^{+}14, FH13, FXG^{+}19, FLL^{+}19, FRZ19, GQXL18, GAI^{+}18, GSR^{+}19, IOV^{+}18, JCMPPC^{+}18, JYY^{+}17, JLC18, JLX^{+}19, KKB18, KC19a, KLMB19, KH18b, LLC^{+}14b, LCL^{+}16, LLH^{+}17, LYY^{+}18, LLY^{+}18, LXMW15, LLM^{+}16, LDY^{+}18, LNY^{+}18, LLAW17, MML^{+}18, MQL^{+}19, NK17, NLS19, PSY^{+}19, QZD^{+}18, QGX18, RBA17, SYY^{+}17, SLW11, SLL^{+}17, TAS^{+}18, VCD^{+}18, Wan18b, XCS^{+}18, YLN15, YZG^{+}18, YZL^{+}19, YCXW18, YHH^{+}19, ZYK17, ZZXL18, ZZX^{+}19, ZCDV19, ZRZ^{+}14]. pressure [RAA^{+}19]. pretense [NNC^{+}19]. preterm [PSI19]. prevalence [RAKJ18]. prevent [LTN10]. Preventing [WTTH19, SM18, TWG^{+}19]. prevention [HA19, KIAD17, SMF^{+}19]. prevention-based [SMF^{+}19]. preventive [PPMM^{+}18]. Price [CLRL18, FEÁ19, GGLW18, JTB13, JHC18, SSHC19, WBR19]. pricing [ADA^{+}19, DVB14, DEG^{+}17, JBR^{+}16, RA12, YLJL18, YVCB10, ZXZL18, GVA^{+}16]. primitives [BH13]. principal [MZYA19]. Principles [VRGR16, GS15]. printed [YCZJ18]. priorities [GS13]. prioritization [SRN^{+}18]. prioritized [YH19]. priority [AS18a, ÇBCA15, DSCJ18, VAdlP12, ZZZC19]. priority-based [ZZZC19]. Privacy [AWYJ16, ACC^{+}19a, ABB^{+}19a, ALL^{+}18, CBC^{+}19, CRYG18, FTA^{+}14, FRZ19, GQXL18, JYY^{+}17, JLX^{+}19, KH18b, LLC^{+}14b, LYY^{+}18, LLY^{+}18, LLGY18, LLAW17, MHZK18, MQL^{+}19, NK17, PMBS14, PSY^{+}19, QGX18, SAK19, TAS^{+}18, URC19, WSN18, XDWL15, YDNV16, YZG^{+}18, ZYK17, ZWS^{+}12, ZCL^{+}19, ACL^{+}18, AdVAGF18, AIA^{+}18b, AM10, CD16, DA18, DGY^{+}18, FH13, FXG^{+}19, FLL^{+}19, GAI^{+}18, GBKJ18, GLD^{+}19b, GSN^{+}18, HAJ^{+}19, HHW^{+}19, HHK^{+}16, HPL^{+}19, IOV^{+}18, JCMPPC^{+}18, JLC18, KKB18, KYB^{+}19, KRZ^{+}19, KC19a, KLMB19, LCL^{+}16, LLH^{+}17, LYC18, LLW^{+}19a, LXMW15, LLM^{+}16, LSHW17, LDY^{+}18, LNY^{+}18, LRBW17, MML^{+}18, NJB19, NWMG17, NLS19, PARMF14, PLW^{+}19, QZD^{+}18, RHH^{+}16, RBA17, SYJA19, SYY^{+}17, SCL14, SCZ^{+}14, SLW11, SCH^{+}17, SLL^{+}17, TBR^{+}19, VCD^{+}18, Wan18b, WZE19, XCS^{+}18, XXQ^{+}19, YLN15, YZW^{+}18, YXA^{+}18, YZL^{+}19, YCXW18, YKÖ17, ZCLW18, ZZXL18, ZCYZ18, ZLT^{+}19, ZZX^{+}19, ZZY^{+}19, ZCDV19, ZRZ^{+}14, FW19]. Privacy-aware [MHZK18, DA18, SCH^{+}17]. Privacy-based [LLGY18]. Privacy-friendly [ABB^{+}19a]. privacy-preservation [KRZ^{+}19]. Privacy-preserved [CRYG18, ZCL^{+}19]. Privacy-Preserving [YDNV16, ALL^{+}18, FRZ19, GQXL18, JLX^{+}19, KH18b, LLC^{+}14b, LLY^{+}18, LLAW17, MQL^{+}19, PSY^{+}19, QGX18, TAS^{+}18, YZG^{+}18, ZYK17, GAI^{+}18, JLC18, KC19a, KLMB19, LCL^{+}16, LLH^{+}17, LXMW15, LLM^{+}16, LDY^{+}18, LNY^{+}18, MML^{+}18, QZD^{+}18, SYY^{+}17, Wan18b, XCS^{+}18, YZL^{+}19, YCXW18, ZZXL18, ZZX^{+}19, ZCDV19]. Privacy-protected [WSN18]. privacy-protecting [CD16]. Privacy/performance [PMBS14]. Private [RBLvM14, BDCC19, CLL^{+}14, FDPR17, GSLI12, HYF18, KKL11, PMBS14, SGKC10, WLL^{+}19a, ZLXZ18]. privatized [SB17b]. PrivBox [ABH18]. privilege [QRW^{+}18, XWRZ19]. PRNU [VOCHC17]. PRNU-based [VOCHC17]. Proactive [GHYK18, LW18a, NHH^{+}19, QGT^{+}18, WZML18, ABDH19, CdRRdCB19, KAW12, LMA^{+}19, NJH^{+}18, SC19]. Probabilistic [AB19c, RT16, SGBK19, Bag16, Bag19, EMM12, LZJL19, SB18, WPY19]. probabilities [DK14]. Probability [HLC16, LYYY18]. Probability-based [HLC16]. Problem [WH19, ABMESM18, ABMMC18, ABMMC22, AK18a, BZMY10, BP13, DV13, DC19, DSCJ18, GMMM18, HXWW18, KPS18, OB19, SJR13, SK18, SV15, TV16, WLH16]. Problems [SSG17, CMT16, FFPS10, GOLL19, LPS19, MGMT18, SJV12, WWX^{+}17, XA10, XRPT18, YDD^{+}18, ZRZL18, ZTKF17]. procedure [AS19a, LYS12]. procedures [RB12]. Process [ANG^{+}19, SJV^{+}15, BGRBA19, CRVZ15, CAS^{+}18, DLZ^{+}14, EG18, HXC^{+}18, JXZ^{+}19, KYZ19, LZLL18a, LWS^{+}12, LHW^{+}18, LCCM18, MKM11, PTD^{+}18, dRRdQGR^{+}18, SSSJ19b, TIHT14, WZ16, YSC^{+}19, YJL^{+}19]. process-based [LHW^{+}18]. process-independent [EG18]. processes [Bag16, BBT19, BKB18a, CHSA18, LDJL19, MAB^{+}15, WMLS14]. Processing [CLNR18, HCB16, LJ17a, MLBS11, WXGM18, YZI18, ATH^{+}19, ADLM18, dRADFG18, BKB11, CQW^{+}19, CMX^{+}16, CLZ18, CLDC19, CGM^{+}18, CBBdL16, EMHE18, FPR18, GXW^{+}19, HB19, HRVW18, Hsu14, HZ19, JSZ^{+}19, KKI14, Kim18, KVHT10, LZY^{+}19a, LKM14, LGW^{+}17, LHW^{+}18, LRC^{+}18, LJW^{+}19a, LWR^{+}19, MLC18b, MTD18, MAPA19, MDM^{+}19, OMKM^{+}19, PKY^{+}17, PPPS18, PPS^{+}18, QNM^{+}19, STC15, Śle14, SSL13, SRN^{+}18, TWZP18, TDBR18, VEET18, WPGN^{+}18, WZW^{+}19a, WWCN13, WMY^{+}18, WZY^{+}19, WdL16, WLW^{+}19, WWZ^{+}19, XB14, XTF^{+}19, YFY^{+}13, ZYB^{+}18, ZFS^{+}18, ZQB^{+}18, ZSBB19, ZSP17]. Processing-as-a-Service [HB19]. processor [GS15, MRN19, MW12, WWZC19]. processors [ASW11, ADH^{+}16, BVFGWA15, GVA^{+}16, KR14, Li18, LG16b, PAB^{+}14, WSH^{+}16, WLRL18, YDT19]. procurement [VVB15]. Producing [Wei11, LHJC18]. product [SPS18, WHCZ18, XLL^{+}19b]. production [BDMO11, CFH^{+}19, EG18, GSV^{+}10, LLC^{+}16, MSS^{+}13, RGM^{+}19, RSRA18, RWV^{+}13, YMW^{+}18]. productive [IHA18]. products [BAB13, GGLW18, KHG^{+}18, VK17]. products-Towards [KHG^{+}18]. professionalized [ZCH^{+}17]. ProFID [ÇÖ13]. Profile [Qur19, VTTK17, CLM^{+}16, LLAW17, SL11]. Profile-based [Qur19, VTTK17]. profile-matching [LLAW17]. profiled [RAKJ18]. profiles [PARMF14]. Profiling [ADLM18, TWdLZ19, EET18, JCD^{+}13, LZHY19, PFS^{+}13, SL11, YZLQ14]. Profit [AB17, AB18c, SSJ19, CBLS13, GBF^{+}12]. Profit-aware [AB17, AB18c]. profit-driven [GBF^{+}12]. prognostics [PLLP19]. program [ASV^{+}13, LCH^{+}11, MWYC12, TBK^{+}10]. Programmable [BRH18, XLL18b]. Programming [Lok12, ADAAD12, Bal16, CMS^{+}18, DSCJ18, GQLX18, GdVC10, GGA^{+}17, JLQ18, NF13, SFR15, WWC14, WM14, WHYZ17, YLG^{+}16, YDQC19, ZME^{+}15]. programs [CAB^{+}18, TDC^{+}14, WWZC19]. Progress [WGX^{+}19, Ano19k, Ano19l, GK18, MAÇ17]. progression [DRC^{+}19]. project [ALFR16, DSCJ18, FEB^{+}19]. projection [BWG19, CGN18]. projects [KS19]. Promises [FFC^{+}18]. promote [Mat18]. prone [AGKZ18]. Proof [LTC^{+}19, LNY^{+}18, SOM^{+}19]. Proof-of-Activity [LTC^{+}19]. Propagation [CsZW14, SKS^{+}18, FX10, LNB14, OA17, WWSL19, WTTH19, XHL^{+}19]. proper [SMBMT^{+}18]. properties [DA16, KJI11, SAM^{+}19]. property [YSZW18]. proportions [kHsZwJW18]. proposal [SCAC^{+}19, dCRL^{+}19]. prospect [CPE^{+}17]. prospects [URC19]. protected [AGBR19, WSN18]. Protecting [CYW^{+}19, Sip12, SSB13, CD16, HHW^{+}19, TBR^{+}19]. Protection [AIM^{+}19, AIA^{+}18b, ELAEAVAM19, HZW^{+}18, HHXL13, HAAR^{+}19, HPL^{+}19, KYB^{+}19, LTJK12, LLW^{+}19a, LZL^{+}12, PvSS17, RS17b, TMJH19, WDJC18, XWRZ19, ZZY^{+}19]. protein [CGV10, DNJG17, MMVS19, SSLF^{+}10, VR12, ZZBZ19]. proteinuria [QXZ^{+}19]. Protocol [GLXF17, AMKC19, AGR19, AQRH^{+}18, AT19b, AKB^{+}18a, ASAA18, AH11, BP13, BRB19a, BGNI19, CHJS^{+}10, FG18, GCK18, GZL^{+}18, HZW^{+}18, IOV^{+}18, JBM^{+}18, KYB^{+}19, KC19a, LCL^{+}19, LSD11, LWW^{+}16, LWL^{+}18, LEW19, LM12, LLAW17, MLC^{+}18a, MZD^{+}16, RACA18, RRU^{+}18, SCS^{+}18, SL19, SAVS19, TKA18b, TZD^{+}19, Wan18b, WDZ19, WMA^{+}19, XZZ^{+}14, ZXW^{+}18, ZYW^{+}18, ZXWA18, JL14]. Protocols [ZYA^{+}18, BSRR18, CJG^{+}18, DZH18, ELAEAVAM19, GBKJ18, LTC12, RC18, RC19, RMDB18, TLSC17, TVV13, TAS^{+}18, URKM19, YH18, ZGS^{+}13, ZAI^{+}18, dCRL^{+}19]. prototype [Ans11, WWVJ17]. prototyping [GMD19, XWW19]. prova [GMB19]. provable [GZQ^{+}19, WXYL15]. Provably [ODK^{+}17, WWW^{+}16, Wan19, ECE^{+}19, KLW^{+}17, TX14]. provably-secure [ECE^{+}19]. Provenance [GCM^{+}11, GM11, Mil11, MCF^{+}11, SB11, ATdC^{+}16, ABG17, Asu13, CPA14, CDL18, CMD^{+}14, DCMB15, DMMM11, FKOC11, GSR^{+}19, HMM18, IHK^{+}18, LCHW14, LLCF11, LFH^{+}15, MSS^{+}13, MG11, SGM11, WHW17, XFTZ16]. provenance-based [GSR^{+}19]. provide [dSGD19]. Provider [GVdBdL15, BMU16, CLM^{+}16, GBRM18, IGB^{+}14, PRSR14]. Providers [ABTA18, FMN^{+}17, SS17, AM19b, CMB17, CAC^{+}15, CBLS13, FG14, LLY^{+}18, MG14, SSRQ19, SKS17, TMMVL12, YDQC19]. Providing [CCCT14, MLM16, BBC^{+}17]. proving [RB12]. provision [FWB13a, GPVN19, Mat18, WLZ^{+}16, ZZH^{+}16]. provisioned [ABN17, MG10]. provisioner [JNR12]. Provisioning [MMVP13, MBV^{+}15, AK19, AAAQJ^{+}18, Ano12r, BKKM11, CVKB12, CVT19, CFVP12, CBS17, FEÁ19, FSM^{+}18a, GAJP18, HC17, HSBE19, HB19, IDCJ11, IKLL12, ISS^{+}15, KPA17, KMK^{+}14, Len16, LSYC18, MJDN15, PDH18, PS19, RML^{+}19, RBN13, RP18, dRRRR^{+}18, SBAD^{+}18, SGJ18, SPSP17, SSKK13, SLB^{+}17, SLZ^{+}18, TKR^{+}15, TSB18, VHML11, VCKB12, VOS12, XTT18]. proximal [HHK18]. proximity [Alp18, GDJ^{+}13, LLAW17, WWVJ17, YCXW18]. proximity-aware [GDJ^{+}13]. Proxy [LAL^{+}15, BGP^{+}17, HCZW17, HYF18, LL16, PAL^{+}19, TX14, Wan18a, ZDW^{+}16]. pruning [YK17]. PS [XRPT18]. PS-CTPSO [XRPT18]. Pseudo [LZCX19, RNR18]. Pseudonym [ACC^{+}19a]. pseudorandom [BGC19b]. PSO [LZL^{+}19a, SGRT19, ZWL13]. PSO-based [SGRT19]. PSO-SVR [LZL^{+}19a]. psychology [AC16]. PTAS [JLX^{+}19]. Public [GMM18, WSQ^{+}16, AAA^{+}19, CC19, DDMPG17, JTB13, LMZ^{+}14, LLW^{+}18a, LH13b, LL16, PPPS18, SCCS11, SCEC18, TVB18, TCH19, VGD^{+}19, WWDF18, WHS^{+}18, ZWQ^{+}19, dOOO^{+}13]. public-resource [SCCS11]. publication [GB10, TSTL16, WSN18]. Publish [BGP^{+}17, AMPZ16, EBCP18, GGC17, MWQ^{+}14, OKF10, XWJ^{+}16]. publish/subscribe [AMPZ16, EBCP18, MWQ^{+}14, OKF10, XWJ^{+}16]. published [PFS^{+}13]. publishing [FTA^{+}14, PSY^{+}19, SLW11, ZWS^{+}12]. PUF [BDL^{+}19]. PUF-based [BDL^{+}19]. Pulse [OMPSPL^{+}19]. purchase [GGLW18]. purpose [BRNR15, GXW^{+}19, VVB11]. pursuit [YZWG18, YSZW18]. pvp [WWP20, WWP19]. PWLM [TMDZ15]. pyramid [DGA18]. Pyxis [vKvWD^{+}13]. Pyxis-DT [vKvWD^{+}13].

Q [KBVH14, NUPA19]. QaMeC [WLH^{+}19]. QCG [ACH^{+}11]. QCG-OMPI [ACH^{+}11]. QFD [ABMMC22, ABMMC18]. QoS [AB18c, BFS^{+}17a, AJY12, AJY15b, ACCD17, AMT^{+}12, AB17, BFS^{+}17b, CVKB12, CCIP18, CM17, CAC^{+}15, CSL17, CSL19, CCCT14, CRB^{+}16, FG18, FD12, FA11b, GMMM18, GHMX10, GJGB19, GJF^{+}12, HB19, HYX^{+}19, HMW14, JS12, JGFB18, KC14, KARP14, LD17, LCL14, LWS^{+}12, LLW^{+}12b, LHPC^{+}19, MCL^{+}16, MZH^{+}17, MCT^{+}15, OWX19, PSJ^{+}12, PNZ14, QCY^{+}19, QCD16, SC16, SSKK13, SLS^{+}19, SK19, VAdlP12, VSKS19, WLZ^{+}16, WPY19, WZ13, WYH^{+}17, WYL^{+}18, WLH^{+}19, XTT18, ZFW14, ZWYH19]. QoS-aware [KC14, ACCD17, AMT^{+}12, CCIP18, CCCT14, FG18, GMMM18, GJGB19, JGFB18, MCL^{+}16, SK19, WLZ^{+}16, WPY19, WZ13]. QoS-based [FA11b, SSKK13]. QoS-driven [BFS^{+}17a, AJY12, BFS^{+}17b, CVKB12, WLH^{+}19]. QoS-enabled [QCY^{+}19]. QoS-guaranteed [XTT18]. QoS-oriented [HYX^{+}19]. QR [SGS^{+}18]. QRD [ZWL13]. QRD-based [ZWL13]. QRE [WSZH18]. QRFence [SGS^{+}18]. QTI [DWJM18]. quadtree [ZWL^{+}16]. Quake [BSE^{+}13]. Quake-Catcher [BSE^{+}13]. Quality [AJY15b, HB19, HXL^{+}18, JY15, MCR^{+}16, Sta17b, VOS12, ZWJ^{+}19a, ACSV18, CMG^{+}19, Che13a, DWJM18, GAFFOG12, GGA^{+}17, KSW^{+}13, LWS^{+}12, RHH^{+}19, SMPC12, WdL16, WLW^{+}19, WPS^{+}18, Zin18, dSGD19]. quality-assurance [WLW^{+}19]. Quantification [LKCS18]. Quantified [LG16a]. Quantifying [CN17, ZTL^{+}19, AB19a, ABGMC19, AB21]. quantitative [CRWZ19, LWW^{+}18, XAW^{+}10]. quantization [CCZ^{+}19]. Quantum [QCX18, ELAEAVAM19, LCW^{+}18, LSV^{+}18]. quantum-based [ELAEAVAM19]. quantum-induced [LSV^{+}18]. quantum-inspired [LCW^{+}18]. queriable [SJTN18]. queries [ATS14, ARP14, CLR16, JLD^{+}19, KYB^{+}19, LXZ^{+}18, MLW^{+}18a, SLC^{+}17, YAO14, ZWZ18, ZLXZ18]. Query [MVG18, ZMP10, JLC18, LZY^{+}19a, LXK^{+}14, LRMS19, MCL^{+}16, MTD18, MW12, PdAF12, QNM^{+}19, WZW^{+}19a, YFY^{+}13]. Querying [NS17a, Ans11, LZL^{+}16, LLCF11, SB17b, ZSZ14]. queue [AS18a, RS17a, ZA14]. queuing [KDG^{+}19, Pon19, YZ12]. quick [KGLY18, WSZH18]. quick-response [KGLY18].

rack [LFP^{+}17]. radar [KS17a, LXT^{+}19]. radiation [GZZ^{+}18]. radicalisation [LCGPC19]. Radio [TKA18b, ASAA18, BMK^{+}14b, CdRRdCB19, LBM18]. RAID [WZML18]. rail [LSZ^{+}18]. railway [CLZ18, GCCL18]. Random [LZL^{+}12, ArMS19, BYL^{+}18, HHW^{+}19, LXD17, LXM^{+}18, MLW^{+}18b, RNR18, ZSFZ19]. randomized [PdASM18]. Range [STA17a, CLR16, CSL19, LWX13, PM14, SJTN18, WZW^{+}19a]. Range- [STA17a]. range-based [LWX13]. range-queriable [SJTN18]. Rangelands [TSTL16]. rank [BTG19, HZC10, KRZ12, RS16]. ranked [LXK^{+}14, NJB19, YQZ^{+}19]. Ranking [YG18, DFG^{+}19, DA16, GVB13, LWD^{+}14, LYW^{+}18a, MCT^{+}15, SKS17, SRKS18, WN10, ZSJ19, ZSGJ19]. ransomware [ArMS19, HDA^{+}19, ZXM^{+}19]. Rapid [AM19b, GMD19, WWCN13, SB19b, XWW19]. RapIoT [GMD19]. RAR [HLL^{+}11]. rate [HJA^{+}19, KHJ10, KCM19, KMJ18, SPT^{+}18, ZXW19]. rather [LSAM13]. rating [DV13, TQL^{+}19, ZZBP19, vKLA^{+}19]. ratio [RCTY19]. Raw [SLC^{+}17]. ray [DRZ^{+}19, SYT^{+}19]. RBAC [LXZ^{+}18]. RConf [PKI^{+}18]. rCUDA [IPG^{+}18]. RDF [DA16, Gra15]. RDMA [ABF^{+}15a, LBM18, WHZ19]. RDMA-enabled [ABF^{+}15a]. re [BGP^{+}17, DEL19, HYF18, JCL^{+}15, LYW^{+}18b, LL16, PFRC16, WHZL10, Wan18a, ZDW^{+}16, LAL^{+}15]. re-decomposition [WHZL10]. re-encryption [BGP^{+}17, HYF18, LL16, Wan18a, ZDW^{+}16, LAL^{+}15]. re-enrollment [DEL19]. re-fusion [LYW^{+}18b]. re-optimization [PFRC16]. re-outsourcing [JCL^{+}15]. read [IDCJ11, TMDZ15]. ready [MMC^{+}18]. Real [AKG^{+}17, BKB11, CDH^{+}19, DGD^{+}16, EET18, HKU^{+}11, HEES19, Kim18, LRZ^{+}18, MG18, NJ17, OKF10, RPA^{+}18, VSBN19, AB19b, BMP^{+}16, CPGBC16, CXZ^{+}19, CCT13, CRRC18, CFF14, FAMA^{+}17, GVURIVBV14, GKW^{+}12, GRX19, HJA^{+}19, HNCJ13, JOPW14, LWZ^{+}19a, LJ19b, LG16b, LHY^{+}19, MPCAF15, MGA^{+}18, MOFGP18, MRN19, OPT^{+}17, PKF14, PTD^{+}18, PWP^{+}18, PGTBC18, SK12, SK19, TSRG17, TCCW19, WQG15, WSH^{+}16, WOPW13, Wu16, YAO14, YNSM12, YCL^{+}19, ZMTT16, ZWJ^{+}19a, ZCK^{+}15, ZSP17, Zin18]. real-life [JOPW14, LWZ^{+}19a]. Real-Time [MG18, AKG^{+}17, BKB11, CDH^{+}19, DGD^{+}16, EET18, HKU^{+}11, HEES19, Kim18, LRZ^{+}18, NJ17, OKF10, RPA^{+}18, VSBN19, AB19b, BMP^{+}16, CXZ^{+}19, CRRC18, GVURIVBV14, GKW^{+}12, GRX19, HNCJ13, LG16b, LHY^{+}19, MPCAF15, MOFGP18, MRN19, OPT^{+}17, PKF14, PTD^{+}18, PGTBC18, SK12, SK19, TSRG17, TCCW19, WQG15, WSH^{+}16, Wu16, YNSM12, YCL^{+}19, ZMTT16, ZWJ^{+}19a, ZCK^{+}15, ZSP17, Zin18]. realistic [CPGBC16, KMST19]. reality [GGH^{+}19, LLS^{+}19, RMSPP17]. Realizing [YH19]. rear [PWP^{+}18]. rear-end [PWP^{+}18]. Reasoning [TVV13, DKFKF18, LLCF11, NS19, SCN^{+}14, ZGZ^{+}10]. reassignment [LXZ^{+}18, STMV18]. recall [HMMW19]. received [AKM18]. Receiver [WCWC20, WCWC19]. rechargeable [CZM^{+}18, YHL^{+}19]. Recognition [GTEL^{+}18, SLTK19, TCCW19, BG12, Bae14, CPP^{+}18, CZ12, EU19, GPJA^{+}14, HLZ^{+}19, HUMA18, HLL12, HLT^{+}18, HMMW19, IFD^{+}19, JLQZ18, KAP19, KSS19, Kha12, KPG19, LLSL18, LZL19b, MCRB19, OCW14, PSS^{+}18, RSY^{+}18, RM19, SCZ^{+}19, TJ18, UMUB19, YPCK12, YXY18]. recognizing [ZDL^{+}19]. Recommend [GGLW18]. Recommendation [GCCPGBGS10, RMA^{+}18, WLL^{+}19b, ABZK15, AK18a, CQW^{+}19, CZL^{+}18a, CCJ16, CHY^{+}18, FLR^{+}16, GJ18, GGMS18, JTL^{+}19, JXZ^{+}19, KZA^{+}18, LCH^{+}11, LCW^{+}18, LZH^{+}18, LLW^{+}19b, LXMW15, LQLX10, LLGY18, LW18b, LWXY19, LSV^{+}18, MML^{+}18, MLW^{+}18b, QZD^{+}18, RW18, WWH^{+}19, WMA18, dOWdAS^{+}18, XLL^{+}18a, XZ16, YWL^{+}17, ZZJY16, ZSGJ19, ZL18]. recommendation-based [CQW^{+}19]. recommendations [SDK19, ZCH^{+}17]. Recommender [EGVT18, YL18, AMPP19, DV13, KKB18, MQL^{+}19, TNY17, TAKV12]. Recommending [ACCM19]. reconfigurable [ASW11, FFC12, HZDS19, LKTC14, YP12]. reconfiguration [EG18, GHY^{+}18, GVURIVBV14, MNV12]. reconsolidation [SHRE16]. reconstructed [AD18]. reconstruction [LLU^{+}18, WWA19, YSZW18]. record [CLC^{+}19, BSRR18]. recordable [LWW^{+}16]. Recording [FM10b]. Records [Rao17, AIM^{+}19, LZLL18b, LHL15]. Recovery [SPJ17, WZML18, BNJ16, DYY^{+}19, LPMY18, LZS18, MSI^{+}12, MDB^{+}18b, QMCX19, RRU^{+}18, SA14, TSAER18, YIA17]. recruitment [SOM^{+}19]. recurrence [ARP^{+}19]. Recurrent [HDKC18, HUY^{+}19, RCW^{+}19, AAYL19]. recurring [ZWZ18]. recursion [BBH18]. redesign [WLB11]. reduce [BEWZ10, Gra15, LLZ^{+}19, MVL18b, SCCS11, TJZ^{+}15]. Reducing [FEÁ19, SYQ^{+}19]. reduction [FdAGdAFV19, MZYA19, VVC^{+}12]. redundancies [LYJ10]. redundancy [AMR18, DZZ^{+}15, PJDO13, ZHL^{+}18]. redundancy-based [DZZ^{+}15]. reefs [FEPC18]. Refactoring [KTB18]. Reference [CBS17, RCOP^{+}11, SCAC^{+}19, XWX^{+}17]. refined [Bag19, WPY19]. Refinery [CSJ^{+}17]. regarding [MLGGB^{+}17]. regeneration [ZYCZ19]. Region [ST11, LXRS19, LZT^{+}19, LZL^{+}12]. region-based [LZT^{+}19]. register [MWYC12]. registration [XPL19]. Registries [SZK18]. Regression [CAPG18, FSM^{+}18a, RRKA19, AQAR^{+}18, ZZC18]. regular [LSG18, LKTC14, THA^{+}17]. regular-topology [LKTC14]. regularization [CWJD19, QCZH19]. regulating [PYH17]. regulation [LTN10, MSS^{+}16]. regulations [SDK19]. regulatory [HHL11]. reinforcement [BW19, LYYW19, NLV^{+}19, QCY^{+}19, SHL^{+}19b, WXZL11, YYW^{+}19, ZZ19]. related [FPR18, GAFFOG12, HXA^{+}17, LKK^{+}16, PKA19, TBS^{+}18]. relation [LLS^{+}14]. relational [GVDT16, LLCF11]. relations [NS19, XLZ^{+}14]. Relationship [YTQ19, BFP18, FPL^{+}19, HCC^{+}14, WLLF16, YZW^{+}18, YTQ20]. relationships [SM10, SCZ^{+}19, YWY^{+}17]. Relaxation [KJI11, MCL^{+}16]. relaxing [KJI11]. relay [DPK^{+}19, ROK19, ZWL13]. relaying [SYJA19]. release [JLL17, WLL^{+}19a, ZLXZ18]. Releasing [WZE19]. Relevance [BBD^{+}19, BCdV^{+}19, PC17]. Relevance-based [BBD^{+}19, PC17]. relevant [NHH^{+}19, SHL^{+}19b]. Reliability [KCK16, AEM10, GIM16, KHG13, KH18a, LYJ10, MSO18, PRL^{+}19, SYQ^{+}19, VK17, WYBS11, ZTC^{+}19]. Reliability-driven [KCK16]. Reliable [GVBG17, KIMR15, CEP19a, CBT^{+}19, DKV14, GSC^{+}19, JCL^{+}15, KB18, LZ10, LPK17, LPK18, LXMW15, LSL^{+}18, MCAS19, MRH17, MKM11, RWO^{+}19, TAS^{+}18, Tor13, VSBN19, WQG15, WHS^{+}17, WLP18, YW12]. relieve [KPS18]. remaining [RCW^{+}19]. remapping [AS14, LJL12]. REMEDI [SZG^{+}19]. remediation [KAW12]. remnants [QC13]. Remote [MWW^{+}15, MDA^{+}19, OPT^{+}17, SYY^{+}17, WXZ^{+}18a, YZN^{+}15, AASI17, BDZ13, CCD^{+}19, CHS11, DCC^{+}14, FCD^{+}14, GHEB^{+}18, GHEB^{+}23, GZW18, HKU^{+}11, HEES19, LWK^{+}18, LH13b, OPO13, SKF^{+}11, SWW^{+}18, WWCN13, WMY^{+}18, YMW^{+}18]. remotely [XAW^{+}10]. removing [CLS^{+}19b]. Renal [QXZ^{+}19]. renewable [GDS18, TuIS^{+}19]. renewed [MDA^{+}19]. Renovating [MQN19]. reordering [WC14]. reorganization [SRN^{+}18]. repackages [MH19]. repackaging [GMCM16, GMCM18]. repair [GDR^{+}14]. repartitioning [KMB16]. RePCoN [PLA18]. replaceability [ZGS^{+}13]. replica [AWN^{+}13, LWTL19b, LWTL19a, MVCC10, MKRD19, SYL18]. replicated [HKS18]. Replicating [MBV^{+}15]. replication [ASD12, AMT^{+}12, CRVZ15, CTR^{+}17, GS16b, GIM16, HNKÖ18, KKAS19, KIS11, LLpC12, LJY12, MLG13, MPC^{+}18, PSJ^{+}12, PGCC^{+}10, PB18, SR12, ST11, SYQ^{+}19, TZBK13]. replicative [SRN^{+}18]. repositories [ARP14, NAM^{+}19, SCBK^{+}16, SSLF^{+}10]. repository [BDG^{+}19, GHJ^{+}19, Ham17]. representation [BCdV^{+}19, BCF16, CPA14, sGbKS19, Ima19, LZL^{+}12, LLSL18, ML19, PBL^{+}18, SHS^{+}19, XLL^{+}18a, XJY^{+}18, XLL^{+}19c, YJY^{+}18]. Representing [GM11, AHL11]. Reprint [AB18c, GMCM18, KKKM18, LPK18, PBC^{+}17, ZFC18]. reproduce [HMM18]. Reproducibility [SPdSR^{+}17, BCG^{+}19, CBBC^{+}17]. Reproducible [GMB19, QWCW19]. Reputation [PS10, ABH18, FMRS18, GA13, GL19, LXMW15, LHX^{+}18, LLJ^{+}11, PVGD^{+}19, RWY^{+}18, TY11, TMP15, WN10, WYBS11, ZZH^{+}18, ZCZ^{+}18]. Reputation-based [PS10, PVGD^{+}19]. request [LLYW19, YZ12]. requests [HFM19, SYAL13]. requirement [SMRM13]. requirements [GDP^{+}18, HSS17, LLW^{+}12b, MKS^{+}19, QCD16, YHA^{+}19, YS16]. rerouting [RRU^{+}18]. resampling [HLZ^{+}19].

Rescheduling [LZ10, dRRdQGR^{+}18]. Research [ACC^{+}19c, CCIP18, CXZC18, CZH^{+}18, CDFZ16, CSC18, CMZ^{+}18, DJJ^{+}18, HGM15, LBJ^{+}18, LBJ^{+}24, LGW^{+}17, MYHZ18, VCE^{+}19, WWZZ18, WZ18, WWZ18, YLL^{+}19, YKÖ17, AZH18, BOWD^{+}19, CLM^{+}16, CLCY18, CDB^{+}19, CGL^{+}10, CMD^{+}14, FEB^{+}19, GSGPP^{+}19, GML^{+}13, GHJ^{+}19, GSN^{+}18, HPP^{+}18, KZ17, KCH^{+}13, MLC^{+}11, MRT^{+}19, MED16, MFL18, MM18, MND^{+}19, MCWP16, OFMZ18, SDWS13, SVN^{+}10a, URC19, VB18, WLHH18, YS16, wZcZN^{+}19, wZcZN^{+}20, ZL18]. researchers [HSB^{+}18, SDWS13]. researches [PSI19]. reservation [CJK^{+}18, DVB14]. reservations [CKP^{+}19]. reserved [DEG^{+}17, VVB15, WCHL10]. reshaping [DMZ12]. residential [GAA19, LIH^{+}19]. residual [KSS19, RT15]. resilience [DLS14]. Resilient [AAAQJ^{+}18, APR^{+}19, CM17, DLZ16, DCF19, EBCP18, LBB^{+}19, PGTBC18, RAdARP19, Wan18a, YAX^{+}18, dSFD^{+}19]. resistance [XYML19]. Resisting [WSU^{+}10]. resolution [ACSdRR17, DCC^{+}14, GMMM18, sGbKS19, HNQ^{+}18, JRJ^{+}11, JP18, PDW^{+}11, RMA^{+}16, Wei11, YDK11, YTQ19, YTQ20]. resolving [SK18]. Resonance [RDSA18, HZX^{+}19, HZX^{+}20]. Resource [COC10, DPK^{+}19, DXL^{+}18, EZTL19, KK11, LWD^{+}14, PFPJ18, PPL^{+}15, PPA18, SMRM13, SAK^{+}10, TCBPR16, TSB18, VPT^{+}15, WCW18, WY17, AJR^{+}19, AGKZ18, AMHJ10, AC18, AEME^{+}18, AS14, BAA^{+}19, BS11, BCF^{+}10, BAB12, BMK^{+}14b, BCR^{+}12, BR10, CVKB12, CVT19, CFM19, CH10, CLC11, CLL18b, CXC^{+}18, CS12, Ciu10b, CCD^{+}10, DVD12, DSCJ18, EG18, Erd13, EMJ^{+}13, EA17, FEÁ19, Fer13, FEPC18, FSM^{+}18b, GEG14, GDJ^{+}13, GAJP18, GBRM18, GBF^{+}12, GJF^{+}12, HAP15, HZZ^{+}14, HCJ14, HB19, HH19, HDLW13, IDM^{+}16, IDCJ11, IKLL12, ISS^{+}15, IAL10, JLCC12, JMAG19, KS18a, KC14, KHG^{+}18, KMI11, KADJ14, KACN16, KV12, KKW^{+}14, KGdL11, KK10b, KMK^{+}14, Lea13, Lea15, LCBF13, LTN10, LPK17, LPK18, Li10, LYYY18, LLZ^{+}18a, LSCL19, LLZ^{+}19, LZCX19, LJY10, LC15, LLS^{+}14, LSL^{+}18, LWZ18, MBS13, NRR^{+}15, NCS12, NK15, PFRC16]. resource [PdASM18, Pip10, PSW^{+}19, PPB16, PKI^{+}18, PRN14, QMSG12, QPTGG^{+}12, RPH19, RGAT18, RC13, RML^{+}19, RLP12, RT15, RP18, dRRRR^{+}18, RGC^{+}10, RSJ^{+}14, RM11, SHBP10, SPD^{+}19, SBAD^{+}18, SIL^{+}13, SCMS12, SMS14b, SSSJ19b, SCCS11, TLC^{+}15, TKR^{+}15, TTB^{+}13, TVB18, TSBH11, URKM19, VAdlP12, VHML11, VVB13b, WCHL10, WCF^{+}15, WLA18a, XWZ^{+}19, XTT18, YC13, YW12, YPHZ14, YMD^{+}13, YMY^{+}17, ZCM19, ZAB15, ZL13, ZZH^{+}16, ZCS^{+}16, ZGL^{+}18, ZXZL18, ZZZ17, ZT19, ZB19, ZA14, dSFD^{+}19, dACNC16]. Resource-aware [DPK^{+}19, PFPJ18, VPT^{+}15, NK15, PSW^{+}19, SSSJ19b]. resource-conscious [ZA14]. resource-constrained [NK15]. resource-impelled [PRN14]. resource-level [GJF^{+}12]. resource-limited [JLCC12]. resources [AMMC18, ACCM19, ABP16, ABN17, BC15, BT17, BNJ16, BKKM11, CMZ^{+}12, CXDM18, DVB14, EG18, EA13, HMH17, KHG13, KVR15, KRD^{+}19, KTTK17, Kyr19, LN13, MDA^{+}19, MBMTJR18, MDD15, RMCMD12, SGJ18, SSKK13, SDF^{+}19, TCBC18, VCKB12, VPT^{+}10, Wu16, YDQC19, ZCW11, dCTVC18]. respect [kHsZwJW18]. responders [RCMT18]. Response [CAB^{+}18, GGDM^{+}18, GGH^{+}19, KZA11, KGLY18, TDC^{+}14, WCWC19, WCWC20, ZFC17, ZFC18]. REST [WZ16]. RESTful [KH19, KH18a]. restoration [XWX^{+}17]. restricted [YSZW18]. restrictions [LDJL19]. result [CB10, ZCW11]. results [GSR^{+}19, Man15, RHMGC14, SZR18, TKTG19]. resuming [ALM^{+}10]. retail [AB18b]. retailers [SSHC19]. Retelab [CGL^{+}10]. Rethink [XJWW15]. retinal [JP18]. Retraction [AB21, ABMMC22, Bo20, DP21a, DP21b, GHEB^{+}23, HZX^{+}20, JLC^{+}20, LBJ^{+}24, WWP20, WCWC20, YWG^{+}20, YTQ20, ZMZ^{+}20, wZcZN^{+}20]. retrieval [AMBB18, AR15, ARP14, DQXW19, FLR^{+}16, FMV14, GWW^{+}19, HZC10, KZCW13, MFC^{+}19, SBD^{+}18, SSZ^{+}17, XAW^{+}10, XXB19, YJS18, ZZLZ18]. retrievals [MPP13]. return [TCH19, YAA^{+}19]. reusable [AEK^{+}18, LTZ15, RS17b, SK18]. reuse [CTR^{+}17, TSTD16]. reuse-by-composition [TSTD16]. Revamp [TKA18b]. reveals [ZZF^{+}19]. revenue [MG14, YDQC19]. reverse [MLW^{+}18b]. reversible [HLC16]. Review [AZH18, KS18b, KT17, QKC19, AHD^{+}19, EGVT18, HHZ19, LNM^{+}19, MND^{+}19, MCWP16, PSI19, SCAC^{+}19, WZH^{+}19, WLHH18, ZZBP19]. Reviewers [Ano10, Ano11a, Ano12a]. reviews [RHH^{+}19]. revision [LCC19]. revisited [RNR18]. Revisiting [WZY^{+}19, ZLL^{+}17b]. revocable [LDZW19]. Reward [CLL18b, RYH^{+}19, WHS^{+}18]. Reward-based [CLL18b, RYH^{+}19]. rewrite [SJ19]. rewrite-based [SJ19]. rewriting [XWX^{+}17]. RF [ASAAM^{+}19]. RF-based [ASAAM^{+}19]. RFID [ABC^{+}18, AMKC19, ABS^{+}18, BGC19b, CWSW14, GAI^{+}18, LCL^{+}19, LEW19, RBA17, SLK17, SJL^{+}17, YH18, YZW^{+}18]. RFID-WSN [ABC^{+}18]. RHKV [WHZ19]. Ridge [CAPG18, SVN^{+}10a]. riding [WSU^{+}10]. right [RPH19, SLD^{+}15]. right-sizing [SLD^{+}15]. rights [MJGW18]. rightsizing [ECA^{+}18]. rigid [WHW16]. Rigorous [Man15, RWY^{+}18]. RIMS [LRZ^{+}18]. ring [ZA13]. RingCube [YGYW16]. Risk [CBC^{+}19, DMM^{+}18, KADJ14, PRSR14, WZZ16, AB19a, ABGMC19, AB21, AdVAGF18, DLDTGMMP16, FTH16, GGDM^{+}18, HUY^{+}19, HLL^{+}11, JBC16, LCGPC19, LZL^{+}19a, NJ16, TY11, WCWC19, WCWC20, YPJ19]. Risk-aware [WZZ16, NJ16]. risk-based [DLDTGMMP16]. risks [AB19a, ABGMC19, AB21, BDA19, FG14, HKP10, SM10]. RLRD [LZL^{+}12]. RLWE [ZXJ^{+}14]. RM-BDP [PPA18]. RNN [CZ19, TSAER18]. road [Eng14, FBS18, LHJC18, LLN^{+}18]. roadmap [LCY19a]. roadside [MCSA18]. robot [DdSdN^{+}19, HMW^{+}19, Zhu18, DHC^{+}17]. robotic [AJR^{+}19, ADH^{+}16]. robotics [DHC^{+}17, VRGR16]. robots [CFPC17, FGW^{+}19]. Robust [PBC^{+}16, PBC^{+}17, PF17, SAVS19, AIA^{+}18a, AIB^{+}18, CM17, HZL18a, HUMA18, IOV^{+}18, LPK17, LPK18, LNK^{+}18, NWMG17, RS17b, WWW^{+}16, WLS^{+}18, HLV^{+}16]. Robustness [WSZH18, HYG^{+}19, SMS14b, SSW^{+}19]. ROC [WCWC19, WCWC20]. Role [SHJS^{+}10, YLG^{+}16, CGJ^{+}10, HLL^{+}11, LXZ^{+}18, SOA17, ZWJ19b, uRYS^{+}19]. role-and-risk [HLL^{+}11]. role-permission [LXZ^{+}18]. rollback [LBB^{+}19]. rolling [JYZ^{+}18]. root [CPP16, LWR^{+}19]. root-cause [LWR^{+}19]. ROS [WSH^{+}16]. ROSI [YAA^{+}19]. rough [LO19, YAO14]. round [CJS19, MG11, ZaTZ^{+}17]. round-optimal [CJS19, ZaTZ^{+}17]. round-trip [MG11]. routable [DK14]. route [TSS^{+}19, YPLZ17]. Router [JL14]. routes [LG18]. Routing [TKA18b, AGR19, AT19b, ASAA18, BRB19a, BRB^{+}19b, BMK^{+}14b, CA13, CJ14, CES^{+}19, Che13a, DSBC19, DK14, ESPN17, GBKJ18, HLL18, HZW^{+}18, HHW^{+}19, HYX^{+}19, HLL^{+}19, IGB^{+}14, JLQ^{+}17, KID^{+}16, LTC12, MFSV19, RQN^{+}19, RWZ^{+}19, SCY^{+}18, SMS14a, SHL^{+}19b, TZD^{+}19, VSKS19, WSZC18, WMA^{+}19, XXX^{+}19, YHL^{+}19, ZF16, ZYW^{+}18, ZTKX19, ZAI^{+}18, LC17]. RPC [KB18]. RPL [AGR19, BRB^{+}19b]. RRNS [CBT^{+}19]. RRSD [SYQ^{+}19]. RSA [YXA^{+}16]. RT [HNCJ13, WSH^{+}16]. RT-MOVICAB-IDS [HNCJ13]. RT-ROS [WSH^{+}16]. RTO [CJG^{+}18]. RTOS [JK17]. Rule [FTP14, LJ17a, XYLZ18, CCIP18, KMI11, LLC^{+}16, LHW^{+}18, MRH17, PMK18, SCN^{+}14, WQZ19, ZCX^{+}18, HCNT14]. rule-based [SCN^{+}14]. Rule-driven [FTP14]. rules [DSCJ18, EBOY14, GHGP19, LGP^{+}19, MCL^{+}16, XL19, ZDW^{+}18]. run [BC15, HIA^{+}18c, TF17]. run-time [HIA^{+}18c, TF17]. running [RM11, SHP^{+}16]. Runtime [GOLL19, JOPW14, SL19, AFB^{+}10, EP12, JPB17, Kyr19, LMZ^{+}14, LKJ17, MZC10, MSE19, PcFP^{+}17, SGL^{+}19, USK16, UDvdW^{+}18, YÁJG^{+}15]. runtime-aware [SGL^{+}19].

S [Che18, PBA18, XLW^{+}17, ZDL^{+}19]. S-ABC [XLW^{+}17]. S-InTime [PBA18]. SA [LHM14]. SaaS [CLAL19, CGL15, CPP16, FHYH15, FPPD14, IS18, LWTL19b, MBA19, RHH^{+}19, VK17, ZRZR19]. SaaS-based [FHYH15]. safe [OMKM^{+}19]. Safety [WDW^{+}19, BMR15, CY12, RSRA18, SPR^{+}10, SCEC18, TF17]. safety-critical [BMR15]. safety-focused [CY12]. SAGE [KID^{+}16]. SAGE-based [KID^{+}16]. SAGE2 [RMA^{+}16]. salesman [DC19]. samples [MR19]. sampling [MK19a, PPLL17, YKK13]. SAND [DCF19]. SAR [FCD^{+}14]. Satellite [RRKA19, CFVP12]. satisfaction [FHHM19, HMA18b, MLSF16]. saturation [Lea15]. save [SHP^{+}16]. saving [CJ14, JKLK17, QCD16, TDSH16, ZZZ17]. savings [LN13]. SCADA [FTA^{+}14, ZSL^{+}19b]. SCAI [HYC^{+}18]. SCAI-SVSC [HYC^{+}18]. Scalability [ASV^{+}13, MG18, LSB^{+}18, MROD10]. Scalable [AGA16, BKB18a, CMX^{+}16, FS18, GZLZ16, GZWQ13, LWTL19a, LLW^{+}19c, MWQ^{+}14, Pal13, PN13, THA^{+}17, TJWS10, TSRG17, ZTKF17, ABG18, AB19b, AFB^{+}10, BCC^{+}17, CLL^{+}14, FWB13a, FWB13b, GGN17, GS15, JSC^{+}15, KMB16, KK16, LCBF13, LN18, LIH^{+}19, ÖEE13, PMLVLS^{+}13, RMA^{+}16, SRZD15, SAGL10, SB17a, SGS^{+}18, SBL18, WJZ^{+}17, YARH18, ZAB15, ZXW19, ZLR^{+}15, dSK^{+}19, HMS15]. Scale [EGVT18, TKRA14, AB19c, BRL19, BAPS14, BCD^{+}18, CZT^{+}15, CWL^{+}18, CZY^{+}18, CRVZ15, CA13, CA15b, CZXL18, CCZ^{+}19, CGM^{+}18, CGJ^{+}10, CSP13, DEL19, DC19, DKJ19, DPL14, FQBCF15, FRB^{+}14, FDPR17, FPGK18, FWB13a, GDZ^{+}19, GLNT13, GSV^{+}10, GLVC18, GDAS18, GIM16, HB19, HST^{+}18, HLCL16, HZ10, Ima19, IPG^{+}18, IS18, JHC10, JYZ^{+}18, JTBS15, LKM14, LTN10, LSZ^{+}16, LFP^{+}17, LZLL18a, LSH^{+}11, LM12, MPR^{+}16, MMVS19, MLW^{+}18b, MKRD19, NS10, PPZ12, PLA18, PB17, PGCC^{+}10, PF17, PPLL17, RCW^{+}19, RMCMD12, SSHC19, SJL^{+}18, TDFZ18, TJWS10, TY11, THT12, WLLF16, WWH^{+}17, WRCC17, WCH^{+}18, WDD18, WCW18, WY19, WWG^{+}19b, WBJM14, WS10, XYLZ18, XWL^{+}15, YGYW16, ZG19, ZWW^{+}13, ZSFZ19, ZW10, ZYTC15, ZXL14, ZA14, dSFP^{+}17, dSK^{+}19]. scale-free [CWL^{+}18, LZLL18a]. scale-out [MKRD19, YGYW16]. Scaling [Eng14, HSV^{+}17, PPG19, SEMJ11, AB16, CTVB12, CJHH13, DWS12, EMJ^{+}13, EPB18, IPCA^{+}16, KTTK17, KAEC^{+}18, LMA^{+}19, SSI19, WPGN^{+}18, WWZC19, dACNC16]. Scaling-out [SEMJ11]. Scatter [FD12]. Scatter/Gather [FD12]. scattering [LSG18]. scenario [CGCB^{+}12, HZLH19, KNI^{+}18, MFC^{+}19, SGRT19]. scenario-based [KNI^{+}18]. scenario-driven [MFC^{+}19]. Scenarios [BNJ16, DFGR14, FAMA^{+}17, FPP^{+}18, GMLGB^{+}17, PZA18, RCMT18, WLB11, ZAA^{+}14]. scene [ASY^{+}18, SHL^{+}19a]. schedule [SK12]. scheduler [AJY15a, ASB18, CTR^{+}17, GdCP19, JLD^{+}19, KCS14, KPJ19, NNRA19, RD14, EMHE18]. schedulers [BBI13]. schedules [vdLLE19]. Scheduling [AHP16, AL18, ACK^{+}15, ABN17, BDNP13, CJHH13, Frî14, GVA^{+}16, KV12, KV17, Li18, LAH10, LC13, LSMVML13, PIP18b, RBJ^{+}13, RB18, RMHMG17, SVK19, SIL^{+}13, SK12, WHW16, ABMESM18, AOIS10, ANA16, ANE13, AHEM17, ADAAD12, ALK15, AGKZ18, ABP16, AB17, AB18c, ABN19, AMS19, AEM10, Bag16, BZMY10, BKS^{+}18, BCC^{+}17, BBB16, BT17, BBI13, CLRL17, CVT19, CXZ^{+}19, ÇBCA15, CA13, CA15b, CM17, CXL^{+}17, CLR18, CWJ^{+}18b, CBK^{+}17, CSL18, CGSJ18, CCCT14, DQLW15, DdM10, DR18, DXL^{+}18, DPL14, DCMW17, DSCJ18, DBS14, DNP14, EMM12, FK12, GBS10, GJ15, GOBL16, GD10, GS15, GPS13, GDS18, GKTK15, GCZ^{+}19, GGS13, HCMJ19, HHW11, HLL^{+}19, HBN^{+}13, HZLH19, JJH19, JLI^{+}13, JEB18, KR19, KS18a, KCK16, KSC^{+}19, KLP19, KX11, KK10b, LLC11, LLC14a, LGY^{+}16, LJGW18, LTTL19, LZZ19]. scheduling [LZYC13, LG16b, LCMX16, LPV^{+}16, LCdPMCT19, LYH^{+}19, MJM^{+}16, MHC14, MLSF16, MSE19, MRN19, Nag16, NRR^{+}15, NPP12, ÖEE13, PNGFJ13, PBV^{+}13, PK11, PKF14, PT16, PAB^{+}14, QCY^{+}19, QPTGG^{+}12, Qur19, RRB10, RC13, RMHCMG15, SS13, SC16, SDTA19, SDST18, SHLJ13, SWG^{+}16, SZK16, SOIS12, SH19, SGJ18, SV15, SG13, SAK^{+}10, SG14, SK19, SSP17, SGL^{+}19, TWdLZ19, TZBK13, TDFZ18, TLL^{+}11, TJZ^{+}15, TKR^{+}15, TPBS14, TdPF^{+}17, TV16, TCCC11, VVB11, VVB13a, VPT^{+}15, VVB13b, VMN^{+}18, WWX^{+}17, WHZL10, WKC^{+}13, WWC14, WQG15, WCC^{+}16, WMQ^{+}16, WCH^{+}18, WSC^{+}19, WWZ^{+}19, WS10, WXZL11, WCC14, Wu16, WWT^{+}16, WPJ16, WHCW19, XA10, XZW^{+}19, XY15, XDH^{+}17, YPLZ17, YYW^{+}19, YKK13, YIA17, YZ12, YSL19, ZCL^{+}14, ZYB^{+}18, ZGL^{+}18, ZXL^{+}18, ZZZC19, ZQB^{+}18, ZZS^{+}19, ZZZ17, Zhu18, ZT19]. schema [GTM19, ZCQ^{+}16, ZS10]. Scheme [AMSPL19, JLX^{+}19, LNLA19, AMN18, AAQ^{+}19, APK^{+}18, AM19a, AK18b, BDL^{+}19, BBH18, BDM^{+}19, BMK^{+}14b, CRRC18, CJS19, CBPP18, DLLZ17, DA18, DC18a, FSY^{+}19, FXG^{+}19, FLT^{+}19, FLL^{+}19, FHZW18, GAI^{+}18, HKA^{+}18, HAAWH^{+}18, HJA^{+}19, HZL^{+}19, HHW^{+}19, HMW14, HLC16, HNQ^{+}18, HZW^{+}16, JNHL18, JKAU19, JLQ^{+}17, KKB18, KHMB13, Kol18, KLMB19, KLW^{+}16, KLW^{+}17, LTJK12, LK12, LLW^{+}18a, LNK^{+}18, LWK^{+}18, LLW^{+}19b, LH13b, LNY^{+}18, LDZW19, LL16, LZY^{+}16, MCN^{+}18, MGN^{+}16, ODK^{+}17, OSANAM19, PGCC^{+}10, QGT^{+}18, ROK19, SGGCR^{+}16, SLS10, SGB^{+}18, SGBK19, SCZ^{+}14, SLH^{+}19, TX14, TLL^{+}19, Tor13, VGD^{+}19, WWDF18, Wan18a, WHS^{+}18, WDKV19, WZCH17, WZS^{+}18, WLS^{+}18, WLH^{+}19, XLL^{+}14, XTZ^{+}19, XXX^{+}19, YWJ^{+}19, YPLZ17, YCT15, YD18, YXD18, ZXJ^{+}14, ZCL^{+}18, ZZSZ18, ZZH^{+}18, ZZXL18, ZLT^{+}19, ZZY^{+}19, ZFC17, ZFC18, ZFH^{+}18, ZCDV19, ZLY^{+}19, ZaTZ^{+}17]. Schemes [YDNV16, BMZ10, CDDR17, CJXX19, CHS11, DGA18, HSP^{+}13, NZL^{+}15, OPO13, RA12, SAR18b, VCD^{+}18]. Schlouder [MGG^{+}17]. Scholarly [BH13, LZH^{+}18]. Schwarz [SMC18]. Science [AC10, BDP11b, CBN16, FEB^{+}19, GLD^{+}19a, LSAM13, AHP^{+}18, ACCM19, ALFR16, ACC^{+}19b, BOWD^{+}19, CAC^{+}10, CRSdS10, CDB^{+}19, CGJ^{+}10, DRZ^{+}19, DVJ^{+}15, DX14, GRZ^{+}19, GDP^{+}18, GRCP^{+}17, KZS^{+}19, KS19, MM18, MCWP16, NJKF18, NLM^{+}16, SAGL10, SPdSR^{+}17, SVB^{+}19, SG14, TBK^{+}10, TBdL16, WCKW10, XLL18b, Zhu14, AC10, BH13, GSGPP^{+}19, GDZ^{+}19, KA13]. sciences [CBBC^{+}17, GCBM17, GRL11, GPS^{+}17]. Scientific [AL18, AGMT17, CBBC^{+}17, LPV^{+}16, PN13, PS13, SG17, Ans11, ABN17, BP10, CTR^{+}17, CZ14, CPA14, CdSDS15, dCCDFdO15, DQC^{+}19, EHT10, FTP14, GAB^{+}14, GSR^{+}19, HMM18, HZP^{+}14, HSB^{+}18, HLCL16, JCD^{+}13, KKBK19, KPM^{+}18, KC19b, LKN^{+}13, LGY^{+}16, LLCF11, LPS19, LZYC13, MJDN15, MdOO^{+}17, MDO^{+}15, MED16, MFL18, NF13, PS19, PPLL17, PAC^{+}17, RKB18, RB18, SV16, SJV12, SCBK^{+}16, TKK^{+}14, TdPF^{+}17, VCKB12, WWT^{+}16, YMW13, YYLC10, ZSX^{+}15, ZLR^{+}15, dSCD^{+}19, dSFD^{+}19]. scientists [BBD^{+}13, PFS^{+}13]. SciSpace [KKBK19]. SCoPE [PP10]. scores [Bo19, Bo20]. scoring [FJA^{+}18, RS16, KKA18]. scrambling [KZA11]. Scratch [FFGP^{+}19]. screening [WZH^{+}19]. script [MBB10]. script-based [MBB10]. scripted [dlFVPSHL^{+}14]. SDBPR [ZSGJ19]. SDLC [AAJ17]. SDM [WLA17b]. SDMS [LLW^{+}12b]. SDMS-O [LLW^{+}12b]. SDN [SME^{+}21, ZFC18, ASA19, BLO^{+}18, CWLZ19, KdGP^{+}19, KBB^{+}16, SME^{+}19, SPT^{+}18, VSKS19, XDH^{+}17, ZWDP18, ZFC17, dSBN19, vdPGZ^{+}16]. SDN-aware [KBB^{+}16]. SDN-Based [SME^{+}21, SME^{+}19, dSBN19]. Seal [CWUS19]. Seamless [SLS10, CPDJ13, LLF^{+}18b]. Search [ANA16, CT19b, FPL^{+}19, MCT^{+}15, TL19, XRPT18, ABMESM18, AMÇ19, CZZ^{+}18, CCZ^{+}19, CWJ^{+}18a, CZL^{+}18b, DLH^{+}17, ESW^{+}17, EA17, FH13, FLR^{+}16, HZL^{+}19, KMT14, LWZ^{+}19a, NS17a, NJB19, PPZ12, PMBS14, RCMT18, RHMGC14, STA17a, SCBK^{+}16, TM19, TWW^{+}18, WXLY16, WCL^{+}17a, WH19, WW13, XLZ^{+}14, XTZ^{+}19, YD18, YXD18, YQZ^{+}19, ZL13, ZWZ19, vW19]. Search-based [MCT^{+}15, XRPT18, RCMT18]. search-optimizing [vW19]. searchable [CLC^{+}19, CXWT19, DLZ16, DLLZ17, HQZH14, LLL^{+}18, WXLY16]. searches [HQH16]. searching [LFH^{+}15, RTS^{+}16, WHMO13]. Sec [SSRQ19]. Sec-SLA [SSRQ19]. SecLAP [AMKC19]. secret [AIM^{+}19, AGBR19, LEW19, SAR18b, ZZ15]. Section [CMA11, ChK11, CC11, HKPT10, JS12, LS10, NFK10, SGM11, WSB^{+}15, XZ11, AC10, AR10, AM10, BB12, BR10, BRHH18, CRSdS10, CGD10, FA11a, FM10a, GVTdL18, HJC10, JO11, Lee12, OM10, RW13, Śle14, TBK^{+}10, TBdL16, XLZ18, ZSH12, Zhu10]. sector [HHW^{+}19, GG10]. sector-based [HHW^{+}19]. SecTrust [AGR19]. SecTrust-RPL [AGR19]. Secure [AD19, AMKC19, AEK^{+}18, BK16, CDDR17, Ciu10a, HQH16, HYS17, HYS18, LLQS14, LNLA19, LHL15, LZLL18b, LLZ^{+}18b, LHBC16, MLW^{+}18a, MCAS19, Mer13, MAA^{+}19, RHH^{+}16, RTS^{+}16, SAR18b, SYW17, SPKG18, WXLY16, WDZ19, XSMS15, YQZ^{+}19, ZMH^{+}18, ABMM18, AGR19, APK^{+}18, BCN^{+}19, CPD^{+}15, CFG^{+}19, CZZ^{+}18, CDL18, DJZ^{+}15, DEL19, ECE^{+}19, FLT^{+}19, FLL^{+}19, FS18, GCTLA^{+}19, GAI^{+}18, GZL^{+}18, GZQ^{+}19, HAAWH^{+}18, HZL^{+}19, HPP^{+}18, HMA^{+}18a, HLL^{+}11, JSC^{+}15, JCL^{+}15, KZA11, KKKM13, KPA17, KFK19, KLW^{+}17, LTJK12, LJS17, LJLW13, LCHW14, LWK^{+}18, LAL^{+}15, LZYC13, LL16, LWF^{+}17, LZY^{+}16, MLC^{+}18a, MK17, MHY^{+}18, NRV^{+}17, NLS19, ODK^{+}17, OSANAM19, PGCML^{+}19, POJ^{+}16, QGT^{+}18, RPH19, Rao17, RSK16, RPA^{+}18, RACA18, RHPV17, SRZD15, SGGCR^{+}16, SB17b, SBL18, SAVS19, TX14, TMB^{+}19, TCN^{+}14, TAS^{+}18, VCDK18, VGD^{+}19, WWW^{+}16, WMX^{+}17, WWDF18, WLGL19, Wan19, WDKV19, WZCH17]. secure [XZP^{+}19, XYML19, YY11, YNY^{+}14, ZZ15, ZWX^{+}19, ZCDV19, ZSW^{+}18b]. Secured [LC17, MVL^{+}18a]. securely [ZZC18]. SecureSense [RHPV17]. Securing [CMNK19, KP12, MKS^{+}19, PWA^{+}19, CDMR19, SKS^{+}18]. Security [AWYJ16, ACL^{+}18, AM10, BHE^{+}19, CCRV13, CHS11, DZZ^{+}15, FJKK17, HAJ^{+}19, HLT^{+}19, HHZ19, KKW^{+}14, LKCS18, LCL^{+}19, LYC18, LSHW17, SME^{+}21, WZWW18, XDWL15, YYS^{+}19, YAA^{+}19, ZZF18, AMN18, ABTF16, AAJ17, AMMC18, AM19a, AGBR19, BFS^{+}17a, BFS^{+}17b, BBvdB^{+}11, BWR12, CKR16, CBT^{+}19, CBPP18, CSC18, CDFW18, CMVA18, CPP16, DC18b, DZH18, DT16, ELAEAVAM19, EZTL19, FTH16, GMLGB^{+}17, GHD19, GKTK15, GJKP18, GSN^{+}18, HCJ14, HCNT14, HPP^{+}18, HIA^{+}18b, KR19, KS18b, KSK^{+}19, KBdLG18, KKKM17, KKKM18, KCCL18, LLW^{+}12a, LNB14, LGY^{+}16, LGL^{+}17, LLWZ18, MZL^{+}19, MGN^{+}16, MSM^{+}18b, MWL^{+}18b, NV11, NWMG17, OA17, PSS13, PGCML^{+}19, PSW^{+}19, RR18, RLM18, RM16, SMF^{+}19, SME^{+}19, SAGGB17, SYJ^{+}19a, SWY^{+}18, SSL^{+}19, SSZ13, SGS^{+}18, SDK19, SSB13, TZL^{+}18, TBR^{+}19, TVV13, TMS^{+}17, VPP^{+}19, WXYL15, WSL^{+}19, WZ18, WS10, YH18, YN18, YS16, YNY^{+}14, YZN^{+}15]. security [YKÖ17, ZCYZ18, ZG18, Zin18, ZL12]. security-aware [GHD19, HCJ14, WS10]. security-by-design [VPP^{+}19]. security-enhanced [AMN18, KKKM17, KKKM18]. Security-Oriented [YAA^{+}19, WZWW18]. segmentation [DH16, LSZ^{+}16, SBCF16, TCCW19, WZF^{+}19]. segmented [WBR19]. SEIM [TL19]. Selected [TCG14, CC11, CCRL18]. Selecting [CLC11, GBKJ18, WMC19]. selecting-and-pairing [WMC19]. Selection [DLH^{+}17, ABMMC18, ABMMC22, AKP^{+}18, AFSH^{+}18, AFSH^{+}19, AT18b, ArMS19, AWN^{+}13, CZY^{+}19, CMI^{+}19, DPK^{+}19, FTK^{+}14, FCY18, FJJ^{+}18, FA11b, GHY^{+}18, GB10, GAA19, KWB19, KGT15, KIC12, KP18, LLWW18, LD17, LDJL19, LWS^{+}12, LWX13, MCJ19, MHZK18, MYK16, MVCC10, PNZ14, RCM17, ROK19, RGC^{+}10, SRKS18, SMZ^{+}16, SLD^{+}18, SDH^{+}19, TDLC17, TZD^{+}19, TMP15, TAKV12, WCHL10, WHS^{+}17, WLA18a, WLQ10, XYLZ18, XLW^{+}17, XWL^{+}18, YPJ19]. Selective [LNY^{+}18, LM12, LHCC18, LHY^{+}19]. Self [AFB^{+}10, CRM^{+}16, DMZ12, GZL^{+}18, KK16, Lea13, LG16a, LHM14, MG10, NAD^{+}18, PYH17, SEPV19, TBB^{+}17, WXZ^{+}18b, dSGD13, AEME^{+}18, BMH10, CLNR18, CCT13, CFMC19, DDD18, EKSDN19, EKGS14, GDR^{+}14, GNGG17, KKB14, LH13b, LWH^{+}18, LW18b, LHPC^{+}19, MPR^{+}16, NJHT11, RP18, dRRRR^{+}18, dRRdQGR^{+}18, RWZ^{+}19, SYT^{+}19, SJL^{+}18, TJZ^{+}15, TCCC11, VRGR16, Vin16, YP12]. self-* [Vin16]. self-adaptation [CLNR18]. self-adapting [EKSDN19]. Self-adaptive [LHM14, SEPV19, WXZ^{+}18b, AEME^{+}18, CFMC19, DDD18, EKGS14, KKB14, LW18b, RWZ^{+}19, TJZ^{+}15]. Self-adjusting [Lea13]. self-certified [LH13b]. self-deployment [LWH^{+}18]. Self-evolving [NAD^{+}18]. Self-healing [AFB^{+}10, GZL^{+}18, dSGD13]. Self-managed [CRM^{+}16, LHPC^{+}19, NJHT11]. Self-managing [TBB^{+}17]. self-organized [RP18]. self-organizing [MPR^{+}16, dRRRR^{+}18, dRRdQGR^{+}18, SJL^{+}18, VRGR16]. self-paced [SYT^{+}19]. Self-provisioned [MG10]. self-reconfigurable [YP12]. Self-regulating [PYH17]. self-repair [GDR^{+}14]. Self-scalable [KK16]. Self-similarity [DMZ12]. self-sizing [GDR^{+}14]. self-stabilization [CCT13]. self-structured [BMH10]. self-tuning [TCCC11]. Self-Understanding [LG16a]. selfish [JLQ^{+}17, JXC^{+}19, PBV^{+}13]. selling [SSHC19]. SemantEco [PSW^{+}14]. Semantic [ARP14, BPC^{+}14, DNW^{+}19, HC17, HQ10, KARP14, MRS^{+}18a, PBC^{+}11, RVST17, Sun10, WZL18, AD18, ABTF16, AHYF19, ACMM19, AGA16, AGA18, ACC^{+}16, BDH14, CPSRG14, DLZ^{+}14, DT16, DPS16, DMMM11, FLR^{+}16, GACM17, HSB^{+}18, LA19, LXL^{+}17, LvW14, MPCAF15, NAM^{+}19, ODC19, PTT12, QGT^{+}18, SBCF16, SHBP10, SCN^{+}14, SSZ^{+}17, TOD17, UZ11, VPT^{+}10, WLLF16, XCGD10, XLZ^{+}14, XWL^{+}15, YCZJ18, ZS10, BDF^{+}16, CAS^{+}16, LKA^{+}19, WFQ^{+}10, Zhu10]. Semantic-aware [BPC^{+}14]. Semantic-based [ARP14, KARP14, PBC^{+}11, ACMM19, LA19]. semantically [AAF18, Li10, PSW^{+}14, RGVGGSSZ14]. semantically-enhanced [RGVGGSSZ14]. Semantics [SPdSR^{+}17, XB14, ZS16, DFLO17, LLG^{+}16, RTHB17, SOA17, TCCW19]. semantics-based [TCCW19]. Semi [ADBO18, AAF18, ArMS19, DBS14, HLL12, WHCZ18, YWY^{+}17]. semi-automated [AAF18]. semi-autonomous [DBS14]. Semi-online [ADBO18]. semi-random [ArMS19]. semi-supervised [HLL12, YWY^{+}17]. semi-tensor [WHCZ18]. semistructured [GZS14]. SemPI [DPS16]. senile [WWP19, WWP20]. sensed [Mat18, XAW^{+}10]. Sensing [AT18b, SST18, AMPZ16, CCD^{+}19, CCC19, DP19, DP20, DP21a, DP21b, DCC^{+}14, FCD^{+}14, GHEB^{+}18, GHEB^{+}23, GZW18, LWH^{+}18, LLGY18, LZY^{+}16, MWW^{+}15, MAD^{+}16, MAPA19, RYH^{+}19, SOM^{+}19, VNAMM19, WWCN13, WMY^{+}18, XLZ18, YMW^{+}18, ZDL^{+}19, KS11]. sensitive [CW13a, DK17, JOSD19, LCL14, QZD^{+}18, QCX18, VSKS19, WYL^{+}18, ZSQ^{+}19]. sensitivities [MHZK18]. sensitivity [LCH^{+}18, SWW^{+}13]. Sensor [AS18b, AZO^{+}19, FPPD14, RRS10, RM19, SYJ^{+}19b, SMS16, TKA18b, YWJ^{+}18, ARSMY19, ASO14, APK^{+}18, ACHP19, AIB^{+}18, BLAN^{+}16, CZY^{+}18, CJ14, DJPM18, FG18, FJJ^{+}18, FP14, GRTV10, GBKJ18, GLXF17, GZL^{+}18, HKA^{+}18, HAAWH^{+}18, HST^{+}18, HDH^{+}18, HZZ^{+}18, IASK14, KWK^{+}18, KS11, KLW^{+}16, LZY^{+}19a, LC17, LLQS14, LCZR12, LZXG12, LWX13, Lok12, LHBC16, LFY^{+}19, LZY^{+}16, PC17, PPS^{+}18, RAA^{+}19, RWZ^{+}19, SJ14, SCY^{+}18, SMS14b, SYCH18, THA^{+}17, TSD18, TYWZ18, TCCW19, WCB^{+}18, WWG19a, WLS^{+}18, XKJ^{+}18, YHL^{+}19, ZWS^{+}12, ZBL^{+}14, ZZLH18, dFPFG19, RVST17]. Sensor-based [RM19]. sensor-cloudlets [Lok12]. sensorized [TCB^{+}17]. sensors [BDE17, BSE^{+}13, CFMC19, GRX19, HUMA18, KMU19, LRBW17, Mat18, OMPSPL^{+}19, SPD^{+}19]. sensory [KLMB19]. sentence [MZP^{+}19]. Sentiment [AAYL19, MWQ^{+}19, ZWWL18, CC19, DMPS19, GGMS18, MZP^{+}19, ZZLZ18]. sentiment-based [GGMS18]. separation [LZL19b]. September [Ano19l]. seq2seq [WLZ^{+}19]. Sequence [LJ17a, KKvdB^{+}17, NGB18, VR12, WSL^{+}19]. sequence-based [NGB18]. sequences [LJW^{+}19a, ML19]. sequential [CGL15, DPL14, FZT^{+}18, KMC18, LBU^{+}10, MWYC12, TNY17, WHCZ18, YGY^{+}19]. SERAC3 [LLZ^{+}18a]. series [Che14, HHS^{+}18, MZH^{+}17, MFT^{+}17, NA19]. serious [AW19]. ServBGP [IGB^{+}14]. Server [FNCR11, LYL15, BK19, CWSW14, EDD^{+}10, FHZW18, KLW^{+}17, Len16, LXD17, LLWW18, LH13b, LWW^{+}18, MP17, MD12, OWX19, RT15, XDHL12, ZZBZ19, ZFC17, ZFC18, BSRR18]. Server-aided [LYL15]. Serverless [PMCC18, GAMC19, KSS19]. servers [BR18, CLH10, DZZ^{+}15, JJH19, JOPW14, JZWL17, KAEC^{+}18, ZWW^{+}13, vKLA^{+}19]. Service [ANE13, AJY15b, ACC^{+}19c, BZS18, Cha14a, CBS17, FMN^{+}17, HSBE19, HB19, JY15, JGB19, KK19, KT17, LJ17b, LNLA19, LZJL19, LYW^{+}16, MJRM16, PRS12, SCL18, SST18, SMG18, SS17, SZK18, SZW^{+}19, TSBH11, VDTK12, YG18, ZWJ^{+}19a, AdI14, AAB^{+}10, AAAQJ^{+}18, AFSH^{+}18, AFSH^{+}19, AWN^{+}13, AK14, AB18b, Ano12r, AM19b, BBC^{+}17, BAA^{+}19, BS17, BML18, BBT19, BCDP12, CMZ^{+}12, CPDJ13, CCT13, CMB17, CCIP18, CPE^{+}17, CMG^{+}19, Cha14b, CSL17, CSL19, CRWZ19, DVD12, DM12, EK11, EAED18, EG18, FLR^{+}16, FD12, FTP14, FWB13a, FWB13b, FA11b, GMMM18, GHY^{+}18, GVURIVBV14, GJGB19, GA13, GAJP18, GHJ^{+}19, GGA^{+}17, GLJ19, GPVN19, HIA18a, HA16, HA18, HC17, HLT^{+}18, HCX^{+}19, HLT^{+}19, IGB^{+}14, JGFB18, JTL^{+}19, JSS^{+}12, KZA^{+}18, KSF^{+}13, KTKN11, KK10a, KKB14, KPA17, KuRAk^{+}18, KSW^{+}13, KK16, KIC12, KK14]. service [KCCL18, LPMY18, Li15, LLW^{+}18a, LD17, LDJL19, LWS^{+}12, LLW^{+}12b, LX13, LDS^{+}18, LSL^{+}18, LXL^{+}19, LSMVML13, LLW^{+}18b, MLL15, MCL^{+}16, MZH^{+}17, MCJ19, ML11, MLM16, MHZK18, MAD^{+}16, MEW^{+}19, MRS^{+}18a, NK15, ÖE13, Pal13, PC18a, PRS^{+}13, PRSR14, PBA18, PTM^{+}18, PTT12, QZD^{+}18, RZ16, dRRRR^{+}18, RMVG^{+}10, RB18, RHMGC14, SBLW14, SSJ19, SLSS19, SPSP17, SKS17, SRKS18, SDC11, SMBMT^{+}18, SMZ^{+}16, SLL^{+}18, SLY^{+}19, SDH^{+}19, SAC11, TDLC17, TJWS10, TZL^{+}18, TGM^{+}19b, TAKV12, UGBM^{+}17, UDST19, VCE^{+}19, VDK12, VOS12, WLZ^{+}16, WHS^{+}17, WWH^{+}19, WPY19, WZ13, XSMS15, XWW19, XCGD10, XLW^{+}17, XRPT18, XCZ^{+}19, XWjZyF19, YCL^{+}19, YW12, YH19, YVCB10, YKK13, YCZJ18, YLA18, ZSL^{+}19a, ZGS^{+}13, ZZLH18, Zin18, vdPGZ^{+}16, CLM^{+}16, EMJ^{+}13, GVdBdL15, IDKD19, LPD^{+}13, RA12, SCH^{+}17]. Service-aware [SMG18]. service-based [BBT19, GAJP18]. Service-Oriented [BZS18, SCL18, AB18b, BAA^{+}19, CPDJ13, EAED18, EG18, GVURIVBV14, HLT^{+}18, LX13, MRS^{+}18a, NK15, ÖE13, SSJ19, SAC11, TAKV12]. Services [ABP18, AM17, CFP^{+}19, DA18, HM19, HXA^{+}17, LKA^{+}19, PZY16, YAP16, ABMC18, APBdI17, ATF11, AHP16, ARB12, AMM^{+}19b, AEM10, BKS^{+}14, BAV16, BCN^{+}19, BZ19, Bel16, CSV^{+}12, CPD^{+}15, CMNK19, Che13a, CAC^{+}15, CYW^{+}19, CDB^{+}19, CGL15, DQC^{+}19, DDMPG17, DX14, EAS^{+}18, EYY19, EBCP18, FHYH15, FTD17, FFPS10, GVB13, GVdBdL15, GAYTC18, GAI^{+}18, HZC10, IFD^{+}19, KIJ^{+}19, KADJ14, KH18a, KK16, Kyr19, LLMP13, LJS17, LKJ^{+}19, LPY^{+}18, LMA^{+}19, MCT^{+}15, MVG^{+}14, MM18, MMLO18, NV11, NZL^{+}15, ODK^{+}17, OB17, PVN^{+}12, PLW^{+}19, PZY17, PKI^{+}18, PFS^{+}13, QCD16, RZ16, RBN13, RHH^{+}19, RW18, RGVGGSSZ14, SRZD15, SPMC10, SMPC12, SBK^{+}16, SB19c, SSK^{+}19, STB^{+}19, SCN^{+}14, SCG^{+}18, TGM11, VAR14, VMSRM12, VSP^{+}14, VPT^{+}10, WRCC17, WDKV19, WG13, WYH^{+}17, WHYZ18, XB14, YMLT13, YAO14, YKÖ17, ZCW11, ZCLW18, ZLT^{+}19, ZWQ^{+}19]. services [ZaTZ^{+}17, dSGD19, dACNC16, ABS11, RVST17, CJS19]. serving [CZM^{+}18]. Session [LZYC13, KMC18, dSBN19]. Set [WH19, CSdCM^{+}17, HXWW18, LO19, NK18, OB19, YLG^{+}16]. Set-Union [WH19, HXWW18, OB19]. Sets [BRH18, ABMMC18, ABM19, ABMMC22, CPE^{+}17, ECPF17b, EA17, WTG^{+}14, YAO14]. settings [FFL^{+}19]. setup [MD12]. severity [ATX13]. SFaaS [KCCL18]. SFC [CWL^{+}19, LN18]. SFC-based [CWL^{+}19, LN18]. SGX [CDMR19]. SHA [LC17]. SHA-3 [LC17]. Shack [MAPA19]. shades [CPMG19]. shape [XJZ^{+}19]. share [DSS19]. shared [BOP^{+}14, BBC^{+}12, Ciu10b, DPS16, GMP^{+}16, KMB16, LTN10, LBU^{+}10, RCOP^{+}11, Wu16, YIA17, YDT19]. shared-nothing [KMB16]. shares [AIM^{+}19]. Sharing [QWCW19, AM19b, BJM^{+}17, BDCC19, CD16, CCS^{+}10, CLC^{+}19, ELAEAVAM19, Erd13, FSM^{+}18b, GGC17, GCTLA^{+}19, GG10, GLB^{+}18, HLYW17, HMA^{+}18a, HYF18, KS11, KLMB19, Lea13, LJLW13, LLZ^{+}19, LAL^{+}15, LHL15, LWYS18, LLL^{+}18, LEW19, LL16, LWZ^{+}19b, MM18, MMLO18, PIP18a, Rao17, RBC^{+}15, SHBP10, SAGL10, SSL^{+}19, SAR18b, SSB13, TKK^{+}14, TCN^{+}14, TCH19, VGD^{+}19, Wan19, WDZ19, WZ18, XCS^{+}18, YLN15, YZW14, YZZC19, ZGL19, ZZ15]. sheet [FZT^{+}18]. shelf [SMM^{+}14]. Sheltering [DR15]. ship [BRL19]. Ships [SME^{+}21, SME^{+}19]. SHM [SCL18]. shockable [AFO^{+}18]. shocks [MID16]. shop [ABMESM18, BZMY10]. Short [KLJS19, JLS19, LCY^{+}19b, ZHW19, ZNC^{+}18]. Short-term [KLJS19, JLS19, LCY^{+}19b]. Shortest [TTC^{+}14]. Shortest-linkage-based [TTC^{+}14]. Showstopper [PBC^{+}17, PBC^{+}16]. shrinkage [LSCL19]. shuffle [WWQ^{+}18]. shuffled [LCW^{+}18]. shutdown [WWZC19]. Sickness [SHH^{+}19, GGH^{+}19]. Side [YN18, SM10, SH19]. Side-channel [YN18]. Sidera [EDD^{+}10]. sieves [SBCF16]. sign [HLZ^{+}19, HYC^{+}18, LLWZ18, XJZ^{+}19, YLH^{+}19]. sign-on [LLWZ18]. signal [AKM18, BLL^{+}19, CLZ18, CMI^{+}19, GCCL18, Ima19, Kim18, LJW^{+}19a, PSI19, WZY^{+}19, WPS^{+}18, XKBA18, YSZW18]. signaling [HSBE19]. signals [AHD^{+}19, AIA^{+}18a, AIA^{+}18b, LQF19, NUPA19, RJN^{+}19, WZY^{+}19]. Signature [XZZ^{+}18, CJS19, LTMW19, QCX18, SCZ^{+}14, TX14, YN18, ZLL^{+}19, ZaTZ^{+}17]. signature-based [LTMW19]. signatures [AMM16, Che13b, MHdlS19, Tso19, WWW^{+}16]. Signcryption [Rao17, LK12, LHL15]. signed [YLH^{+}19]. significant [PSI19]. silicon [YLJ^{+}17]. SIMD [ZHHQ18]. similarities [CRYG18]. Similarity [TIHT14, WDW^{+}19, CCZ^{+}19, CSL19, DMZ12, GVI13, GA13, KMC18, LXL^{+}17, LLF^{+}18a, LLWW18, RAKJ18, SCBK^{+}16, XXX^{+}19, ZFH^{+}18]. similarity-aware [ZFH^{+}18]. Similarity-based [TIHT14]. simple [WMC19, dSFD^{+}19]. Simplified [KVHT10]. Simplifying [DDD18]. simulated [AJY12, Cha11, JXZ^{+}19]. Simulating [BL15, FPGK18, MSS^{+}16]. Simulation [AB16, CFPC17, Wri19, BDP11a, BGC^{+}19a, BKB18a, BGMLS17, CPGdS^{+}13, CPGBC16, CFP^{+}19, ÇBCA15, CWW^{+}13, DCC^{+}14, Eng14, FDPR17, HCB16, HX19, HKG^{+}16, IPG^{+}18, JH16, KNI^{+}18, KMB^{+}17, KVK^{+}18, KKP19, KGVW14, LCdPMCT19, MGT18, MND^{+}19, RLRC13, TKA^{+}18a, Tur18, VL19, WHBC19, ZYTC15, ZMN19]. simulation-based [JH16]. Simulations [FBS18, NS10, BSE^{+}13, BMP^{+}16, FQBCF15, GDP^{+}18, SSLF^{+}10, TKA^{+}18a, dSK^{+}19]. Simulator [HKT^{+}19, LLAH13]. Simultaneous [BTG19, Bag11, MFSV19, TZBK13]. simultaneously [CLH10, MZP^{+}19]. Sina [MNC^{+}18, WRCC17, WNR19]. Single [KN10, LY18a, SSG19, CJXX19, GD10, JBP^{+}18, LLWZ18, SCLC19, SYCH18]. Single- [SSG19]. single-class [GD10]. single-generation [CJXX19]. Single-pass [LY18a]. single-sensor [SYCH18]. single-stage [SCLC19]. sink [CCC19, QZM^{+}18, WCL^{+}17b]. Sinkhole [JL14]. sinks [FG18, PC17]. sinks-based [FG18]. site [SA14, TCN^{+}16]. siteDriverID [SGGCR^{+}16]. sites [CYJ19, GCV^{+}14, KSK^{+}11]. sitting [TMB^{+}19]. Situated [ZBF14]. situation [YYS^{+}19]. situational [AL14, RR18, ZRZR19]. situations [BDM^{+}19, BKB18b, RGSL18]. size [JLL17, YKK13]. sized [BAKB19, SCCS11]. sizing [GDR^{+}14, SLD^{+}15]. skeletal [LZT^{+}19]. skeletons [KMZJ16]. Sketching [CEP19a]. skew [LZW^{+}18]. skills [CFM19, SR19]. sky [DRZ^{+}19, DGD^{+}16]. SkyDrive [QC13]. skyline [LLM^{+}16]. SLA [AB16, BAA^{+}19, BS11, CMZ^{+}12, ENC^{+}12, GEG14, JTBS15, KKB14, MG14, MG16, MEBA12, NJH^{+}18, NHH^{+}19, PVGD^{+}19, SBK^{+}16, SSL12, SSRQ19, TMS^{+}17, ZAC^{+}18]. SLA-aware [BAA^{+}19]. SLA-based [AB16, CMZ^{+}12, KKB14]. SLA-driven [GEG14]. Slack [HLL^{+}17]. SLAs [GJF^{+}12, HHS^{+}18]. slave [HYG^{+}19]. Sleep [DDD^{+}19, AKM18, BJM^{+}17, CFMC19, HJA^{+}19, KLP19, KCH^{+}13]. slicing [DBP19]. Sliding [YL16, AQB15]. slotted [ZWJ^{+}19a]. SM [ATA19]. small [JLC^{+}20, PM14, SSHC19, TDC^{+}14, YWG^{+}19, YWG^{+}20]. small-world [PM14]. Smart [APRC16, AFSH^{+}19, AT18b, ACD^{+}19, BA17, BBCN18, CMNK19, CAB^{+}18, CDH^{+}19, DP17, DP19, DP20, DP21a, DP21b, DLL^{+}19, HYC^{+}18, HSS17, KPA17, LLZ^{+}18a, LNLA19, MPI^{+}18, PMDS18, PC18b, RMSPP17, RPA^{+}18, SAGGB17, SDDG17, Sta17b, SJSA19, TF17, VPA^{+}18, WLA17b, WLA18b, WSY^{+}19, XJZ^{+}19, ZYW^{+}18, AMN18, ABMM18, ABC^{+}18, AJR^{+}19, AR18, AK19, AFSH^{+}18, AT19a, ATA19, AAQ^{+}19, AHYF19, AC18, BAJ^{+}19, BBC^{+}17, Bae14, BLO^{+}18, BCN^{+}19, BWG19, BW19, CBC^{+}19, CHS11, CLH^{+}18, CFH^{+}19, CYW^{+}19, CBPP18, CS19, CCC19, CZL^{+}18b, DFRW17, DDMPG17, DGR^{+}19, EAED18, FG18, FTK17, FSM^{+}18b, GHD19, GAI^{+}18, Ham19, HC17, HHH^{+}19, HCZW17, HMA^{+}18a, HMA18b, HX19, JBC16, KGS^{+}19, KLJS19, KAS^{+}18, KFK19, KS18d, LRJG19, LNK^{+}18, LKFB18, LML^{+}19, LNM^{+}19, LSV^{+}18, LCY19a, LCY^{+}19b, LLW^{+}18b, MCN^{+}18, MLC^{+}18a, MK17, MVL^{+}18a, Mat18, NAGD18, NWMG17, NWL17]. smart [OCW14, Osm19, PC17, PRL^{+}19, PPS^{+}18, QCZH19, RGN^{+}18, RGSL18, RAdARP19, SRdlPG19, SB19c, SP18b, SAG19, SAVS19, TGM^{+}19a, TOD17, TZL^{+}18, TLL^{+}19, TWZP18, TCB^{+}17, UGBM^{+}17, WDJC18, WHBC19, WGC19, XYLZ18, XLL^{+}18a, YJS18, YSHM19, ZZY^{+}19, ZXD^{+}19, dSK^{+}19, AR18, AL14, CGSV17, FAMA^{+}17, GXL^{+}12, GMLGB^{+}17, HSS17, hKRM17, KADJ14, LSL^{+}15, TAS^{+}18, UPP17]. Smart-citie [AT18b]. Smart-Cities [Sta17b]. Smart-grid [AT18b]. smart-spaces [AT19a]. smarter [APBdI17, FRM^{+}18]. smartification [BZ19]. smartphone [AB19b, HUMA18, VOCHC17, OMD^{+}18]. smartphones [WWVJ17, YYD^{+}14]. SmartSantander [DDMPG17]. smooth [LSW^{+}19]. Smoothing [DV13]. SMP [BTM10]. SMS [CYZK15, PCK19]. SMT [PAB^{+}14]. SMT-based [PAB^{+}14]. SN [RMDB18]. snake [RDSA18]. Snort [SI18]. SNR [RJN^{+}19]. SNUSE [DEL19]. SOA [BDP11b]. SoC [DQC^{+}19, QZM^{+}18]. SoC-based [DQC^{+}19]. Social [AC16, AMPP19, BCJT13, DMPS19, DCF19, EGVT18, KZA^{+}18, LSAM13, NJ18, NJ19, PLLP19, PYM18, RQN^{+}19, SP18a, SRP19, WZWW18, ZCYZ18, ZG18, ZSGJ19, AUSA19, AKP^{+}18, AQRH^{+}18, AQAR^{+}18, AT19a, ABG18, ACM^{+}18, AMPS19, ATM^{+}19, ACD^{+}19, BCdV^{+}19, BTP19, CWJ16, CGM^{+}18, CHY^{+}18, DFLO17, DYC^{+}18, FPL^{+}19, GJ18, GBY16, GTSAR^{+}14, GWC^{+}16, GGC18, GSN^{+}18, GGMS18, HA19, HCC^{+}14, HAM18, HMC19, HYF18, JLQ^{+}17, JXC^{+}19, JWW14, JLD^{+}19, KS19, KAS^{+}18, LRL^{+}14, LCW^{+}18, LZP^{+}18, LLW^{+}19c, LLG^{+}16, LWL^{+}18, LWXY19, LLAW17, LJW^{+}19b, MML^{+}18, MYHZ18, MNC^{+}18, MVG18, MBL^{+}19, MLW^{+}18b, MSM^{+}18b, MCG^{+}15, NBB18, NZOCJ^{+}19, NJ17, NLM^{+}16, NO19, ODC19, PBL^{+}18, PPPS18, PBA18, QGX18, SOD18, SB19c, SCZ^{+}19, SA19, SB18, SKS^{+}18, SGBK19, TL19, WWVJ17, WRCC17, WJLW18, WDD18, WSN18, WMA^{+}19, WMA18, WLA18b, XCS^{+}18, XXX^{+}19, XWjZyF19, YMLT13, YZL^{+}18]. social [YXZG18b, YCXW18, YWY^{+}17, YL18, YLA18, YLH^{+}19, ZSJ19, ZLL17a, ZZH^{+}18, ZNC^{+}18, ZCL^{+}19, ZZC14, ZZLZ18, ZSS^{+}18, ZCDV19, ZWJ19b, AMQS^{+}19, GCCPGBGS10, HZW^{+}18, HAT19, IJCR19, JBM^{+}18, LKCS18, MWQ^{+}19, RC18, RC19, SYJA19, SSW^{+}19]. Social-aware [DCF19, JLQ^{+}17]. social-based [JXC^{+}19]. social-IoT [AT19a]. Socially [LAQ^{+}19]. Socially-aware [LAQ^{+}19]. society [APRC16, LLS^{+}14, PKF14, Zhu14, Zhu14]. Socio [ABD^{+}19, HMC19, KKS18b, CC19]. Socio-cyber [ABD^{+}19]. socio-economic [CC19]. socio-inspired [KKS18b]. Socio-spatial [HMC19]. SocioScope [NJ18]. Soft [EA17, TA19, BBB16, GKW^{+}12]. Soft-biometrics [TA19]. Software [CWL^{+}19, CS19, DO15, GRCP^{+}17, GTSP^{+}19, JAAD^{+}16, KB18, KKvdB^{+}17, MKRD19, NAGD18, SMG18, TDSH16, TSTD16, WRK^{+}15, ACHP19, CMZ^{+}12, CY12, CJK^{+}18, CKP^{+}19, CCDP19, CDMR19, GZLZ16, GVBG17, GHYK18, GXL^{+}18, HA16, HA18, HYG^{+}19, KHG^{+}18, Kol18, KK16, Lee12, LBM18, LGP^{+}19, LLW^{+}19c, LLS^{+}14, LRC^{+}18, PLL^{+}18, PSLZ18, QCY^{+}19, RA12, SF19, SMS14a, VSDD13, WHS^{+}17, WZWC18, YLJ^{+}17, ZGL19, ZWJ^{+}19a, ZLZ13, ZBCT17, ZZSZ18, dSK^{+}19, BRH18, EMJ^{+}13, IDKD19, RGSL18, RA12]. Software-as-a-Service [EMJ^{+}13, RA12]. Software-Defined [CWL^{+}19, SMG18, GTSP^{+}19, ACHP19, CJK^{+}18, CKP^{+}19, GHYK18, GXL^{+}18, HYG^{+}19, QCY^{+}19, ZGL19, ZWJ^{+}19a, ZBCT17]. Software-Intensive [DO15]. soils [SDST18]. solar [TTC^{+}14, GCCPGBGS10]. solid [PLZX19, MPCAF15]. solid-state [PLZX19]. solubility [HHM^{+}19]. solution [ATH^{+}19, BS17, DSS19, DMM^{+}18, GHY^{+}18, HCL^{+}17, HXY13, IDKD19, KNI^{+}18, LZX16, MGV^{+}18, PIP18a, RPMG10, SJR13, TDSH16, VRS^{+}19, WTG^{+}19, YLN15]. solutioning [MNY^{+}19]. solutions [AR17, ACL^{+}18, AAC^{+}19, BCP18, CSdCM^{+}17, CCRL18, EGVT18, KS18b, LKA^{+}19, OdI14, OMPSPL^{+}19, PAL^{+}19, PYM18, PDH18, PWA^{+}19, QCD16, dCRL^{+}19, dCTVC18]. solve [SSG17, ZZC18]. solvers [BEWZ10]. solving [CMT16, FFPS10]. Some [Pal16, KAW12, Man15, SZR18]. Sort [MAÇ17, ZZS^{+}19, GK18]. sorting [LQK^{+}16]. SOS [AMPP19]. sound [DH16, RJN^{+}19]. soundness [KYZ19]. sounds [WLZ^{+}14]. Source [QC18, TKA18b, ASAAM^{+}19, GBKJ18, HZW^{+}18, HHW^{+}19, KTTK17, MQN19, PWA^{+}19, PSK^{+}10, VOCHC17]. sources [CCJ16, ZMP10]. South [HPP^{+}18]. SP [LZW^{+}18]. SP-Partitioner [LZW^{+}18]. Space [WZL18, DBD^{+}14, GQLX18, JHC10, KI19, LGZY18, MSM^{+}18b, YPHZ14, ZZ15]. Spaces [DP17, AT19a, BR18, LLS^{+}14, SNC18, SAG19]. Spam [EAA16, CYZK15, CWZ^{+}17, SB18]. SPAMINE [AVPV17]. spammer [MK19a]. spammers [AM19b]. Spanish [SGGCR^{+}16]. spanning [CFL^{+}18]. spare [sGbKS19]. spark [GLVC18, LZW^{+}18, SJ18, SJ19, CSG^{+}18, HSV^{+}17, LWR^{+}19, TZLL18, ZHW19]. Spark-on-Yarn [ZHW19]. sparse [FFL^{+}19, LLSL18, XJY^{+}18, YJY^{+}18, ZWL^{+}16]. Sparsity [YSZW18, HDO16, MQL^{+}19, YZWG18]. sparsity-aware [MQL^{+}19]. Spatial [ATM^{+}19, CMNK19, GRL11, ATS14, CGN18, CAL^{+}18, GFD14, HMC19, LWZ^{+}19a, LLS^{+}19, LZL19b, MTD18, NWD^{+}18, NZL^{+}15, SSZ^{+}17, YXZG18b, YDT19, ZSZ14, ZLT^{+}19, dlFVPSHL^{+}14]. spatial-temporal [CAL^{+}18, NWD^{+}18, YXZG18b]. Spatio [SSP17, ZWYH19, MLW^{+}18a, QNM^{+}19]. Spatio-temporal [SSP17, ZWYH19, MLW^{+}18a, QNM^{+}19]. spatiotemporal [WZW19b]. speaker [ZDM^{+}19]. Spec [MHZK18]. Special [ADALZ14, AC10, AR10, AM10, BB13, BDF^{+}16, BB12, BR10, BRHH18, CMA11, CRSdS10, ChK11, CC11, CSYY18, CGD10, CPSRG14, DFRW17, DPDS14, DDB14, DO15, FA11a, FM10a, GMP^{+}18, HYZS16, HKPT10, HJC10, JO11, JS12, JY15, KZ14, Lee12, LS10, LNB14, NFK10, OM10, RW13, SZV19, SGM11, Śle14, TCG14, TBK^{+}10, TKRA14, TBdL16, WSB^{+}15, XZ11, XLZ18, YGS16, YJHZ14, ZSH12, Zhu10, ADLW12, ARB12, CRW^{+}16, GVTdL18, KJ12, PC18b]. Specific [JC15, KZA^{+}18, KANS18, SCZ^{+}19]. specifically [RLRC13]. specification [MCF^{+}11]. specifications [ATdC^{+}16]. Spectral [RJN^{+}19, DNW^{+}19, GHEB^{+}18, GHEB^{+}23, LJJ18, RCM17]. spectrogram [WZY^{+}19]. Spectrum [DSBC19, LSG18, JSS^{+}12]. speculative [LAL^{+}14]. speech [WZY^{+}19, YXY18]. Speed [CZM^{+}18, HGM15, SF19, ZTKX19, DJJ^{+}18, HDB18, JH16, KC19b, LSH^{+}11, LYS12, MWL18a, MAPA19, RPA^{+}18, TYWZ18, VSBN19, DSBC19]. Speed-adaptive [ZTKX19]. SpEED-IoT [DSBC19]. speedup [PJDO13]. SPFCNN [ZSQ^{+}19]. SPFCNN-Miner [ZSQ^{+}19]. spill [FCD^{+}14]. spinal [WWP19, WWP20]. spine [WWP19, WWP20]. spiral [TM19]. Split [MK19a, MYK16]. split-layer [MYK16]. SPMD [MMRL17]. SpO2 [HJA^{+}19]. sport [KMU19]. spot [CLRL18, DR18, FEÁ19, JTB13, LXJD18]. spot-checking [LXJD18]. spread [WJLW18]. spreading [DYC^{+}18, SSZ^{+}17]. spreadsheets [Asu13]. SQL [LZL^{+}16, LRMS19, SW17]. SSDs [LQK^{+}16]. SSE [Lin18]. SSL [DC17]. SSL/TLS [DC17]. SSO [MLM16, YPJ19]. SSOR [BTM10]. stabilization [CCT13]. stable [FTK^{+}14, SI19, ZF16]. stack [PDK10]. stacked [LCL^{+}18, RSY^{+}18]. Stackelberg [LN13, SSHC19, SSJ19]. stacking [LYC^{+}19]. Stag [ZSS^{+}18]. stage [QZD^{+}18, SCLC19]. stages [AAJ17, KN10, UZ11]. stakes [CND^{+}19]. stalking [EA13]. standard [RNR18]. standardized [BAB13, MEBA12, WBKL16]. standards [ACWJ19, SDK19]. star [LLpC12]. star-topology [LLpC12]. start [BKY18, KMT14, TJZ^{+}15]. State [CsZzG^{+}13, GBY16, SJV^{+}15, SGL^{+}19, WW13, XWW19, dCTVC18, ANG^{+}19, GHLW18, Kyr19, LCC19, LTC12, LSL^{+}15, Mér17, PLZX19]. State-based [WW13, Mér17]. State-of-the-art [dCTVC18]. Static [ABG17, BBB16, SDTA19, SMS14b]. station [XTL^{+}19]. stationary [CZT^{+}15, GHEB^{+}18, GHEB^{+}23, UMUB19]. statistic [GWW^{+}19]. Statistical [LCGPC19, MSA^{+}19, MR19, ZMZ^{+}19, ZMZ^{+}20, BMP^{+}16, HFM19, KHWZ18, RS16, SCH^{+}17]. statistics [WSN18]. Status [CBBC^{+}17, KARP14, LAQ^{+}19]. steel [YLL^{+}19]. Steering [LBU^{+}10, WCKW10, MDO^{+}15, RMCN^{+}10, SSC^{+}19]. steganography [MSM^{+}18b, SAM^{+}19]. stencil [GMB19, YDT19]. stent [ZMZ^{+}19, ZMZ^{+}20]. step [KRZ12, YXA^{+}18, JLQZ18]. stereoscopic [OMKM^{+}19]. still [BKY18]. stimulation [NDZ^{+}18a, NDZ^{+}18b, NDZ^{+}19]. Stochastic [FBM19, FAL^{+}19, SLD^{+}15, BKB18a, CLRL17, Li15, LWZ18, SMS14b, TLL^{+}11, ZZF^{+}19]. stock [Che14, JHC18, Lin18]. stone [HZX^{+}19, HZX^{+}20]. Storage [SRN^{+}18, AR15, AKCY^{+}17, AGA16, ADDV16, BSRR18, Bha18, BBB^{+}19, CW16, Che13b, CBT^{+}19, CDL18, DLZ^{+}14, DLLZ17, DKK^{+}13, DLH^{+}17, ED16, ED19, FH13, FLT^{+}19, GTSAR^{+}14, GTSP^{+}19, GG10, GLB^{+}18, GZQ^{+}19, HSM13, HDO16, HNKÖ18, JSC^{+}15, JCL^{+}15, LZL^{+}17, LSH^{+}11, LFH^{+}15, MPP13, MYW^{+}19, MPC^{+}18, MAA^{+}19, MM18, MKRD19, PWMX17, RP18, RTS^{+}16, RAA^{+}18, SJTN18, SYY^{+}17, STC15, SGBK19, SSKK13, SMS13, SCH^{+}19, SYQ^{+}19, TSWL17, WCF^{+}15, WZH^{+}18, WLA17a, WLA17b, WHW17, WLML17, XFLL16, XLL^{+}14, XJWW15, XWM18, XYML19, XXB19, YZW14, YNY^{+}14, YZN^{+}15, ZFW14, ZWL^{+}16, ZYCZ19, ZW10]. store [WHZ19, YFY^{+}13]. stories [VPA^{+}18]. Storing [LLCF11, LZL^{+}17, SW17]. STORK [TOS18]. storm [WSQ^{+}18, FPR18]. story [ACM^{+}18, LJ19a]. stragglers [OWX19]. strategic [ECPF17a]. Strategies [MGV^{+}18, AS19b, ASD12, BBT19, CPGdS^{+}13, HCJ14, LWTL19b, LSMVML13, RGC^{+}10, VGC^{+}13, dOOO^{+}13]. Strategy [ZZLR18, ABMESM18, CWJ^{+}18b, CS12, GJGB19, GS16b, HA19, LFL^{+}17, LSG^{+}19, NQQL13, SSL12, SYL18, TWG^{+}19, WZZ16, WCM^{+}19, WYJ^{+}19, WMC19, WS10, WW13, XWL^{+}18, XWjZyF19, YLWW18, YYLC10, YMY^{+}17, YDQC19, YK17]. STRATFram [BBT19]. Stream [CLNR18, dRADFG18, ACC^{+}16, BGC^{+}19a, BVFGWA15, EMHE18, FPR18, GRX19, HRVW18, LOR^{+}18, LSD11, SSL13, SGL^{+}19, WPGN^{+}18, WHMO13, WLW^{+}19, YL16, ZSBB19, BGC^{+}19a]. streaming [BDNP13, DHW^{+}17, kHsZwJW18, KWK16, KSK^{+}11, LZW^{+}18, MBC^{+}11, MCAS19, MYK16, RGGH18, SKF^{+}11, TCN^{+}16, WLL^{+}19a, WdL16, Wei11, YMD^{+}13, ZCK^{+}15]. Streamlining [WBKL16]. streams [CFM17, DGD^{+}16, FSV^{+}19, GACM17, LCCM18, MPP13, MTD18, PvSS17, RBLvM14, TSRG17, TCBPR16, UMUB19, WSN18, YARH18, ZSP17]. Street [AR18]. strength [AKM18]. string [LHCC18]. string-wise [LHCC18]. Stroke [TJ18]. Stroke-order-free [TJ18]. Structural [ARP14, XPL19, ZZF^{+}19]. Structure [KS18a, PC18a, WZL18, ASV^{+}13, CWL^{+}18, CD16, CLS^{+}19b, DZZ^{+}15, LY17, LY18a, PLL^{+}18, SGBK19, YK17, YNLY19, ZFY18, ZZBZ19, ZB19]. Structure-aware [PC18a]. structure-based [SGBK19]. Structured [CDG^{+}14, BMH10, CHSA18, KX11, RHMGC14, SB17a, VMSRM12]. structures [KWB19, PPG19, RGCCL18, SB18, ZZ15]. student [CMEA^{+}19, FFGP^{+}19]. students [FJA^{+}18, Kim18]. studies [Asu13, CPP^{+}18, LLES19, SHS^{+}19]. Study [SAGGB17, dLB10, APR^{+}19, CGN18, CLAL19, CCT13, CMG^{+}19, CT19c, CPMG19, FGM11, GSV^{+}10, GRS^{+}19, GMM18, HMA18b, IPG^{+}18, KHJ10, KC19b, LCY19a, MYBMM18, MNC^{+}18, MKS18, MOFGP18, PDDS10, PM14, RMJ^{+}18, SB11, SHH^{+}19, WLH16, WRCC17, XLL^{+}19b, ZDL^{+}13, ZN12, ZTC^{+}19, ZXL14, dOOO^{+}13]. Studying [FAMA^{+}17]. sub [YPCK12]. sub-directional [YPCK12]. Subevents [NO19]. subjective [LLJ^{+}11, XJY^{+}18]. subjects [Wei11]. Subscribe [BGP^{+}17, AMPZ16, EBCP18, MWQ^{+}14, OKF10, XWJ^{+}16]. subsets [QC18]. subsetting [PPZ12]. SubSift [PFS^{+}13]. subspace [ArMS19]. subsystem [CPSD18]. subtalar [Bo19, Bo20]. subtraction [DGA18]. successful [KS19]. successive [LWXY19, ZWS^{+}12]. sufferers [NDZ^{+}18a, NDZ^{+}18b, NDZ^{+}19]. sufficient [XWL^{+}18]. suite [SPMC10]. summarisation [BCR^{+}12]. Summarization [CZ14, CZ19]. SunwayMR [WHYZ17]. Super [GHO^{+}11, JP18]. Super-resolution [JP18].

SuperFine [NS17b]. superscalar [RMCN^{+}10]. superscheduling [MJRM16]. supertree [NS17b]. Supervised [hKBB11, HLL12, YWY^{+}17]. supplier [ABMMC18, ABMMC22]. supply [ABMM18, AB19a, ABGMC19, AB21, BDA19, JLQ18, LBJ^{+}18, LBJ^{+}24, PYH17, YLL^{+}19]. supply-demand [PYH17]. Support [CLAL19, CLY14, DJPM18, SCJ^{+}19a, ADBM19, BBWB^{+}18, BRB19a, BKB18b, BMP^{+}16, CAS^{+}16, CFMC19, Che13a, EP12, FK11, GGTRRC16, GHLW18, GKTK15, GSY^{+}19, HHL11, Ham17, HAA^{+}16, JNS^{+}19, KDG^{+}19, KFBKD14, LZ10, LPK17, LLW^{+}19c, LDY^{+}18, MDB^{+}18b, PWB^{+}13, SCEC18, TKTG19, VDK12, YÁJG^{+}15, ZZDM^{+}18, dlFVPSHL^{+}14, LPK18]. Supported [DPDS14, JO11, HJA^{+}19, PPS^{+}19, ZAI^{+}18]. Supporting [GJF^{+}12, JHC10, Lok12, LXZ^{+}18, LHCC18, CES^{+}19, CMP^{+}17, CDB^{+}19, CPLH19, DMPP16, GdCP19, HZL18a, SB11, SAK^{+}10, YD18, ZBF14]. surface [NUPA19]. surgery [WWP19, WWP20]. surrounding [SLTK19]. Surveillance [MPI^{+}18, AMBB18, HST^{+}18, HZZ^{+}18, LDX19, RHH^{+}16, TWZP18, UMUB19]. Survey [DRGC^{+}19, GBB18, JAAD^{+}16, AMI16, ABP18, ASD12, AMS19, BMK^{+}14a, BdDPP16, BRNR15, BBI13, FSV^{+}19, FLR13, GXL^{+}12, HQH16, HAP11, HMZ18, HDB18, IAM^{+}18, KKKM13, KAH^{+}19, KRZ^{+}19, KAW12, LY19, MDO^{+}15, RLM18, TVB18, VNAMM19, WLW^{+}19, WGM15, ZZF18, ZAI^{+}18]. surveying [YWZ^{+}18]. surveys [DGD^{+}16]. Suspending [ALM^{+}10]. Sustainability [GLD^{+}19a, AK18b, KPB18, LIC18, SA19]. Sustainable [KHG^{+}18, HYC^{+}18, KZS^{+}19]. sustaining [BFN18, SZG^{+}19]. SVD [AM19a, CHY^{+}18]. SVM [CTU19, RW18]. SVM-based [RW18]. SVMs [FZHH14].

SVR [JP18, LZL^{+}19a]. SVSC [HYC^{+}18]. SW [XZL^{+}19]. SWAN [PTM^{+}18]. swap [CSJ^{+}17]. SwapBench [ZLL^{+}17b]. swapping [ZLL^{+}17b]. Swarm [CT19b, XRPT18, GHEB^{+}18, GHEB^{+}23, JNR12, LAH10, LSV^{+}18, OB19, SJL^{+}18, WCL^{+}17b, XZW^{+}19, ZSMS18, HAT19]. SWARM-based [HAT19]. sweep [SIL^{+}13, SVN10b]. SWI [CAS^{+}16]. swift [WCF^{+}15, PLCGS11, GCM^{+}11]. switch [STC15, ZGL19, KSC^{+}19, SCJ^{+}19a, ŠCJ^{+}19b]. Switched [CGD10, CGJ^{+}10, KK10b]. switches [LKTC14]. switching [CJHH13, CWL^{+}19, HAP15, LLL^{+}19, MSM^{+}18a]. switching-frequency [CJHH13]. Sybil [AQRH^{+}18, AQAR^{+}18, JNHL18]. Sybil-precaution [AQRH^{+}18]. Symbiotic [ANA16, EA17, Sko19]. symbolic [CsZzG^{+}13]. Symmetric [LL18, DLZ16, FH13, GMdFPLC17]. symmetrical [RS17b, USK16]. symmetry [XJZ^{+}19]. symptom [YZL^{+}18]. symptom-matching [YZL^{+}18]. synchronisation [PDDS10]. synchronization [ANG^{+}19, CZT^{+}15, FSY^{+}19, HXY13, MM18, MMLO18]. synchronization-based [ANG^{+}19]. synchronizations [KYZ19]. synchronized [LKK^{+}16]. Synchronous [dRRdQGR^{+}18, GGLD10, OPT^{+}17, XZZ^{+}19]. syndrome [Bo19, Bo20]. synergetic [XFM16]. synergistic [KSS11]. synoptic [DGD^{+}16]. syntactic [GMMM18]. Synthesis [wZcZN^{+}19, wZcZN^{+}20, BGNI19, RS17b, SK18, TA19]. synthesists [SOA17]. synthesizing [LIH^{+}19]. synthetic [MR19]. Syst [AB19a, AB21, ABMMC22, BFS^{+}17a, Bo20, Cha14b, DP20, DP21a, DP21b, GHEB^{+}23, HZX^{+}20, HYS18, JLC^{+}20, NDZ^{+}18a, NDZ^{+}19, SME^{+}21, WWP20, WCWC20, YWG^{+}20, YTQ20, ZMZ^{+}20, wZcZN^{+}20]. System [AMPP19, BZS18, CBC^{+}19, CFH^{+}19, CCM^{+}14, DP20, DP21a, DP21b, GCD^{+}18, KV17, LSD^{+}17, LRZ^{+}18, LSG^{+}19, MPI^{+}18, ML17, RT16, RMA^{+}18, SVN^{+}10a, SS17, ZYA^{+}18, AD18, ABZK15, ABDH19, ABC^{+}18, ABD^{+}19, AKCY^{+}17, AQAR^{+}18, AASI17, AHYF19, AHM^{+}18, AHMS18, AMR^{+}19, AIB^{+}18, ATM^{+}19, AS18b, ACC^{+}19c, ABH18, BAJ^{+}19, BDE17, BG12, BPC^{+}14, BW19, CCT13, CM17, CWSW14, CLL^{+}14, CYZK15, CLH^{+}18, CZH^{+}18, CLS19a, CLDC19, CW13b, CSL18, CAS^{+}18, CRC^{+}19, DVJ^{+}15, DP19, DLL^{+}19, DCF19, DdSdN^{+}19, FK11, FMV14, FAL^{+}19, FAMA^{+}17, FPR18, FNA11, GLM^{+}12, GVBG17, GML^{+}13, GCCL18, GGDM^{+}18, GJKP18, GG10, HUMA18, HDA^{+}19, HIA^{+}18b, HZM14, JBC16, JAAD^{+}16, JXC^{+}19, JSC^{+}15, JXZ^{+}19, JSZ^{+}19, JOSD19, KKB18, KIAD17, KDG^{+}19, KS17a, KLJS19, KKL11, KVHT10, KLV^{+}18, LCH^{+}11, LRL^{+}14, LBJ^{+}18, LBJ^{+}24, LXD17, LZXW13, LLC^{+}16, LLN^{+}18, LSZ^{+}18, LWTL19a, LJY12]. system [LLW^{+}12b, LDY^{+}18, LLL^{+}18, LYW^{+}16, LCY^{+}19b, MVL^{+}18a, MKK13, Mat18, MTD18, MPR^{+}16, MAY18, MOFGP18, MFSV19, MQN19, NDA^{+}19, NLS19, PZC19, PdASM18, PBA18, RBGA18, RHH^{+}16, RRP^{+}14, RML^{+}19, RB12, RGM^{+}19, RGSL18, RWV^{+}13, SSG17, SAK19, SI18, SPSP17, SCL14, SZD^{+}17, SWW^{+}18, SM18, SB16, SYCH18, SAVS19, TTC^{+}14, TNY17, TMM^{+}13, TCCC11, UPP17, USK16, UDvdW^{+}18, VSP^{+}14, VPT^{+}10, WCF^{+}15, WMX^{+}17, WWZC19, WY19, WGC19, WDZ19, WLP18, XCS^{+}18, XKBA18, XKJ^{+}18, XWL^{+}15, XJY^{+}18, XYML19, XXB19, YMW^{+}18, YCY10, YWL^{+}17, YZL^{+}18, YJS18, YWZ^{+}18, YZL^{+}19, YCD^{+}19, YH19, YS16, YZ12, YZW14, ZWL13, ZFW14, ZSX^{+}15, ZGL^{+}18, ZWZ18, ZZXL18, ZRZR19, ZCQ^{+}16, ZWX^{+}19, ZW10, Zhu18, ZAI^{+}18, Zin18, ZXL14, APRC16, BBB^{+}11, GCCPGBGS10, WXYL15, ZPPE17]. System-level [SVN^{+}10a]. Systematic [KT17, MND^{+}19, MCWP16, PVHTP19, SAC11, WZH^{+}19, WLHH18]. systemic [ABF^{+}15b]. Systems [AWYJ16, AMSPL19, ADALZ14, CCRL18, DPDS14, DO15, EGVT18, LBJ^{+}24, LJC^{+}19, PIP18b, SHS^{+}19, SAPA17, TCG14, VLAC^{+}13, Wri19, YGS16, YZI18, ZZLR18, ABMM18, ASAB^{+}18, AB18a, AB19c, ATdC^{+}16, ABF^{+}15b, AR10, AB17, AB18c, AM10, ADLM18, ADBO18, AMR18, AH11, BK19, BMR15, BK16, BBWB^{+}18, BDL^{+}19, BFP18, BDNP13, BDWM17, BBMG10, BBB^{+}19, BRNR15, BR10, CCL19, CPGdS^{+}13, CMEA^{+}19, CXZ^{+}19, CSdCM^{+}17, CDL^{+}16, CTR^{+}17, CLM^{+}14a, CCJ16, CXL^{+}17, CSJ^{+}17, CLDC19, CBT^{+}19, CLP^{+}14, CSC18, CCDP19, CR14, DPK^{+}19, DKV14, DEL19, DFRW17, DV13, DLDTGMMP16, DYI^{+}16, DCMW17, EAED18, EG18, EMHE18, ECPF17b, FDPR17, FM10a, GHD19, GDJ^{+}13, GJ15, GA13, GL19, GEAR13, GOLL19, GM11, GDR^{+}14, HAF^{+}16, HNKÖ18, HRVW18, HCB16, HZP^{+}14, HPP^{+}18, HHXL13, HKS18, HDLW13, HLL^{+}17, HLT^{+}18, HCX^{+}19, HLNM11, IPG^{+}18]. systems [JLCC12, JZWL17, JLY^{+}18, KHWZ18, KANS18, KBVH14, Kha12, KB18, KKB^{+}19, Kim14, KCS14, KARP14, KB16, KFBKD14, LBD18, LLES19, LKN^{+}13, LRYJ17, LLF^{+}18b, LWK^{+}18, LCL14, LFH^{+}15, LLGY18, LSAM13, Lok12, LKJ17, LXZ^{+}18, LM12, LZY^{+}16, LHCC18, LHY^{+}19, MWQ^{+}14, MBM18, MGV^{+}18, MM10, MQL^{+}19, MFL18, MRN19, MMPF19, MKRD19, MGA^{+}19, NPH19, NKB19, NFK10, NK15, NSSA^{+}14, NQQL13, OFD17, OKF10, OPO13, OA17, OCCK14, PLL^{+}18, PdASM18, PMK18, PPS^{+}19, PKF14, PARMF14, PWMX17, PLP^{+}19, PNZ14, PQBP17, PSW^{+}19, PB18, PYH17, PSBB15, QMCX19, RBA17, RD14, RVC16b, RVC16a, SZV19, SB19a, SRdlPG19, SI18, SMS14a, SSSJ19a, SBLW14, SZK16, SOIS12, SFR15, Sko19, SMM^{+}14, SLS^{+}19, SMS13, SK19, SCH^{+}19, SGL^{+}19, SVN10b, TKR^{+}15, TAS^{+}18, THT12, TCH19, Tur18, URC19, URKM19, VF18, VDK12, Vin16, VMN^{+}18, WHZL10]. systems [WLZ^{+}16, WHS^{+}17, WZWC18, WPJ16, WZS^{+}18, WLS^{+}18, WZML18, XLZ18, YLJ^{+}17, YYS^{+}19, YP12, YL18, YMD^{+}13, ZAB15, ZMTT16, ZXW19, ZA13, ZME^{+}15, ZYB^{+}18, ZYC^{+}19, ZLG^{+}14, ZYTC15, ZLL^{+}17b, dSFP^{+}17, HYZS16, HLV^{+}16, LWW^{+}16, TKRA14].

T [Che18, QMSG12]. T-Alloc [QMSG12]. T-S [Che18]. T1000 [LYW^{+}18b]. T3 [EMHE18]. T3-Scheduler [EMHE18]. Table [JL14, GXL^{+}18, LTC12, LPY^{+}18, MWYC12]. tables [CHS^{+}18]. tablet [CPP^{+}18]. tablet-based [CPP^{+}18]. Tac [SLH^{+}19]. Tac-U [SLH^{+}19]. tacit [MED16]. tactics [LLES19]. tactile [AAS^{+}19, SLH^{+}19, VKT^{+}19]. Tag [SJL^{+}17]. tagging [WMA18]. tags [BGC19b, LEW19]. tailored [BCN^{+}19, RLRC13]. Tailoring [GNVST14]. Tale [BCG^{+}19, LLES19]. TAM [CT19c]. Taming [TGM^{+}19b]. target [ASA19, HSV^{+}17, LWX13, LXT^{+}19, WFQ^{+}10, ZBL^{+}14]. targeted [NNC^{+}19]. targeting [SBCF16, wZcZN^{+}19, wZcZN^{+}20]. Task [AMS19, AEM10, MVC^{+}13, SL11, ZZ19, ANA16, ADBO18, AMR18, BKS^{+}18, CLRL17, CXZ^{+}19, CA13, CLC11, CdSDS15, CXL^{+}17, CWJ^{+}18b, DKV14, EMM12, FGW^{+}19, FDPR17, GCZ^{+}19, IDM^{+}16, JLD^{+}19, JEB18, KOT18, KRZ^{+}19, KA19, LTTL19, LCMX16, LPL^{+}16, LWR^{+}19, MGMT18, Nag16, NPP12, PLLA18, PB18, QZM^{+}18, RS17a, SV16, SDTA19, TdPF^{+}17, TVB18, WKC^{+}13, WZM^{+}18, WSC^{+}19, WWG19a, WDR^{+}19, WPJ16, XAW^{+}10, YYW^{+}19, ZMTT16, ZCK^{+}15, ZGL^{+}18, ZZZC19, vKvWD^{+}13]. task-based [JEB18]. task-efficient [QZM^{+}18]. task-level [PLLA18, WDR^{+}19, ZCK^{+}15]. tasks [APAZ17, BDS^{+}10, CFGM16, CA15b, CLR18, GVA^{+}16, GGS13, LGY^{+}16, Li18, LLS^{+}19, LG16b, MLBS11, MRN19, Nag16, PKA19, SHP^{+}16, SC19, TLL^{+}11, WHW16, Wu16, ZXW19, ZCL^{+}14]. TASSER [ZZBZ19]. Taverna [ABG17]. Taxonomies [SV15]. Taxonomy [ALK15, DZH18, SYK^{+}17, ATS14, BMK^{+}14a, DC18b, DJ13, LCC19, LLWZ18, QKC19, TVB18, YHA^{+}19]. TBRS [LLW^{+}19b]. TCARS [RMA^{+}18]. TCKPT [KKJJ10]. TCP [KHJ10, NLLC19, WWD^{+}14]. TCP/IP [KHJ10]. team [UZ11]. teams [DBS14, GLD^{+}19a]. technical [UNM^{+}16]. technique [AKM18, DC17, GSC11, KHWZ18, KKAS19, SMRM13, SKS^{+}18, SLA^{+}16, WCC14, WTS14, YSL19, dFPFG19]. Techniques [Gra15, WRK^{+}15, ZYA^{+}18, AD18, AAN^{+}18, AMI16, ACMM19, AMS19, CGN18, CLAL19, DMPS19, DGR^{+}19, HSC15, JYY^{+}17, MVG18, SC19, SMS16, SV15, SKS17, SK12, SJSA19, TSBH11, VOCHC17, WMX^{+}17, WLB11]. Technological [dlFVPSHL^{+}14]. Technologies [BDF^{+}16, DPDS14, ZPPE17, AZH18, ABP18, CPSRG14, DFRW17, GVBG17, HSB^{+}18, HJC10, HSS17, IJCR19, KPS18, KS17b, LLMP13, MGA^{+}19, PPS^{+}19, PGCML^{+}19, RVC16b, RVC16a, Śle14, SHH^{+}19, WCKW10, HMS15]. Technology [ChK11, LCY19a, ZZLR18, AB19b, AAC^{+}19, ACPI19, Ham19, HCZW17, LBJ^{+}18, LBJ^{+}24, NLM^{+}16, SSL^{+}19, WZH^{+}18, YLL^{+}19, vdPGZ^{+}16]. TEE [DJZ^{+}15]. telecare [LWK^{+}18]. telecollaboration [AKB18b]. telecommunication [WWZZ18]. telehealth [WQG15]. telemetry [KMC18]. teleophthalmology [DRC^{+}19]. telerehabilitation [AKG^{+}17]. Telescope [CSV^{+}19, LCdPMCT19]. tell [GMM18, VPA^{+}18]. Temperature [RMRSA19, MLGGB^{+}17, ZXD^{+}19]. Temperature-aware [RMRSA19]. template [LTJK12, LZL^{+}12, MCRB19]. templates [AHM^{+}18, ACCM19, AGBR19]. Temporal [CPA14, GGMS18, AVPV17, CAL^{+}18, KZCW13, LFL^{+}17, LZL19b, MLW^{+}18a, ML19, NWD^{+}18, OCW14, QNM^{+}19, RAKJ18, RKB18, SSP17, WXPL17, WMBV17, XLZ^{+}14, XWL^{+}18, XHL^{+}19, YXZG18b, YDT19, ZZZC19, ZWYH19]. temporal-spatial [LZL19b]. temporary [LZL^{+}17]. tenacity [LSG18]. tenancy [BPC^{+}14, TCBPR16]. tenant [EMJ^{+}13, GTSP^{+}19, MDD15, PMLVLS^{+}13, RB18, SWW^{+}18, ZGB^{+}17]. tenant-based [EMJ^{+}13]. Tensor [BHE^{+}19, CSV^{+}12, DJPM18, EBOY14, SAG19, WHCZ18]. tensor-based [EBOY14]. TeraGrid [VHML10]. Terascale [WSB^{+}15]. term [DLS^{+}12, ECPF17b, HKG^{+}16, JLS19, KLJS19, LCY^{+}19b, MLL15, PBL^{+}18]. terms [KMK^{+}14]. terrain [OPO13]. terrorist [LO19]. Test [SAPA17, CsZzG^{+}13, FdAGdAFV19, GMM18, MSA^{+}19, WZCH17]. testbed [ACHP19, KSW^{+}13, MLC^{+}11, ZWDP18]. Testing [GAYTC18, MFT^{+}17, CCL19, MFC^{+}19, SZ12]. text [ASY^{+}18, BD18, FMV14, FS19, KP18, NS19, RHH^{+}19, WLLF16, XZ16, ZWWL18, ZZLZ18]. textile [WPS^{+}18]. texts [ZNC^{+}18]. textual [CFM17, JTL^{+}19]. theaters [KO11]. their [AMÇ19, JBP^{+}18, PFS^{+}13, SS13, VPA^{+}18, XYLZ18]. theoretic [DPL14, FJL^{+}16, GBKJ18, Hua10, JLQ^{+}17, JLCC12, JCL^{+}19, KK14, TLSC17, WWRS16]. Theoretical [CSL18, LY19, ZZ15]. Theory [ASW11, CDFZ16, FEPC18, YLWW18, AS18b, AY16, CPE^{+}17, Fer13, IS18, LKCS18, LZL^{+}19a, PMPC13, SSHC19, SBK18, WQZ19, XL19, YMY^{+}17, ZGZ^{+}10, ZDL^{+}19, ZCDV19, ZS10]. Theory-based [FEPC18, LKCS18, YMY^{+}17]. therapy [GZZ^{+}18, NDZ^{+}18a, NDZ^{+}18b, NDZ^{+}19, ZMZ^{+}19, ZMZ^{+}20]. there [DA18]. Thermal [SSG19, TZQ18, SSP17, ZTD^{+}18]. Thermal-aware [SSG19, SSP17, ZTD^{+}18]. Thin [JLX^{+}19, DSD^{+}11, KuRAk^{+}18]. Thin-client [JLX^{+}19]. Things [AT18a, AHYF19, ASAA18, CCRL18, DP20, DP21a, DP21b, GTEL^{+}18, GBMP13, KK19, LSD^{+}17, PPMM^{+}18, RMDB18, ZYA^{+}18, LLMP13, NWT19, ABMM18, AKP^{+}18, ACWJ19, AR18, AGR19, AMQS^{+}19, AT19b, AVPV17, AMPZ16, BA17, BAJ^{+}19, BZ19, BRH18, BGS^{+}19, BdDPP16, BRB19a, BRB^{+}19b, BWG19, CT19a, CMG^{+}19, CMI^{+}19, CBT^{+}19, CMP^{+}17, CBPP18, CDFW18, CMZ^{+}18, CDH^{+}19, DPK^{+}19, DZH18, DGR^{+}19, DP19, DC18a, ESW^{+}17, EAS^{+}18, FG18, FJJ^{+}18, FTK17, FRM^{+}18, FMRS18, FPL^{+}19, GQXL18, GLC19, GMLGB^{+}17, GBB18, GHYK18, GMD19, GCK18, GZW18, GLD^{+}19b, HDKC18, HKA^{+}18, Ham19, HZL18a, HZW^{+}18, HAT19, HHW^{+}19, HHK^{+}16, HPP^{+}18, HIA^{+}18b, HNQ^{+}18, HXL^{+}18, HSS17, IJCR19, JKAU19, JYY^{+}17, JYZ^{+}19, KWK^{+}18, KOT18, KSK^{+}19, KCM19, Kim18, KLH^{+}18, KMST19, LKCS18, LKJ^{+}19, LBD^{+}19, LHO17, LYC18, LW19, LDS^{+}18, LYYW19, LLW^{+}19d, LZJL19, LRBW17, MWQ^{+}19, MK17, MVL^{+}18a, MGL^{+}18]. Things [MLGGB^{+}17, MMC^{+}18, MDM^{+}19, MGN^{+}16, MPLM18, NSR^{+}19, NJH^{+}18, NHH^{+}19, NWT19, OFD17, PC18a, PTD^{+}18, PLGMCdF18, PC18b, QZM^{+}18, QCZH19, RGN^{+}18, RMSPP17, RJS^{+}19, RACA18, RHPV17, RGM^{+}19, RC18, RC19, SAGGB17, SYJ^{+}19a, SHS^{+}19, SST18, SWY^{+}18, SYJA19, SHL^{+}19a, SDDG17, SSW^{+}19, SCZ^{+}14, SCG^{+}18, TLSC17, TLL^{+}19, URKM19, WWX^{+}17, WZW^{+}19a, WYJ^{+}19, WZ18, WLZ^{+}19, XWW19, YLVY15, YWZ^{+}18, YCT15, YWLL19, YHA^{+}19, YN18, YAP16, ZPPE17, ZAI^{+}18, dFBP^{+}17, uRYS^{+}19]. Things-based [HIA^{+}18b, KOT18, SYJ^{+}19a]. Things-load [CT19a]. thinking [NLM^{+}16]. third [ED19, SGM11]. THOR [RAdARP19]. thoracolumbar [RBGA18]. though [KMC18]. thoughts [KAW12]. thousand [RM19]. thread [CMS^{+}18, GS15, PAB^{+}14]. threaded [MAC14, MCA^{+}18, PWWD18]. threads [LAL^{+}14]. Threat [CSYY18, AJ19, ALL^{+}18, GHP^{+}18, HDKC18, HDA^{+}19, KAW12, NAAC19, NAM^{+}19, WZH^{+}18]. threatening [AFO^{+}18]. threats [AHS^{+}18, GGDM^{+}18, RR18, RLM18, SMF^{+}19, XWRZ19]. Three [ABMMC18, ABMMC22, AMSPL19, EHT10, OSANAM19, WZWC18, LLES19, LSD^{+}17, LNK^{+}18, LWT18, ZFY18, ZZZ17]. three-dimensional [ZZZ17]. Three-Factor [AMSPL19, LNK^{+}18]. three-layer [ZFY18]. Three-level [WZWC18]. three-tier [LSD^{+}17]. Three-way [ABMMC18, ABMMC22]. threshold [HMW14, XJZ^{+}19, ZXJ^{+}14]. threshold-based [HMW14]. Throughput [TSD18, CJPC19, HAF^{+}16, PMMAM13, SCY^{+}18, TCN^{+}16, YKK13, ZSX^{+}15, ZBCT17]. throughput-oriented [CJPC19]. thru [SYW17]. Thwarting [VS13]. Tibidabo [RRP^{+}14]. ticket [XZZ^{+}18]. TIDE [SHL^{+}19b]. tier [GDR^{+}14, HGG^{+}14, IDCJ11, KIS11, LSD^{+}17, LPD^{+}13, SRN^{+}18]. tier-aware [SRN^{+}18]. Tiered [CLDC19]. tightly [BC15]. TILAA [VKT^{+}19]. TILE64 [LC14]. tiled [KWK16, PDK10, YDK11, KID^{+}16]. tiling [YDT19, vWMBS14]. Time [DFG^{+}19, GBS10, JCMPPC^{+}18, JCA^{+}19, MG18, NRV^{+}17, RMA^{+}18, SHL^{+}19b, ZWJ^{+}19a, AB19b, AQB15, AKG^{+}17, ABP16, BKB11, BMP^{+}16, CXZ^{+}19, CND^{+}19, Che14, CRRC18, CBK^{+}17, CDH^{+}19, DFLO17, DGD^{+}16, EET18, GVURIVBV14, GKW^{+}12, GMdFPLC17, GRX19, GGLW18, HKU^{+}11, HJA^{+}19, HEES19, HNCJ13, HIA^{+}18c, HHS^{+}18, KCK16, Kim18, KSC^{+}19, Li15, LRZ^{+}18, Li18, LLL^{+}19, LJ19b, LG16b, LPL^{+}16, LWW^{+}16, LHY^{+}19, MZH^{+}17, MPCAF15, MGA^{+}18, MFT^{+}17, MOFGP18, MRN19, MFSV19, MCG^{+}15, NF13, NJ17, NA19, OKF10, OWX19, OPT^{+}17, PKF14, PTD^{+}18, PGTBC18, RPA^{+}18, SHP^{+}16, SSL12, STB^{+}19, SK12, SK19, ŠCJ^{+}19b, TZST14, TJZ^{+}15, TF17, TSRG17, TQL^{+}19, TCCW19, VSBN19, WLZ^{+}14, WQG15, WTM^{+}17, WMA^{+}19, WSH^{+}16, Wu16, XLL18b, YNSM12, YCL^{+}19, ZMTT16, ZCK^{+}15, ZFC17, ZFC18, ZSP17, Zin18, WPJ16]. Time- [RMA^{+}18]. Time-aware [DFG^{+}19]. Time-based [JCMPPC^{+}18]. time-constrained [SSL12]. time-constraints [LPL^{+}16]. time-critical [KSC^{+}19, OWX19, STB^{+}19, ŠCJ^{+}19b]. time-efficient [XLL18b]. time-invariant [GMdFPLC17]. Time-Of-Flight [JCA^{+}19]. time-recordable [LWW^{+}16]. Time-relevant [SHL^{+}19b]. time-series [MFT^{+}17]. Time-slotted [ZWJ^{+}19a]. time-varying [Li15]. time/cost [KCK16]. time/cost-constrained [KCK16]. Timed [BP13, SZW^{+}19]. timely [QC18]. times [BBI13, CLRL17, JLL17, SPSP17]. Timetable [WSZC18]. Timetable-aware [WSZC18]. TimeTrustSVD [TQL^{+}19]. timing [CCDP19]. TIRIAC [NJ16]. TLS [DC17]. TMWSNs [MLW^{+}18a]. TOA [HXY13]. today [MK16a, MK16b]. token [Ciu10a, Ciu10b, OMPSPL^{+}19]. tokens [CJK^{+}18]. TOLA [SZD^{+}17]. tolerance [AMR18, GdVC10, LXJD18, NKP16, WDR^{+}19]. Tolerant [ACC^{+}19a, SE19, AFB^{+}10, CXZ^{+}19, CCL11, DZZ^{+}15, DK14, GCV^{+}14, LCBF13, LYW^{+}16, PIP18a, RWY^{+}18, RWZ^{+}19, SPR^{+}10, WHCW19, ZTKX19]. Tomogravity [LGZY18]. tomorrow [MK16a, MK16b]. tool [AB19a, ABGMC19, AB21, BGMLS17, FTH16, Ham17, MMAA19, MCG^{+}15, NS17b, SB19b, SDST18]. toolkit [AAI^{+}19, GMD19, ZYB^{+}18]. Tools [CBS17, TBK^{+}10, WRK^{+}15, BRNR15, CAC^{+}10, GLM^{+}12, KKvdB^{+}17, SP18a]. toolset [PTD^{+}18, Tao10]. Top [CCJ16, MLW^{+}18a, KMT14, MAD^{+}16, TDBR18, ZZC14, RW18]. Top- [CCJ16, MLW^{+}18a, TDBR18, ZZC14, RW18]. tophi [YTQ19, YTQ20]. Topic [AK18a, SZD^{+}17, KH19, LDS^{+}18, LZJL19, NO19, WJS^{+}18, ZZBP19, ZZJY16]. topic-based [KH19]. Topic-oriented [SZD^{+}17]. Topics [FM17, PZY16, HO17, PPB16, PZY17]. topological [Bag19, SNXB17]. topology [AAD^{+}13, LLpC12, LKTC14, PRC^{+}14, SHRE16, XWJ^{+}16, EMHE18]. topology-aware [PRC^{+}14, SHRE16]. Tor [WLYL11]. TOSCA [SCJ^{+}19a, WBKL16]. touch [Alp18, GMCM16, GMCM18, PDW^{+}11]. touchstroke [Alp18]. Tourism [CT19c]. Towns [GMLGB^{+}17]. trace [PD11]. traceability [Che18, WHS^{+}18, XLL^{+}19b]. Traceable [LDZW19, QRW^{+}18]. Traceable-then-revocable [LDZW19]. Traceback [JL14]. traces [DMMM11, PKB19]. tracing [ZLL^{+}19, ZSBB19]. track [SSC^{+}19]. tracker [MKS18]. tracking [CEP19a, FAMA^{+}17, JHC18, LWX13, SLK17, WJS^{+}18, WWTF18, YYD^{+}14, ZBL^{+}14]. tracking-based [JHC18]. Trade [DGCGH^{+}17, BDA19, DMM14, GBS10, KCM19, PMBS14]. Trade-off [DGCGH^{+}17, BDA19, GBS10, KCM19, PMBS14]. trade-offs [DMM14]. traded [VPT^{+}10]. Tradeoff [JCL^{+}19, GLD^{+}19b]. tradeoffs [KAEC^{+}18]. trading [Che14, GFW^{+}18, MVG^{+}14, RT16]. traditional [GK18, QMSG12]. Traffic [AT18a, BLO^{+}18, DK17, EMHE18, ACHP19, AC18, BMK^{+}14b, Che13a, CAL^{+}18, CS19, FTK^{+}14, FAMA^{+}17, HH19, HZL18b, JBM^{+}18, JH16, KXS^{+}16, KLJS19, KBdLG18, LLZ^{+}19, LDX19, LSL^{+}15, LWT18, LLZ^{+}18b, LGZY18, LMCSE19, LFY^{+}19, MG19, MRS18b, PGTBC18, RGSL18, SLH^{+}19, SLY^{+}19, Tur18, WLZ^{+}16, WHBC19, WSY^{+}19, XJZ^{+}19, YCL^{+}19, YSL19, ZCS^{+}16, ZYK17, ZJW^{+}14, ZMN19, ZWMC19]. traffic-aware [BMK^{+}14b, SLY^{+}19]. traffic-indexed [LSL^{+}15]. Traffic-sensitive [DK17]. Train [MK19a, MWL18a]. training [dLB10]. trajectories [CZXL18, NWD^{+}18]. Trajectory [WMBV17, DHW^{+}17, KXS^{+}16, WSN18]. transaction [GXW^{+}19, KJI11, KVvE18, LGW^{+}17]. Transactional [WZ13, GAYTC18, LAL^{+}14]. transactions [CLS^{+}19b, DR15, LAL^{+}14, WLGL19]. transbronchial [BDS^{+}10]. transceivers [ABS^{+}18]. transcription [MMF16]. Transfer [AMSPL19, HKU^{+}11, KHJ10, KB16, LCL^{+}19, LRYJ17, LSD11, LKFB18, MFL18, RSK16, RACA18, RLL^{+}17, TZST14]. transferability [CLM^{+}14a]. Transferring [KLW^{+}18]. transfers [DPBK16, LBM18, MWPVB12]. transform [BWR12, BW13, GHEB^{+}18, GHEB^{+}23, NUPA19]. transformation [GHGP19, HQZH14, LZCX19, LZL^{+}16, PKP19, SR19]. transformations [CJXX19, DGA18, LTZ15, SSFFR19]. transformed [SC19]. transformed-domain-based [SC19]. Transforming [PSK^{+}10]. Transient [DRZ^{+}19, AMR18, EA13]. transit [AHP16]. transition [GBY16, LXM^{+}18]. transitional [KRZ12]. transitions [PM14]. translation [ABF^{+}15a, CGZL19]. Transmission [KESL17, WDW^{+}19, YZZC19, BWR12, HMA^{+}18a, HYX^{+}19, LLL^{+}19, OSANAM19, PZA18, PWMX17, PLP^{+}19, TZD^{+}19, YWL^{+}17, ZZZC19]. transmissions [HSP^{+}13]. Transparent [MSI^{+}12, DW11, GTMZ17, KKJJ10, MGLPPJ13, ZWW^{+}13, vKvWD^{+}13]. transplanting [XKJ^{+}18]. transport [ZMN19]. Transportation [GCD^{+}18, LNLA19, AAA^{+}19, NWL17, PLP^{+}19]. transporting [CHJS^{+}10]. travel [BAKB19, JXZ^{+}19, dSSCdL19]. traveling [DC19]. travellers [NWD^{+}18]. treatment [Bo19, Bo20, GP11, WCWC19, WCWC20]. tree [AMM19a, BWR12, BW13, CD16, CH10, CLY14, GOBL16, HLL18, HHXL13, KMI11, LRMS19, TWW^{+}18, WWQ^{+}18, HCNT14, KK11, WTG^{+}19, CZXL18]. Tree-Rule [HCNT14]. tree-verified [HHXL13]. trees [CLR16, CY12, CFL^{+}18, EBCP18, PMK18]. tremor [AAN^{+}18]. trend [PLA18]. trending [HO17, HAM18]. trends [AVPV17, CLCMG^{+}18, GMP^{+}18, HPP^{+}18, KLH^{+}18, TKR^{+}15, VB18]. Tri [MBC^{+}11]. Tri-continental [MBC^{+}11]. trial [BRL19]. Trident [SB11]. triggered [KID^{+}16]. trip [LSV^{+}18, MG11, XCZ^{+}19]. triple [GWW^{+}19, ZZJ17]. triple-constrain [GWW^{+}19]. triple-way [ZZJ17]. TRIPOD [Pip10]. Triva [SHN10]. Trojan [CLK11]. trolleys [MOFGP18]. trouble [XZZ^{+}18]. trucks [PWP^{+}18]. TrueID [HCL^{+}17]. TruGRC [WLL^{+}19b]. TruSMS [CYZK15]. Trust [AWYJ16, ABTA18, EKSDN19, ECPF17b, KGS^{+}19, SS17, WLL^{+}19b, WCVL12, YDNV16, AGR19, ACL^{+}18, AZO^{+}19, AM10, CYZK15, FMRS18, GL19, GVdBdL15, HA19, JWW14, KZA^{+}18, LLW^{+}19b, LMM19, MML^{+}18, MG16, MPR^{+}16, NV11, NJ16, PVGD^{+}19, SYJA19, SSRQ19, SCL14, SRKS18, TDLC17, TAHS14, TY11, TQL^{+}19, WW11, ZYK17, ZZF18, ZSS^{+}18, TY11]. Trust-Aware [WLL^{+}19b, AGR19, MPR^{+}16]. trust-based [MML^{+}18]. trust-driven [NJ16]. trustable [CCCT14]. Trusted [BCP18, XZ11, CWJ16, ED19, HLNM11, JWW14, KuRAk^{+}18, NV11, WW11, ZZL^{+}10, ZZQ^{+}13]. Trusting [Lee12]. trustless [KGS^{+}19]. Trustworthiness [ABTA18, SZ12, CFGM16, LHX^{+}18, SS17, ZG18]. Trustworthy [DLMS15, YLVY15, CYZK15, FY19, ZWGC19]. truth [ZZX^{+}19]. truthful [DWJM18]. Tsallis [RDSA18]. TSH [WWG^{+}19b]. Tubular [QXZ^{+}19]. tumor [ASYF18, JLC^{+}20, YWG^{+}19, YWG^{+}20, YPJ19]. tumour [KMB^{+}17]. tunable [NUPA19]. tunable-Q [NUPA19]. tuner [vW19, vW19]. Tuning [JLRS18, CQW^{+}19, EBCP18, LCL^{+}18, RSJ^{+}14, Tao10, TCCC11, WLH16]. tuple [BR18]. Turbo [XRPT18]. turnaround [CRTN17]. TV [LCH^{+}11]. tweets [DFG^{+}19]. Twister4Azure [GZWQ13]. Twitter [AAYL19, BOHCC17, CWZ^{+}17, LCGPC19, SCZ^{+}19]. Two [DST14, ED19, GDAS18, JLI^{+}13, LW19, LIH^{+}19, TTB^{+}13, Tso19, YG18, YDNV16, AWN^{+}13, CHS11, CCD^{+}10, CHY^{+}18, EMHE18, GS13, HZW19, HZ19, MTD18, OMPSPL^{+}19, PM14, QZD^{+}18, QPTGG^{+}12, SJL^{+}17, TJ18, WWX^{+}17, WZF^{+}19, WH19, WLS^{+}18, YPCK12, YCH19, ZGV19]. two-agent [WWX^{+}17]. Two-dimensional [GDAS18, HZW19, PM14, SJL^{+}17, YPCK12]. Two-factor [ED19, LW19, WLS^{+}18]. Two-in-one [Tso19]. two-layer [TJ18]. Two-level [JLI^{+}13, CCD^{+}10, CHY^{+}18, EMHE18, MTD18, QPTGG^{+}12]. two-parallel-branch [WZF^{+}19]. two-phase [HZ19, WH19]. two-stage [QZD^{+}18]. Two-way [YG18, YCH19]. twofold [LDJL19]. TXOP [HMW14]. type [GGH^{+}19, IS18, JLS19]. type-2 [IS18]. types [SW17, YH18].

u [Kim14, SLH^{+}19]. u-healthcare [Kim14]. UASNs [KKN18]. UAV [HLL^{+}19, SGRT19, UKK^{+}19, YYW^{+}19].

UAV-enabled [UKK^{+}19]. uBaaS [LXL^{+}19]. Ubiquitous [ADALZ14, YGS16, AK19, ADLW12, CFPC17, Lok12, OdI14, Pal13, PPB16, QMCX19, VFHB14, WQ14]. UDP [CHJS^{+}10, LBM18]. UGV [ZDL^{+}19]. UHF [ABS^{+}18, BGC19b]. UHF-RFID [BGC19b]. ultra [GSC^{+}19, PDW^{+}11]. ultra-high-resolution [PDW^{+}11]. ultra-reliable [GSC^{+}19]. ultrasound [DNW^{+}19, YTQ19, YTQ20]. unbiased [Bag19]. uncanny [CPMG19]. Uncertain [BN17, AMM^{+}19b, CAL^{+}18, CLM14b, LY17]. uncertainty [CVT19, CAPG18, DM12, FZT^{+}18, HFL^{+}19, Kol18, LPBB^{+}18, FW19]. uncompressed [BD18, HKU^{+}11]. uncorrelated [MSM^{+}18b]. Under-fitting [MK19a]. understand [Eng14]. Understanding [KC19b, LG16a, LO19, SEMJ11, DM12, LMZ^{+}14, NJ18]. underutilized [KHG13]. underwater [CJG^{+}18, LSZ^{+}16, LLU^{+}18]. undo [LHCC18, LHY^{+}19]. Unfolding [DFLO17, SSLF^{+}10]. unforgeability [SCZ^{+}14]. Unicode [RM16]. UNICORE [BBvdB^{+}11]. Unicorn [XWZ^{+}19]. unification [BBH18, FLN^{+}18]. Unified [MWMA10, XWZ^{+}19, DJ13, HIA18a, HNQ^{+}18, LXL^{+}19, MRT^{+}19, RSRA18, WZF^{+}19, dOWdAS^{+}18, WG13]. uniform [CJ14, MZD^{+}16, ZCLW18]. Unifying [XFTZ16]. Unikernel [WMJW18]. Union [WH19, HXWW18, OB19]. Unit [RCW^{+}19, MCSA18, PAB^{+}14]. units [GXW^{+}19]. Universal [RACA18]. unlicensed [SLH^{+}19]. unmanned [MND^{+}19]. Unobtrusive [AKM18]. unpredictability [YZW^{+}18, ZZZC19]. unpredictability-based [YZW^{+}18]. unpredictable [ACBM15]. unresectable [ZMZ^{+}19, ZMZ^{+}20]. Unsharp [BMU18]. unstable [YZC^{+}19]. unstructured [AR15, CEP19b, JTL^{+}19, WW13]. Unsupervised [QCZH19, SCLC19, FGM11]. untraceability [YHL16]. unwanted [ZYK17]. UOV [YN18]. update [FS18, GZQ^{+}19, LTC12, LYY^{+}18, LGP^{+}19, MCRB19, PWMX17, ZCL^{+}18]. updates [Kol18, LPY^{+}18, SL19]. Updating [QCYJ17, LYL^{+}19, WCM^{+}19, ZDW^{+}18]. Uploading [CSQL17]. upon [PLLA18]. upstream [WLP10]. Urban [BAJ^{+}19, KXS^{+}16, ATM^{+}19, BA17, GACM17, LSV^{+}18, SSFFR19, SAG19, Tur18]. Urgent [BKB18b, BBWB^{+}18, KNI^{+}18, KKS^{+}18a, LPK17, LPK18]. URL [LYS^{+}19, LYC^{+}19]. usability [APRC16, GMM18]. Usage [CDL^{+}16, EH10, JOPW14, KPJ19, MM10, MWMA10, RGAT18]. USB [CLK11]. USB-based [CLK11]. Use [MCR^{+}16, Pip10, AR18, BSE^{+}13, CGCB^{+}12, CT19c, JCA^{+}19, KMZJ16, MRT^{+}19, PPG19, VHML11, VFHB14, WWA19, vdPGZ^{+}16]. used [KMU19]. USEE [MZD^{+}16]. useful [RCW^{+}19]. User [AKB18b, LJW^{+}19b, PNZ14, SZK18, WZM^{+}18, YHH^{+}19, ZWJ19b, vKvWD^{+}13, ABZK15, ACL^{+}18, AMQS^{+}19, APK^{+}18, BCN^{+}19, BDWM17, àCKPM19, CFGM16, CHS11, DEL19, DLLZ17, DGGH11, DMM^{+}18, DYC^{+}18, DGdL15, EKGS14, FK11, FTK17, FTD17, GGH^{+}19, HCL^{+}17, JRJ^{+}11, JXZ^{+}19, KZA^{+}18, KJI11, KID^{+}16, KKP19, KGLY18, KLW^{+}16, LZP^{+}18, LJY10, LH13b, LLW^{+}12b, LLGY18, LASL16, LXZ^{+}18, LWZ^{+}19b, MCRB19, MBL^{+}19, MLBS11, OAMS18, PLA18, PARMF14, PQBP17, PIP18a, PDW^{+}11, QGT^{+}18, RHH^{+}19, RAA^{+}18, RRH16, SBCF16, SSA^{+}19, SRP19, SCN^{+}14, SHH^{+}19, SSC^{+}19, WTG^{+}14, WMQ^{+}16, WSN18, WDKV19, WMA18, YAO14, YHL16, ZDL^{+}13, ZLZ13, ZXZL18, ZCZ^{+}18, dACNC16]. user-based [WMA18]. User-centric [SZK18, ACL^{+}18, DMM^{+}18, DGdL15, PIP18a, QGT^{+}18]. user-defined [KJI11]. user-generated [DGGH11]. user-level [MLBS11, ZLZ13]. user-orientated [ZDL^{+}13]. users [GCD^{+}18, IG12, JBR^{+}16, Lok12, SCZ^{+}19, SCH^{+}17, VGD^{+}19, YD18]. Using [CPP^{+}18, Che13b, CdSDS15, DBS14, Eng14, FDGR14, FMRS18, GTEL^{+}18, GRZ^{+}19, GMdFPLC17, GJKP18, HPGMM18, HGM15, KGdL11, KKL11, KKA18, LTN10, LFH^{+}15, OAMS18, PCC18, RMSPP17, RWV^{+}13, RMCMD12, SA19, STC15, VFHB14, ZBF14, dSFD^{+}19, AD18, AAN^{+}18, AsRA^{+}19, AFO^{+}18, AdI14, ABD^{+}19, ATF11, AQAR^{+}18, ASAAM^{+}19, ArMS19, AEK^{+}18, AHM^{+}18, APK^{+}18, AB19c, AMPS19, AIB^{+}18, AKM18, ASA19, AEM10, AY16, BTG19, BK19, BA17, BAJ^{+}19, BDE17, BG12, BLL^{+}19, BDP11a, BGI14, BDA19, BYL^{+}18, BBC^{+}12, BR19, BTP19, BMK^{+}14b, BDG^{+}19, BMP^{+}16, CT19b, CJPC19, CMX^{+}16, CWJD19, CLAL19, CSdCM^{+}17, CHJS^{+}10, CHSA18, CY12, CPW19, CH10, CFL^{+}15, CLY14, CHS11, CWJ16, Che18, COC10, CXC^{+}18, CTU19, CJK^{+}18, CRC^{+}19, CSQL17, DRZ^{+}19, DNJG17, DP19, DP20, DP21a, DP21b, DLL^{+}19, DC18a, DMM14, DNP14, EU19, ECPF17b, FEÁ19, FLT17]. using [FZW^{+}18, FJA^{+}18, FdAGdAFV19, FJ18, FZT^{+}18, FAMA^{+}17, FFPS10, FS19, FCD^{+}14, GQLX18, GXW^{+}19, GHGP19, GJGB19, GHP^{+}18, GHEB^{+}18, GHEB^{+}23, GBY16, GVA^{+}16, GCV^{+}14, GM11, GRS^{+}19, GMB19, GZWQ13, Ham19, HJA^{+}19, HSV^{+}17, HUMA18, HCNT14, HSB^{+}18, HMA18b, HYX^{+}19, HNQ^{+}18, HXY13, HBN^{+}13, HZLH19, HKP10, IDM^{+}16, JGFB18, JGB19, JP18, JH16, JHC18, JL14, JNR12, JOSD19, KR19, KHG13, KMB^{+}17, KMI11, KADJ14, KHO^{+}19, KBdLG18, KFK19, KVHT10, KCH^{+}13, KLV^{+}18, KS18d, KS18c, KLW^{+}16, LC17, LQK^{+}16, LY17, LY18a, LYJ10, LZXW13, LXL^{+}17, LJJ18, LLWW18, LYC^{+}19, LD17, LH13b, LLCF11, Lin18, LAH10, LSD11, LLGY18, LLZ^{+}18b, LJW^{+}19a, LMM19, LSW^{+}19, LYW^{+}16, LLU^{+}18, LKTC14, LJW^{+}19b, LTZ15, MLL15, MWL18a, MSS^{+}13, MJM^{+}16, MK17, MBM18, MVG18, MDB^{+}18a, Mat18, MSBA16, MRH17, MKM11, MDD15, MGA^{+}18, MGLPPJ13, MBA19, MSM^{+}18b, MRN19, NK18]. using [NNC^{+}19, NUPA19, NAAC19, OMKM^{+}19, OMD^{+}18, PZC19, PSI19, PSLZ18, PC18a, PPPS18, Pon19, PRW14, PDDS10, QGX18, QGT^{+}18, RBGA18, RRKA19, RJN^{+}19, RLP12, RTHB17, RSK16, RS16, RAA^{+}19, RGC^{+}10, RKB18, RS17b, SCL18, SPT^{+}18, SC19, SB19a, SF19, SAGGB17, SPdSR^{+}17, SMRM13, ST11, SLK17, SSA^{+}19, SEHS19, SMS14a, SPR^{+}10, SGM11, STB^{+}19, SB18, SKS^{+}18, SAR18b, SSL13, SCMS12, dSSCdL19, SHH^{+}19, SKS17, SJV12, SSB13, SJSA19, TBS^{+}18, TZQ18, TM19, TF17, TKA18b, TV16, TSB18, TCH19, UMUB19, VS13, VS19, VSKS19, WLZ^{+}14, WMX^{+}17, WCB^{+}18, WCM^{+}19, WNR19, WY19, WMC19, WMA^{+}19, WBKL16, WLA18a, WCC14, WTS14, WLS^{+}18, WLRL18, XLZ^{+}14, YWJ^{+}19, YWJ^{+}18, YLG^{+}16, YSL19, YCZJ18, YWF^{+}10, YK17, ZG19, ZGZ^{+}10, ZLZ13, ZWL^{+}16, ZXW^{+}18, ZGV19, ZZ19, ZZJY16, ZAC^{+}18, ZZS^{+}19, Zin18, dFPFG19, vWMBS14]. uterine [PSI19]. Utility [Pal16, PdAF12, ACHP19, BB12, FP13, GLD^{+}19b, HIA18a, MPR^{+}16, NK17, RZ16, XY15, YVCB10, YK17, YNLY19, ZLL17a, GBS10]. Utility-driven [PdAF12]. utility-privacy [GLD^{+}19b]. utilization [GXL^{+}18, LLC^{+}14b, MDB^{+}18a, RSJ^{+}14, TCH19]. Utilizing [GVI13, CZL^{+}18a, HFM19, SK19].

v1.1 [MCF^{+}11]. V2 [CCL11]. V2I [SAH19]. vaccination [BOHCC17]. Validation [GHGP19, ZYA^{+}18]. validations [DC17]. valley [CPMG19]. Value [ACHP19, BT17, HA16, WBR19, GAYTC18, HPZL18, VVB11, WHZ19]. Value-based [ACHP19, WBR19, VVB11]. valued [MZH^{+}17]. VANET [HHZ19, SBK18]. VANETs [IOV^{+}18, LWYS18, PRL^{+}19, ROK19, SAH19, TA18]. variability [ASB18, HJA^{+}19]. Variable [SCCS11, DRZ^{+}19, FZT^{+}18]. Variable-sized [SCCS11]. variant [ZXWA18]. variants [AAYL19]. variation [BBI13]. various [DGA18, JOPW14]. varying [Li15]. vault [KHMB13, RHH^{+}16]. VCPU [WCC^{+}16]. VDB [YWF^{+}10]. VDC [SLB^{+}17]. Vector [CCZ^{+}19, CLY14, HAA^{+}16, KI19, ZZ15, ZWL^{+}16]. Vehicle [SAH19, DJH^{+}19, LLL^{+}19, TLL^{+}19, XHW19, XZZ^{+}19, HLL^{+}19]. Vehicle-Assisted [HLL^{+}19]. vehicle-to-infrastructure [XHW19]. Vehicle-to-Internet [SAH19]. vehicle-to-vehicle [DJH^{+}19]. vehicles [KMJ18, MND^{+}19, HAJ^{+}19, JBM^{+}18, KLMB19, KLJS19, LWZ^{+}19a, LXF19, NLS19, WDW^{+}19, XXQ^{+}19]. Vehicular [KMJ18, KIMR15, RQN^{+}19, WSQ^{+}18, BK16, FW19, LZY^{+}19a, LLYW19, LLW^{+}19b, NLLC19, SGGCR^{+}16, VCD^{+}18, WLZ^{+}16, WSZC18, ZTKX19, LWW^{+}16]. vein [LWF^{+}17]. vending [SDDG17]. ventricular [AFO^{+}18, DNW^{+}19]. Verifiable [ABH18, CZZ^{+}18, LLL^{+}18, NJB19, WCL^{+}17a, WLXZ18, ZZ15]. Verification [KYZ19, BK19, CY12, CWUS19, CMVA18, DL19, EPB18, FLT^{+}19, LPMY18, LYZC15, LEW19, MKM11, NSR^{+}19, ZDM^{+}19, ZW10]. verified [HKA^{+}18, HHXL13]. Verifying [AH11]. versatile [SBD^{+}18, ZWX^{+}19]. version [QCYJ17, ZDW^{+}16]. versioned [ED16]. versioning [HDO16]. versus [FLN^{+}18, MGMT18]. vertebroplasty [WWP19, WWP20]. vertex [WMC19]. vertical [FJL^{+}16, KAEC^{+}18, LLF^{+}18b, MCdA16]. vertices [LXM^{+}18, WCM^{+}19]. very [CWJ^{+}18b]. VHT [MYBMM18]. via [ATA19, AkBAL^{+}19, AS18b, BBH18, BBCN18, CZL^{+}18a, CZY^{+}19, CWUS19, DH16, DLS14, DGA18, FBM19, GCBM17, GDP^{+}18, GLD^{+}19a, GJF^{+}12, GDAS18, JHC10, LZT^{+}19, LLSL18, LEW19, MZH^{+}17, MBC^{+}11, MYK16, NS19, QCYJ17, RGM^{+}19, SC16, SYJA19, SAG19, WZW^{+}19a, WJLW18, XFTZ16, XLL^{+}18a, YHH^{+}19]. viability [dACAM13]. VIALACTEA [SVB^{+}19]. Victoria [PWB^{+}13]. Video [TWZP18, DQXW19, GGH^{+}19, HKU^{+}11, HSP^{+}13, HMA^{+}18a, JLCC12, KWK16, KSK^{+}11, KSW^{+}13, KHO^{+}19, LYS12, LYXT14, MYK16, PSPP16, SKF^{+}11, TZD^{+}19, YARH18]. videos [HWWT12, SLTK19]. VIDIA [SG19]. view [AD18, BTP19, DCMB15, LYS12, LTZ15, QCZH19, ZTC^{+}19]. viewpoint [PD11]. violation [HHS^{+}18, OWX19, XWL^{+}18, ZAC^{+}18]. violations [ENC^{+}12, NJH^{+}18, NHH^{+}19]. Viper [WPGN^{+}18]. Virtual [ACC^{+}19c, BCF^{+}10, CFVP12, DK17, LYYY17, LPBB^{+}18, MG10, RCTY19, SSFFR19, SBP^{+}17, WLL^{+}19b, ZLL^{+}16, AD19, ADA^{+}19, ABF^{+}15a, ALM^{+}10, AAM^{+}16, AS14, BJ12, BOWD^{+}19, CGN18, CRM^{+}16, CRC13, CNR19, dCCDFdO15, DJZ^{+}15, DPBK16, DEG^{+}17, DQLW15, DCMW17, FBM19, GVI13, GCV^{+}14, GGH^{+}19, HMH17, HZZ^{+}14, HH19, HSC15, HHZ16, JNR12, JDW^{+}14, KSSG16, KTKN11, KCS14, LJS17, Li10, LJLW13, LYYY18, LLF^{+}18a, LLWW18, LFHQ19, LXRS19, LLZ^{+}19, LLS^{+}19, LJY10, LJL12, LSYC18, LC13, LSW^{+}19, Man15, MRT^{+}19, MK19b, MFG^{+}14, MROD10, NWT19, PFPJ18, Pon19, RMHMG17, SS13, SB14, SDWS13, SHLJ13, SMSF18, SAK^{+}10, SYAL13, SLA^{+}16, SLZ^{+}18, TMMVL12, UZ11, VVB13b, WQG15, WTM^{+}17, WY19, XJWW15, XWX^{+}17, YLHJ14, YPLZ17, YWF^{+}10, ZYZ^{+}18, ZCS^{+}16, ZWHC17, ZHHC17, ZFY18, ZZZ17, ZZL^{+}10, HLCL16, LWW^{+}13]. virtual-to-physical [ABF^{+}15a]. virtualization [CMG^{+}19, FJKK17, KKB14, LLW^{+}12a, LCL14, MG19, MSI^{+}12, SVN^{+}10a, SWW^{+}18, VOS12, WCC^{+}16, YCL^{+}19, ZLZ13, MCJ19]. Virtualized [BB17, ASB18, BdM11, BBC^{+}12, CP17, FNCR11, GKW^{+}12, GBF^{+}12, GSP^{+}17, JLI^{+}13, JK17, LC13, PPL^{+}15, TTB^{+}13]. virus [LFL^{+}17]. VisGenome [JHC10]. Visibility [BHE^{+}19]. Visible [PCK19]. Vision [AUSA19, GMLGB^{+}17, GBMP13]. Visual [DM12, ADLM18, BCD^{+}18, HHZ16, LSZ^{+}18, LDX19, LYXT14, MSM^{+}18b, PC17, PKA19, XJY^{+}18]. Visualization [YDK11, MID16, NOF18, OPO13, ODC19, SHN10, WCKW10]. visualizations [Gra15]. Visualizing [BAPS14, GRMSOG18, PDK10]. vital [HYC^{+}18, XKBA18]. vitality [LFL^{+}17]. vivo [WXZ^{+}18a]. VLBI [CHJS^{+}10, WWD^{+}14]. VLBI UDP [CHJS^{+}10]. VM [AB16, CJHH13, CFF14, FCY18, HZW^{+}16, JFZL17, KSF^{+}13, MYW^{+}19, STMV18, SSSJ19a, SE19, WLA18a, YZLQ14, YLJL18, ZZJ17]. VM-scaling [AB16]. VMI [LLF^{+}18a]. VMM [AD18]. VMP [JNR12]. VMs [KMT14, SEPV19]. VNF [CYW^{+}19]. vocabulary [LYXT14]. vocal [AHMS18]. VoD [LJY12]. VoIP [Che13a]. volatile [CCL11]. volatility [Lin18]. voltage [LG16b, WWZC19]. volumes [WHMO13]. volunteer [ALFR16, CGCB^{+}12, CMT16, CCCT14, GDJ^{+}13, GJ15, JCA^{+}19, KJ12, KMV^{+}15, MKK13, PdASM18, SG13]. Voronoi [STA17a, WZW^{+}19a]. Voronoi-based [STA17a]. VOs [VKK14]. voting [ZCZ^{+}18]. VPN [MSI^{+}12, vdPGZ^{+}16]. VR [SHH^{+}19]. vs [IPG^{+}18, NLLC19]. VSA [SHLJ13]. VSI [HM19]. Vulnerabilities [YS16, GHYK18, GGC18]. Vulnerability [LKCS18, NJKH13].

W3P [FKOC11]. wait [BKY18]. Wale [DSS19]. walk [MLW^{+}18b, ZSFZ19]. walks [LXM^{+}18]. wall [NRR^{+}15, KID^{+}16]. wallet [WLGL19]. wallets [VSM^{+}19]. Wallis [MSA^{+}19]. walls [YDK11]. WAN [SISGS18, ZFY18]. wandering [Ciu10b]. want [VSM^{+}19]. wardens [MWCK19]. Warning [ZPPE17, VMN^{+}18]. warp [ZXL^{+}18]. WASAM [AL14]. wasn’t [DA18]. watchdog [ABF^{+}15b, JXC^{+}19]. water [BNJ16]. watermark [PvSS17]. watermarking [AIA^{+}18a, AIA^{+}18b, AIM^{+}19, AM19a, BW13, HKA^{+}18, HPL^{+}19, RS17b, SKS^{+}18, ZSMS18]. wavefront [MAPA19]. Wavelet [KKB^{+}19, BWR12, BW13, GHEB^{+}18, GHEB^{+}23, NUPA19, WLZ^{+}14]. wavelet-time [WLZ^{+}14]. wavelets [PRW14]. way [ABMMC18, ABMMC22, dRADFG18, GFW^{+}18, LAQ^{+}19, WDR^{+}19, YG18, YCH19, ZZJ17, SVB^{+}19]. WBAN [GCK18]. weakening [CWL^{+}18]. weaker [Wan18b]. weakness [JLS19]. wear [WZS^{+}18]. wear-aware [WZS^{+}18]. wearable [CRRC18, CRC^{+}19, DP19, DP20, DP21a, DP21b, GRX19, KMU19, LNY^{+}18, PSW^{+}19]. weather [LCL^{+}18, ZXL14]. Web [CAS^{+}16, LKA^{+}19, AAF18, AS19b, ACPI19, AAI^{+}19, CFM19, CSL19, FTH16, HC17, KH18a, LPMY18, LXL^{+}17, LYS^{+}19, MCT^{+}15, NWT19, NK18, RW18, SB19b, ZJW^{+}14, ATF11, BRXdS11, BDF^{+}16, CMZ^{+}12, CAC^{+}10, CSL17, CGL^{+}10, DMMM11, FD12, FKOC11, FCD^{+}14, GP11, HZC10, KS11, LLMP13, LASL16, ML11, MGLPPJ13, PFS^{+}13, RHMGC14, TGM11, VSP^{+}14, WCVL12, XLZ^{+}14, XRPT18, YMLT13, YAO14, ZCW11, ZWW^{+}13]. web-based [FTH16, DMMM11]. Web-centred [LASL16]. web-resource [CFM19]. webpage [LYC^{+}19]. websites [SZ12]. weekly [MHW^{+}16]. Weibo [MNC^{+}18, WRCC17, WNR19]. weight [AKB^{+}18a, Eng14, LTC12]. weighted [BRL19, LLSL18, LJW^{+}19b, PLL^{+}18, SSRQ19, TJ18, YPCK12, YZ12, YL16, ZCW19, RCM17]. weighted-fair-queuing [YZ12]. WFQ [YZ12]. whale [ABMESM18]. wheelchair [SLTK19]. Where [MGR11]. while [BKY18, LLW^{+}12b]. white [ZSW^{+}18a]. white-box [ZSW^{+}18a]. whitelists [HLNM11]. Whole [BCG^{+}19, CMX^{+}16]. whole-exome [CMX^{+}16]. Wi [AKM18, DCBF19, KWB19, SLS10]. Wi-Fi [AKM18, DCBF19, KWB19, SLS10]. wide [AL14, BBI13, GG10, LCBF13, ZGV19]. wide-area [AL14, LCBF13]. WiFi [AMRM18, NS10]. WiFi-based [NS10]. Wiki [LG16a]. Wiki-Health [LG16a]. Wikipedia [BAPS14]. wildfire [JNHL18]. WiMAX [SLS10]. WiMAX/heterogeneous [SLS10]. Wind [SMC18, CT19b, YGY^{+}19]. window [YL16]. windows [AQB15, RJN^{+}19]. wings [GTCZG^{+}18]. wire [BKY18]. Wireless [AS18b, AZO^{+}19, CWL^{+}19, PC17, SYJ^{+}19b, SMS16, TKA18b, WWTF18, ARSMY19, APK^{+}18, AIB^{+}18, BLAN^{+}16, CZY^{+}18, CJ14, CZM^{+}18, FG18, FJJ^{+}18, FAL^{+}19, GBKJ18, GLXF17, GZL^{+}18, HKA^{+}18, HMW14, HDH^{+}18, IASK14, KMU19, KLW^{+}16, LC17, LLQS14, LW18a, LCZR12, LZXG12, LHBC16, LFY^{+}19, LZY^{+}16, MYK16, NS10, PNZ14, QGX18, RYH^{+}19, RWZ^{+}19, SJ14, SCY^{+}18, SCS^{+}18, SHL^{+}19a, TSD18, WDJC18, WCW18, WLS^{+}18, WGX^{+}19, XZ14b, YHL^{+}19, ZBL^{+}14, ZZLH18, ZWJ^{+}18, AS18a]. wise [LHCC18, XTL^{+}19, SEPV19]. withholding [HS19]. within [AJY15b, DMPP16, EP12, EBCP18, LLS^{+}19, MZP^{+}19, MT17, WDR^{+}19, ZDL^{+}13]. without [DH16, DSM^{+}19, LY17, SCLC19]. Witness [CXWT19]. Witness-based [CXWT19]. WLAN [MYBMM18]. WOA [HEES19]. WoBinGO [ISS^{+}15]. wolf [CT19a]. words [RM19, SAM^{+}19]. Work [JO11, TF17, FP13, PPS^{+}19]. workbench [SB11, ŠCJ^{+}19b]. worker [LC14]. Workflow [AL18, AM17, GKW^{+}12, HMS15, MJM^{+}16, MSE19, MDM^{+}19, SV16, ANE13, AJR^{+}19, AKCY^{+}17, AB18a, ALK15, ABP16, ABN19, CHC^{+}17, CdSDS15, CLR18, CBK^{+}17, CGSJ18, DVJ^{+}15, DMMM11, DNP14, EKSDN19, EHT10, FK11, GHY^{+}18, GGC17, GJ15, GRCP^{+}17, HHD^{+}12, HZP^{+}14, HHW11, HLT^{+}19, JPB17, KVHT10, LGY^{+}16, LLCF11, LSH^{+}11, MSS^{+}13, MJDN15, NF13, OdOD^{+}13, PAC^{+}17, Qur19, RRB10, RB12, RLRC13, RB18, SLC^{+}17, SB11, SGJ18, Sip12, SIL^{+}13, SV15, SDC11, SCBK^{+}16, SK19, SLL^{+}18, TMM^{+}13, TKK^{+}14, WYBS11, WWT^{+}16, WSY^{+}19, XZW^{+}19, XWL^{+}18, XCZ^{+}19, ZCW11, ZME^{+}15, ZLR^{+}15, ZZS^{+}19, ZT19, dSGD13, dSFP^{+}17]. Workflow-and-Platform [SV16]. Workflow-based [MDM^{+}19, LSH^{+}11, MSS^{+}13]. workflow-oriented [SLL^{+}18]. Workflows [LPV^{+}16, MG11, AHP16, AHP^{+}18, ABG17, AB17, ABN17, AB18c, AGMT17, Bal16, BBB16, BKKM11, CLRL17, CMX^{+}16, CTR^{+}17, CBBC^{+}17, dCCDFdO15, DGR^{+}15, DM12, GAB^{+}14, GSR^{+}19, HMM18, HCJ14, JTS13, JTBS15, JCD^{+}13, KS18a, KCK16, KS17a, KOP^{+}17, KTTK17, MT17, MdOO^{+}17, MDO^{+}15, Nag16, PS19, PFS^{+}13, QWCW19, RKB18, SG17, SSC^{+}19, TWdLZ19, TdPF^{+}17, TGM11, WZZ16, WHW16, WWZ^{+}19, WDR^{+}19, YYLC10, ZYB^{+}18, dSCD^{+}19, dSFD^{+}19, dOOO^{+}13, ACBM15]. Workforce [MMF16]. Workforce-efficient [MMF16]. Workload [KMB16, KS18c, WCF^{+}15, XAW^{+}10, AOIS10, AMKM18, ADBO18, CWW^{+}16, DMZ12, KVR15, LJGW18, LSCL19, PZA18, PdAF12, RSJ^{+}14, SS13, SRN^{+}18, WMLS14, ZWW^{+}13]. Workload-aware [KMB16]. Workloads [BB17, BBI13, CJPC19, DGdL15, HRVW18, LFHQ19, MMVP13, RB18, VVB13a, VVB13b, WMQ^{+}16]. workpiece [OPT^{+}17]. works [PFS^{+}13]. workshop [NFK10]. workspace [CSV^{+}19, KKBK19, WHYZ18]. workspaces [GMP^{+}16, PDW^{+}11, RCOP^{+}11]. world [ACBM15, CPLH19, LYC18, LNM^{+}19, PWP^{+}18, PM14, YAO14]. worms [AMM16, CsZW14, FX10]. worth [RM19]. WoT [AMM^{+}19b]. WPAN [vdLLE19]. WPANs [GHMX10]. WPD [WLZ^{+}14]. wrapper [HAA^{+}16, HIA^{+}18c]. write [LFHQ19, LXRS19, MPC^{+}18, SCH^{+}19, TMDZ15]. write-enabled [MPC^{+}18]. write-intensive [LFHQ19]. WSADF [MSE19]. WSN [ABC^{+}18, BARMB14, MOFGP18, RRU^{+}18, SCL18, SMS14a]. WSN-based [SCL18]. WSNs [ABB19b, FFL^{+}19, HZW^{+}18, HHW^{+}19, WCL^{+}17b]. WSNs-assisted [FFL^{+}19].

X [DRZ^{+}19, SYT^{+}19]. X-ray [DRZ^{+}19, SYT^{+}19]. XDMoD [SDF^{+}19]. XE6 [KBVH14]. Xen [SHLJ13]. XIA [ZZJ17]. XML [ARP14, CLAL19, SW17]. XSACd [FSM^{+}18b]. XSACd-Cross-domain [FSM^{+}18b]. XSEDE [ZZBZ19]. XSS [GGC18].

Yarn [ZHW19]. Yin [QXZ^{+}19]. Yin-nourishing [QXZ^{+}19]. yoking [LNY^{+}18]. yoking-proof [LNY^{+}18]. YOLOv2 [SLTK19]. yourself [LMZ^{+}14].

Zebra [PRW14]. zero [AIA^{+}18a, AIA^{+}18b, HKA^{+}18]. zero-watermarking [AIA^{+}18a,

AIA^{+}18b].  Zeus [SPD^{+}19].  ZigBee
[ADBM19].  zone [FJJ^{+}18].  zones

[JCMPPC^{+}18]. zoo [MCRB19].

References

Adhikari:2018:HBL

[AA18] Mainak Adhikari and Tarachand Amgoth. Heuristicbased load-balancing algorithm for IaaS cloud. Future Generation Computer Systems, 81(??):156–165, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1730732X.

Alic:2019:BBD

[AAA^{+}19] Andy S. Alic, Jussara Almeida, Giovanni Aloisio, Nazareno Andrade, Nuno Antunes, Danilo Ardagna, Rosa M. Badia, Tania Basso, Ignacio Blanquer, Tarciso Braz, Andrey Brito, Donatello Elia, Sandro Fiore, Dorgival Guedes, Marco Lattuada, Daniele Lezzi, Matheus Maciel, Wagner Meira, Demetrio Mestre, Regina Moraes, Fabio Morais, Carlos Eduardo Pires, Nádia P. Kozievitch, Walter dos Santos, Paulo Silva, and Marco Vieira. BIGSEA: a big data analytics platform for public transportation information. Future Generation Computer Systems, 96(??):243–269, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18304448.

Al-Ayyoub:2018:RSP

[AAAQJ^{+}18] Mahmoud Al-Ayyoub, Muneera Al-Quraan, Yaser Jararweh, Elhadj Benkhelifa, and Salim Hariri. Resilient service provisioning in cloud based data centers. Future Generation Computer Systems, 86(??):765–774, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17314681.

Aiftimiei:2010:DIG

[AAB^{+}10] Cristina Aiftimiei, Paolo Andreetto, Sara Bertocco, Simone Dalla Fina, Alvise Dorigo, Eric Frizziero, Alessio Gianelle, Moreno Marzolla, Mirco Mazzucato, Massimo Sgaravatto, Sergio Traldi, and Luigi Zangrando. Design and implementation of the gLite CREAM job management service. Future Generation Computer Systems, 26(4): 654–667, April 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ali:2019:CFT

[AAC^{+}19] Mohammed Aamir Ali, Muhammad Ajmal Azad, Mario Parreno Centeno, Feng Hao, and Aad van Moorsel. Consumerfacing technology fraud:

Economics, attack methods and potential solutions. Future Generation Computer Systems, 100(??):408–427, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18316984.

Aoun:2013:TOA

[AAD^{+}13] Rosy Aoun, Chinwe E. Abosi, Elias A. Doumith, Reza Nejabati, Maurice Gagnaire, and Dimitra Simeonidou. Towards an optimized abstracted topology design in cloud environment. Future Generation Computer Systems, 29(1):46–60, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000891.

Abdou:2018:SAF

[AAF18] Mohamed Abdou, Sayed AbdelGaber, and Marwa Farhan. A semi-automated framework for semantically annotating web content. Future Generation Computer Systems, 81(??):94– 102, April 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17307343.

Asim:2019:ATF

[AAI^{+}19] Muhammad Asim, Muhammad Faisal Amjad, Waseem Iqbal, Hammad Afzal, Haider Abbas, and Yin Zhang. AndroKit: a toolkit for forensics analysis of web browsers on Android platform. Future Generation Computer Systems, 94(??): 781–794, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18308070.

Aljawarneh:2017:CSE

[AAJ17] Shadi A. Aljawarneh, Ali Alawneh, and Reem Jaradat. Cloud security engineering: Early stages of SDLC. Future Generation Computer Systems, 74(??):385–392, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303788.

Aroca:2016:PEA

[AAM^{+}16] Jordi Arjona Aroca, Antonio Fernández Anta, Miguel A. Mosteiro, Christopher Thraves, and Lin Wang. Power-efficient assignment of virtual machines to physical machines. Future Generation Computer Systems, 54(??): 82–94, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000072.

Amin:2019:DLE

[AAM^{+}19] Syed Umar Amin, Mansour Alsulaiman, Ghulam Muhammad, Mohamed Amine Mekhtiche, and M. Shamim Hossain. Deep learning for EEG motor imagery classification based on multi-layer CNNs feature fusion. Future Generation Computer Systems, 101(??):542–554, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19306077.

Abdulhay:2018:GTI

[AAN^{+}18] Enas Abdulhay, N. Arunkumar, Kumaravelu Narasimhan, Elamaran Vellaiappan, and V. Venkatraman. Gait and tremor investigation using machine learning techniques for the diagnosis of Parkinson disease. Future Generation Computer Systems, 83(??):366–373, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1731631X.

Alam:2019:JPC

[AAQ^{+}19] Sheraz Alam, Naveed Aqdas, I. M. Qureshi, Sajjad A. Ghauri, and Mubashar Sarfraz. Joint power and channel allocation scheme for IEEE 802.11af based smart grid communication network. Future Generation Computer Systems, 95(??):694– 712, June 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18315863.

Arshad:2019:PAC

[AAS^{+}19] Junaid Arshad, Muhammad Ajmal Azad, Khaled Salah, Razi Iqbal, Muhammad Imran Tariq, and Tariq Umer. Performance analysis of content discovery for ad-hoc tactile networks. Future Generation Computer Systems, 94(??): 726–739, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831834X.

AlMamun:2017:CBF

[AASI17] Khondaker Abdullah Al Mamun, Musaed Alhussein, Kashfia Sailunaz, and Mohammad Saiful Islam.

Cloud based framework for Parkinson’s disease diagnosis and monitoring system for remote healthcare applications. Future Generation Computer Systems, 66(??):36–47, January 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003568.

Abid:2019:SAT

[AAYL19] Fazeel Abid, Muhammad Alam, Muhammad Yasir, and Chen Li. Sentiment analysis through recurrent variants latterly on convolutional neural network of Twitter. Future Generation Computer Systems, 95(??):292–308, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18324944.

Antonescu:2016:SSB

[AB16]  Alexandru-Florian  An-

tonescu and Torsten Braun. Simulation of SLA-based VM-scaling algorithms for cloud-distributed applications. Future Generation Computer Systems, 54(??):260–273, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000321.

Arabnejad:2017:MQC

[AB17] Hamid Arabnejad and Jorge G. Barbosa. Multi-QoS constrained and profitaware scheduling approach for concurrent workflows on heterogeneous systems. Future Generation Computer Systems, 68(??):211– 221, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303740.

See reprint [AB18c].

Ali:2018:DHG

[AB18a] Ihtisham Ali and Susmit Bagchi. Designing hybrid graph model and algorithmic analysis of workflow decomposition in mobile distributed systems. Future Generation Computer Systems, 86(??):145–161, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1731943X.

Anderson:2018:SOF

[AB18b] Mark Anderson and Joseph Bolton. A service-oriented framework for collating retail intelligence. Future Generation Computer Systems, 80(??):400–408, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17314693.

Arabnejad:2018:RMQ

[AB18c] Hamid Arabnejad and Jorge G. Barbosa. Reprint of “Multi-QoS constrained and Profit-aware scheduling approach for concurrent workflows on heterogeneous systems”. Future Generation Computer Systems, 78 (part 1)(??):402– 412, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306379.

See [AB17].

Abdel-Basset:2019:CFR

[AB19a] Mohamed Abdel-Basset. Corrigendum to “A framework for risk assessment, management and evaluation: Economic tool for quantifying risks in supply chain” [Future Gener. Comput. Syst. 90 (2019) 489–502]. Future Generation Computer Systems, 93(??):1076–1077, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/

science/article/pii/S0167739X18331959.

See [ABGMC19] and retraction notice [AB21].

Alavi:2019:OST

[AB19b] Amir H. Alavi and William G.

Buttlar. An overview of smartphone technology for citizen-centered, real-time and scalable civil infrastructure monitoring. Future Generation Computer Systems, 93(??):651–672, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18315875.

Ali:2019:PNL

[AB19c] Moazam Ali and Susmit Bagchi. Probabilistic normed load monitoring in large scale distributed systems using mobile agents. Future Generation Computer Systems, 96(??):148– 167, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18309841.

Abdel-Basset:2021:RNC

[AB21] Mohamed Abdel-Basset.

Retraction notice to “Corrigendum to “A framework for risk assessment, management and evaluation: Economic tool for quantifying risks in supply chain”’ [Future Gener. Comput. Syst. (2019) 1076–1077]. Future Generation Computer Systems, 120(??): 127, July 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000972.

See [AB19a].

AlOmar:2019:PFP

[ABB^{+}19a] Abdullah Al Omar, Md Zakirul Alam Bhuiyan, Anirban Basu, Shinsaku Kiyomoto, and Mohammad Shahriar Rahman. Privacy-friendly platform for healthcare data in cloud based on blockchain environment. Future Generation Computer Systems, 95(??):511– 521, June 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314201.

Athmani:2019:EED

[ABB19b] Samir Athmani, Azeddine Bilami, and Djallel Eddine Boubiche. EDAK: An efficient dynamic authentication and key management mechanism for heterogeneous WSNs. Future Generation Computer Systems, 92(??):789–799, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17315388.

Adame:2018:CRW

[ABC^{+}18] Toni Adame, Albert Bel, Anna Carreras, Joan Melià-Seguí, Miquel Oliver, and Rafael Pous. CUIDATS: an RFID-WSN hybrid monitoring system for smart health care environments. Future Generation Computer Systems, 78 (part 2)(??):602–615, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308007.

Ahmad:2019:SCN

[ABD^{+}19] Awais Ahmad, Muhammad Babar, Sadia Din, Shehzad Khalid, Muhammad Mazhar Ullah, Anand Paul, Alavalapati Goutham Reddy, and Nasro Min-Allah. Socio-cyber network: The potential of cyber-physical system to define human behaviors using big data analytics. Future Generation Computer Systems, 92(??):868–878, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17307781.

Abusitta:2019:DLA

[ABDH19] Adel Abusitta, Martine Bellaiche, Michel Dagenais, and Talal Halabi. A deep learning approach for proactive multi-cloud cooperative intrusion detection system. Future Generation Computer Systems, 98(??):308–318, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18326992.

Ammendola:2015:AAV

[ABF^{+}15a] Roberto Ammendola, Andrea Biagioni, Ottorino Frezza, Werner Geurts, Gert Goossens, Francesca Lo Cicero, Alessandro Lonardo, Pier Stanislao Paolucci, Davide Rossetti, Francesco Simula, Laura Tosoratto, and Piero Vicini. ASIP acceleration for virtual-to-physical address translation on RDMA-enabled FPGAbased network interfaces. Future Generation Computer Systems, 53(??):109– 118, December 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002726.

Ammendola:2015:HWM

[ABF^{+}15b] Roberto Ammendola, Andrea Biagioni, Ottorino Frezza, Francesca Lo Cicero, Alessandro Lonardo, Pier Stanislao Paolucci, Davide Rossetti, Francesco Simula, Laura Tosoratto, and Piero Vicini. A hierarchical watchdog mechanism for systemic fault awareness on distributed systems. Future Generation Computer Systems, 53(??):90–99, December 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002751.

Alper:2017:SAT

[ABG17] Pinar Alper, Khalid Belhajjame, and Carole A. Goble. Static analysis of Taverna workflows to predict provenance patterns. Future Generation Computer Systems, 75(??):310– 329, October 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300225.

Alabdulkarim:2018:BSB

[ABG18] Yazeed Alabdulkarim, Sumita Barahmand, and Shahram Ghandeharizadeh. BG: A scalable benchmark for interactive social networking actions. Future Generation Computer Systems, 85(??):29–38, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17323373.

Abdel-Basset:2019:FRA

[ABGMC19] Mohamed Abdel-Basset, M. Gunasekaran, Mai Mohamed, and Naveen Chilamkurti. A framework for risk assessment, management and evaluation: Economic tool for quantifying risks in supply chain. Future Generation Computer Systems, 90(??):489– 502, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18312172.

See corrigendum [AB19a] and retraction notice [AB21].

Azad:2018:PVD

[ABH18] Muhammad Ajmal Azad, Samiran Bag, and Feng Hao. PrivBox: Verifiable decentralized reputation system for online marketplaces. Future Generation Computer Systems, 89(??):44–57, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17330315.

Abdel-Basset:2019:NPF

[ABM19] Mohamed Abdel-Basset and Mai Mohamed. A novel and powerful framework based on neutrosophic sets to aid patients with cancer. Future Generation Computer Systems, 98(??):144–153, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18325664.

See retraction notice [AB21].

Abdel-Basset:2018:NFE

[ABMC18] Mohamed Abdel-Basset, Mai Mohamed, and Victor Chang. NMCDA: A framework for evaluating cloud computing services. Future Generation Computer Systems, 86(??):12–29, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17327814.

Abdel-Basset:2018:HWO

[ABMESM18] Mohamed Abdel-Basset, Gunasekaran Manogaran, Doaa El-Shahat, and Seyedali Mirjalili. A hybrid whale optimization algorithm based on local search strategy for the permutation flow shop scheduling problem. Future Generation Computer Systems, 85(??):129–145, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18302735.

See retraction notice [ABMMC22].

Abdel-Basset:2018:ITI

[ABMM18] Mohamed Abdel-Basset, Gunasekaran Manogaran, and Mai Mohamed. Internet of Things (IoT) and its impact on supply chain: A framework for building smart, secure and efficient systems. Future Generation Computer Systems, 86(??):614–628, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1830400X.

See retraction notice [ABMMC22].

Abdel-Basset:2018:TWD

[ABMMC18] Mohamed Abdel-Basset, Gunasekaran Manogaran, Mai Mohamed, and Naveen Chilamkurti. Three-way decisions based on neutrosophic sets and AHP– QFD framework for supplier selection problem. Future Generation Computer Systems, 89(??):19–30, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18310859.

See retraction notice [ABMMC22].

Abdel-Basset:2022:RNT

[ABMMC22] Mohamed Abdel-Basset, Gunasekaran Manogaran, Mai Mohamed, and Naveen Chilamkurti. Retraction notice to “Three-way decisions based on neutrosophic sets and AHP– QFD framework for supplier selection problem” [Future Gener. Comput. Syst. 89 (2018) 19–30]. Future Generation Computer Systems, 128(??): 569, March 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004477.

See [ABMMC18].

Arabnejad:2017:SDC

[ABN17] Vahid Arabnejad, Kris Bubendorfer, and Bryan Ng. Scheduling deadline constrained scientific workflows on dynamically provisioned cloud resources. Future Generation Computer Systems, 75(??):348– 364, October 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300201.

Arabnejad:2019:DMW

[ABN19] Vahid Arabnejad, Kris Bubendorfer, and Bryan Ng. Dynamic multiworkflow scheduling: a deadline and cost-aware approach for commercial clouds. Future Generation Computer Systems, 100(??):98–108, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18307386.

Arabnejad:2016:LTC

[ABP16] Hamid Arabnejad, Jorge G. Barbosa, and Radu Prodan. Low-time complexity budget-deadline constrained workflow scheduling on heterogeneous resources. Future Generation Computer Systems, 55(??): 29–40, February 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002587.

Altmann:2018:ECS

[ABP18] Jörn Altmann, José Ángel Bañares, and Ioan Petri. Economics of computing services: A literature survey about technologies for an economy of fungible cloud services. Future Generation Computer Systems, 87(??):828–830, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18313062.

Afgan:2011:AIS

[ABS11] Enis Afgan, Purushotham Bangalore, and Karolj Skala. Application Information Services for distributed computing environments. Future Generation Computer Systems, 27(2):173–181, February 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

AlAjrawi:2018:BDC

[ABS^{+}18] Shams Al Ajrawi, Hayden Bialek, Mahasweta Sarkar, Ramesh Rao, and Syed Hassan Ahmed. Bidirectional channel modeling for implantable UHF– RFID transceivers in braincomputer interface applications. Future Generation Computer Systems, 88(??):683–692, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17327772.

Alhanahnah:2018:CAM

[ABTA18] Mohannad Alhanahnah, Peter Bertok, Zahir Tari, and Sahel Alouneh. Contextaware multifaceted trust framework for evaluating trustworthiness of cloud providers. Future Generation Computer Systems, 79 (part 2)(??):488– 499, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17300717.

Alamri:2016:MAS

[ABTF16] Abdullah Alamri, Peter Bertok, James A. Thom, and Adil Fahad. The mediator authorization-security model for heterogeneous semantic knowledge bases. Future Generation Computer Systems, 55(??):227– 237, February 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500059X.

Abbas:2015:CBH

[ABZK15] Assad Abbas, Kashif Bilal, Limin Zhang, and Samee U. Khan. A cloud based health insurance plan recommendation system: a user centered approach. Future Generation Computer Systems, 43–44(??):99–109, February 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001587.

Annett:2010:SSC

[AC10] James F. Annett and P. V.

Coveney. Special section: CompSci07, Computational Science 2007: Interdisciplinary challenges and perspectives, from the Grid to e-Science. Future Generation Computer Systems, 26(3):504–505, March 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Aime:2016:SPI

[AC16] X. Aimé and J. Charlet.

Social psychology insights into ontology engineering. Future Generation Computer Systems, 54(??):348– 351, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001752.

Asghari:2018:ECA

[AC18] Vahid Asghari and Mohamed Cheriet. Energy and connectivity aware resource optimization of nodes traffic distribution in smart home networks. Future Generation Computer Systems, 88(??):559– 570, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17316564.

Allen:2015:WDW

[ACBM15] M. David Allen, Adriane Chapman, Barbara Blaustein, and Lisa Mak. What do we do now? Workflows for an unpredictable world. Future Generation Computer Systems, 42(??):1–10, January 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001526.

Aufaure:2016:BIS

[ACC^{+}16] Marie-Aude Aufaure, Raja Chiky, Olivier Curé, Houda Khrouf, and Gabriel Kepeklian. From business intelligence to semantic data stream management. Future Generation Computer Systems, 63(??):100–107, October 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003635.

Ahmad:2019:PAP

[ACC^{+}19a] Naveed Ahmad, Haitham Cruickshank, Yue Cao, Fakhri Alam Khan, Mohammad Asif, Awais Ahmad, and Gwanggil Jeon.

Privacy by architecture pseudonym framework for delay tolerant network. Future Generation Computer Systems, 93(??):979–992, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17308427.

Assante:2019:EOS

[ACC^{+}19b] M. Assante, L. Candela, D. Castelli, R. Cirillo, G. Coro, L. Frosini, L. Lelii, F. Mangiacrapa, P. Pagano, G. Panichi, and F. Sinibaldi. Enacting open science by D4Science. Future Generation Computer Systems, 101(??):555–563, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831464X.

Assante:2019:GSD

[ACC^{+}19c] Massimiliano Assante, Leonardo Candela, Donatella Castelli, Roberto Cirillo, Gianpaolo Coro, Luca Frosini, Lucio Lelii, Francesco Mangiacrapa, Valentina Marioli, Pasquale Pagano, Giancarlo Panichi, Costantino Perciante, and Fabio Sinibaldi. The gCube system: Delivering virtual research environments as-a-service. Future Generation Computer Systems, 95(??): 445–453, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17328364.

Anastasi:2017:QAG

[ACCD17] Gaetano F. Anastasi, Emanuele Carlini, Massimo Coppola, and Patrizio Dazzi. QoS-aware genetic cloud brokering. Future Generation Computer Systems, 75(??):1–13, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17306659.

Antequera:2019:RHR

[ACCM19] Ronny Bazan Antequera, Prasad Calyam, Arjun Ankathatti Chandrashekara, and Reshmi Mitra. Recommending heterogeneous resources for science gateway applications based on custom templates composition. Future Generation Computer Systems, 100(??):281–297, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18314663.

Atzori:2019:SDS

[ACD^{+}19] Luigi Atzori, Claudia Campolo, Bin Da, Roberto Girau, Antonio Iera, Giacomo Morabito, and Salvatore Quattropani. Smart devices in the social loops: Criteria and algorithms for the creation of the social links. Future Generation Computer Systems, 97(??): 327–339, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18323598.

Agullo:2011:QOM

[ACH^{+}11] Emmanuel Agullo, Camille Coti, Thomas Herault, Julien Langou, Sylvain Peyronnet, Ala Rezmerita, Franck Cappello, and Jack Dongarra. QCG-OMPI: MPI applications on grids. Future Generation Computer Systems, 27(4):357– 369, April 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Alipio:2019:VBU

[ACHP19] Melchizedek I. Alipio, Alex Gyran A. Co, Melinda Faye C. Hilario, and Christian Miguel C. Pama. Value-based utility implementation in softwaredefined testbed for sensor data traffic management. Future Generation Computer Systems, 101(??):737–746, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19307873.

Anjos:2015:MSD

[ACK^{+}15] Julio C. S. Anjos, Iván Carrera, Wagner Kolberg, Andre Luis Tibola, Luciana B. Arantes, and Claudio R. Geyer. MRA++: Scheduling and data placement on MapReduce for heterogeneous environments. Future Generation Computer Systems, 42(??):22– 35, January 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001642.

Campo:2019:CHU

[àCKPM19] Simon à Campo, Vasssilis-Javed Khan, Konstantinos Papangelis, and Panos Markopoulos. Community heuristics for user interface evaluation of crowdsourcing platforms. Future Generation Computer Systems, 95(??):775–789, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17307136.

Akram:2018:SPT

[ACL^{+}18] Raja Naeem Akram, Hsiao-Hwa Chen, Javier Lopez, Damien Sauveron, and Laurence T. Yang. Security, privacy and trust of user-centric solutions. Future Generation Computer Systems, 80(??):417– 420, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17326146.

Amato:2018:MSC

[ACM^{+}18] Flora Amato, Aniello Castiglione, Fabio Mercorio, Mario Mezzanzanica, Vincenzo Moscato, Antonio Picariello, and Giancarlo Sperlì. Multimedia story creation on social networks. Future Generation Computer Systems, 86(??):412–420, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17322483.

Amato:2019:ADF

[ACMM19] F. Amato, G. Cozzolino, V. Moscato, and F. Moscato. Analyse digital forensic evidences through a semantic-based methodology and NLP techniques. Future Generation Computer Systems, 98(??):297– 307, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19301906.

Asensio:2019:EHA

[ACPI19] J. A. Asensio, J. Criado, N. Padilla, and L. Iribarne. Emulating home automation installations through component-based web technology. Future Generation Computer Systems, 93(??):777–791, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17306088.

Alberti:2017:NNR

[ACSdRR17] Antonio Marcos Alberti, Marco Aurelio Favoreto Casaroli, Dhananjay Singh, and Rodrigo da Rosa Righi. Naming and name resolution in the future Internet: Introducing the NovaGenesis approach. Future Generation Computer Systems, 67(??):163–179, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302643.

Ardagna:2018:CAD

[ACSV18] Danilo Ardagna, Cinzia Cappiello, Walter Samá, and Monica Vitali. Contextaware data quality assessment for big data. Future Generation Computer Systems, 89(??):548–562, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17329151.

Ahmad:2019:IAS

[ACWJ19] Awais Ahmad, Salvatore Cuomo, Wei Wu, and Gwanggil Jeon. Intelligent algorithms and standards for interoperability in Internet of Things. Future Generation Computer Systems, 92(??):1187–1191, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18328516.

A:2018:AML

[AD18] Ajay Kumara M. A. and Jaidhar C. D. Automated multi-level malware detection system based on reconstructed semantic view of executables using machine learning techniques at VMM. Future Generation Computer Systems, 79 (part 1)(??):431– 446, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17311809.

Agarwal:2019:SVM

[AD19] Amit Agarwal and Ta Nguyen Binh Duong. Secure virtual machine placement in cloud data centers. Future Generation Computer Systems, 100(??):210–222, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18326116.

Aldossary:2019:EAC

[ADA^{+}19]  Mohammad  Aldossary,

Karim Djemame, Ibrahim Alzamil, Alexandros Kostopoulos, Antonis Dimakis, and Eleni Agiatzidou. Energyaware cost prediction and pricing of virtual machines in cloud computing environments. Future Generation Computer Systems, 93(??):442–459, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18310288.

Al-Daoud:2012:PAL

[ADAAD12] Hadil Al-Daoud, Issam Al-Azzoni, and Douglas G. Down. Power-aware linear programming based scheduling for heterogeneous computer clusters. Future Generation Computer Systems, 28(5):745– 754, May 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000525.

Al-Dubi:2014:SIU

[ADALZ14] Ahmed Al-Dubi, Shoukat Ali, Lei Liu, and Dakai Zhu. Special issue on ubiquitous computing and future communication systems. Future Generation Computer Systems, 39(??): 1–2, October 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000971.

Abane:2019:NZZ

[ADBM19] Amar Abane, Mehammed Daoui, Samia Bouzefrane, and Paul Muhlethaler. NDN-over-ZigBee: A Zig-Bee support for Named Data Networking. Future Generation Computer Systems, 93(??):792–798, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17303928.

Armant:2018:SOT

[ADBO18] Vincent Armant, Milan De Cauwer, Kenneth N. Brown, and Barry O’Sullivan. Semi-online task assignment policies for workload consolidation in cloud computing systems. Future Generation Computer Systems, 82(??):89– 103, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17319143.

Ateniese:2016:ECS

[ADDV16] Giuseppe Ateniese, Özgür Dagdelen, Ivan Damgård, and Daniele Venturi. Entangled cloud storage. Future Generation Computer Systems, 62(??):104–118, September 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16000108.

Almenar:2016:EGM

[ADH^{+}16] Francisco Almenar, Carlos Domínguez, Houcine Hassan, Juan-Miguel Martínez, and Pedro López. Embedded GPU and multicore processors for emotionalbased mobile robotic agents. Future Generation Computer Systems, 56(??):192– 201, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001855.

Aguilera:2014:AAS

[AdI14] Unai Aguilera and Diego López de Ipiña. An architecture for automatic service composition in MANET using a distributed service graph. Future Generation Computer Systems, 34(??): 176–189, May 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002434.

Arleo:2018:PDG

[ADLM18] Alessio Arleo, Walter Didimo, Giuseppe Liotta, and Fabrizio Montecchiani. Profiling distributed graph processing systems through visual analytics. Future Generation Computer Systems, 87(??):43–57, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17323828.

Al-Dubai:2012:ESI

[ADLW12] Ahmed Al-Dubai, Xue Liu, and Yulei Wu. Editorial to special issue: Recent advances in mobile and ubiquitous computing. Future Generation Computer Systems, 28(6):900– 901, June 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000404.

Alemany:2018:EPR

[AdVAGF18] J. Alemany, E. del Val, J. Alberola, and A. García-Fornes. Estimation of privacy risk through centrality metrics. Future Generation Computer Systems, 82(??):63–76, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17310956.

Alam:2018:SPE

[AEK^{+}18] Masoom Alam, Naina Emmanuel, Tanveer Khan, Abid Khan, Nadeem Javaid, Kim-Kwang Raymond Choo, and Rajkumar Buyya. Secure policy execution using reusable garbled circuit in the cloud. Future Generation Computer Systems, 87(??):488–501, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17315273.

Azgomi:2010:TSM

[AEM10] Mohammad Abdollahi Azgomi and Reza Entezari-Maleki. Task scheduling modelling and reliability evaluation of grid services using coloured Petri nets. Future Generation Computer Systems, 26(8): 1141–1150, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ataie:2018:PAP

[AEME^{+}18] Ehsan Ataie, Reza Entezari-Maleki, Sayed Ehsan Etesami, Bernhard Egger, Danilo Ardagna, and Ali Movaghar. Poweraware performance analysis of self-adaptive resource management in IaaS clouds. Future Generation Computer Systems, 86(??):134–144, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17326638.

Angskun:2010:SHN

[AFB^{+}10] Thara Angskun, Graham Fagg, George Bosilca, Jelena Pješivac-Grbović, and Jack Dongarra. Selfhealing network for scalable fault-tolerant runtime environments. Future Generation Computer Systems, 26(3):479–485, March 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Acharya:2018:AIS

[AFO^{+}18] U. Rajendra Acharya, Hamido Fujita, Shu Lih Oh, U. Raghavendra, Jen Hong Tan, Muhammad Adam, Arkadiusz Gertych, and Yuki Hagiwara. Automated identification of shockable and nonshockable life-threatening ventricular arrhythmias using convolutional neural network. Future Generation Computer Systems, 79 (part 3)(??):952–959, February 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17315248.

Agosti:2016:DLI

[AFS16] Maristella Agosti, Nicola Ferro, and Gianmaria Silvello. Digital library interoperability at high level of abstraction. Future Generation Computer Systems, 55(??):129–146, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003003.

Al-Faifi:2018:PPM

[AFSH^{+}18] Abdullah Mohammed Al-Faifi, Biao Song, Mohammad Mehedi Hassan, Atif Alamri, and Abdu Gumaei. Performance prediction model for cloud service selection from smart data. Future Generation Computer Systems, 85(??):97– 106, August 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18300141.

Al-Faifi:2019:HMC

[AFSH^{+}19] Abdullah Al-Faifi, Biao Song, Mohammad Mehedi Hassan, Atif Alamri, and Abdu Gumaei. A hybrid multi criteria decision method for cloud service selection from smart data. Future Generation Computer Systems, 93(??):43– 57, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18313918.

Antunes:2016:SSA

[AGA16] Mário Antunes, Diogo Gomes, and Rui L. Aguiar. Scalable semantic aware context storage. Future Generation Computer Systems, 56(??):675–683, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002885.

Antunes:2018:TID

[AGA18] Mário Antunes, Diogo Gomes, and Rui L. Aguiar. Towards IoT data classification through semantic features. Future Generation Computer Systems, 86(??):792–798, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17327553.

Atighehchi:2019:GHC

[AGBR19] Kevin Atighehchi, Loubna Ghammam, Morgan Barbier, and Christophe Rosenberger. GREYC-Hashing: Combining biometrics and secret for enhancing the security of protected templates. Future Generation Computer Systems, 101(??):819–830, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1833125X.

Anta:2018:CAF

[AGKZ18] Antonio Fernández Anta, Chryssis Georgiou, Dariusz R. Kowalski, and Elli Zavou. Competitive analysis of fundamental scheduling algorithms on a faultprone machine and the impact of resource augmentation. Future Generation Computer Systems, 78 (part 1)(??):245–256, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301674.

Atkinson:2017:SWP

[AGMT17] Malcolm Atkinson, Sandra Gesing, Johan Montagnat, and Ian Taylor. Scientific workflows: Past, present and future. Future Generation Computer Systems, 75(??):216–227, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17311202.

Airehrour:2019:SRS

[AGR19] David Airehrour, Jairo A.

Gutierrez, and Sayan Kumar Ray. SecTrust-RPL: A secure trust-aware RPL routing protocol for Internet of Things. Future Generation Computer Systems, 93(??):860–876, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17306581.

Aziz:2011:VDP

[AH11] Benjamin Aziz and Geoff Hamilton. Verifying a delegation protocol for grid systems. Future Generation Computer Systems, 27(5):476–485, May 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Acharya:2019:CFE

[AHD^{+}19] U. Rajendra Acharya, Yuki Hagiwara, Sunny Nitin Deshpande, S. Suren, Joel En Wei Koh, Shu Lih Oh, N. Arunkumar, Edward J. Ciaccio, and Choo Min Lim. Characterization of focal EEG signals: A review. Future Generation Computer Systems, 91(??):290–299, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18318818.

Acevedo:2017:CPF

[AHEM17] César Acevedo, Porfidio Hernández, Antonio Espinosa, and Víctor Méndez. A critical path file location (CPFL) algorithm for data-aware multiworkflow scheduling on HPC clusters. Future Generation Computer Systems, 74(??):51–62, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17306507.

Alawneh:2011:EFR

[AHL11] Luay Alawneh and Abdelwahab Hamou-Lhadj. An exchange format for representing dynamic information generated from High Performance Computing applications. Future Generation Computer Systems, 27(4):381–394, April 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ali:2018:ECM

[AHM^{+}18] Zulfiqar Ali, M. Shamim Hossain, Ghulam Muhammad, Ihsan Ullah, Hamid Abachi, and Atif Alamri. Edge-centric multimodal authentication system using encrypted biometric templates. Future Generation Computer Systems, 85(??):76–87, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17328741.

Ali:2018:IHS

[AHMS18] Zulfiqar Ali, M. Shamim Hossain, Ghulam Muhammad, and Arun Kumar Sangaiah. An intelligent healthcare system for detection and classification to discriminate vocal fold disorders. Future Generation Computer Systems, 85(??):19–28, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324421.

Aktas:2016:SFC

[AHP16] Mehmet Fatih Aktas, Georgiana Haldeman, and Manish Parashar. Scheduling and flexible control of bandwidth and in-transit services for end-to-end application workflows. Future Generation Computer Systems, 56(??):284–294, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002915.

Alali:2018:CBC

[AHP^{+}18] Fatma Alali, Nathan Hanford, Eric Pouyoul, Raj Kettimuthu, Mariam Kiran, Ben Mack-Crane, Brian Tierney, Yatish Kumar, and Dipak Ghosal. Calibers: a bandwidth calendaring paradigm for science workflows. Future Generation Computer Systems, 89(??):736–745, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18302437.

Abawajy:2018:ICT

[AHS^{+}18] Jemal Abawajy, Shamsul Huda, Shaila Sharmeen, Mohammad Mehedi Hassan, and Ahmad Almogren. Identifying cyber threats to mobile-IoT applications in edge computing paradigm. Future Generation Computer Systems, 89(??):525– 538, December 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18300906.

Alam:2019:ACO

[AHU^{+}19] Md Golam Rabiul Alam, Mohammad Mehedi Hassan, Md. Zia Uddin, Ahmad Almogren, and Giancarlo Fortino. Autonomic computation offloading in mobile edge for IoT applications. Future Generation Computer Systems, 90(??):149–157, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18303996.

Alhakbani:2019:EEM

[AHYF19] Noura Alhakbani, Mohammad Mehedi Hassan, Mourad Ykhlef, and Giancarlo Fortino. An efficient event matching system for semantic smart data in the Internet of Things (IoT) environment. Future Generation Computer Systems, 95(??):163– 174, June 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831478X.

Ali:2018:CBR

[AIA^{+}18a] Zulfiqar Ali, Muhammad Imran, Mansour Alsulaiman, Muhammad Shoaib, and Sana Ullah. Chaos-based robust method of zerowatermarking for medical signals. Future Generation Computer Systems, 88(??):400–412, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18304709.

Ali:2018:ZWA

[AIA^{+}18b] Zulfiqar Ali, Muhammad Imran, Mansour Alsulaiman, Tanveer Zia, and Muhammad Shoaib. A zero-watermarking algorithm for privacy protection in biomedical signals. Future Generation Computer Systems, 82(??):290– 303, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17322975.

Amin:2018:RAP

[AIB^{+}18] Ruhul Amin, SK Hafizul Islam, G. P. Biswas, Muhammad Khurram Khan, and Neeraj Kumar. A robust and anonymous patient monitoring system using wireless medical sensor networks. Future Generation Computer Systems, 80(??): 483–495, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301509.

Ali:2019:PRD

[AIM^{+}19] Zulfiqar Ali, Muhammad Imran, Sally Mc-Clean, Naveed Khan, and Muhammad Shoaib. Protection of records and data authentication based on secret shares and watermarking. Future Generation Computer Systems, 98(??):331–341, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18328802.

AlSallal:2019:IAI

[AIP^{+}19] Muna AlSallal, Rahat Iqbal, Vasile Palade, Saad Amin, and Victor Chang. An integrated approach for intrinsic plagiarism detection. Future Generation Computer Systems, 96(??): 700–712, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17326018.

Akatyev:2019:EII

[AJ19] Nikolay Akatyev and Joshua I. James. Evidence identification in IoT networks based on threat assessment. Future Generation Computer Systems, 93(??):814–821, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17300857.

Afrin:2019:MOR

[AJR^{+}19] Mahbuba Afrin, Jiong Jin, Ashfaqur Rahman, Yu-Chu Tian, and Ambarish Kulkarni. Multi-objective resource allocation for edge cloud based robotic workflow in smart factory. Future Generation Computer Systems, 97(??):119–130, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326785.

Albodour:2012:HLQ

[AJY12] Reda Albodour, Anne James, and Norlaily Yaacob. High level QoS-driven model for grid applications in a simulated environment. Future Generation Computer Systems, 28(7): 1133–1144, July 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002159.

Abraham:2015:GBP

[AJY15a] Goodhead T. Abraham, Anne James, and Norlaily Yaacob. Group-based parallel multi-scheduler for grid computing. Future Generation Computer Systems, 50(??):140–153, September 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000291.

Albodour:2015:QWB

[AJY15b] Reda Albodour, Anne James, and Norlaily Yaacob. QoS within business grid quality of service (BGQoS). Future Generation Computer Systems, 50(??):22–37, September 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002210.

Altmann:2014:CMB

[AK14] Jörn Altmann and Mohammad Mahdi Kashef. Cost model based service placement in federated hybrid clouds. Future Generation Computer Systems, 41(??):79–90, December 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001629.

Aslan:2018:TRA

[AK18a] Serpil Aslan and Mehmet Kaya. Topic recommendation for authors as a link prediction problem. Future Generation Computer Systems, 89(??):249–264, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18302711.

Aujla:2018:MES

[AK18b] Gagangeet Singh Aujla and Neeraj Kumar. MEnSuS: An efficient scheme for energy management with sustainability of cloud data centers in edge-cloud environment. Future Generation Computer Systems, 86(??):1279–1300, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17321581.

Ahuja:2019:NFD

[AK19] Kiran Ahuja and Arun Khosla. A novel framework for data acquisition and ubiquitous communication provisioning in smart cities. Future Generation Computer Systems, 101(??):785–803, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19306211.

Amin:2018:LWA

[AKB^{+}18a] Ruhul Amin, Neeraj Kumar, G. P. Biswas, R. Iqbal, and Victor Chang. A light weight authentication protocol for IoTenabled devices in distributed cloud computing environment. Future Generation Computer Systems, 78 (part 3)(??):1005–1019, January 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630824X.

Anton:2018:UEI

[AKB18b] David Anton, Gregorij Kurillo, and Ruzena Bajcsy. User experience and interaction performance in 2D/3D telecollaboration. Future Generation Computer Systems, 82(??):77– 88, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17323385.

Al-khafajiy:2019:IFC

[AkBAL^{+}19] Mohammed Al-khafajiy, Thar Baker, Hilal Al-Libawy, Zakaria Maamar, Moayad Aloqaily, and Yaser Jararweh. Improving fog computing performance via Fog-2-Fog collaboration. Future Generation Computer Systems, 100(??):266–280, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331868.

Al-Kiswany:2017:CLO

[AKCY^{+}17] Samer Al-Kiswany, Lauro B. Costa, Hao Yang, Emalayan Vairavanathan, and Matei Ripeanu. A cross-layer optimized storage system for workflow applications. Future Generation Computer Systems, 75(??):423– 437, October 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302959.

Anton:2017:RTC

[AKG^{+}17] David Antón, Gregorij Kurillo, Alfredo Goñi, Arantza Illarramendi, and Ruzena Bajcsy. Realtime communication for Kinect-based telerehabilitation. Future Generation Computer Systems, 75(??): 72–81, October 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303806.

Ammae:2018:UDB

[AKM18] Osamu Ammae, Joseph Korpela, and Takuya Maekawa. Unobtrusive detection of body movements during sleep using Wi-Fi received signal strength with model adaptation technique. Future Generation Computer Systems, 78 (part 2)(??):616– 625, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17304454.

Ahmad:2018:TMO

[AKP^{+}18] Awais Ahmad, Murad Khan, Anand Paul, Sadia Din, M. Mazhar Rathore, Gwanggil Jeon, and Gyu Sang Choi. Toward modeling and optimization of features selection in Big Data based social Internet of Things. Future Generation Computer Systems, 82(??):715–726, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17304533.

Alcaraz:2014:WDW

[AL14] Cristina Alcaraz and Javier Lopez. WASAM: a dynamic wide-area situational awareness model for critical domains in Smart Grids. Future Generation Computer Systems, 30(??):146–154, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001441.

Alkhanak:2018:HHC

[AL18] Ehab Nabiel Alkhanak and Sai Peck Lee. A hyperheuristic cost optimisation approach for scientific workflow scheduling in cloud computing. Future Generation Computer Systems, 86(??):480–506, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17328297.

Arcanjo:2016:MEV

[ALFR16] Jeferson S. Arcanjo, Eduardo F. P. Luz, Álvaro L. Fazenda, and Fernando M. Ramos. Methods for evaluating volunteers’ contributions in a deforestation detection citizen science project. Future Generation Computer Systems, 56(??): 550–557, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002289.

Alkhanak:2015:CAC

[ALK15] Ehab Nabiel Alkhanak, Sai Peck Lee, and Saif Ur Rehman Khan. Costaware challenges for workflow scheduling approaches in cloud computing environments: Taxonomy and opportunities. Future Generation Computer Systems, 50(??):3–21, September 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000242.

Au:2018:PPP

[ALL^{+}18] Man Ho Au, Kaitai Liang, Joseph K. Liu, Rongxing Lu, and Jianting Ning. Privacy-preserving personal data operation on mobile cloud: Chances and challenges over advanced persistent threat. Future Generation Computer Systems, 79 (part 1)(??):337– 349, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17312864.

Anedda:2010:SMR

[ALM^{+}10] Paolo Anedda, Simone Leo, Simone Manca, Massimo Gaggero, and Gianluigi Zanetti. Suspending, migrating and resuming HPC virtual clusters. Future Generation Computer Systems, 26(8):1063– 1072, October 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Alpar:2018:BTA

[Alp18] Orcan Alpar. Biometric touchstroke authentication by fuzzy proximity of touch locations. Future Generation Computer Systems, 86(??):71–80, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17326055.

Afgan:2019:CDD

[ALTG19] Enis Afgan, Andrew Lonie, James Taylor, and Nuwan Goonasekera. Cloud-Launch: Discover and deploy cloud applications. Future Generation Computer Systems, 94(??):802– 810, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17307793.

Arenas:2010:SSS

[AM10] Alvaro E. Arenas and Philippe Massonet. Special section: Security, trust and privacy in Grid systems. Future Generation Computer Systems, 26(7): 1024–1025, July 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Amato:2017:ECW

[AM17] Flora Amato and Francesco Moscato. Exploiting cloud and workflow patterns for the analysis of composite cloud services. Future Generation Computer Systems, 67(??):255–265, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302291.

Araghi:2019:EHI

[AM19a] Tanya Koohpayeh Araghi and Azizah Abd Manaf.

An enhanced hybrid image watermarking scheme for security of medical and non-medical images based on DWT and 2-D SVD. Future Generation Computer Systems, 101 (??):1223–1246, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310842.

Azad:2019:RDS

[AM19b] Muhammad Ajmal Azad and Ricardo Morla. Rapid detection of spammers through collaborative information sharing across multiple service providers. Future Generation Computer Systems, 95(??):841– 854, June 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17307525.

Aisilahong:2019:EEC

[AMAY19] Guhaer Aisilahong, Palida Maimaiti, Abuduxukuer Abulimiti, and Wu Yun Yun. Establishment and evaluation of contracture model of knee joint. Future Generation Computer Systems, 98(??):252–258, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1832689X.

Ahmad:2018:OOC

[AMBB18] Jamil Ahmad, Khan Muhammad, Sambit Bakshi, and Sung Wook Baik. Objectoriented convolutional features for fine-grained image retrieval in large surveillance datasets. Future Generation Computer Systems, 81(??):314–330, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17318575.

Azar:2019:EEI

[AMBC19] Joseph Azar, Abdallah Makhoul, Mahmoud Barhamgi, and Raphaël Couturier. An energy efficient IoT data compression approach for edge machine learning. Future Generation Computer Systems, 96(??):168–175, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331716.

Amrahov:2019:NIS

[AMÇ19] Sahin Emrah Amrahov, Adnan Saher Mohammed, and Fatih V. Çelebi. New and improved search algorithms and precise analysis of their average-case complexity. Future Generation Computer Systems, 95(??): 743–753, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18319307.

Alonso-Monsalve:2018:HMC

[AMGCC18] Saúl Alonso-Monsalve, Félix García-Carballeira, and Alejandro Calderón. A heterogeneous mobile cloud computing model for hybrid clouds. Future Generation Computer Systems, 87(??):651–666, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17313894.

Arnedo-Moreno:2010:JRA

[AMHJ10] Joan Arnedo-Moreno and Jordi Herrera-Joancomartí. JXTA resource access control by means of advertisement encryption. Future Generation Computer Systems, 26(1):21–28, January 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ahmed:2016:SAD

[AMI16] Mohiuddin Ahmed, Abdun Naser Mahmood, and Md. Rafiqul Islam. A survey of anomaly detection techniques in financial domain. Future Generation Computer Systems, 55(??):278–288, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000023.

Aghili:2019:SSL

[AMKC19] Seyed Farhad Aghili, Hamid Mala, Pallavi Kaliyar, and Mauro Conti. SecLAP: Secure and lightweight RFID authentication protocol for Medical IoT. Future Generation Computer Systems, 101(??):621–634, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304777.

Amiri:2018:OLM

[AMKM18] Maryam Amiri, Leyli Mohammad-Khanli, and Raffaela Mirandola. An online learning model based on episode mining for workload prediction in cloud. Future Generation Computer Systems, 87(??):83–101, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18300712.

Aljawarneh:2016:IAM

[AMM16] Shadi A. Aljawarneh, Raja A. Moftah, and Abdelsalam M. Maatuk. Investigations of automatic methods for detecting the polymorphic worms signatures. Future Generation Computer Systems, 60(??): 67–77, July 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300048.

Amory:2019:DCT

[AMM19a] Abduljawad A. Amory, Ghulam Muhammad, and Hassan Mathkour. Deep convolutional tree networks. Future Generation Computer Systems, 101(??):152–168, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19306168.

Awad:2019:CWS

[AMM^{+}19b] Samir Awad, Abdelhamid Malki, Mimoun Malki, Mahmoud Barhamgi, and Djamal Benslimane. Composing WoT services with uncertain data. Future Generation Computer Systems, 101(??):940–950, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304467.

Amato:2018:ISC

[AMMC18] Flora Amato, Francesco Moscato, Vincenzo Moscato, and Francesco Colace. Improving security in cloud by formal modeling of IaaS resources. Future Generation Computer Systems, 87(??):754–764, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17305964.

Abbasinezhad-Mood:2018:DHI

[AMN18] Dariush Abbasinezhad-Mood and Morteza Nikooghadam. Design and hardware implementation of a securityenhanced elliptic curve cryptography based lightweight authentication scheme for smart grid communications. Future Generation Computer Systems, 84(??): 47–57, July 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17315376.

Amato:2019:SMR

[AMPP19] Flora Amato, Vincenzo Moscato, Antonio Picariello, and Francesco Piccialli. SOS: A multimedia recommender system for online social networks. Future Generation Computer Systems, 93(??):914– 923, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17301693.

Amato:2019:EEM

[AMPS19] Flora Amato, Vincenzo Moscato, Antonio Picariello, and Giancarlo Sperli’ì. Extreme events management using multimedia social networks. Future Generation Computer Systems, 94(??):444–452, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18312159.

Antonic:2016:MCS

[AMPZ16] Aleksandar Antonić, Martina Marjanović, Kresimir Pripuzić, and Ivana Podnar Zarko. A mobile crowd sensing ecosystem enabled by CUPUS: Cloud-based publish/subscribe middleware for the Internet of Things. Future Generation Computer Systems, 56(??): 607–622, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002575.

Al-Muhtadi:2019:MCL

[AMQS^{+}19] Jalal Al-Muhtadi, Ma Qiang, Kashif Saleem, Manan Al-Musallam, and Joel J. P. C. Rodrigues. Misty clouds — a layered cloud platform for online user anonymity in Social Internet of Things. Future Generation Computer Systems, 92(??):812– 820, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312797.

Asghari:2018:ETF

[AMR18] Seyyed Amir Asghari, Mohammadreza Binesh Marvasti, and Amir M. Rahmani. Enhancing transient fault tolerance in embedded systems through an OS task level redundancy approach. Future Generation Computer Systems, 87(??):58–65, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18300190.

Almeida:2019:CAI

[AMR^{+}19] Aitor Almeida, Rubén Mulero, Piercosimo Rametta, Vladimir Urošević, Marina Andrić, and Luigi Patrono. A critical analysis of an IoT-aware AAL system for elderly monitoring. Future Generation Computer Systems, 97(??): 598–619, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18321769.

Atkinson:2018:YWL

[AMRM18] John S. Atkinson, John E. Mitchell, Miguel Rio, and George Matich. Your WiFi is leaking: What do your mobile apps gossip about you? Future Generation Computer Systems, 80(??): 546–557, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301480.

Arunarani:2019:TST

[AMS19] AR. Arunarani, D. Manjula, and Vijayan Sugumaran. Task scheduling techniques in cloud computing: A literature survey. Future Generation Computer Systems, 91(??):407–415, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17321519.

Aghili:2019:LLT

[AMSPL19] Seyed Farhad Aghili, Hamid Mala, Mohammad Shojafar, and Pedro Peris-Lopez. LACO: Lightweight threefactor authentication, access control and ownership transfer scheme for e-health systems in IoT. Future Generation Computer Systems, 96(??):410– 424, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331297.

Andronikou:2012:DQA

[AMT^{+}12] Vassiliki Andronikou, Konstantinos Mamouras, Konstantinos Tserpes, Dimosthenis Kyriazis, and Theodora Varvarigou. Dynamic QoS-aware data replication in grid environments based on data “importance”. Future Generation Computer Systems, 28 (3):544–553, March 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X11000069.

Abdullahi:2016:SOS

[ANA16]  Mohammed  Abdullahi,

Md Asri Ngadi, and Shafi’i Muhammad Abdulhamid. Symbiotic organism search optimization based task scheduling in cloud computing environment. Future Generation Computer Systems, 56(??): 640–650, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002599.

Abrishami:2013:DCW

[ANE13] Saeid Abrishami, Mahmoud Naghibzadeh, and Dick H. J. Epema. Deadlineconstrained workflow scheduling algorithms for Infrastructure as a Service Clouds. Future Generation Computer Systems, 29(1):158–169, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001008.

Ahmed:2019:PSS

[ANG^{+}19] Ejaz Ahmed, Anjum Naveed, Abdullah Gani, Siti Hafizah Ab Hamid, Muhammad Imran, and Mohsen Guizani. Process state synchronizationbased application execution management for mobile edge/cloud computing. Future Generation Computer Systems, 91(??):579–589, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18307970.

Anonymous:2010:AR

[Ano10]  Anonymous.  Acknowl-

edgement of reviewers over 2009. Future Generation Computer Systems, 26(4): 531–532, April 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2011:AR

[Ano11a]  Anonymous.  Acknowl-

edgement of reviewers over 2010. Future Generation Computer Systems, 27(5): 427–429, May 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2011:C

[Ano11b] Anonymous. Contents. Future Generation Computer Systems, 27(8):iii–iv, October 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001270.

Anonymous:2011:EB

[Ano11c]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 27(8): ??, October 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001245.

Anonymous:2012:AR

[Ano12a]  Anonymous.  Acknowl-

edgement of reviewers over 2011. Future Generation Computer Systems, 28(8):1155–1157, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001215.

Anonymous:2012:Ca

[Ano12b] Anonymous. Contents. Future Generation Computer Systems, 28(1):iii–iv, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001567.

Anonymous:2012:Cb

[Ano12c] Anonymous. Contents. Future Generation Computer Systems, 28(2):??, February 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001786.

Anonymous:2012:Cc

[Ano12d]  Anonymous.  Contents.

Future Generation Computer Systems, 28(3):??, March 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002007.

Anonymous:2012:Cd

[Ano12e]  Anonymous.  Contents.

Future Generation Computer Systems, 28(4):??, April 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000064.

Anonymous:2012:Ce

[Ano12f]  Anonymous.  Contents.

Future Generation Computer Systems, 28(5):??, May 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1200026X.

Anonymous:2012:Cf

[Ano12g]  Anonymous.  Contents.

Future Generation Computer Systems, 28(6):iii–iv, June 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000453.

Anonymous:2012:Cg

[Ano12h]  Anonymous.  Contents.

Future Generation Computer Systems, 28(7):iii–iv, July 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000726.

Anonymous:2012:Ch

[Ano12i] Anonymous. Contents. Future Generation Computer Systems, 28(8):iii–iv, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001203.

Anonymous:2012:EBa

[Ano12j]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 28(1): ??, January 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001531.

Anonymous:2012:EBb

[Ano12k]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 28(2): ??, February 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001750.

Anonymous:2012:EBc

[Ano12l]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 28 (3):??, March 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1100197X.

Anonymous:2012:EBd

[Ano12m]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 28 (4):??, April 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000039.

Anonymous:2012:EBe

[Ano12n]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 28 (5):??, May 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000234.

Anonymous:2012:EBf

[Ano12o]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 28 (6):??, June 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000428.

Anonymous:2012:EBg

[Ano12p]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 28 (7):??, July 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000696.

Anonymous:2012:EBh

[Ano12q]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 28(8): ??, October 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001173.

Anonymous:2012:OHA

[Ano12r] Anonymous. OPTIMIS: a holistic approach to cloud service provisioning. Future Generation Computer Systems, 28(1):66–77, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1100104X.

Anonymous:2013:Ca

[Ano13a]  Anonymous.  Contents.

Future Generation Computer Systems, 29(2):iii– iv, February 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12002026.

Anonymous:2013:Cb

[Ano13b]  Anonymous.  Contents.

Future Generation Computer Systems, 29(3):iii–iv, March 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12002130.

Anonymous:2013:Cc

[Ano13c]  Anonymous.  Contents.

Future Generation Computer Systems, 29(5):iii–iv, July 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300040X.

Anonymous:2013:Cd

[Ano13d] Anonymous. Contents. Future Generation Computer Systems, 29(6):iii–iv, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000563.

Anonymous:2013:Ce

[Ano13e]  Anonymous.  Contents.

Future Generation Computer Systems, 29(7):iii– iv, September 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000915.

Anonymous:2013:EBa

[Ano13f]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 29(1): ??, January 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001914.

Anonymous:2013:EBb

[Ano13g]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 29(2): ??, February 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001999.

Anonymous:2013:EBc

[Ano13h]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 29 (3):??, March 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12002105.

Anonymous:2013:EBd

[Ano13i]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 29 (4):??, June 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000095.

Anonymous:2013:EBe

[Ano13j]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 29(4): iii–iv, June 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000125.

Anonymous:2013:EBf

[Ano13k]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 29 (5):??, July 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300037X.

Anonymous:2013:EBg

[Ano13l]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 29 (6):??, August 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000538.

Anonymous:2013:EBh

[Ano13m]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 29(7): ??, September 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000885.

Anonymous:2014:Ca

[Ano14a]  Anonymous.  Contents.

Future Generation Computer Systems, 38(??):iii, September 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000922.

Anonymous:2014:Cb

[Ano14b] Anonymous. Contents. Future Generation Computer Systems, 39(??):iii, October 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001137.

Anonymous:2014:Cc

[Ano14c] Anonymous. Contents. Future Generation Computer Systems, 40(??):iii, November 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001459.

Anonymous:2014:Cd

[Ano14d] Anonymous. Contents. Future Generation Computer Systems, 41(??):iii, December 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001745.

Anonymous:2014:EBa

[Ano14e]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 38(??): ifc, September 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000892.

Anonymous:2014:EBb

[Ano14f]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 39(??): ifc, October 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001101.

Anonymous:2014:EBc

[Ano14g]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 40(??): ifc, November 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001423.

Anonymous:2014:EBd

[Ano14h]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 41(??): ifc, December 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1400171X.

Anonymous:2015:Ca

[Ano15a] Anonymous. Contents. Future Generation Computer Systems, 42(??):iii, January 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001952.

Anonymous:2015:Cb

[Ano15b]  Anonymous.  Contents.

Future Generation Computer Systems, 43–44(??): iii, February 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002337.

Anonymous:2015:Cc

[Ano15c]  Anonymous.  Contents.

Future Generation Computer Systems, 45(??):iii, April 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002659.

Anonymous:2015:Cd

[Ano15d]  Anonymous.  Contents.

Future Generation Computer Systems, 46(??):iii, May 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000126.

Anonymous:2015:Ce

[Ano15e]  Anonymous.  Contents.

Future Generation Computer Systems, 47(??):iii, June 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000461.

Anonymous:2015:Cf

[Ano15f]  Anonymous.  Contents.

Future Generation Computer Systems, 48(??):ibc, July 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000722.

Anonymous:2015:Cg

[Ano15g] Anonymous. Contents. Future Generation Computer Systems, 49(??):iii–iv, August 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000916.

Anonymous:2015:Ch

[Ano15h]  Anonymous.  Contents.

Future Generation Computer Systems, 50(??):iii, September 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15001673.

Anonymous:2015:Ci

[Ano15i] Anonymous. Contents. Future Generation Computer Systems, 51(??):iii–iv, October 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15001946.

Anonymous:2015:Cj

[Ano15j]  Anonymous.  Contents.

Future Generation Computer Systems, 52(??):iii– iv, November 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002125.

Anonymous:2015:EBa

[Ano15k]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 42(??): IFC, January 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001927.

Anonymous:2015:EBb

[Ano15l]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 43–44(??):IFC, February 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002301.

Anonymous:2015:EBc

[Ano15m]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 45 (??):IFC, April 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002623.

Anonymous:2015:EBd

[Ano15n]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 46 (??):IFC, May 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000096.

Anonymous:2015:EBe

[Ano15o]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 47 (??):IFC, June 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000436.

Anonymous:2015:EBf

[Ano15p]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 48 (??):ibc, July 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000692.

Anonymous:2015:EBg

[Ano15q]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 49(??): ifc, August 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000886.

Anonymous:2015:EBh

[Ano15r]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 50(??): ifc, September 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15001697.

Anonymous:2015:EBi

[Ano15s]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 51(??): ifc, October 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001910.

Anonymous:2015:EBj

[Ano15t]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 52(??): ifc, November 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002095.

Anonymous:2015:EBk

[Ano15u]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 53(??): ifc, December 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500240X.

Anonymous:2016:Ca

[Ano16a] Anonymous. Contents. Future Generation Computer Systems, 54(??):iii–vi, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002708.

Anonymous:2016:Cb

[Ano16b]  Anonymous.  Contents.

Future Generation Computer Systems, 55(??):iii– v, February 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003489.

Anonymous:2016:Cc

[Ano16c]  Anonymous.  Contents.

Future Generation Computer Systems, 56(??):iii– vi, March 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003696.

Anonymous:2016:Cd

[Ano16d]  Anonymous.  Contents.

Future Generation Computer Systems, 57(??):iii, April 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15004045.

Anonymous:2016:Ce

[Ano16e]  Anonymous.  Contents.

Future Generation Computer Systems, 58(??):iii, May 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16000182.

Anonymous:2016:Cf

[Ano16f]  Anonymous.  Contents.

Future Generation Computer Systems, 59(??):iii, June 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300139.

Anonymous:2016:Cg

[Ano16g]  Anonymous.  Contents.

Future Generation Computer Systems, 60(??):iii, July 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300395.

Anonymous:2016:Ch

[Ano16h] Anonymous. Contents. Future Generation Computer Systems, 61(??):iii, August 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300784.

Anonymous:2016:Ci

[Ano16i]  Anonymous.  Contents.

Future Generation Computer Systems, 62(??):iii– iv, September 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300991.

Anonymous:2016:Cj

[Ano16j] Anonymous. Contents. Future Generation Computer Systems, 63(??):iii–iv, October 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301613.

Anonymous:2016:Ck

[Ano16k]  Anonymous.  Contents.

Future Generation Computer Systems, 64(??):iii– iv, November 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302229.

Anonymous:2016:Cl

[Ano16l] Anonymous. Contents. Future Generation Computer Systems, 65(??):iii–iv, December 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302461.

Anonymous:2016:EBa

[Ano16m]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 54(??): ifc, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002678.

Anonymous:2016:EBb

[Ano16n]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 55(??): ifc, February 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003453.

Anonymous:2016:EBc

[Ano16o]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 56 (??):ifc, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003660.

Anonymous:2016:EBd

[Ano16p]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 57 (??):ifc, April 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500401X.

Anonymous:2016:EBe

[Ano16q]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 58 (??):ifc, May 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16000157.

Anonymous:2016:EBf

[Ano16r]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 59 (??):ifc, June 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300103.

Anonymous:2016:EBg

[Ano16s]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 60 (??):ifc, July 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630036X.

Anonymous:2016:EBh

[Ano16t]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 61(??): ifc, August 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300759.

Anonymous:2016:EBi

[Ano16u]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 62(??): ifc, September 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300966.

Anonymous:2016:EBj

[Ano16v]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 63(??): ifc, October 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301583.

Anonymous:2016:EBk

[Ano16w]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 64(??): ifc, November 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302199.

Anonymous:2016:EBl

[Ano16x]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 65(??): ifc, December 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302436.

Anonymous:2017:Ca

[Ano17a] Anonymous. Contents. Future Generation Computer Systems, 66(??):iii–iv, January 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302886.

Anonymous:2017:Cb

[Ano17b]  Anonymous.  Contents.

Future Generation Computer Systems, 67(??):iii– v, February 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16305246.

Anonymous:2017:Cc

[Ano17c]  Anonymous.  Contents.

Future Generation Computer Systems, 68(??):iii– v, March 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16306306.

Anonymous:2017:Cd

[Ano17d]  Anonymous.  Contents.

Future Generation Computer Systems, 69(??):iii, April 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16307853.

Anonymous:2017:Ce

[Ano17e]  Anonymous.  Contents.

Future Generation Computer Systems, 70(??):iii– iv, May 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301024.

Anonymous:2017:Cf

[Ano17f]  Anonymous.  Contents.

Future Generation Computer Systems, 71(??):iii– iv, June 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302868.

Anonymous:2017:Cg

[Ano17g]  Anonymous.  Contents.

Future Generation Computer Systems, 72(??):iii– iv, July 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17305307.

Anonymous:2017:Ch

[Ano17h] Anonymous. Contents. Future Generation Computer Systems, 73(??):iii, August 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17306210.

Anonymous:2017:Ci

[Ano17i] Anonymous. Contents. Future Generation Computer Systems, 75(??):iii–iv, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17311251.

Anonymous:2017:EBa

[Ano17j]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 66(??): ifc, January 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302850.

Anonymous:2017:EBb

[Ano17k]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 67(??): ifc, February 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16305210.

Anonymous:2017:EBc

[Ano17l]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 68 (??):ifc, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16306276.

Anonymous:2017:EBd

[Ano17m]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 69 (??):ifc, April 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16307828.

Anonymous:2017:EBe

[Ano17n]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 70 (??):ifc, May 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300997.

Anonymous:2017:EBf

[Ano17o]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 71 (??):ifc, June 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302832.

Anonymous:2017:EBg

[Ano17p]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 72 (??):ifc, July 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17305277.

Anonymous:2017:EBh

[Ano17q]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 73(??): ifc, August 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17306180.

Anonymous:2017:EBi

[Ano17r]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 74(??): iii–v, September 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17309779.

Anonymous:2017:EBj

[Ano17s]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 74(??): ifc, September 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17309743.

Anonymous:2017:EBk

[Ano17t]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 75(??): ifc, October 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17311226.

Anonymous:2017:EBl

[Ano17u]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 76(??): ifc, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17313936.

Anonymous:2017:EBm

[Ano17v]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 77(??): ifc, December 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1731765X.

Anonymous:2018:EBa

[Ano18a]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 78 (part 1)(??):ifc, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17319878.

Anonymous:2018:EBb

[Ano18b]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 78 (part 2)(??):ifc, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1731974X.

Anonymous:2018:EBc

[Ano18c]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 78 (part 3)(??):ifc, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17319945.

Anonymous:2018:EBd

[Ano18d]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 79 (part 1)(??):IFC– , 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17324937.

Anonymous:2018:EBe

[Ano18e]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 79 (part 2)(??):IFC– , 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17325578.

Anonymous:2018:EBf

[Ano18f]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 79 (part 3)(??):ifc, February 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17325645.

Anonymous:2018:EBg

[Ano18g]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 80 (??):ifc, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17326171.

Anonymous:2018:EBh

[Ano18h]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 81 (??):ii, April 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17328662.

Anonymous:2018:EBi

[Ano18i]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 82(??): ii, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18301195.

Anonymous:2018:EBj

[Ano18j]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 83(??): ii, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18304667.

Anonymous:2018:EBk

[Ano18k]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 84(??): ii, July 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18306824.

Anonymous:2018:EBl

[Ano18l]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 85(??): ii, August 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18309002.

Anonymous:2018:EBm

[Ano18m]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 86(??):ii, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18311324.

Anonymous:2018:EBn

[Ano18n]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 87(??): ii, October 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18313104.

Anonymous:2018:EBo

[Ano18o]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 88(??): ii, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317795.

Anonymous:2018:EBp

[Ano18p]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 89(??): ii, December 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831865X.

Anonymous:2019:EBa

[Ano19a]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 90 (??):ii, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18321563.

Anonymous:2019:EBb

[Ano19b]  Anonymous.  Editorial

board. Future Generation Computer Systems, 91(??): ii, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18324671.

Anonymous:2019:EBc

[Ano19c]  Anonymous.  Editorial

board. Future Generation Computer Systems, 92(??): ii, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18328218.

Anonymous:2019:EBd

[Ano19d]  Anonymous.  Editorial

board. Future Generation Computer Systems, 93(??): ii, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18331996.

Anonymous:2019:EBe

[Ano19e]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 94 (??):ii, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19303541.

Anonymous:2019:EBf

[Ano19f]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 95 (??):ii, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19305059.

Anonymous:2019:EBg

[Ano19g]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 96 (??):ii, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19309094.

Anonymous:2019:EBh

[Ano19h]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 97 (??):ii, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19311148.

Anonymous:2019:EBi

[Ano19i]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 100(??):ii, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1932984X.

Anonymous:2019:EBj

[Ano19j]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 101(??):ii, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19329905.

Anonymous:2019:PO

[Ano19k] Anonymous. In progress (October 2019). Future Generation Computer Systems, 99(??):??, October 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2019:PS

[Ano19l] Anonymous. In progress (September 2019). Future Generation Computer Systems, 98(??):??, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2019:PN

[Ano19m] Anonymous. Pages 1–1102 (November 2019). Future Generation Computer Systems, 100(??):1–1102, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2019:PD

[Ano19n] Anonymous. Pages 1–1294 (December 2019). Future Generation Computer Systems, 101(??):1–1294, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2019:PJb

[Ano19o] Anonymous. Pages 1–750 (July 2019). Future Generation Computer Systems, 96(??):1–750, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2019:PJa

[Ano19p] Anonymous. Pages 1–890 (June 2019). Future Generation Computer Systems, 95(??):1–890, June 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2019:PA

[Ano19q] Anonymous. Pages 1–910 (August 2019). Future Generation Computer Systems, 97(??):1–910, August 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2019:PM

[Ano19r] Anonymous. Pages 1–956 (May 2019). Future Generation Computer Systems, 94(??):1–956, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ansell:2011:MPQ

[Ans11] Peter Ansell. Model and prototype for querying multiple linked scientific datasets. Future Generation Computer Systems, 27 (3):329–333, March 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Abdullah:2010:OWA

[AOIS10] Monir Abdullah, Mohamed Othman, Hamidah Ibrahim, and Shamala Subramaniam. Optimal workload allocation model for scheduling divisible data grid applications. Future Generation Computer Systems, 26(7):971–978, July 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Abdi:2017:CMD

[APAZ17] Somayeh Abdi, Latif PourKarimi, Mahmood Ahmadi, and Farzad Zargari. Cost minimization for deadlineconstrained bag-of-tasks applications in federated hybrid clouds. Future Generation Computer Systems, 71(??):113–128, June 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301735.

Aguilera:2017:CCD

[APBdI17] Unai Aguilera, Oscar Peña, Oscar Belmonte, and Diego López de Ipiña. Citizen-centric data services for smarter cities. Future Generation Computer Systems, 76(??):234–247, November 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16304770.

Ali:2018:SUA

[APK^{+}18] Rifaqat Ali, Arup Kumar Pal, Saru Kumari, Marimuthu Karuppiah, and Mauro Conti. A secure user authentication and key-agreement scheme using wireless sensor networks for agriculture monitoring. Future Generation Computer Systems, 84(??):200–215, July 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17303862.

Azimi:2019:MDR

[APR^{+}19] Iman Azimi, Tapio Pahikkala, Amir M. Rahmani, Hannakaisa Niela-Vilén, Anna Axelin, and Pasi Liljeberg. Missing data resilient decision-making for healthcare IoT through personalization: a case study on maternal health. Future Generation Computer Systems, 96(??):297–308, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18316480.

Ahmad:2016:SCS

[APRC16] Awais Ahmad, Anand Paul, M. Mazhar Rathore, and Hangbae Chang. Smart cyber society: Integration of capillary devices with high usability based on Cyber-Physical System. Future Generation Computer Systems, 56(??):493– 503, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002563.

Alowayyed:2019:PHP

[APS^{+}19] S. Alowayyed, T. Piontek, J. L. Suter, O. Hoenen, D. Groen, O. Luk, B. Bosak, P. Kopta, K. Kurowski, O. Perks, K. Brabazon, V. Jancauskas, D. Coster, P. V. Coveney, and A. G. Hoekstra. Patterns for high performance multiscale computing. Future Generation Computer Systems, 91(??):335–346, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18300669.

Al-Qurishi:2018:PSS

[AQAR^{+}18] Muhammad Al-Qurishi, Majed Alrubaian, Sk Md Mizanur Rahman, Atif Alamri, and Mohammad Mehedi Hassan. A prediction system of Sybil attack in social network using deepregression model. Future Generation Computer Systems, 87(??):743–753, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17300821.

Aniello:2015:HFB

[AQB15] Leonardo Aniello, Leonardo Querzoni, and Roberto Baldoni. High frequency batch-oriented computations over large sliding time windows. Future Generation Computer Systems, 43–44(??):1–11, February 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001782.

Al-Qurishi:2018:EKA

[AQRH^{+}18] Muhammad Al-Qurishi, Sk Md Mizanur Rahman, M. Shamim Hossain, Ahmad Almogren, Majed Alrubaian, Atif Alamri, Mabrook Al-Rakhami, and B. B. Gupta. An efficient key agreement protocol for Sybil-precaution in online social networks. Future Generation Computer Systems, 84(??):139– 148, July 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17306623.

Antonopoulos:2010:SSM

[AR10] Nick Antonopoulos and Omer F. Rana. Special section: Management and optimisation of P2P and Grid systems with network economics. Future Generation Computer Systems, 26(8): 1282–1284, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Akande:2015:TES

[AR15] Oyindamola O. Akande and Philip J. Rhodes. Towards an efficient storage and retrieval mechanism for large unstructured grids. Future Generation Computer Systems, 45(??): 53–69, April 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002180.

Ahmed:2017:MEC

[AR17] Ejaz Ahmed and Mubashir Husain Rehmani. Mobile edge computing: Opportunities, solutions, and challenges. Future Generation Computer Systems, 70(??):59– 63, May 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303260.

Ahmed:2018:HAS

[AR18] Syed Hassan Ahmed and Shalli Rani. A hybrid approach, Smart Street use case and future aspects for Internet of Things in smart cities. Future Generation Computer Systems, 79 (part 3)(??):941–951, February 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17309949.

Altmann:2012:PSI

[ARB12] Jörn Altmann, Omer Rana, and Rajkumar Buyya. Preface of special issue on the economics of computing services. Future Generation Computer Systems, 28(8):1283–1284, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1200088X.

Al-rimy:2019:CRE

[ArMS19] Bander Ali Saleh Al-rimy, Mohd Aizaini Maarof, and Syed Zainudeen Mohd Shaid. Crypto-ransomware early detection model using novel incremental bagging with enhanced semirandom subspace selection. Future Generation Computer Systems, 101(??):476–491, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18321101.

Alghamdi:2014:SBS

[ARP14] Norah Saleh Alghamdi, Wenny Rahayu, and Eric Pardede. Semantic-based structural and content indexing for the efficient retrieval of queries over large XML data repositories. Future Generation Computer Systems, 37(??):212–231, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000351.

Afonso:2019:RPB

[ARP^{+}19] Luis C. S. Afonso, Gustavo H. Rosa, Clayton R. Pereira, Silke A. T. Weber, Christian Hook, Victor Hugo C. Albuquerque, and João P. Papa. A recurrence plot-based approach for Parkinson’s disease identification. Future Generation Computer Systems, 94(??):282–292, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18322507.

Al-Roubaiey:2019:EEA

[ARSMY19] Anas Al-Roubaiey, Tarek Sheltami, Ashraf Mahmoud, and Ansar Yasar. EATDDS: Energy-aware middleware for wireless sensor and actuator networks. Future Generation Computer Systems, 96(??): 196–206, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331777.

Atif:2014:APA

[AS14] Muhammad Atif and Peter Strazdins. Adaptive parallel application resource remapping through the live migration of virtual machines. Future Generation Computer Systems, 37(??): 148–161, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001428.

Ambigavathi:2018:EEL

[AS18a] M. Ambigavathi and D. Sridharan. Energy efficient and load balanced priority queue algorithm for Wireless Body Area Network. Future Generation Computer Systems, 88(??):586– 593, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18308458.

Anni:2018:WIS

[AS18b] Jerline Sheebha Anni and Arun Kumar Sangaiah. Wireless integrated sensor network: Boundary intellect system for elephant detection via cognitive theory and fuzzy cognitive maps. Future Generation Computer Systems, 83(??):522– 534, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302170.

Ahmed:2019:CHM

[AS19a] Tanveer Ahmed and Abhishek Srivastava. Combining humans and machines for the future: a novel procedure to predict human interest. Future Generation Computer Systems, 96(??):

713–730, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18301353.

Aldwyan:2019:LAF

[AS19b] Yasser Aldwyan and Richard O.

Sinnott. Latency-aware failover strategies for containerized web applications in distributed clouds. Future Generation Computer Systems, 101(??):1081– 1095, December 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304224.

Aydeger:2019:MTD

[ASA19] Abdullah Aydeger, Nico Saputro, and Kemal Akkaya. A moving target defense and network forensics framework for ISP networks using SDN and NFV. Future Generation Computer Systems, 94(??): 496–509, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18307817.

Anamalamudi:2018:ARP

[ASAA18] Satish Anamalamudi, Abdur Rashid Sangi, Mohammed Alkatheiri, and Ahmedin Mohammed Ahmed.

AODV routing protocol for cognitive radio access based Internet of Things (IoT). Future Generation Computer Systems, 83(??):228–238, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17319209.

Al-Sad:2019:RBD

[ASAAM^{+}19] Mohammad F. Al-Sa’d, Abdulla Al-Ali, Amr Mohamed, Tamer Khattab, and Aiman Erbad. RFbased drone detection and identification using deep learning approaches: an initiative towards a large open source drone database. Future Generation Computer Systems, 100(??):86–97, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18330760.

Al-Shara:2018:CEE

[ASAB^{+}18] Zakarea Al-Shara, Frederico Alvares, Hugo Bruneliere, Jonathan Lejeune, Charles Prud’Homme, and Thomas Ledoux. CoMe4ACloud: An end-to-end framework for autonomic cloud systems. Future Generation Computer Systems, 86(??):339–354, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17320605.

Asyabi:2018:PHC

[ASB18] Esmail Asyabi, Mohsen Sharifi, and Azer Bestavros. ppXen: A hypervisor CPU scheduler for mitigating performance variability in virtualized clouds. Future Generation Computer Systems, 83(??):75– 84, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17310555.

Amjad:2012:SDR

[ASD12] Tehmina Amjad, Muhammad Sher, and Ali Daud. A survey of dynamic replication strategies for improving data availability in data grids. Future Generation Computer Systems, 28(2):337–349, February 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001208.

Ali:2014:ALD

[ASO14] Syed Taha Ali, Vijay Sivaraman, and Diethelm Ostry. Authentication of lossy data in bodysensor networks for cloudbased healthcare monitoring. Future Generation Computer Systems, 35(??): 80–90, June 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001866.

Abuhamad:2019:CAI

[AsRA^{+}19] Mohammed Abuhamad, Ji su Rhim, Tamer AbuHmed, Sana Ullah, Sanggil Kang, and DaeHun Nyang. Code authorship identification using convolutional neural networks. Future Generation Computer Systems, 95(??):104–115, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18315528.

Asuncion:2013:ADP

[Asu13] Hazeline U. Asuncion. Automated data provenance capture in spreadsheets, with case studies. Future Generation Computer Systems, 29(8):2169–2181, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000691.

Aguilar:2013:SAD

[ASV^{+}13] Xavier Aguilar, Michael Schliephake, Olav Vahtras, Judit Gimenez, and Erwin Laure. Scalability analysis of Dalton, a molecular structure program. Future Generation Computer Systems, 29(8):2197–2204, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000733.

Ahmadi:2011:CRP

[ASW11] Mahmood Ahmadi, Asadollah Shahbahrami, and Stephan Wong. Collaboration of reconfigurable processors in grid computing: Theory and application. Future Generation Computer Systems, 27(6): 850–859, June 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ansari:2018:NML

[ASY^{+}18] Ghulam Jillani Ansari, Jamal Hussain Shah, Mussarat Yasmin, Muhammad Sharif, and Steven Lawrence Fernandes. A novel machine learning approach for scene text extraction. Future Generation Computer Systems, 87(??):328–340, October 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17321520.

Amin:2018:BDA

[ASYF18] Javeria Amin, Muhammad Sharif, Mussarat Yasmin, and Steven Lawrence Fernandes. Big data analysis for brain tumor detection: Deep convolutional neural networks. Future Generation Computer Systems, 87(??):290–297, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1732229X.

Al-Turjman:2018:ICF

[AT18a] Fadi Al-Turjman. Informationcentric framework for the Internet of Things (IoT): Traffic modeling & optimization. Future Generation Computer Systems, 80(??):63–75, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16307610.

Al-Turjman:2018:MCS

[AT18b] Fadi Al-Turjman. Mobile couriers’ selection for the smart-grid in smartcities’ pervasive sensing. Future Generation Computer Systems, 82(??):327– 341, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17310737.

Al-Turjman:2019:EDS

[AT19a] Fadi Al-Turjman. 5Genabled devices and smartspaces in social-IoT: An overview. Future Generation Computer Systems, 92(??):732–744, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17311962.

Al-Turjman:2019:CRP

[AT19b] Fadi Al-Turjman. Cognitive routing protocol for disaster-inspired Internet of Things. Future Generation Computer Systems, 92(??):1103–1115, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X16307439.

Al-Turjman:2019:IES

[ATA19] Fadi Al-Turjman and Mohammad Abujubbeh. IoTenabled smart grid via SM: an overview. Future Generation Computer Systems, 96(??):579–590, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1831759X.

Almeida:2016:PMB

[ATdC^{+}16] Fernanda Nascimento Almeida, Gisela Tunes, Julio Cezar Brettas da Costa, Ester Cerdeira Sabino, Alfredo Mendrone-Júnior, and João Eduardo Ferreira. A provenance model based on declarative specifications for intensive data analyses in hemotherapy information systems. Future Generation Computer Systems, 59(??):105– 113, June 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1500299X.

Ai:2011:PCW

[ATF11] Lifeng Ai, Maolin Tang, and Colin Fidge. Partitioning composite Web services for decentralized execution using a genetic algorithm. Future Generation Computer Systems, 27(2): 157–172, February 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

AlShaer:2019:IHS

[ATH^{+}19] Mohammed AlShaer, Yehia Taher, Rafiqul Haque, Mohand-Saïd Hacid, and Mohamed Dbouk. IB-RIDIA: a hybrid solution for processing big logistics data. Future Generation Computer Systems, 97(??): 792–804, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1830606X.

Anejionu:2019:SUD

[ATM^{+}19] Obinna C. D. Anejionu, Piyushimita (Vonu) Thakuriah, Andrew McHugh, Yeran Sun, David McArthur, Phil Mason, and Rod Walpole. Spatial urban data system: a cloud-enabled big data infrastructure for social and economic urban analytics. Future Generation Computer Systems, 98(??):456–473, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18319046.

Alamri:2014:TMO

[ATS14] Sultan Alamri, David Taniar, and Maytham Safar. A taxonomy for moving object queries in spatial databases. Future Generation Computer Systems, 37(??):232–242, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000326.

Arshad:2013:NIS

[ATX13] Junaid Arshad, Paul Townend, and Jie Xu. A novel intrusion severity analysis approach for Clouds. Future Generation Computer Systems, 29(1):416– 428, January 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001488.

Afzal:2019:EIP

[AUSA19] Bilal Afzal, Muhammad Umair, Ghalib Asadullah Shah, and Ejaz Ahmed. Enabling IoT platforms for social IoT applications: Vision, feature mapping, and challenges. Future Generation Computer Systems, 92(??):718–731, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17312724.

Aljawarneh:2017:GSA

[AVPV17] Shadi A. Aljawarneh, Radhakrishna Vangipuram, Veereswara Kumar Puligadda, and Janaki Vinjamuri. G-SPAMINE: an approach to discover temporal association patterns and trends in Internet of Things. Future Generation Computer Systems, 74(??):430–443, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300328.

Argasinski:2019:APS

[AW19] Jan K. Argasiński and Pawel Wegrzyn. Affective patterns in serious games. Future Generation Computer Systems, 92(??):526– 538, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312190.

Almuttairi:2013:TPS

[AWN^{+}13] Rafah M. Almuttairi, Rajeev Wankar, Atul Negi, C. R. Rao, Arun Agarwal, and Rajkumar Buyya. A two phased service oriented broker for replica selection in data grids. Future Generation Computer Systems, 29(4):953– 972, June 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1200180X.

Abawajy:2016:TSP

[AWYJ16] Jemal Abawajy, Guojun Wang, Laurence T. Yang, and Bahman Javadi. Trust, security and privacy in emerging distributed systems. Future Generation Computer Systems, 55(??):224–226, February 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003507.

Azzedin:2016:MBC

[AY16] Farag Azzedin and Mohammed Yahaya. Modeling BitTorrent choking algorithm using game theory. Future Generation Computer Systems, 55(??):255– 265, February 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000515.

Aazam:2018:OFC

[AZH18] Mohammad Aazam, Sherali Zeadally, and Khaled A. Harras. Offloading in fog computing for IoT: Review, enabling technologies, and research opportunities. Future Generation Computer Systems, 87(??):278–289, October 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18301973.

Anwar:2019:BBB

[AZO^{+}19] Raja Waseem Anwar, Anazida Zainal, Fatma Outay, Ansar Yasar, and Saleem Iqbal. BTEM: Belief based trust evaluation mechanism for wireless sensor networks. Future Generation Computer Systems, 96(??):605–616, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331042.

Babar:2017:SUP

[BA17] Muhammad Babar and Fahim Arif. Smart urban planning using big data analytics to contend with the interoperability in Internet of Things. Future Generation Computer Systems, 77(??):65–76, December 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17308993.

Baktir:2019:SAO

[BAA^{+}19] Ahmet Cihat Baktir, Betül Ahat, Necati Aras, Atay Özgövde, and Cem Ersoy. SLA-aware optimal resource allocation for service-oriented networks. Future Generation Computer Systems, 101(??):959–974, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19306430.

Beloglazov:2012:EAR

[BAB12] Anton Beloglazov, Jemal Abawajy, and Rajkumar Buyya. Energy-aware resource allocation heuristics for efficient management of data centers for Cloud computing. Future Generation Computer Systems, 28(5):755–768, May 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11000689.

Breskovic:2013:CSP

[BAB13] Ivan Breskovic, Jörn Altmann, and Ivona Brandic. Creating standardized products for electronic markets. Future Generation Computer Systems, 29(4): 1000–1011, June 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001434.

Bae:2014:OBA

[Bae14] Ihn-Han Bae. An ontologybased approach to ADL recognition in smart homes. Future Generation Computer Systems, 33(??):32– 41, April 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000642.

Bagchi:2011:DAO

[Bag11] Susmit Bagchi. A distributed algorithm for ordered, atomic and simultaneous group communication. Future Generation Computer Systems, 27(5): 466–475, May 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bagchi:2016:DSP

[Bag16] Susmit Bagchi. Distributed scheduling with probabilistic and fuzzy classifications of processes. Future Generation Computer Systems, 62(??):1–16, September 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300450.

Bagchi:2019:DTA

[Bag19] Susmit Bagchi. Design and topological analysis of probabilistic distributed mutual exclusion algorithm with unbiased refined ordering. Future Generation Computer Systems, 95(??): 175–186, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18323288.

Babar:2019:UDM

[BAJ^{+}19] Muhammad Babar, Fahim Arif, Mian Ahmad Jan, Zhiyuan Tan, and Fazlullah Khan. Urban data management system: Towards big data analytics for Internet of Things based smart urban environment using customized Hadoop. Future Generation Computer Systems, 96(??):398– 409, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18321095.

Baqueri:2019:ABM

[BAKB19] Syed Fazal Abbas Baqueri, Muhammad Adnan, Bruno Kochan, and Tom Bellemans. Activity-based model for medium-sized cities considering external activity-travel: Enhancing FEATHERS framework. Future Generation Computer Systems, 96(??):51– 63, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314316.

Balis:2016:HMC

[Bal16]  Bartosz Balis.  Hyper-

Flow: a model of computation, programming approach and enactment engine for complex distributed workflows. Future Generation Computer Systems, 55(??):147–162, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002770.

Bellatreche:2017:MDEa

[BAP17a] Ladjel Bellatreche, Yamine Ait Ameur, and George Angelos Papadopoulos. Models and data engineering. Future Generation Computer Systems, 68(??):442– 444, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306392.

Bellatreche:2017:MDEb

[BAP17b] Ladjel Bellatreche, Yamine Ait Ameur, and George Angelos Papadopoulos. Models and data engineering. Future Generation Computer Systems, 70(??):1–3, May 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16304204.

Biuk-Aghai:2014:VLS

[BAPS14] Robert P. Biuk-Aghai, Cheong-Iao Pang, and Yain-Whar Si. Visualizing large-scale human collaboration in Wikipedia. Future Generation Computer Systems, 31(??):120– 133, February 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000617.

Barthes:2011:OFM

[Bar11] Jean-Paul A. Barthès. OMAS: a flexible multiagent environment for CSCWD. Future Generation Computer Systems, 27(1):78–87, January 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Baroudi:2014:DCP

[BARMB14] Uthman Baroudi, Anas Al-Roubaiey, Samir Mekid, and Abdelhafid Bouhraoua. Delay characterization and performance evaluation of cluster-based WSN with different deployment distributions. Future Generation Computer Systems, 39(??):100–110, October 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000363.

Banares:2016:ECS

[BAV16] José Ángel Bañares, Jörn Altmann, and Kurt Vanmechelen. Economics of computing services. Future Generation Computer Systems, 55(??):401–402, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003106.

Brandic:2012:SSR

[BB12] Ivona Brandic and Rajkumar Buyya. Special section: Recent advances in utility and cloud computing. Future Generation Computer Systems, 28(1): 36–38, January 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001117.

Balaji:2013:GEI

[BB13] Pavan Balaji and Rajkumar Buyya. Guest Editors’ introduction: Special issue on cluster, grid, and cloud computing. Future Generation Computer Systems, 29(8):2220–2221, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000812.

Bassem:2017:MCP

[BB17] Christine Bassem and Azer Bestavros. Multi-capacity bin packing with dependent items and its application to the packing of brokered workloads in virtualized environments. Future Generation Computer Systems, 72(??):129–144, July 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302771.

Begeman:2011:LIS

[BBB^{+}11] K. Begeman, A. N. Belikov, D. R. Boxhoorn, F. Dijkstra, H. Holties, Z. Meyer-Zhao, G. A. Renting, E. A. Valentijn, and W.-J. Vriend. LO-FAR Information System. Future Generation Computer Systems, 27(3):319– 328, March 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bochenina:2016:SSM

[BBB16] Klavdiya Bochenina, Nikolay Butakov, and Alexander Boukhanovsky. Static scheduling of multiple workflows with soft deadlines in non-dedicated heterogeneous environments. Future Generation Computer Systems, 55(??):51– 61, February 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002629.

Boukhelef:2019:OCD

[BBB^{+}19] Djillali Boukhelef, Jalil Boukhobza, Kamel Boukhalfa, Hamza Ouarnoughi, and Laurent Lemarchand. Optimizing the cost of DBaaS object placement in hybrid storage systems. Future Generation Computer Systems, 93(??):176–187, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17325785.

Bertran:2012:EAS

[BBC^{+}12] Ramon Bertran, Yolanda Becerra, David Carrera, Vicenç Beltran, Marc Gonzàlez, Xavier Martorell, Nacho Navarro, Jordi Torres, and Eduard Ayguadé. Energy accounting for shared virtualized environments under DVFS using PMCbased power models. Future Generation Computer Systems, 28(2):457–468, February 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11000355.

Bentley:2013:HDA

[BBC^{+}13] Robert Bentley, John Brooke, André Csillaghy, Donal Fellows, Anja Le Blanc, Mauro Messerotti, David Pérez-Suárez, Gabriele Pierantoni, and Marco Soldati. HELIO: Discovery and analysis of data in heliophysics. Future Generation Computer Systems, 29(8):2157–2168, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000666.

Badii:2017:AAK

[BBC^{+}17] C. Badii, P. Bellini, D. Cenni, A. Difino, P. Nesi, and M. Paolucci. Analysis and assessment of a knowledge based smart city architecture providing service APIs. Future Generation Computer Systems, 75(??):14–29, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302273.

Bellini:2018:MCS

[BBCN18] Pierfrancesco Bellini, Ivan Bruno, Daniele Cenni, and Paolo Nesi. Managing cloud via smart cloud engine and knowledge base. Future Generation Computer Systems, 78 (part 1)(??):142–154, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303867.

Bechhofer:2013:WLD

[BBD^{+}13] Sean Bechhofer, Iain Buchan, David De Roure, Paolo Missier, John Ainsworth, Jiten Bhagat, Philip Couch, Don Cruickshank, Mark Delderfield, Ian Dunlop, Matthew Gamble, Danius Michaelides, Stuart Owen, David Newman, Shoaib Sufi, and Carole Goble. Why linked data is not enough for scientists. Future Generation Computer Systems, 29(2):599–611, February 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001439.

Bagozi:2019:RBA

[BBD^{+}19] Ada Bagozi, Devis Bianchini, Valeria De Antonellis, Massimiliano Garda, and Alessandro Marini. A relevance-based approach for Big Data exploration. Future Generation Computer Systems, 101(??):51– 69, December 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18320867.

Barwell:2018:FPF

[BBH18] Adam D. Barwell, Christopher Brown, and Kevin Hammond. Finding parallel functional pearls: automatic parallel recursion scheme detection in Haskell functions via antiunification. Future Generation Computer Systems, 79 (part 2)(??):669– 686, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17315200.

Burkimsher:2013:SSM

[BBI13] Andrew Burkimsher, Iain Bate, and Leandro Soares Indrusiak. A survey of scheduling metrics and an improved ordering policy for list schedulers operating on workloads with dependencies and a wide variation in execution times. Future Generation Computer Systems, 29(8):2009– 2025, October 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12002257.

Bey:2010:MAS

[BBMG10] Kadda Beghdad Bey, Farid Benhammadi, Aicha Mokhtari, and Zahia Gessoum. Mixture of ANFIS systems for CPU load prediction in metacomputing environment. Future Generation Computer Systems, 26(7): 1003–1011, July 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

BenJrad:2019:SFD

[BBT19] Aicha Ben Jrad, Sami Bhiri, and Samir Tata. STRATFram: a framework for describing and evaluating elasticity strategies for service-based business processes in the cloud. Future Generation Computer Systems, 97(??):69–89, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18306125.

Benedyczak:2011:KAU

[BBvdB^{+}11] Krzysztof Benedyczak, Piotr Bała, Sven van den Berghe, Roger Menday, and Bernd Schuller. Key aspects of the UNICORE 6 security model. Future Generation Computer Systems, 27(2):195–201, February 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Balis:2018:HAM

[BBWB^{+}18] Bartosz Balis, Robert Brzoza-Woch, Marian Bubak, Marek Kasztelnik, Bartosz Kwolek, Piotr Nawrocki, Piotr Nowakowski, Tomasz Szydlo, and Krzysztof Zielinski. Holistic approach to management of IT infrastructure for environmental monitoring and decision support systems with urgent computing capabilities. Future Generation Computer Systems, 79 (part 1)(??):128– 143, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X16302576.

Belgacem:2015:HHC

[BC15] Mohamed Ben Belgacem and Bastien Chopard. A hybrid HPC/cloud distributed infrastructure: Coupling EC2 cloud resources with HPC clusters to run large tightly coupled multiscale applications. Future Generation Computer Systems, 42(??): 11–21, January 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001514.

Belgacem:2017:MHN

[BC17] Mohamed Ben Belgacem and Bastien Chopard. MUSCLE-HPC: a new high performance API to couple multiscale parallel applications. Future Generation Computer Systems, 67(??):72–82, February 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630259X.

Bellavista:2017:GGA

[BCC^{+}17] Paolo Bellavista, Marcello Cinque, Antonio Corradi, Luca Foschini, Flavio Frattini, and Javier Povedano-Molina. GAMESH: a grid architecture for scalable monitoring and enhanced dependable job scheduling. Future Generation Computer Systems, 71(??):192– 201, June 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16304460.

Brook:2018:LCL

[BCD^{+}18] James Brook, Felix Cuadrado, Eric Deliot, Julio Guijarro, Rycharde Hawkes, Marco Lotz, Romaric Pascal, Suksant Sae-Lor, Luis M. Vaquero, Joan Varvenne, and Lawrence Wilcock. Loom: Complex large-scale visual insight for large hybrid IT infrastructure management. Future Generation Computer Systems, 80(??): 47–62, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303843.

Borgetto:2012:EAS

[BCDP12] Damien Borgetto, Henri Casanova, Georges Da Costa, and Jean-Marc Pierson. Energy-aware service allocation. Future Generation Computer Systems, 28(5):769–779, May 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000690.

Benavent:2019:FBK

[BCdV^{+}19] Xaro Benavent, Angel Castellanos, Esther de Ves, Ana García-Serrano, and Juan Cigarrán. FCA-based knowledge representation and local generalized linear models to address relevance and diversity in diverse social images. Future Generation Computer Systems, 100(??):250–265, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18307271.

Belmonte:2010:FRM

[BCF^{+}10] Ó. Belmonte, M. Castañeda, D. Fernández, J. Gil, S. Aguado, E. Varella, M. Nuñez, and J. Segarra. Federate resource management in a Distributed Virtual Environment. Future Generation Computer Systems, 26(3):308–317, March 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Boratto:2016:DRP

[BCF16] Ludovico Boratto, Salvatore Carta, and Gianni Fenu. Discovery and representation of the preferences of automatically detected groups: Exploiting the link between group modeling and clustering. Future Generation Computer Systems, 64(??):165– 174, November 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003209.

Brinckman:2019:CER

[BCG^{+}19] Adam Brinckman, Kyle Chard, Niall Gaffney, Mihael Hategan, Matthew B. Jones, Kacper Kowalik, Sivakumar Kulasekaran, Bertram Ludäscher, Bryce D. Mecum, Jarek Nabrzyski, Victoria Stodden, Ian J. Taylor, Matthew J. Turk, and Kandace Turner. Computing environments for reproducibility: Capturing the “Whole Tale”. Future Generation Computer Systems, 94(??):854–867, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310695.

Bubendorfer:2013:ESC

[BCJT13] Kris Bubendorfer, Kyle Chard, Koshy John, and Ashfag M. Thaufeeg. eScience in the social cloud. Future Generation Computer Systems, 29(8):2143–2156, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000630.

Boselli:2018:COJ

[BCMM18] Roberto Boselli, Mirko Cesarini, Fabio Mercorio, and Mario Mezzanzanica. Classifying online job advertisements through machine learning. Future Generation Computer Systems, 86(??):319–328, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17321830.

Barra:2019:BDE

[BCN^{+}19] Silvio Barra, Aniello Castiglione, Fabio Narducci, Maria De Marsico, and Michele Nappi. Biometric data on the edge for secure, smart and user tailored access to cloud services. Future Generation Computer Systems, 101(??):534–541, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19303188.

Bouazzouni:2018:TMC

[BCP18] Mohamed Amine Bouazzouni, Emmanuel Conchon, and Fabrice Peyrard. Trusted mobile computing: an overview of existing solutions. Future Generation Computer Systems, 80(??): 596–612, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301510.

Brunner:2012:NAS

[BCR^{+}12] René Brunner, Agustín C.

Caminero, Omer F. Rana, Felix Freitag, and Leandro Navarro. Networkaware summarisation for resource discovery in P2Pcontent networks. Future Generation Computer Systems, 28(3):563–572, March 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1100032X.

Bhushan:2018:CCU

[BD18] S. N. Bharath Bhushan and Ajit Danti. Classification of compressed and uncompressed text documents. Future Generation Computer Systems, 88(??):614– 623, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17320708.

Baryannis:2019:PSC

[BDA19] George Baryannis, Samir Dani, and Grigoris Antoniou. Predicting supply chain risks using machine learning: the trade-off between performance and interpretability. Future Generation Computer Systems, 101(??):993–1004, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308003.

Bianchi:2019:ICC

[BDCC19] Giuseppe Bianchi, Tooska Dargahi, Alberto Caponi, and Mauro Conti. Intelligent conditional collaborative private data sharing. Future Generation Computer Systems, 96(??): 1–10, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1830774X.

Botta:2016:ICC

[BdDPP16] Alessio Botta, Walter de Donato, Valerio Persico, and Antonio Pescapé. Integration of cloud computing and Internet of Things: a survey. Future Generation Computer Systems, 56(??): 684–700, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003015.

Badawi:2017:MCB

[BDE17] Hawazin Faiz Badawi, Haiwei Dong, and Abdulmotaleb El Saddik. Mobile cloud-based physical activity advisory system using biofeedback sensors. Future Generation Computer Systems, 66(??):59–70, January 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003428.

Bonacin:2016:SIS

[BDF^{+}16] Rodrigo Bonacin, Nicoletta Dessì, Maria Grazia Fugini, Olga Nabuco, and Marcos Renato da Silveira. Special issue on Semantic Technologies for Collaborative Web. Future Generation Computer Systems, 54(??):344–347, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002204.

Brinckman:2019:CCD

[BDG^{+}19] Adam Brinckman, Ewa Deelman, Sandeep Gupta, Jarek Nabrzyski, Soowang Park, Rafael Ferreira da Silva, Ian J. Taylor, and Karan Vahi. Collaborative circuit designs using the CRAFT repository. Future Generation Computer Systems, 94(??): 841–853, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17310683.

Brandis:2014:TFG

[BDH14] Knud Brandis, Srdan Dzombeta, and Knut Haufe. Towards a framework for governance architecture management in cloud environments: a semantic perspective. Future Generation Computer Systems, 32(??):274–281, March 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300201X.

Barbareschi:2019:PBM

[BDL^{+}19] Mario Barbareschi, Alessandra De Benedictis, Erasmo La Montagna, Antonino Mazzeo, and Nicola Mazzocca. A PUF-based mutual authentication scheme for cloud-edges IoT systems. Future Generation Computer Systems, 101(??):246–261, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301293.

Berl:2011:ECM

[BdM11] Andreas Berl and Hermann de Meer. An energy consumption model for virtualized office environments. Future Generation Computer Systems, 27(8):1047– 1055, October 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Belkhouja:2019:BBA

[BDM^{+}19] Taha Belkhouja, Xiaojiang Du, Amr Mohamed, Abdulla K. Al-Ali, and Mohsen Guizani. Biometricbased authentication scheme for implantable medical devices during emergency situations. Future Generation Computer Systems, 98(??):109–119, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18325792.

Brock:2011:CCM

[BDMO11] Nathan Brock, Michelle Daniels, Steve Morris, and Peter Otto. A collaborative computing model for audio post-production. Future Generation Computer Systems, 27(7):935–943, July 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Benoit:2013:SLC

[BDNP13] Anne Benoit, Alexandru Dobrila, Jean-Marc Nicod, and Laurent Philippe. Scheduling linear chain streaming applications on heterogeneous systems with failures. Future Generation Computer Systems, 29(5): 1140–1151, July 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1200235X.

Balke:2011:AEI

[BDP11a] Tina Balke, Marina De Vos, and Julian Padget. Analysing energyincentivized cooperation in next generation mobile networks using normative frameworks and an agentbased simulation. Future Generation Computer Systems, 27(8):1092–1102, October 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bosin:2011:ESP

[BDP11b] Andrea Bosin, Nicoletta Dessì, and Barbara Pes. Extending the SOA paradigm to e-science environments. Future Generation Computer Systems, 27(1):20– 31, January 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bulat:2010:CTC

[BDS^{+}10] J. Bułat, K. Duda, M. Socha, P. Turcza, T. Zieliński, and M. Duplaga. Computational tasks in computerassisted transbronchial biopsy. Future Generation Computer Systems, 26 (3):455–461, March 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bent:2017:MUB

[BDWM17] Oliver Bent, Prasenjit Dey, Komminist Weldemariam, and Mukesh K. Mohania. Modeling user behavior data in systems of engagement. Future Generation Computer Systems, 68(??): 456–464, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630156X.

Berruti:2013:PEM

[BDZ13] Luca Berruti, Franco Davoli, and Sandro Zappatore. Performance evaluation of measurement data acquisition mechanisms in a distributed computing environment integrating remote laboratory instrumentation. Future Generation Computer Systems, 29(2):460–471, February 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001550.

Beltran:2016:BNA

[Bel16]  Marta Beltrán.  BE-

Cloud: a new approach to analyse elasticity enablers of cloud services. Future Generation Computer Systems, 64(??):39– 49, November 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630125X.

Bungartz:2010:PRM

[BEWZ10] H.-J. Bungartz, W. Eckhardt, T. Weinzierl, and C. Zenger. A precompiler to reduce the memory footprint of multiscale PDE solvers in C++. Future Generation Computer Systems, 26(1):175–182, January 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Baig:2018:CGN

[BFN18] Roger Baig, Felix Freitag, and Leandro Navarro. Cloudy in guifi.net: Establishing and sustaining a community cloud as open commons. Future Generation Computer Systems, 87(??):868–887, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1732856X.

Bartolini:2018:ERB

[BFP18] Sandro Bartolini, Pierfrancesco Foglia, and Cosimo Antonio Prete. Exploring the relationship between architectures and management policies in the design of NUCA-based chip multicore systems. Future Generation Computer Systems, 78 (part 2)(??):481– 501, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17311792.

Batista:2017:CQD

[BFS^{+}17a] Bruno Guazzelli Batista, Carlos Henrique Gomes Ferreira, Danilo Costa Marim Segura, Dionisio Machado Leite Filho, and Maycon Leone Maciel Peixoto. Corrigendum to “A QoS-driven approach for cloud computing addressing attributes of performance and security” [Future Gener. Comput. Syst. 68 (March) (2017) 260–274]. Future Generation Computer Systems, 73 (??):115, August 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306513.

See [BFS^{+}17b].

Batista:2017:QDA

[BFS^{+}17b] Bruno Guazzelli Batista, Carlos Henrique Gomes Ferreira, Danilo Costa Marim Segura, Dionisio Machado Leite Filho, and Maycon Leone Maciel Peixoto. A QoSdriven approach for cloud computing addressing attributes of performance and security. Future Generation Computer Systems, 68(??):260–274, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303491.

See corrigendum [BFS^{+}17a].

Badrinath:2012:PBR

[BG12] G. S. Badrinath and Phalguni Gupta. Palmprint based recognition system using phase-difference information. Future Generation Computer Systems, 28(1):287–305, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X10002499.

Barika:2019:ISM

[BGC^{+}19a] Mutaz Barika, Saurabh Garg, Andrew Chan, Rodrigo N. Calheiros, and Rajiv Ranjan. IoTSim-Stream: Modelling stream graph application in cloud simulation. Future Generation Computer Systems, 99(??):86–105, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18317047.

Blanco:2019:AIP

[BGC19b] Jesús Blanco, Andrés García, and Valentín Cañas. Analysis and improvements of the pseudorandom number generation in passive UHF-RFID tags. Future Generation Computer Systems, 99(??):115– 123, October 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18317849.

Barbierato:2014:PEN

[BGI14] Enrico Barbierato, Marco Gribaudo, and Mauro Iacono. Performance evaluation of NoSQL big-data applications using multiformalism models. Future Generation Computer Systems, 37(??):345–353, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000028.

Borges:2017:CHH

[BGMLS17] Francisco Borges, Albert Gutierrez-Milla, Emilio Luque, and Remo Suppi. Care HPS: a high performance simulation tool for parallel and distributed agent-based modeling. Future Generation Computer Systems, 68(??):59– 73, March 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302758.

Bouloukakis:2019:ASM

[BGNI19] Georgios Bouloukakis, Nikolaos Georgantas, Patient Ntumba, and Valérie Issarny. Automated synthesis of mediators for middleware-layer protocol interoperability in the IoT. Future Generation Computer Systems, 101 (??):1271–1294, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18323586.

Borcea:2017:PEE

[BGP^{+}17] Cristian Borcea, Arnab ‘Bobby” Deb Gupta, Yuriy Polyakov, Kurt Rohloff, and Gerard Ryan. PI-CADOR: End-to-end encrypted publish–subscribe information distribution with proxy re-encryption. Future Generation Computer Systems, 71(??):177– 191, June 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16303983.

Boukadi:2019:BPO

[BGRBA19] Khouloud Boukadi, Rima Grati, Molka Rekik, and Hanêne Ben-Abdallah. Business process outsourcing to cloud containers: How to find the optimal deployment? Future Generation Computer Systems, 97(??): 397–408, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18303467.

Benedetto:2019:TPF

[BGS^{+}19] José I. Benedetto, Luis A. González, Pablo Sanabria, Andrés Neyem, and Jaime Navón. Towards a practical framework for code offloading in the Internet of Things. Future Generation Computer Systems, 92(??):424–437, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18302310.

Blanke:2013:SPB

[BH13] Tobias Blanke and Mark Hedges. Scholarly primitives: Building institutional infrastructure for humanities e-Science. Future Generation Computer Systems, 29(2):654–661, February 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001178.

Bhat:2018:BDC

[Bha18] Wasim Ahmad Bhat. Bridging data-capacity gap in big data storage. Future Generation Computer Systems, 87(??):538–548, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312876.

Baskaran:2019:ENV

[BHE^{+}19] Muthu M. Baskaran, Thomas Henretty, James Ezick, Richard Lethin, and David Bruns-Smith. Enhancing network visibility and security through tensor analysis. Future Generation Computer Systems, 96(??): 207–215, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18302073.

Bai:2012:MMP

[BJ12] Yuebin Bai and Yanwen Ju. MOVE: a mobile personalized virtual computing environment. Future Generation Computer Systems, 28(6):890–899, June 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1000258X.

Beier:2017:MDS

[BJM^{+}17] Maximilian Beier, Christoph Jansen, Geert Mayer, Thomas Penzel, Andrea Rodenbeck, René Siewert, Michael Witt, Jie Wu, and Dagmar Krefting. Multicenter data sharing for collaboration in sleep medicine. Future Generation Computer Systems, 67(??):466–480, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300693.

Bali:2016:SCE

[BK16] Rasmeet S. Bali and Neeraj Kumar. Secure clustering for efficient data dissemination in vehicular cyberphysical systems. Future Generation Computer Systems, 56(??):476–492, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002836.

Babamir:2019:DDB

[BK19] Faezeh Sadat Babamir and Murvet Kirci. Dynamic digest based authentication for client-server systems using biometric verification. Future Generation Computer Systems, 101(??):112–126, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304480.

Balis:2011:RTG

[BKB11] Bartosz Balis, Bartosz Kowalewski, and Marian Bubak. Real-time Grid monitoring based on complex event processing. Future Generation Computer Systems, 27(8):1103–1112, October 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bochenina:2018:SPS

[BKB18a] Klavdiya Bochenina, Sergey Kesarev, and Alexander Boukhanovsky. Scalable parallel simulation of dynamical processes on large stochastic Kronecker graphs. Future Generation Computer Systems, 78 (part 2)(??):502–515, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301504.

Boukhanovsky:2018:UCD

[BKB18b] Alexander V. Boukhanovsky, Valeria V. Krzhizhanovskaya, and Marian Bubak. Urgent computing for decision support in critical situations. Future Generation Computer Systems, 79 (part 1)(??):111– 113, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17324895.

Byun:2011:COP

[BKKM11] Eun-Kyu Byun, Yang-Suk Kee, Jin-Soo Kim, and Seungryoul Maeng. Cost optimized provisioning of elastic resources for application workflows. Future Generation Computer Systems, 27(8):1011–1026, October 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Baars:2014:CCS

[BKS^{+}14] Thijs Baars, Ravi Khadka, Hristo Stefanov, Slinger Jansen, Ronald Batenburg, and Eugene van Heusden. Chargeback for cloud services. Future Generation Computer Systems, 41(??):91–103, December 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001502.

Basu:2018:ICM

[BKS^{+}18] Sayantani Basu, Marimuthu Karuppiah, K. Selvakumar, Kuan-Ching Li, S. K. Hafizul Islam, Mohammad Mehedi Hassan, and Md. Zakirul Alam Bhuiyan. An intelligent/ cognitive model of task scheduling for IoT applications in cloud computing environment. Future Generation Computer Systems, 88(??):254–261, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18308926.

Bhattacharyya:2018:WWL

[BKY18] Shilpi Bhattacharyya, Dimitrios Katramatos, and Shinjae Yoo. Why wait? Let us start computing while the data is still on the wire. Future Generation Computer Systems, 89(??):563–574, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18302450.

Blanc:2013:IMA

[BL13] M. Blanc and J.-F. Lalande. Improving Mandatory Access Control for HPC clusters. Future Generation Computer Systems, 29(3):876–885, March 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X12000854.

Bux:2015:DSH

[BL15] Marc Bux and Ulf Leser. DynamicCloudSim: Simulating heterogeneity in computational clouds. Future Generation Computer Systems, 46(??):85– 99, May 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001770.

Benzaid:2016:FAW

[BLAN^{+}16] Chafika Benzaid, Karim Lounis, Ameer Al-Nemrat, Nadjib Badache, and Mamoun Alazab. Fast authentication in wireless sensor networks. Future Generation Computer Systems, 55(??):362–375, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001393.

Bai:2019:LMD

[BLL^{+}19] Tong Bai, Jinzhao Lin, Guoquan Li, Huiqian Wang, Peng Ran, Zhangyong Li, Dan Li, Yu Pang, Wei Wu, and Gwanggil Jeon. A lightweight method of data encryption in BANs using electrocardiogram signal. Future Generation Computer Systems, 92(??):800–811, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17310361.

Baldassarre:2019:MPM

[BLMU19] Giorgio Baldassarre, Paolo Lo Giudice, Lorenzo Musarella, and Domenico Ursino. The MIoT paradigm: Main features and an “ad-hoc” crawler. Future Generation Computer Systems, 92(??):29–42, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17326274.

Bahnasse:2018:NSA

[BLO^{+}18] Ayoub Bahnasse, Fatima Ezzahraa Louhab, Hafsa Ait Oulahyane, Mohamed Talea, and Assia Bakali. Novel SDN architecture for smart MPLS traffic engineering-DiffServ aware management. Future Generation Computer Systems, 87(??):115–126, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17323725.

Brocco:2010:EEI

[BMH10] Amos Brocco, Apostolos Malatras, and Béat Hirsbrunner. Enabling efficient information discovery in a self-structured grid. Future Generation Computer Systems, 26(6):838–846, June 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bilal:2014:TSG

[BMK^{+}14a] Kashif Bilal, Saif Ur Rehman Malik, Osman Khalid, Abdul Hameed, Enrique Alvarez, Vidura Wijaysekara, Rizwana Irfan, Sarjan Shrestha, Debjyoti Dwivedy, Mazhar Ali, Usman Shahid Khan, Assad Abbas, Nauman Jalil, and Samee U. Khan. A taxonomy and survey on Green Data Center Networks. Future Generation Computer Systems, 36(??): 189–208, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001519.

Bourdena:2014:RIT

[BMK^{+}14b] Athina Bourdena, Constandinos X. Mavromoustakis, George Kormentzas, Evangelos Pallis, George Mastorakis, and Muneer Bani Yassein. A resource intensive traffic-aware scheme using energy-aware routing in cognitive radio networks. Future Generation Computer Systems, 39(??): 16–28, October 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000387.

Barakat:2018:ACD

[BML18] Lina Barakat, Simon Miles, and Michael Luck. Adaptive composition in dynamic service environments. Future Generation Computer Systems, 80(??): 215–228, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16307233.

Budgaga:2016:PAU

[BMP^{+}16] Walid Budgaga, Matthew Malensek, Sangmi Pallickara, Neil Harvey, F. Jay Breidt, and Shrideep Pallickara. Predictive analytics using statistical, learning, and ensemble methods to support realtime exploration of discrete event simulations. Future Generation Computer Systems, 56(??):360–374, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1500223X.

Baldoni:2015:LFP

[BMR15] Roberto Baldoni, Luca Montanari, and Marco Rizzuto. On-line failure prediction in safetycritical systems. Future Generation Computer Systems, 45(??):123–132, April 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002490.

Bendjoudi:2012:HBB

[BMT12] A. Bendjoudi, N. Melab, and E.-G. Talbi. Hierarchical branch and bound algorithm for computational grids. Future Generation Computer Systems, 28(8):1168–1176, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000544.

Banditwattanawong:2016:MPC

[BMU16] Thepparit Banditwattanawong, Masawee Masdisornchote, and Putchong Uthayopas. Multi-provider cloud computing network infrastructure optimization. Future Generation Computer Systems, 55(??):116–128, February 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002800.

Bhateja:2018:UMA

[BMU18] Vikrant Bhateja, Mukul Misra, and Shabana Urooj. Unsharp masking approaches for HVS based enhancement of mammographic masses: A comparative evaluation. Future Generation Computer Systems, 82(??):176–189, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17322112.

Bicocchi:2010:HDD

[BMZ10] Nicola Bicocchi, Marco Mamei, and Franco Zambonelli. Handling dynamics in diffusive aggregation schemes: An evaporative approach. Future Generation Computer Systems, 26(6):877–889, June 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bobek:2017:UCD

[BN17] Szymon Bobek and Grzegorz J. Nalepa. Uncertain context data management in dynamic mobile environments. Future Generation Computer Systems, 66(??):110–124, January 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630187X.

Bonacin:2016:OMI

[BNJ16] Rodrigo Bonacin, Olga Fernanda Nabuco, and Ivo Pierozzi Junior. Ontology models of the impacts of agriculture and climate changes on water resources: Scenarios on interoperability and information recovery. Future Generation Computer Systems, 54(??):423– 434, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15001028.

Bo:2019:ASC

[Bo19] Lei Bo. AOFAS scores for curative effect analysis on arthroscopic treatment of subtalar ankle instability with osteochondral injury syndrome. Future Generation Computer Systems, 91(??):506–510, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18318119.

See retraction notice [Bo20].

Bo:2020:RNA

[Bo20] Lei Bo. Retraction notice to “AOFAS scores for curative effect analysis on arthroscopic treatment of subtalar ankle instability with osteochondral injury syndrome” [future gener. comput. syst. 91 (2019) 506–510]. Future Generation Computer Systems, 107(??):1144, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306373.

See [Bo19].

Bello-Orgaz:2017:DDC

[BOHCC17] Gema Bello-Orgaz, Julio Hernandez-Castro, and David Camacho. Detecting discussion communities on vaccination in Twitter. Future Generation Computer Systems, 66(??):125– 136, January 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302175.

Berkhoff:2014:CFI

[BOP^{+}14] Christian Berkhoff, Sergio F. Ochoa, José A. Pino, Jesus Favela, Jonice Oliveira, and Luis A. Guerrero. Clairvoyance: a framework to integrate shared displays and mobile computing devices. Future Generation Computer Systems, 34(??):190–200, May 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002306.

Barker:2019:GIS

[BOWD^{+}19] Michelle Barker, Silvia Delgado Olabarriaga, Nancy Wilkins-Diehr, Sandra Gesing, Daniel S. Katz, Shayan Shahand, Scott Henwood, Tristan Glatard, Keith Jeffery, Brian Corrie, Andrew Treloar, Helen Glaves, Lesley Wyborn, Neil P. Chue Hong, and Alessandro Costa. The global impact of science gateways, virtual research environments and virtual laboratories. Future Generation Computer Systems, 95(??): 240–248, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314018.

Brooke:2010:ESC

[BP10] John M. Brooke and Michael S. Parkin. Enabling scientific collaboration on the Grid. Future Generation Computer Systems, 26(3):521–530, March 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Benerecetti:2013:TPI

[BP13] Massimo Benerecetti and Adriano Peron. Timed protocol insecurity problem is NP-complete. Future Generation Computer Systems, 29(3):843–862, March 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1100224X.

Bernabe:2014:SAM

[BPC^{+}14] Jorge Bernal Bernabe, Juan M. Marin Perez, Jose M. Alcaraz Calero, Felix J. Garcia Clemente, Gregorio Martinez Perez, and Antonio F. Gomez Skarmeta. Semanticaware multi-tenancy authorization system for cloud architectures. Future Generation Computer Systems, 32(??):154–167, March 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001070.

Buyya:2010:SSF

[BR10] Rajkumar Buyya and Rajiv Ranjan. Special section: Federated resource management in Grid and cloud computing systems. Future Generation Computer Systems, 26(8):1189–1191, October 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Belyaev:2018:COA

[BR18] Kirill Belyaev and Indrakshi Ray. Componentoriented access controlapplication servers meet tuple spaces for the masses. Future Generation Computer Systems, 86(??):726– 739, September 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167739X17309275.

Bhatia:2019:DOC

[BR19] Vandana Bhatia and Rinkle Rani. A distributed overlapping community detection model for large graphs using autoencoder. Future Generation Computer Systems, 94(??):16– 26, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17327863.

Bouaziz:2019:EMA

[BRB19a] Maha Bouaziz, Abderrezak Rachedi, and Abdelfettah Belghith. EKF-MRPL: Advanced mobility support routing protocol for Internet of Mobile Things: Movement prediction approach. Future Generation Computer Systems, 93(??):822–832, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17306805.

Bouaziz:2019:ERE

[BRB^{+}19b] Maha Bouaziz, Abderrezak Rachedi, Abdelfettah Belghith, Marion Berbineau, and Saad Al-Ahmadi. EMA-RPL: Energy and mobility aware routing for the Internet of Mobile Things. Future Generation Computer Systems, 97(??):247–258, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18302541.

Bendouda:2018:PAB

[BRH18] Djamila Bendouda, Abderrezak Rachedi, and Hafid Haffaf. Programmable architecture based on Software Defined Network for Internet of Things: Connected dominated sets approach. Future Generation Computer Systems, 80(??): 188–197, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17314590.

Byrski:2018:SSF

[BRHH18] Aleksander Byrski, Katarzyna Rycerz, John Hughes, and Kevin Hammond. Special section on functional paradigm for high performance computing. Future Generation Computer Systems, 79 (part 2)(??):643– 644, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17320836.

Bai:2019:MGO

[BRL19] Weiwei Bai, Junsheng Ren, and Tieshan Li. Modified genetic optimization-based locally weighted learning identification modeling of ship maneuvering with full scale trial. Future Generation Computer Systems, 93(??):1036–1045, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17308385.

Buchert:2015:SGP

[BRNR15] Tomasz Buchert, Cristian Ruiz, Lucas Nussbaum, and Olivier Richard. A survey of general-purpose experiment management tools for distributed systems. Future Generation Computer Systems, 45(??): 1–12, April 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002003.

Broumandnia:2019:MCM

[Bro19] Ali Broumandnia. The 3D modular chaotic map to digital color image encryption. Future Generation Computer Systems, 99(??):489–499, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19300214.

Barros:2011:CAB

[BRXdS11] Ricardo Barros, José A.

Rodrigues Nt., Geraldo B. Xexéo, and Jano M. de Souza. A collaborative approach to build evaluated Web page datasets. Future Generation Computer Systems, 27(1):119– 126, January 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Balakrishnan:2011:SEC

[BS11] P. Balakrishnan and Thamarai Selvi Somasundaram. SLA enabled CARE resource broker. Future Generation Computer Systems, 27(3):265–279, March 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Balakrishnan:2017:MMS

[BS17] Senthil Murugan Balakrishnan and Arun Kumar Sangaiah. MIFIMmiddleware solution for service centric anomaly in future Internet models. Future Generation Computer Systems, 74(??):349–365, September 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302564.

Benson:2013:PUS

[BSE^{+}13] K. Benson, S. Schlachter, T. Estrada, M. Taufer, J. Lawrence, and E. Cochran. On the powerful use of simulations in the Quake-Catcher Network to efficiently position low-cost earthquake sensors. Future Generation Computer Systems, 29(8):2128–2142, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000721.

Belyaev:2018:DAP

[BSRR18] Kirill Belyaev, Wuliang Sun, Indrakshi Ray, and Indrajit Ray. On the design and analysis of protocols for Personal Health Record storage on Personal Data Server devices. Future Generation Computer Systems, 80(??):467–482, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301388.

Boloni:2017:VIB

[BT17] Ladislau Bölöni and Damla Turgut. Value of information based scheduling of cloud computing resources. Future Generation Computer Systems, 71(??):212– 220, June 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16304472.

B:2019:SDM

[BTG19] Shijila B., Anju Jose Tom, and Sudhish N. George. Simultaneous denoising and moving object detection using low rank approximation. Future Generation Computer Systems, 90(??):198–210, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18306046.

Butrylo:2010:PSP

[BTM10] Boguslaw Butrylo, Marek Tudruj, and Lukasz Masko. Parallel SSOR preconditioning implemented on dynamic SMP clusters with communication on the fly. Future Generation Computer Systems, 26(3):491– 497, March 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bhattacharjee:2019:IMS

[BTP19] Sreyasee Das Bhattacharjee, William J. Tolone, and Ved Suhas Paranjape. Identifying malicious social media contents using multi-view contextaware active learning. Future Generation Computer Systems, 100(??):365–379, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18307349.

Bu:2018:EFM

[Bu18]  Fanyu Bu.  An effi-

cient fuzzy c-means approach based on canonical polyadic decomposition for clustering big data in IoT. Future Generation Computer Systems, 88(??):675– 682, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18306472.

Basanta-Val:2015:IPD

[BVFGWA15] P. Basanta-Val, N. Fernández-García, A. J. Wellings, and N. C. Audsley. Improving the predictability of distributed stream processors. Future Generation Computer Systems, 52(??):22– 36, November 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000850.

Bhatnagar:2013:BIW

[BW13] Gaurav Bhatnagar and Q. M. Jonathan Wu. Biometrics inspired watermarking based on a fractional dual tree complex wavelet transform. Future Generation Computer Systems, 29(1):182–195, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1200129X.

Bu:2019:SAI

[BW19] Fanyu Bu and Xin Wang. A smart agriculture IoT system based on deep reinforcement learning. Future Generation Computer Systems, 99(??):500–507, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19307277.

Bu:2019:MPD

[BWG19] Fanyu Bu, Xin Wang, and Bo Gao. A multiprojection deep computation model for smart data in Internet of Things. Future Generation Computer Systems, 93(??):68– 76, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18319393.

Bhatnagar:2012:FDT

[BWR12] Gaurav Bhatnagar, Jonathan Wu, and Balasubramanian Raman. Fractional dual tree complex wavelet transform and its application to biometric security during communication and transmission. Future Generation Computer Systems, 28(1):254–267, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X10002323.

Bei:2018:CMC

[BYL^{+}18] Zhendong Bei, Zhibin Yu, Ni Luo, Chuntao Jiang, Chengzhong Xu, and Shengzhong Feng. Configuring in-memory cluster computing using random forest. Future Generation Computer Systems, 79 (part 1)(??):1– 15, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X16305167.

Bello:2019:TES

[BZ19] Oladayo Bello and Sherali Zeadally. Toward efficient smartification of the Internet of Things (IoT) services. Future Generation Computer Systems, 92(??):663–673, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17317326.

Bartos:2019:NEC

[BZHV19] Vaclav Bartos, Martin Zadnik, Sheikh Mahbub Habib, and Emmanouil Vasilomanolakis. Network entity characterization and attack prediction. Future Generation Computer Systems, 97(??):674–686, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18307799.

Bagheri:2010:AIA

[BZMY10] A. Bagheri, M. Zandieh, Iraj Mahdavi, and M. Yazdani. An artificial immune algorithm for the flexible job-shop scheduling problem. Future Generation Computer Systems, 26(4):533–541, April 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bessis:2018:SOS

[BZS18] Nik Bessis, Xiaojun Zhai, and Stelios Sotiriadis. Service-oriented system engineering. Future Generation Computer Systems, 80(??):211–214, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17326122.

Celaya:2013:TRA

[CA13] Javier Celaya and Unai Arronategui. A task routing approach to largescale scheduling. Future Generation Computer Systems, 29(5):1097–1111, July 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12002294.

Calero:2015:CAA

[CA15a] Jose M. Alcaraz Calero and Juan Gutiérrez Aguado. Comparative analysis of architectures for monitoring cloud computing infrastructures. Future Generation Computer Systems, 47 (??):16–30, June 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002684.

Celaya:2015:FSB

[CA15b] Javier Celaya and Unai Arronategui. Fair scheduling of bag-of-tasks applications on large-scale platforms. Future Generation Computer Systems, 49(??): 28–44, August 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000552.

Cioara:2018:OFM

[CAB^{+}18] Tudor Cioara, Ionut Anghel, Massimo Bertoncini, Ioan Salomie, Diego Arnone, Marzia Mammina, Terpsichori-Helen Velivassaki, and Marcel Antal. Optimized flexibility management enacting data centres participation in smart demand response programs. Future Generation Computer Systems, 78 (part 1)(??):330– 342, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301200.

Campos:2010:ACM

[CAC^{+}10] Ricardo Silva Campos, Ronan Mendonca Amorim, Caroline Mendonça Costa, Bernardo Lino de Oliveira, Ciro de Barros Barbosa, Joakim Sundnes, and Rodrigo Weber dos Santos. Approaching cardiac modeling challenges to computer science with CellMLbased Web tools. Future Generation Computer Systems, 26(3):462–470, March 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chen:2015:HMC

[CAC^{+}15] Jen-Hsiang Chen, Fahmida Abedin, Kuo-Ming Chao, Nick Godwin, Yinsheng Li, and Chen-Fang Tsai. A hybrid model for cloud providers and consumers to agree on QoS of cloud services. Future Generation Computer Systems, 50(??):38– 48, September 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002568.

Chen:2018:NFD

[CAL^{+}18] Weihong Chen, Jiyao An, Renfa Li, Li Fu, Guoqi Xie, Md Zakirul Alam Bhuiyan, and Keqin Li. A novel fuzzy deep-learning approach to traffic flow prediction with uncertain spatial-temporal data features. Future Generation Computer Systems, 89(??):78–88, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18307398.

Chen:2018:EMI

[CAPG18] Bo-Wei Chen, Nik Nailah Binti Abdullah, Sangoh Park, and Y. Gu. Efficient multiple incremental computation for kernel ridge regression with Bayesian uncertainty modeling. Future Generation Computer Systems, 82(??):679–688, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17308117.

Cardoso:2016:SSW

[CAS^{+}16] Silvio D. Cardoso, Flor K. Amanqui, Kleberson J. A. Serique, José L. C. dos Santos, and Dilvan A. Moreira. SWI: a Semantic Web Interactive Gazetteer to support Linked Open Data. Future Generation Computer Systems, 54(??):389– 398, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15001818.

Cioara:2018:ESN

[CAS^{+}18] Tudor Cioara, Ionut Anghel, Ioan Salomie, Lina Barakat, Simon Miles, Dianne Reidlinger, Adel Taweel, Ciprian Dobre, and Florin Pop. Expert system for nutrition care process of older adults. Future Generation Computer Systems, 80(??): 368–383, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17311056.

Choi:2010:GBA

[CB10] SungJin Choi and Rajkumar Buyya. Group-based adaptive result certification mechanism in Desktop Grids. Future Generation Computer Systems, 26(5):776–786, May 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cohen-Boulakia:2017:SWC

[CBBC^{+}17]  Sarah  Cohen-Boulakia,

Khalid Belhajjame, Olivier Collin, Jérôme Chopard, Christine Froidevaux, Alban Gaignard, Konrad Hinsen, Pierre Larmande, Yvan Le Bras, Frédéric Lemoine, Fabien Mareuil, Hervé Ménager, Christophe Pradal, and Christophe Blanchet. Scientific workflows for computational reproducibility in the life sciences: Status, challenges and opportunities. Future Generation Computer Systems, 75(??):284–298, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300316.

Cushing:2016:TDP

[CBBdL16] Reginald Cushing, Adam Belloum, Marian Bubak, and Cees de Laat. Towards a data processing plane: an automatabased distributed dynamic data processing model. Future Generation Computer Systems, 59(??):21– 32, June 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003647.

Chaaya:2019:CAS

[CBC^{+}19] Karam Bou Chaaya, Mahmoud Barhamgi, Richard Chbeir, Philippe Arnould, and Djamal Benslimane. Context-aware system for dynamic privacy risk inference: Application to smart IoT environments. Future Generation Computer Systems, 101(??):1096–1111, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19301311.

Cavdar:2015:SFP

[ÇBCA15] Derya Çavdar, Robert Birke, Lydia Y. Chen, and Fatih Alagöz. A simulation framework for priority scheduling on heterogeneous clusters. Future Generation Computer Systems, 52(??):37– 48, November 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15001004.

Chirkin:2017:ETE

[CBK^{+}17] Artem M. Chirkin, Adam S. Z. Belloum, Sergey V. Kovalchuk, Marc X. Makkes, Mikhail A. Melnik, Alexander A. Visheratin, and Denis A. Nasonov. Execution time estimation for workflow scheduling. Future Generation Computer Systems, 75(??):376–387, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300304.

Costa:2013:AIE

[CBLS13] Rostand Costa, Francisco Brasileiro, Guido Lemos, and Dênio Sousa. Analyzing the impact of elasticity on the profit of cloud computing providers. Future Generation Computer Systems, 29(7):1777–1785, September 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000058.

Chard:2016:NHS

[CBN16] Ryan Chard, Kris Bubendorfer, and Bryan Ng. Network health and e-science in commercial clouds. Future Generation Computer Systems, 56(??):595–604, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002022.

Chifor:2018:SAS

[CBPP18] Bogdan-Cosmin Chifor, Ion Bica, Victor-Valeriu Patriciu, and Florin Pop. A security authorization scheme for smart home Internet of Things devices. Future Generation Computer Systems, 86(??):740–749, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17311020.

Chauhan:2017:RAP

[CBS17] Muhammad Aufeef Chauhan, Muhammad Ali Babar, and Quan Z. Sheng. A reference architecture for provisioning of tools as a service: Meta-model, ontologies and design elements. Future Generation Computer Systems, 69(??):41– 65, April 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16307221.

Chervyakov:2019:ARC

[CBT^{+}19] Nikolay Chervyakov, Mikhail Babenko, Andrei Tchernykh, Nikolay Kucherov, Vanessa Miranda-López, and Jorge M. Cortés-Mendoza. AR-RRNS: Configurable reliable distributed data storage systems for Internet of Things to ensure security. Future Generation Computer Systems, 92(??):1080–1092, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17306015.

Chang:2011:SSG

[CC11] Ruay-Shiung Chang and Han-Chieh Chao. Special section: Grid and pervasive computing (selected papers from 2010 International Conference on Grid and Pervasive Computing). Future Generation Computer Systems, 27(6):820– 822, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chen:2019:MPM

[CC19] Mu-Yen Chen and Ting-Hsuan Chen. Modeling public mood and emotion: Blog and news sentiment and socio-economic phenomena. Future Generation Computer Systems, 96(??):692–699, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17323750.

Cruz:2019:ASP

[CCC19] Pedro Cruz, Rodrigo S. Couto, and Luís Henrique M. K. Costa. An algorithm for sink positioning in busassisted smart city sensing. Future Generation Computer Systems, 93(??):761– 769, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17304247.

Conejero:2014:VTC

[CCCT14] Javier Conejero, Blanca Caminero, Carmen Carrión, and Luis Tomás. From volunteer to trustable computing: Providing QoSaware scheduling mechanisms for multi-grid computing environments. Future Generation Computer Systems, 34(??):76– 93, May 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002690.

Clematis:2010:JRM

[CCD^{+}10] A. Clematis, A. Corana, D. D’Agostino, A. Galizia, and A. Quarati. Jobresource matchmaking on Grid through two-level benchmarking. Future Generation Computer Systems, 26(8):1165–1179, October 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Carpentieri:2019:OPL

[CCD^{+}19] Bruno Carpentieri, Arcangelo Castiglione, Alfredo De Santis, Francesco Palmieri, and Raffaele Pizzolante. One-pass lossless data hiding and compression of remote sensing data. Future Generation Computer Systems, 90(??):222– 239, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18306629.

Cinque:2019:FLT

[CCDP19] Marcello Cinque, Domenico Cotroneo, Raffaele Della Corte, and Antonio Pecchia. A framework for on-line timing error detection in software systems. Future Generation Computer Systems, 90(??):521– 538, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18309609.

Casalicchio:2018:RCL

[CCIP18] Emiliano Casalicchio, Valeria Cardellini, Gianluca Interino, and Monica Palmirani. Research challenges in legal-rule and QoSaware cloud service brokerage. Future Generation Computer Systems, 78 (part 1)(??):211–223, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306641.

Chen:2016:TFR

[CCJ16] Hanhua Chen, Xiaolong Cui, and Hai Jin. Topk followee recommendation over microblogging systems by exploiting diverse information sources. Future Generation Computer Systems, 55(??):534–543, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1400096X.

Charr:2011:JVF

[CCL11] Jean-Claude Charr, Raphaël Couturier, and David Laiymani. JACEP2P-V2: a fully decentralized and fault tolerant environment for executing parallel iterative asynchronous applications on volatile distributed architectures. Future Generation Computer Systems, 27(5):606–613, May 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cabana:2019:MMM

[CCL19] Antoine Cabana, Christophe Charrier, and Alain Louis. Mono and multi-modal biometric systems assessment by a common black box testing framework. Future Generation Computer Systems, 101(??):293–303, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1833111X.

Cinquini:2014:ESG

[CCM^{+}14] Luca Cinquini, Daniel Crichton, Chris Mattmann, John Harney, Galen Shipman, Feiyi Wang, Rachana Ananthakrishnan, Neill Miller, Sebastian Denvil, Mark Morgan, Zed Pobre, Gavin M. Bell, Charles Doutriaux, Robert Drach, Dean Williams, Philip Kershaw, Stephen Pascoe, Estanislao Gonzalez, Sandro Fiore, and Roland Schweitzer. The Earth System Grid Federation: an open infrastructure for access to distributed geospatial data. Future Generation Computer Systems, 36(??):400– 417, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001477.

Camacho:2018:BAC

[CCMGF18] David Camacho, Carlos Cotta, J. J. Merelo-Guervós, and Francisco Fernández. Bioinspired algorithms in complex ephemeral environments. Future Generation Computer Systems, 88(??):732– 734, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18317680.

Castiglione:2018:CBD

[CCMP18] Aniello Castiglione, Francesco Colace, Vincenzo Moscato, and Francesco Palmieri. CHIS: A big data infrastructure to manage digital cultural items. Future Generation Computer Systems, 86(??):1134–1145, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17305605.

Chang:2018:ITB

[CCRL18] Victor Chang, Dickson K. W. Chiu, Muthu Ramachandran, and Chung-Sheng Li. Internet of Things, Big Data and Complex Information Systems: Challenges, solutions and outputs from IoTBD 2016, COMPLEXIS 2016 and CLOSER 2016 selected papers and CLOSER 2015 keynote. Future Generation Computer Systems, 79 (part 3)(??):973–974, February 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17319660.

Casola:2013:CAS

[CCRV13] Valentina Casola, Antonio Cuomo, Massimiliano Rak, and Umberto Villano. The CloudGrid approach: Security analysis and performance evaluation. Future Generation Computer Systems, 29(1):387–401, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001476.

Chen:2010:DBF

[CCS^{+}10] Po-Cheng Chen, Jyh-Biau Chang, Ce-Kuen Shieh, Chia-Han Lin, and Yi-Chang Zhuang. A DSMbased fragmented data sharing framework for grids. Future Generation Computer Systems, 26(4): 668–677, April 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Caron:2013:WSS

[CCT13] Eddy Caron, Florent Chuffart, and Cédric Tedeschi. When self-stabilization meets real platforms: an experimental study of a peer-to-peer service discovery system. Future Generation Computer Systems, 29(6):1533–1543, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001860.

Chen:2018:INM

[CCY^{+}18] Chong Chen, Dan Chen, Yingnan Yan, Gaofeng Zhang, Qingguo Zhou, and Rui Zhou. Integration of numerical model and cloud computing. Future Generation Computer Systems, 79 (part 1)(??):396– 407, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312347.

Chen:2019:VLQ

[CCZ^{+}19] Wei Chen, Jincai Chen, Fuhao Zou, Yuan-Fang Li, Ping Lu, Qiang Wang, and Wei Zhao. Vector and line quantization for billion-scale similarity search on GPUs. Future Generation Computer Systems, 99(??):295–307, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1833084X.

Canard:2016:HPP

[CD16] S. Canard and J. Devigne.

Highly privacy-protecting data sharing in a tree structure. Future Generation Computer Systems, 62(??):119–127, September 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300036.

Chung:2019:NNS

[CDB^{+}19] Joaquin Chung, Sean Donovan, Jeronimo Bezerra, Heidi Morgan, Julio Ibarra, Russ Clark, and Henry Owen. Novel network services for supporting big data science research. Future Generation Computer Systems, 98(??):512–521, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314651.

Castiglione:2017:SGC

[CDDR17] Arcangelo Castiglione, Paolo D’Arco, Alfredo De Santis, and Rosario Russo. Secure group communication schemes for dynamic heterogeneous distributed computing. Future Generation Computer Systems, 74(??):313–324, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003830.

Conti:2018:ITS

[CDFW18] Mauro Conti, Ali Dehghantanha, Katrin Franke, and Steve Watson. Internet of Things security and forensics: Challenges and opportunities. Future Generation Computer Systems, 78 (part 2)(??):544–546, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17316667.

Choo:2016:CCT

[CDFZ16] Kim-Kwang Raymond Choo, Josep Domingo-Ferrer, and Lei Zhang. Cloud cryptography: Theory, practice and future research directions. Future Generation Computer Systems, 62(??):51–53, September 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300942.

Campa:2014:PPH

[CDG^{+}14] Sonia Campa, Marco Danelutto, Mehdi Goli, Horacio González-Vélez, Alina Madalina Popescu, and Massimo Torquati. Parallel patterns for heterogeneous CPU/GPU architectures: Structured parallelism from cluster to cloud. Future Generation Computer Systems, 37(??): 354–366, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000041.

Curry:2019:RTL

[CDH^{+}19] Edward Curry, Wassim Derguech, Souleiman Hasan, Christos Kouroupetroglou, and Umair ul Hassan. A real-time linked dataspace for the Internet of Things: Enabling “Pay-As-You-Go” data management in smart environments. Future Generation Computer Systems, 90(??):405–422, January 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1732887X.

Carniani:2016:UCC

[CDL^{+}16] Enrico Carniani, Davide D’Arenzo, Aliaksandr Lazouski, Fabio Martinelli, and Paolo Mori. Usage control on cloud systems. Future Generation Computer Systems, 63(??):37– 55, October 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300875.

Cui:2018:ABC

[CDL18] Hui Cui, Robert H. Deng, and Yingjiu Li. Attributebased cloud storage with secure provenance over encrypted data. Future Generation Computer Systems, 79 (part 2)(??):461– 472, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17322835.

Coppolino:2019:CAE

[CDMR19] Luigi Coppolino, Salvatore D’Antonio, Giovanni Mazzeo, and Luigi Romano. A comparative analysis of emerging approaches for securing Java software with Intel SGX. Future Generation Computer Systems, 97(??):620– 633, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18315942.

Costa:2019:NPC

[CdRRdCB19] Felipe Rabuske Costa, Rodrigo da Rosa Righi, Cristiano André da Costa, and Cristiano Bonato Both. Nuoxus: A proactive caching model to manage multimedia content distribution on fog radio access networks. Future Generation Computer Systems, 93(??):143–155, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18302449.

Chen:2015:UIM

[CdSDS15] Weiwei Chen, Rafael Ferreira da Silva, Ewa Deelman, and Rizos Sakellariou. Using imbalance metrics to optimize task clustering in scientific workflow executions. Future Generation Computer Systems, 46 (??):69–84, May 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001848.

Cafaro:2019:CSB

[CEP19a] Massimo Cafaro, Italo Epicoco, and Marco Pulimeno. CMSS: Sketching based reliable tracking of large network flows. Future Generation Computer Systems, 101(??):770–784, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1930490X.

Cafaro:2019:MFI

[CEP19b] Massimo Cafaro, Italo Epicoco, and Marco Pulimeno. Mining frequent items in unstructured P2P networks. Future Generation Computer Systems, 95 (??):1–16, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18315838.

Chemodanov:2019:AAA

[CES^{+}19] Dmitrii Chemodanov, Flavio Esposito, Andrei Sukhov, Prasad Calyam, Huy Trinh, and Zakariya Oraibi. AGRA: AI-augmented geographic routing approach for IoT-based incidentsupporting applications. Future Generation Computer Systems, 92(??): 1051–1065, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        https://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X17303965.

Corradi:2014:VCR

[CFF14] Antonio Corradi, Mario Fanelli, and Luca Foschini.

VM consolidation: a real case based on OpenStack Cloud. Future Generation Computer Systems, 32(??): 118–127, March 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001082.

Celesti:2019:ASM

[CFG^{+}19] Antonio Celesti, Maria Fazio, Antonino Galletta, Lorenzo Carnevale, Jiafu Wan, and Massimo Villari. An approach for the secure management of hybrid cloud-edge environments. Future Generation Computer Systems, 90(??): 1–19, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18300682.

Castano:2016:CCC

[CFGM16] Silvana Castano, Alfio Ferrara, Lorenzo Genta, and Stefano Montanelli. Combining crowd consensus and user trustworthiness for managing collective tasks. Future Generation Computer Systems, 54(??):378–388, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15001065.

Chen:2019:EBP

[CFH^{+}19] Xiaodao Chen, Junqing Fan, Qing He, Yuewei Wang, Dongbo Liu, and Shiyan Hu. Economical and balanced production in smart petroleum cyber-physical system. Future Generation Computer Systems, 95(??):364–371, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17312062.

Chang:2015:MCB

[CFL^{+}15] Yue-Shan Chang, Chih-Tien Fan, Win-Tsung Lo, Wan-Chun Hung, and Shyan-Ming Yuan. Mobile cloud-based depression diagnosis using an ontology and a Bayesian network. Future Generation Computer Systems, 43–44(??): 87–98, February 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1400137X.

Cheng:2018:CIS

[CFL^{+}18] Baolei Cheng, Jianxi Fan, Qiang Lyu, Jingya Zhou, and Zhao Liu. Constructing independent spanning trees with height n on the n-dimensional crossed cube. Future Generation Computer Systems, 87(??):404–415, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18302723.

Castano:2017:EAT

[CFM17] Silvana Castano, Alfio Ferrara, and Stefano Montanelli. Exploratory analysis of textual data streams. Future Generation Computer Systems, 68(??):391– 406, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302357.

Castano:2019:LCS

[CFM19] Silvana Castano, Alfio Ferrara, and Stefano Montanelli. Leveraging crowd skills and consensus for collaborative web-resource labeling. Future Generation Computer Systems, 95(??): 790–801, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17307392.

Cecchinel:2019:LLM

[CFMC19] Cyril Cecchinel, François Fouquet, Sébastien Mosser, and Philippe Collet. Leveraging live machine learning and deep sleep to support a self-adaptive efficient configuration of battery powered sensors. Future Generation Computer Systems, 92(??):225–240, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18305740.

Casadei:2019:MSO

[CFP^{+}19] Roberto Casadei, Giancarlo Fortino, Danilo Pianini, Wilma Russo, Claudio Savaglio, and Mirko Viroli. Modelling and simulation of opportunistic IoT services with aggregate computing. Future Generation Computer Systems, 91(??):252–262, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18307246.

Cho:2017:SFU

[CFPC17] Seoungjae Cho, Simon Fong, Yong Woon Park, and Kyungeun Cho. Simulation framework of ubiquitous network environments for designing diverse network robots. Future Generation Computer Systems, 76(??):468–473, November 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300607.

Celesti:2012:VMP

[CFVP12] Antonio Celesti, Maria Fazio, Massimo Villari, and Antonio Puliafito. Virtual machine provisioning through satellite communications in federated Cloud environments. Future Generation Computer Systems, 28(1):85–93, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001038.

Carretero:2018:NDM

[CGBAP18] Jesus Carretero, Javier Garcia-Blas, Gabriel Antoniu, and Dana Petcu. New directions in mobile, hybrid, and heterogeneous clouds for cyberinfrastructures. Future Generation Computer Systems, 87(??):615–617, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18313049.

Calderon:2012:EVC

[CGCB^{+}12] Alejandro Calderón, Felix García-Carballeira, Borja Bergua, Luis Miguel Sánchez, and Jesús Carretero. Expanding the volunteer computing scenario: a novel approach to use parallel applications on volunteer computing. Future Generation Computer Systems, 28(6):881–889, June 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000550.

Clarke:2010:SSS

[CGD10] P. E. L. Clarke, C. Greenwood, and A. C. Davenhall. Special section: Switched lightpaths. Future Generation Computer Systems, 26(1):97–98, January 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Castiglione:2014:EMF

[CGIP14] Aniello Castiglione, Marco Gribaudo, Mauro Iacono, and Francesco Palmieri. Exploiting mean field analysis to model performances of big data architectures. Future Generation Computer Systems, 37(??):203– 211, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001611.

Coveney:2010:LSC

[CGJ^{+}10] P. V. Coveney, G. Giupponi, S. Jha, S. Manos, J. MacLaren, S. M. Pickles, R. S. Saksena, T. Soddemann, J. L. Suter, M. Thyveetil, and S. J. Zasada. Large scale computational science on federated international grids: The role of switched optical networks. Future Generation Computer Systems, 26 (1):99–110, January 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cotelo:2010:RGG

[CGL^{+}10] Carmen Cotelo, Andrés Gómez, J. Ignacio López, David Mera, José M. Cotos, J. Pérez Marrero, and Constantino Vázquez. Retelab: a geospatial Grid Web laboratory for the oceanographic research community. Future Generation Computer Systems, 26(8):1157–1164, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Church:2015:EHS

[CGL15] Philip Church, Andrzej Goscinski, and Christophe Lefèvre. Exposing HPC and sequential applications as services through the development and deployment of a SaaS cloud. Future Generation Computer Systems, 43–44(??):24–37, February 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1400185X.

Corbellini:2018:DND

[CGM^{+}18] Alejandro Corbellini, Daniela Godoy, Cristian Mateos, Silvia Schiaffino, and Alejandro Zunino. DPM: a novel distributed largescale social graph processing framework for link prediction algorithms. Future Generation Computer Systems, 78 (part 1)(??):474– 480, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302352.

Cicirelli:2019:IMA

[CGM^{+}19] Franco Cicirelli, Antonio Guerrieri, Alessandro Mercuri, Giandomenico Spezzano, and Andrea Vinci. ITEMa: A methodological approach for cognitive edge computing IoT ecosystems. Future Generation Computer Systems, 92(??):189– 197, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17330224.

Caggianese:2018:ESI

[CGN18] Giuseppe Caggianese, Luigi Gallo, and Pietro Neroni. Evaluation of spatial interaction techniques for virtual heritage applications: a case study of an interactive holographic projection. Future Generation Computer Systems, 81(??): 516–527, April 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17316047.

Choudhary:2018:GBH

[CGSJ18] Anubhav Choudhary, Indrajeet Gupta, Vishakha Singh, and Prasanta K. Jana. A GSA based hybrid algorithm for biobjective workflow scheduling in cloud computing. Future Generation Computer Systems, 83(??):14– 26, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17303217.

Cicirelli:2017:EBP

[CGSV17] Franco Cicirelli, Antonio Guerrieri, Giandomenico Spezzano, and Andrea Vinci. An edge-based platform for dynamic Smart City applications. Future Generation Computer Systems, 76(??):106–118, November 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16308342.

Cannataro:2010:IDP

[CGV10] Mario Cannataro, Pietro H. Guzzi, and Pierangelo Veltri. IMPRECO: Distributed prediction of protein complexes. Future Generation Computer Systems, 26(3):434–440, March 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chen:2019:MLG

[CGZL19] Xu Chen, Judith Gelernter, Han Zhang, and Jin Liu. Multi-lingual geoparsing based on machine translation. Future Generation Computer Systems, 96(??):667–677, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17316217.

Chang:2010:RDT

[CH10] Ruay-Shiung Chang and Min-Shuo Hu. A resource discovery tree using bitmap for grids. Future Generation Computer Systems, 26(1):29–37, January 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chang:2011:SAF

[Cha11] Yang-Lang Chang. A simulated annealing feature extraction approach for hyperspectral images. Future Generation Computer Systems, 27(4):419–426, April 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chang:2014:BIS

[Cha14a] Victor Chang. The business intelligence as a service in the cloud. Future Generation Computer Systems, 37(??):512–534, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002926.

See corrigendum [Cha14b].

Chang:2014:CBI

[Cha14b]  Victor Chang.  Corri-

gendum to “The business intelligence as a service in the cloud” [Future Gener. Comput. Syst. 37C (2014) 512–534]. Future Generation Computer Systems, 41(??):16, December 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/S0167739X14001496.

See [Cha14a].

Chandra:2015:BAN

[Cha15] Deka Ganesh Chandra. BASE analysis of NoSQL database. Future Generation Computer Systems, 52(??):13–21, November 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001788.

Chang:2017:PAO

[CHC^{+}17] Cheng Chang, Ligang He, Nadeem Chaudhary, Songling Fu, Hao Chen, Jianhua Sun, Kenli Li, Zhangjie Fu, and Ming-Liang Xu. Performance analysis and optimization for workflow authorization. Future Generation Computer Systems, 67(??):194– 205, February 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303089.

Chen:2013:CQS

[Che13a] Chin-Ling Chen. Combining quality of services path first routing and admission control to support VoIP traffic. Future Generation Computer Systems, 29(7):1742–1750, September 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000982.

Chen:2013:UAS

[Che13b] Lanxiang Chen. Using algebraic signatures to check data possession in cloud storage. Future Generation Computer Systems, 29(7):1709–1715, September 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000167.

Chen:2014:HOF

[Che14] Mu-Yen Chen. A highorder fuzzy time series forecasting model for Internet stock trading. Future Generation Computer Systems, 37(??):461–467, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002045.

Chen:2018:TCA

[Che18] Rui-Yang Chen. A traceability chain algorithm for artificial neural networks using T-S fuzzy cognitive maps in blockchain. Future Generation Computer Systems, 80(??):198–210, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1730064X.

Casey:2010:VAT

[CHJS^{+}10] Simon Casey, Richard Hughes-Jones, Ralph Spencer, Matthew Strong, Paul Burgess, Arpad Szomoru, and Colin Greenwood. VLBI UDP: An application for transporting VLBI data using the UDP protocol. Future Generation Computer Systems, 26(1): 120–127, January 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chang:2011:SSF

[ChK11] Ruay-Shiung Chang and Tai hoon Kim. Special section: Future generation information technology. Future Generation Computer Systems, 27(4): 370–371, April 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chen:2011:SEI

[CHS11] Tien-Ho Chen, Han-Cheng Hsiang, and Wei-Kuan Shih. Security enhancement on an improvement on two remote user authentication schemes using smart cards. Future Generation Computer Systems, 27(4):377–380, April 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chen:2018:CHT

[CHS^{+}18] Zhiwen Chen, Xin He, Jianhua Sun, Hao Chen, and Ligang He. Concurrent hash tables on multicore machines: Comparison, evaluation and implications. Future Generation Computer Systems, 82(??):127–141, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17317715.

Castro:2018:ADC

[CHSA18] David Castro, Kevin Hammond, Susmit Sarkar, and Yasir Alguwaifli. Automatically deriving cost models for structured parallel processes using hylomorphisms. Future Generation Computer Systems, 79 (part 2)(??):653– 668, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17307288.

Cala:2013:CCF

[CHWW13] Jacek Cala, Hugo Hiden, Simon Woodman, and Paul Watson. Cloud computing for fast prediction of chemical activity. Future Generation Computer Systems, 29(7):1860–1869, September 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000253.

Cui:2018:NCA

[CHY^{+}18] Laizhong Cui, Wenyuan Huang, Qiao Yan, F. Richard Yu, Zhenkun Wen, and Nan Lu. A novel context-aware recommendation algorithm with twolevel SVD in social networks. Future Generation Computer Systems, 86(??):1459–1470, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17314905.

Chadwick:2010:CMA

[CIK10] David W. Chadwick, George Inman, and Nate Klingenstein. A conceptual model for attribute aggregation. Future Generation Computer Systems, 26(7):1043– 1052, July 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ciuffoletti:2010:STP

[Ciu10a] Augusto Ciuffoletti. Secure token passing at application level. Future Generation Computer Systems, 26 (7):1026–1031, July 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ciuffoletti:2010:WTC

[Ciu10b] Augusto Ciuffoletti. The wandering token: Congestion avoidance of a shared resource. Future Generation Computer Systems, 26 (3):473–478, March 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chang:2014:ESR

[CJ14] Jau-Yang Chang and Pei-Hao Ju. An energysaving routing architecture with a uniform clustering algorithm for wireless body sensor networks. Future Generation Computer Systems, 35(??):128–140, June 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300191X.

Chen:2018:ARH

[CJG^{+}18] YanKun Chen, Fei Ji, Quansheng Guan, Yide Wang, Fangjiong Chen, and Hua Yu. Adaptive RTO for handshakingbased MAC protocols in underwater acoustic networks. Future Generation Computer Systems, 86(??):1185–1192, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17318356.

Chen:2013:SOV

[CJHH13] Huacai Chen, Hai Jin, Kan Hu, and Jian Huang. Scheduling overcommitted VM: Behavior monitoring and dynamic switchingfrequency scaling. Future Generation Computer Systems, 29(1):341–351, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001452.

Chung:2018:ARA

[CJK^{+}18] Joaquin Chung, Eun-Sung Jung, Rajkumar Kettimuthu, Nageswara S. V. Rao, Ian T. Foster, Russ Clark, and Henry Owen. Advance reservation access control using software-defined networking and tokens. Future Generation Computer Systems, 79 (part 1)(??):225– 234, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17303606.

Chao:2015:CLM

[CJN^{+}15] Kuo-Ming Chao, Anne E.

James, Antonios G. Nanos, Jen-Hsiang Chen, Sergiu-Dan Stan, Ionut Muntean, Giorgio Figliolini, Pierluigi Rea, Chedli B. Bouzgarrou, Pavel Vitliemov, Joshua Cooper, and Jürgen van Capelle. Cloud Elearning for Mechatronics: CLEM. Future Generation Computer Systems, 48 (??):ibc, July 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002271.

Cadenelli:2019:CUO

[CJPC19] Nicola Cadenelli, Zoran Jakšić, Jordà Polo, and David Carrera. Considerations in using OpenCL on GPUs and FPGAs for throughput-oriented genomics workloads. Future Generation Computer Systems, 94(??):148–159, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314183.

Cheon:2019:CRO

[CJS19] Jung Hee Cheon, Jin-Hyuck Jeong, and Ji Sun Shin. Cryptoanalysis on ‘A round-optimal latticebased blind signature scheme for cloud services’. Future Generation Computer Systems, 95(??):100– 103, June 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317667.

See [ZaTZ^{+}17].

Chang:2019:GTS

[CJXX19] Jinyong Chang, Yanyan Ji, Maozhi Xu, and Rui Xue. General transformations from single-generation to multi-generation for homomorphic message authentication schemes in network coding. Future Generation Computer Systems, 91(??):416–425, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17330170.

Chung:2019:OIA

[CKP^{+}19] Joaquin Chung, Rajkumar Kettimuthu, Nam Pho, Russ Clark, and Henry Owen. Orchestrating intercontinental advance reservations with softwaredefined exchanges. Future Generation Computer Systems, 95(??):534–547, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18301687.

Chang:2016:CCA

[CKR16] Victor Chang, Yen-Hung Kuo, and Muthu Ramachandran. Cloud computing adoption framework: a security framework for business clouds. Future Generation Computer Systems, 57(??):24– 41, April 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003118.

Cao:2019:SMC

[CLAL19] Yang Cao, Chung-Horng Lung, Samuel A. Ajila, and Xiaolin Li. Support mechanisms for cloud configuration using XML filtering techniques: a case study in SaaS. Future Generation Computer Systems, 95 (??):52–67, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18319502.

Chang:2011:SMF

[CLC11] Ruay-Shiung Chang, Chun-Fu Lin, and Jen-Jom Chen. Selecting the most fitting resource for task execution. Future Generation Computer Systems, 27(2): 227–231, February 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chen:2019:BBS

[CLC^{+}19] Lanxiang Chen, Wai-Kong Lee, Chin-Chen Chang, Kim-Kwang Raymond Choo, and Nan Zhang. Blockchain based searchable encryption for electronic health record sharing. Future Generation Computer Systems, 95(??): 420–429, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314134.

Camacho:2018:ECD

[CLCMG^{+}18] David Camacho, Raúl Lara-Cabrera, J. J. Merelo-Guervós, Pedro A. Castillo, Carlos Cotta, Antonio J. Fernández-Leiva, Francisco Fernández de Vega, and Francisco Chávez. From ephemeral computing to deep bioinspired algorithms: New trends and applications. Future Generation Computer Systems, 88(??):735–746, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317692.

Choo:2018:FRF

[CLCY18] Kim-Kwang Raymond Choo, Rongxing Lu, Liqun Chen, and Xun Yi. A foggy research future: Advances and future opportunities in fog computing research. Future Generation Computer Systems, 78 (part 2)(??):677–679, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17319672.

Cheng:2019:TDM

[CLDC19] Peng Cheng, Yutong Lu, Yunfei Du, and Zhiguang Chen. Tiered data management system: Accelerating data processing on HPC systems. Future Generation Computer Systems, 101(??):894–908, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18320351.

Chang:2010:ADM

[CLH10] Ruay-Shiung Chang, Chun-Fu Lin, and Shih-Chun Hsi. Accessing data from many servers simultaneously and adaptively in data grids. Future Generation Computer Systems, 26(1):63– 71, January 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chen:2018:ECC

[CLH^{+}18] Min Chen, Wei Li, Yixue Hao, Yongfeng Qian, and Iztok Humar. Edge cognitive computing based smart healthcare system. Future Generation Computer Systems, 86(??):403–411, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17326766.

Clark:2011:CTU

[CLK11] John Clark, Sylvain Leblanc, and Scott Knight. Compromise through USB-based Hardware Trojan Horse device. Future Generation Computer Systems, 27(5): 555–563, May 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chen:2014:SBO

[CLL^{+}14] Lung-Pin Chen, Jien-An Lin, Kuan-Ching Li, Ching-Hsien Hsu, and Zhi-Xian Chen. A scalable blackbox-oriented elearning system based on desktop grid over private cloud. Future Generation Computer Systems, 38(??): 1–10, September 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000429.

Carie:2018:HDC

[CLL^{+}18a] Anil Carie, Mingchu Li, Chang Liu, Prakasha Reddy, and Waseef Jamal. Hybrid directional CR– MAC based on Q-learning with directional power control. Future Generation Computer Systems, 81(??): 340–347, April 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1731806X.

Chang:2018:RBM

[CLL18b] Ben-Jye Chang, Yu-Wei Lee, and Ying-Hsin Liang. Reward-based Markov chain analysis adaptive global resource management for inter-cloud computing. Future Generation Computer Systems, 79 (part 2)(??): 588–603, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1730612X.

Chen:2014:NEC

[CLM^{+}14a] Xiaofeng Chen, Jin Li, Jianfeng Ma, Wenjing Lou, and Duncan S. Wong. New and efficient conditional e-payment systems with transferability. Future Generation Computer Systems, 37(??):252–258, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300160X.

Cuzzocrea:2014:MCF

[CLM14b] Alfredo Cuzzocrea, Carson Kai-Sang Leung, and Richard Kyle MacKinnon. Mining constrained frequent itemsets from distributed uncertain data. Future Generation Computer Systems, 37(??):117– 126, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002458.

Chard:2016:GNP

[CLM^{+}16] Kyle Chard, Mattias Lidman, Brendan McCollam, Josh Bryan, Rachana Ananthakrishnan, Steven Tuecke, and Ian Foster. Globus Nexus: a Platformas-a-Service provider of research identity, profile, and group management. Future Generation Computer Systems, 56(??):571–583, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500285X.

Cardellini:2018:DSA

[CLNR18] Valeria Cardellini, Francesco Lo Presti, Matteo Nardelli, and Gabriele Russo Russo. Decentralized self-adaptation for elastic data stream processing. Future Generation Computer Systems, 87(??):171–185, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17326821.

Chetsa:2014:EPC

[CLP^{+}14] G. L. Tsafack Chetsa, L. Lefèvre, J. M. Pierson, P. Stolf, and G. Da Costa. Exploiting performance counters to predict and improve energy performance of HPC systems. Future Generation Computer Systems, 36(??): 287–298, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001556.

Carlini:2016:DMR

[CLR16] Emanuele Carlini, Alessandro Lulli, and Laura Ricci. dragon: Multidimensional range queries on distributed aggregation trees. Future Generation Computer Systems, 55(??):101– 115, February 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002526.

Chen:2018:IBB

[CLR18] Long Chen, Xiaoping Li, and Rubén Ruiz. Idle block based methods for cloud workflow scheduling with preemptive and nonpreemptive tasks. Future Generation Computer Systems, 89(??):659–669, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18300761.

Cai:2017:DBD

[CLRL17] Zhicheng Cai, Xiaoping Li, Rubén Ruiz, and Qianmu Li. A delay-based dynamic scheduling algorithm for bag-of-task workflows with stochastic task execution times in clouds. Future Generation Computer Systems, 71(??):57– 72, June 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300870.

Cai:2018:PFS

[CLRL18] Zhicheng Cai, Xiaoping Li, Rubén Ruiz, and Qianmu Li. Price forecasting for spot instances in cloud computing. Future Generation Computer Systems, 79 (part 1)(??):38– 53, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312621.

Chen:2019:DPB

[CLS19a] Jian Chen, Zhihan Lv, and Houbing Song. Design of personnel big data management system based on blockchain. Future Generation Computer Systems, 101(??):1122–1129, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313354.

Cheng:2019:PBM

[CLS^{+}19b] Lichen Cheng, Jiqiang Liu, Chunhua Su, Kaitai Liang, Guangquan Xu, and Wei Wang. Polynomial-based modifiable blockchain structure for removing fraud transactions. Future Generation Computer Systems, 99(??):154–163, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19303012.

Chau:2014:SVM

[CLY14] Asdrúbal López Chau, Xiaoou Li, and Wen Yu. Support vector machine classification for large datasets using decision tree and Fisher linear discriminant. Future Generation Computer Systems, 36(??):57– 65, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001350.

Cao:2018:PPA

[CLZ18] Yuan Cao, Peng Li, and Yuzhuo Zhang. Parallel processing algorithm for railway signal fault diagnosis data based on cloud computing. Future Generation Computer Systems, 88(??):279–283, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18306927.

Chejerla:2017:QGR

[CM17] Brijesh Kashyap Chejerla and Sanjay. K. Madria. QoS guaranteeing robust scheduling in attack resilient cloud integrated cyber physical system. Future Generation Computer Systems, 75(??):145–157, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302650.

Cafaro:2011:SSG

[CMA11] Massimo Cafaro, Henning Müller, and Nabil Abdennadher. Special section: Grid and pervasive computing 2009. Future Generation Computer Systems, 27(5):587–589, May 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Carvalho:2017:CPI

[CMB17] Marcus Carvalho, Daniel A.

Menascé, and Francisco Brasileiro. Capacity planning for IaaS cloud providers offering multiple service classes. Future Generation Computer Systems, 77(??):97–111, December 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308561.

Curcin:2014:IIP

[CMD^{+}14] V. Curcin, S. Miles, R. Danger, Y. Chen, R. Bache, and A. Taweel. Implementing interoperable provenance in biomedical research. Future Generation Computer Systems, 34(??):1–16, May 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002653.

Cantabella:2019:ASB

[CMEA^{+}19]  Magdalena  Cantabella,

Raquel Martínez-España, Belén Ayuso, Juan Antonio Yáñez, and Andrés Muñoz. Analysis of student behavior in learning management systems through a big data framework. Future Generation Computer Systems, 90(??):262–272, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17329217.

Celesti:2019:SCV

[CMG^{+}19] Antonio Celesti, Davide Mulfari, Antonino Galletta, Maria Fazio, Lorenzo Carnevale, and Massimo Villari. A study on container virtualization for guarantee quality of service in Cloud-of-Things. Future Generation Computer Systems, 99(??):356–364, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18325615.

Chatterjee:2019:NML

[CMI^{+}19] Rajdeep Chatterjee, Tanmoy Maitra, SK Hafizul Islam, Mohammad Mehedi Hassan, Atif Alamri, and Giancarlo Fortino. A novel machine learning based feature selection for motor imagery EEG signal classification in Internet of Medical Things environment. Future Generation Computer Systems, 98(??):419– 434, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18327699.

Chaturvedi:2019:SSD

[CMNK19] Kanishk Chaturvedi, Andreas Matheus, Son H. Nguyen, and Thomas H. Kolbe. Securing spatial data infrastructures for distributed smart city applications and services. Future Generation Computer Systems, 101(??):723–736, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18330024.

Chianese:2017:AEB

[CMP^{+}17] Angelo Chianese, Fiammetta Marulli, Francesco Piccialli, Paolo Benedusi, and Jai E. Jung. An associative engines based approach supporting collaborative analytics in the Internet of Cultural Things. Future Generation Computer Systems, 66(??):187– 198, January 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300929.

Castello:2018:ALT

[CMS^{+}18] Adrián Castelló, Rafael Mayo, Kevin Sala, Vicenç Beltran, Pavan Balaji, and Antonio J. Peña. On the adequacy of lightweight thread approaches for highlevel parallel programming models. Future Generation Computer Systems, 84(??): 22–31, July 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17328819.

Cesario:2016:DVC

[CMT16] Eugenio Cesario, Carlo Mastroianni, and Domenico Talia. Distributed volunteer computing for solving ensemble learning problems. Future Generation Computer Systems, 54(??): 68–78, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002332.

Costa:2018:ASV

[CMVA18] Gabriele Costa, Alessio Merlo, Luca Verderame, and Alessandro Armando. Automatic security verification of mobile app configurations. Future Generation Computer Systems, 80(??):519–536, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301996.

Cala:2016:SEW

[CMX^{+}16] J. Cala, E. Marei, Y. Xu, K. Takeda, and P. Missier. Scalable and efficient wholeexome data processing using workflows on the cloud. Future Generation Computer Systems, 65(??):153– 168, December 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16000030.

Cacciari:2012:SBM

[CMZ^{+}12] Claudio Cacciari, Daniel Mallmann, Csilla Zsigri, Francesco D’Andria, Björn Hagemeier, Angela Rumpl, Wolfgang Ziegler, and Josep Martrat. SLA-based management of software licenses as Web service resources in distributed computing infrastructures. Future Generation Computer Systems, 28(8):1340–1349, October 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002287.

Cui:2018:RDF

[CMZ^{+}18] Yongfeng Cui, Yuankun Ma, Zhongyuan Zhao, Ya Li, Wei Lu, and Wanneng Shu. Research on data fusion algorithm and anti-collision algorithm based on Internet of Things. Future Generation Computer Systems, 85(??):107–115, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17329035.

Carneiro:2017:QEE

[CN17] Davide Carneiro and Paulo Novais. Quantifying the effects of external factors on individual performance. Future Generation Computer Systems, 66(??):171– 186, January 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301303.

Carneiro:2019:PCT

[CND^{+}19] Davide Carneiro, Paulo Novais, Dalila Durães, José Miguel Pego, and Nuno Sousa. Predicting completion time in high-stakes exams. Future Generation Computer Systems, 92(??):549–559, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17313092.

Ciobanu:2019:DCA

[CNP^{+}19] Radu-Ioan Ciobanu, Catalin Negru, Florin Pop, Ciprian Dobre, Constandinos X. Mavromoustakis, and George Mastorakis. Drop computing: Ad-hoc dynamic collaborative computing. Future Generation Computer Systems, 92(??):889–899, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17305678.

Cotroneo:2019:OCV

[CNR19] Domenico Cotroneo, Roberto Natella, and Stefano Rosiello. Overload control for virtual network functions under CPU contention. Future Generation Computer Systems, 99(??):164–176, October 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18308562.

Cem:2013:PPF

[ÇÖ13] Emrah Çem and Öznur Özkasap. ProFID: Practical frequent items discovery in peer-to-peer networks. Future Generation Computer Systems, 29(6):1544–1560, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001859.

Cheng:2010:RFG

[COC10] Wai-Khuen Cheng, Boon-Yaik Ooi, and Huah-Yong Chan. Resource federation in Grid using automated intelligent agent negotiation. Future Generation Computer Systems, 26(8): 1116–1126, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Carabas:2017:EEV

[CP17] Mihai Carabas and Pantelimon George Popescu. Energy-efficient virtualized clusters. Future Generation Computer Systems, 74(??):151–157, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003313.

Chen:2014:TRM

[CPA14] Peng Chen, Beth Plale, and Mehmet S. Aktas. Temporal representation for mining scientific data provenance. Future Generation Computer Systems, 36(??):363–378, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002355.

Castiglione:2015:CBA

[CPD^{+}15] Arcangelo Castiglione, Raffaele Pizzolante, Alfredo De Santis, Bruno Carpentieri, Aniello Castiglione, and Francesco Palmieri. Cloud-based adaptive compression and secure management services for 3D healthcare data. Future Generation Computer Systems, 43–44(??):120–134, February 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001332.

Cai:2013:NSO

[CPDJ13] Haibin Cai, Chao Peng, Robert H. Deng, and Linhua Jiang. A novel serviceoriented intelligent seamless migration algorithm and application for pervasive computing environments. Future Generation Computer Systems, 29(8):1919–1930, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000459.

Castiglione:2017:CCA

[CPE^{+}17] Arcangelo Castiglione, Raffaele Pizzolante, Christian Esposito, Alfredo De Santis, Francesco Palmieri, and Aniello Castiglione. A collaborative clinical analysis service based on theory of evidence, fuzzy linguistic sets and prospect theory and its application to craniofacial disorders in infants. Future Generation Computer Systems, 67(??):230– 241, February 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302515.

Camarasu-Pop:2016:CAM

[CPGBC16] Sorina Camarasu-Pop, Tristan Glatard, and Hugues Benoit-Cattin. Combining analytical modeling, realistic simulation and real experimentation for the optimization of Monte-Carlo applications on the European Grid Infrastructure. Future Generation Computer Systems, 57(??):13– 23, April 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003398.

Camarasu-Pop:2013:MCS

[CPGdS^{+}13] Sorina Camarasu-Pop, Tristan Glatard, Rafael Ferreira da Silva, Pierre Gueth, David Sarrut, and Hugues Benoit-Cattin. Monte Carlo simulation on heterogeneous distributed systems: a computing framework with parallel merging and checkpointing strategies. Future Generation Computer Systems, 29(3):728–738, March 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001768.

Compton:2019:PDG

[CPLH19] Andrew J. Compton, John M.

Pecarina, Alan C. Lin, and Kenneth M. Hopkinson. PeerAppear: a distributed geospatial index supporting collaborative world model construction and maintenance. Future Generation Computer Systems, 95(??):802–815, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17307409.

Ciechanowski:2019:SUV

[CPMG19] Leon Ciechanowski, Aleksandra Przegalinska, Mikolaj Magnuski, and Peter Gloor. In the shades of the uncanny valley: An experimental study of humanchatbot interaction. Future Generation Computer Systems, 92(??):539–548, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17312268.

Cotroneo:2016:ARC

[CPP16] Domenico Cotroneo, Andrea Paudice, and Antonio Pecchia. Automated root cause identification of security alerts: Evaluation in a SaaS Cloud. Future Generation Computer Systems, 56(??):375– 387, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002897.

Carneiro:2018:UBF

[CPP^{+}18] Davide Carneiro, Ana P.

Pinheiro, Marta Pereira, Inês Ferreira, Miguel Domingues, and Paulo Novais. Using behavioral features in tablet-based auditory emotion recognition studies. Future Generation Computer Systems, 89(??):646– 658, December 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17328534.

Candel:2018:AMC

[CPSD18] Francisco Candel, Salvador Petit, Julio Sahuquillo, and José Duato. Accurately modeling the onchip and off-chip GPU memory subsystem. Future Generation Computer Systems, 82(??):510–519, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17302091.

Colomo-Palacios:2014:SIE

[CPSRG14] Ricardo Colomo-Palacios, Vladimir Stantchev, and Alejandro Rodríguez-González. Special issue on exploiting semantic technologies with particularization on linked data over grid and cloud architectures. Future Generation Computer Systems, 32(??):260–262, March 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002392.

Chai:2019:BPI

[CPW19] Tingting Chai, Shitala Prasad, and Shenghui Wang. Boosting palmprint identification with gender information using Deep-Net. Future Generation Computer Systems, 99(??): 41–53, October 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18331261.

Cai:2019:MRB

[CQW^{+}19] Lin Cai, Yong Qi, Wei Wei, Jinsong Wu, and Jingwei Li. mrMoulder: A recommendation-based adaptive parameter tuning approach for big data processing platform. Future Generation Computer Systems, 93(??):570–582, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17318526.

Corradi:2014:MCS

[CR14] Antonio Corradi and Omer F. Rana. The management of cloud systems. Future Generation Computer Systems, 32(??):24– 26, March 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002343.

Conejero:2016:AHP

[CRB^{+}16] Javier Conejero, Omer Rana, Peter Burnap, Jeffrey Morgan, Blanca Caminero, and Carmen Carrión. Analyzing Hadoop power consumption and impact on application QoS. Future Generation Computer Systems, 55(??):213–223, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000643.

Carlini:2013:FLD

[CRC13] Emanuele Carlini, Laura Ricci, and Massimo Coppola. Flexible load distribution for hybrid distributed virtual environments. Future Generation Computer Systems, 29(6):1561–1572, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1200177X.

Costa:2019:EDA

[CRC^{+}19] Angelo Costa, Jaime A. Rincon, Carlos Carrascosa, Vicente Julian, and Paulo Novais. Emotions detection on an ambient intelligent system using wearable devices. Future Generation Computer Systems, 92(??):479–489, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17310269.

Calatrava:2016:SMC

[CRM^{+}16] Amanda Calatrava, Eloy Romero, Germán Moltó, Miguel Caballer, and Jose Miguel Alonso. Selfmanaged cost-efficient virtual elastic clusters on hybrid cloud infrastructures. Future Generation Computer Systems, 61(??):13– 25, August 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300024.

Chen:2018:LLR

[CRRC18] Zitao Chen, Wei Ren, Yi Ren, and Kim-Kwang Raymond Choo. LiReK: A lightweight and realtime key establishment scheme for wearable embedded devices by gestures or motions. Future Generation Computer Systems, 84(??):126–138, July 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17305666.

Cannataro:2010:SSB

[CRSdS10] Mario Cannataro, Mathilde Romberg, Joakim Sundnes, and Rodrigo Weber dos Santos. Special section: Biomedical and bioinformatics challenges to computer science. Future Generation Computer Systems, 26(3):421–423, March 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cunha:2017:JPA

[CRTN17] Renato L. F. Cunha, Eduardo R. Rodrigues, Leonardo P. Tizzei, and Marco A. S. Netto. Job placement advisor based on turnaround predictions for HPC hybrid clouds. Future Generation Computer Systems, 67(??):35–46, February 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302655.

Casanova:2015:IPR

[CRVZ15] Henri Casanova, Yves Robert, Frédéric Vivien, and Dounia Zaidouni. On the impact of process replication on executions of large-scale parallel applications with coordinated checkpointing. Future Generation Computer Systems, 51(??):7–19, October 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000953.

Chang:2016:EFS

[CRW^{+}16] Victor Chang, Muthu Ramachandran, Gary Wills, Robert John Walters, Chung-Sheng Li, and Paul Watters. Editorial for FGCS special issue: Big Data in the cloud. Future Generation Computer Systems, 65(??):73–75, December 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630084X.

Cui:2019:PCL

[CRWZ19] Jin Cui, Lei Ren, Xiaokang Wang, and Lin Zhang. Pairwise comparison learning based bearing health quantitative modeling and its application in service life prediction. Future Generation Computer Systems, 97(??):578–586, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18328097.

Chen:2018:PPB

[CRYG18] Bo-Wei Chen, Seungmin Rho, Laurence T. Yang, and Yu Gu. Privacypreserved big data analysis based on asymmetric imputation kernels and multiside similarities. Future Generation Computer Systems, 78 (part 2)(??):859– 866, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16305623.

Chinnaiah:2012:GRB

[CS12] Valliyammai Chinnaiah and Thamarai Selvi Somasundaram. A Grid resource brokering strategy based on resource and network performance in Grid. Future Generation Computer Systems, 28(3):491– 499, March 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001683.

Cokic:2019:SDN

[CS19] Mita Cokic and Ivan Seskar. Software defined network management for dynamic smart GRID traffic. Future Generation Computer Systems, 96(??): 270–282, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18311063.

Choo:2018:MSC

[CSC18] Kim-Kwang Raymond Choo, Yongtang Shi, and Zengqiang Chen. Measurements and security of complex networks and systems: Research advances and challenges. Future Generation Computer Systems, 83(??):374–375, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18304618.

Carastan-Santos:2017:FEH

[CSdCM^{+}17] Danilo Carastan-Santos, Raphael Y. de Camargo, David C. Martins, Jr., Siang W. Song, and Luiz C. S. Rozante. Finding exact hitting set solutions for systems biology applications using heterogeneous GPU clusters. Future Generation Computer Systems, 67(??):418–429, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630022X.

Chao:2018:GBP

[CSG^{+}18] Zemin Chao, Shengfei Shi, Hong Gao, Jizhou Luo, and Hongzhi Wang. A graybox performance model for Apache Spark. Future Generation Computer Systems, 89(??):58–67, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17323233.

Chen:2017:RSE

[CSJ^{+}17] Xianzhang Chen, Edwin H.-M. Sha, Weiwen Jiang, Chaoshu Yang, Ting Wu, and Qingfeng Zhuge. Refinery swap: an efficient swap mechanism for hybrid DRAM–NVM systems. Future Generation Computer Systems, 77(??):52–64, December 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630588X.

Chen:2017:EWS

[CSL17] Zhen Chen, Limin Shen, and Feng Li. Exploiting Web service geographical neighborhood for collaborative QoS prediction. Future Generation Computer Systems, 68(??):248– 259, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303533.

Chongdarakul:2018:THA

[CSL18] Waralak Chongdarakul, Peraphon Sophatsathit, and Chidchanok Lursinsap. Theoretical and heuristic aspects of heterogeneous system scheduling with constraints on client’s multiple I/O ports. Future Generation Computer Systems, 78 (part 3)(??):901– 919, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17316072.

Chen:2019:YNM

[CSL19] Zhen Chen, Limin Shen, and Feng Li. Your neighbors are misunderstood: On modeling accurate similarity driven by data range to collaborative web service QoS prediction. Future Generation Computer Systems, 95(??):404–419, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831611X.

Cui:2013:GEP

[CSP13] Xiaohui Cui, Jesse St.

Charles, and Thomas Potok. GPU enhanced parallel computing for large scale data clustering. Future Generation Computer Systems, 29(7):1736–1741, September 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001707.

Cui:2017:UMD

[CSQL17] Baojiang Cui, Peilin Shi, Weikong Qi, and Ming Li. Uploading multiply deferrable big data to the cloud platform using costeffective online algorithms. Future Generation Computer Systems, 67(??):276– 285, February 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301054.

Caan:2012:EGB

[CSV^{+}12] M. W. A. Caan, S. Shahand, F. M. Vos, A. H. C. van Kampen, and S. D. Olabarriaga. Evolution of grid-based services for Diffusion Tensor Image analysis. Future Generation Computer Systems, 28(8):1194–1204, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1200060X.

Costa:2019:IWC

[CSV^{+}19] Alessandro Costa, Eva Sciacca, Fabio Vitello, Ugo Becciani, Pietro Massimino, Simone Riggi, and David Sanchez. An integrated workspace for the Cherenkov Telescope Array. Future Generation Computer Systems, 94(??): 811–819, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17309585.

Chen:2018:SIA

[CSYY18] Jiageng Chen, Chunhua Su, Kuo-Hui Yeh, and Moti Yung. Special issue on advanced persistent threat. Future Generation Computer Systems, 79 (part 1)(??):243– 246, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17324913.

Chen:2014:FFF

[CsZW14] Ting Chen, Xiao song Zhang, and Yue Wu. FPM: Four-factors Propagation Model for passive P2P worms. Future Generation Computer Systems, 36(??): 133–141, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001398.

Chen:2013:SAD

[CsZzG^{+}13] Ting Chen, Xiao song Zhang, Shi ze Guo, Hong yuan Li, and Yue Wu. State of the art: Dynamic symbolic execution for automated test generation. Future Generation Computer Systems, 29(7):1758– 1773, September 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000398.

C:2019:IMT

[CT19a] Shilaja C. and Arunprasath T. Internet of Medical Things-load optimization of power flow based on hybrid enhanced grey wolf optimization and Dragonfly algorithm. Future Generation Computer Systems, 98(??):319–330, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326360.

C:2019:OPF

[CT19b] Shilaja C. and Arunprasath T. Optimal power flow using moth swarm algorithm with gravitational search algorithm considering wind power. Future Generation Computer Systems, 98(??):708–715, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18327250.

Chen:2019:DBI

[CT19c] Chia-Chen Chen and Jia-Lun Tsai. Determinants of behavioral intention to use the personalized location-based mobile tourism application: an empirical study by integrating TAM with ISSM. Future Generation Computer Systems, 96(??):628– 638, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302534.

Casas:2017:BSD

[CTR^{+}17] Israel Casas, Javid Taheri, Rajiv Ranjan, Lizhe Wang, and Albert Y. Zomaya. A balanced scheduler with data reuse and replication for scientific workflows in cloud computing systems. Future Generation Computer Systems, 74(??):168– 178, September 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500388X.

Choudhary:2019:AIC

[CTU19]  Meenakshi  Choudhary,

Vivek Tiwari, and Venkanna U. An approach for iris contact lens detection and classification using ensemble of customized DenseNet and SVM. Future Generation Computer Systems, 101(??):1259– 1270, December 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331911.

Calheiros:2012:CSE

[CTVB12] Rodrigo N. Calheiros, Adel Nadjaran Toosi, Christian Vecchiola, and Rajkumar Buyya. A coordinator for scaling elastic applications across multiple clouds. Future Generation Computer Systems, 28(8):1350–1362, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000635.

Cuzzocrea:2014:IMA

[Cuz14] Alfredo Cuzzocrea. Innovative methods and algorithms for advanced data-intensive computing. Future Generation Computer Systems, 37(??):60– 63, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1400065X.

Calheiros:2012:APQ

[CVKB12] Rodrigo N. Calheiros, Christian Vecchiola, Dileban Karunamoorthy, and Rajkumar Buyya. The Aneka platform and QoSdriven resource provisioning for elastic applications on hybrid Clouds. Future Generation Computer Systems, 28(6):861– 870, June 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001397.

Calzarossa:2019:MFC

[CVT19] Maria Carla Calzarossa, Marco L. Della Vedova, and Daniele Tessera. A methodological framework for cloud resource provisioning and scheduling of data parallel applications under uncertainty. Future Generation Computer Systems, 93(??):212–223, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18305752.

Cheng:2013:NPI

[CW13a] Luwei Cheng and Cho-Li Wang. Network performance isolation for latency-sensitive cloud applications. Future Generation Computer Systems, 29(4):1073–1084, June 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001331.

Chmaj:2013:PCS

[CW13b] Grzegorz Chmaj and Krzysztof Walkowiak. A P2P computing system for overlay networks. Future Generation Computer Systems, 29(1):242–249, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X10002293.

Chang:2016:MCC

[CW16] Victor Chang and Gary Wills. A model to compare cloud and non-cloud storage of Big Data. Future Generation Computer Systems, 57(??):56–76, April 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003167.

Chen:2016:CGB

[CWJ16] Shuhong Chen, Guojun Wang, and Weijia Jia. Cluster-group based trusted computing for mobile social networks using implicit social behavioral graph. Future Generation Computer Systems, 55(??):391–400, February 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001265.

Cheng:2018:CSA

[CWJ^{+}18a] Jiatang Cheng, Lei Wang, Qiaoyong Jiang, Zijian Cao, and Yan Xiong. Cuckoo search algorithm with dynamic feedback information. Future Generation Computer Systems, 89(??):317–334, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18300244.

Cheng:2018:HCE

[CWJ^{+}18b] Yongli Cheng, Fang Wang, Hong Jiang, Yu Hua, Dan Feng, Yunxiang Wu, Tingwei Zhu, and Wenzhong Guo. A highly cost-effective task scheduling strategy for very large graph computation. Future Generation Computer Systems, 89(??):698–712, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17323683.

Cao:2019:CLN

[CWJD19] Jinxin Cao, Hongcui Wang, Di Jin, and Jianwu Dang. Combination of links and node contents for community discovery using a graph regularization approach. Future Generation Computer Systems, 91(??):361–370, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17327711.

Cai:2018:ENC

[CWL^{+}18] Jun Cai, Yu Wang, Yan Liu, Jian-Zhen Luo, Wenguo Wei, and Xiaoping Xu. Enhancing network capacity by weakening community structure in scalefree network. Future Generation Computer Systems, 87(??):765–771, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1730393X.

Chien:2019:SBA

[CWL^{+}19] Wei-Che Chien, Hung-Yen Weng, Chin-Feng Lai, Zhang Fan, Han-Chieh Chao, and Ying Hu. A SFC-based access point switching mechanism for software-defined wireless network in IoV. Future Generation Computer Systems, 98(??):577–585, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1832082X.

Cui:2019:DDD

[CWLZ19] Jie Cui, Mingjun Wang, Yonglong Luo, and Hong Zhong. DDoS detection and defense mechanism based on cognitive-inspired computing in SDN. Future Generation Computer Systems, 97(??):275–283, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18324336.

Chen:2014:CRS

[CWSW14] Shuai-Min Chen, Mu-En Wu, Hung-Min Sun, and King-Hang Wang. CR-FID: an RFID system with a cloud database as a back-end server. Future Generation Computer Systems, 30(??):155–161, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001039.

Chung:2019:SIV

[CWUS19] Wei-Ho Chung, Mu-En Wu, Yeong-Luh Ueng, and Yu-Hsuan Su. Seal imprint verification via feature analysis and classifications. Future Generation Computer Systems, 101(??):458–466, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19300299.

Chen:2013:HMS

[CWW^{+}13] Dan Chen, Lizhe Wang, Xiaomin Wu, Jingying Chen, Samee U. Khan, Joanna Kolodziej, Mingwei Tian, Fang Huang, and Wangyang Liu. Hybrid modelling and simulation of huge crowd over a hierarchical Grid architecture. Future Generation Computer Systems, 29(5): 1309–1317, July 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000593.

Chen:2016:MCC

[CWW^{+}16] Liwei Chen, Yipeng Wang, Huandong Wang, Wenxiang Wang, Hua Jing, and Guangfei Zhang. Multiplecombinational-channel: a network architecture for workload balance and deadlock free. Future Generation Computer Systems, 56(??):238–246, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1500271X.

Chen:2017:IDI

[CWZ^{+}17] Chao Chen, Sheng Wen, Jun Zhang, Yang Xiang, Jonathan Oliver, Abdulhameed Alelaiwi, and Mohammad Mehedi Hassan. Investigating the deceptive information in Twitter spam. Future Generation Computer Systems, 72(??): 319–326, July 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301546.

Cheng:2018:ECR

[CXC^{+}18] Yuxia Cheng, Yang Xiang, Wenzhi Chen, Houcine Hassan, and Abdulhameed Alelaiwi. Efficient cache resource aggregation using adaptive multi-level exclusive caching policies. Future Generation Computer Systems, 86(??):964–974, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17320952.

Castane:2018:OHR

[CXDM18] Gabriel G. Castañé, Huanhuan Xiong, Dapeng Dong, and John P. Morrison. An ontology for heterogeneous resources management interoperability and HPC in the cloud. Future Generation Computer Systems, 88(??):373–384, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17330467.

Chen:2017:ETS

[CXL^{+}17] Weihong Chen, Guoqi Xie, Renfa Li, Yang Bai, Chunnian Fan, and Keqin Li. Efficient task scheduling for budget constrained parallel applications on heterogeneous cloud computing systems. Future Generation Computer Systems, 74(??): 1–11, September 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16304411.

Chen:2019:WBS

[CXWT19] Yu-Chi Chen, Xin Xie, Peter Shaojui Wang, and Raylin Tso. Witness-based searchable encryption with optimal overhead for cloudedge computing. Future Generation Computer Systems, 100(??):715–723, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19301086.

Cao:2019:LAR

[CXZ^{+}19] Kun Cao, Guo Xu, Junlong Zhou, Mingsong Chen, Tongquan Wei, and Keqin Li. Lifetime-aware real-time task scheduling on fault-tolerant mixedcriticality embedded systems. Future Generation Computer Systems, 100(??):165–175, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1832315X.

Chang:2018:RDM

[CXZC18] Wenbing Chang, Zhenzhong Xu, Shenghan Zhou, and Wen Cao. Research on detection methods based on doc2vec abnormal comments. Future Generation Computer Systems, 86(??):656–662, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18303042.

Cha:2012:SFV

[CY12] Sungdeok Cha and Junbeom Yoo. A safetyfocused verification using software fault trees. Future Generation Computer Systems, 28(8):1272–1282, October 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000070.

Chen:2019:EMB

[CYJ19] Jun Chen, Chen Yu, and Hai Jin. Evaluation model for business sites planning based on online and offline datasets. Future Generation Computer Systems, 91(??):465–474, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18306605.

Cheng:2019:PVS

[CYW^{+}19] Yuxia Cheng, Huijuan Yao, Yu Wang, Yang Xiang, and Hongpei Li. Protecting VNF services with smart online behavior anomaly detection method. Future Generation Computer Systems, 95(??):265–276, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18310744.

Chen:2015:TTS

[CYZK15] Liang Chen, Zheng Yan, Weidong Zhang, and Raimo Kantola. TruSMS: a trustworthy SMS spam control system based on trust management. Future Generation Computer Systems, 49(??):77–93, August 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001320.

Chen:2012:NMM

[CZ12] Xi Chen and Jiashu Zhang. A novel maximum margin neighborhood preserving embedding for face recognition. Future Generation Computer Systems, 28(1):212–217, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X10002128.

Chen:2014:SSD

[CZ14] Jingqiang Chen and Hai Zhuge. Summarization of scientific documents by detecting common facts in citations. Future Generation Computer Systems, 32(??): 246–252, March 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001635.

Chen:2019:ESD

[CZ19] Jingqiang Chen and Hai Zhuge. Extractive summarization of documents with images based on multi-modal RNN. Future Generation Computer Systems, 99(??):186–196, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326876.

Chen:2018:RAM

[CZH^{+}18] Jinbo Chen, Huiling Zhou, Hongyu Hu, Yan Song, Daniela Gifu, Youzhu Li, and Ye Huang. Research on agricultural monitoring system based on convolutional neural network. Future Generation Computer Systems, 88(??):271– 278, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18310756.

Cao:2018:MRO

[CZL^{+}18a] Bin Cao, Jianwei Zhao, Xin Liu, Xinyuan Kang, Shan Yang, Kai Kang, and Ming Yu. Multiobjective recommendation optimization via utilizing distributed parallel algorithm. Future Generation Computer Systems, 86(??):1259–1268, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17319489.

Cui:2018:SAB

[CZL^{+}18b] Laizhong Cui, Kai Zhang, Genghui Li, Xizhao Wang, Shu Yang, Zhong Ming, Joshua Zhexue Huang, and Nan Lu. A smart artificial bee colony algorithm with distance-fitness-based neighbor search and its application. Future Generation Computer Systems, 89(??):478–493, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1831152X.

Chen:2018:SCM

[CZM^{+}18] Feiyu Chen, Zhiwei Zhao, Geyong Min, Weifeng Gao, Jinjun Chen, Hancong Duan, and Po Yang. Speed control of mobile chargers serving wireless rechargeable networks. Future Generation Computer Systems, 80(??):242–249, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16307506.

Cai:2015:TAS

[CZT^{+}15] Chang Cai, Ke Zeng, Lin Tang, Dan Chen, Weizhou Peng, Jiaqing Yan, and Xiaoli Li. Towards adaptive synchronization measurement of large-scale non-stationary non-linear data. Future Generation Computer Systems, 43–44(??):110–119, February 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001241.

Chen:2018:HTE

[CZXL18] Xu Chen, Jie Zhang, Zheng Xu, and Jin Liu. HIB-tree: An efficient index method for the big data analytics of large-scale human activity trajectories. Future Generation Computer Systems, 86(??):1269–1278, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18300116.

Cao:2018:DPC

[CZY^{+}18] Bin Cao, Jianwei Zhao, Po Yang, Zhihan Lv, Xin Liu, Xinyuan Kang, Shan Yang, Kai Kang, and Amjad Anvari-Moghaddam. Distributed parallel cooperative coevolutionary multi-objective large-scale immune algorithm for deployment of wireless sensor networks. Future Generation Computer Systems, 82(??):256–267, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17313523.

Cao:2019:MFS

[CZY^{+}19] Bin Cao, Jianwei Zhao, Po Yang, Peng Yang, Xin Liu, Jun Qi, Andrew Simpson, Mohamed Elhoseny, Irfan Mehmood, and Khan Muhammad. Multiobjective feature selection for microarray data via distributed parallel algorithms. Future Generation Computer Systems, 100(??):952–981, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18318995.

Chen:2018:VKS

[CZZ^{+}18] Zehong Chen, Fangguo Zhang, Peng Zhang, Joseph K. Liu, Jiwu Huang, Hanbang Zhao, and Jian Shen. Verifiable keyword search for secure big data-based mobile healthcare networks with fine-grained authorization control. Future Generation Computer Systems, 87(??):712–724, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17300730.

Dessi:2016:MLA

[DA16] Andrea Dessi and Maurizio Atzori. A machinelearning approach to ranking RDF properties. Future Generation Computer Systems, 54(??):366–377, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001764.

Dimitriou:2018:WTD

[DA18] Tassos Dimitriou and Naser Al Ibrahim. “I wasn’t there” — deniable, privacy-aware scheme for decentralized locationbased services. Future Generation Computer Systems, 86(??):253–265, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17304491.

deAlfonso:2013:EEA

[dACAM13] Carlos de Alfonso, Miguel Caballer, Fernando Alvarruiz, and Germán Moltó. An economic and energyaware analysis of the viability of outsourcing cluster computing to a cloud. Future Generation Computer Systems, 29(3):704– 712, March 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001720.

deAssuncao:2016:IUP

[dACNC16] Marcos Dias de Assunção, Carlos H. Cardonha, Marco A. S. Netto, and Renato L. F. Cunha. Impact of user patience on autoscaling resource capacity for cloud services. Future Generation Computer Systems, 55(??):41–50, February 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002794.

Dosanjh:2014:EDS

[DBD^{+}14] S. S. Dosanjh, R. F. Barrett, D. W. Doerfler, S. D. Hammond, K. S. Hemmert, M. A. Heroux, P. T. Lin, K. T. Pedretti, A. F. Rodrigues, T. G. Trucano, and J. P. Luitjens. Exascale design space exploration and co-design. Future Generation Computer Systems, 30(??):46– 58, January 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000782.

Dawaliby:2019:ADN

[DBP19] Samir Dawaliby, Abbas Bradai, and Yannis Pousset. Adaptive dynamic network slicing in LoRa networks. Future Generation Computer Systems, 98(??):697–707, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18318429.

Durfee:2014:UHS

[DBS14] Edmund H. Durfee, James C.

Boerkoel, Jr., and Jason Sleight. Using hybrid scheduling for the semi-autonomous formation of expert teams. Future Generation Computer Systems, 31(??):200–212, February 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1300068X.

DOrazio:2017:TCS

[DC17] Christian J. D’Orazio and Kim-Kwang Raymond Choo. A technique to circumvent SSL/TLS validations on iOS devices. Future Generation Computer Systems, 74(??):366–374, September 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302801.

Diro:2018:DAD

[DC18a] Abebe Abeshu Diro and Naveen Chilamkurti. Distributed attack detection scheme using deep learning approach for Internet of Things. Future Generation Computer Systems, 82(??):761–768, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17308488.

DOrazio:2018:CIS

[DC18b] Christian J. D’Orazio and Kim-Kwang Raymond Choo. Circumventing iOS security mechanisms for APT forensic investigations: A security taxonomy for cloud apps. Future Generation Computer Systems, 79 (part 1)(??): 247–261, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X16305647.

Dong:2019:NGA

[DC19] Xueshi Dong and Yongle Cai. A novel genetic algorithm for large scale colored balanced traveling salesman problem. Future Generation Computer Systems, 95(??):727–742, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831728X.

Davoli:2019:DEP

[DCBF19] Luca Davoli, Antonio Cilfone, Laura Belli, and Gianluigi Ferrari. Design and experimental performance analysis of a B.A.T.M.A.N.-based double Wi-Fi interface mesh network. Future Generation Computer Systems, 92(??):593–603, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17323452.

Dou:2013:CBF

[DCC13] Wanchun Dou, Qi Chen, and Jinjun Chen. A confidence-based filtering method for DDoS attack defense in cloud environment. Future Generation Computer Systems, 29(7):1838–1850, September 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12002312.

Dou:2014:MSN

[DCC^{+}14] Minggang Dou, Jingying Chen, Dan Chen, Xiaodao Chen, Ze Deng, Xuguang Zhang, Kai Xu, and Jian Wang. Modeling and simulation for natural disaster contingency planning driven by high-resolution remote sensing images. Future Generation Computer Systems, 37(??):367–377, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002823.

Coutinho:2015:OVM

[dCCDFdO15] Rafaelli de C. Coutinho, Lúcia M. A. Drummond, Yuri Frota, and Daniel de Oliveira. Optimizing virtual machine allocation for parallel scientific workflows in federated clouds. Future Generation Computer Systems, 46(??):51– 68, May 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002027.

Ding:2019:SSA

[DCF19] Ding Ding, Mauro Conti, and Renato Figueiredo. SAND: Social-aware, networkfailure resilient, and decentralized microblogging system. Future Generation Computer Systems, 93(??): 637–650, April 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0167739X18304370.

Danger:2015:ACV

[DCMB15] Roxana Danger, Vasa Curcin, Paolo Missier, and Jeremy Bryans. Access control and view generation for provenance graphs. Future Generation Computer Systems, 49(??):8– 27, August 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500031X.

Duan:2017:EAS

[DCMW17] Hancong Duan, Chao Chen, Geyong Min, and Yu Wu. Energy-aware scheduling of virtual machines in heterogeneous cloud computing systems. Future Generation Computer Systems, 74(??):142– 150, September 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300292.

daCruz:2019:PBA

[dCRL^{+}19] Mauro A. A. da Cruz, Joel J. P. C. Rodrigues, Pascal Lorenz, Petar Solic, Jalal Al-Muhtadi, and Victor Hugo C. Albuquerque. A proposal for bridging application layer protocols to HTTP on IoT solutions. Future Generation Computer Systems, 97(??):145– 152, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18322763.

deCarvalho:2018:ESR

[dCTVC18] Juliana Oliveira de Carvalho, Fernando Trinta, Dario Vieira, and Omar Andres Carmona Cortes. Evolutionary solutions for resources management in multiple clouds: Stateof-the-art and future directions. Future Generation Computer Systems, 88(??):284–296, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18300025.

Dorronsoro:2014:SIE

[DDB14] Bernabé Dorronsoro, Grégoire Danoy, and Pascal Bouvry. Special issue: Energyefficiency in large distributed computing architectures. Future Generation Computer Systems, 36(??):187–188, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1400051X.

DeSensi:2018:SSA

[DDD18] Daniele De Sensi, Tiziano De Matteis, and Marco Danelutto. Simplifying self-adaptive and poweraware computing with Nornir. Future Generation Computer Systems, 87(??):136–151, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17326699.

DeFalco:2019:EBC

[DDD^{+}19] Ivanoe De Falco, Giuseppe De Pietro, Antonio Della Cioppa, Giovanna Sannino, Umberto Scafuri, and Ernesto Tarantino. Evolution-based configuration optimization of a deep neural network for the classification of obstructive sleep apnea episodes. Future Generation Computer Systems, 98(??):377– 391, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1832805X.

Dalman:2013:CMM

[DDJ^{+}13] Tolga Dalman, Tim Dörnemann, Ernst Juhnke, Michael Weitzel, Wolfgang Wiechert, Katharina Nöh, and Bernd Freisleben. Cloud MapReduce for Monte Carlo bootstrap applied to Metabolic Flux Analysis. Future Generation Computer Systems, 29(2):582–590, February 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001920.

Dodonov:2010:NAD

[DdM10] Evgueni Dodonov and Rodrigo Fernandes de Mello. A novel approach for distributed application scheduling based on prediction of communication events. Future Generation Computer Systems, 26(5): 740–752, May 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Diaz-Diaz:2017:BMA

[DDMPG17] Raimundo Díaz-Díaz, Luis Muñoz, and Daniel Pérez-González. Business model analysis of public services operating in the smart city ecosystem: the case of SmartSantander. Future Generation Computer Systems, 76(??):198–214, November 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301553.

Dourado:2019:NAM

[DdSdN^{+}19] Carlos M. J. M. Dourado, Suane P. P. da Silva, Raul V. M. da Nóbrega, Antonio C. S. Barros, Arun K. Sangaiah, Pedro P. Rebouças Filho, and Victor Hugo C. de Albuquerque. A new approach for mobile robot localization based on an online IoT system. Future Generation Computer Systems, 100(??):859–881, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18317400.

Diaz:2017:OAV

[DEG^{+}17] José Luis Díaz, Joaquín Entrialgo, Manuel García, Javier García, and Daniel Fernando García. Optimal allocation of virtual machines in multi-cloud environments with reserved and on-demand pricing. Future Generation Computer Systems, 71(??):129– 144, June 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301954.

DeOliveiraNunes:2019:SSC

[DEL19] Ivan De Oliveira Nunes, Karim Eldefrawy, and Tancrède Lepoint. SNUSE: a secure computation approach for large-scale user re-enrollment in biometric authentication systems. Future Generation Computer Systems, 98(??):259– 273, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1833098X.

deFarias:2017:CDE

[dFBP^{+}17] Claudio M. de Farias, Italo C. Brito, Luci Pirmez, Flávia C. Delicato, Paulo F. Pires, Taniro C.

Rodrigues, Igor L. dos Santos, Luiz F. R. C. Carmo, and Thais Batista. COM-FIT: a development environment for the Internet of Things. Future Generation Computer Systems, 75(??):128–144, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302163.

DeMaio:2019:TAA

[DFG^{+}19] Carmen De Maio, Giuseppe Fenza, Mariacristina Gallo, Vincenzo Loia, and Mimmo Parente. Time-aware adaptive tweets ranking through deep learning. Future Generation Computer Systems, 93(??):924–932, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17308087.

Dede:2014:BMI

[DFGR14] Elif Dede, Zacharia Fadika, Madhusudhan Govindaraju, and Lavanya Ramakrishnan. Benchmarking MapReduce implementations under different application scenarios. Future Generation Computer Systems, 36(??):389–399, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000065.

DeMaio:2017:USC

[DFLO17] Carmen De Maio, Giuseppe Fenza, Vincenzo Loia, and Francesco Orciuoli. Unfolding social content evolution along time and semantics. Future Generation Computer Systems, 66(??):146–159, January 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301649.

deFarias:2019:MSD

[dFPFG19] Claudio M. de Farias, Luci Pirmez, Giancarlo Fortino, and Antonio Guerrieri. A multi-sensor data fusion technique using data correlations among multiple applications. Future Generation Computer Systems, 92(??):109–118, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18307283.

Dessi:2017:SIA

[DFRW17] Nicoletta Dessì, Mariagrazia Fugini, Ismail Bouassida Rodriguez, and Usman Wajid. Special issue on advanced technologies enabling adaptive and collaborative smart systems. Future Generation Computer Systems, 68(??):361– 364, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302539.

Dogra:2018:ODS

[DGA18] Ayush Dogra, Bhawna Goyal, and Sunil Agrawal. Osseous and digital subtraction angiography image fusion via various enhancement schemes and Laplacian pyramid transformations. Future Generation Computer Systems, 82(??):149–157, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17320812.

Dodero:2017:TBI

[DGCGH^{+}17] Juan Manuel Dodero, Enrique Juan González-Conejero, Guillermo Gutiérrez-Herrera, Sonia Peinado, José Tomás Tocino, and Iván Ruiz-Rube. Tradeoff between interoperability and data collection performance when designing an architecture for learning analytics. Future Generation Computer Systems, 68(??):31–37, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302813.

Djorgovski:2016:RTD

[DGD^{+}16] S. G. Djorgovski, M. J. Graham, C. Donalek, A. A. Mahabal, A. J. Drake, M. Turmon, and T. Fuchs. Real-time data mining of massive data streams from synoptic sky surveys. Future Generation Computer Systems, 59(??):95–104, June 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1500326X.

Dumitru:2015:UCE

[DGdL15] Cosmin Dumitru, Paola Grosso, and Cees de Laat. A user-centric execution environment for CineGrid workloads. Future Generation Computer Systems, 53(??):55–62, December 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000837.

Diaz:2011:MUG

[DGGH11] Laura Díaz, Carlos Granell, Michael Gould, and Joaquín Huerta. Managing usergenerated information in geospatial cyberinfrastructures. Future Generation Computer Systems, 27(3): 304–314, March 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dias:2015:DCI

[DGR^{+}15] Jonas Dias, Gabriel Guerra, Fernando Rochinha, Alvaro L. G. A. Coutinho, Patrick Valduriez, and Marta Mattoso. Datacentric iteration in dynamic workflows. Future Generation Computer Systems, 46(??):114–126, May 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002155.

Din:2019:MLI

[DGR^{+}19] Ikram Ud Din, Mohsen Guizani, Joel J. P. C. Rodrigues, Suhaidi Hassan, and Valery V. Korotaev. Machine learning in the Internet of Things: Designed techniques for smart cities. Future Generation Computer Systems, 100(??):826–843, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19304030.

Dong:2018:LEN

[DGY^{+}18] Kai Dong, Taolin Guo, Haibo Ye, Xuansong Li, and Zhen Ling. On the limitations of existing notions of location privacy. Future Generation Computer Systems, 86(??):1513–1522, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1731052X.

Deng:2016:THS

[DH16] Shi-Wen Deng and Ji-Qing Han. Towards heart sound classification without segmentation via autocorrelation feature and diffusion maps. Future Generation Computer Systems, 60(??): 13–21, July 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16000121.

Du:2017:RCB

[DHC^{+}17] Zhihui Du, Ligang He, Yinong Chen, Yu Xiao, Peng Gao, and Tongzhou Wang. Robot Cloud: Bridging the power of robotics and cloud computing. Future Generation Computer Systems, 74(??):337–348, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16000042.

Ding:2018:ACM

[DHL18] Jianhao Ding, Lansheng Han, and Dan Li. An adaptive control momentum method as an optimizer in the cloud. Future Generation Computer Systems, 89(??):192–200, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17329989.

Deng:2017:EOD

[DHW^{+}17] Ze Deng, Wei Han, Lizhe Wang, Rajiv Ranjan, Albert Y. Zomaya, and Wei Jie. An efficient online direction-preserving compression approach for trajectory streaming data. Future Generation Computer Systems, 68(??):150–162, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303508.

Dukaric:2013:TUT

[DJ13] Robert Dukaric and Matjaz B. Juric. Towards a unified taxonomy and architecture of cloud frameworks. Future Generation Computer Systems, 29(5): 1196–1210, July 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001793.

Du:2019:ABD

[DJH^{+}19] Derong Du, Xin Jian, Long Hu, Yong Tan, Xiaoping Zeng, and Xiaoheng Tan. Angular beta distribution for 3D vehicle-to-vehicle channel modeling. Future Generation Computer Systems, 91(??):238–243, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18316832.

Du:2018:RDL

[DJJ^{+}18] Xinwu Du, Jiangtao Ji, Xin Jin, Can Li, and Xulong Yang. Research on divider losses with highspeed photography for foxtail millet harvesting. Future Generation Computer Systems, 88(??):55– 60, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18306368.

Deng:2018:SHO

[DJPM18] Xiaowu Deng, Peng Jiang, Xiaoning Peng, and Chunqiao Mi. Support highorder tensor data description for outlier detection in high-dimensional big sensor data. Future Generation Computer Systems, 81(??):177–187, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301528.

Dai:2015:TVD

[DJZ^{+}15] Weiqi Dai, Hai Jin, Deqing Zou, Shouhuai Xu, Weide Zheng, Lei Shi, and Laurence Tianruo Yang. TEE: a virtual DRTM based execution environment for secure cloud-end computing. Future Generation Computer Systems, 49(??):47– 57, August 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001538.

Duong:2014:FTR

[DK14] Dinh Thuy Duong and Keiichi Kaneko. Fault-tolerant routing based on approximate directed routable probabilities for hypercubes. Future Generation Computer Systems, 37(??): 88–96, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002677.

Deshpande:2017:TSL

[DK17] Umesh Deshpande and Kate Keahey. Trafficsensitive live migration of virtual machines. Future Generation Computer Systems, 72(??):118–128, July 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301133.

Doumbouya:2018:AGC

[DKFKF18] Mamadou Bilo Doumbouya, Bernard Kamsu-Foguem, Hugues Kenfack, and Clovis Foguem. Argumentation graphs with constraint-based reasoning for collaborative expertise. Future Generation Computer Systems, 81(??):16– 29, April 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17311706.

Dorri:2019:MBM

[DKJ19] Ali Dorri, Salil S. Kanhere, and Raja Jurdak. MOF-BC: A memory optimized and flexible blockchain for large scale networks. Future Generation Computer Systems, 92(??):357–373, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17329552.

Doh:2013:TGD

[DKK^{+}13] In Hwan Doh, Young Jin Kim, Eunsam Kim, Jongmoo Choi, Donghee Lee, and Sam H. Noh. Towards greener data centers with storage class memory. Future Generation Computer Systems, 29(8):1969– 1980, October 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001143.

Das:2014:CME

[DKV14] Anup Das, Akash Kumar, and Bharadwaj Veeravalli. Communication and migration energy aware task mapping for reliable multiprocessor systems. Future Generation Computer Systems, 30(??):216–228, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001301.

DeBenedictis:2019:IVD

[DL19] Marco De Benedictis and Antonio Lioy. Integrity verification of Docker containers for a lightweight cloud environment. Future Generation Computer Systems, 97(??):236–246, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18327201.

deLlano:2010:SNN

[dLB10] Rafael Menéndez de Llano and José Luis Bosque. Study of neural net training methods in parallel and distributed architectures. Future Generation Computer Systems, 26(2):267– 275, February 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Diaz-Lopez:2016:DCM

[DLDTGMMP16] Daniel Díaz-López, Ginés Dólera-Tormo, Félix Gómez-Mármol, and Gregorio Martínez-Pérez. Dynamic counter-measures for riskbased access control systems: an evolutive approach. Future Generation Computer Systems, 55(??):321–335, February 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002052.

de-la-Fuente-Valentin:2014:TSE

[dlFVPSHL^{+}14] Luis de-la Fuente-Valentín, Mar Pérez-Sanagustín, Davinia Hernández-Leo, Abelardo Pardo, Josep Blat, and Carlos Delgado Kloos. Technological support for the enactment of collaborative scripted learning activities across multiple spatial locations. Future Generation Computer Systems, 31(??):223–237, February 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001064.

Dong:2017:EED

[DLH^{+}17] Xinhua Dong, Ruixuan Li, Heng He, Xiwu Gu, Mudar Sarem, Meikang Qiu, and Keqin Li. EDS: an efficient data selection policy for search engine storage architectures. Future Generation Computer Systems, 74(??):220–231, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300279.

Ding:2019:SEG

[DLL^{+}19] Shuai Ding, Ling Li, Zhenmin Li, Hao Wang, and Yanchun Zhang. Smart electronic gastroscope system using a cloud-edge collaborative framework. Future Generation Computer Systems, 100(??):395–407, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18324324.

Deng:2017:MUS

[DLLZ17] Zuojie Deng, Kenli Li, Keqin Li, and Jingli Zhou. A multi-user searchable encryption scheme with keyword authorization in a cloud storage. Future Generation Computer Systems, 72(??):208–218, July 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301285.

DiPietro:2015:ATC

[DLMS15] Roberto Di Pietro, Flavio Lombardi, Fabio Martinelli, and Daniele Sgandurra. AntiCheetah: Trustworthy computing in an outsourced (cheating) environment. Future Generation Computer Systems, 48(??):ibc, July 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000394.

Dong:2012:EDA

[DLS^{+}12] Fang Dong, Junzhou Luo, Aibo Song, Jiuxin Cao, and Jun Shen. An effective data aggregation based adaptive long term CPU load prediction mechanism on computational grid. Future Generation Computer Systems, 28(7):1030–1044, July 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002202.

DiPietro:2014:CCR

[DLS14] Roberto Di Pietro, Flavio Lombardi, and Matteo Signorini. CloRExPa: Cloud resilience via execution path analysis. Future Generation Computer Systems, 32(??):168–179, March 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001069.

Dong:2014:TMD

[DLXR14] Bin Dong, Xiuqiao Li, Limin Xiao, and Li Ruan. Towards minimizing disk I/O contention: a partitioned file assignment approach. Future Generation Computer Systems, 37(??): 178–190, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002860.

Deng:2014:MAP

[DLZ^{+}14] Yuhui Deng, Lijuan Lu, Qiang Zou, Shuqiang Huang, and Jipeng Zhou. Modeling the aging process of flash storage by leveraging semantic I/O. Future Generation Computer Systems, 32(??):338– 344, March 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001817.

Dai:2016:MLR

[DLZ16] Shuguang Dai, Huige Li, and Fangguo Zhang. Memory leakage-resilient searchable symmetric encryption. Future Generation Computer Systems, 62(??):76–84, September 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003404.

Doulamis:2012:VUI

[DM12] Anastasios Doulamis and Nikolaos Matsatsinis. Visual understanding industrial workflows under uncertainty on distributed service oriented architectures. Future Generation Computer Systems, 28(3): 605–617, March 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000252.

DAgostino:2019:CEC

[DMC^{+}19] Daniele D’Agostino, Lucia Morganti, Elena Corni, Daniele Cesini, and Ivan Merelli. Combining edge and cloud computing for low-power, cost-effective metagenomics analysis. Future Generation Computer Systems, 90(??):79– 85, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18300293.

Drozdowski:2014:ETO

[DMM14] Maciej Drozdowski, Jedrzej M. Marszalkowski, and Jakub Marszalkowski. Energy trade-offs analysis using equal-energy maps. Future Generation Computer Systems, 36(??):311– 321, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001490.

Dini:2018:RAA

[DMM^{+}18] Gianluca Dini, Fabio Martinelli, Ilaria Matteucci, Marinella Petrocchi, Andrea Saracino, and Daniele Sgandurra. Risk analysis of Android applications: a user-centric solution. Future Generation Computer Systems, 80(??):505–518, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301534.

Ding:2011:LPD

[DMMM11] Li Ding, James Michaelis, Jim McCusker, and Deborah L. McGuinness. Linked provenance data: a semantic Web-based approach to interoperable workflow traces. Future Generation Computer Systems, 27(6): 797–805, June 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dessi:2016:CCB

[DMPP16] Nicoletta Dessì, Gabriele Milia, Emanuele Pascariello, and Barbara Pes. COWB: a cloud-based framework supporting collaborative knowledge management within biomedical communities. Future Generation Computer Systems, 54(??):399–408, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001041.

Diamantini:2019:SID

[DMPS19] Claudia Diamantini, Alex Mircoli, Domenico Potena, and Emanuele Storti. Social information discovery enhanced by sentiment analysis techniques. Future Generation Computer Systems, 95(??):816–828, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17308634.

Deng:2012:SSB

[DMZ12] Yuhui Deng, Xiaohua Meng, and Jipeng Zhou. Self-similarity: Behind workload reshaping and prediction. Future Generation Computer Systems, 28(2):350–357, February 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1100121X.

Dias:2017:MGB

[DNJG17] Sérgio E. D. Dias, Quoc T.

Nguyen, Joaquim A. Jorge, and Abel J. P. Gomes. Multi-GPU-based detection of protein cavities using critical points. Future Generation Computer Systems, 67(??):430–440, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302394.

Durillo:2014:MOE

[DNP14] Juan J. Durillo, Vlad Nae, and Radu Prodan. Multiobjective energy-efficient workflow scheduling using list-based heuristics. Future Generation Computer Systems, 36(??):221–236, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001507.

Dongdong:2019:SCF

[DNW^{+}19] Jiao Dongdong, Arunkumar N., Zhang Wenyu, Li Beibei, Zhang Xinlei, and Zhu Guangjian. Semantic clustering fuzzy c means spectral model based comparative analysis of cardiac color ultrasound and electrocardiogram in patients with left ventricular heart failure and cardiomyopathy. Future Generation Computer Systems, 92(??):324– 328, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18322192.

Drira:2015:SIA

[DO15] Khalil Drira and Flavio Oquendo. Special issue on advanced architectures for the future generation of software-intensive systems. Future Generation Computer Systems, 47(??): 60–61, June 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000412.

deOliveira:2013:PEP

[dOOO^{+}13] Daniel de Oliveira, Kary A. C. S. Ocaña, Eduardo Ogasawara, Jonas Dias, João Gonçalves, Fernanda Baião, and Marta Mattoso. Performance evaluation of parallel strategies in public clouds: a study with phylogenomic workflows. Future Generation Computer Systems, 29(7):1816–1825, September 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000034.

Werneck:2018:KUF

[dOWdAS^{+}18] Rafael de Oliveira Werneck, Waldir Rodrigues de Almeida, Bernardo Vecchia Stein, Daniel Vatanabe Pazinato, Pedro Ribeiro Mendes Júnior, Otávio Augusto Bizetto Penatti, Anderson Rocha, and Ricardo da Silva Torres. Kuaa: a unified framework for design, deployment, execution, and recommendation of machine learning experiments. Future Generation Computer Systems, 78 (part 1)(??): 59–76, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301565.

Dessi:2017:SSA

[DP17] Nicoletta Dessì and Barbara Pes. Smart spaces for adaptive information integration in bioinformatics. Future Generation Computer Systems, 68(??):407– 415, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302345.

Din:2019:SHM

[DP19] Sadia Din and Anand Paul. Smart health monitoring and management system: Toward autonomous wearable sensing for Internet of Things using big data analytics. Future Generation Computer Systems, 91(??):611–619, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17315078.

See erratum [DP20] and retraction notice [DP21b].

Din:2020:ESH

[DP20] Sadia Din and Anand Paul.

Erratum to “Smart Health Monitoring and Management System: Toward autonomous wearable sensing for Internet of Things using Big Data Analytics” [Future Gener. Comput. Syst. 91 (2019) 611–619]. Future Generation Computer Systems, 111(??):939, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19317005.

See [DP19] and retraction notice [DP21a].

Din:2021:RNE

[DP21a] Sadia Din and Anand Paul.

Retraction notice to Erratum to “Smart Health Monitoring and Management System: Toward autonomous wearable sensing for Internet of Things using Big Data Analytics [Future Gener. Comput. Syst. 91 (2019) 611–619]”’ [Future Gener. Comput. Syst. 108 (2019) 1350–1359]. Future Generation Computer Systems, 124(??):496, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002600.

See [DP20].

Din:2021:RNS

[DP21b] Sadia Din and Anand Paul. Retraction notice to “Smart Health Monitoring and Management System: Toward autonomous wearable sensing for Internet of Things using Big Data Analytics” [Future Gener. Comput. Syst. 91 (2019) 611–619]. Future Generation Computer Systems, 124(??):495, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002612.

See [DP19].

DeMaio:2016:MEC

[DPBK16] Vincenzo De Maio, Radu Prodan, Shajulin Benedict, and Gabor Kecskemeti. Modelling energy consumption of network transfers and virtual machine migration. Future Generation Computer Systems, 56(??):388–406, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002307.

Divoli:2014:SIA

[DPDS14] Anna Divoli, Domenico Potena, Claudia Diamantini, and Waleed W. Smari. Special issue on advances in computer supported collaboration: Systems and technologies. Future Generation Computer Systems, 31(??):105–110, February 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002549.

Dao:2019:RAR

[DPK^{+}19] Nhu-Ngoc Dao, Minho Park, Joongheon Kim, Jeongyeup Paek, and Sungrae Cho. Resource-aware relay selection for intercell interference avoidance in 5G heterogeneous network for Internet of Things systems. Future Generation Computer Systems, 93(??):877–887, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17304028.

Duan:2014:SCG

[DPL14] Rubing Duan, Radu Prodan, and Xiaorong Li. A sequential cooperative game theoretic approach to scheduling multiple largescale applications in grids.

Future Generation Computer Systems, 30(??):27– 43, January 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001805.

Diamantini:2016:SSF

[DPS16] Claudia Diamantini, Domenico Potena, and Emanuele Storti. SemPI: a semantic framework for the collaborative construction and maintenance of a shared dictionary of performance indicators. Future Generation Computer Systems, 54(??):352–365, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500103X.

DAgostino:2019:SBC

[DQC^{+}19] Daniele D’Agostino, Alfonso Quarati, Andrea Clematis, Lucia Morganti, Elena Corni, Valentina Giansanti, Daniele Cesini, and Ivan Merelli. SoCbased computing infrastructures for scientific applications and commercial services: Performance and economic evaluations. Future Generation Computer Systems, 96(??):11– 22, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18311622.

Ding:2015:EES

[DQLW15] Youwei Ding, Xiaolin Qin, Liang Liu, and Taochun Wang. Energy efficient scheduling of virtual machines in cloud with deadline constraint. Future Generation Computer Systems, 50(??):62– 74, September 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000369.

Ding:2019:LVC

[DQXW19] Songtao Ding, Shiru Qu, Yuling Xi, and Shaohua Wan. A long video caption generation algorithm for big video data retrieval. Future Generation Computer Systems, 93(??):583– 595, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18319782.

Diegues:2015:BSD

[DR15] Nuno Diegues and Paolo Romano. Bumper: Sheltering distributed transactions from conflicts. Future Generation Computer Systems, 51(??):20–35, October 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000941.

Domanal:2018:ECO

[DR18] Shridhar G. Domanal and G. Ram Mohana Reddy. An efficient cost optimized scheduling for spot instances in heterogeneous cloud environment. Future Generation Computer Systems, 84(??):11–21, July 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17303667.

Astorga:2018:PWT

[dRADFG18] David del Rio Astorga, Manuel F. Dolz, Javier Fernández, and J. Daniel García. Paving the way towards high-level parallel pattern interfaces for data stream processing. Future Generation Computer Systems, 87(??):228–241, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17324524.

Das:2019:DML

[DRC^{+}19] Arun Das, Paul Rad, Kim-Kwang Raymond Choo, Babak Nouhi, Jonathan Lish, and James Martel. Distributed machine learning cloud teleophthalmology IoT for predicting AMD disease progression. Future Generation Computer Systems, 93(??):486– 498, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18317941.

DelaPrieta:2019:SAB

[DRGC^{+}19] Fernando De la Prieta, Sara Rodríguez-González, Pablo Chamoso, Juan Manuel Corchado, and Javier Bajo. Survey of agent-based cloud computing applications. Future Generation Computer Systems, 100(??):223–236, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302717.

Righi:2018:MTS

[dRRdQGR^{+}18] Rodrigo da Rosa Righi, Roberto de Quadros Gomes, Vinicius Facco Rodrigues, Cristiano André da Costa, Antonio Marcos Alberti, Laércio Lima Pilla, and Philippe Olivier Alexandre Navaux. MigPF: Towards on self-organizing process rescheduling of bulk-synchronous parallel applications. Future Generation Computer Systems, 78 (part 1)(??):272–286, January 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301145.

Righi:2018:LPP

[dRRRR^{+}18] Rodrigo da Rosa Righi, Vinicius Facco Rodrigues, Gustavo Rostirolla, Cristiano André da Costa, Eduardo Roloff, and Philippe Olivier Alexandre Navaux. A lightweight plug-and-play elasticity service for self-organizing resource provisioning on parallel applications. Future Generation Computer Systems, 78 (part 1)(??):176–190, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302339.

DAgostino:2019:SGE

[DRZ^{+}19] Daniele D’Agostino, Luca Roverelli, Gabriele Zereik, Giuseppe La Rocca, Andrea De Luca, Ruben Salvaterra, Andrea Belfiore, Gianni Lisini, Giovanni Novara, and Andrea Tiengo. A science gateway for exploring the X-ray transient and variable sky using EGI Federated Cloud. Future Generation Computer Systems, 94(??):868– 878, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310051.

Debroy:2019:SIS

[DSBC19] Saptarshi Debroy, Priyanka Samanta, Amina Bashir, and Mainak Chatterjee. SpEED-IoT: Spectrum aware energy efficient routing for device-to-device IoT communication. Future Generation Computer Systems, 93(??):833–848, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17307148.

daSilva:2019:CLM

[dSBN19] Helber Wagner da Silva, Felipe Rocha Barbalho, and Augusto Venâncio Neto. Cross-layer multiuser session control for optimized communications on SDN-based cloud platforms. Future Generation Computer Systems, 92(??):1116–1130, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17306453.

daSilva:2019:MIB

[dSCD^{+}19] Rafael Ferreira da Silva, Scott Callaghan, Tu Mai Anh Do, George Papadimitriou, and Ewa Deelman. Measuring the impact of burst buffers on data-intensive scientific workflows. Future Generation Computer Systems, 101(??):208–220, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18308100.

Dumic:2018:EPR

[DSCJ18] Mateja Dumić, Dominik Sisejković, Rebeka Corić, and Domagoj Jakobović. Evolving priority rules for resource constrained project scheduling problem with genetic programming. Future Generation Computer Systems, 86(??):211–221, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1732441X.

Deboosere:2011:GDM

[DSD^{+}11] L. Deboosere, P. Simoens, J. De Wachter, B. Vankeirsbilck, F. De Turck, B. Dhoedt, and P. Demeester. Grid design for mobile thin client computing. Future Generation Computer Systems, 27(6):681–693, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

daSilva:2019:USP

[dSFD^{+}19] Rafael Ferreira da Silva, Rosa Filgueira, Ewa Deelman, Erola Pairo-Castineira, Ian M. Overton, and Malcolm P. Atkinson. Using simple PID-inspired controllers for online resilient resource management of distributed scientific workflows. Future Generation Computer Systems, 95(??): 615–628, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300481.

daSilva:2017:CWM

[dSFP^{+}17] Rafael Ferreira da Silva, Rosa Filgueira, Ilia Pietri, Ming Jiang, Rizos Sakellariou, and Ewa Deelman. A characterization of workflow management systems for extremescale applications. Future Generation Computer Systems, 75(??):228–238, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302510.

daSilva:2013:SHW

[dSGD13] Rafael Ferreira da Silva, Tristan Glatard, and Frédéric Desprez. Self-healing of workflow activity incidents on distributed computing infrastructures. Future Generation Computer Systems, 29(8):2284–2294, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1300126X.

daSilva:2019:CMQ

[dSGD19] Madalena Pereira da Silva, Alexandre Leopoldo Gonçalves, and Mário Antônio Ribeiro Dantas. A conceptual model for quality of experience management to provide context-aware eHealth services. Future Generation Computer Systems, 101(??):1041–1061, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18314079.

Dabbagh:2014:FDI

[DSK^{+}14] Mehiar Dabbagh, Naoum Sayegh, Ayman Kayssi, Imad Elhajj, and Ali Chehab. Fast dynamic Internet mapping. Future Generation Computer Systems, 39(??):55–66, October 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000818.

deMDelEsposte:2019:DES

[dSK^{+}19] Arthur de M.Del Esposte, Eduardo F. Z. Santana, Lucas Kanashiro, Fabio M. Costa, Kelly R. Braghetto, Nelson Lago, and Fabio Kon. Design and evaluation of a scalable smart city software platform with large-scale simulations. Future Generation Computer Systems, 93(??):427–441, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18307301.

Devi:2019:HMI

[DSM^{+}19] Swagatika Devi, Manmath Narayan Sahoo, Khan Muhammad, Weiping Ding, and Sambit Bakshi. Hiding medical information in brain MR images without affecting accuracy of classifying pathological brain. Future Generation Computer Systems, 99(??):235–246, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18322647.

Machado:2017:EHC

[dSMAdR^{+}17] Roger da Silva Machado, Ricardo Borges Almeida, Diórgenes Yuri Leal da Rosa, João Ladislau Barbará Lopes, Ana Marilza Pernas, and Adenauer Corrêa Yamin. EXEHDA–HM: a compositional approach to explore contextual information on hybrid models. Future Generation Computer Systems, 73(??):1– 12, August 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630509X.

Douzis:2018:MGI

[DSPA18] Konstantinos Douzis, Stelios Sotiriadis, Euripides G. M. Petrakis, and Cristiana Amza. Modular and generic IoT management on the cloud. Future Generation Computer Systems, 78 (part 1)(??):369–378, January 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301662.

DUrso:2019:WSS

[DSS19] Fabio D’Urso, Corrado Santoro, and Federico Fausto Santoro. Wale: a solution to share libraries in Docker containers. Future Generation Computer Systems, 100(??):513–522, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18327511.

Soares:2019:OTM

[dSSCdL19] Elton F. de S. Soares, Carlos Alberto V. Campos, and Sidney C. de Lucena. Online travel mode detection method using automated machine learning and feature engineering. Future Generation Computer Systems, 101 (??):1201–1212, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19305874.

DeFalco:2010:AMM

[DST10] Ivanoe De Falco, Umberto Scafuri, and Ernesto Tarantino. An adaptive multisite mapping for computationally intensive grid applications. Future Generation Computer Systems, 26(6):857–867, June 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DeFalco:2014:TNF

[DST14] I. De Falco, U. Scafuri, and E. Tarantino. Two new fast heuristics for mapping parallel applications on cloud computing. Future Generation Computer Systems, 37 (??):1–13, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000442.

DiModica:2016:MSS

[DT16] Giuseppe Di Modica and Orazio Tomarchio. Matchmaking semantic security policies in heterogeneous clouds. Future Generation Computer Systems, 55(??):176–185, February 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000631.

Devi:2013:SAA

[DV13] M. K. Kavitha Devi and P. Venkatesh. Smoothing approach to alleviate the meager rating problem in collaborative recommender systems. Future Generation Computer Systems, 29(1):262–270, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000938.

Depoorter:2014:ARC

[DVB14] W. Depoorter, K. Vanmechelen, and J. Broeckhove. Advance reservation, co-allocation and pricing of network and computational resources in grids. Future Generation Computer Systems, 41(??):1– 15, December 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001368.

Desmet:2012:DSO

[DVD12] Stein Desmet, Bruno Volckaert, and Filip De Turck. Design of a service oriented architecture for efficient resource allocation in media environments. Future Generation Computer Systems, 28(3):527– 532, March 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000264.

Deelman:2015:PWM

[DVJ^{+}15] Ewa Deelman, Karan Vahi, Gideon Juve, Mats Rynge, Scott Callaghan, Philip J. Maechling, Rajiv Mayani, Weiwei Chen, Rafael Ferreira da Silva, Miron Livny, and Kent Wenger. Pegasus, a workflow management system for science automation. Future Generation Computer Systems, 46 (??):17–35, May 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002015.

deVrieze:2011:BEM

[dVXB^{+}11] Paul de Vrieze, Lai Xu, Athman Bouguettaya, Jian Yang, and Jinjun Chen. Building enterprise mashups. Future Generation Computer Systems, 27(5):637– 642, May 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Deng:2011:LAT

[DW11] Yuhui Deng and Frank Wang. LAG: Achieving transparent access to legacy data by leveraging grid environment. Future Generation Computer Systems, 27(1):32–39, January 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dai:2018:GQQ

[DWJM18] Wei Dai, Yufeng Wang, Qun Jin, and Jianhua Ma. Geo-QTI: A quality aware truthful incentive mechanism for cyber physical enabled geographic crowdsensing. Future Generation Computer Systems, 79 (part 1)(??): 447–459, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17307070.

Dougherty:2012:MDA

[DWS12] Brian Dougherty, Jules White, and Douglas C. Schmidt. Model-driven auto-scaling of green cloud computing infrastructure. Future Generation Computer Systems, 28(2):371– 378, February 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000902.

Dobre:2014:ISB

[DX14] C. Dobre and F. Xhafa.

Intelligent services for Big Data science. Future Generation Computer Systems, 37(??):267–281, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001593.

Deng:2014:CCC

[DXA14] Robert H. Deng, Yang Xiang, and Man Ho Au. Cryptography in cloud computing. Future Generation Computer Systems, 30 (??):90, January 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002197.

Dou:2018:RCA

[DXL^{+}18] Wanchun Dou, Xiaolong Xu, Xiang Liu, Laurence T. Yang, and Yiping Wen. A resource coallocation method for loadbalance scheduling over big data platforms. Future Generation Computer Systems, 86(??):1064–1075, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17314735.

Du:2018:MIM

[DYC^{+}18] Zhanwei Du, Yongjian Yang, Qing Cai, Chijun Zhang, and Yuan Bai. Modeling and inferring mobile phone users’ negative emotion spreading in social networks. Future Generation Computer Systems, 78 (part 3)(??):933–942, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17305939.

Dorier:2016:EFM

[DYI^{+}16] Matthieu Dorier, Orcun Yildiz, Shadi Ibrahim, Anne-Cécile Orgerie, and Gabriel Antoniu. On the energy footprint of I/O management in Exascale HPC systems. Future Generation Computer Systems, 62(??):17–28, September 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300462.

Deng:2019:DRA

[DYY^{+}19] Song Deng, Changan Yuan, Lechan Yang, Xiao Qin, and Aihua Zhou. Data recovery algorithm under intrusion attack for Energy Internet. Future Generation Computer Systems, 100(??):109–121, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18310276.

Das:2018:TAS

[DZH18] Ashok Kumar Das, Sherali Zeadally, and Debiao He. Taxonomy and analysis of security protocols for Internet of Things. Future Generation Computer Systems, 89(??):110–125, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18308112.

Du:2019:LMB

[DZLA19] Xiaokun Du, Rongbo Zhu, Yanhong Li, and Ashiq Anjum. Language modelbased automatic prefix abbreviation expansion method for biomedical big data analysis. Future Generation Computer Systems, 98(??):238–251, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18326529.

Dai:2015:SEC

[DZZ^{+}15] Hongjun Dai, Shulin Zhao, Jiutian Zhang, Meikang Qiu, and Lixin Tao. Security enhancement of cloud servers with a redundancybased fault-tolerant cache structure. Future Generation Computer Systems, 52(??):147–155, November 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000540.

Exarchakos:2013:CST

[EA13] Georgios Exarchakos and Nick Antonopoulos. Cooperative stalking of transient nomadic resources on overlay networks. Future Generation Computer Systems, 29(6):1473–1484, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12002282.

Ezugwu:2017:SSB

[EA17] Absalom E. Ezugwu and Aderemi O. Adewumi. Soft sets based symbiotic organisms search algorithm for resource discovery in cloud computing environment. Future Generation Computer Systems, 76(??):33– 50, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301838.

El-Alfy:2016:SFF

[EAA16] El-Sayed M. El-Alfy and Ali A. AlHasan. Spam filtering framework for multimodal mobile communication based on dendritic cell algorithm. Future Generation Computer Systems, 64(??):98–107, November 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300310.

Elshenawy:2018:TSO

[EAED18] Mohamed Elshenawy, Baher Abdulhai, and Mohamed El-Darieby. Towards a service-oriented cyber physical systems of systems for smart city mobility applications. Future Generation Computer Systems, 79 (part 2)(??):575– 587, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17307471.

Elhoseny:2018:HMI

[EAS^{+}18] Mohamed Elhoseny, Ahmed Abdelaziz, Ahmed S. Salama, A. M. Riad, Khan Muhammad, and Arun Kumar Sangaiah. A hybrid model of Internet of Things and cloud computing to manage big data in health services applications. Future Generation Computer Systems, 86(??):1383–1394, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17322021.

Esposito:2018:OTP

[EBCP18] Christian Esposito, Andrea Bruno, Giuseppe Cattaneo, and Francesco Palmieri. On the optimal tuning and placement of FEC codecs within multicasting trees for resilient publish/subscribe services in edge-IoT architectures. Future Generation Computer Systems, 88(??):140– 150, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18301912.

Elayyadi:2014:TBD

[EBOY14] Isam Elayyadi, Salima Benbernou, Mourad Ouziri, and Muhammad Younas. A tensor-based distributed discovery of missing association rules on the cloud. Future Generation Computer Systems, 35(??):49– 56, June 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002628.

Evangelinou:2018:EAC

[ECA^{+}18] Athanasia Evangelinou, Michele Ciavotta, Danilo Ardagna, Aliki Kopaneli, George Kousiouris, and Theodora Varvarigou. Enterprise applications cloud rightsizing through a joint benchmarking and optimization approach. Future Generation Computer Systems, 78 (part 1)(??):102– 114, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630512X.

ElKassem:2019:MEP

[ECE^{+}19] Nada El Kassem, Liqun Chen, Rachid El Bansarkhani, Ali El Kaafarani, Jan Camenisch, Patrick Hough, Paulo Martins, and Leonel Sousa. More efficient, provably-secure direct anonymous attestation from lattices. Future Generation Computer Systems, 99(??):425–458, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19300536.

Esposito:2017:IGE

[ECPF17a] Christian Esposito, Aniello Castiglione, Francesco Palmieri, and Massimo Ficco. Improving the gossiping effectiveness with distributed strategic learning (invited paper). Future Generation Computer Systems, 71(??): 221–233, June 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16305520.

Esposito:2017:TMD

[ECPF17b] Christian Esposito, Aniello Castiglione, Francesco Palmieri, and Massimo Ficco. Trust management for distributed heterogeneous systems by using linguistic term sets and hierarchies, aggregation operators and mechanism design. Future Generation Computer Systems, 74(??):325–336, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003878.

Esiner:2016:AVD

[ED16] Ertem Esiner and Anwitaman Datta. Auditable versioned data storage outsourcing. Future Generation Computer Systems, 55(??):17–28, February 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002538.

Esiner:2019:TFA

[ED19] Ertem Esiner and Anwitaman Datta. Twofactor authentication for trusted third party free dispersed storage. Future Generation Computer Systems, 90(??):291–306, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17322859.

Eavis:2010:POS

[EDD^{+}10] Todd Eavis, George Dimitrov, Ivan Dimitrov, David Cueva, Alex Lopez, and Ahmad Taleb. Parallel OLAP with the Sidera server. Future Generation Computer Systems, 26(2): 259–266, February 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Esteban:2013:DAE

[EDH^{+}13] Francisco J. Esteban, David Díaz, Pilar Hernández, Juan A. Caballero, Gabriel Dorado, and Sergio Gálvez. Direct approaches to exploit many-core architecture in bioinformatics. Future Generation Computer Systems, 29(1):15–26, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000830.

Enes:2018:BRT

[EET18] Jonatan Enes, Roberto R. Expósito, and Juan Touriño. BDWatchdog: Real-time monitoring and profiling of big data applications and frameworks. Future Generation Computer Systems, 87(??):420–437, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17316096.

Engelsberger:2018:DRS

[EG18] Maximilian Engelsberger and Thomas Greiner. Dynamic reconfiguration of service-oriented resources in cyber-physical production systems by a processindependent approach with multiple criteria and multiple resource management operations. Future Generation Computer Systems, 88(??):424–441, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17320538.

Eirinaki:2018:RSL

[EGVT18] Magdalini Eirinaki, Jerry Gao, Iraklis Varlamis, and Konstantinos Tserpes. Recommender systems for large-scale social networks: a review of challenges and solutions. Future Generation Computer Systems, 78 (part 1)(??):413–418, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17319684.

Elmroth:2010:DUL

[EH10] Erik Elmroth and Daniel Henriksson. Distributed usage logging for Federated Grids. Future Generation Computer Systems, 26(8): 1215–1225, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Elmroth:2010:TFD

[EHT10] Erik Elmroth, Francisco Hernández, and Johan Tordsson. Three fundamental dimensions of scientific workflow interoperability: Model of computation, language, and execution environment. Future Generation Computer Systems, 26(2):245–256, February 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ebadi:2011:NDH

[EK11] Saeed Ebadi and Leyli Mohammad Khanli. A new distributed and hierarchical mechanism for service discovery in a grid environment. Future Generation Computer Systems, 27(6):836–842, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Evers:2014:ULE

[EKGS14] Christoph Evers, Romy Kniewel, Kurt Geihs, and Ludger Schmidt. The user in the loop: Enabling user participation for selfadaptive applications. Future Generation Computer Systems, 34(??):110–123, May 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002744.

El-Kassabi:2019:TET

[EKSDN19] Hadeel T. El-Kassabi, M. Adel Serhani, Rachida Dssouli, and Alramzana N. Navaz. Trust enforcement through self-adapting cloud workflow orchestration. Future Generation Computer Systems, 97(??): 462–481, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18313529.

EL-Latif:2019:EQB

[ELAEAVAM19] Ahmed A. Abd EL-Latif, Bassem Abd-El-Atty, Salvador E. Venegas-Andraca, and Wojciech Mazurczyk. Efficient quantum-based security protocols for information sharing and data protection in 5G networks. Future Generation Computer Systems, 100(??):893–906, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19307010.

Eskandari:2018:TST

[EMHE18] Leila Eskandari, Jason Mair, Zhiyi Huang, and David Eyers. T3-Scheduler: a Topology and Traffic aware two-level Scheduler for stream processing systems in a heterogeneous cluster. Future Generation Computer Systems, 89(??):617–632, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17326432.

Espadas:2013:TBR

[EMJ^{+}13] Javier Espadas, Arturo Molina, Guillermo Jiménez, Martín Molina, Raúl Ramírez, and David Concha. A tenant-based resource allocation model for scaling Software-as-a-Service applications over cloud computing infrastructures. Future Generation Computer Systems, 29(1):273–286, January 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1100210X.

Entezari-Maleki:2012:PTS

[EMM12] Reza Entezari-Maleki and Ali Movaghar. A probabilistic task scheduling method for grid environments. Future Generation Computer Systems, 28(3): 513–524, March 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001713.

Emeakaroha:2012:TAD

[ENC^{+}12] Vincent C. Emeakaroha, Marco A. S. Netto, Rodrigo N. Calheiros, Ivona Brandic, Rajkumar Buyya, and César A. F. De Rose. Towards autonomic detection of SLA violations in Cloud infrastructures. Future Generation Computer Systems, 28(7):1017–1029, July 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002184.

Engelmann:2014:SMC

[Eng14] Christian Engelmann. Scaling to a million cores and beyond: Using light-weight simulation to understand the challenges ahead on the road to exascale. Future Generation Computer Systems, 30(??):59–65, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000745.

Ericson:2012:AHL

[EP12] Kathleen Ericson and Shrideep Pallickara. Adaptive heterogeneous language support within a cloud runtime. Future Generation Computer Systems, 28(1):128–135, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1100094X.

Ericson:2013:PHD

[EP13] Kathleen Ericson and Shrideep Pallickara. On the performance of high dimensional data clustering and classification algorithms. Future Generation Computer Systems, 29(4): 1024–1034, June 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001343.

Evangelidis:2018:PMV

[EPB18] Alexandros Evangelidis, David Parker, and Rami Bahsoon. Performance modelling and verification of cloud-based autoscaling policies. Future Generation Computer Systems, 87(??):629–638, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312475.

Erdil:2013:ACR

[Erd13] D. Cenk Erdil. Autonomic cloud resource sharing for intercloud federations. Future Generation Computer Systems, 29(7):1700–1708, September 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000908.

Erickson:2017:IRA

[ESPN17] Alejandro Erickson, Iain A.

Stewart, Jose A. Pascual, and Javier Navaridas. Improved routing algorithms in the dual-port datacenter networks HCN and BCN. Future Generation Computer Systems, 75(??):58– 71, October 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300134.

Ebrahimi:2017:AMH

[ESW^{+}17] Mohammad Ebrahimi, Elaheh ShafieiBavani, Raymond K. Wong, Simon Fong, and Jinan Fiaidhi. An adaptive meta-heuristic search for the Internet of Things. Future Generation Computer Systems, 76(??):486–494, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003908.

Exposito:2013:PAH

[ETR^{+}13] Roberto R. Expósito, Guillermo L. Taboada, Sabela Ramos, Juan Touriño, and Ramón Doallo. Performance analysis of HPC applications in the cloud. Future Generation Computer Systems, 29(1):218– 229, January 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001458.

Elmahmudi:2019:DFR

[EU19] Ali Elmahmudi and Hassan Ugail. Deep face recognition using imperfect facial data. Future Generation Computer Systems, 99(??):213–225, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18331133.

Espada:2019:CCS

[EYY19] Jordán Pascual Espada, Ronald Yager, and Zhiyong Yu. Communications, collaborations and services in networks of embedded devices. Future Generation Computer Systems, 92(??):560–563, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1832853X.

Elgendy:2019:RAC

[EZTL19] Ibrahim A. Elgendy, Weizhe Zhang, Yu-Chu Tian, and Keqin Li. Resource allocation and computation offloading with data security for mobile edge computing. Future Generation Computer Systems, 100(??):531–541, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18328346.

Fiore:2011:SSD

[FA11a] S. Fiore and G. Aloisio.

Special section: Data management for eScience. Future Generation Computer Systems, 27(3):290–291, March 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fudzee:2011:QBA

[FA11b] Mohd Farhan Md Fudzee and Jemal H. Abawajy.

QoS-based adaptation service selection broker. Future Generation Computer Systems, 27(3):256–264, March 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fahd:2018:CPA

[FAA^{+}18] Shah Fahd, Mehreen Afzal, Haider Abbas, Waseem Iqbal, and Salman Waheed. Correlation power analysis of modes of encryption in AES and its countermeasures. Future Generation Computer Systems, 83(??):496–509, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17311822.

Feng:2019:SGB

[FAL^{+}19] Li Feng, Amjad Ali, Hannan Bin Liaqat, Muhammad Aksam Iftikhar, Ali Kashif Bashir, and Sangheon Pack. Stochastic gamebased dynamic information delivery system for wireless cooperative networks. Future Generation Computer Systems, 95(??):277– 291, June 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18318417.

Fernandez-Ares:2017:SRT

[FAMA^{+}17] A. Fernández-Ares, A. M.

Mora, M. G. Arenas, P. García-Sanchez, G. Romero, V. Rivas, P. A. Castillo, and J. J. Merelo. Studying real traffic and mobility scenarios for a Smart City using a new monitoring and tracking system. Future Generation Computer Systems, 76(??):163– 179, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306604.

Farkiani:2019:SVN

[FBM19] Behrooz Farkiani, Bahador Bakhshi, and S. Ali MirHassani. Stochastic virtual network embedding via accelerated Benders decomposition. Future Generation Computer Systems, 94(??):199–213, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317163.

Fiore:2018:REA

[FBS18] Sandro Fiore, Mohamed Bakhouya, and Waleed W. Smari. On the road to exascale: Advances in high performance computing and simulations — an overview and editorial. Future Generation Computer Systems, 82(??):450–458, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18301146.

Fustes:2014:CIW

[FCD^{+}14] Diego Fustes, Diego Cantorna, Carlos Dafonte, Bernardino Arcay, Alfonso Iglesias, and Minia Manteiga. A cloud-integrated Web platform for marine monitoring using GIS and remote sensing. Application to oil spill detection through SAR images. Future Generation Computer Systems, 34(??):155–160, May 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001994.

Fan:2018:VIS

[FCY18] Chih-Tien Fan, Yue-Shan Chang, and Shyan-Ming Yuan. VM instance selection for deadline constraint job on agent-based interconnected cloud. Future Generation Computer Systems, 87(??):470–487, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17314449.

Farsandaj:2012:SGB

[FD12] Kian Farsandaj and Chen Ding. Scatter/Gather browsing of Web service QoS data. Future Generation Computer Systems, 28(7):1145–1154, July 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002214.

Faria:2019:UCE

[FdAGdAFV19] Kenyo Abadio Crosara Faria, Raphael de Aquino Gomes, Eduardo Noronha de Andrade Freitas, and Auri Marcelo Rizzo Vincenzi. On using collaborative economy for test cost reduction in high fragmented environments. Future Generation Computer Systems, 95(??):502–510, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18310410.

Fadika:2014:MUM

[FDGR14] Zacharia Fadika, Elif Dede, Madhusudhan Govindaraju, and Lavanya Ramakrishnan. MARIANE: Using MApReduce in HPC environments. Future Generation Computer Systems, 36(??):379–388, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002719.

Ficco:2017:OTA

[FDPR17] Massimo Ficco, Beniamino Di Martino, Roberto Pietrantuono, and Stefano Russo. Optimized task allocation on private cloud for hybrid simulation of large-scale critical systems. Future Generation Computer Systems, 74(??):104– 118, September 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300061.

Fabra:2019:RPR

[FEÁ19] Javier Fabra, Joaquín Ezpeleta, and Pedro Álvarez. Reducing the price of resource provisioning using EC2 spot instances with prediction models. Future Generation Computer Systems, 96(??):348–367, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831166X.

Fiore:2019:BSG

[FEB^{+}19] Sandro Fiore, Donatello Elia, Ignacio Blanquer, Francisco V. Brasileiro, Alessandra Nuzzo, Paola Nassisi, Iana A. A. Rufino, Arie C. Seijmonsbergen, Niels S. Anders, Carlos de O. Galvão, John E. de B. L. Cunha, Miguel Caballer, Mariane S. Sousa-Baena, Vanderlei P. Canhos, and Giovanni Aloisio. BioClimate: a science gateway for climate change and biodiversity research in the EUBrazilCloud-Connect project. Future Generation Computer Systems, 94(??):895–909, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17312591.

Ficco:2018:CRG

[FEPC18] Massimo Ficco, Christian Esposito, Francesco Palmieri, and Aniello Castiglione. A coral-reefs and game theory-based approach for optimizing elastic cloud resource allocation. Future Generation Computer Systems, 78 (part 1)(??):343–352, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301364.

Ferretti:2013:GRD

[Fer13] Stefano Ferretti. Gossiping for resource discovering: an analysis based on complex network theory. Future Generation Computer Systems, 29(6):1631– 1644, August 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001367.

Fons:2012:BBC

[FFC12] M. Fons, F. Fons, and E. Cantó. Biometricsbased consumer applications driven by reconfigurable hardware architectures. Future Generation Computer Systems, 28(1):268–286, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X10002177.

Farahani:2018:TFD

[FFC^{+}18] Bahar Farahani, Farshad Firouzi, Victor Chang, Mustafa Badaroglu, Nicholas Constant, and Kunal Mankodiya. Towards fog-driven IoT eHealth: Promises and challenges of IoT in medicine and healthcare. Future Generation Computer Systems, 78 (part 2)(??):659–676, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17307677.

Filva:2019:CLA

[FFGP^{+}19] Daniel Amo Filvà, Marc Alier Forment, Francisco José García-Peñalvo, David Fonseca Escudero, and María José Casañ. Clickstream for learning analytics to assess students’ behavior with scratch. Future Generation Computer Systems, 93(??): 673–686, April 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0167739X18313499.

Fu:2019:WAO

[FFL^{+}19] Xiuwen Fu, Giancarlo Fortino, Wenfeng Li, Pasquale Pace, and Yongsheng Yang. WSNs-assisted opportunistic network for low-latency message forwarding in sparse settings. Future Generation Computer Systems, 91(??):223–237, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18307350.

Folino:2010:GPS

[FFPS10] Gianluigi Folino, Agostino Forestiero, Giuseppe Papuzzo, and Giandomenico Spezzano. A grid portal for solving geoscience problems using distributed knowledge discovery services. Future Generation Computer Systems, 26(1): 87–96, January 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fito:2014:BDM

[FG14] J. Oriol Fitó and Jordi Guitart. Business-driven management of infrastructurelevel risks in Cloud providers. Future Generation Computer Systems, 32(??):41– 53, March 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001045.

Faheem:2018:MMS

[FG18] Muhammad Faheem and V. C. Gungor. MQRP: Mobile sinks-based QoSaware data gathering protocol for wireless sensor networks-based smart grid applications in the context of industry 4.0-based on Internet of Things. Future Generation Computer Systems, 82(??):358–374, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17314541.

Feng:2013:EDM

[FGG13] Dawei Feng, Cécile Germain, and Tristan Glatard. Efficient distributed monitoring with active Collaborative Prediction. Future Generation Computer Systems, 29(8):2272–2283, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001155.

Faro:2011:MMD

[FGM11] Alberto Faro, Daniela Giordano, and Francesco Maiorana. Mining massive datasets by an unsupervised parallel clustering on a GRID: Novel algorithms and case study. Future Generation Computer Systems, 27(6):711–724, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fang:2019:CTA

[FGW^{+}19] Baofu Fang, Xiaoping Guo, Zaijun Wang, Yong Li, Mohamed Elhoseny, and Xiaohui Yuan. Collaborative task assignment of interconnected, affective robots towards autonomous healthcare assistant. Future Generation Computer Systems, 92(??):241–251, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18316844.

Fan:2013:CPP

[FH13] Chun-I Fan and Shi-Yuan Huang. Controllable privacy preserving search based on symmetric predicate encryption in cloud storage. Future Generation Computer Systems, 29(7):1716–1724, September 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1200101X.

Feng:2019:EDC

[FHHM19] Dandan Feng, Lijun Hu, Jing Hao, and Zhengliang Ma. Effect of dexmedetomidine on comfort and satisfaction of patients. Future Generation Computer Systems, 98(??):167–170, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18326955.

Fan:2015:IPF

[FHYH15] Haolong Fan, Farookh Khadeer Hussain, Muhammad Younas, and Omar Khadeer Hussain. An integrated personalization framework for SaaS-based cloud services. Future Generation Computer Systems, 53(??):157– 173, December 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001867.

Feng:2018:ABB

[FHZW18] Qi Feng, Debiao He, Sherali Zeadally, and Huaqun Wang. Anonymous biometrics-based authentication scheme with key distribution for mobile multi-server environment. Future Generation Computer Systems, 84(??):239– 251, July 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17309020.

Faroughi:2018:CCA

[FJ18] Azadeh Faroughi and Reza Javidan. CANF: Clustering and anomaly detection method using nearest and farthest neighbor. Future Generation Computer Systems, 89(??):166–177, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17325128.

Farhan:2018:IBS

[FJA^{+}18] Muhammad Farhan, Sohail Jabbar, Muhammad Aslam, Mohammad Hammoudeh, Mudassar Ahmad, Shehzad Khalid, Murad Khan, and Kijun Han. IoT-based students interaction framework using attention-scoring assessment in eLearning. Future Generation Computer Systems, 79 (part 3)(??):909– 919, February 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17310920.

Farman:2018:MCB

[FJJ^{+}18] Haleem Farman, Bilal Jan, Huma Javed, Naveed Ahmad, Javed Iqbal, Muhammad Arshad, and Shaukat Ali. Multi-criteria based zone head selection in Internet of Things based wireless sensor networks. Future Generation Computer Systems, 87(??):364– 371, October 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167739X17320174.

Firoozjaei:2017:SCN

[FJKK17] Mahdi Daghmehchi Firoozjaei, Jaehoon (Paul) Jeong, Hoon Ko, and Hyoungshick Kim. Security challenges with network functions virtualization. Future Generation Computer Systems, 67(??):315–324, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302321.

Farokhi:2016:HCC

[FJL^{+}16] Soodeh Farokhi, Pooyan Jamshidi, Ewnetu Bayuh Lakew, Ivona Brandic, and Erik Elmroth. A hybrid cloud controller for vertical memory elasticity: a control-theoretic approach. Future Generation Computer Systems, 65(??):57– 72, December 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630139X.

Farkas:2011:PGP

[FK11] Z. Farkas and P. Kacsuk. P-GRADE Portal: a generic workflow system to support user communities. Future Generation Computer Systems, 27(5): 454–465, May 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Farkas:2012:EHD

[FK12] Z. Farkas and P. Kacsuk. Evaluation of hierarchical desktop grid scheduling algorithms. Future Generation Computer Systems, 28(6):871–880, June 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X10002748.

Farkas:2010:IBE

[FKBG10] Z. Farkas, P. Kacsuk, Z. Balaton, and G. Gombás. Interoperability of BOINC and EGEE. Future Generation Computer Systems, 26(8):1092–1103, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Freitas:2011:WBO

[FKOC11] Andre Freitas, Tomas Knap, Sean O’Riain, and Edward Curry. W3P: Building an OPM based provenance model for the Web. Future Generation Computer Systems, 27(6): 766–774, June 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Forkan:2014:CCO

[FKT14] Abdur Forkan, Ibrahim Khalil, and Zahir Tari. Co-CaMAAL: a cloud-oriented context-aware middleware in ambient assisted living. Future Generation Computer Systems, 35(??):114– 127, June 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001544.

Fan:2019:SPP

[FLL^{+}19] Yongkai Fan, Xiaodong Lin, Wei Liang, Gang Tan, and Priyadarsi Nanda. A secure privacy preserving deduplication scheme for cloud computing. Future Generation Computer Systems, 101(??):127–135, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18329649.

Fu:2018:MVU

[FLN^{+}18] Jingcheng Fu, Jianwen Li, Yawei Niu, Guanghui Wang, and Jianliang Wu. Multipolarization versus unification in community networks. Future Generation Computer Systems, 83(??):454–460, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17310142.

Fernando:2013:MCC

[FLR13] Niroshinie Fernando, Seng W. Loke, and Wenny Rahayu. Mobile cloud computing: a survey. Future Generation Computer Systems, 29(1): 84–106, January 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001318.

Fang:2016:AOF

[FLR^{+}16] Daren Fang, Xiaodong Liu, Imed Romdhani, Pooyan Jamshidi, and Claus Pahl. An agility-oriented and fuzziness-embedded semantic model for collaborative cloud service search, retrieval and recommendation. Future Generation Computer Systems, 56(??): 11–26, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003052.

Fan:2017:DFC

[FLT17] Wenhao Fan, Yuan’an Liu, and Bihua Tang. DEXIN: a fast content-based multiattribute event matching algorithm using dynamic exclusive and inclusive methods. Future Generation Computer Systems, 68(??):289–303, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630437X.

Fan:2019:OSD

[FLT^{+}19] Yongkai Fan, Xiaodong Lin, Gang Tan, Yuqing Zhang, Wei Dong, and Jing Lei. One secure data integrity verification scheme for cloud storage. Future Generation Computer Systems, 96(??):376–385, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18311002.

Folino:2010:SSB

[FM10a] Gianluigi Folino and Carlo Mastroianni. Special section: Bio-inspired algorithms for distributed systems. Future Generation Computer Systems, 26(6): 835–837, June 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Furlinger:2010:RCF

[FM10b] Karl Fürlinger and Shirley Moore. Recording the control flow of parallel applications to determine iterative and phase-based behavior. Future Generation Computer Systems, 26 (1):162–166, January 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fortis:2017:TCI

[FM17] Teodor-Florin Fortis and Victor Ion Munteanu. Topics in cloud incident management. Future Generation Computer Systems, 72(??):163–164, July 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16305179.

Farris:2017:MMI

[FMN^{+}17] I. Farris, L. Militano, M. Nitti, L. Atzori, and A. Iera. MIFaaS: a mobile-IoT-federation-as-a-service model for dynamic cooperation of IoT cloud providers. Future Generation Computer Systems, 70(??):126–137, May 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302138.

Fortino:2018:UTL

[FMRS18] G. Fortino, F. Messina, D. Rosaci, and G. M. L. Sarné. Using trust and local reputation for group formation in the Cloud of Things. Future Generation Computer Systems, 89(??):804–815, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18308707.

Feller:2012:ICH

[FMSSM12] Eugen Feller, John Mehnert-Spahn, Michael Schoettner, and Christine Morin. Independent checkpointing in a heterogeneous grid environment. Future Generation Computer Systems, 28(1):163–170, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000409.

Farruggia:2014:TBI

[FMV14] Alfonso Farruggia, Rosario Magro, and Salvatore Vitabile. A text based indexing system for mammographic image retrieval and classification. Future Generation Computer Systems, 37(??):243–251, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000338.

Fiore:2011:DAL

[FNA11] Sandro Fiore, Alessandro Negro, and Giovanni Aloisio. The data access layer in the GRelC system architecture. Future Generation Computer Systems, 27(3):334–340, March 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fiore:2012:CGP

[FNA12] Sandro Fiore, Alessandro Negro, and Giovanni Aloisio. The Climate-G Portal: The context, key features and a multidimensional analysis. Future Generation Computer Systems, 28(1):1–8, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11000975.

Ferreto:2011:SCM

[FNCR11] Tiago C. Ferreto, Marco A. S. Netto, Rodrigo N. Calheiros, and César A. F. De Rose. Server consolidation with migration control for virtualized data centers. Future Generation Computer Systems, 27(8): 1027–1034, October 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fox:2013:RWU

[FP13] Geoffrey Fox and Shrideep Pallickara. Recent work in utility and cloud computing. Future Generation Computer Systems, 29(4): 986–987, June 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000071.

Fortino:2014:ICC

[FP14] Giancarlo Fortino and Mukaddim Pathan. Integration of Cloud computing and body sensor networks. Future Generation Computer Systems, 35(??): 57–61, June 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1400017X.

Filelis-Papadopoulos:2018:FSL

[FPGK18] Christos K. Filelis-Papadopoulos, George A. Gravvanis, and Panagiotis E. Kyziropoulos. A framework for simulating large scale cloud infrastructures. Future Generation Computer Systems, 79 (part 2)(??):703– 714, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17303230.

Fu:2019:SES

[FPL^{+}19] Cai Fu, Chenchen Peng, Xiao-Yang Liu, Laurence T. Yang, Jia Yang, and Lansheng Han. Search engine: The social relationship driving power of Internet of Things. Future Generation Computer Systems, 92(??):972–986, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17307884.

Furfaro:2018:CBP

[FPP^{+}18] Angelo Furfaro, Antonio Piccolo, Andrea Parise, Luciano Argento, and Domenico Saccà. A cloudbased platform for the emulation of complex cybersecurity scenarios. Future Generation Computer Systems, 89(??):791–803, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630704X.

Fortino:2014:BSA

[FPPD14] Giancarlo Fortino, Daniele Parisi, Vincenzo Pirrone, and Giuseppe Di Fatta. BodyCloud: a SaaS approach for community body sensor networks. Future Generation Computer Systems, 35(??):62– 79, June 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002793.

Ficco:2018:ARP

[FPR18] Massimo Ficco, Roberto Pietrantuono, and Stefano Russo. Aging-related performance anomalies in the Apache Storm stream processing system. Future Generation Computer Systems, 86(??):975–994, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17319027.

Fernandez-Quiruelas:2015:LSC

[FQBCF15] V. Fernández-Quiruelas, C. Blanco, A. S. Cofiño, and J. Fernández. Largescale climate simulations harnessing clusters, grid and cloud infrastructures. Future Generation Computer Systems, 51(??):36– 44, October 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001016.

Ferreira:2014:AIC

[FRB^{+}14] Kurt B. Ferreira, Rolf Riesen, Patrick Bridges, Dorian Arnold, and Ron Brightwell. Accelerating incremental checkpointing for extreme-scale computing. Future Generation Computer Systems, 30(??): 66–77, January 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000770.

Frincu:2014:SHA

[Frî14] Marc Eduard Frîncu. Scheduling highly available applications on cloud environments. Future Generation Computer Systems, 32(??): 138–153, March 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001136.

Firouzi:2018:ITB

[FRM^{+}18] Farshad Firouzi, Amir M. Rahmani, K. Mankodiya, M. Badaroglu, G. V. Merrett, P. Wong, and Bahar Farahani. Internetof-Things and big data for smarter healthcare: From device to architecture, applications and analytics. Future Generation Computer Systems, 78 (part 2)(??):583–586, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17319726.

Filho:2019:PPA

[FRZ19] Walter Priesnitz Filho, Carlos Ribeiro, and Thomas Zefferer. Privacy-preserving attribute aggregation in eID federations. Future Generation Computer Systems, 92(??):1–16, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17327966.

Fugkeaw:2018:SSA

[FS18] Somchart Fugkeaw and Hiroyuki Sato. Scalable and secure access control policy update for outsourced big data. Future Generation Computer Systems, 79 (part 1)(??):364– 373, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312426.

Francis:2019:LDT

[FS19] Leena Mary Francis and N. Sreenath. Live detection of text in the natural environment using convolutional neural network. Future Generation Computer Systems, 98(??):444– 455, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18322180.

Farias:2018:RBP

[FSM^{+}18a] Victor A. E. Farias, Flávio R. C. Sousa, José Gilvan R. Maia, João Paulo P. Gomes, and Javam C. Machado. Regression based performance modeling and provisioning for NoSQL cloud databases. Future Generation Computer Systems, 79 (part 1) (??):72–81, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301607.

Fysarakis:2018:XCD

[FSM^{+}18b] Konstantinos Fysarakis, Othonas Soultatos, Charalampos Manifavas, Ioannis Papaefstathiou, and Ioannis Askoxylakis. XSACdcross-domain resource sharing & access control for smart environments. Future Generation Computer Systems, 80(??):572–582, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301340.

Fylaktopoulos:2018:DMP

[FSP^{+}18] G. Fylaktopoulos, M. Skolarikis, I. Papadopoulos, G. Goumas, A. Sotiropoulos, and I. Maglogiannis. A distributed modular platform for the development of cloud based applications. Future Generation Computer Systems, 78 (part 1)(??):127–141, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302716.

Fei:2019:CDS

[FSV^{+}19] Xiang Fei, Nazaraf Shah, Nandor Verba, Kuo-Ming Chao, Victor Sanchez-Anguix, Jacek Lewandowski, Anne James, and Zahid Usman. CPS data streams analytics based on machine learning for cloud and fog computing: a survey. Future Generation Computer Systems, 90(??):435–450, January 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17330613.

Fan:2019:BBC

[FSY^{+}19] Kai Fan, Shili Sun, Zheng Yan, Qiang Pan, Hui Li, and Yintang Yang. A blockchain-based clock synchronization scheme in IoT. Future Generation Computer Systems, 101(??):524–533, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18326657.

Fahad:2014:PPP

[FTA^{+}14] Adil Fahad, Zahir Tari, Abdulmohsen Almalawi, Andrzej Goscinski, Ibrahim Khalil, and Abdun Mahmood. PPFSCADA: Privacy preserving framework for SCADA data publishing. Future Generation Computer Systems, 37(??): 496–511, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000478.

Fki:2017:AFC

[FTD17] Emna Fki, Said Tazi, and Khalil Drira. Automated and flexible composition based on abstract services for a better adaptation to user intentions. Future Generation Computer Systems, 68(??):376–390, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302382.

Fugini:2016:WBC

[FTH16] Mariagrazia Fugini, Mahsa Teimourikia, and George Hadjichristofi. A webbased cooperative tool for risk management with adaptive security. Future Generation Computer Systems, 54(??):409–422, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001077.

Fahad:2014:OSF

[FTK^{+}14] Adil Fahad, Zahir Tari, Ibrahim Khalil, Abdulmohsen Almalawi, and Albert Y. Zomaya. An optimal and stable feature selection approach for traffic classification based on multi-criterion fusion. Future Generation Computer Systems, 36(??):156–169, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001945.

Fensel:2017:CAE

[FTK17] Anna Fensel, Dana Kathrin Tomic, and Andreas Koller. Contributing to appliances’ energy efficiency with Internet of Things, smart data and user engagement. Future Generation Computer Systems, 76(??):329– 338, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306653.

Fernandez:2014:RDS

[FTP14] Hector Fernandez, Cédric Tedeschi, and Thierry Priol. Rule-driven service coordination middleware for scientific applications. Future Generation Computer Systems, 35 (??):1–13, June 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002872.

Feng:2019:PPA

[FW19] Xia Feng and Liangmin Wang. PAU: Privacy Assessment method with Uncertainty consideration for cloud-based vehicular networks. Future Generation Computer Systems, 96(??): 368–375, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18326608.

Fu:2013:GSA

[FWB13a] Yongquan Fu, Yijie Wang, and Ernst Biersack. A general scalable and accurate decentralized level monitoring method for largescale dynamic service provision in hybrid clouds. Future Generation Computer Systems, 29(5):1235–1253, July 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12002075.

Fu:2013:HAS

[FWB13b] Yongquan Fu, Yijie Wang, and Ernst Biersack. HybridNN: an accurate and scalable network location service based on the inframetric model. Future Generation Computer Systems, 29(6):1485–1504, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12002221.

Fan:2010:MPP

[FX10] Xiang Fan and Yang Xiang. Modeling the propagation of Peer-to-Peer worms. Future Generation Computer Systems, 26(8): 1433–1443, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fan:2019:EPP

[FXG^{+}19] Kai Fan, Huiyue Xu, Longxiang Gao, Hui Li, and Yintang Yang. Efficient and privacy preserving access control scheme for fog-enabled IoT. Future Generation Computer Systems, 99(??):134–142, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18323367.

Feng:2019:MCD

[FY19] Wei Feng and Zheng Yan. MCS-Chain: Decentralized and trustworthy mobile crowdsourcing based on blockchain. Future Generation Computer Systems, 95(??):649–666, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18326967.

Feng:2014:MND

[FZHH14] Wenying Feng, Qinglei Zhang, Gongzhu Hu, and Jimmy Xiangji Huang. Mining network data for intrusion detection through combining SVMs with ant colony networks. Future Generation Computer Systems, 37(??):127–140, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001416.

Feng:2018:DDA

[FZT^{+}18] Yixiong Feng, Zhifeng Zhang, Guangdong Tian, Zhihan Lv, Shaoxu Tian, and Hongfei Jia. Datadriven accurate design of variable blank holder force in sheet forming under interval uncertainty using sequential approximate multi-objective optimization. Future Generation Computer Systems, 86(??):1242–1250, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17303114.

Fang:2018:EPP

[FZW^{+}18] Yuyi Fang, Zhiwei Zhao, Zi Wang, Geyong Min, Yue Cao, Haojun Huang, and Hao Yin. Eavesdrop with PoKeMon: Position free keystroke monitoring using acoustic data. Future Generation Computer Systems, 87(??):704–711, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17300493.

Ghaffarinejad:2013:ICD

[GA13] Ali Ghaffarinejad and Mohammad Kazem Akbari. An incentive compatible and distributed reputation mechanism based on context similarity for service oriented systems. Future Generation Computer Systems, 29(3):863–875, March 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000866.

Gungor:2019:ASC

[GAA19] Onat Güngör, Baris Aksanli, and Reyhan Aydogan. Algorithm selection and combining multiple learners for residential energy prediction. Future Generation Computer Systems, 99(??):391–400, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19305795.

Garijo:2014:CMS

[GAB^{+}14] Daniel Garijo, Pinar Alper, Khalid Belhajjame, Oscar Corcho, Yolanda Gil, and Carole Goble. Common motifs in scientific workflows: an empirical analysis. Future Generation Computer Systems, 36(??): 338–351, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001970.

Gao:2017:ADI

[GACM17] Feng Gao, Muhammad Intizar Ali, Edward Curry, and Alessandra Mileo. Automated discovery and integration of semantic urban data streams: the ACEIS middleware. Future Generation Computer Systems, 76(??):561–581, November 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1730331X.

Galbally:2012:HPF

[GAFFOG12] Javier Galbally, Fernando Alonso-Fernandez, Julian Fierrez, and Javier Ortega-Garcia. A high performance fingerprint liveness detection method based on quality related features. Future Generation Computer Systems, 28(1):311– 321, January 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1000244X.

Gope:2018:LPP

[GAI^{+}18] Prosanta Gope, Ruhul Amin, S. K. Hafizul Islam, Neeraj Kumar, and Vinod Kumar Bhalla. Lightweight and privacypreserving RFID authentication scheme for distributed IoT infrastructure with secure localization services for smart city environment. Future Generation Computer Systems, 83(??):629–637, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17313043.

Ghobaei-Arani:2018:ARP

[GAJP18] Mostafa Ghobaei-Arani, Sam Jabbehdari, and Mohammad Ali Pourmina. An autonomic resource provisioning approach for service-based cloud applications: a hybrid approach. Future Generation Computer Systems, 78 (part 1)(??):191–210, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302327.

Gimenez-Alventosa:2019:FPA

[GAMC19]  V.  Giménez-Alventosa,

Germán Moltó, and Miguel Caballer. A framework and a performance assessment for serverless MapReduce on AWS Lambda. Future Generation Computer Systems, 97(??):259–274, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18325172.

Garg:2018:CCB

[GAW^{+}18] Saurabh Garg, Jagannath Aryal, Hao Wang, Tejal Shah, Gabor Kecskemeti, and Rajiv Ranjan. Cloud computing based bushfire prediction for cyber physical emergency applications. Future Generation Computer Systems, 79 (part 1)(??):354– 363, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17302042.

Gonzalez-Aparicio:2018:TTS

[GAYTC18] María Teresa González-Aparicio, Muhammad Younas, Javier Tuya, and Rubén Casado. Testing of transactional services in NoSQL key-value databases. Future Generation Computer Systems, 80(??):384–399, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1731467X.

Goscinski:2010:TDA

[GB10] Andrzej Goscinski and Michael Brock. Toward dynamic and attribute based publication, discovery and selection for cloud computing. Future Generation Computer Systems, 26(7): 947–970, July 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ge:2018:BDI

[GBB18] Mouzhi Ge, Hind Bangui, and Barbora Buhnova. Big data for Internet of Things: A survey. Future Generation Computer Systems, 87(??):601–614, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17316953.

Goiri:2012:EEM

[GBF^{+}12] Íñigo Goiri, Josep Ll. Berral, J. Oriol Fitó, Ferran Julià, Ramon Nou, Jordi Guitart, Ricard Gavaldà, and Jordi Torres. Energy-efficient and multifaceted resource management for profit-driven virtualized data centers. Future Generation Computer Systems, 28(5):718– 731, May 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11002366.

Gu:2018:DTF

[GBKJ18] Chen Gu, Matthew Bradbury, Jack Kirton, and Arshad Jhumka. A decision theoretic framework for selecting source location privacy aware routing protocols in wireless sensor networks. Future Generation Computer Systems, 87(??):514–526, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17317028.

Gubbi:2013:ITI

[GBMP13] Jayavardhana Gubbi, Rajkumar Buyya, Slaven Marusic, and Marimuthu Palaniswami. Internet of Things (IoT): a vision, architectural elements, and future directions. Future Generation Computer Systems, 29(7):1645–1660, September 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000241.

Goher:2018:CPC

[GBRM18] Syeda ZarAfshan Goher, Peter Bloodsworth, Raihan Ur Rasool, and Richard McClatchey. Cloud provider capacity augmentation through automated resource bartering. Future Generation Computer Systems, 81(??):203–218, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17304867.

Garg:2010:TCT

[GBS10] Saurabh Kumar Garg, Rajkumar Buyya, and Howard Jay Siegel. Time and cost trade-off management for scheduling parallel applications on Utility Grids. Future Generation Computer Systems, 26(8): 1344–1355, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ghosh:2016:STC

[GBY16] Goldina Ghosh, Soumya Banerjee, and Neil Y. Yen. State transition in communication under social network: an analysis using fuzzy logic and density based clustering towards big data paradigm. Future Generation Computer Systems, 65(??):207– 220, December 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300309.

Gesing:2017:BAL

[GCBM17] Sandra Gesing, Jesus Carretero, Javier Garcia Blas, and Johan Montagnat. Boosting analyses in the life sciences via clusters, grids and clouds. Future Generation Computer Systems, 67(??):325–328, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16305064.

Gong:2018:MIL

[GCCL18] Peng Gong, Yuan Cao, Baigen Cai, and Kun Li. Multi-information location data fusion system of railway signal based on cloud computing. Future Generation Computer Systems, 88(??):594–598, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18307015.

Garcia-Crespo:2010:SSL

[GCCPGBGS10] Ángel García-Crespo, Ricardo Colomo-Palacios, Juan Miguel Gómez-Berbís, and Francisco García-Sánchez. SOLAR: Social Link Advanced Recommendation System. Future Generation Computer Systems, 26(3):374–380, March 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gosman:2018:CFU

[GCD^{+}18] Catalin Gosman, Tudor Cornea, Ciprian Dobre, Florin Pop, and Aniello Castiglione. Controlling and filtering users data in intelligent transportation system. Future Generation Computer Systems, 78 (part 2)(??):807–816, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16307555.

Gohar:2018:CBG

[GCK18] Moneeb Gohar, Jin-Ghoo Choi, and Seok-Joo Koh. CoAP-based group mobility management protocol for the Internet-of-Things in WBAN environment. Future Generation Computer Systems, 88(??):309– 318, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17324743.

Gadelha:2011:PMS

[GCM^{+}11] Luiz M. R. Gadelha, Jr., Ben Clifford, Marta Mattoso, Michael Wilde, and Ian Foster. Provenance management in Swift. Future Generation Computer Systems, 27(6):775–780, June 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gonzalez-Compean:2019:PBC

[GCTLA^{+}19] J. L. Gonzalez-Compean, Oscar Telles, Ivan Lopez-Arevalo, Miguel Morales-Sandoval, Victor J. Sosa-Sosa, and Jesus Carretero. A policy-based containerized filter for secure information sharing in organizational environments. Future Generation Computer Systems, 95(??):430–444, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18311774.

Gomez:2014:FTV

[GCV^{+}14] A. Gómez, L. M. Carril, R. Valin, J. C. Mouriño, and C. Cotelo. Faulttolerant virtual cluster experiments on federated sites using BonFIRE. Future Generation Computer Systems, 34(??):17– 25, May 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002914.

Gu:2019:EET

[GCZ^{+}19] Lin Gu, Jingjing Cai, Deze Zeng, Yu Zhang, Hai Jin, and Weiqi Dai. Energy efficient task allocation and energy scheduling in green energy powered edge computing. Future Generation Computer Systems, 95(??): 89–99, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18310550.

Ghosh:2010:MAC

[GD10] Preetam Ghosh and Sajal K. Das. Mobility-aware cost-efficient job scheduling for single-class grid jobs in a generic mobile grid architecture. Future Generation Computer Systems, 26(8): 1356–1367, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Goyal:2018:TDG

[GDAS18] Bhawna Goyal, Ayush Dogra, Sunil Agrawal, and B. S. Sohi. Twodimensional gray scale image denoising via morphological operations in NSST domain & bitonic filtering. Future Generation Computer Systems, 82(??):158– 175, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17319106.

Garcia:2019:ECS

[GdCP19] Álvaro López García, Enol Fernández del Castillo, and Isabel Campos Plasencia. An efficient cloud scheduler design supporting preemptible instances. Future Generation Computer Systems, 95(??):68– 78, June 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1830270X.

Ghafarian:2013:CPA

[GDJ^{+}13] Toktam Ghafarian, Hossein Deldari, Bahman Javadi, Mohammad H. Yaghmaee, and Rajkumar Buyya. CycloidGrid: a proximity-aware P2Pbased resource discovery architecture in volunteer computing systems. Future Generation Computer Systems, 29(6):1583–1595, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001665.

Gesing:2018:GRA

[GDP^{+}18] Sandra Gesing, Rion Dooley, Marlon Pierce, Jens Krüger, Richard Grunzke, Sonja Herres-Pawlis, and Alexander Hoffmann. Gathering requirements for advancing simulations in HPC infrastructures via science gateways. Future Generation Computer Systems, 82(??):544–554, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17303011.

Gueye:2014:CSS

[GDR^{+}14] Soguy Mak-Karé Gueye, Noël De Palma, Éric Rutten, Alain Tchana, and Nicolas Berthier. Coordinating self-sizing and self-repair managers for multi-tier systems. Future Generation Computer Systems, 35(??):14–26, June 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1400003X.

Grange:2018:GIS

[GDS18] Léo Grange, Georges Da Costa, and Patricia Stolf. Green IT scheduling for data center powered with renewable energy. Future Generation Computer Systems, 86(??):99–120, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17300092.

Gonzalez:2010:CFT

[GdVC10] Daniel Lombraña González, Francisco Fernández de Vega, and Henri Casanova. Characterizing fault tolerance in genetic programming. Future Generation Computer Systems, 26(6):847– 856, June 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gesing:2019:SGC

[GDZ^{+}19] Sandra Gesing, Maytal Dahan, Michael Zentner, Nancy Wilkins-Diehr, and Katherine Lawrence. The Science Gateways Community Institute: Collaborations and efforts on international scale. Future Generation Computer Systems, 101(??):951–958, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18315176.

Giot:2013:FCP

[GEAR13] Romain Giot, Mohamad El-Abed, and Christophe Rosenberger. Fast computation of the performance evaluation of biometric systems: Application to multibiometrics. Future Generation Computer Systems, 29(3):788– 799, March 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000362.

Garcia:2014:SDD

[GEG14] Andrés García García, Ignacio Blanquer Espert, and Vicente Hernández García. SLA-driven dynamic cloud resource management. Future Generation Computer Systems, 31(??):1–11, February 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1300215X.

Granell:2014:GII

[GFD14] Carlos Granell, Óscar Belmonte Fernández, and Laura Díaz. Geospatial information infrastructures to address spatial needs in health: Collaboration, challenges and opportunities. Future Generation Computer Systems, 31(??):213–222, February 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000629.

Gu:2018:GCD

[GFW^{+}18] Chonglin Gu, Longxiang Fan, Wenbin Wu, Hejiao Huang, and Xiaohua Jia. Greening cloud data centers in an economical way by energy trading with power grid. Future Generation Computer Systems, 78 (part 1)(??):89–101, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308330.

Gu:2010:SHP

[GG10] Yunhong Gu and Robert L.

Grossman. Sector: a high performance wide area community data storage and sharing system. Future Generation Computer Systems, 26(5):720–728, May 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Guerout:2017:MIL

[GGA^{+}17] Tom Guérout, Yacine Gaoua, Christian Artigues, Georges Da Costa, Pierre Lopez, and Thierry Monteil. Mixed integer linear programming for quality of service optimization in clouds. Future Generation Computer Systems, 71 (??):1–17, June 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630869X.

Garijo:2017:ALP

[GGC17] Daniel Garijo, Yolanda Gil, and Oscar Corcho. Abstract, link, publish, exploit: an end to end framework for workflow sharing.

Future Generation Computer Systems, 75(??):271– 283, October 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300274.

Gupta:2018:HDB

[GGC18] Shashank Gupta, B. B. Gupta, and Pooja Chaudhary. Hunting for DOMbased XSS vulnerabilities in mobile cloud-based online social network. Future Generation Computer Systems, 79 (part 1)(??): 319–336, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17311068.

Gonzalez-Granadillo:2018:DRM

[GGDM^{+}18] G. Gonzalez-Granadillo, S. Dubus, A. Motzek, J. Garcia-Alfaro, E. Alvarez, M. Merialdo, S. Papillon, and H. Debar. Dynamic risk management response system to handle cyber threats. Future Generation Computer Systems, 83(??):535–552, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17311433.

Guna:2019:IVC

[GGH^{+}19] Joze Guna, Gregor Gersak, Iztok Humar, Jeungeun Song, Janko Drnovšek, and Matevž Pogačnik. Influence of video content type on users’ virtual reality sickness perception and physiological response. Future Generation Computer Systems, 91(??):263–276, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18316546.

Gentzsch:2013:HPC

[GGJ13] Wolfgang Gentzsch, Lucio Grandinetti, and Gerhard Joubert. High performance computing in the cloud. Future Generation Computer Systems, 29(6): 1407, August 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001021.

Gesbert:2010:BSP

[GGLD10] Louis Gesbert, Frédéric Gava, Frédéric Loulergue, and Frédéric Dabrowski. Bulk synchronous parallel ML with exceptions. Future Generation Computer Systems, 26(3):486– 490, March 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Guo:2018:RPC

[GGLW18] Junpeng Guo, Zeng Gao, Na Liu, and Yi Wu. Recommend products with consideration of multicategory inter-purchase time and price. Future Generation Computer Systems, 78 (part 1)(??):451– 461, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1730256X.

Gurini:2018:TPP

[GGMS18] Davide Feltoni Gurini, Fabio Gasparetti, Alessandro Micarelli, and Giuseppe Sansonetti. Temporal people-to-people recommendation on social networks with sentimentbased matrix factorization. Future Generation Computer Systems, 78 (part 1)(??):430–439, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17304077.

Ghadiri:2017:BFP

[GGN17] Nasser Ghadiri, Meysam Ghaffari, and Mohammad Amin Nikbakht. BigFCM: Fast, precise and scalable FCM on Hadoop. Future Generation Computer Systems, 77(??):29– 39, December 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17312359.

Gutierrez-Garcia:2013:FHA

[GGS13] J. Octavio Gutierrez-Garcia and Kwang Mong Sim. A family of heuristics for agent-based elastic cloud bag-of-tasks concurrent scheduling. Future Generation Computer Systems, 29(7):1682–1699, September 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000179.

Garcia-Galan:2016:ACS

[GGTRRC16] Jesús García-Galán, Pablo Trinidad, Omer F. Rana, and Antonio Ruiz-Cortés. Automated configuration support for infrastructure migration to the cloud. Future Generation Computer Systems, 55(??):200–212, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000618.

Genge:2019:ESA

[GHD19] Béla Genge, Piroska Haller, and Adrian-Vasile Duka. Engineering security-aware control applications for data authentication in smart industrial cyberphysical systems. Future Generation Computer Systems, 91(??):206–222, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18303431.

Gharbia:2018:MSP

[GHEB^{+}18] Reham Gharbia, Aboul Ella Hassanien, Ali Hassan El-Baz, Mohamed Elhoseny, and M. Gunasekaran. Multi-spectral and panchromatic image fusion approach using stationary wavelet transform and swarm flower pollination optimization for remote sensing applications. Future Generation Computer Systems, 88(??):501– 511, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18308318.

See retraction notice [GHEB^{+}23].

Gharbia:2023:RNM

[GHEB^{+}23] Reham Gharbia, Aboul Ella Hassanien, Ali Hassan El-Baz, Mohamed Elhoseny, and M. Gunasekaran. Retraction notice to “Multispectral and panchromatic image fusion approach using stationary wavelet transform and swarm flower pollination optimization for remote sensing applications” [Future Gener. Comput. Syst. 88 (2018) 501–511]. Future Generation Computer Systems, 148(??):636, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002339.

See [GHEB^{+}18].

Garcia-Holgado:2019:VLE

[GHGP19] Alicia García-Holgado and Francisco José García-Peñalvo. Validation of the learning ecosystem metamodel using transformation rules. Future Generation Computer Systems, 91(??):300–310, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18312767.

Grunzke:2019:MRS

[GHJ^{+}19] Richard Grunzke, Volker Hartmann, Thomas Jejkal, Helen Kollai, Ajinkya Prabhune, Hendrik Herold, Aline Deicke, Christiane Dressler, Julia Dolhoff, Julia Stanek, Alexander Hoffmann, Ralph Müller-Pfefferkorn, Torsten Schrade, Gotthard Meinel, Sonja Herres-Pawlis, and Wolfgang E. Nagel. The MASi repository service — comprehensive, metadatadriven and multi-community research data management. Future Generation Computer Systems, 94(??):879– 894, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17305344.

Geiger:2018:BSS

[GHLW18] Matthias Geiger, Simon Harrer, Jörg Lenhard, and Guido Wirtz. BPMN 2.0: the state of support and implementation. Future Generation Computer Systems, 80(??):250–262, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17300250.

Gao:2010:QAM

[GHMX10] Jianliang Gao, Jia Hu, Geyong Min, and Li Xu. QoS analysis of medium access control in LR-WPANs under bursty error channels. Future Generation Computer Systems, 26(8):

1426–1432, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Grosso:2011:CSH

[GHO^{+}11] Paola Grosso, Laurin Herr, Naohisa Ohta, Paul Hearty, and Cees de Laat. CineGrid: Super high definition media over optical networks. Future Generation Computer Systems, 27(7):881–885, July 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ghafir:2018:DAP

[GHP^{+}18] Ibrahim Ghafir, Mohammad Hammoudeh, Vaclav Prenosil, Liangxiu Han, Robert Hegarty, Khaled Rabie, and Francisco J. Aparicio-Navarro. Detection of advanced persistent threat using machinelearning correlation analysis. Future Generation Computer Systems, 89(??):349–359, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18307532.

Gao:2018:TSS

[GHY^{+}18] Honghao Gao, Wanqiu Huang, Xiaoxian Yang, Yucong Duan, and Yuyu Yin. Toward service selection for workflow reconfiguration:an interface-based computing solution. Future Generation Computer Systems, 87(??):298–311, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17320575.

Ge:2018:PDM

[GHYK18] Mengmeng Ge, Jin B. Hong, Simon Enoch Yusuf, and Dong Seong Kim. Proactive defense mechanisms for the softwaredefined Internet of Things with non-patchable vulnerabilities. Future Generation Computer Systems, 78 (part 2)(??):568–582, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17314723.

Gribaudo:2018:PMF

[GIK18] M. Gribaudo, M. Iacono, and M. Kiran. A performance modeling framework for lambda architecture based applications. Future Generation Computer Systems, 86(??):1032–1041, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17315364.

Gribaudo:2016:IRP

[GIM16] Marco Gribaudo, Mauro Iacono, and Daniele Manini. Improving reliability and performances in large scale distributed applications with erasure codes and replication. Future Generation Computer Systems, 56(??):773–782, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002290.

Ghafarian:2015:CAD

[GJ15] Toktam Ghafarian and Bahman Javadi. Cloudaware data intensive workflow scheduling on volunteer computing systems. Future Generation Computer Systems, 51(??):87– 97, October 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002416.

Gan:2018:FFG

[GJ18] Mingxin Gan and Rui Jiang. FLOWER: Fusing global and local associations towards personalized social recommendation. Future Generation Computer Systems, 78 (part 1)(??):462–473, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302522.

Goiri:2012:SCB

[GJF^{+}12] Íñigo Goiri, Ferran Julià, J. Oriol Fitó, Mario Macías, and Jordi Guitart. Supporting CPUbased guarantees in cloud SLAs via resource-level QoS metrics. Future Generation Computer Systems, 28(8):1295–1302, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11002275.

Gavvala:2019:QAC

[GJGB19] Siva Kumar Gavvala, Chandrashekar Jatoth, G. R. Gangadharan, and Rajkumar Buyya. QoSaware cloud service composition using eagle strategy. Future Generation Computer Systems, 90(??):273– 290, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17320630.

Grzonka:2018:UMA

[GJKP18] Daniel Grzonka, Agnieszka Jakóbik, Joanna Kolodziej, and Sabri Pllana. Using a multi-agent system and artificial intelligence for monitoring and improving the cloud performance and security. Future Generation Computer Systems, 86(??):1106–1117, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17310531.

Ghanbari:2018:CCF

[GJY18] Ryan Ghanbari, Mahdi Jalili, and Xinghuo Yu. Correlation of cascade failures and centrality measures in complex networks. Future Generation Computer Systems, 83(??):390– 400, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17319507.

Goel:2018:BMC

[GK18] Shubham Goel and Ravinder Kumar. Brownian Motus and Clustered Binary Insertion Sort methods: An efficient progress over traditional methods. Future Generation Computer Systems, 86(??):266–280, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17318423.

Grzonka:2015:ANN

[GKTK15] Daniel Grzonka, Joanna Kolodziej, Jie Tao, and Samee Ullah Khan. Artificial Neural Network support to monitoring of the evolutionary driven security aware scheduling in computational distributed environments. Future Generation Computer Systems, 51(??):72–86, October 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002258.

Gogouvitis:2012:WMS

[GKW^{+}12] Spyridon Gogouvitis, Kleopatra Konstanteli, Stefan Waldschmidt, George Kousiouris, Gregory Katsaros, Andreas Menychtas, Dimosthenis Kyriazis, and Theodora Varvarigou. Workflow management for soft realtime interactive applications in virtualized environments. Future Generation Computer Systems, 28(1):193–209, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11000999.

Ghasempouri:2019:MTR

[GL19] Seyed Asgary Ghasempouri and Behrouz Tork Ladani. Modeling trust and reputation systems in hostile environments. Future Generation Computer Systems, 99(??):571–592, October 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326013.

Guo:2018:KAA

[GLB^{+}18] Cheng Guo, Ningqi Luo, Md Zakirul Alam Bhuiyan, Yingmo Jie, Yuanfang Chen, Bin Feng, and Muhammad Alam. Keyaggregate authentication cryptosystem for data sharing in dynamic cloud storage. Future Generation Computer Systems, 84(??):190–199, July 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17307926.

Gao:2019:CPD

[GLC19] Jing Gao, Peng Li, and Zhikui Chen. A canonical polyadic deep convolutional computation model for big data feature learning in Internet of Things. Future Generation Computer Systems, 99(??):508– 516, October 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19307393.

Gesing:2019:SGS

[GLD^{+}19a] Sandra Gesing, Katherine Lawrence, Maytal Dahan, Marlon E. Pierce, Nancy Wilkins-Diehr, and Michael Zentner. Science gateways: Sustainability via on-campus teams. Future Generation Computer Systems, 94(??):97–102, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18315395.

Guan:2019:ADU

[GLD^{+}19b] Zhitao Guan, Zefang Lv, Xiaojiang Du, Longfei Wu, and Mohsen Guizani. Achieving data utilityprivacy tradeoff in Internet of Medical Things: a machine learning approach. Future Generation Computer Systems, 98(??):60– 68, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18327195.

Guerrero:2019:EEC

[GLJ19] Carlos Guerrero, Isaac Lera, and Carlos Juiz. Evaluation and efficiency comparison of evolutionary algorithms for service placement optimization in fog architectures. Future Generation Computer Systems, 97(??):131–144, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18325147.

Gallard:2012:ANG

[GLM^{+}12] Jérôme Gallard, Adrien Lèbre, Christine Morin, Thomas Naughton, Stephen L. Scott, and Geoffroy Vallée. Architecture for the next generation system management tools. Future Generation Computer Systems, 28(1):136–146, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001142.

Ghosh:2013:MPA

[GLNT13] Rahul Ghosh, Francesco Longo, Vijay K. Naik, and Kishor S. Trivedi. Modeling and performance analysis of large scale IaaS clouds. Future Generation Computer Systems, 29(5): 1216–1234, July 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001410.

Gonzalez-Lopez:2018:DNN

[GLVC18] Jorge Gonzalez-Lopez, Sebastián Ventura, and Alberto Cano. Distributed nearest neighbor classification for large-scale multilabel data on spark. Future Generation Computer Systems, 87(??):66–82, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17327759.

Guo:2017:MMM

[GLXF17] Qi Guo, Xiaohong Li, Guangquan Xu, and Zhiyong Feng. MP-MID: Multi-protocol oriented middleware-level intrusion detection method for wireless sensor networks. Future Generation Computer Systems, 70(??):42– 47, May 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301959.

Groth:2011:RDS

[GM11] Paul Groth and Luc Moreau. Representing distributed systems using the Open Provenance Model. Future Generation Computer Systems, 27(6):757– 765, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Guerrera:2019:RSC

[GMB19] Danilo Guerrera, Antonio Maffia, and Helmar Burkhart. Reproducible stencil compiler benchmarks using prova! Future Generation Computer Systems, 92(??):933–946, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17304612.

Gurulian:2016:YCT

[GMCM16] Iakovos Gurulian, Konstantinos Markantonakis, Lorenzo Cavallaro, and Keith Mayes. You can’t touch this: Consumercentric Android application repackaging detection. Future Generation Computer Systems, 65(??):1– 9, December 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301327.

See reprint [GMCM18].

Gurulian:2018:RYC

[GMCM18] Iakovos Gurulian, Konstantinos Markantonakis, Lorenzo Cavallaro, and Keith Mayes. Reprint of “You can’t touch this: Consumer-centric Android application repackaging detection”. Future Generation Computer Systems, 80(??):537–545, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17325451.

See [GMCM16].

Gianni:2019:RTR

[GMD19] Francesco Gianni, Simone Mora, and Monica Divitini. RapIoT toolkit: Rapid prototyping of collaborative Internet of Things applications. Future Generation Computer Systems, 95(??):867–879, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17308555.

Gonzalez-Manzano:2017:EHE

[GMdFPLC17] L. González-Manzano, José M. de Fuentes, P. Peris-Lopez, and C. Camara. Encryption by Heart (EbH) — using ECG for time-invariant symmetric key generation. Future Generation Computer Systems, 77(??):136– 148, December 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16307798.

Gidding:2013:ADC

[GML^{+}13] Aaron Gidding, Yuma Matsui, Thomas E. Levy, Tom DeFanti, and Falko Kuester. ArchaeoSTOR: a data curation system for research on the archeological frontier. Future Generation Computer Systems, 29(8):2117–2127, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000678.

Garcia:2017:MDP

[GMLGB^{+}17] Cristian González García, Daniel Meana-Llorián, B. Cristina Pelayo G-Bustelo, Juan Manuel Cueva Lovelle, and Nestor Garcia-Fernandez. Midgar: Detection of people through computer vision in the Internet of Things scenarios to improve the security in Smart Cities, Smart Towns, and Smart Homes. Future Generation Computer Systems, 76(??):301– 313, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308652.

Guerar:2018:CAP

[GMM18] Meriem Guerar, Alessio Merlo, and Mauro Migliardi. Completely automated public physical test to tell computers and humans apart: A usability study on mobile devices. Future Generation Computer Systems, 82(??):617–630, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17303709.

Gabrel:2018:QAA

[GMMM18] Virginie Gabrel, Maude Manouvrier, Kamil Moreau, and Cécile Murat. QoSaware automatic syntactic service composition problem: Complexity and resolution. Future Generation Computer Systems, 80(??): 311–321, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17305575.

Goy:2016:ODC

[GMP^{+}16] Anna Goy, Diego Magro, Giovanna Petrone, Claudia Picardi, and Marino Segnan. Ontology-driven collaborative annotation in shared workspaces. Future Generation Computer Systems, 54(??):435–449, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001053.

Gravina:2017:CBA

[GMP^{+}17] Raffaele Gravina, Congcong Ma, Pasquale Pace, Gianluca Aloi, Wilma Russo, Wenfeng Li, and Giancarlo Fortino. Cloudbased Activity-aaService cyber-physical framework for human activity monitoring in mobility. Future Generation Computer Systems, 75(??):158–171, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303016.

Gravvanis:2018:SIR

[GMP^{+}18] George A. Gravvanis, John P. Morrison, Dana Petcu, Theo Lynn, and Christos K. Filelis-Papadopoulos. Special issue: Recent trends in cloud computing. Future Generation Computer Systems, 79 (part 2)(??):700–702, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324998.

Gunupudi:2017:CSC

[GNGG17] Rajesh Kumar Gunupudi, Mangathayaru Nimmala, Narsimha Gugulothu, and Suresh Reddy Gali. CLAPP: a self constructing feature clustering approach for anomaly detection. Future Generation Computer Systems, 74(??):417–429, September 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16308718.

Gutierrez-Nolasco:2014:TCG

[GNVST14] Sebastian Gutierrez-Nolasco, Nalini Venkatasubramanian, Mark-Oliver Stehr, and Carolyn Talcott. Tailoring consistency in group membership for mobile networks. Future Generation Computer Systems, 31(??):134–146, February 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001283.

Ghanbari:2016:MOM

[GOBL16] Shamsollah Ghanbari, Mohamed Othman, Mohd Rizam Abu Bakar, and Wah June Leong. Multiobjective method for divisible load scheduling in multi-level tree network.

Future Generation Computer Systems, 54(??):132– 143, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000771.

Gopalakrishna:2019:REC

[GOLL19] Aravind Kota Gopalakrishna, Tanir Ozcelebi, Johan J. Lukkien, and Antonio Liotta. Runtime evaluation of cognitive systems for non-deterministic multiple output classification problems. Future Generation Computer Systems, 100(??):1005–1016, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18328358.

Grandinetti:2011:WBP

[GP11] Lucio Grandinetti and Ornella Pisacane. Web based prediction for diabetes treatment. Future Generation Computer Systems, 27(2):139–147, February 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Garcia-Penalvo:2014:ILR

[GPJA^{+}14] Francisco José García-Peñalvo, Mark Johnson, Gustavo Ribeiro Alves, Miroslav Minović, and Miguel Ángel Conde-González. Informal learning recognition through a cloud ecosystem. Future Generation Computer Systems, 32(??):282–294, March 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001696.

Gonzalez-Pardo:2017:ABC

[GPJC17] Antonio Gonzalez-Pardo, Jason J. Jung, and David Camacho. ACO-based clustering for Ego Network analysis. Future Generation Computer Systems, 66(??):160–170, January 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302278.

Grandinetti:2013:ACM

[GPS13] L. Grandinetti, O. Pisacane, and M. Sheikhalishahi. An approximate ϵ-constraint method for a multiobjective job scheduling in the cloud. Future Generation Computer Systems, 29(8):1901–1908, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000836.

Guzzetti:2017:PAE

[GPS^{+}17] Sofia Guzzetti, Tiziano Passerini, Jaroslaw Slawinski, Umberto Villa, Alessandro Veneziani, and Vaidy Sunderam. Platform and algorithm effects on computational fluid dynamics applications in life sciences. Future Generation Computer Systems, 67(??):382–396, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300681.

Goswami:2014:FFD

[GPV^{+}14] Gaurav Goswami, Brian M.

Powell, Mayank Vatsa, Richa Singh, and Afzel Noore. FaceDCAPTCHA: Face detection based color image CAPTCHA. Future Generation Computer Systems, 31(??):59–68, February 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001690.

Guijarro:2019:CDB

[GPVN19] Luis Guijarro, Vicent Pla, Jose R. Vidal, and Maurizio Naldi. Competition in data-based service provision: Nash equilibrium characterization.

Future Generation Computer Systems, 96(??):35– 50, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18315140.

Gai:2018:MBD

[GQLX18] Keke Gai, Meikang Qiu, Meiqin Liu, and Zenggang Xiong. In-memory big data analytics under space constraints using dynamic programming. Future Generation Computer Systems, 83(??):219–227, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17316011.

Gai:2018:PPM

[GQXL18] Keke Gai, Meikang Qiu, Zenggang Xiong, and Meiqin Liu. Privacypreserving multi-channel communication in Edgeof-Things. Future Generation Computer Systems, 85(??):190–200, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18300037.

Graves:2015:TRC

[Gra15]  Alvaro Graves.  Tech-

niques to reduce cluttering of RDF visualizations. Future Generation Computer Systems, 53(??):152– 156, December 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002398.

Glatard:2017:SAI

[GRCP^{+}17] Tristan Glatard, Marc-Étienne Rousseau, Sorina Camarasu-Pop, Reza Adalat, Natacha Beck, Samir Das, Rafael Ferreira da Silva, Najmeh Khalili-Mahani, Vladimir Korkhov, Pierre-Olivier Quirion, Pierre Rioux, Sílvia D. Olabarriaga, Pierre Bellec, and Alan C. Evans. Software architectures to integrate workflow engines in science gateways. Future Generation Computer Systems, 75(??):239–255, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300249.

Giuliani:2011:GES

[GRL11] Gregory Giuliani, Nicolas Ray, and Anthony Lehmann. Grid-enabled Spatial Data Infrastructure for environmental sciences: Challenges and opportunities. Future Generation Computer Systems, 27(3): 292–303, March 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gomez-Romero:2018:VLK

[GRMSOG18] Juan Gómez-Romero, Miguel Molina-Solana, Axel Oehmichen, and Yike Guo. Visualizing large knowledge graphs: a performance analysis. Future Generation Computer Systems, 89(??):224–238, December 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17323610.

Gudigar:2019:AMA

[GRS^{+}19] Anjan Gudigar, U. Raghavendra, Tan Ru San, Edward J. Ciaccio, and U. Rajendra Acharya. Application of multiresolution analysis for automated detection of brain abnormality using MR images: a comparative study. Future Generation Computer Systems, 90(??):359–367, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18314560.

Garbinato:2010:IPA

[GRTV10] Benoît Garbinato, Denis Rochat, Marco Tomassini, and François Vessaz. Injecting power-awareness into epidemic information dissemination in sensor networks. Future Generation Computer Systems, 26(6):868–876, June 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Greco:2019:ESC

[GRX19] Luca Greco, Pierluigi Ritrovato, and Fatos Xhafa. An edge-stream computing infrastructure for realtime analysis of wearable sensors data. Future Generation Computer Systems, 93(??):515–528, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18314031.

Galizia:2019:UAA

[GRZ^{+}19] Antonella Galizia, Luca Roverelli, Gabriele Zereik, Emanuele Danovaro, Andrea Clematis, and Daniele D’Agostino. Using Apache Airavata and EasyGateway for the creation of complex science gateway frontend. Future Generation Computer Systems, 94(??): 910–919, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17310671.

Garofalakis:2013:AMP

[GS13] John Garofalakis and Eleftherios Stergiou. An analytical model for the performance evaluation of multistage interconnection networks with two class priorities. Future Generation Computer Systems, 29(1):114–129, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001100.

Giorgi:2015:STS

[GS15] R. Giorgi and A. Scionti. A scalable thread scheduling co-processor based on data-flow principles. Future Generation Computer Systems, 53(??):100–108, December 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1400274X.

Gil:2016:MMB

[GS16a] David Gil and Il-Yeol Song. Modeling and management of Big Data: Challenges and opportunities. Future Generation Computer Systems, 63(??):96–99, October 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002514.

Gill:2016:DCA

[GS16b] Navneet Kaur Gill and Sarbjeet Singh. A dynamic, cost-aware, optimized data replication strategy for heterogeneous cloud data centers. Future Generation Computer Systems, 65(??):10–32, December 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301273.

Ghaeb:2011:HPD

[GSC11] J. A. Ghaeb, M. A. Smadi, and J. Chebil. A high performance data integrity assurance based on the determinant technique. Future Generation Computer Systems, 27(5):614–619, May 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Guo:2019:DNN

[GSC^{+}19] Jie Guo, Bin Song, Yuhao Chi, Lahiru Jayasinghe, Chau Yuen, Yong Liang Guan, Xiaojiang Du, and Mohsen Guizani. Deep neural network-aided Gaussian message passing detection for ultra-reliable low-latency communications. Future Generation Computer Systems, 95(??): 629–638, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18318028.

Garcia-Silva:2019:EFR

[GSGPP^{+}19] Andres Garcia-Silva, Jose Manuel Gomez-Perez, Raul Palma, Marcin Krystek, Simone Mantovani, Federica Foglini, Valentina Grande, Francesco De Leo, Stefano Salvi, Elisa Trasatti, Vito Romaniello, Mirko Albani, Cristiano Silvagni, Rosemarie Leone, Fulvio Marelli, Sergio Albani, Michele Lazzarini, Hazel J. Napier, Helen M. Glaves, Timothy Aldridge, Charles Meertens, Fran Boler, Henry W. Loescher, Christine Laney, Melissa A. Genazzio, Daniel Crawl, and Ilkay Altintas. Enabling FAIR research in Earth Science through research objects. Future Generation Computer Systems, 98(??):550–564, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314638.

Ghanbari:2012:FBO

[GSLI12] Hamoun Ghanbari, Bradley Simmons, Marin Litoiu, and Gabriel Iszlai. Feedbackbased optimization of a private cloud. Future Generation Computer Systems, 28(1):104–111, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001014.

Gupta:2018:RRC

[GSN^{+}18] B. B. Gupta, A. K.

Sangaiah, Nadia Nedjah, Shingo Yamaguchi, Zhiyong Zhang, and Michael Sheng. Recent research in computational intelligence paradigms into security and privacy for online social networks (OSNs). Future Generation Computer Systems, 86(??):851–854, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18311233.

Gomes:2017:ECM

[GSP^{+}17] Andre S. Gomes, Bruno Sousa, David Palma, Vitor Fonseca, Zhongliang Zhao, Edmundo Monteiro, Torsten Braun, Paulo Simoes, and Luis Cordeiro. Edge caching with mobility prediction in virtualized LTE mobile networks. Future Generation Computer Systems, 70(??):148– 162, May 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302072.

Guerine:2019:PBH

[GSR^{+}19] Marcos Guerine, Murilo B. Stockinger, Isabel Rosseti, Luidi G. Simonetti, Kary A. C. S. Ocaña, Alexandre Plastino, and Daniel de Oliveira. A provenancebased heuristic for preserving results confidentiality in cloud-based scientific workflows. Future Generation Computer Systems, 97(??):697–713, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18307787.

Glatard:2010:LSF

[GSV^{+}10] Tristan Glatard, Remi S. Soleman, Dick J. Veltman, Aart J. Nederveen, and Sílvia D. Olabarriaga. Large-scale functional MRI study on a production grid. Future Generation Computer Systems, 26(4):685– 692, April 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Guo:2017:ODA

[GSY^{+}17] Jie Guo, Bin Song, Fei Richard Yu, Zheng Yan, and Laurence T. Yang. Object detection among multimedia big data in the compressive measurement domain under mobile distributed architecture. Future Generation Computer Systems, 76(??):519–527, November 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17303333.

Guan:2019:EMF

[GSY^{+}19] Jianfeng Guan, Vishal Sharma, Ilsun You, Mohammed Atiquzzaman, and Muhammad Imran. Extension of MIH for FP-MIPv6 (EMIH-FPMIPv6) to support optimized heterogeneous handover. Future Generation Computer Systems, 97(??):775–791, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18323410.

Gracia-Tinedo:2018:GWY

[GTCZG^{+}18] Raúl Gracia-Tinedo, Cristian Cotes, Edgar Zamora-Gómez, Genís Ortiz, Adrián Moreno-Martínez, Marc Sánchez-Artigas, Pedro García-López, Raquel Sánchez, Alberto Gómez, and Anastasio Illana. Giving wings to your data: a first experience of personal cloud interoperability. Future Generation Computer Systems, 78 (part 3)(??):1055– 1070, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301334.

Gad:2018:IRU

[GTEL^{+}18] Ramadan Gad, Muhammad Talha, Ahmed A. Abd El-Latif, M. Zorkany, Ayman El-Sayed, Nawal El-Fishawy, and Ghulam Muhammad. Iris recognition using multialgorithmic approaches for Cognitive Internet of Things (CIoT) framework. Future Generation Computer Systems, 89(??):178– 191, December 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18306915.

Gutierrez:2019:DOS

[GTM19] Yoan Gutiérrez, David Tomás, and Isabel Moreno. Developing an ontology schema for enriching and linking digital media assets. Future Generation Computer Systems, 101(??):381–397, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18323859.

Guo:2017:ODF

[GTMZ17] Kehua Guo, Yayuan Tang, Jianhua Ma, and Yaoxue Zhang. Optimized dependent file fetch middleware in transparent computing platform. Future Generation Computer Systems, 74(??):199–207, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003234.

Gracia-Tinedo:2014:GFS

[GTSAR^{+}14] Raúl Gracia-Tinedo, Marc Sánchez-Artigas, Aleix Ramírez, Adrián Moreno-Martínez, Xavier León, and Pedro García-López. Giving form to social cloud storage through experimentation: Issues and insights. Future Generation Computer Systems, 40(??): 1–16, November 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000958.

Gracia-Tinedo:2019:SDO

[GTSP^{+}19] Raúl Gracia-Tinedo, Josep Sampé, Gerard París, Marc Sánchez-Artigas, Pedro García-López, and Yosef Moatti. Software-defined object storage in multitenant environments. Future Generation Computer Systems, 99(??):54–72, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18322167.

Gogos:2016:SIT

[GVA^{+}16] Christos Gogos, Christos Valouxis, Panayiotis Alefragis, George Goulas, Nikolaos Voros, and Efthymios Housos. Scheduling independent tasks on heterogeneous processors using heuristics and Column Pricing. Future Generation Computer Systems, 60(??):48– 66, July 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16000297.

Garg:2013:FRC

[GVB13] Saurabh Kumar Garg, Steve Versteeg, and Rajkumar Buyya. A framework for ranking of cloud computing services. Future Generation Computer Systems, 29(4):1012–1023, June 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001422.

Garcia-Valls:2017:RST

[GVBG17] Marisol García-Valls, Paolo Bellavista, and Aniruddha Gokhale. Reliable software technologies and communication middleware: a perspective and evolution directions for cyberphysical system, mobility, and cloud computing. Future Generation Computer Systems, 71(??):171–176, June 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302819.

Gommans:2015:SPG

[GVdBdL15] Leon Gommans, John Vollbrecht, Betty Gommans de Bruijn, and Cees de Laat. The Service Provider Group framework: a framework for arranging trust and power to facilitate authorization of network services. Future Generation Computer Systems, 45(??):176–192, April 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1400123X.

Gadiraju:2016:BPM

[GVDT16] Krishna Karthik Gadiraju, Manik Verma, Karen C. Davis, and Paul G. Talaga. Benchmarking performance for migrating a relational application to a parallel implementation. Future Generation Computer Systems, 63(??):148– 156, October 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003994.

Gerofi:2013:UMC

[GVI13] Balazs Gerofi, Zoltan Vass, and Yutaka Ishikawa. Utilizing memory content similarity for improving the performance of highly available virtual machines. Future Generation Computer Systems, 29(4):1085– 1095, June 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001446.

Grosso:2018:EIS

[GVTdL18] Paola Grosso, Malathi Veeraraghavan, Brian Tierney, and Cees de Laat. Editorial: INDIS special section FGCS. Future Generation Computer Systems, 79 (part 1)(??):166– 168, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324901.

Garcia-Valls:2014:LCR

[GVURIVBV14] Marisol García-Valls, Patricia Uriol-Resuela, Felipe Ibáñez-Vázquez, and Pablo Basanta-Val. Low complexity reconfiguration for real-time data-intensive service-oriented applications. Future Generation Computer Systems, 37(??): 191–200, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002379.

Guo:2016:PEF

[GWC^{+}16] Rui Guo, Hongzhi Wang, Mengwen Chen, Jianzhong Li, and Hong Gao. Parallelizing the extraction of fresh information from online social networks. Future Generation Computer Systems, 59(??):33– 46, June 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003787.

Gao:2019:CIC

[GWW^{+}19] Z. Gao, D. Y. Wang, S. H.

Wan, H. Zhang, and Y. L. Wang. Cognitive-inspired class-statistic matching with triple-constrain for camera free 3D object retrieval. Future Generation Computer Systems, 94(??): 641–653, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18323549.

Gao:2012:SCN

[GXL^{+}12] Jingcheng Gao, Yang Xiao, Jing Liu, Wei Liang, and C. L. Philip Chen. A survey of communication/networking in Smart Grids. Future Generation Computer Systems, 28(2):391–404, February 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000653.

Guo:2018:BFT

[GXL^{+}18] Zehua Guo, Yang Xu, Ruoyan Liu, Andrey Gushchin, Kuan yin Chen, Anwar Walid, and H. Jonathan Chao. Balancing flow table occupancy and link utilization in software-defined networks. Future Generation Computer Systems, 89(??):213–223, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18306666.

Gao:2019:AMT

[GXW^{+}19] Lan Gao, Yunlong Xu, Rui Wang, Hailong Yang, Zhongzhi Luan, and Depei Qian. Accelerating inmemory transaction processing using general purpose graphics processing units. Future Generation Computer Systems, 97(??): 836–848, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18328681.

Guo:2018:SHG

[GZL^{+}18] Hua Guo, Yandong Zheng, Xiong Li, Zhoujun Li, and Chunhe Xia. Self-healing group key distribution protocol in wireless sensor networks for secure IoT communications. Future Generation Computer Systems, 89(??):713–721, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18304102.

Gao:2016:SCP

[GZLZ16] Shuai Gao, Yujing Zeng, Hongbin Luo, and Hongke Zhang. Scalable control plane for intra-domain communication in software defined information centric networking. Future Generation Computer Systems, 56(??):110–120, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003301.

Guo:2019:ODP

[GZQ^{+}19] Wei Guo, Hua Zhang, Sujuan Qin, Fei Gao, Zhengping Jin, Wenmin Li, and Qiaoyan Wen. Outsourced dynamic provable data possession with batch update for secure cloud storage. Future Generation Computer Systems, 95(??): 309–322, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18319861.

Gassler:2014:GCS

[GZS14] Wolfgang Gassler, Eva Zangerle, and Günther Specht. Guided curation of semistructured data in collaborativelybuilt knowledge bases. Future Generation Computer Systems, 31(??):111–119, February 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001076.

Gong:2018:RAM

[GZW18] Bei Gong, Yu Zhang, and Yubo Wang. A remote attestation mechanism for the sensing layer nodes of the Internet of Things. Future Generation Computer Systems, 78 (part 3)(??):867–886, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17315352.

Gunarathne:2013:SPC

[GZWQ13] Thilina Gunarathne, Bingjing Zhang, Tak-Lon Wu, and Judy Qiu. Scalable parallel computing on clouds using Twister4Azure iterative MapReduce. Future Generation Computer Systems, 29(4):1035–1048, June 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001379.

Guo:2018:AOM

[GZZ^{+}18] Caiping Guo, Pengcheng Zhang, Liyuan Zhang, Zhiguo Gui, and Huazhong Shu. Application of optimization model with piecewise penalty to intensitymodulated radiation therapy. Future Generation Computer Systems, 81(??):

280–290, April 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17313201.

Haile:2016:VCS

[HA16] Netsanet Haile and Jörn Altmann. Value creation in software service platforms. Future Generation Computer Systems, 55(??):495– 509, February 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1500309X.

Haile:2018:EIP

[HA18] Netsanet Haile and Jörn Altmann. Evaluating investments in portability and interoperability between software service platforms. Future Generation Computer Systems, 78 (part 1)(??):224–241, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17307719.

Hamzelou:2019:MSP

[HA19] Nasrin Hamzelou and Mehrdad Ashtiani. A mitigation strategy for the prevention of cascading trust failures in social networks. Future Generation Computer Systems, 94(??):564– 586, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18305065.

Huda:2016:HSV

[HAA^{+}16] Shamsul Huda, Jemal Abawajy, Mamoun Alazab, Mali Abdollalihian, Rafiqul Islam, and John Yearwood. Hybrids of support vector machine wrapper and filter based framework for malware detection. Future Generation Computer Systems, 55(??):376–390, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001228.

Huda:2019:AEI

[HAAR^{+}19] Shamsul Huda, Jemal Abawajy, Baker Al-Rubaie, Lei Pan, and Mohammad Mehedi Hassan. Automatic extraction and integration of behavioural indicators of malware for protection of cyber-physical networks. Future Generation Computer Systems, 101(??):1247–1258, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19306776.

Hamid:2018:KDS

[HAAWH^{+}18] Md. Abdul Hamid, M. Abdullah-Al-Wadud, Mohammad Mehedi Hassan, Ahmad Almogren, Atif Alamri, Abu Raihan M. Kamal, and Md. Mamun-Or-Rashid. A key distribution scheme for secure communication in acoustic sensor networks. Future Generation Computer Systems, 86(??):1209–1217, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17315212.

Hanford:2016:INP

[HAF^{+}16] Nathan Hanford, Vishal Ahuja, Matthew Farrens, Dipak Ghosal, Mehmet Balman, Eric Pouyoul, and Brian Tierney. Improving network performance on multicore systems: Impact of core affinities on high throughput flows. Future Generation Computer Systems, 56(??):277–283, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002927.

Habib:2019:SPB

[HAJ^{+}19] Muhammad Asif Habib, Mudassar Ahmad, Sohail Jabbar, Shehzad Khalid, Junaid Chaudhry, Kashif Saleem, Joel J. P. C. Rodrigues, and Muhammad Sayim Khalil. Security and privacy based access control model for Internet of Connected Vehicles. Future Generation Computer Systems, 97(??): 687–696, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831656X.

Hamid:2017:MDA

[Ham17] Brahim Hamid. A modeldriven approach for developing a model repository: Methodology and tool support. Future Generation Computer Systems, 68(??): 473–490, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301042.

Hossain:2018:CAM

[HAM18] M. Shamim Hossain, Mohammed F. Alhamid, and Ghulam Muhammad. Collaborative analysis model for trending images on social networks. Future Generation Computer Systems, 86(??):855–862, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301383.

Hamidi:2019:ADS

[Ham19] Hodjat Hamidi. An approach to develop the smart health using Internet of Things and authentication based on biometric technology. Future Generation Computer Systems, 91(??):434–449, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18313517.

Haque:2011:SEM

[HAP11] Aminul Haque, Saadat M. Alhashmi, and Rajendran Parthiban. A survey of economic models in grid computing. Future Generation Computer Systems, 27(8): 1056–1069, October 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Haque:2015:OBA

[HAP15] Aminul Haque, Saadat M. Alhashmi, and Rajendran Parthiban. An optimization-based adaptive resource management framework for economic grids: a switching mechanism. Future Generation Computer Systems, 47(??): 48–59, June 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002167.

Hasan:2019:SBD

[HAT19] Mohammed Zaki Hasan and Fadi Al-Turjman. SWARM-based data delivery in Social Internet of Things. Future Generation Computer Systems, 92(??):821–836, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312281.

Heidari:2019:QSQ

[HB19] Safiollah Heidari and Rajkumar Buyya. Quality of service (QoS)- driven resource provisioning for large-scale graph processing in cloud computing environments: Graph Processing-as-a-Service (GPaaS). Future Generation Computer Systems, 96(??):490–501, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314109.

Huang:2013:EDD

[HBN^{+}13] Ye Huang, Nik Bessis, Peter Norrington, Pierre Kuonen, and Beat Hirsbrunner. Exploring decentralized dynamic scheduling for grids and clouds using the community-aware scheduling algorithm. Future Generation Computer Systems, 29(1):402–415, January 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11000872.

Han:2017:SSP

[HC17] Son N. Han and Noel Crespi. Semantic service provisioning for smart objects: Integrating IoT applications into the web. Future Generation Computer Systems, 76(??):180– 197, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16308676.

Higashino:2016:CMS

[HCB16] Wilson A. Higashino, Miriam A. M. Capretz, and Luiz F. Bittencourt. CEP-Sim: Modelling and simulation of complex event processing systems in cloud environments. Future Generation Computer Systems, 65(??):122–139, December 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003362.

Han:2014:ESR

[HCC^{+}14] Xiao Han, Ángel Cuevas, Noël Crespi, Rubén Cuevas, and Xiaodi Huang. On exploiting social relationship and personal background for content discovery in P2P networks. Future Generation Computer Systems, 40(??):17– 29, November 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001290.

Huynh:2019:ECO

[HCHH19] De-Thu Huynh, Min Chen, Trong-Thua Huynh, and Chu Hong Hai. Energy consumption optimization for green deviceto-device multimedia communications. Future Generation Computer Systems, 92(??):1131–1141, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17304405.

He:2014:DSA

[HCJ14] Ligang He, Nadeem Chaudhary, and Stephen A. Jarvis. Developing securityaware resource management strategies for workflows. Future Generation Computer Systems, 38(??):61–68, September 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002094.

Hu:2017:TPS

[HCL^{+}17] Guangwu Hu, Wenlong Chen, Qi Li, Yong Jiang, and Ke Xu. TrueID: a practical solution to enhance Internet accountability by assigning packets with creditable user identity code. Future Generation Computer Systems, 72(??): 219–226, July 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16303004.

Hao:2019:AEA

[HCMJ19] Yongsheng Hao, Jie Cao, Tinghuai Ma, and Sai Ji. Adaptive energy-aware scheduling method in a meteorological cloud. Future Generation Computer Systems, 101(??):1142–1157, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19309471.

He:2014:ICN

[HCNT14] Xiangjian He, Thawatchai Chomsiri, Priyadarsi Nanda, and Zhiyuan Tan. Improving cloud network security using the Tree-Rule firewall. Future Generation Computer Systems, 30(??):116–126, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001386.

Huang:2018:BBF

[HCW^{+}18] Hui Huang, Xiaofeng Chen, Qianhong Wu, Xinyi Huang, and Jian Shen. Bitcoin-based fair payments for outsourcing computations of fog devices. Future Generation Computer Systems, 78 (part 2)(??):850–858, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16307579.

Huang:2019:DDO

[HCX^{+}19] Chengxi Huang, Hongming Cai, Lida Xu, Boyi Xu, Yizhi Gu, and Lihong Jiang. Data-driven ontology generation and evolution towards intelligent service in manufacturing systems. Future Generation Computer Systems, 101(??):197–207, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317904.

He:2017:DPC

[HCZW17] Hui He, Lijie Cui, Fenglan Zhou, and Dong Wang. Distributed proxy cache technology based on autonomic computing in smart cities. Future Generation Computer Systems, 76(??):370–383, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300590.

Homayoun:2019:DDR

[HDA^{+}19] Sajad Homayoun, Ali Dehghantanha, Marzieh Ahmadzadeh, Sattar Hashemi, Raouf Khayami, Kim-Kwang Raymond Choo, and David Ellis Newton. DRTHIS: Deep ransomware threat hunting and intelligence system at the fog layer. Future Generation Computer Systems, 90(??):94–104, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17328467.

Huang:2018:CCH

[HDB18] Shan Huang, Dezun Dong, and Wei Bai. Congestion control in high-speed lossless data center networks: a survey. Future Generation Computer Systems, 89(??):360–374, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18300323.

Hu:2018:CIP

[HDH^{+}18] Enwen Hu, Zhongliang Deng, Mudan Hu, Lu Yin, and Wen Liu. Cooperative indoor positioning with factor graph based on FIM for wireless sensor network. Future Generation Computer Systems, 89(??):126– 136, December 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18306885.

HaddadPajouh:2018:DRN

[HDKC18] Hamed HaddadPajouh, Ali Dehghantanha, Raouf Khayami, and Kim-Kwang Raymond Choo. A deep recurrent neural network based approach for Internet of Things malware threat hunting. Future Generation Computer Systems, 85(??):88–96, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1732486X.

Hu:2013:EIC

[HDLW13] Yusuo Hu, Dafan Dong, Jiang Li, and Feng Wu. Efficient and incentivecompatible resource allocation mechanism for P2P-assisted content delivery systems. Future Generation Computer Systems, 29(6):1611–1620, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001641.

Harshan:2016:DEC

[HDO16] J. Harshan, Anwitaman Datta, and Frédérique Oggier. DiVers: an erasure code based storage architecture for versioning exploiting sparsity. Future Generation Computer Systems, 59(??):47– 62, June 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16000078.

Hassan:2019:HRT

[HEES19] Mohammed K. Hassan, Ali I. El Desouky, Sally M. Elghamrawy, and Amany M. Sarhan. A hybrid real-time remote monitoring framework with NB-WOA algorithm for patients with chronic diseases. Future Generation Computer Systems, 93(??):77–95, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17326596.

Helfert:2016:NHC

[Hel16] Markus Helfert. New horizons of cloud computing. Future Generation Computer Systems, 55(??):163– 164, February 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003490.

Hong:2019:IAM

[HFL^{+}19] Zhaoxi Hong, Yixiong Feng, Zhiwu Li, Yong Wang, Hao Zheng, Zhongkai Li, and Jianrong Tan. An integrated approach for multi-objective optimisation and MCDM of energy Internet under uncertainty. Future Generation Computer Systems, 97(??): 90–104, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18310112.

Habibi:2019:EDR

[HFM19] Moslem Habibi, MohammadAmin Fazli, and Ali Movaghar. Efficient distribution of requests in federated cloud computing environments utilizing statistical multiplexing. Future Generation Computer Systems, 90(??):451–460, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18309038.

Huang:2016:EDP

[HFT16] Shi-Yuan Huang, Chun-I Fan, and Yi-Fan Tseng. Enabled/disabled predicate encryption in clouds. Future Generation Computer Systems, 62(??):148– 160, September 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003921.

Han:2014:ECA

[HGG^{+}14] Rui Han, Moustafa M. Ghanem, Li Guo, Yike Guo, and Michelle Osmond. Enabling cost-aware and adaptive elasticity of multi-tier cloud applications. Future Generation Computer Systems, 32(??): 82–98, March 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001148.

Hinrich:2015:CRU

[HGM15] P. Hinrich, P. Grosso, and I. S. Monga. Collaborative research using eScience infrastructure and high speed networks. Future Generation Computer Systems, 45 (??):161, April 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1400260X.

Hejja:2019:EIT

[HH19] Khaled Hejja and Xavier Hesselbach. Evaluating impacts of traffic migration and virtual network functions consolidation on power aware resource allocation algorithms. Future Generation Computer Systems, 101(??):83–98, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304108.

He:2012:MAI

[HHD^{+}12] Ligang He, Chenlin Huang, Kewei Duan, Kenli Li, Hao Chen, Jianhua Sun, and Stephen A. Jarvis. Modeling and analyzing the impact of authorization on workflow executions. Future Generation Computer Systems, 28(8):1177–1193, October 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000568.

Hassan:2019:NCD

[HHH^{+}19] Md Rafiul Hassan, Md Sarwar M. Haque, Muhammad Imtiaz Hossain, Mohammad Mehedi Hassan, and Abdulhameed Alelaiwi. A novel cascaded deep neural network for analyzing smart phone data for indoor localization. Future Generation Computer Systems, 101(??):760–769, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19308696.

Henze:2016:CAP

[HHK^{+}16] Martin Henze, Lars Hermerschmidt, Daniel Kerpen, Roger Häußling, Bernhard Rumpe, and Klaus Wehrle. A comprehensive approach to privacy in the cloudbased Internet of Things. Future Generation Computer Systems, 56(??):701– 718, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002964.

Hasan:2018:AID

[HHK18] Ragib Hasan, Mahmud Hossain, and Rasib Khan. Aura: An incentive-driven ad-hoc IoT cloud framework for proximal mobile computation offloading. Future Generation Computer Systems, 86(??):821–835, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1732602X.

Hamdaqa:2011:ABC

[HHL11] Mohammad Hamdaqa and Abdelwahab Hamou-Lhadj. An approach based on citation analysis to support effective handling of regulatory compliance. Future Generation Computer Systems, 27(4):395–410, April 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hashim:2019:HGS

[HHM^{+}19] Fatma A. Hashim, Essam H. Houssein, Mai S. Mabrouk, Walid Al-Atabany, and Seyedali Mirjalili. Henry gas solubility optimization: a novel physicsbased algorithm. Future Generation Computer Systems, 101(??):646–667, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19306557.

Hussain:2018:CTS

[HHS^{+}18] Walayat Hussain, Farookh Khadeer Hussain, Morteza Saberi, Omar Khadeer Hussain, and Elizabeth Chang. Comparing time series with machine learningbased prediction approaches for violation management in cloud SLAs. Future Generation Computer Systems, 89(??):464–477, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17330406.

Hsu:2011:OSW

[HHW11] Chih-Chiang Hsu, Kuo-Chan Huang, and Feng-Jian Wang. Online scheduling of workflow applications in grid environments. Future Generation Computer Systems, 27(6):860– 870, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

He:2019:SBR

[HHW^{+}19] Yu He, Guangjie Han, Hao Wang, James Adu Ansere, and Whenbo Zhang. A sector-based random routing scheme for protecting the source location privacy in WSNs for the Internet of Things. Future Generation Computer Systems, 96(??): 438–448, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18316406.

Hou:2013:BMP

[HHXL13] Fangyong Hou, Hongjun He, Nong Xiao, and Fang Liu. Bus and memory protection through chaingenerated and tree-verified IV for multiprocessors systems. Future Generation Computer Systems, 29(3): 901–912, March 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000817.

Huang:2016:NVN

[HHZ16] Mao Lin Huang, Tze-Haw Huang, and Xuyun Zhang. A novel virtual node approach for interactive visual analytics of big datasets in parallel coordinates. Future Generation Computer Systems, 55(??):510–523, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000382.

Hussain:2019:IVS

[HHZ19] Rasheed Hussain, Fatima Hussain, and Sherali Zeadally. Integration of VANET and 5G security: a review of design and implementation issues. Future Generation Computer Systems, 101(??):843–864, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19306909.

Habiba:2018:CBU

[HIA18a] Mansura Habiba, Md. Rafiqul Islam, and A. B. M. Shawkat Ali. A component based unified architecture for utility service in cloud. Future Generation Computer Systems, 87(??):725–742, October 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17302935.

Hossain:2018:ITB

[HIA^{+}18b] Mahmud Hossain, S. M. Riazul Islam, Farman Ali, Kyung-Sup Kwak, and Ragib Hasan. An Internet of Things-based health prescription assistant and its security system design. Future Generation Computer Systems, 82(??):422– 439, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17314085.

Huda:2018:HMF

[HIA^{+}18c] Shamsul Huda, Rafiqul Islam, Jemal Abawajy, John Yearwood, Mohammad Mehedi Hassan, and Giancarlo Fortino. A hybrid-multi filter-wrapper framework to identify runtime behaviour for fast malware detection. Future Generation Computer Systems, 83(??):193–207, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17325049.

Haoyu:2019:ICB

[HJA^{+}19] Li Haoyu, Li Jianxing, N. Arunkumar, Ahmed Faeq Hussein, and Mustafa Musa Jaber. An IoMT cloudbased real time sleep apnea detection scheme by using the SpO2 estimation supported by heart rate variability. Future Generation Computer Systems, 98(??):69–77, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18326980.

Hsu:2010:SSP

[HJC10] Ching-Hsien Hsu, Hai Jin, and Franck Cappello. Special section: Peer-to-Peer Grid technologies. Future Generation Computer Systems, 26(5):701–703, May 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hameed:2018:TFV

[HKA^{+}18] Khizar Hameed, Abid Khan, Mansoor Ahmed, Alavalapati Goutham Reddy, and M. Mazhar Rathore. Towards a formally verified zero watermarking scheme for data integrity in the Internet of Things basedwireless sensor networks. Future Generation Computer Systems, 82(??):274– 289, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17322756.

Kim:2011:SCC

[hKBB11] Tai hoon Kim, Debnath Bhattacharyya, and Samir Kumar Bandyopadhyay. Supervised chromosome clustering and image classification. Future Generation Computer Systems, 27(4):372–376, April 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hussain:2016:OBM

[HKG^{+}16] Iftikhar Hussain, Luk Knapen, Stéphane Galland, Ansar-Ul-Haque Yasar, Tom Bellemans, Davy Janssens, and Geert Wets. Organizational-based model and agent-based simulation for long-term carpooling. Future Generation Computer Systems, 64(??):125– 139, November 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300322.

Hwang:2010:MRO

[HKP10] Junseok Hwang, Hak-Jin Kim, and Jihyoun Park. Managing risks in an open computing environment using mean absolute deviation portfolio optimization. Future Generation Computer Systems, 26(8): 1381–1390, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Herrero:2010:SSG

[HKPT10] Pilar Herrero, Daniel S. Katz, María S. Pérez, and Domenico Talia. Special section: Grid computing, high-performance and distributed applications. Future Generation Computer Systems, 26(2): 257–258, February 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kim:2017:SCI

[hKRM17] Tai hoon Kim, Carlos Ramos, and Sabah Mohammed. Smart City and IoT. Future Generation Computer Systems, 76(??):159–162, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17305253.

Hsu:2018:CCA

[HKS18] Ta-Yuan Hsu, Ajay D. Kshemkalyani, and Min Shen. Causal consistency algorithms for partially replicated and fully replicated systems. Future Generation Computer Systems, 86(??):1118–1133, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17308166.

He:2019:LLP

[HKT^{+}19] Zaihong He, Jishun Kuang, Yanjie Tan, Shihui Peng, and Huailiang Tan. LPCMsim: a lightweight phase change memory simulator. Future Generation Computer Systems, 97(??):661– 673, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18311890.

Halak:2011:RTL

[HKU^{+}11] Jirí Halák, Michal Krsek, Sven Ubik, Petr Zejdl, and Felix Nevrela. Realtime long-distance transfer of uncompressed 4K video for remote collaboration. Future Generation Computer Systems, 27(7):886–892, July 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hu:2016:PBR

[HLC16] Yu-Chen Hu, Chun-Chi Lo, and Wu-Lin Chen. Probability-based reversible image authentication scheme for image demosaicking. Future Generation Computer Systems, 62(??):92– 103, September 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630070X.

Hu:2016:VDM

[HLCL16] Changjun Hu, Yang Li, Xin Cheng, and Zhenyu Liu. A Virtual Dataspaces Model for large-scale materials scientific data access. Future Generation Computer Systems, 54(??):456– 468, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1500179X.

Han:2018:OBM

[HLH^{+}18] Lansheng Han, Songsong Liu, Shuxia Han, Wenjing Jia, and Jingwei Lei. Owner based malware discrimination. Future Generation Computer Systems, 80(??):496–504, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301315.

Hu:2011:RRR

[HLL^{+}11] Jinwei Hu, Ruixuan Li, Zhengding Lu, Jianfeng Lu, and Xiaopu Ma. RAR: a role-and-risk based flexible framework for secure collaboration. Future Generation Computer Systems, 27(5):574–586, May 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Huang:2012:ESS

[HLL12] Hong Huang, Jianwei Li, and Jiamin Liu. Enhanced semi-supervised local Fisher discriminant analysis for face recognition. Future Generation Computer Systems, 28(1):244–253, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X10002153.

Hu:2017:SAA

[HLL^{+}17] Yikun Hu, Chubo Liu, Kenli Li, Xuedi Chen, and Keqin Li. Slack allocation algorithm for energy minimization in cluster systems. Future Generation Computer Systems, 74(??):119–131, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302941.

Han:2018:NRA

[HLL18] Zhijie Han, Yaqiong Li, and Jie Li. A novel routing algorithm for IoT cloud based on hash offset tree. Future Generation Computer Systems, 86(??):456–463, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17324512.

Hu:2019:JDR

[HLL^{+}19] Menglan Hu, Weidong Liu, Junqiu Lu, Rui Fu, Kai Peng, Xiaoqiang Ma, and Jiangchuan Liu. On the joint design of routing and scheduling for Vehicle-Assisted Multi-UAV inspection. Future Generation Computer Systems, 94(??):214–223, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18324567.

Huh:2011:MAW

[HLNM11] Jun Ho Huh, John Lyle, Cornelius Namiluko, and Andrew Martin. Managing application whitelists in trusted distributed systems. Future Generation Computer Systems, 27(2): 211–226, February 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Huang:2018:BEI

[HLT^{+}18] Zhenqiu Huang, Kwei-Jay Lin, Bo-Lung Tsai, Surong Yan, and Chi-Sheng Shih. Building edge intelligence for online activity recognition in service-oriented IoT systems. Future Generation Computer Systems, 87(??):557–567, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17306520.

Huang:2019:SME

[HLT^{+}19] Binbin Huang, Zhongjin Li, Peng Tang, Shangguang Wang, Jun Zhao, Haiyang Hu, Wanqing Li, and Victor Chang. Security modeling and efficient computation offloading for service workflow in mobile edge computing. Future Generation Computer Systems, 97(??):755–774, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326773.

Hu:2016:RCP

[HLV^{+}16] Fei Hu, Yu Lu, Athanasios V. Vasilakos, Qi Hao, Rui Ma, Yogendra Patil, Ting Zhang, Jiang Lu, Xin Li, and Neal N. Xiong. Robust Cyber-Physical Systems: Concept, models, and implementation. Future Generation Computer Systems, 56(??):449–475, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002071.

Hao:2012:ELB

[HLW12] Yongsheng Hao, Guanfeng Liu, and Na Wen. An enhanced load balancing mechanism based on deadline control on GridSim. Future Generation Computer Systems, 28(4):657– 665, April 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001956.

Hassan:2017:MHD

[HLYW17] Mohammad Mehedi Hassan, Kai Lin, Xuejun Yue, and Jiafu Wan. A multimedia healthcare data sharing approach through cloudbased body area network. Future Generation Computer Systems, 66(??):48– 58, January 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15004070.

Huo:2018:IMC

[HLZ18] Jiuyuan Huo, Liqun Liu, and Yaonan Zhang. An improved multi-cores parallel artificial bee colony optimization algorithm for parameters calibration of hydrological model. Future Generation Computer Systems, 81(??):492–504, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17304417.

Han:2019:HRM

[HLZ^{+}19] Guanghui Han, Xiabi Liu, Heye Zhang, Guangyuan Zheng, Nouman Qadeer Soomro, Murong Wang, and Weihua Liu. Hybrid resampling and multi-feature fusion for automatic recognition of cavity imaging sign in lung CT. Future Generation Computer Systems, 99(??):558–570, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19306806.

Helfert:2019:VES

[HM19] Markus Helfert and Víctor Méndez Muñoz. VSI: Emerging services cloud. Future Generation Computer Systems, 95(??):500–501, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19301153.

Hossain:2018:CAS

[HMA^{+}18a] M. Shamim Hossain, Ghulam Muhammad, Wadood Abdul, Biao Song, and B. B. Gupta. Cloudassisted secure video transmission and sharing framework for smart cities. Future Generation Computer Systems, 83(??):596–606, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17305198.

Hossain:2018:ICS

[HMA18b] M. Shamim Hossain, Ghulam Muhammad, and Syed Umar Amin. Improving consumer satisfaction in smart cities using edge computing and caching: a case study of date fruits classification. Future Generation Computer Systems, 88(??):333– 341, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1733056X.

Hosseinpour:2019:SSI

[HMC19] Mohammad Hosseinpour, Mohammad Reza Malek, and Christophe Claramunt. Socio-spatial influence maximization in location-based social networks. Future Generation Computer Systems, 101(??):304–314, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19305849.

Heidari:2019:HHO

[HMF^{+}19] Ali Asghar Heidari, Seyedali Mirjalili, Hossam Faris, Ibrahim Aljarah, Majdi Mafarja, and Huiling Chen. Harris hawks optimization: Algorithm and applications. Future Generation Computer Systems, 97(??): 849–872, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18313530.

Hallawi:2017:MCC

[HMH17] Huda Hallawi, Jörn Mehnen, and Hongmei He. Multicapacity combinatorial ordering GA in application to cloud resources allocation and efficient virtual machines consolidation. Future Generation Computer Systems, 69(??): 1–10, April 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16304630.

Hasham:2018:CIP

[HMM18] Khawar Hasham, Kamran Munir, and Richard Mc-Clatchey. Cloud infrastructure provenance collection and management to reproduce scientific workflows execution. Future Generation Computer Systems, 86(??):799–820, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17314917.

Huang:2019:AMR

[HMMW19] Runhe Huang, Peter Kimani Mungai, Jianhua Ma, and Kevin I-Kai Wang. Associative memory and recall model with KID model for human activity recognition. Future Generation Computer Systems, 92(??):312–323, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18307337.

Hidders:2015:RAS

[HMS15] Jan Hidders, Paolo Missier, and Jacek Sroka. Recent advances in Scalable Workflow Enactment Engines and Technologies.

Future Generation Computer Systems, 46(??):1–2, May 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000047.

Hu:2014:PAT

[HMW14] Jia Hu, Geyong Min, and Mike E. Woodward. Performance analysis of a threshold-based dynamic TXOP scheme for intra-AC QoS in wireless LANs. Future Generation Computer Systems, 38(??):69– 74, September 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002008.

Hu:2019:IFI

[HMW^{+}19] Long Hu, Yiming Miao, Gaoxiang Wu, Mohammad Mehedi Hassan, and Iztok Humar. iRobot-Factory: an intelligent robot factory based on cognitive manufacturing and edge computing. Future Generation Computer Systems, 90(??):569–577, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831183X.

Hirsch:2018:ACC

[HMZ18] Matías Hirsch, Cristian Mateos, and Alejandro Zunino. Augmenting computing capabilities at the edge by jointly exploiting mobile devices: a survey. Future Generation Computer Systems, 88(??):644– 662, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1732767X.

Herrero:2013:RMI

[HNCJ13] Álvaro Herrero, Martí Navarro, Emilio Corchado, and Vicente Julián. RT-MOVICAB-IDS: Addressing real-time intrusion detection. Future Generation Computer Systems, 29(1):250–261, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X10002785.

Hassanzadeh-Nazarabadi:2018:DLA

[HNKÖ18] Yahya Hassanzadeh-Nazarabadi, Alptekin Küpçü, and Öznur Özkasap. Decentralized and locality aware replication method for DHT-based P2P storage systems. Future Generation Computer Systems, 84(??):32–46, July 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17326973.

Hu:2018:UFI

[HNQ^{+}18] Pengfei Hu, Huansheng Ning, Tie Qiu, Yue Xu, Xiong Luo, and Arun Kumar Sangaiah. A unified face identification and resolution scheme using cloud computing in Internet of Things. Future Generation Computer Systems, 81(??): 582–592, April 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17305228.

Hachaj:2017:CTT

[HO17] Tomasz Hachaj and Marek R.

Ogiela. Clustering of trending topics in microblogging posts: a graphbased approach. Future Generation Computer Systems, 67(??):297–304, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300863.

Hernandez:2018:UML

[HPGMM18] Álvaro Brandón Hernández, María S. Perez, Smrati Gupta, and Victor Muntés-Mulero. Using machine learning to optimize parallelism in big data applications. Future Generation Computer Systems, 86(??):1076–1092, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17314668.

Hurrah:2019:DWF

[HPL^{+}19] Nasir N. Hurrah, Shabir A. Parah, Nazir A. Loan, Javaid A. Sheikh, Mohammad Elhoseny, and Khan Muhammad. Dual watermarking framework for privacy protection and content authentication of multimedia. Future Generation Computer Systems, 94(??): 654–673, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317096.

Hong:2018:ASS

[HPP^{+}18] Seungwan Hong, Sangho Park, Lee Won Park, Minseo Jeon, and Hangbae Chang. An analysis of security systems for electronic information for establishing secure Internet of Things environments: Focusing on research trends in the security field in South Korea.

Future Generation Computer Systems, 82(??):769– 782, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17308567.

Han:2018:CPM

[HPZL18] Shi-Zhuan Han, Wen-Tsao Pan, Ying-Ying Zhou, and Zong-Li Liu. Construct the prediction model for China agricultural output value based on the optimization neural network of fruit fly optimization algorithm. Future Generation Computer Systems, 86(??):663–669, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18302802.

Huang:2010:MPA

[HQ10] Zhixing Huang and Yuhui Qiu. A multiple-perspective approach to constructing and aggregating Citation Semantic Link Network. Future Generation Computer Systems, 26(3):400– 407, March 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Han:2016:SSC

[HQH16] Fei Han, Jing Qin, and Jiankun Hu. Secure searches in the cloud: a survey. Future Generation Computer Systems, 62(??):66–75, September 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16000091.

Han:2014:GTK

[HQZH14] Fei Han, Jing Qin, Huawei Zhao, and Jiankun Hu. A general transformation from KP-ABE to searchable encryption. Future Generation Computer Systems, 30(??):107–115, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001921.

Hendrix:2014:CCA

[HRR^{+}14] Valerie Hendrix, Lavanya Ramakrishnan, Youngryel Ryu, Catharine van Ingen, Keith R. Jackson, and Deborah Agarwal. CAMP: Community Access MODIS Pipeline. Future Generation Computer Systems, 36(??):418–429, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002021.

Hidalgo:2018:MSP

[HRVW18] Nicolas Hidalgo, Erika Rosas, Cristobal Vasquez, and Daniel Wladdimiro. Measuring stream processing systems adaptability under dynamic workloads. Future Generation Computer Systems, 88(??):413– 423, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17326304.

Haghighat:2019:BWG

[HS19] Alireza Toroghi Haghighat and Mehdi Shajari. Block withholding game among Bitcoin mining pools. Future Generation Computer Systems, 97(??):482–491, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17330686.

Horta:2018:ASC

[HSB^{+}18] Vitor Horta, Victor Ströele, Regina Braga, José Maria N. David, and Fernanda Campos. Analyzing scientific context of researchers and communities by using complex network and semantic technologies. Future Generation Computer Systems, 89(??):584–605, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17328431.

Hayat:2019:SGB

[HSBE19] Routaib Hayat, Essaid Sabir, Elarbi Badidi, and Mohammed ElKoutbi. A signaling game-based approach for data-as-a-service provisioning in IoT-cloud. Future Generation Computer Systems, 92(??): 1040–1050, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        https://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X17306891.

Hsu:2015:LLA

[HSC15] Ching-Hsien Hsu, Kenn D. Slagter, and Yeh-Ching Chung. Locality and loading aware virtual machine mapping techniques for optimizing communications in MapReduce applications. Future Generation Computer Systems, 53(??):43–54, December 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000989.

Han:2013:IBD

[HSM13] Jinguang Han, Willy Susilo, and Yi Mu. Identity-based data storage in cloud computing. Future Generation Computer Systems, 29(3): 673–681, March 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001719.

Holub:2013:GAD

[HSP^{+}13] Petr Holub, Martin Srom, Martin Pulec, Jirí Matela, and Martin Jirman. GPUaccelerated DXT and JPEG compression schemes for low-latency network transmissions of HD, 2K, and 4K video. Future Generation Computer Systems, 29(8):1991–2006, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001209.

Hui:2017:MRB

[HSS17] Terence K. L. Hui, R. Simon Sherratt, and Daniel Díaz Sánchez. Major requirements for building Smart Homes in smart cities based on Internet of Things technologies. Future Generation Computer Systems, 76(??):358–369, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16304721.

Hilal:2018:DSM

[HST^{+}18] Allaa R. Hilal, Aya Sayedelahl, Arash Tabibiazar, Mohamed S. Kamel, and Otman A. Basir. A distributed sensor management for large-scale IoT indoor acoustic surveillance. Future Generation Computer Systems, 86(??):1170–1184, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18300529.

Hsu:2014:IBD

[Hsu14] Ching-Hsien Hsu. Intelligent big data processing. Future Generation Computer Systems, 36(??):16– 18, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000284.

Harnie:2017:SML

[HSV^{+}17] Dries Harnie, Mathijs Saey, Alexander E. Vapirev, Jörg Kurt Wegner, Andrey Gedich, Marvin Steijaert, Hugo Ceulemans, Roel Wuyts, and Wolfgang De Meuter. Scaling machine learning for target prediction in drug discovery using Apache Spark. Future Generation Computer Systems, 67(??):409– 417, February 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630111X.

Haggui:2018:HCD

[HTL^{+}18] Olfa Haggui, Claude Tadonki, Lionel Lacassagne, Fatma Sayadi, and Bouraoui Ouni. Harris corner detection on a NUMA manycore. Future Generation Computer Systems, 88(??):442–452, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1732188X.

Huang:2010:CTA

[Hua10]  H. Howie Huang.  A

control-theoretic approach to automated local policy enforcement in computational grids. Future Generation Computer Systems, 26(6):787–796, June 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hassan:2018:RHA

[HUMA18] Mohammed Mehedi Hassan, Md. Zia Uddin, Amr Mohamed, and Ahmad Almogren. A robust human activity recognition system using smartphone sensors and deep learning. Future Generation Computer Systems, 81(??):307–313, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17317351.

Hao:2019:RCN

[HUY^{+}19] Yixue Hao, Mohd Usama, Jun Yang, M. Shamim Hossain, and Ahmed Ghoneim. Recurrent convolutional neural network based multimodal disease risk prediction. Future Generation Computer Systems, 92(??):76–83, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18318843.

Hsu:2012:AEF

[HWWT12] Chih-Yu Hsu, Hao-Feng Wang, Hui-Ching Wang, and Kuo-Kun Tseng. Automatic extraction of face contours in images and videos. Future Generation Computer Systems, 28(1):322–335, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X10002281.

Hu:2019:DDM

[HX19] Tao Hu and Chunxia Xiao.

Data-driven main color map feature learning, design and simulation for smart ethnic cloth. Future Generation Computer Systems, 97(??):153–164, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18323070.

Hossain:2017:CBM

[HXA^{+}17] M. Shamim Hossain, Changsheng Xu, A. M. Artoli, Manzur Murshed, and Stefan Göbel. Cloud-based multimedia services for healthcare and other related applications. Future Generation Computer Systems, 66(??):27–29, January 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302412.

Huang:2018:FCB

[HXC^{+}18] Tao Huang, Boyi Xu, Hongming Cai, Jiawei Du, Kuo-Ming Chao, and Chengxi Huang. A fog computing based concept drift adaptive process mining framework for mobile APPs. Future Generation Computer Systems, 89(??):670–684, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17330273.

Huang:2018:IQE

[HXL^{+}18] Xiaohong Huang, Kun Xie, Supeng Leng, Tingting Yuan, and Maode Ma. Improving quality of experience in multimedia Internet of Things leveraging machine learning on big data. Future Generation Computer Systems, 86(??):1413–1423, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324500.

He:2018:NBA

[HXWW18] Yichao He, Haoran Xie, Tak-Lam Wong, and Xizhao Wang. A novel binary artificial bee colony algorithm for the set-union knapsack problem. Future Generation Computer Systems, 78 (part 1)(??):77–86, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310415.

Huang:2013:ECF

[HXY13] Jun Huang, Yanbo Xue, and Le Yang. An efficient closed-form solution for joint synchronization and localization using TOA. Future Generation Computer Systems, 29(3): 776–781, March 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001847.

Hu:2018:SSS

[HYC^{+}18] Long Hu, Jun Yang, Min Chen, Yongfeng Qian, and Joel J. P. C. Rodrigues. SCAI-SVSC: Smart clothing for effective interaction with a sustainable vital sign collection. Future Generation Computer Systems, 86(??):329–338, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17328285.

Huang:2018:PIB

[HYF18] Qinlong Huang, Yixian Yang, and Jingyi Fu. PRECISE: Identitybased private data sharing with conditional proxy reencryption in online social networks. Future Generation Computer Systems, 86(??):1523–1533, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17310257.

Hu:2019:DSC

[HYG^{+}19] Tao Hu, Peng Yi, Zehua Guo, Julong Lan, and Yuxiang Hu. Dynamic slave controller assignment for enhancing control plane robustness in software-defined networks. Future Generation Computer Systems, 95(??):681– 693, June 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1832288X.

Huang:2017:SED

[HYS17] Qinlong Huang, Yixian Yang, and Mansuo Shen. Secure and efficient data collaboration with hierarchical attribute-based encryption in cloud computing. Future Generation Computer Systems, 72(??): 239–249, July 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16303521.

See corrigendum [HYS18].

Huang:2018:CSE

[HYS18] Qinlong Huang, Yixian Yang, and Mansuo Shen.

Corrigendum to “Secure and efficient data collaboration with hierarchical attribute-based encryption in cloud computing” [Future Gener. Comput. Syst. 72 (2017) 239–249]. Future Generation Computer Systems, 86(??):1534, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18312238.

See [HYS17].

Hossain:2019:QOM

[HYX^{+}19] M. Shamim Hossain, Xinghui You, Wenjing Xiao, Jiayi Lu, and Enmin Song. QoS-oriented multimedia transmission using multipath routing. Future Generation Computer Systems, 99(??):226–234, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19305667.

Hassan:2016:SIM

[HYZS16] Houcine Hassan, Laurence T. Yang, Haibo Zhang, and Marco D. Santambrogio. Special issue on: Multicore and Many-core Architectures for Future Generation Embedded Systems. Future Generation Computer Systems, 56(??):169–170, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003702.

Huang:2010:DBL

[HZ10] Kun Huang and Dafang Zhang. DHT-based lightweight broadcast algorithms in large-scale computing infrastructures. Future Generation Computer Systems, 26(3):291–303, March 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hu:2019:PBG

[HZ19] Kekun Hu and Guosun Zeng. Placing big graph into cloud for parallel processing with a twophase community-aware approach. Future Generation Computer Systems, 101(??):1187–1200, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19303103.

Hao:2010:WSD

[HZC10] Yanan Hao, Yanchun Zhang, and Jinli Cao. Web services discovery and rank: An information retrieval approach. Future Generation Computer Systems, 26(8):1053–1062, October 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hou:2019:DIR

[HZDS19] Junjie Hou, Yongxin Zhu, Sen Du, and Shijin Song. Design and implementation of reconfigurable acceleration for in-memory distributed big data computing. Future Generation Computer Systems, 92(??):68–75, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1831210X.

Han:2018:ERA

[HZL18a] Qi Han, Yinghui Zhang, and Hui Li. Efficient and robust attribute-based encryption supporting access policy hiding in Internet of Things. Future Generation Computer Systems, 83(??):269–277, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1731868X.

Hou:2018:FNN

[HZL18b] Yue Hou, Long Zhao, and Huaiwei Lu. Fuzzy neural network optimization and network traffic forecasting based on improved differential evolution. Future Generation Computer Systems, 81(??):425–432, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17304582.

He:2019:LSC

[HZL^{+}19] Heng He, Ji Zhang, Peng Li, Yu Jin, and Tao Zhang. A lightweight secure conjunctive keyword search scheme in hybrid cloud. Future Generation Computer Systems, 93(??):727– 736, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18311749.

Huang:2019:OSA

[HZLH19] Xiaohong Huang, Yong Zhang, Dandan Li, and Lu Han. An optimal scheduling algorithm for hybrid EV charging scenario using consortium blockchains. Future Generation Computer Systems, 91(??):555–562, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18313578.

Huang:2014:CHI

[HZM14] Runhe Huang, Xin Zhao, and Jianhua Ma. The contours of a human individual model based empathetic u-pillbox system for humanistic geriatric healthcare. Future Generation Computer Systems, 37(??): 404–416, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002057.

Holl:2014:NOP

[HZP^{+}14] Sonja Holl, Olav Zimmermann, Magnus Palmblad, Yassene Mohammed, and Martin Hofmann-Apitius. A new optimization phase for scientific workflow management systems. Future Generation Computer Systems, 36(??):352–362, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001842.

Huang:2016:ADI

[HZW^{+}16] Tian Huang, Yongxin Zhu, Yafei Wu, Stéphane Bressan, and Gillian Dobbie. Anomaly detection and identification scheme for VM live migration in cloud infrastructure. Future Generation Computer Systems, 56(??):736–745, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500206X.

Han:2018:SLP

[HZW^{+}18] Guangjie Han, Lina Zhou, Hao Wang, Wenbo Zhang, and Sammy Chan. A source location protection protocol based on dynamic routing in WSNs for the Social Internet of Things. Future Generation Computer Systems, 82(??):689– 697, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17310348.

Hammad:2019:NTD

[HZW19] Mohamed Hammad, Shanzhuo Zhang, and Kuanquan Wang. A novel twodimensional ECG feature extraction and classification algorithm based on convolution neural network for human authentication. Future Generation Computer Systems, 101(??):180–196, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18329923.

Huajun:2019:CAB

[HZX^{+}19] Lin Huajun, Liu Zhao, Ma Xuemei, Wang Tingting, Xu Bodong, Zhou Xiaona, and Zhang Zhongtao. Correlation analysis of biochemical indicators in common bile duct stone patients with negative magnetic resonance cholangiopancreatography. Future Generation Computer Systems, 98(??):530– 535, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326232.

See retraction notice [HZX^{+}20].

Huajun:2020:RNC

[HZX^{+}20] Lin Huajun, Liu Zhao, Ma Xuemei, Wang Tingting, Xu Bodong, Zhou Xiaona, and Zhang Zhongtao. Retraction notice to “Correlation analysis of biochemical indicators in common bile duct stone patients with negative magnetic resonance cholangiopancreatography” [Future Gener. Comput. Syst. 98 (2019) 530–535]. Future Generation Computer Systems, 107(??):1146, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/S0167739X20306488.

See [HZX^{+}19].

He:2014:DRC

[HZZ^{+}14] Ligang He, Deqing Zou, Zhang Zhang, Chao Chen, Hai Jin, and Stephen A. Jarvis. Developing resource consolidation frameworks for moldable virtual machines in clouds. Future Generation Computer Systems, 32(??):69– 81, March 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001112.

Hu:2018:DSB

[HZZ^{+}18] Tao Hu, Hao Zhang, Xinyan Zhu, Julaine Clunis, and Gelan Yang. Depth sensor based human detection for indoor surveillance. Future Generation Computer Systems, 88(??):540–551, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18308537.

Izakian:2010:AMR

[IAL10] Hesam Izakian, Ajith Abraham, and Behrouz Tork Ladani. An auction method for resource allocation in computational grids. Future Generation Computer Systems, 26(2): 228–235, February 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Inoubli:2018:ESB

[IAM^{+}18] Wissem Inoubli, Sabeur Aridhi, Haithem Mezni, Mondher Maddouri, and Engelbert Mephu Nguifo. An experimental survey on big data frameworks. Future Generation Computer Systems, 86(??):546–564, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17327450.

Ibaida:2014:CEF

[IASK14] Ayman Ibaida, Dhiah Al-Shammary, and Ibrahim Khalil. Cloud enabled fractal based ECG compression in wireless body sensor networks. Future Generation Computer Systems, 35(??): 91–101, June 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002896.

Iwashita:2019:LCD

[IdAP19] Adriana Sayuri Iwashita, Victor Hugo C. de Albuquerque, and João Paulo Papa. Learning concept drift with ensembles of optimum-path forestbased classifiers. Future Generation Computer Systems, 95(??):198–211, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1832257X.

Iqbal:2011:ARP

[IDCJ11] Waheed Iqbal, Matthew N. Dailey, David Carrera, and Paul Janecek. Adaptive resource provisioning for read intensive multi-tier applications in the cloud. Future Generation Computer Systems, 27(6):871– 879, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Imran:2019:TOS

[IDKD19] Muhammad Imran, Muhammad Hanif Durad, Farrukh Aslam Khan, and Abdelouahid Derhab. Toward an optimal solution against Denial of Service attacks in Software Defined Networks. Future Generation Computer Systems, 92(??):444–453, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18302930.

Iglesias:2016:ITC

[IDM^{+}16] Jesus Omana Iglesias, Milan De Cauwer, Deepak Mehta, Barry O’Sullivan, and Liam Murphy. Increasing task consolidation efficiency by using more accurate resource estimations. Future Generation Computer Systems, 56(??): 407–420, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002873.

Islam:2019:BBF

[IFD^{+}19] Naveed Islam, Yasir Faheem, Ikram Ud Din, Muhammad Talha, Mohsen Guizani, and Mudassir Khalil. A blockchain-based fog computing framework for activity recognition as an application to ehealthcare services. Future Generation Computer Systems, 100(??):569–578, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19309860.

Islam:2012:GUE

[IG12] Salekul Islam and Jean-Charles Grégoire. Giving users an edge: a flexible Cloud model and its application for multimedia.

Future Generation Computer Systems, 28(6):823– 832, June 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000143.

Itani:2014:SBI

[IGB^{+}14] Wassim Itani, Cesar Ghali, Ramzi Bassil, Ayman Kayssi, and Ali Chehab. ServBGP: BGP-inspired autonomic service routing for multi-provider collaborative architectures in the cloud. Future Generation Computer Systems, 32(??): 99–117, March 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001094.

Ismail:2018:MPP

[IHA18] Walaa N. Ismail, Mohammad Mehedi Hassan, and Hessah A. Alsalamah. Mining of productive periodicfrequent patterns for IoT data analytics. Future Generation Computer Systems, 88(??):512–523, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17326481.

Imran:2018:API

[IHK^{+}18] Muhammad Imran, Helmut Hlavacs, Fakhri Alam Khan, Saima Jabeen, Fiaz Gul Khan, Sajid Shah, and Mafawez Alharbi. Aggregated provenance and its implications in clouds. Future Generation Computer Systems, 81(??):348– 358, April 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17320927.

Imran:2019:ETS

[IJCR19] Muhammad Imran, Sohail Jabbar, Naveen Chilamkurti, and Joel J. P. C. Rodrigues. Enabling technologies for Social Internet of Things. Future Generation Computer Systems, 92(??):715–717, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18328541.

Islam:2012:EPM

[IKLL12] Sadeka Islam, Jacky Keung, Kevin Lee, and Anna Liu. Empirical prediction models for adaptive resource provisioning in the cloud. Future Generation Computer Systems, 28(1):155–162, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001129.

Imani:2019:ASR

[Ima19] Maryam Imani. Adaptive signal representation and multi-scale decomposition for panchromatic and multispectral image fusion. Future Generation Computer Systems, 99(??):410– 424, October 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18323641.

Islam:2018:REP

[IOV^{+}18] SK Hafizul Islam, Mohammad S. Obaidat, Pandi Vijayakumar, Enas Abdulhay, Fagen Li, and M. Krishna Chaitanya Reddy. A robust and efficient password-based conditional privacy preserving authentication and group-key agreement protocol for VANETs. Future Generation Computer Systems, 84(??):216–227, July 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17308439.

Ibrahim:2016:GEC

[IPCA^{+}16] Shadi Ibrahim, Tien-Dat Phan, Alexandra Carpen-Amarie, Houssem-Eddine Chihoub, Diana Moise, and Gabriel Antoniu. Governing energy consumption in Hadoop through CPU frequency scaling: an analysis. Future Generation Computer Systems, 54(??):219–232, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000060.

Imbernon:2018:ELS

[IPG^{+}18] Baldomero Imbernón, Javier Prades, Domingo Giménez, José M. Cecilia, and Federico Silla. Enhancing large-scale docking simulation on heterogeneous systems: An MPI vs rCUDA study. Future Generation Computer Systems, 79 (part 1)(??):26– 37, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17309974.

Iranpour:2018:DLB

[IS18] Eiman Iranpour and Saeed Sharifian. A distributed load balancing and admission control algorithm based on fuzzy type-2 and game theory for largescale SaaS cloud architectures. Future Generation Computer Systems, 86(??):81–98, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1732006X.

Ivanovic:2015:EGR

[ISS^{+}15] Milos Ivanovic, Visnja Simic, Boban Stojanovic, Ana Kaplarevic-Malisic, and Branko Marovic. Elastic grid resource provisioning with WoBinGO: a parallel framework for genetic algorithm based optimization. Future Generation Computer Systems, 42(??): 44–54, January 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001678.

Jararweh:2016:SDC

[JAAD^{+}16] Yaser Jararweh, Mahmoud Al-Ayyoub, Ala’ Darabseh, Elhadj Benkhelifa, Mladen Vouk, and Andy Rindos. Software defined cloud: Survey, system and evaluation. Future Generation Computer Systems, 58 (??):56–74, May 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003283.

Jacobsson:2016:RAS

[JBC16] Andreas Jacobsson, Martin Boldt, and Bengt Carlsson. A risk analysis of a smart home automation system. Future Generation Computer Systems, 56(??): 719–733, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002812.

Jain:2018:CLP

[JBM^{+}18] Bindiya Jain, Gursewak Brar, Jyoteesh Malhotra, Shalli Rani, and Syed Hassan Ahmed. A cross layer protocol for traffic management in Social Internet of Vehicles. Future Generation Computer Systems, 82(??):707–714, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312694.

Johnston:2018:CSB

[JBP^{+}18] Steven J. Johnston, Philip J.

Basford, Colin S. Perkins, Herry Herry, Fung Po Tso, Dimitrios Pezaros, Robert D. Mullins, Eiko Yoneki, Simon J. Cox, and Jeremy Singer. Commodity single board computer clusters and their applications. Future Generation Computer Systems, 89(??):201– 212, December 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18301833.

Javed:2016:CMM

[JBR^{+}16] Barkha Javed, Peter Bloodsworth, Raihan Ur Rasool, Kamran Munir, and Omer Rana. Cloud Market Maker: an automated dynamic pricing marketplace for cloud users. Future Generation Computer Systems, 54(??):52–67, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002058.

James:2015:BIS

[JC15] Anne James and Jen-Yao Chung. Business and industry specific cloud: Challenges and opportunities. Future Generation Computer Systems, 48(??): ibc, July 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002593.

Jenviriyakul:2019:ACV

[JCA^{+}19] Patcharaporn Jenviriyakul, Gantaphon Chalumporn, Tiranee Achalakul, Filippo Costa, and Khajonpong Akkarajitsakul. ALICE Connex: a volunteer computing platform for the time-of-flight calibration of the ALICE experiment. An opportunistic use of CPU cycles on Android devices. Future Generation Computer Systems, 94(??):510– 523, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1731141X.

Juve:2013:CPS

[JCD^{+}13] Gideon Juve, Ann Chervenak, Ewa Deelman, Shishir Bharathi, Gaurang Mehta, and Karan Vahi. Characterizing and profiling scientific workflows. Future Generation Computer Systems, 29(3):682–692, March 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001732.

Jiang:2015:TSR

[JCL^{+}15] Tao Jiang, Xiaofeng Chen, Jin Li, Duncan S. Wong, Jianfeng Ma, and Joseph K. Liu. Towards secure and reliable cloud storage against data re-outsourcing. Future Generation Computer Systems, 52(??):86– 94, November 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002362.

Jie:2019:TGL

[JCL^{+}19] Yingmo Jie, Kim-Kwang Raymond Choo, Mingchu Li, Ling Chen, and Cheng Guo. Tradeoff gain and loss optimization against man-in-the-middle attacks based on game theoretic model. Future Generation Computer Systems, 101(??):169–179, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18315541.

Jardi-Cedo:2018:TBL

[JCMPPC^{+}18] Roger Jardí-Cedó, Macià Mut-Puigserver, M. Magdalena Payeras-Capellà, Jordi Castellà-Roca, and Alexandre Viejo. Timebased low emission zones preserving drivers’ privacy. Future Generation Computer Systems, 80(??):558– 571, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301972.

Jin:2014:MLM

[JDW^{+}14] Hai Jin, Li Deng, Song Wu, Xuanhua Shi, Hanhua Chen, and Xiaodong Pan. MECOM: Live migration of virtual machines by adaptively compressing memory pages. Future Generation Computer Systems, 38(??):23– 35, September 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002100.

Juarez:2018:DEA

[JEB18] Fredy Juarez, Jorge Ejarque, and Rosa M. Badia. Dynamic energy-aware scheduling for parallel taskbased application in cloud computing. Future Generation Computer Systems, 78 (part 1)(??):257–271, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630214X.

Jiang:2017:DFA

[JFZL17] Jianhua Jiang, Yunzhao Feng, Jia Zhao, and Keqin Li. DataABC: a fast ABC based energy-efficient live VM consolidation policy with data-intensive energy evaluation model. Future Generation Computer Systems, 74(??):132–141, September 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301236.

Jatoth:2019:OFA

[JGB19]  Chandrashekar  Jatoth,

G. R. Gangadharan, and Rajkumar Buyya. Optimal fitness aware cloud service composition using an adaptive genotypes evolution based genetic algorithm. Future Generation Computer Systems, 94(??): 185–198, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18308501.

Jatoth:2018:QAB

[JGFB18]  Chandrashekar  Jatoth,

G. R. Gangadharan, Ugo Fiore, and Rajkumar Buyya. QoS-aware big service composition using MapReduce based evolutionary algorithm with guided mutation. Future Generation Computer Systems, 86(??):1008–1018, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17315637.

Jeon:2016:MCS

[JH16] Seungwoo Jeon and Bonghee Hong. Monte Carlo simulation-based traffic speed forecasting using historical big data. Future Generation Computer Systems, 65(??):182–195, December 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003799.

Jakubowska:2010:VCS

[JHC10] Joanna Jakubowska, Ela Hunt, and Matthew Chalmers. VisGenome with Cartoon-Plus: Supporting large scale genomic analyses via physical space deformation. Future Generation Computer Systems, 26(3): 441–454, March 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jeon:2018:PGT

[JHC18] Seungwoo Jeon, Bonghee Hong, and Victor Chang. Pattern graph trackingbased stock price prediction using big data. Future Generation Computer Systems, 80(??):171–187, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301991.

Jang:2019:KLE

[JJH19] Joonhyouk Jang, Jinman Jung, and Jiman Hong. K-LZF: an efficient and fair scheduling for edge computing servers. Future Generation Computer Systems, 98(??):44– 53, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18330000.

Joe:2017:EDI

[JK17] Hyunwoo Joe and Hyungshin Kim. Effects of dynamic isolation for full virtualized RTOS and GPOS guests. Future Generation Computer Systems, 70(??): 26–41, May 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16308020.

Jan:2019:PBM

[JKAU19] Mian Ahmad Jan, Fazlullah Khan, Muhammad Alam, and Muhammad Usman. A payloadbased mutual authentication scheme for Internet of Things. Future Generation Computer Systems, 92(??):1028–1039, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17303898.

Joe:2017:OOP

[JKLK17] Hyunwoo Joe, Jungseok Kim, Jemin Lee, and Hyungshin Kim. Outputoriented power saving mode for mobile devices. Future Generation Computer Systems, 72(??):49– 64, July 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301224.

Jeong:2014:ITP

[JL14] EunHee Jeong and ByungKwan Lee. An IP Traceback Protocol using a compressed hash table, a sinkhole router and data mining based on network forensics against network attacks. Future Generation Computer Systems, 33(??): 42–52, April 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002410.

Jiang:2018:AHP

[JLC18] Rong Jiang, Rongxing Lu, and Kim-Kwang Raymond Choo. Achieving high performance and privacypreserving query over encrypted multidimensional big metering data. Future Generation Computer Systems, 78 (part 1)(??):392– 401, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301157.

Ji:2020:RNI

[JLC^{+}20] Yanli Ji, Wei Lu, Guowei Che, Mei Yang, and Lunxu Liu. Retraction notice to “Inhibition of KRAS gene mutation on non-small cell lung cancer and its effect on circulating tumor cells” [future gener. comput. syst. 98 (2019) 104–108]. Future Generation Computer Systems, 111(??):937, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19332996.

See [YWG^{+}19].

Ji:2012:PEV

[JLCC12] Wen Ji, Jiangchuan Liu, Min Chen, and Yiqiang Chen. Power-efficient video encoding on resourcelimited systems: a gametheoretic approach. Future Generation Computer Systems, 28(2):427–436, February 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000537.

Jin:2019:DLA

[JLD^{+}19] Jiahui Jin, Junzhou Luo, Mingyang Du, Yongcheng Dang, Feng Li, Jinghui Zhang, and Aibo Song. A data-locality-aware task scheduler for distributed social graph queries. Future Generation Computer Systems, 93(??):1010–1022, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17309548.

Jin:2013:FTL

[JLI^{+}13] Hai Jin, Xiao Ling, Shadi Ibrahim, Wenzhi Cao, Song Wu, and Gabriel Antoniu. Flubber: Two-level disk scheduling in virtualized environment. Future Generation Computer Systems, 29(8):2222–2238, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001246.

Jia:2017:MMA

[JLL17] Zhaohong Jia, Xiaohao Li, and Joseph Y.-T. Leung. Minimizing makespan for arbitrary size jobs with release times on p-batch machines with arbitrary capacities. Future Generation Computer Systems, 67(??): 22–34, February 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302680.

Jedari:2017:GTI

[JLQ^{+}17] Behrouz Jedari, Li Liu, Tie Qiu, Azizur Rahim, and Feng Xia. A gametheoretic incentive scheme for social-aware routing in selfish mobile social networks. Future Generation Computer Systems, 70(??): 178–190, May 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302059.

Ji:2018:MOL

[JLQ18] Ying Ji, Meng Li, and Shaojian Qu. Multiobjective linear programming games and applications in supply chain competition. Future Generation Computer Systems, 86(??):591–597, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18304266.

Jian:2018:INB

[JLQZ18] Chengfeng Jian, Miao Li, Keyi Qiu, and Meiyu Zhang. An improved NBAbased STEP design intention feature recognition. Future Generation Computer Systems, 88(??):357– 362, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18302954.

Jakobs:2018:TLA

[JLRS18] Thomas Jakobs, Jens Lang, Gudula Rünger, and Paul Stöcker. Tuning linear algebra for energy efficiency on multicore machines by adapting the ATLAS library. Future Generation Computer Systems, 82(??):555–564, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1730359X.

Jeong:2019:DTI

[JLS19] Junho Jeong, Joong Yeon Lim, and Yunsik Son. A data type inference method based on long short-term memory by improved feature for weakness analysis in binary code. Future Generation Computer Systems, 100(??):1044–1052, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18329145.

Jiang:2019:PPP

[JLX^{+}19] Wenbo Jiang, Hongwei Li, Guowen Xu, Mi Wen, Guishan Dong, and Xiaodong Lin. PTAS: Privacy-preserving thinclient authentication scheme in blockchain-based PKI. Future Generation Computer Systems, 96(??):185– 195, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18315097.

Jiang:2018:FNF

[JLY^{+}18] Wenbin Jiang, Min Long, Laurence T. Yang, Xiaobai Liu, Hai Jin, Alan L. Yuille, and Ye Chi. FIPIP: a novel fine-grained parallel partition based intraframe prediction on heterogeneous many-core systems. Future Generation Computer Systems, 78 (part 1)(??):316–329, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301194.

Jlassi:2019:TCC

[JMAG19] Sindyana Jlassi, Amel Mammar, Imed Abbassi, and Mohamed Graiet. Towards correct cloud resource allocation in FOSS applications. Future Generation Computer Systems, 91(??):392–406, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18306095.

Jan:2018:SAD

[JNHL18] Mian Ahmad Jan, Priyadarsi Nanda, Xiangjian He, and Ren Ping Liu. A Sybil attack detection scheme for a forest wildfire monitoring application. Future Generation Computer Systems, 80(??):613–626, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301522.

Jeyarani:2012:DIA

[JNR12] R. Jeyarani, N. Nagaveni, and R. Vasanth Ram. Design and implementation of adaptive poweraware virtual machine provisioner (APA–VMP) using swarm intelligence. Future Generation Computer Systems, 28(5):811–821, May 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001130.

Jayaratne:2019:DIP

[JNS^{+}19] Madhura Jayaratne, Dinithi Nallaperuma, Daswin De Silva, Damminda Alahakoon, Brian Devitt, Kate E. Webster, and Naveen Chilamkurti. A data integration platform for patient-centered ehealthcare and clinical decision support. Future Generation Computer Systems, 92(??):996–1008, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17308142.

James:2011:SSC

[JO11] Anne James and Jonice Oliviera. Special section: Computer supported cooperative work in design. Future Generation Computer Systems, 27(1):56–57, January 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jarus:2014:RPU

[JOPW14] M. Jarus, A. Oleksiak, T. Piontek, and J. Weglarz. Runtime power usage estimation of HPC servers for various classes of reallife applications. Future Generation Computer Systems, 36(??):299–310, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300157X.

Junior:2019:CSO

[JOSD19] Warley Junior, Eduardo Oliveira, Albertinin Santos, and Kelvin Dias. A context-sensitive offloading system using machinelearning classification algorithms for mobile cloud environment. Future Generation Computer Systems, 90(??):503–520, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17326729.

Jebadurai:2018:SRR

[JP18] Jebaveerasingh Jebadurai and J. Dinesh Peter. Super-resolution of retinal images using multikernel SVR for IoT healthcare applications. Future Generation Computer Systems, 83(??):338–346, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17322136.

Janetschek:2017:WRE

[JPB17] M. Janetschek, R. Prodan, and S. Benedict. A workflow runtime environment for manycore parallel architectures. Future Generation Computer Systems, 75(??):330–347, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302571.

Jagodic:2011:EMU

[JRJ^{+}11] Ratko Jagodic, Luc Renambot, Andrew Johnson, Jason Leigh, and Sachin Deshpande. Enabling multi-user interaction in large high-resolution distributed environments. Future Generation Computer Systems, 27(7):914–923, July 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

James:2012:SSQ

[JS12] Anne James and Weiming Shen. Special section: QoS in Grid and Cloud. Future Generation Computer Systems, 28(7):1003–1004, July 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000672.

Jin:2013:PCU

[JS13] Hui Jin and Xian-He Sun.

Performance comparison under failures of MPI and MapReduce: an analytical approach. Future Generation Computer Systems, 29(7):1808–1815, September 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000290.

Jiang:2015:SSS

[JSC^{+}15] Hai Jiang, Feng Shen, Su Chen, Kuan-Ching Li, and Young-Sik Jeong. A secure and scalable storage system for aggregate data in IoT. Future Generation Computer Systems, 49(??): 133–141, August 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1400243X.

Jiang:2018:CPA

[JSMG18] Yinhao Jiang, Willy Susilo, Yi Mu, and Fuchun Guo. Ciphertext-policy attribute-based encryption against key-delegation abuse in fog computing. Future Generation Computer Systems, 78 (part 2)(??):720–729, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301322.

Jin:2012:HAS

[JSS^{+}12] Yingwei Jin, Lulu Sun, Yanming Shen, Keqiu Li, and Geyong Min. A highly available spectrum allocation service model in dynamic spectrum market. Future Generation Computer Systems, 28(6):940– 946, June 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001889.

Jin:2019:CEP

[JSZ^{+}19] Dawei Jin, Si Shi, Yin Zhang, Haider Abbas, and Tiong-Thye Goh. A complex event processing framework for an adaptive language learning system. Future Generation Computer Systems, 92(??):857– 867, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17312815.

Javadi:2013:CSP

[JTB13] Bahman Javadi, Ruppa K. Thulasiram, and Rajkumar Buyya. Characterizing spot price dynamics in public cloud environments. Future Generation Computer Systems, 29(4): 988–999, June 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001483.

Jrad:2015:SEL

[JTBS15] Foued Jrad, Jie Tao, Ivona Brandic, and Achim Streit. SLA enactment for largescale healthcare workflows on multi-cloud. Future Generation Computer Systems, 43–44(??):135–148, February 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001381.

Jiang:2019:CSR

[JTL^{+}19] Yuanchun Jiang, Dandan Tao, Yezheng Liu, Jianshan Sun, and Haifeng Ling. Cloud service recommendation based on unstructured textual information. Future Generation Computer Systems, 97(??): 387–396, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326943.

Javadi:2013:DOD

[JTS13] Bahman Javadi, Martin Tomko, and Richard O.

Sinnott. Decentralized orchestration of data-centric workflows in cloud environments. Future Generation Computer Systems, 29(7):1826–1837, September 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000228.

Jung:2017:CCI

[Jun17] Jason J. Jung. Computational collective intelligence with big data: Challenges and opportunities. Future Generation Computer Systems, 66(??):87– 88, January 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302837.

Jun:2018:NGC

[Jun18] Gao Jun. Newton–Gauss curvature matrix based cDBN for online edible fungus drying prediction model. Future Generation Computer Systems, 81(??): 273–279, April 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17315157.

Jiang:2014:GTG

[JWW14] Wenjun Jiang, Guojun Wang, and Jie Wu. Generating trusted graphs for trust evaluation in online social networks. Future Generation Computer Systems, 31(??):48–58, February 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1200146X.

Jedari:2019:SBW

[JXC^{+}19] Behrouz Jedari, Feng Xia, Honglong Chen, Sajal K. Das, Amr Tolba, and Zafer Al-Makhadmeh. A social-based watchdog system to detect selfish nodes in opportunistic mobile networks. Future Generation Computer Systems, 92(??):777–788, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312803.

Jiao:2019:NNN

[JXZ^{+}19] Xu Jiao, Yingyuan Xiao, Wenguang Zheng, Hongya Wang, and Ching-Hsien Hsu. A novel next new point-of-interest recommendation system based on simulated user travel decision-making process. Future Generation Computer Systems, 100(??): 982–993, November 2019.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X19302432.

James:2015:SIQ

[JY15] Anne James and Norlaily Yaacob. Special issue: Quality of service in grid and cloud 2015. Future Generation Computer Systems, 50(??):1–2, September 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15001107.

Jayaraman:2017:PPI

[JYY^{+}17] Prem Prakash Jayaraman, Xuechao Yang, Ali Yavari, Dimitrios Georgakopoulos, and Xun Yi. Privacy preserving Internet of Things: From privacy techniques to a blueprint architecture and efficient implementation. Future Generation Computer Systems, 76(??):540–549, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17303308.

Jian:2018:MSI

[JYZ^{+}18] Lihua Jian, Xiaomin Yang, Zhili Zhou, Kai Zhou, and Kai Liu. Multiscale image fusion through rolling guidance filter. Future Generation Computer Systems, 83(??):310–325, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17317892.

Jin:2019:PAO

[JYZ^{+}19] Yue Jin, Han Yu, Yin Zhang, Ning Pan, and Mohsen Guizani. Predictive analysis in outpatients assisted by the Internet of Medical Things. Future Generation Computer Systems, 98(??):219– 226, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18327717.

Jin:2018:FIF

[JZJ^{+}18] Hao Jin, Ke Zhou, Hong Jiang, Dongliang Lei, Ronglei Wei, and Chunhua Li. Full integrity and freshness for cloud data. Future Generation Computer Systems, 80(??):640–652, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301984.

Jiang:2017:MMM

[JZWL17] Yiwei Jiang, Yuqing Zhu, Weili Wu, and Deying Li. Makespan minimization for MapReduce systems with different servers. Future Generation Computer Systems, 67(??):13–21, February 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302618.

Katz:2013:RAS

[KA13] Daniel S. Katz and David Abramson. Recent advances in e-Science. Future Generation Computer Systems, 29(2):548, February 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000969.

Kolomvatsos:2019:MCO

[KA19] Kostas Kolomvatsos and Christos Anagnostopoulos. Multi-criteria optimal task allocation at the edge. Future Generation Computer Systems, 93(??):358–372, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18313736.

Koch:2016:ORC

[KACN16] Fernando Koch, Marcos D.

Assunção, Carlos Cardonha, and Marco A. S. Netto. Optimising resource costs of cloud computing for education. Future Generation Computer Systems, 55(??):473–479, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000758.

Kirkham:2014:RDS

[KADJ14] Tom Kirkham, Django Armstrong, Karim Djemame, and Ming Jiang. Risk driven Smart Home resource management using cloud services. Future Generation Computer Systems, 38(??):13– 22, September 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001714.

Krzywda:2018:PPT

[KAEC^{+}18] Jakub Krzywda, Ahmed Ali-Eldin, Trevor E. Carlson, Per-Olov Östberg, and Erik Elmroth. Powerperformance tradeoffs in data center servers: DVFS, CPU pinning, horizontal, and vertical scaling. Future Generation Computer Systems, 81(??):114–128, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17304910.

Khan:2019:ECS

[KAH^{+}19] Wazir Zada Khan, Ejaz Ahmed, Saqib Hakak, Ibrar Yaqoob, and Arif Ahmed. Edge computing: a survey. Future Generation Computer Systems, 97(??):219–235, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18319903.

Karpowicz:2018:DIE

[KANS18] Michal P. Karpowicz, Piotr Arabas, and Ewa Niewiadomska-Szynkiewicz. Design and implementation of energy-aware application-specific CPU frequency governors for the heterogeneous distributed computing systems. Future Generation Computer Systems, 78 (part 1)(??):302– 315, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301212.

Katzouris:2019:POE

[KAP19] Nikos Katzouris, Alexander Artikis, and Georgios Paliouras. Parallel online event calculus learning for complex event recognition. Future Generation Computer Systems, 94(??): 468–478, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17320216.

Kourtesis:2014:SBQ

[KARP14] Dimitrios Kourtesis, Jose María Alvarez-Rodríguez, and Iraklis Paraskakis. Semanticbased QoS management in cloud systems: Current status and future challenges. Future Generation Computer Systems, 32(??): 307–323, March 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300232X.

Kousiouris:2018:IIL

[KAS^{+}18] George Kousiouris, Adnan Akbar, Juan Sancho, Paula Ta-shma, Alexandros Psychas, Dimosthenis Kyriazis, and Theodora Varvarigou. An integrated information lifecycle management framework for exploiting social network data to identify dynamic large crowd concentration events in smart cities applications. Future Generation Computer Systems, 78 (part 2)(??):516–530, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17300055.

Khorshed:2012:SGT

[KAW12] Md. Tanzim Khorshed, A. B. M. Shawkat Ali, and Saleh A. Wasimi. A survey on gaps, threat remediation challenges and some thoughts for proactive attack detection in cloud computing. Future Generation Computer Systems, 28(6):833–851, June 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000180.

Kroiss:2016:LBM

[KB16] Christian Kroiß and Tomás Bures. Logic-based modeling of information transfer in cyber-physical multiagent systems. Future Generation Computer Systems, 56(??):124–139, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002939.

Khan:2018:SAA

[KB18] Asmat Ullah Khan and Susmit Bagchi. Software architecture and algorithm for reliable RPC for geodistributed mobile computing systems. Future Generation Computer Systems, 86(??):185–198, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18300451.

Koulouzis:2016:SAF

[KBB^{+}16] Spiros Koulouzis, Adam S. Z. Belloum, Marian T. Bubak, Zhiming Zhao, Miroslav Zivković, and Cees T. A. M. de Laat. SDN-aware federation of distributed data. Future Generation Computer Systems, 56(??):64–76, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500312X.

Koning:2018:CEB

[KBdLG18] Ralph Koning, Nick Buraglio, Cees de Laat, and Paola Grosso. CoreFlow: Enriching Bro security events using network traffic monitoring data. Future Generation Computer Systems, 79 (part 1)(??):235– 242, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17305952.

Kerbyson:2014:PCC

[KBVH14] Darren J. Kerbyson, Kevin J. Barker, Abhinav Vishnu, and Adolfy Hoisie. A performance comparison of current HPC systems: Blue Gene/Q, Cray XE6 and InfiniBand systems. Future Generation Computer Systems, 30(??):291– 304, January 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001337.

Kaur:2014:REF

[KC14] Pankaj Deep Kaur and Inderveer Chana. A resource elasticity framework for QoS-aware execution of cloud applications. Future Generation Computer Systems, 37(??): 14–25, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000430.

Kim:2019:APC

[KC19a] Soohyung Kim and Yon Dohn Chung. An anonymization protocol for continuous and dynamic privacypreserving data collection. Future Generation Computer Systems, 93(??):1065–1073, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17319520.

Kiran:2019:UFH

[KC19b] Mariam Kiran and Anshuman Chhabra. Understanding flows in highspeed scientific networks: a Netflow data study. Future Generation Computer Systems, 94(??):72– 79, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18302322.

Kuo:2018:SKE

[KCCL18] Chien-Ting Kuo, Po-Wen Chi, Victor Chang, and Chin-Laung Lei. SFaaS: Keeping an eye on IoT fusion environment with security fusion as a service. Future Generation Computer Systems, 86(??):1424–1436, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324834.

Krefting:2013:GBS

[KCH^{+}13] Dagmar Krefting, Sebastian Canisius, Andreas Hoheisel, Helena Loose, Thomas Tolxdorff, and Thomas Penzel. Grid based sleep research — analysis of polysomnographies using a grid infrastructure. Future Generation Computer Systems, 29(7):1671– 1679, September 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X10000531.

Kianpisheh:2016:RDS

[KCK16] Somayeh Kianpisheh, Nasrolah Moghadam Charkari, and Mehdi Kargahi. Reliabilitydriven scheduling of time/ cost-constrained grid workflows. Future Generation Computer Systems, 55(??): 1–16, February 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1500237X.

Khernane:2019:OPR

[KCM19] Nesrine Khernane, Jean-François Couchot, and Ahmed Mostefaoui. Optimal power/rate tradeoff for Internet of Multimedia Things lifetime maximization under dynamic links capacity. Future Generation Computer Systems, 93(??):737–750, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17304302.

Kim:2014:ECS

[KCS14] Nakku Kim, Jungwook Cho, and Euiseong Seo. Energy-credit scheduler: an energy-aware virtual machine scheduler for cloud systems. Future Generation Computer Systems, 32(??):128–137, March 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1200115X.

Kokkinos:2011:EDC

[KCV11] P. Kokkinos, K. Christodoulopoulos, and E. Varvarigos. Efficient data consolidation in grid networks and performance analysis. Future Generation Computer Systems, 27(2):182–194, February 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kiss:2019:CAQ

[KDG^{+}19] Tamas Kiss, James DesLauriers, Gregoire Gesmier, Gabor Terstyanszky, Gabriele Pierantoni, Osama Abu Oun, Simon J. E. Taylor, Anastasia Anagnostou, and Jozsef Kovacs. A cloud-agnostic queuing system to support the implementation of deadlinebased application execution policies. Future Generation Computer Systems, 101(??):99–111, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19303814.

Koning:2019:MES

[KdGP^{+}19] R. Koning, B. de Graaff, G. Polevoy, R. Meijer, C. de Laat, and P. Grosso. Measuring the efficiency of SDN mitigations against attacks on computer infrastructures. Future Generation Computer Systems, 91(??):144–156, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18302255.

Kabat:2016:HPF

[KDHP16] Milan Kabát, Vojtech David, Petr Holub, and Martin Pulec. Highperformance forward error correction: Enabling multi-gigabit flows and beyond on commodity GPU and CPU hardware in presence of packet loss. Future Generation Computer Systems, 54(??):326–335, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000990.

Kotian:2017:ITP

[KESL17] R. Kotian, G. Exarchakos, S. Stavros, and A. Liotta. Impact of transmission power control in multi-hop networks. Future Generation Computer Systems, 75(??):94–107, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303910.

Krzywicki:2014:CAD

[KFBKD14] D. Krzywicki, L. Faber, A. Byrski, and M. Kisiel-Dorohinicki. Computing agents for decision support systems. Future Generation Computer Systems, 37(??):390–400, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000272.

Kramer:2019:ISA

[KFK19] Michel Krämer, Sven Frese, and Arjan Kuijper. Implementing secure applications in smart city clouds using microservices. Future Generation Computer Systems, 99(??):308–320, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18304539.

Koning:2011:UOR

[KGdL11] Ralph Koning, Paola Grosso, and Cees de Laat. Using ontologies for resource description in the CineGrid Exchange. Future Generation Computer Systems, 27(7):960–965, July 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kuang:2018:QRF

[KGLY18] Zhikai Kuang, Songtao Guo, Jiadi Liu, and Yuanyuan Yang. A quick-response framework for multi-user computation offloading in mobile cloud computing. Future Generation Computer Systems, 81(??):166–176, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1730465X.

Kochovski:2019:TMB

[KGS^{+}19] Petar Kochovski, Sandi Gec, Vlado Stankovski, Marko Bajec, and Pavel D. Drobintsev. Trust management in a blockchain based fog computing platform with trustless smart oracles. Future Generation Computer Systems, 101(??):747–759, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301281.

Kougka:2015:PAE

[KGT15] Georgia Kougka, Anastasios Gounaris, and Kostas Tsichlas. Practical algorithms for execution engine selection in data flows. Future Generation Computer Systems, 45(??):133– 148, April 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002453.

Kolodziej:2014:ADI

[KGVW14] Joanna Kolodziej, Horacio González-Vélez, and Lizhe Wang. Advances in data-intensive modelling and simulation. Future Generation Computer Systems, 37(??):282–283, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000314.

Kobusinska:2018:TIR

[KH18a] Anna Kobusińska and Ching-Hsien Hsu. Towards increasing reliability of clouds environments with RESTful web services. Future Generation Computer Systems, 87(??):502–513, October 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17316709.

Koo:2018:PPD

[KH18b] Dongyoung Koo and Junbeom Hur. Privacypreserving deduplication of encrypted data with dynamic ownership management in fog computing. Future Generation Computer Systems, 78 (part 2)(??):739–752, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301309.

Khaled:2019:ICF

[KH19] Ahmed E. Khaled and Sumi Helal. Interoperable communication framework for bridging RESTful and topic-based communication in IoT. Future Generation Computer Systems, 92(??):628–643, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17317387.

Khan:2012:HPB

[Kha12] Muhammad Khurram Khan.

High performance biometrics recognition algorithms and systems. Future Generation Computer Systems, 28(1):210–211, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1100149X.

Kang:2013:AHJ

[KHG13] Woochul Kang, H. Howie Huang, and Andrew Grimshaw. Achieving high job execution reliability using underutilized resources in a computational economy. Future Generation Computer Systems, 29(3):763– 775, March 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001811.

Kern:2018:SSP

[KHG^{+}18] Eva Kern, Lorenz M. Hilty, Achim Guldner, Yuliyan V. Maksimov, Andreas Filler, Jens Gröger, and Stefan Naumann. Sustainable software products-towards assessment criteria for resource and energy efficiency. Future Generation Computer Systems, 86(??):199–210, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17314188.

Kershaw:2010:SCB

[KHJ10] Stephen Kershaw and Richard Hughes-Jones. A study of constant bit-rate data transfer over TCP/IP. Future Generation Computer Systems, 26(1):128– 134, January 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Khalil-Hani:2013:BEB

[KHMB13]  Mohamed  Khalil-Hani,

Muhammad N. Marsono, and Rabia Bakhteri. Biometric encryption based on a fuzzy vault scheme with a fast chaff generation algorithm. Future Generation Computer Systems, 29(3): 800–810, March 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000350.

Koh:2019:ADC

[KHO^{+}19] Joel En Wei Koh, Yuki Hagiwara, Shu Lih Oh, Jen Hong Tan, Edward J. Ciaccio, Peter H. Green, Suzanne K. Lewis, and U. Rajendra Acharya. Automated diagnosis of celiac disease using DWT and nonlinear features with video capsule endoscopy images. Future Generation Computer Systems, 90(??): 86–93, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18313633.

Hu:2018:PBG

[kHsZwJW18] Ke kun Hu, Guo sun Zeng, Huo wen Jiang, and Wei Wang. Partitioning big graph with respect to arbitrary proportions in a streaming manner. Future Generation Computer Systems, 80(??): 1–11, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17303424.

Kabir:2018:NST

[KHWZ18] Enamul Kabir, Jiankun Hu, Hua Wang, and Guangping Zhuo. A novel statistical technique for intrusion detection systems. Future Generation Computer Systems, 79 (part 1)(??):303–318, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301371.

Kastrati:2019:PAM

[KI19] Zenun Kastrati and Ali Shariq Imran. Performance analysis of machine learning classifiers on improved concept vector space models. Future Generation Computer Systems, 96(??):552– 562, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1831745X.

Khan:2017:DPS

[KIAD17] Farrukh Aslam Khan, Muhammad Imran, Haider Abbas, and Muhammad Hanif Durad. A detection and prevention system against collaborative attacks in mobile ad hoc networks. Future Generation Computer Systems, 68(??):416– 427, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302400.

Kumar:2012:CLA

[KIC12] Neeraj Kumar, Rahat Iqbal, and Naveen Chilamkurti. Capacity and load-aware service discovery with service selection in peer-to-peer grids. Future Generation Computer Systems, 28(7):1090–1099, July 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002317.

Kido:2016:SBT

[KID^{+}16] Yoshiyuki Kido, Kohei Ichikawa, Susumu Date, Yasuhiro Watashiba, Hirotake Abe, Hiroaki Yamanaka, Eiji Kawai, Haruo Takemura, and Shinji Shimojo. SAGE-based Tiled Display Wall enhanced with dynamic routing functionality triggered by user interaction. Future Generation Computer Systems, 56(??):303–314, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003131.

Khan:2019:HCS

[KIJ^{+}19] Sana Ullah Khan, Naveed Islam, Zahoor Jan, Ikram Ud Din, Atif Khan, and Yasir Faheem. An e-health care services framework for the detection and classification of breast cancer in breast cytology images as an IoMT application. Future Generation Computer Systems, 98(??):286–296, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18325536.

Kim:2014:CAM

[Kim14] Haeng-Kon Kim. Convergence agent model for developing u-healthcare systems. Future Generation Computer Systems, 35(??): 39–48, June 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002446.

Kim:2018:RTB

[Kim18] Pyoung Won Kim. Realtime bio-signal-processing of students based on an intelligent algorithm for Internet of Things to assess engagement levels in a classroom. Future Generation Computer Systems, 86(??):716–722, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17325037.

Kumar:2015:BCG

[KIMR15] Neeraj Kumar, Rahat Iqbal, Sudip Misra, and Joel J. P. C. Rodrigues. Bayesian coalition game for contention-aware reliable data forwarding in vehicular mobile cloud. Future Generation Computer Systems, 48(??):ibc, July 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002064.

Khanli:2011:PDR

[KIS11] Leyli Mohammad Khanli, Ayaz Isazadeh, and Tahmuras N. Shishavan. PHFS: a dynamic replication method, to decrease access latency in the multi-tier data grid. Future Generation Computer Systems, 27 (3):233–244, March 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kondo:2012:PSI

[KJ12] Derrick Kondo and Bahman Javadi. Preface to the special issue on volunteer computing and desktop grids. Future Generation Computer Systems, 28(6): 852–853, June 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002329.

Kyriazis:2018:CFD

[KJ18] Dimosthenis Kyriazis and Keith G. Jeffery. Cloud forward: From distributed to complete computing. Future Generation Computer Systems, 78 (part 1)(??): 87–88, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17319647.

Kipp:2012:LGP

[KJFS12] Alexander Kipp, Tao Jiang, Mariagrazia Fugini, and Ioan Salomie. Layered Green Performance Indicators. Future Generation Computer Systems, 28(2):478–489, February 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11000860.

Khachana:2011:RAP

[KJI11] Rose Tshoganetso Khachana, Anne James, and Rahat Iqbal. Relaxation of ACID properties in Au-TrA, the adaptive userdefined transaction relaxing approach. Future Generation Computer Systems, 27(1):58–66, January 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kertesz:2010:GNM

[KK10a] Attila Kertész and Péter Kacsuk. GMBS: a new middleware service for making grids interoperable. Future Generation Computer Systems, 26(4):542– 553, April 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Koseoglu:2010:JRN

[KK10b] Mehmet Koseoglu and Ezhan Karasan. Joint resource and network scheduling with adaptive offset determination for optical burst switched grids. Future Generation Computer Systems, 26(4):576– 589, April 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Khanli:2011:FFR

[KK11] Leyli Mohammad Khanli and Saeed Kargar. FRDT: Footprint Resource Discovery Tree for grids. Future Generation Computer Systems, 27(2):148–156, February 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kunsemoller:2014:GTA

[KK14] Jörn Künsemöller and Holger Karl. A game-theoretic approach to the financial benefits of infrastructureas-a-service. Future Generation Computer Systems, 41(??):44–52, December 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000508.

Krol:2016:SSS

[KK16] Dariusz Król and Jacek Kitowski. Self-scalable services in service oriented software for cost-effective data farming. Future Generation Computer Systems, 54(??):1–15, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002265.

Karatas:2019:FBD

[KK19] Firat Karatas and Ibrahim Korpeoglu. Fog-Based Data Distribution Service (F-DAD) for Internet of Things (IoT) applications. Future Generation Computer Systems, 93(??):156– 169, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18315310.

Kumar:2018:FFA

[KKA18] Ajit Kumar, K. S. Kuppusamy, and G. Aghila. FAMOUS: Forensic Analysis of MObile devices Using Scoring of application permissions. Future Generation Computer Systems, 83(??):158–172, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17323257.

Kaseb:2019:ITI

[KKAS19] Mostafa R. Kaseb, Mohamed H. Khafagy, Ihab A. Ali, and ElSayed M. Saad. An improved technique for increasing availability in Big Data replication. Future Generation Computer Systems, 91(??):493–505, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18311129.

Kertesz:2014:ISA

[KKB14] A. Kertesz, G. Kecskemeti, and I. Brandic. An interoperable and self-adaptive approach for SLA-based service virtualization in heterogeneous Cloud environments. Future Generation Computer Systems, 32(??): 54–68, March 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001124.

Kaur:2018:EMP

[KKB18] Harmanjeet Kaur, Neeraj Kumar, and Shalini Batra. An efficient multiparty scheme for privacy preserving collaborative filtering for healthcare recommender system. Future Generation Computer Systems, 86(??):297–307, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17327012.

Khan:2019:WOO

[KKB^{+}19] Arsla Khan, Saud Khan, Sobia Baig, Hafiz Muhammad Asif, and Soo Young Shin. Wavelet OFDM with overlap FDE for non-Gaussian channels in precoded NOMA based systems. Future Generation Computer Systems, 97(??): 165–179, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18318272.

Khan:2019:SSC

[KKBK19] Awais Khan, Taeuk Kim, Hyunki Byun, and Youngjae Kim. SciSpace: a scientific collaboration workspace for geo-distributed HPC data centers. Future Generation Computer Systems, 101(??):398–409, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326025.

Kajdanowicz:2014:PPL

[KKI14] Tomasz Kajdanowicz, Przemyslaw Kazienko, and Wojciech Indyk. Parallel processing of large graphs. Future Generation Computer Systems, 32(??):324–337, March 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001726.

Kovacs:2010:AMT

[KKJJ10] József Kovács, Peter Kacsuk, Radoslaw Januszewski, and Gracjan Jankowski. Application and middleware transparent checkpointing with TCKPT on ClusterGrids. Future Generation Computer Systems, 26(3):498–503, March 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kiss:2019:MMB

[KKK^{+}19] Tamas Kiss, Peter Kacsuk, Jozsef Kovacs, Botond Rakoczi, Akos Hajnal, Attila Farkas, Gregoire Gesmier, and Gabor Terstyanszky. MiCADO — Microservice-based Cloud Application-level Dynamic Orchestrator. Future Generation Computer Systems, 94(??):937–946, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310506.

Khan:2013:TSM

[KKKM13] Abdul Nasir Khan, M. L. Mat Kiah, Samee U. Khan, and Sajjad A. Madani. Towards secure mobile cloud computing: a survey. Future Generation Computer Systems, 29(5):1278–1299, July 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001598.

Kritikos:2017:TSE

[KKKM17] Kyriakos Kritikos, Tom Kirkham, Bartosz Kryza, and Philippe Massonet. Towards a security-enhanced PaaS platform for multicloud applications. Future Generation Computer Systems, 67(??):206–226, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303880.

See reprint [KKKM18].

Kritikos:2018:RTS

[KKKM18] Kyriakos Kritikos, Tom Kirkham, Bartosz Kryza, and Philippe Massonet. Reprint of “Towards a security-enhanced PaaS platform for multi-cloud applications”. Future Generation Computer Systems, 78 (part 1)(??):155–175, January 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306343.

See [KKKM17].

Kovacs:2011:UPD

[KKL11] József Kovács, Peter Kacsuk, and Andre Lomaka. Using a private desktop grid system for accelerating drug discovery. Future Generation Computer Systems, 27(6):657–666, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Karmakar:2018:EDD

[KKN18] Gour Karmakar, Joarder Kamruzzaman, and Nusrat Nowsheen. An efficient data delivery mechanism for AUV-based ad hoc UASNs. Future Generation Computer Systems, 86(??):1193–1208, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17323439.

Ko:2019:UEE

[KKP19] Ilju Ko, Dongho Kim, and Jong Hyuk Park. A user experience environment model for human activity simulation. Future Generation Computer Systems, 96(??):660–666, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17317338.

Kovalchuk:2018:DDD

[KKS^{+}18a] Sergey V. Kovalchuk, Evgeniy Krotov, Pavel A. Smirnov, Denis A. Nasonov, and Alexey N. Yakovlev. Distributed data-driven platform for urgent decision making in cardiological ambulance control. Future Generation Computer Systems, 79 (part 1)(??):144– 154, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630348X.

Kumar:2018:SEL

[KKS18b] Meeta Kumar, Anand J.

Kulkarni, and Suresh Chandra Satapathy. Socio evolution & learning optimization algorithm: a socio-inspired optimization methodology. Future Generation Computer Systems, 81(??):252–272, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17317259.

Kierzynka:2017:EES

[KKvdB^{+}17] Michal Kierzynka, Lars Kosmann, Micha vor dem Berge, Stefan Krupop, Jens Hagemeyer, René Griessl, Meysam Peykanu, and Ariel Oleksiak. Energy efficiency of sequence alignment tools — software and hardware perspectives. Future Generation Computer Systems, 67(??):455– 465, February 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301169.

Kolodziej:2014:SEP

[KKW^{+}14] Joanna Kolodziej, Samee Ullah Khan, Lizhe Wang, Marek Kisiel-Dorohinicki, Sajjad A. Madani, Ewa Niewiadomska-Szynkiewicz, Albert Y. Zomaya, and Cheng-Zhong Xu. Security, energy, and performanceaware resource allocation mechanisms for computational grids. Future Generation Computer Systems, 31(??):77–92, February 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001823.

Kobusinska:2018:ETI

[KLH^{+}18] Anna Kobusińska, Carson Leung, Ching-Hsien Hsu, Raghavendra S., and Victor Chang. Emerging trends, issues and challenges in Internet of Things, big data and cloud computing. Future Generation Computer Systems, 87(??):416–419, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18311270.

Kong:2019:STT

[KLJS19] Fanhui Kong, Jian Li, Bin Jiang, and Houbing Song. Short-term traffic flow prediction in smart multimedia system for Internet of Vehicles based on deep belief network. Future Generation Computer Systems, 93(??):460–472, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18320326.

Kong:2019:PPS

[KLMB19] Qinglei Kong, Rongxing Lu, Maode Ma, and Haiyong Bao. A privacypreserving sensory data sharing scheme in Internet of Vehicles. Future Generation Computer Systems, 92(??):644–655, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17317004.

Ko:2019:CEC

[KLP19] Haneul Ko, Jaewook Lee, and Sangheon Pack. CG-E2S2: Consistency-guaranteed and energy-efficient sleep scheduling algorithm with data aggregation for IoT. Future Generation Computer Systems, 92(??): 1093–1102, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

https://www.sciencedirect.
com/science/article/pii/
S0167739X17305824.

Kumar:2018:CIB

[KLV^{+}18] Priyan Malarvizhi Kumar, S. Lokesh, R. Varatharajan, Gokulnath Chandra Babu, and P. Parthasarathy. Cloud and IoT based disease prediction and diagnosis system for healthcare using fuzzy neural classifier. Future Generation Computer Systems, 86(??):527–534, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18303753.

Kumari:2016:UFM

[KLW^{+}16] Saru Kumari, Xiong Li, Fan Wu, Ashok Kumar Das, Hamed Arshad, and Muhammad Khurram Khan. A user friendly mutual authentication and key agreement scheme for wireless sensor networks using chaotic maps. Future Generation Computer Systems, 63(??):56–75, October 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300930.

Kumari:2017:DPS

[KLW^{+}17] Saru Kumari, Xiong Li, Fan Wu, Ashok Kumar Das, Kim-Kwang Raymond Choo, and Jian Shen. Design of a provably secure biometricsbased multi-cloud-server authentication scheme. Future Generation Computer Systems, 68(??):320–330, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303776.

Kettimuthu:2018:TPD

[KLW^{+}18] Rajkumar Kettimuthu, Zhengchun Liu, David Wheeler, Ian Foster, Katrin Heitmann, and Franck Cappello. Transferring a petabyte in a day. Future Generation Computer Systems, 88(??):191–198, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18302280.

Kamal:2016:WAI

[KMB16] Joarder Kamal, Manzur Murshed, and Rajkumar Buyya. Workload-aware incremental repartitioning of shared-nothing distributed databases for scalable OLTP applications. Future Generation Computer Systems, 56(??):421– 435, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003040.

Kang:2017:CTS

[KMB^{+}17] Guiyeom Kang, Claudio Márquez, Ana Barat, Annette T. Byrne, Jochen H. M. Prehn, Joan Sorribes, and Eduardo César. Colorectal tumour simulation using agent based modelling and high performance computing. Future Generation Computer Systems, 67(??):397–408, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300723.

Kohwalter:2018:FIS

[KMC18] Troy Kohwalter, Leonardo Murta, and Esteban Clua. Filtering irrelevant sequential data out of game session telemetry though similarity collapses. Future Generation Computer Systems, 84(??):108–122, July 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17320897.

Khanli:2011:ARL

[KMI11] Leyli Mohammad Khanli, Farnaz Mahan, and Ayaz Isazadeh. Active rule learning using decision tree for resource management in Grid computing. Future Generation Computer Systems, 27(6):703–710, June 2011. CODEN FGSEVI.

ISSN 0167-739X (print), 1872-7115 (electronic).

Kim:2018:VDM

[KMJ18] Taesik Kim, Hong Min, and Jinman Jung. Vehicular datacenter modeling for cloud computing: Considering capacity and leave rate of vehicles. Future Generation Computer Systems, 88(??):363–372, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18300487.

Kousiouris:2014:DBB

[KMK^{+}14] George Kousiouris, Andreas Menychtas, Dimosthenis Kyriazis, Spyridon Gogouvitis, and Theodora Varvarigou. Dynamic, behavioral-based estimation of resource provisioning based on highlevel application terms in Cloud platforms. Future Generation Computer Systems, 32(??):27–40, March 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001057.

K:2019:ODL

[KMK^{+}19] Lakshmanaprabu S. K., Sachi Nandan Mohanty, Shankar K., Arunkumar N., and Gustavo Ramirez. Optimal deep learning model for classification of lung cancer on CT images. Future Generation Computer Systems, 92(??):374– 382, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18317011.

Koroniotis:2019:TDR

[KMST19] Nickolaos Koroniotis, Nour Moustafa, Elena Sitnikova, and Benjamin Turnbull. Towards the development of realistic botnet dataset in the Internet of Things for network forensic analytics: Bot-IoT dataset. Future Generation Computer Systems, 100(??):779–796, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18327687.

Kessaci:2014:MSL

[KMT14] Yacine Kessaci, Nouredine Melab, and El-Ghazali Talbi. A multi-start local search heuristic for an energy efficient VMs assignment on top of the OpenNebula cloud manager. Future Generation Computer Systems, 36(??): 237–256, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001520.

Kos:2019:CWC

[KMU19] Anton Kos, Veljko Milutinović, and Anton Umek. Challenges in wireless communication for connected sensors and wearable devices used in sport biofeedback applications. Future Generation Computer Systems, 92(??):582–592, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17316692.

Kovacs:2015:BGC

[KMV^{+}15] József Kovács, Attila Csaba Marosi, Ádám Visegrádi, Zoltán Farkas, Peter Kacsuk, and Róbert Lovas. Boosting gLite with cloud augmented volunteer computing. Future Generation Computer Systems, 43–44(??):12–23, February 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001897.

Katz:2016:ASC

[KMZJ16] Daniel S. Katz, Andre Merzky, Zhao Zhang, and Shantenu Jha. Application skeletons: Construction and use in eScience. Future Generation Computer Systems, 59(??):114– 124, June 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003143.

Kijsipongse:2010:PPS

[KN10] Ekasit Kijsipongse and Sudsanguan Ngamsuriyaroj. Placing pipeline stages on a Grid: Single path and multipath pipeline execution. Future Generation Computer Systems, 26(1):50–62, January 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kalyuzhnaya:2018:TSB

[KNI^{+}18] Anna V. Kalyuzhnaya, Denis Nasonov, Sergey V. Ivanov, Sergey S. Kosukhin, and Alexander V. Boukhanovsky. Towards a scenario-based solution for extreme metocean event simulation applying urgent computing. Future Generation Computer Systems, 79 (part 2)(??):604– 617, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17311081.

Kaneko:2011:DIL

[KO11] Kunitake Kaneko and Naohisa Ohta. Design and implementation of live image file feeding to dome theaters. Future Generation Computer Systems, 27(7):944–951, July 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kolomvatsos:2018:IUD

[Kol18] Kostas Kolomvatsos. An intelligent, uncertainty driven management scheme for software updates in pervasive IoT applications. Future Generation Computer Systems, 83(??):116– 131, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X16305301.

Kranjc:2017:COW

[KOP^{+}17] Janez Kranjc, Roman Orac, Vid Podpecan, Nada Lavrac, and Marko Robnik-Sikonja. Clowd-Flows: Online workflows for distributed big data mining. Future Generation Computer Systems, 68(??): 38–58, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302709.

Khalil:2018:ETA

[KOT18] Enan A. Khalil, Suat Ozdemir, and Suleyman Tosun. Evolutionary task allocation in Internet of Things-based application domains. Future Generation Computer Systems, 86(??):121–133, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1732071X.

Kim:2012:SHM

[KP12] Kyungtae Kim and Changwoo Pyo. Securing heap memory by data pointer encoding. Future Generation Computer Systems, 28(8):1252–1257, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000094.

Kushwaha:2018:LBB

[KP18] Neetu Kushwaha and Millie Pant. Link based BPSO for feature selection in big data text clustering. Future Generation Computer Systems, 82(??):190–199, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17321854.

Khan:2017:TSS

[KPA17] Zaheer Khan, Zeeshan Pervez, and Abdul Ghafoor Abbasi. Towards a secure service provisioning framework in a smart city environment. Future Generation Computer Systems, 77(??):112–135, December 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17313687.

Kobusinska:2018:BDF

[KPB18] Anna Kobusińska, Kamil Pawluczuk, and Jerzy Brzeziński. Big data fingerprinting information analytics for sustainability. Future Generation Computer Systems, 86 (??):1321–1337, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17329965.

Krishnamurthi:2019:APD

[KPG19] Rajalakshmi Krishnamurthi, Rizwan Patan, and Amir H. Gandomi. Assistive pointer device for limb impaired people: a novel frontier point method for hand movement recognition. Future Generation Computer Systems, 98(??):650–659, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18323082.

Kim:2019:AJA

[KPJ19] Hyun-Woo Kim, Jong Hyuk Park, and Young-Sik Jeong. Adaptive job allocation scheduler based on usage pattern for computing offloading of IoT. Future Generation Computer Systems, 98(??):18– 24, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18327638.

Kiran:2018:EIC

[KPM^{+}18] Mariam Kiran, Eric Pouyoul, Anu Mercian, Brian Tierney, Chin Guok, and Inder Monga. Enabling intent to configure scientific networks for high performance demands. Future Generation Computer Systems, 79 (part 1)(??):205– 214, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1730626X.

Khanduzi:2018:DEA

[KPS18] Raheleh Khanduzi, M. Reza Peyghami, and Arun Kumar Sangaiah. Data envelopment analysis and interdiction median problem with fortification for enabling IoT technologies to relieve potential attacks. Future Generation Computer Systems, 79 (part 3)(??):928–940, February 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17313110.

Kim:2014:AIM

[KR14] Minwoo Kim and Won Woo Ro. Architectural investigation of matrix data layout on multicore processors. Future Generation Computer Systems, 37(??): 64–75, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002380.

K:2019:OSM

[KR19] Senthilkumar K. and Ramesh Ramadoss. Optimized scheduling of multicore ECU architecture with biosecurity CAN network using AUTOSAR. Future Generation Computer Systems, 98(??):1–11, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18322362.

Kayes:2019:CAA

[KRD^{+}19] A. S. M. Kayes, Wenny Rahayu, Tharam Dillon, Elizabeth Chang, and Jun Han. Context-aware access control with imprecise context characterization for cloud-based data resources. Future Generation Computer Systems, 93(??):237– 255, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18300785.

Khanli:2012:NST

[KRZ12] Leyli Mohammad Khanli, Shiva Razzaghzadeh, and Sadegh Vahabzadeh Zargari. A new step toward load balancing based on competency rank and transitional phases in Grid networks. Future Generation Computer Systems, 28(4): 682–688, April 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11002299.

Khan:2019:MCS

[KRZ^{+}19] Fazlullah Khan, Ateeq Ur Rehman, Jiangbin Zheng, Mian Ahmad Jan, and Muhammad Alam. Mobile crowdsensing: a survey on privacy-preservation, task management, assignment models, and incentives mechanisms. Future Generation Computer Systems, 100(??):456–472, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18315632.

Kanbayashi:2011:DAS

[KS11] Ryo Kanbayashi and Mitsuhisa Sato. A distributed architecture of Sensing Web for sharing open sensor nodes. Future Generation Computer Systems, 27(5):643–648, May 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Knepper:2017:FOS

[KS17a] Richard Knepper and Matthew Standish. Forward Observer system for radar data workflows: Big data management in the field. Future Generation Computer Systems, 76(??):92–97, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310567.

Kozhirbayev:2017:PCC

[KS17b] Zhanibek Kozhirbayev and Richard O. Sinnott. A performance comparison of container-based technologies for the Cloud. Future Generation Computer Systems, 68(??):175–182, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303041.

Kanagaraj:2018:SAR

[KS18a] K. Kanagaraj and S. Swamynathan. Structure aware resource estimation for effective scheduling and execution of data intensive workflows in cloud. Future Generation Computer Systems, 79 (part 3)(??):878– 891, February 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308111.

Khan:2018:ISR

[KS18b] Minhaj Ahmad Khan and Khaled Salah. IoT security: Review, blockchain solutions, and open challenges. Future Generation Computer Systems, 82(??):395– 411, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17315765.

Kumar:2018:WPC

[KS18c] Jitendra Kumar and Ashutosh Kumar Singh. Workload prediction in cloud using artificial neural network and adaptive differential evolution. Future Generation Computer Systems, 81(??): 41–52, April 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300444.

Kumar:2018:MPA

[KS18d] Santosh Kumar and Sanjay Kumar Singh. Monitoring of pet animal in smart cities using animal biometrics. Future Generation Computer Systems, 83(??):553–563, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X16307385.

Kee:2019:BSO

[KS19] Kerk F. Kee and Andrew R. Schrock. Best social and organizational practices of successful science gateways and cyberinfrastructure projects. Future Generation Computer Systems, 94(??):795–801, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17310725.

Knight:2019:TES

[KSC^{+}19] Louise Knight, Polona Stefanic, Matej Cigale, Andrew C. Jones, and Ian Taylor. Towards extending the SWITCH platform for time-critical, cloud-based CUDA applications: Job scheduling parameters influencing performance. Future Generation Computer Systems, 100(??):542–556, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18311014.

Katsaros:2013:SFE

[KSF^{+}13] Gregory Katsaros, Josep Subirats, J. Oriol Fitó, Jordi Guitart, Pierre Gilet, and Daniel Espling. A service framework for energyaware monitoring and VM management in clouds. Future Generation Computer Systems, 29(8):2077–2091, October 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12002269.

Kitamura:2011:BLV

[KSK^{+}11] Masahiko Kitamura, Daisuke Shirai, Kunitake Kaneko, Takahiro Murooka, Tomoko Sawabe, Tatsuya Fujii, and Atsushi Takahara. Beyond 4K: 8K 60p live video streaming to multiple sites. Future Generation Computer Systems, 27(7):952– 959, July 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Khattak:2019:PLS

[KSK^{+}19] Hasan Ali Khattak, Munam Ali Shah, Sangeen Khan, Ihsan Ali, and Muhammad Imran. Perception layer security in Internet of Things. Future Generation Computer Systems, 100(??):144–164, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304194.

Kakarontzas:2011:ACC

[KSS11] G. Kakarontzas, I. K.

Savvas, and I. Stamelos. Agents, clusters and components: a synergistic approach to the GSP. Future Generation Computer Systems, 27(8):999–1010, October 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Keshavarzian:2019:MDR

[KSS19] Alireza Keshavarzian, Saeed Sharifian, and Sanaz Seyedin.

Modified deep residual network architecture deployed on serverless framework of IoT platform based on human activity recognition application. Future Generation Computer Systems, 101(??):14–28, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304376.

Katsaros:2016:EFE

[KSSG16] Gregory Katsaros, Pascal Stichler, Josep Subirats, and Jordi Guitart. Estimation and forecasting of ecological efficiency of virtual machines. Future Generation Computer Systems, 55(??):480–494, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000035.

Kluge:2013:PQS

[KSW^{+}13] Michael Kluge, Stephen Simms, Thomas William, Robert Henschel, Andy Georgi, Christian Meyer, Matthias S. Mueller, Craig A. Stewart, Wolfgang Wünsch, and Wolfgang E. Nagel. Performance and quality of service of data and video movement over a 100 Gbps testbed. Future Generation Computer Systems, 29(1):230–240, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001380.

Koksal:2017:ODD

[KT17] Ömer Köksal and Bedir Tekinerdogan. Obstacles in data distribution service middleware: a systematic review. Future Generation Computer Systems, 68(??): 191–210, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630351X.

Kozsik:2018:FCR

[KTB18] Tamás Kozsik, Melinda Tóth, and István Bozó. Free the conqueror! Refactoring divide-and-conquer functions. Future Generation Computer Systems, 79 (part 2)(??):687– 699, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17309822.

Kecskemeti:2011:AVA

[KTKN11] Gabor Kecskemeti, Gabor Terstyanszky, Peter Kacsuk, and Zsolt Neméth. An approach for virtual appliance distribution for service deployment. Future Generation Computer Systems, 27(3):280–289, March 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Krieger:2017:BOS

[KTTK17] Michael T. Krieger, Oscar Torreno, Oswaldo Trelles, and Dieter Kranzlmüller. Building an open source cloud environment with auto-scaling resources for executing bioinformatics and biomedical workflows. Future Generation Computer Systems, 67(??):329– 340, February 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300218.

Khan:2018:TCF

[KuRAk^{+}18] Imran Khan, Habib ur Rehman, Mohammad Hussein Fayiz Al-khatib, Zahid Anwar, and Masoom Alam. A thin client friendly trusted execution framework for infrastructure-as-a-service clouds. Future Generation Computer Systems, 89(??):239–248, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17329709.

Kokkinos:2012:SER

[KV12] P. Kokkinos and E. A.

Varvarigos. Scheduling efficiency of resource information aggregation in grid networks. Future Generation Computer Systems, 28(1):9–23, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001191.

Kumar:2017:EAC

[KV17] Neetesh Kumar and Deo Prakash Vidyarthi. An energy aware cost effective scheduling framework for heterogeneous cluster system. Future Generation Computer Systems, 71(??):73– 88, June 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300341.

Krefting:2010:SIM

[KVHT10] Dagmar Krefting, Michal Vossberg, Andreas Hoheisel, and Thomas Tolxdorff. Simplified implementation of medical image processing algorithms into a Grid using a workflow management system. Future Generation Computer Systems, 26(4):681– 684, April 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Karbovskii:2018:MAB

[KVK^{+}18] Vladislav Karbovskii, Daniil Voloshin, Andrey Karsakov, Alexey Bezgodov, and Carlos Gershenson. Multimodel agent-based simulation environment for massgatherings and pedestrian dynamics. Future Generation Computer Systems, 79 (part 1)(??):155– 165, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16303739.

Karpenko:2015:AGW

[KVR15] Dmytro Karpenko, Roman Vitenberg, and Alexander L. Read. ATLAS grid workload on NDGF resources: Analysis, modeling, and workload generation. Future Generation Computer Systems, 47(??): 31–47, June 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002581.

Kiljan:2018:ETA

[KVvE18] Sven Kiljan, Harald Vranken, and Marko van Eekelen. Evaluation of transaction authentication methods for online banking. Future Generation Computer Systems, 80(??):430–447, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301352.

Khan:2019:WFF

[KWB19] Muhammad Faizan Khan, Guojun Wang, and Md Zakirul Alam Bhuiyan. Wi-Fi frequency selection concept for effective coverage in collapsed structures. Future Generation Computer Systems, 97(??):409–424, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326591.

Kimball:2016:LBD

[KWK16] Jason Kimball, Tom Wypych, and Falko Kuester. Low bandwidth desktop and video streaming for collaborative tiled display environments. Future Generation Computer Systems, 54(??):336–343, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002320.

Kamruzzaman:2018:ASN

[KWK^{+}18] Joarder Kamruzzaman, Guojun Wang, Gour Karmakar, Iftekhar Ahmad, and Md Zakirul Alam Bhuiyan. Acoustic sensor networks in the Internet of Things applications. Future Generation Computer Systems, 86(??):1167–1169, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18311257.

Kaiiali:2013:GAG

[KWR^{+}13] Mustafa Kaiiali, Rajeev Wankar, C. R. Rao, Arun Agarwal, and Rajkumar Buyya. Grid Authorization Graph. Future Generation Computer Systems, 29(8):1909–1918, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000708.

Kolodziej:2011:EGB

[KX11] Joanna Kolodziej and Fatos Xhafa. Enhancing the genetic-based scheduling in computational grids by a structured hierarchical population. Future Generation Computer Systems, 27(8):1035–1046, October 2011. CODEN FGSEVI.

ISSN 0167-739X (print), 1872-7115 (electronic).

Kong:2016:UTC

[KXS^{+}16] Xiangjie Kong, Zhenzhen Xu, Guojiang Shen, Jinzhong Wang, Qiuyuan Yang, and Benshi Zhang. Urban traffic congestion estimation and prediction based on floating car trajectory data. Future Generation Computer Systems, 61(??):97–107, August 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003611.

Kelarev:2019:MPA

[KYB^{+}19] Andrei Kelarev, Xun Yi, Shahriar Badsha, Xuechao Yang, Leanne Rylands, and Jennifer Seberry. A multistage protocol for aggregated queries in distributed cloud databases with privacy protection. Future Generation Computer Systems, 90(??):368– 380, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18306514.

Kyriazis:2019:MPR

[Kyr19] Dimosthenis Kyriazis. Modelling and prediction of resources and services state evolvement for efficient runtime adaptations. Future Generation Computer Systems, 94(??):1–10, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18313177.

Kang:2019:VBS

[KYZ19] Guosheng Kang, Liqin Yang, and Liang Zhang. Verification of behavioral soundness for artifactcentric business process model with synchronizations. Future Generation Computer Systems, 98(??):503–511, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302201.

Katz:2014:SIE

[KZ14] Daniel S. Katz and Zhao Zhang. Special issue on eScience infrastructure and applications. Future Generation Computer Systems, 36(??):335–337, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000521.

Katz:2017:LER

[KZ17] Daniel S. Katz and Xiaobo Zhou. Leadingedge research in cluster, cloud, and grid computing: Best papers from the IEEE/ACM CCGrid 2015 conference. Future Generation Computer Systems, 72 (??):78–80, July 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303557.

Khan:2011:CRB

[KZA11] Muhammad Khurram Khan, Jiashu Zhang, and Khaled Alghathbar. Challengeresponse-based biometric image scrambling for secure personal identification. Future Generation Computer Systems, 27(4): 411–418, April 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kalai:2018:SCS

[KZA^{+}18] Ahlem Kalaï, Corinne Amel Zayani, Ikram Amous, Wafa Abdelghani, and Florence Sèdes. Social collaborative service recommendation approach based on user’s trust and domainspecific expertise. Future Generation Computer Systems, 80(??):355–367, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17311044.

Ke:2013:RME

[KZCW13] Jia Ke, Yongzhao Zhan, Xiaojun Chen, and Manrong Wang. The retrieval of motion event by associations of temporal frequent pattern growth. Future Generation Computer Systems, 29(1):442–450, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001154.

Kalyanam:2019:MSE

[KZS^{+}19] Rajesh Kalyanam, Lan Zhao, Carol Song, Larry Biehl, Derrick Kearney, I. Luk Kim, Jaewoo Shin, Nelson Villoria, and Venkatesh Merwade. My-GeoHub — a sustainable and evolving geospatial science gateway. Future Generation Computer Systems, 94(??):820–832, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17310087.

Lajevardi:2019:SBC

[LA19] Amir Mohammadzade Lajevardi and Morteza Amini.

A semantic-based correlation approach for detecting hybrid and low-level APTs. Future Generation Computer Systems, 96(??): 64–88, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314924.

Liu:2010:SJC

[LAH10] Hongbo Liu, Ajith Abraham, and Aboul Ella Hassanien. Scheduling jobs on computational grids using a fuzzy particle swarm optimization algorithm. Future Generation Computer Systems, 26(8):1336–1343, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Li:2014:EES

[LAL^{+}14] Gongming Li, Hong An, Qi Li, Bobin Deng, and Wenbo Dai. Efficient execution of speculative threads and transactions with hardware transactional memory. Future Generation Computer Systems, 30(??):242–253, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001313.

Liang:2015:SEC

[LAL^{+}15] Kaitai Liang, Man Ho Au, Joseph K. Liu, Willy Susilo, Duncan S. Wong, Guomin Yang, Yong Yu, and Anjia Yang. A secure and efficient Ciphertext-Policy Attribute-Based Proxy Re-Encryption for cloud data sharing. Future Generation Computer Systems, 52(??):95–108, November 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002507.

Liaqat:2019:SAC

[LAQ^{+}19] Hannan Bin Liaqat, Amjad Ali, Junaid Qadir, Ali Kashif Bashir, Muhammad Bilal, and Fiaz Majeed. Socially-aware congestion control in ad-hoc networks: Current status and the way forward. Future Generation Computer Systems, 97(??):634–660, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18319083.

Lizcano:2016:WCE

[LASL16] David Lizcano, Fernando Alonso, Javier Soriano, and Genoveva López. Webcentred end-user component modelling. Future Generation Computer Systems, 54(??):16–40, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002253.

Losada:2019:LRR

[LBB^{+}19] Nuria Losada, George Bosilca, Aurélien Bouteiller, Patricia González, and María J. Martín. Local rollback for resilient MPI applications with applicationlevel checkpointing and message logging. Future Generation Computer Systems, 91(??):450–464, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18303443.

Labba:2018:PAE

[LBD18] Chahrazed Labba, Narjès Bellamine Ben Saoud, and Julie Dugdale. A predictive approach for the efficient distribution of agentbased systems on a hybridcloud. Future Generation Computer Systems, 86(??):750–764, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324329.

Leroux:2019:MFD

[LBD^{+}19] Sam Leroux, Steven Bohez, Elias De Coninck, Pieter Van Molle, Bert Vankeirsbilck, Tim Verbelen, Pieter Simoens, and Bart Dhoedt. Multi-fidelity deep neural networks for adaptive inference in the Internet of Multimedia Things. Future Generation Computer Systems, 97(??):355–360, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17324664.

Leng:2018:RAS

[LBJ^{+}18] Kaijun Leng, Ya Bi, Linbo Jing, Han-Chi Fu, and Inneke Van Nieuwenhuyse. Research on agricultural supply chain system with double chain architecture based on blockchain technology. Future Generation Computer Systems, 86(??):641–649, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18304527.

See retraction notice [LBJ^{+}24].

Leng:2024:RNR

[LBJ^{+}24] Kaijun Leng, Ya Bi, Linbo Jing, Han-Chi Fu, and Inneke Van Nieuwenhuyse.

Retraction notice to “Research on agricultural supply chain system with double chain architecture based on blockchain technology” [Future Generation Computer Systems 86 (2018) 641–649]. Future Generation Computer Systems, 151(??):272, February 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003758.

Lenkiewicz:2018:EED

[LBM18] Przemyslaw Lenkiewicz, P. Chris Broekema, and Bernard Metzler. Energyefficient data transfers in radio astronomy with software UDP RDMA. Future Generation Computer Systems, 79 (part 1)(??): 215–224, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17304715.

Lorenz:2010:SSJ

[LBU^{+}10] Daniel Lorenz, Peter Buchholz, Christian Uebing, Wolfgang Walkowiak, and Roland Wismüller. Steering of sequential jobs with a distributed shared memory based model for online steering. Future Generation Computer Systems, 26 (1):155–161, January 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liu:2013:SVV

[LC13] Denghui Liu and Jinli Cao. Scheduling paravirtualized virtual machines based on events. Future Generation Computer Systems, 29(5):1130–1139, July 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12002348.

Lin:2014:MWM

[LC14] Xuan-Yi Lin and Yeh-Ching Chung. Masterworker model for MapReduce paradigm on the TILE64 many-core platform. Future Generation Computer Systems, 36(??): 19–30, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001003.

Lin:2015:FRE

[LC15] Kawuu W. Lin and Sheng-Hao Chung. A fast and resource efficient mining algorithm for discovering frequent patterns in distributed computing environments. Future Generation Computer Systems, 52(??):49–58, November 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001843.

Laxmi:2017:GGS

[LC17] B. Prathusha Laxmi and A. Chilambuchelvan. GSR: Geographic Secured Routing using SHA-3 algorithm for node and message authentication in wireless sensor networks. Future Generation Computer Systems, 76(??):98–105, November 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1730986X.

Lee:2013:MFS

[LCBF13] Kyungyong Lee, Taewoong Choi, Patrick Oscar Boykin, and Renato J. Figueiredo. MatchTree: Flexible, scalable, and fault-tolerant wide-area resource discovery with distributed matchmaking and aggregation. Future Generation Computer Systems, 29(6):1596–1610, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001653.

Lara-Cabrera:2019:TSA

[LCC19] Raúl Lara-Cabrera and David Camacho. A taxonomy and state of the art revision on affective games. Future Generation Computer Systems, 92(??):516– 525, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17312153.

Loreti:2018:DAC

[LCCM18] Daniela Loreti, Federico Chesani, Anna Ciampolini, and Paola Mello. A distributed approach to compliance monitoring of business process event streams. Future Generation Computer Systems, 82(??):104– 118, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17317909.

Lopez-Casado:2019:OSS

[LCdPMCT19] Carmen López-Casado, Carlos Pérez del Pulgar, Víctor F. Muñoz, and Alberto J. Castro-Tirado. Observation scheduling and simulation in a global telescope network. Future Generation Computer Systems, 95(??):116–125, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317503.

Lara-Cabrera:2019:SAR

[LCGPC19] Raúl Lara-Cabrera, Antonio Gonzalez-Pardo, and David Camacho. Statistical analysis of risk assessment factors and metrics to evaluate radicalisation in Twitter. Future Generation Computer Systems, 93(??):971–978, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17308348.

Lai:2011:CCB

[LCH^{+}11] Chin-Feng Lai, Jui-Hung Chang, Chia-Cheng Hu, Yueh-Min Huang, and Han-Chieh Chao. CPRS: a cloud-based program recommendation system for digital TV platforms. Future Generation Computer Systems, 27(6):823–835, June 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liu:2018:MBS

[LCH^{+}18] Bo Liu, Xiaolin Chang, Zhen Han, Kishor Trivedi, and Ricardo J. Rodríguez. Model-based sensitivity analysis of IaaS cloud availability. Future Generation Computer Systems, 83(??): 1–13, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301796.

Li:2014:DPE

[LCHW14] Jin Li, Xiaofeng Chen, Qiong Huang, and Duncan S. Wong. Digital provenance: Enabling secure data forensics in cloud computing. Future Generation Computer Systems, 37(??):259–266, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002161.

Lin:2014:IQA

[LCL14] Jenn-Wei Lin, Chien-Hung Chen, and Chi-Yi Lin. Integrating QoS awareness with virtualization in cloud computing systems for delay-sensitive applications. Future Generation Computer Systems, 37(??): 478–487, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002987.

Li:2016:AEP

[LCL^{+}16] Dongsheng Li, Chao Chen, Qin Lv, Li Shang, Yingying Zhao, Tun Lu, and Ning Gu. An algorithm for efficient privacypreserving item-based collaborative filtering. Future Generation Computer Systems, 55(??):311–320, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002374.

Lin:2018:DFT

[LCL^{+}18] Szu-Yin Lin, Chi-Chun Chiang, Jung-Bin Li, Zih-Siang Hung, and Kuo-Ming Chao. Dynamic fine-tuning stacked autoencoder neural network for weather forecast. Future Generation Computer Systems, 89(??):446–454, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17329801.

Lee:2019:SER

[LCL^{+}19] Cheng-Chi Lee, Shun-Der Chen, Chun-Ta Li, Chung-Lun Cheng, and Yan-Ming Lai. Security enhancement on an RFID ownership transfer protocol based on cloud. Future Generation Computer Systems, 93(??):266–277, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18315711.

Liu:2016:EET

[LCMX16] Tundong Liu, Fufeng Chen, Yingran Ma, and Yi Xie. An energy-efficient task scheduling for mobile devices based on cloud assistant. Future Generation Computer Systems, 61(??): 1–12, August 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300176.

Li:2018:CFR

[LCW^{+}18] Wenjuan Li, Jian Cao, Jiyi Wu, Changqin Huang, and Rajkumar Buyya. A collaborative filtering recommendation method based on discrete quantuminspired shuffled frog leaping algorithms in social networks. Future Generation Computer Systems, 88(??):262–270, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18300232.

Lu:2019:TRB

[LCY19a] Hsi-Peng Lu, Chiao-Shan Chen, and Hueiju Yu.

Technology roadmap for building a smart city: an exploring study on methodology. Future Generation Computer Systems, 97(??):727–742, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18304011.

Luo:2019:STE

[LCY^{+}19b] Haidong Luo, Hongming Cai, Han Yu, Yan Sun, Zhuming Bi, and Lihong Jiang. A short-term energy prediction system based on edge computing for smart city. Future Generation Computer Systems, 101(??):444–457, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19305515.

Lin:2012:BEC

[LCZR12] Kai Lin, Min Chen, Sherali Zeadally, and Joel J. P. C. Rodrigues. Balancing energy consumption with mobile agents in wireless sensor networks. Future Generation Computer Systems, 28(2):446– 456, February 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1100029X.

Liang:2017:DSS

[LD17] Helan Liang and Yanhua Du. Dynamic service selection with QoS constraints and inter-service correlations using cooperative coevolution. Future Generation Computer Systems, 76(??):119–135, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16305325.

Liang:2019:CMO

[LDJL19] Helan Liang, Yanhua Du, Ting Jiang, and Fanzhang Li. A comprehensive multi-objective approach of service selection for service processes with twofold restrictions. Future Generation Computer Systems, 92(??):119–140, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18313700.

Liu:2018:CBT

[LDS^{+}18] Yezheng Liu, Fei Du, Jianshan Sun, Yuanchun Jiang, Jianmin He, Tingting Zhu, and Chunhua Sun. A crowdsourcingbased topic model for service matchmaking in Internet of Things. Future Generation Computer Systems, 87(??):186–197, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18304138.

Ling:2019:HNC

[LDX19] Chih Wei Ling, Anwitaman Datta, and Jun Xu. On hybrid network coding for visual traffic surveillance. Future Generation Computer Systems, 100(??):440–455, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18328450.

Liu:2018:HPP

[LDY^{+}18] Ximeng Liu, Robert H.

Deng, Yang Yang, Hieu N. Tran, and Shangping Zhong. Hybrid privacypreserving clinical decision support system in fogcloud computing. Future Generation Computer Systems, 78 (part 2)(??):825– 837, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17303874.

Liu:2019:TTR

[LDZW19] Zhenhua Liu, Shuhong Duan, Peilin Zhou, and Baocang Wang. Traceablethen-revocable ciphertextpolicy attribute-based encryption scheme. Future Generation Computer Systems, 93(??):903–913, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17320964.

Leal:2013:SAR

[Lea13] Katia Leal. Self-adjusting resource sharing policies in Federated Grids. Future Generation Computer Systems, 29(2):488–496, February 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001537.

Leal:2015:ARS

[Lea15] Katia Leal. Anticipating resource saturation in Federated Grids. Future Generation Computer Systems, 45(??):114–122, April 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002192.

Lee:2012:SST

[Lee12]  Gyungho Lee.  Special

section: Trusting software behavior. Future Generation Computer Systems, 28(8):1236–1237, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000379.

Lent:2016:ECC

[Len16] Ricardo Lent. Evaluating the cooling and computing energy demand of a datacentre with optimal server provisioning. Future Generation Computer Systems, 57(??):1– 12, April 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003210.

Liu:2019:DVP

[LEW19] Y. Liu, M. F. Ezerman, and H. Wang. Double verification protocol via secret sharing for lowcost RFID tags. Future Generation Computer Systems, 90(??):118–128, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17323518.

Liu:2015:UPE

[LFH^{+}15] Jinjun Liu, Dan Feng, Yu Hua, Bin Peng, and Zhenhua Nie. Using provenance to efficiently improve metadata searching performance in storage systems. Future Generation Computer Systems, 50(??):99– 110, September 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002246.

Li:2019:ELV

[LFHQ19] Chunguang Li, Dan Feng, Yu Hua, and Leihua Qin. Efficient live virtual machine migration for memory write-intensive workloads. Future Generation Computer Systems, 95(??): 126–139, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18320491.

Li:2017:EVI

[LFL^{+}17] Min Li, Cai Fu, Xiao-Yang Liu, Jia Yang, Tianqing Zhu, and Lansheng Han. Evolutionary virus immune strategy for temporal networks based on community vitality. Future Generation Computer Systems, 74(??):276–290, September 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301261.

Li:2017:CAR

[LFP^{+}17] Chung-Sheng Li, Hubertus Franke, Colin Parris, Bulent Abali, Mukil Kesavan, and Victor Chang. Composable architecture for rack scale big data computing. Future Generation Computer Systems, 67(??):180–193, February 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302631.

Lu:2019:AAF

[LFY^{+}19] Jiayi Lu, Luanye Feng, Jun Yang, Mohammad Mehedi Hassan, Abdulhameed Alelaiwi, and Iztok Humar. Artificial agent: the fusion of artificial intelligence and a mobile agent for energy-efficient traffic control in wireless sensor networks. Future Generation Computer Systems, 95(??): 45–51, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18327912.

Li:2016:WHQ

[LG16a] Yang Li and Yike Guo.

Wiki-Health: From quantified self to self-understanding. Future Generation Computer Systems, 56(??):333– 359, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002617.

Liu:2016:EES

[LG16b] Jun Liu and Jinhua Guo. Energy efficient scheduling of real-time tasks on multi-core processors with voltage islands. Future Generation Computer Systems, 56(??):202–210, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002046.

Lujak:2018:CME

[LG18] Marin Lujak and Stefano Giordani. Centrality measures for evacuation: Finding agile evacuation routes. Future Generation Computer Systems, 83(??):401– 412, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17309858.

Li:2017:ECM

[LGL^{+}17] Zhongjin Li, Jidong Ge, Chuanyi Li, Hongji Yang, Haiyang Hu, Bin Luo, and Victor Chang. Energy cost minimization with job security guarantee in Internet data center. Future Generation Computer Systems, 73(??):63–78, August 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16307634.

Li:2019:FCF

[LGP^{+}19] Pan Li, Songtao Guo, Chengsheng Pan, Li Yang, Guiyan Liu, and Yue Zeng. Fast congestion-free consistent flow forwarding rules update in software defined networking. Future Generation Computer Systems, 97(??):743–754, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317850.

Li:2017:RID

[LGW^{+}17] Jianjiang Li, Qian Ge, Jie Wu, Yue Li, Xiaolei Yang, and Zhanning Ma. Research and implementation of a distributed transaction processing middleware. Future Generation Computer Systems, 74(??):232–240, September 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630005X.

Li:2016:SCA

[LGY^{+}16] Zhongjin Li, Jidong Ge, Hongji Yang, Liguo Huang, Haiyang Hu, Hao Hu, and Bin Luo. A security and cost aware scheduling algorithm for heterogeneous tasks of scientific workflow in clouds. Future Generation Computer Systems, 65(??):140–152, December 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003982.

Liu:2018:TSB

[LGZY18] Guiyan Liu, Songtao Guo, Quanjun Zhao, and Yuanyuan Yang. Tomogravity space based traffic matrix estimation in data center networks. Future Generation Computer Systems, 86(??):39–50, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16308469.

Laure:2013:P

[LH13a] Erwin Laure and Sverker Holmgren. Preface. Future Generation Computer Systems, 29(8):2115–2116, October 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300112X.

Liao:2013:NMS

[LH13b] Yi-Pin Liao and Chih-Ming Hsiao. A novel multi-server remote user authentication scheme using self-certified public keys for mobile clients. Future Generation Computer Systems, 29(3): 886–900, March 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000829.

Lounis:2016:HCS

[LHBC16] Ahmed Lounis, Abdelkrim Hadjidj, Abdelmadjid Bouabdallah, and Yacine Challal. Healing on the cloud: Secure cloud architecture for medical wireless sensor networks. Future Generation Computer Systems, 55(??):266–277, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000266.

Lv:2018:SSU

[LHCC18] Xiao Lv, Fazhi He, Weiwei Cai, and Yuan Cheng. Supporting selective undo of string-wise operations for collaborative editing systems. Future Generation Computer Systems, 82(??): 41–62, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1730897X.

Lee:2018:CLM

[LHJC18] Jiwan Lee, Bonghee Hong, Sunghoon Jung, and Victor Chang. Clustering learning model of CCTV image pattern for producing road hazard meteorological information. Future Generation Computer Systems, 86(??):1338–1350, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17318253.

Liu:2015:SSP

[LHL15] Jianghua Liu, Xinyi Huang, and Joseph K. Liu. Secure sharing of Personal Health Records in cloud computing: Ciphertext-Policy Attribute-Based Signcryption. Future Generation Computer Systems, 52(??):67–76, November 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002076.

Lian:2014:SSA

[LHM14] Yanling Lian, Xinyi Huang, and Yi Mu. SA^{3}: Selfadaptive anonymous authentication for dynamic authentication policies. Future Generation Computer Systems, 30(??):133–139, January 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001210.

Li:2017:ECA

[LHO17] Fagen Li, Jiaojiao Hong, and Anyembe Andrew Omala. Efficient certificateless access control for industrial Internet of Things. Future Generation Computer Systems, 76(??):285–292, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308664.

Lopez-Huguet:2019:SMM

[LHPC^{+}19] Sergio López-Huguet, Alfonso Pérez, Amanda Calatrava, Carlos de Alfonso, Miguel Caballer, Germán Moltó, and Ignacio Blanquer. A self-managed Mesos cluster for data analytics with QoS guarantees. Future Generation Computer Systems, 96(??): 449–461, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18311087.

Liu:2018:NPB

[LHW^{+}18] Zelei Liu, Liang Hu, Chunyi Wu, Yan Ding, Quangang Wen, and Jia Zhao. A novel processbased association rule approach through maximal frequent itemsets for big data processing. Future Generation Computer Systems, 81(??):414–424, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17307859.

Lin:2018:DNR

[LHX^{+}18] Hui Lin, Jia Hu, Chuanfeng Xu, Jianfeng Ma, and Mengyang Yu. DTRM: A new reputation mechanism to enhance data trustworthiness for highperformance cloud computing. Future Generation Computer Systems, 83(??):293–302, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17317247.

Lv:2019:ISU

[LHY^{+}19] Xiao Lv, Fazhi He, Xiaohu Yan, Yiqi Wu, and Yuan Cheng. Integrating selective undo of featurebased modeling operations for real-time collaborative CAD systems. Future Generation Computer Systems, 100(??):473–497, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18323033.

Li:2010:GRD

[Li10] Juan Li. Grid resource discovery based on semantically linked virtual organizations. Future Generation Computer Systems, 26 (3):361–373, March 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Li:2015:AFD

[Li15]  Keqin Li.  Analysis

of file download time in peer-to-peer networks with stochastic and timevarying service capacities. Future Generation Computer Systems, 42(??):36– 43, January 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001666.

Li:2018:SPT

[Li18] Keqin Li. Scheduling parallel tasks with energy and time constraints on multiple manycore processors in a cloud computing environment. Future Generation Computer Systems, 82(??):591–605, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17300298.

Lv:2018:BDA

[LIC18] Zhihan Lv, Rahat Iqbal, and Victor Chang. Big data analytics for sustainability. Future Generation Computer Systems, 86(??):1238–1241, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18311269.

Liu:2019:TAS

[LIH^{+}19] Xiufeng Liu, Nadeem Iftikhar, Huan Huo, Rongling Li, and Per Sieverts Nielsen. Two approaches for synthesizing scalable residential energy consumption data. Future Generation Computer Systems, 95(??):586–600, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18316212.

Lin:2018:MFS

[Lin18] Zhe Lin. Modelling and forecasting the stock market volatility of SSE Composite Index using GARCH models. Future Generation Computer Systems, 79 (part 3)(??):960–972, February 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17313067.

Lee:2017:SCB

[LJ17a] O-Joun Lee and Jai E.

Jung. Sequence clusteringbased automated rule generation for adaptive complex event processing. Future Generation Computer Systems, 66(??):100–109, January 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300243.

Lei:2017:SCB

[LJ17b] Yu Lei and Zhang Junxing. Service composition based on multi-agent in the cooperative game. Future Generation Computer Systems, 68(??):128–135, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302783.

Lee:2019:MAC

[LJ19a] O-Joun Lee and Jason J. Jung. Modeling affective character network for story analytics. Future Generation Computer Systems, 92(??):458–478, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17310221.

Limkar:2019:NMP

[LJ19b] Suresh V. Limkar and Rakesh Kumar Jha. A novel method for parallel indexing of real time geospatial big data generated by IoT devices. Future Generation Computer Systems, 97(??):433–452, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18303637.

Lin:2019:ACD

[LJC^{+}19] Yezhi Lin, Xinyuan Jin, Jiuqiang Chen, Ali Hassan Sodhro, and Zhifang Pan. An analytic computationdriven algorithm for decentralized multicore systems. Future Generation Computer Systems, 96(??): 101–110, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18321010.

Li:2018:HEF

[LJGW18] Xiang Li, Xiaohong Jiang, Peter Garraghan, and Zhaohui Wu. Holistic energy and failure aware workload scheduling in cloud datacenters. Future Generation Computer Systems, 78 (part 3)(??):887– 900, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17315650.

Li:2018:OCD

[LJJ18] Weimin Li, Shu Jiang, and Qun Jin. Overlap community detection using spectral algorithm based on node convergence degree. Future Generation Computer Systems, 79 (part 1)(??):408– 416, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17318411.

Liao:2012:TGC

[LJL12] Xiaofei Liao, Hai Jin, and Haikun Liu. Towards a green cluster through dynamic remapping of virtual machines. Future Generation Computer Systems, 28(2):469– 477, February 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11000641.

Li:2013:CSS

[LJLW13] Jianxin Li, Yu Jia, Lu Liu, and Tianyu Wo. Cyber-LiveApp: a secure sharing and migration approach for live virtual desktop applications in a cloud environment. Future Generation Computer Systems, 29(1):330–340, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001403.

Lee:2017:DIS

[LJS17] YangSun Lee, Junho Jeong, and Yunsik Son. Design and implementation of the secure compiler and virtual machine for developing secure IoT services. Future Generation Computer Systems, 76(??):350– 357, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300589.

Liu:2019:ADC

[LJW^{+}19a] Dong-Wei Liu, Run-Ping Jia, Cai-Feng Wang, N. Arunkumar, K. Narasimhan, M. Udayakumar, and V. Elamaran. Automated detection of cancerous genomic sequences using genomic signal processing and machine learning. Future Generation Computer Systems, 98(??):233–237, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326645.

Luo:2019:UBP

[LJW^{+}19b] Xiong Luo, Changwei Jiang, Weiping Wang, Yang Xu, Jenq-Haur Wang, and Wenbing Zhao. User behavior prediction in social networks using weighted extreme learning machine with distribution optimization. Future Generation Computer Systems, 93(??):1023–1035, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17307938.

Liao:2010:EOR

[LJY10] Xiaofei Liao, Hai Jin, and Xiaojie Yuan. ESPM: An optimized resource distribution policy in virtual user environment. Future Generation Computer Systems, 26(8):1393–1402, October 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liao:2012:NDR

[LJY12] Xiaofei Liao, Hai Jin, and Linchen Yu. A novel data replication mechanism in P2P VoD system. Future Generation Computer Systems, 28(6):930– 939, June 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001919.

Li:2012:BIB

[LK12] Fagen Li and Muhammad Khurram Khan. A biometric identity-based signcryption scheme. Future Generation Computer Systems, 28(1):306–310, January 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X10002141.

Louge:2019:SWS

[LKA^{+}19] Thierry Louge, Mohamed Hedi Karray, Bernard Archimède, Zakaria Maamar, and Michael Mrissa. Semantic Web Services Composition in the astrophysics domain: Issues and solutions. Future Generation Computer Systems, 90(??):185– 197, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18301869.

Lee:2018:GTB

[LKCS18] Seokcheol Lee, Sungjin Kim, Ken Choi, and Taeshik Shon. Game theory-based security vulnerability quantification for Social Internet of Things. Future Generation Computer Systems, 82(??):752–760, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17308440.

Liu:2018:TSD

[LKFB18] Zhengchun Liu, Rajkumar Kettimuthu, Ian Foster, and Peter H. Beckman. Toward a smart data transfer node. Future Generation Computer Systems, 89(??):10–18, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18302346.

Loukil:2017:ABR

[LKJ17] Sihem Loukil, Slim Kallel, and Mohamed Jmaiel. An approach based on runtime models for developing dynamically adaptive systems. Future Generation Computer Systems, 68(??): 365–375, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302369.

Lee:2019:FDN

[LKJ^{+}19] Keuntae Lee, Seokhwa Kim, Jaehoon (Paul) Jeong, Sejun Lee, Hyoungshick Kim, and Jung-Soo Park. A framework for DNS naming services for Internet-of-Things devices. Future Generation Computer Systems, 92(??):617– 627, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17317429.

Lohrer:2016:GFS

[LKK^{+}16] Johannes-Y. Lohrer, Daniel Kaltenthaler, Peer Kröger, Christiaan van der Meijden, and Henriette Obermaier. A generic framework for synchronized distributed data management in archaeological related disciplines. Future Generation Computer Systems, 56(??):558–570, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002241.

Lee:2014:LSI

[LKM14] Daewoo Lee, Jin-Soo Kim, and Seungryoul Maeng. Large-scale incremental processing with MapReduce. Future Generation Computer Systems, 36(??): 66–79, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001891.

Li:2013:OCA

[LKN^{+}13] Yuan-Fang Li, Gavin Kennedy, Faith Ngoran, Philip Wu, and Jane Hunter. An ontologycentric architecture for extensible scientific data management systems. Future Generation Computer Systems, 29(2):641–653, February 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1100118X.

Lugones:2014:RRT

[LKTC14] Diego Lugones, Kostas Katrinis, Georgios Theodoropoulos, and Martin Collier. A reconfigurable, regular-topology cluster/ datacenter network using commodity optical switches. Future Generation Computer Systems, 30(??):78–89, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000769.

Lu:2016:PFC

[LL16] Yang Lu and Jiguo Li. A pairing-free certificatebased proxy re-encryption scheme for secure data sharing in public clouds. Future Generation Computer Systems, 62(??):140– 147, September 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1500360X.

Li:2018:SGI

[LL18] Jing Jian Li and Bo Ling. Symmetric graphs and interconnection networks. Future Generation Computer Systems, 83(??):461– 467, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17309871.

Liu:2013:HMS

[LLAH13] Yang Liu, Maozhen Li, Nasullah Khalid Alham, and Suhel Hammoud. HSim: a MapReduce simulator in enabling Cloud Computing. Future Generation Computer Systems, 29(1):300–308, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000884.

Luo:2017:PPM

[LLAW17] Entao Luo, Qin Liu, Jemal H. Abawajy, and Guojun Wang. Privacypreserving multi-hop profilematching protocol for proximity mobile social networks. Future Generation Computer Systems, 68(??): 222–233, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303107.

Lee:2011:IJS

[LLC11] Yun-Han Lee, Seiven Leu, and Ruay-Shiung Chang. Improving job scheduling algorithms in a grid environment. Future Generation Computer Systems, 27 (8):991–998, October 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lee:2014:PAC

[LLC14a] Cheng-Yu Lee, Tzong-Yen Lin, and Rong-Guey Chang. Power-aware code scheduling assisted with power gating and DVS. Future Generation Computer Systems, 34(??):66– 75, May 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002756.

Li:2014:PPD

[LLC^{+}14b] Jingwei Li, Jin Li, Xiaofeng Chen, Zheli Liu, and Chunfu Jia. Privacypreserving data utilization in hybrid clouds. Future Generation Computer Systems, 30(??):98–106, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001258.

Li:2016:EMB

[LLC^{+}16] Ying Li, Weiwei Liu, Bin Cao, Jianwei Yin, and Min Yao. An efficient MapReduce-based rule matching method for production system. Future Generation Computer Systems, 54(??):478–489, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000655.

Lim:2011:SRQ

[LLCF11] Chunhyeok Lim, Shiyong Lu, Artem Chebotko, and Farshad Fotouhi. Storing, reasoning, and querying OPM-compliant scientific workflow provenance using relational databases. Future Generation Computer Systems, 27(6):781– 789, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lewis:2019:TTS

[LLES19] Grace Lewis, Patricia Lago, Sebastián Echeverría, and Pieter Simoens. A tale of three systems: Case studies on the application of architectural tactics for cyber-foraging. Future Generation Computer Systems, 96(??):119– 147, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18312834.

Li:2018:LCS

[LLF^{+}18a] Huixi Li, Wenjun Li, Qilong Feng, Shigeng Zhang, Haodong Wang, and Jianxin Wang. Leveraging content similarity among VMI files to allocate virtual machines in cloud. Future Generation Computer Systems, 79 (part 2)(??):528– 542, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17300146.

Li:2018:NOV

[LLF^{+}18b] Xiaohong Li, Feng Liu, Zhiyong Feng, Guangquan Xu, and Zhangjie Fu. A novel optimized vertical handover framework for seamless networking integration in cyber-enabled systems. Future Generation Computer Systems, 79 (part 1)(??):417– 430, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17305216.

Liu:2016:DCS

[LLG^{+}16] Weidong Liu, Xiangfeng Luo, Zhiguo Gong, Junyu Xuan, Ngai Meng Kou, and Zheng Xu. Discovering the core semantics of event from social media.

Future Generation Computer Systems, 64(??):175– 185, November 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003805.

Liu:2018:PBR

[LLGY18] Rui Liu, Junbin Liang, Wenyu Gao, and Ruiyun Yu. Privacy-based recommendation mechanism in mobile participatory sensing systems using crowdsourced users’ preferences. Future Generation Computer Systems, 80(??):76– 88, March 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306070.

Li:2017:MKP

[LLH^{+}17] Ping Li, Jin Li, Zhengan Huang, Tong Li, Chong-Zhi Gao, Siu-Ming Yiu, and Kai Chen. Multi-key privacy-preserving deep learning in cloud computing. Future Generation Computer Systems, 74(??):76–85, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302005.

Liu:2011:NRC

[LLJ^{+}11] Yining Liu, Keqiu Li, Yingwei Jin, Yong Zhang, and Wenyu Qu. A novel reputation computation model based on subjective logic for mobile ad hoc networks. Future Generation Computer Systems, 27(5):547– 554, May 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liu:2018:VSE

[LLL^{+}18] Zheli Liu, Tong Li, Ping Li, Chunfu Jia, and Jin Li. Verifiable searchable encryption with aggregate keys for data sharing system. Future Generation Computer Systems, 78 (part 2)(??):778–788, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302340.

Liang:2019:EDP

[LLL^{+}19] Wei Liang, Jing Long, Xia Lei, Zhiqiang You, Haibo Luo, Jiahong Cai, and Kuan-Ching Li. Efficient data packet transmission algorithm for IPV6 mobile vehicle network based on fast switching model with time difference. Future Generation Computer Systems, 100(??):132–143, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304716.

Liu:2016:EPP

[LLM^{+}16] Ximeng Liu, Rongxing Lu, Jianfeng Ma, Le Chen, and Haiyong Bao. Efficient and privacy-preserving skyline computation framework across domains. Future Generation Computer Systems, 62(??):161–174, September 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003180.

Lara:2013:DFE

[LLMP13] Juan Alfonso Lara, David Lizcano, María Aurora Martínez, and Juan Pazos. Developing front-end Web 2.0 technologies to access services, content and things in the future Internet. Future Generation Computer Systems, 29(5): 1184–1195, July 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000204.

Li:2018:ARD

[LLN^{+}18] Yujie Li, Huimin Lu, Yoshiki Nakayama, Hyoungseop Kim, and Seiichi Serikawa. Automatic road detection system for an airland amphibious car drone. Future Generation Computer Systems, 85(??):51– 59, August 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1732770X.

Lee:2012:PAD

[LLpC12] Ming-Chang Lee, Fang-Yie Leu, and Ying ping Chen. PFRF: an adaptive data replication algorithm based on startopology data grids. Future Generation Computer Systems, 28(7):1045–1057, July 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002147.

Li:2014:SEE

[LLQS14] Hongjuan Li, Keqiu Li, Wenyu Qu, and Ivan Stojmenovic. Secure and energy-efficient data aggregation with malicious aggregator identification in wireless sensor networks. Future Generation Computer Systems, 37(??):108– 116, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002859.

Liu:2014:ECA

[LLS^{+}14] Jin Liu, Juan Li, Xiaoping Sun, Yuan Xie, Jeff Lei, and Qiping Hu. An Embedded Co-AdaBoost based construction of software document relation coupled resource spaces for cyber-physical society. Future Generation Computer Systems, 32(??):198–210, March 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12002373.

Liang:2019:EEC

[LLS^{+}19] Hai-Ning Liang, Feiyu Lu, Yuwei Shi, Vijayakumar Nanjappan, and Konstantinos Papangelis. Evaluating the effects of collaboration and competition in navigation tasks and spatial knowledge acquisition within virtual reality environments. Future Generation Computer Systems, 95(??):855–866, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17308324.

Liu:2018:NFR

[LLSL18] Xingang Liu, Lingyun Lu, Zhixin Shen, and Kaixuan Lu. A novel face recognition algorithm via weighted kernel sparse representation. Future Generation Computer Systems, 80(??): 653–663, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302370.

Li:2019:GAD

[LLT^{+}19] Junnan Li, Zhihui Lu, Yu Tong, Jie Wu, Shalin Huang, Meikang Qiu, and Wei Du. A general AIdefined attention network for predicting CDN performance. Future Generation Computer Systems, 100(??):759–769, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19307356.

Lu:2018:LIU

[LLU^{+}18] Huimin Lu, Yujie Li, Tomoki Uemura, Hyoungseop Kim, and Seiichi Serikawa. Low illumination underwater light field images reconstruction using deep convolutional neural networks. Future Generation Computer Systems, 82(??):142–148, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17326845.

Li:2012:CVS

[LLW^{+}12a] Jianxin Li, Bo Li, Tianyu Wo, Chunming Hu, Jinpeng Huai, Lu Liu, and K. P. Lam. CyberGuarder: a virtualization security assurance architecture for green cloud computing. Future Generation Computer Systems, 28(2):379– 390, February 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1100063X.

Liu:2012:SSD

[LLW^{+}12b] Tiejiang Liu, Tun Lu, Wei Wang, Qi Wang, Zhenyu Liu, Ning Gu, and Xianghua Ding. SDMS-O: a service deployment management system for optimization in clouds while guaranteeing users’ QoS requirements. Future Generation Computer Systems, 28(7):1100–1109, July 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002226.

Li:2018:HDM

[LLW^{+}18a] Rui Li, Zhaohui Liu, Huayi Wu, Ru Li, Guangsheng Dong, and Jie Jiang. Hierarchical decomposition method and combination forecasting scheme for access load on public map service platforms. Future Generation Computer Systems, 87(??):213–227, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17315923.

Lv:2018:GAS

[LLW^{+}18b] Zhihan Lv, Xiaoming Li, Weixi Wang, Baoyun Zhang, Jinxing Hu, and Shengzhong Feng. Government affairs service platform for smart city. Future Generation Computer Systems, 81(??):443–451, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17311391.

Li:2019:EBF

[LLW^{+}19a] Zhidan Li, Wenmin Li, QiaoYan Wen, Jiageng Chen, Wei Yin, and Kaitai Liang. An efficient blind filter: Location privacy protection and the access control in FinTech. Future Generation Computer Systems, 100(??):797–810, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1833173X.

Liang:2019:TTB

[LLW^{+}19b] Wei Liang, Jing Long, Tien-Hsiung Weng, Xuhui Chen, Kuan-Ching Li, and Albert Y. Zomaya. TBRS: A trust based recommendation scheme for vehicular CPS network. Future Generation Computer Systems, 92(??):383–398, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18305545.

Liao:2019:SDC

[LLW^{+}19c] Lingxia Liao, Chin-Feng Lai, Jaifu Wan, Victor C. M. Leung, and Tien-Chi Huang. Scalable distributed control plane for on-line social networks support cognitive neural computing in software defined networks. Future Generation Computer Systems, 93(??):993–1001, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17307963.

Liu:2019:IIM

[LLW^{+}19d] Yuxin Liu, Anfeng Liu, Tian Wang, Xiao Liu, and Neal N. Xiong. An intelligent incentive mechanism for coverage of data collection in cognitive Internet of Things. Future Generation Computer Systems, 100(??):701–714, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18309208.

Li:2018:OVM

[LLWW18] Huixi Li, Wenjun Li, Haodong Wang, and Jianxin Wang. An optimization of virtual machine selection and placement by using memory content similarity for server consolidation in cloud. Future Generation Computer Systems, 84(??):98–107, July 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17323063.

Liu:2018:ASS

[LLWZ18] Xing Liu, Jiqiang Liu, Wei Wang, and Sencun Zhu. Android single sign-on security: Issues, taxonomy and directions. Future Generation Computer Systems, 89(??):402–420, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18302358.

Li:2018:PPM

[LLY^{+}18] Ping Li, Tong Li, Heng Ye, Jin Li, Xiaofeng Chen, and Yang Xiang. Privacypreserving machine learning with multiple data providers. Future Generation Computer Systems, 87(??):341–350, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17327036.

Lai:2019:EDR

[LLYW19] Yongxuan Lai, Hailin Lin, Fan Yang, and Tian Wang. Efficient data request answering in vehicular adhoc networks based on fog nodes and filters. Future Generation Computer Systems, 93(??):130–142, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18316194.

Li:2018:SSE

[LLZ^{+}18a] Junnan Li, Zhihui Lu, Wei Zhang, Jie Wu, Hao Qiang nd Bo Li, and Patrick C. K.

Hung. SERAC3: Smart and economical resource allocation for big data clusters in community clouds. Future Generation Computer Systems, 85(??):210– 221, August 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1830267X.

Liu:2018:SIT

[LLZ^{+}18b] Jian Liu, Jiangtao Li, Lei Zhang, Feifei Dai, Yuanfei Zhang, Xinyu Meng, and Jian Shen. Secure intelligent traffic light control using fog computing. Future Generation Computer Systems, 78 (part 2)(??):817– 824, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302157.

Li:2019:PSB

[LLZ^{+}19] Huixi Li, Wenjun Li, Shigeng Zhang, Haodong Wang, Yi Pan, and Jianxin Wang. Page-sharing-based virtual machine packing with multi-resource constraints to reduce network traffic in migration for clouds. Future Generation Computer Systems, 96(??): 462–471, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18302619.

Luo:2012:HHO

[LM12] Yi Luo and D. Manivannan. HOPE: a Hybrid Optimistic checkpointing and selective Pessimistic mEssage logging protocol for large scale distributed systems. Future Generation Computer Systems, 28(8):1217–1235, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000659.

Lombardi:2019:PAP

[LMA^{+}19] Federico Lombardi, Andrea Muti, Leonardo Aniello, Roberto Baldoni, Silvia Bonomi, and Leonardo Querzoni. PASCAL: an architecture for proactive auto-scaling of distributed services. Future Generation Computer Systems, 98(??):342–361, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18303728.

Lopez-Martin:2019:NNA

[LMCSE19] Manuel Lopez-Martin, Belen Carro, and Antonio Sanchez-Esguevillas. Neural network architecture based on gradient boosting for IoT traffic prediction. Future Generation Computer Systems, 100(??):656–673, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310064.

Liu:2019:ECE

[LML^{+}19] Xiaolong Liu, Khan Muhammad, Jaime Lloret, Yu-Wen Chen, and Shyan-Ming Yuan. Elastic and cost-effective data carrier architecture for smart contract in blockchain. Future Generation Computer Systems, 100(??):590–599, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18328334.

Liu:2019:DMM

[LMM19] Liang Liu, Zuchao Ma, and Weizhi Meng. Detection of multiple-mix-attack malicious nodes using perceptron-based trust in IoT networks. Future Generation Computer Systems, 101(??):865–879, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301013.

Li:2014:TUR

[LMZ^{+}14] Zheng Li, Karan Mitra, Miranda Zhang, Rajiv Ranjan, Dimitrios Georgakopoulos, Albert Y. Zomaya, Liam O’Brien, and Shengtao Sun. Towards understanding the runtime configuration management of do-it-yourself content delivery network applications over public clouds. Future Generation Computer Systems, 37(??): 297–308, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002835.

Leon:2013:SGD

[LN13] Xavier León and Leandro Navarro. A Stackelberg game to derive the limits of energy savings for the allocation of data center resources. Future Generation Computer Systems, 29(1): 74–83, January 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001306.

Lieber:2018:HSS

[LN18] Matthias Lieber and Wolfgang E. Nagel. Highly scalable SFC-based dynamic load balancing and its application to atmospheric modeling. Future Generation Computer Systems, 82(??):575–590, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17308178.

Li:2014:SIB

[LNB14] Gang Li, Wenjia Niu, and Lynn Batten. Special issue on behavior data security issues in network information propagation. Future Generation Computer Systems, 36(??):120–121, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000302.

Li:2018:RBB

[LNK^{+}18] Xiong Li, Jianwei Niu, Saru Kumari, Fan Wu, and Kim-Kwang Raymond Choo. A robust biometrics based three-factor authentication scheme for Global Mobility Networks in smart city. Future Generation Computer Systems, 83(??):607–618, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1730585X.

Lin:2019:SEL

[LNLA19] Jiaping Lin, Jianwei Niu, Hui Li, and Mohammed Atiquzzaman. A secure and efficient locationbased service scheme for smart transportation. Future Generation Computer Systems, 92(??):694–704, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17317235.

Liu:2019:RSW

[LNM^{+}19] Hong Liu, Huansheng Ning, Qitao Mu, Yumei Zheng, Jing Zeng, Laurence T. Yang, Runhe Huang, and Jianhua Ma. A review of the smart world. Future Generation Computer Systems, 96(??):678– 691, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17319532.

Liu:2018:SDY

[LNY^{+}18] Hong Liu, Huansheng Ning, Yinliang Yue, Yueliang Wan, and Laurence T. Yang. Selective disclosure and yoking-proof based privacy-preserving authentication scheme for cloud assisted wearable devices. Future Generation Computer Systems, 78 (part 3)(??):976–986, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17305873.

Loia:2019:UCE

[LO19] Vincenzo Loia and Francesco Orciuoli. Understanding the composition and evolution of terrorist group networks: a rough set approach. Future Generation Computer Systems, 101(??):983–992, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19307757.

Loke:2012:SUS

[Lok12]  Seng W. Loke.  Sup-

porting ubiquitous sensorcloudlets and contextcloudlets: Programming compositions of contextaware systems for mobile users. Future Generation Computer Systems, 28(4): 619–632, April 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001701.

Li:2018:EEE

[LOR^{+}18] Yunbo Li, Anne-Cécile Orgerie, Ivan Rodero, Betsegaw Lemma Amersho, Manish Parashar, and Jean-Marc Menaud. Endto-end energy models for edge cloud-based IoT platforms: Application to data stream analysis in IoT. Future Generation Computer Systems, 87(??):667–678, October 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17314309.

Lopez-Pires:2018:VMP

[LPBB^{+}18] Fabio López-Pires, Benjamín Barán, Leonardo Benítez, Saúl Zalimben, and Augusto Amarilla. Virtual machine placement for elastic infrastructures in overbooked cloud computing datacenters under uncertainty. Future Generation Computer Systems, 79 (part 3)(??):830–848, February 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17303126.

Lloyd:2013:PIM

[LPD^{+}13] W. Lloyd, S. Pallickara, O. David, J. Lyon, M. Arabi, and K. Rojas. Performance implications of multi-tier application deployments on Infrastructure-as-a-Service clouds: Towards performance modeling. Future Generation Computer Systems, 29(5):1254–1264, July 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12002270.

Leong:2017:RRE

[LPK17] Siew Hoon Leong, Antonio Parodi, and Dieter Kranzlmüller. A robust reliable energy-aware urgent computing resource allocation for flash-flood ensemble forecasting on HPC infrastructures for decision support. Future Generation Computer Systems, 68(??): 136–149, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303119.

See reprint [LPK18].

Leong:2018:RRR

[LPK18] Siew Hoon Leong, Antonio Parodi, and Dieter Kranzlmüller. Reprint of A robust reliable energyaware urgent computing resource allocation for flashflood ensemble forecasting on HPC infrastructures for decision support. Future Generation Computer Systems, 79 (part 1)(??):114– 127, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X16306355.

See [LPK17].

Liu:2016:MDT

[LPL^{+}16] Kaiyang Liu, Jun Peng, Heng Li, Xiaoyong Zhang, and Weirong Liu. Multidevice task offloading with time-constraints for energy efficiency in mobile cloud computing. Future Generation Computer Systems, 64(??):1–14, November 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300905.

Laleh:2018:CVF

[LPMY18] Touraj Laleh, Joey Paquet, Serguei Mokhov, and Yuhong Yan. Constraint verification failure recovery in web service composition. Future Generation Computer Systems, 89(??):387– 401, December 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17320629.

Lintermann:2019:POM

[LPS19] Andreas Lintermann, Dirk Pleiter, and Wolfgang Schröder. Performance of ODROID-MC1 for scientific flow problems. Future Generation Computer Systems, 95(??):149–162, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18322015.

Liu:2016:MOS

[LPV^{+}16] Ji Liu, Esther Pacitti, Patrick Valduriez, Daniel de Oliveira, and Marta Mattoso. Multi-objective scheduling of scientific workflows in multisite clouds. Future Generation Computer Systems, 63(??): 76–95, October 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300917.

Li:2018:EFI

[LPY^{+}18] Jianxin Li, Hao Peng, Erica Yang, Chunming Hu, Shenghai Zhong, and Lihong Wang. Eagle+: a fast incremental approach to automaton and table online updates for cloud services. Future Generation Computer Systems, 80(??): 275–285, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301917.

Lu:2019:FIA

[LQF19] Yu Lu, Yingjian Qi, and Xianghua Fu. A framework for intelligent analysis of digital cardiotocographic signals from IoMT-based foetal monitoring. Future Generation Computer Systems, 101(??):1130–1141, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19311434.

Lee:2016:AEE

[LQK^{+}16] Young-Sik Lee, Luis Cavazos Quero, Sang-Hoon Kim, Jin-Soo Kim, and Seungryoul Maeng. ActiveSort: Efficient external sorting using active SSDs in the MapReduce framework. Future Generation Computer Systems, 65(??):76–89, December 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300474.

Liu:2010:HCF

[LQLX10] Zhaobin Liu, Wenyu Qu, Haitao Li, and Changsheng Xie. A hybrid collaborative filtering recommendation mechanism for P2P networks. Future Generation Computer Systems, 26(8):1409–1417, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lopez:2017:EPS

[LRBW17] Javier Lopez, Ruben Rios, Feng Bao, and Guilin Wang. Evolving privacy: From sensors to the Internet of Things. Future Generation Computer Systems, 75(??):46–57, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16306719.

Liu:2018:PMD

[LRC^{+}18] Lei Liu, Yongjian Ren, Lizhen Cui, Yuliang Shi, and Qingzhong Li. Performance measurement of data flow processing employing software defined architecture. Future Generation Computer Systems, 82(??):235–243, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17317181.

Lago:2019:LMC

[LRJG19] Paula Lago, Claudia Roncancio, and Claudia Jiménez-Guarín. Learning and managing context enriched behavior patterns in smart homes. Future Generation Computer Systems, 91(??):191–205, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18307180.

Lei:2014:TBS

[LRL^{+}14] Jingsheng Lei, Yanghui Rao, Qing Li, Xiaojun Quan, and Liu Wenyin. Towards building a social emotion detection system for online news. Future Generation Computer Systems, 37(??):438–448, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002033.

Lino:2019:ACB

[LRMS19] Adriano Lino, Álvaro Rocha, Luís Macedo, and Amanda Sizo. Application of clustering-based decision tree approach in SQL query error database. Future Generation Computer Systems, 93(??):392–406, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18306319.

Li:2017:ANE

[LRYJ17] Tan Li, Yufei Ren, Dantong Yu, and Shudong Jin. Analysis of NUMA effects in modern multicore systems for the design of high-performance data transfer applications. Future Generation Computer Systems, 74(??):41– 50, September 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16305799.

Li:2018:RRT

[LRZ^{+}18] Yangfan Li, Wei Ren, Tianqing Zhu, Yi Ren, Yue Qin, and Wei Jie. RIMS: A real-time and intelligent monitoring system for live-broadcasting platforms. Future Generation Computer Systems, 87(??):259–266, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18301754.

Li:2010:SSP

[LS10] Keqiu Li and Ivan Stojmenovic. Special section: P2P and Internet computing. Future Generation Computer Systems, 26(8): 1391–1392, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lloyd:2013:ESG

[LSAM13] Ashley D. Lloyd, Terence M. Sloan, Mario Antonioletti, and G. A. McGilvary. Embedded systems for global e-social science: Moving computation rather than data. Future Generation Computer Systems, 29(5):1120–1129, July 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12002336.

Lehrig:2018:CTS

[LSB^{+}18] Sebastian Lehrig, Richard Sanders, Gunnar Brataas, Mariano Cecowski, Simon Ivansek, and Jure Polutnik. CloudStore — towards scalability, elasticity, and efficiency benchmarking and analysis in cloud computing. Future Generation Computer Systems, 78 (part 1)(??):115– 126, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17306246.

Li:2019:ECR

[LSCL19] Chunlin Li, Hezhi Sun, Yi Chen, and Youlong Luo. Edge cloud resource expansion and shrinkage based on workload for minimizing the cost. Future Generation Computer Systems, 101(??):327–340, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331418.

Liu:2011:CPD

[LSD11] Shaofeng Liu, Jürgen P. Schulze, and Thomas A. DeFanti. CSTP: a parallel data transfer protocol using cross-stream coding. Future Generation Computer Systems, 27(7): 977–985, July 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Li:2017:SMP

[LSD^{+}17] Wei Li, Igor Santos, Flavia C. Delicato, Paulo F. Pires, Luci Pirmez, Wei Wei, Houbing Song, Albert Zomaya, and Samee Khan. System modelling and performance evaluation of a three-tier cloud of things. Future Generation Computer Systems, 70(??): 104–125, May 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302047.

Li:2018:SBS

[LSG18] Yinkui Li, Yongtang Shi, and Xiaofeng Gu. Spectrum bounds for the scattering number, integrity, tenacity of regular graphs. Future Generation Computer Systems, 83(??):450– 453, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17301292.

Lu:2019:AML

[LSG^{+}19] Feng Lu, Ziqian Shi, Lin Gu, Hai Jin, and Laurence Tianruo Yang. An adaptive multi-level caching strategy for distributed database system. Future Generation Computer Systems, 97(??):61– 68, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18313037.

Liu:2011:CEW

[LSH^{+}11] Shaofeng Liu, Jurgen P.

Schulze, Laurin Herr, Jeffrey D. Weekley, Bing Zhu, Natalie V. Osdol, Dana Plepys, and Mike Wan. CineGrid Exchange: a workflow-based peta-scale distributed storage platform on a high-speed network. Future Generation Computer Systems, 27(7):

966–976, July 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liu:2017:PSP

[LSHW17] Qin Liu, Avinash Srinivasan, Jiankun Hu, and Guojun Wang. Preface: Security and privacy in big data clouds. Future Generation Computer Systems, 72(??):206–207, July 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17305241.

Lu:2014:MDM

[LSJ^{+}14] Lu Lu, Xuanhua Shi, Hai Jin, Qiuyue Wang, Daxing Yuan, and Song Wu. Morpho: a decoupled MapReduce framework for elastic cloud computing. Future Generation Computer Systems, 36(??): 80–90, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002902.

Liu:2015:ATI

[LSL^{+}15] Ting Liu, Yanan Sun, Yang Liu, Yuhong Gui, Yucheng Zhao, Dai Wang, and Chao Shen. Abnormal trafficindexed state estimation: a cyber-physical fusion approach for Smart Grid attack detection. Future Generation Computer Systems, 49(??):94–103, August 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001861.

Liu:2018:CSA

[LSL^{+}18] Dan Liu, Xin Sui, Li Li, Zhengang Jiang, Huan Wang, Zetian Zhang, and Yan Zeng. A cloud service adaptive framework based on reliable resource allocation. Future Generation Computer Systems, 89(??):455–463, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18304606.

Lucas-Simarro:2013:SSO

[LSMVML13] Jose Luis Lucas-Simarro, Rafael Moreno-Vozmediano, Ruben S. Montero, and Ignacio M. Llorente. Scheduling strategies for optimal service deployment across multiple clouds. Future Generation Computer Systems, 29(6):1431–1441, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000192.

Logesh:2018:HQI

[LSV^{+}18] R. Logesh, V. Subramaniyaswamy, V. Vijayakumar, Xiao-Zhi Gao, and V. Indragandhi. A hybrid quantum-induced swarm intelligence clustering for the urban trip recommendation in smart city. Future Generation Computer Systems, 83(??):653– 673, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17319222.

Liu:2019:OSM

[LSW^{+}19] Weipeng Liu, Yaoguang Su, Wanqing Wu, Chen Xin, Zeng-Guang Hou, and Gui-Bin Bian. An operating smooth man-machine collaboration method for cataract capsulorhexis using virtual fixture. Future Generation Computer Systems, 98(??):522–529, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18322027.

Liao:2018:EEV

[LSYC18] Dan Liao, Gang Sun, Guanghua Yang, and Victor Chang. Energy-efficient virtual content distribution network provisioning in cloud-based data centers. Future Generation Computer Systems, 83(??):347– 357, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1732099X.

Li:2016:MBF

[LSZ^{+}16] Xiu Li, Jingdong Song, Fan Zhang, Xiaogang Ouyang, and Samee U. Khan. MapReduce-based fast fuzzy c-means algorithm for large-scale underwater image segmentation. Future Generation Computer Systems, 65(??):90– 101, December 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300486.

Li:2018:CEV

[LSZ^{+}18] Qingyong Li, Zhiping Shi, Huayan Zhang, Yunqiang Tan, Shengwei Ren, Peng Dai, and Weiyi Li. A cyber-enabled visual inspection system for rail corrugation. Future Generation Computer Systems, 79 (part 1)(??):374– 382, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17307069.

Lee:2012:LWA

[LTC12] Steven S. W. Lee, Po-Kai Tseng, and Alice Chen. Link weight assignment and loop-free routing table update for link state routing protocols in energy-aware internet. Future Generation Computer Systems, 28(2):437–445, February 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11000847.

Liu:2019:FFH

[LTC^{+}19] Zhiqiang Liu, Shuyang Tang, Sherman S. M. Chow, Zhen Liu, and Yu Long. Fork-free hybrid consensus with flexible proof-of-activity. Future Generation Computer Systems, 96(??):515–524, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18326256.

Lee:2012:SBD

[LTJK12] Hyunggu Lee, Andrew Beng Jin Teoh, Ho Gi Jung, and Jaihie Kim. A secure biometric discretization scheme for face template protection. Future Generation Computer Systems, 28(1):218–231, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X10002165.

Li:2019:DCB

[LTMW19] Wenjuan Li, Steven Tug, Weizhi Meng, and Yu Wang. Designing collaborative blockchained signaturebased intrusion detection in IoT environments. Future Generation Computer Systems, 96(??):481–489, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18327237.

Leon:2010:UER

[LTN10] Xavier León, Tuan Anh Trinh, and Leandro Navarro. Using economic regulation to prevent resource congestion in large-scale shared infrastructures. Future Generation Computer Systems, 26(4):599–607, April 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Li:2019:CCA

[LTTL19] Chunlin Li, Jianhang Tang, Hengliang Tang, and Youlong Luo. Collaborative cache allocation and task scheduling for dataintensive applications in edge computing environment. Future Generation Computer Systems, 95(??): 249–264, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18309439.

Lytra:2015:HAD

[LTZ15] Ioanna Lytra, Huy Tran, and Uwe Zdun. Harmonizing architectural decisions with component view models using reusable architectural knowledge transformations and constraints. Future Generation Computer Systems, 47(??):80– 96, June 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002441.

Liu:2014:GKF

[LvW14] Ziyuan Liu and Georg von Wichert. A generalizable knowledge framework for semantic indoor mapping based on Markov logic networks and data driven MCMC. Future Generation Computer Systems, 36 (??):42–56, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001404.

Li:2018:PCE

[LW18a] Xiaomin Li and Jiafu Wan.

Proactive caching for edge computing-enabled industrial mobile wireless networks. Future Generation Computer Systems, 89(??):89–97, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17329588.

Liu:2018:SAP

[LW18b] Shudong Liu and Lei Wang. A self-adaptive point-of-interest recommendation algorithm based on a multi-order Markov model. Future Generation Computer Systems, 89(??):506–514, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18300438.

Li:2019:TFA

[LW19] Wenting Li and Ping Wang. Two-factor authentication in industrial Internet-of-Things: Attacks, evaluation and new construction. Future Generation Computer Systems, 101(??):694–708, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19307642.

Li:2014:RAM

[LWD^{+}14] Xiaoling Li, Huaimin Wang, Bo Ding, Xiaoyong Li, and Dawei Feng. Resource allocation with multi-factor node ranking in data center networks. Future Generation Computer Systems, 32(??):1– 12, March 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002070.

Lu:2017:EFV

[LWF^{+}17] Yu Lu, Shiqian Wu, Zhijun Fang, Naixue Xiong, Sook Yoon, and Dong Sun Park. Exploring finger vein based personal authentication for secure IoT. Future Generation Computer Systems, 77(??):149–160, December 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17314772.

Lin:2018:DCB

[LWH^{+}18] Yaguang Lin, Xiaoming Wang, Fei Hao, Liang Wang, Lichen Zhang, and Ruonan Zhao. An ondemand coverage based self-deployment algorithm for big data perception in mobile sensing networks. Future Generation Computer Systems, 82(??):220– 234, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17313262.

Li:2018:SCM

[LWK^{+}18] Xiong Li, Fan Wu, Muhammad Khurram Khan, Lili Xu, Jian Shen, and Minho Jo. A secure chaotic map-based remote authentication scheme for telecare medicine information systems. Future Generation Computer Systems, 84(??):149–159, July 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1730688X.

Liu:2018:ABH

[LWL^{+}18] Yi Liu, Hao Wang, Tong Li, Ping Li, and Jie Ling. Attribute-based handshake protocol for mobile healthcare social networks. Future Generation Computer Systems, 86(??):873–880, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630749X.

Lu:2019:LLB

[LWR^{+}19] Siyang Lu, Xiang Wei, Bingbing Rao, Byungchul Tak, Long Wang, and Liqiang Wang. LADRA: Log-based abnormal task detection and root-cause analysis in big data processing with Spark. Future Generation Computer Systems, 95(??):392–403, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18307684.

Liu:2012:QSQ

[LWS^{+}12] Min Liu, Mingrui Wang, Weiming Shen, Nan Luo, and Junwei Yan. A quality of service (QoS)-aware execution plan selection approach for a service composition process. Future Generation Computer Systems, 28(7):1080–1089, July 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11002172.

Liu:2018:DME

[LWSY18] Wei Liu, Jing Wang, Arun Kumar Sangaiah, and Jian Yin. Dynamic metric embedding model for point-of-interest prediction. Future Generation Computer Systems, 83(??):183–192, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17320824.

Liu:2018:ELM

[LWT18] Zhen Liu, Ruoyu Wang, and Deyu Tang. Extending labeled mobile network traffic data by three levels traffic identification fusion. Future Generation Computer Systems, 88(??):453– 466, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17309937.

Li:2019:SDR

[LWTL19a] Chunlin Li, Chengyi Wang, Hengliang Tang, and Youlong Luo. Scalable and dynamic replica consistency maintenance for edge-cloud system. Future Generation Computer Systems, 101(??):590–604, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331558.

Li:2019:DMO

[LWTL19b] Chunlin Li, YaPing Wang, Hengliang Tang, and Youlong Luo. Dynamic multiobjective optimized replica placement and migration strategies for SaaS applications in edge cloud. Future Generation Computer Systems, 100(??):921–937, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18323124.

Lu:2013:IBV

[LWW^{+}13] Xicheng Lu, Huaimin Wang, Ji Wang, Jie Xu, and Dongsheng Li. Internet-based Virtual Computing Environment: Beyond the data center as a computer. Future Generation Computer Systems, 29(1):309–322, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001440.

Liu:2016:TRC

[LWW^{+}16] Jianqi Liu, Jiafu Wan, Qinruo Wang, Bi Zeng, and Shaoliang Fang. A timerecordable cross-layer communication protocol for the positioning of Vehicular Cyber-Physical Systems. Future Generation Computer Systems, 56(??): 438–448, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002721.

Lin:2018:EQA

[LWW^{+}18] Weiwei Lin, Wentai Wu, Haoyu Wang, James Z. Wang, and Ching-Hsien Hsu. Experimental and quantitative analysis of server power model for cloud data centers. Future Generation Computer Systems, 86(??):940–950, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X16306872.

Liu:2013:PBS

[LWX13] Zhigang Liu, Jinkuan Wang, and Yanbo Xue. PCRLB-based sensor selection for maneuvering target tracking in rangebased sensor networks. Future Generation Computer Systems, 29(7):1751–1757, September 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000155.

Liu:2019:EGC

[LWXY19] Shenghao Liu, Bang Wang, Minghua Xu, and Laurence T. Yang. Evolving graph construction for successive recommendation in event-based social networks. Future Generation Computer Systems, 96(??): 502–514, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1832332X.

Liu:2018:DSV

[LWYS18] Jianhua Liu, Xin Wang, Guangxue Yue, and Shigen Shen. Data sharing in VANETs based on evolutionary fuzzy game. Future Generation Computer Systems, 81(??):141–155, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17324081.

Liu:2018:DED

[LWZ18] Yang Liu, Wei Wei, and Ruqing Zhang. DESRP: an efficient differential evolution algorithm for stochastic demand-oriented resource placement in heterogeneous clouds. Future Generation Computer Systems, 88(??):234–242, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18303510.

Li:2019:IAK

[LWZ^{+}19a] Yanhong Li, Meng Wang, Rongbo Zhu, Ashiq Anjum, Xiaokun Du, Yuhe Feng, Changyin Luo, and Shasha Tian. Intelligent augmented keyword search on spatial entities in reallife Internet of Vehicles. Future Generation Computer Systems, 94(??):697– 711, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18320223.

Luo:2019:GII

[LWZ^{+}19b] Shuyun Luo, Zhenyu Wen, Xiaomei Zhang, Weiqiang Xu, Albert Y. Zomaya, and Rajiv Ranjan. GoSharing: an intelligent incentive framework based on users’ association for cooperative content sharing in mobile edge networks. Future Generation Computer Systems, 95(??):601– 614, June 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18320065.

Liu:2013:CSO

[LX13] Lu Liu and Jie Xu. Clouds and service-oriented architectures. Future Generation Computer Systems, 29(1):271–272, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000970.

Levitin:2017:ODP

[LXD17] Gregory Levitin, Liudong Xing, and Yuanshun Dai. Optimal data partitioning in cloud computing system with random server assignment. Future Generation Computer Systems, 70(??): 17–25, May 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16308160.

Lin:2019:DDC

[LXF19] Kai Lin, Fuzhen Xia, and Giancarlo Fortino. Datadriven clustering for multimedia communication in Internet of Vehicles. Future Generation Computer Systems, 94(??):610–619, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18317928.

Levitin:2018:ODS

[LXJD18] Gregory Levitin, Liudong Xing, Barry W. Johnson, and Yuanshun Dai. Optimization of dynamic spot-checking for collusion tolerance in grid computing. Future Generation Computer Systems, 86(??):30–38, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17305514.

Li:2014:EMK

[LXK^{+}14] Ruixuan Li, Zhiyong Xu, Wanshang Kang, Kin Choong Yow, and Cheng-Zhong Xu. Efficient multikeyword ranked query over encrypted data in cloud computing. Future Generation Computer Systems, 30(??):179–190, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300143X.

Li:2017:OAM

[LXL^{+}17] Huakang Li, Zheng Xu, Tao Li, Guozi Sun, and Kim-Kwang Raymond Choo. An optimized approach for massive web page classification using entity similarity based on semantic network. Future Generation Computer Systems, 76(??):510–518, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17303321.

Lu:2019:UUB

[LXL^{+}19] Qinghua Lu, Xiwei Xu, Yue Liu, Ingo Weber, Liming Zhu, and Weishan Zhang. uBaaS: a unified blockchain as a service platform. Future Generation Computer Systems, 101(??):564–575, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18319873.

Liu:2018:DCF

[LXM^{+}18] Min Liu, Zhen Xiong, Yue Ma, Peng Zhang, Jianliang Wu, and Xingqin Qi. DPRank centrality: Finding important vertices based on random walks with a new defined transition matrix. Future Generation Computer Systems, 83(??):376–389, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1732407X.

Lin:2015:RRP

[LXMW15] Hui Lin, Li Xu, Yi Mu, and Wei Wu. A reliable recommendation and privacypreserving based crosslayer reputation mechanism for mobile cloud computing. Future Generation Computer Systems, 52(??):125–136, November 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1400226X.

Li:2019:NIM

[LXRS19] Shupan Li, Limin Xiao, Li Ruan, and Shubin Su. A novel integrity measurement method based on copy-on-write for region in virtual machine. Future Generation Computer Systems, 97(??):714–726, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317618.

Liu:2019:BCM

[LXT^{+}19] Ningbo Liu, Yanan Xu, Yonghua Tian, Hongwei Ma, and Shuliang Wen. Background classification method based on deep learning for intelligent automotive radar target detection. Future Generation Computer Systems, 94(??): 524–535, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18318120.

Lu:2018:SUA

[LXZ^{+}18] Jianfeng Lu, Yun Xin, Zhao Zhang, Hao Peng, and Jianmin Han. Supporting user authorization queries in RBAC systems by role-permission reassignment. Future Generation Computer Systems, 88(??):707–717, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17324433.

Lee:2017:NEA

[LY17] Gangin Lee and Unil Yun. A new efficient approach for mining uncertain frequent patterns using minimum data structure without false positives. Future Generation Computer Systems, 68(??):89–110, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630303X.

Lee:2018:SPB

[LY18a] Gangin Lee and Unil Yun. Single-pass based efficient erasable pattern mining using list data structure on dynamic incremental databases. Future Generation Computer Systems, 80(??):12–28, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16307129.

Lin:2018:CEC

[LY18b] Huiqiu Lin and Weihua Yang. A conditional edge connectivity of doubleorbit networks. Future Generation Computer Systems, 83(??):445–449, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17319519.

Liang:2019:SGT

[LY19] Xueqin Liang and Zheng Yan. A survey on game theoretical methods in human–machine networks. Future Generation Computer Systems, 92(??):674– 693, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17317120.

Li:2018:ITS

[LYC18] Jin Li, Qiben Yan, and Victor Chang. Internet of Things: Security and privacy in a connected world. Future Generation Computer Systems, 78 (part 3)(??):931–932, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17319817.

Li:2019:SMU

[LYC^{+}19] Yukun Li, Zhenguo Yang, Xu Chen, Huaping Yuan, and Wenyin Liu. A stacking model using URL and HTML features for phishing webpage detection. Future Generation Computer Systems, 94(??):27– 39, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1830503X.

Luo:2019:CBF

[LYH^{+}19] Juan Luo, Luxiu Yin, Jinyu Hu, Chun Wang, Xuan Liu, Xin Fan, and Haibo Luo. Containerbased fog computing architecture and energybalancing scheduling algorithm for energy IoT. Future Generation Computer Systems, 97(??):50–60, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831358X.

Leu:2010:IRH

[LYJ10] Fang-Yie Leu, Chao-Tung Yang, and Fuu-Cheng Jiang. Improving reliability of a heterogeneous gridbased intrusion detection platform using levels of redundancies. Future Generation Computer Systems, 26(4):554–568, April 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liu:2015:SAA

[LYL15] Zhusong Liu, Hongyang Yan, and Zhike Li. Serveraided anonymous attributebased authentication in cloud computing. Future Generation Computer Systems, 52(??):61– 66, November 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002544.

Liu:2019:FCA

[LYL^{+}19] Duo Liu, Chaoshu Yang, Shiming Li, Xianzhang Chen, Jinting Ren, Renping Liu, Moming Duan, Yujuan Tan, and Liang Liang. FitCNN: A cloud-assisted and low-cost framework for updating CNNs on IoT devices. Future Generation Computer Systems, 91(??):277–289, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18312019.

Liu:2012:HSI

[LYS12] Xingang Liu, Laurence Tianruo Yang, and Kwanghoon Sohn. High-speed interview frame mode decision procedure for multiview video coding. Future Generation Computer Systems, 28(6):947–956, June 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11000951.

Li:2019:IUE

[LYS^{+}19] Bo Li, Guiqin Yuan, Li Shen, Ruoyi Zhang, and Yiyang Yao. Incorporating URL embedding into ensemble clustering to detect web anomalies. Future Generation Computer Systems, 96(??):176–184, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18320004.

Lu:2016:FTS

[LYW^{+}16] Kuan Lu, Ramin Yahyapour, Philipp Wieder, Edwin Yaqub, Monir Abdullah, Bernd Schloer, and Constantinos Kotsokalis.

Fault-tolerant service level agreement lifecycle management in clouds using actor system. Future Generation Computer Systems, 54(??):247–259, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000783.

Li:2018:NPD

[LYW^{+}18a] Jin Li, Yun Yang, Xiaoling Wang, Zhiming Zhao, and Tong Li. A novel parallel distance metric-based approach for diversified ranking on large graphs. Future Generation Computer Systems, 88(??):79– 91, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18301766.

Liu:2018:TMM

[LYW^{+}18b] Changxi Liu, Hailong Yang, Rui Wang, Zhongzhi Luan, and Depei Qian. T1000: Mitigating the memory footprint of convolution neural networks with decomposition and refusion. Future Generation Computer Systems, 84(??): 1–10, July 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17327395.

Liu:2014:FAV

[LYXT14] Yizhi Liu, Ying Yang, Hongtao Xie, and Sheng Tang. Fusing audio vocabulary with visual features for pornographic video detection. Future Generation Computer Systems, 31(??): 69–76, February 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001689.

Li:2018:PPC

[LYY^{+}18] Yannan Li, Yong Yu, Bo Yang, Geyong Min, and Huai Wu. Privacy preserving cloud data auditing with efficient key update. Future Generation Computer Systems, 78 (part 2)(??):789–798, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302989.

Liu:2019:DRL

[LYYW19] Zhuo Liu, Chenhui Yao, Hang Yu, and Taihua Wu. Deep reinforcement learning with its application for lung cancer detection in medical Internet of Things. Future Generation Computer Systems, 97(??):1– 9, August 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19303772.

Li:2017:BNB

[LYYY17] Zhihua Li, Chengyu Yan, Xinrong Yu, and Ning Yu. Bayesian network-based virtual machines consolidation method. Future Generation Computer Systems, 69(??):75–87, April 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16307415.

Li:2018:EAM

[LYYY18] Zhihua Li, Chengyu Yan, Lei Yu, and Xinrong Yu. Energy-aware and multiresource overload probability constraint-based virtual machine dynamic consolidation method. Future Generation Computer Systems, 80(??):139–156, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16307476.

Liu:2015:EIV

[LYZC15] Chang Liu, Chi Yang, Xuyun Zhang, and Jinjun Chen. External integrity verification for outsourced big data in cloud and IoT: a big picture. Future Generation Computer Systems, 49(??):58–67, August 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001551.

Lee:2010:RRJ

[LZ10] Young Choon Lee and Albert Y. Zomaya. Rescheduling for reliable job completion with the support of clouds. Future Generation Computer Systems, 26(8): 1192–1199, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liang:2019:PTM

[LZCX19] Xinle Liang, Shengchao Zhou, Huaping Chen, and Rui Xu. Pseudo transformation mechanism between resource allocation and bin-packing in batching environments. Future Generation Computer Systems, 95(??):79–88, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18304163.

Liang:2018:MCD

[LZH^{+}18] Wei Liang, Xiaokang Zhou, Suzhen Huang, Chunhua Hu, Xuesong Xu, and Qun Jin. Modeling of crossdisciplinary collaboration for potential field discovery and recommendation based on scholarly big data. Future Generation Computer Systems, 87(??):591–600, October 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17316424.

Lu:2019:PMM

[LZHY19] Gangzhao Lu, Weizhe Zhang, Hui He, and Laurence T. Yang. Performance modeling for MPI applications with low overhead fine-grained profiling. Future Generation Computer Systems, 90(??):317– 326, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18308252.

Liu:2019:SMI

[LZJL19] Yezheng Liu, Tingting Zhu, Yuanchun Jiang, and Xiao Liu. Service matchmaking for Internet of Things based on probabilistic topic model. Future Generation Computer Systems, 94(??):272–281, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18320788.

Liu:2012:RLR

[LZL^{+}12] Eryun Liu, Heng Zhao, Jimin Liang, Liaojun Pang, Hongtao Chen, and Jie Tian. Random local region descriptor (RLRD): a new method for fixedlength feature representation of fingerprint image and its application to template protection. Future Generation Computer Systems, 28(1):236–243, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11000021.

Liao:2016:DAQ

[LZL^{+}16] Ying-Ti Liao, Jiazheng Zhou, Chia-Hung Lu, Shih-Chang Chen, Ching-Hsien Hsu, Wenguang Chen, Mon-Fong Jiang, and Yeh-Ching Chung. Data adapter for querying and transformation between SQL and NoSQL database. Future Generation Computer Systems, 65(??):111– 121, December 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300085.

Li:2017:DCB

[LZL^{+}17] Jianjiang Li, Peng Zhang, Yuance Li, Wei Chen, Yajun Liu, and Lizhe Wang. A data-check based distributed storage model for storing hot temporary data. Future Generation Computer Systems, 73(??): 13–21, August 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630557X.

Liang:2019:DED

[LZL^{+}19a] Haibo Liang, Jialing Zou, Zhiling Li, Muhammad Junaid Khan, and Yanjun Lu. Dynamic evaluation of drilling leakage risk based on fuzzy theory and PSO-SVR algorithm. Future Generation Computer Systems, 95(??):454–466, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18321149.

Liu:2019:GRM

[LZL19b] Guanqun Liu, Shaohui Zhong, and Ting Li. Gait recognition method of temporal-spatial HOG features in critical separation of Fourier correction points. Future Generation Computer Systems, 94(??):

11–15, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18318508.

Li:2018:IDC

[LZLL18a] Jinqiang Li, Shuming Zhou, Xuequn Li, and Xiaowang Li. An insertion– deletion–compensation model with Poisson process for scale-free networks. Future Generation Computer Systems, 83(??):425–430, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17305848.

Liu:2018:SFG

[LZLL18b] Yi Liu, Yinghui Zhang, Jie Ling, and Zhusong Liu. Secure and fine-grained access control on e-healthcare records in mobile cloud computing. Future Generation Computer Systems, 78 (part 3)(??):1020–1026, January 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16308251.

Li:2018:MUA

[LZP^{+}18] Yongjun Li, Zhen Zhang, You Peng, Hongzhi Yin, and Quanqing Xu. Matching user accounts based on user generated content across social networks. Future Generation Computer Systems, 83(??):104–115, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324809.

Li:2018:MCP

[LZS18] Bin Li, Qinglei Zhou, and Xueming Si. Mimic computing for password recovery. Future Generation Computer Systems, 84(??): 58–77, July 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17320435.

Liang:2019:DAS

[LZT^{+}19] Baoyu Liang, Yunkai Zhai, Chao Tong, Jie Zhao, Jun Li, Xianying He, and Qianqian Ma. A deep automated skeletal bone age assessment model via regionbased convolutional neural network. Future Generation Computer Systems, 98(??):54–59, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326153.

Liu:2018:SPN

[LZW^{+}18] Guipeng Liu, Xiaomin Zhu, Ji Wang, Deke Guo, Weidong Bao, and Hui Guo. SP-Partitioner: A novel partition method to handle intermediate data skew in spark streaming. Future Generation Computer Systems, 86 (??):1054–1063, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17314784.

Lv:2019:MOF

[LZWF19] Li Lv, Jia Zhao, Jiayuan Wang, and Tanghuai Fan. Multi-objective firefly algorithm based on compensation factor and elite learning. Future Generation Computer Systems, 91(??):37–47, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18313955.

Liu:2016:HSM

[LZX16] Xiaocheng Liu, Ziming Zhong, and Kai Xu. A hybrid solution method for CFD applications on GPU-accelerated hybrid HPC platforms. Future Generation Computer Systems, 56(??):759–765, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500254X.

Liu:2012:DEE

[LZXG12] Zhixin Liu, Qingchao Zheng, Liang Xue, and Xinping Guan. A distributed energy-efficient clustering algorithm with improved coverage in wireless sensor networks. Future Generation Computer Systems, 28(5):780–790, May 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000707.

Li:2013:LBM

[LZXW13] Rui Li, Yinfeng Zhang, Zhengquan Xu, and Huayi Wu. A load-balancing method for network GISs in a heterogeneous clusterbased system using access density. Future Generation Computer Systems, 29(2):528–535, February 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001616.

Luo:2016:KMB

[LZY^{+}16] Xiong Luo, Dandan Zhang, Laurence T. Yang, Ji Liu, Xiaohui Chang, and Huansheng Ning. A kernel machine-based secure data sensing and fusion scheme in wireless sensor networks for the cyber-physical systems. Future Generation Computer Systems, 61(??): 85–96, August 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003350.

Lai:2019:CAC

[LZY^{+}19a] Yongxuan Lai, Lu Zhang, Fan Yang, Lv Zheng, Tian Wang, and Kuan-Ching Li. CASQ: Adaptive and cloud-assisted query processing in vehicular sensor networks. Future Generation Computer Systems, 94(??):237–249, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18302164.

Li:2019:CPC

[LZY^{+}19b] Jianjiang Li, Kai Zhang, Xiaolei Yang, Peng Wei, Jie Wang, Karan Mitra, and Rajiv Ranjan. Category preferred canopy-K-means based collaborative filtering algorithm. Future Generation Computer Systems, 93(??):1046–1054, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17308890.

Liu:2013:CSK

[LZYC13] Chang Liu, Xuyun Zhang, Chi Yang, and Jinjun Chen. CCBKE — session key negotiation for fast and secure scheduling of scientific applications in cloud computing. Future Generation Computer Systems, 29(5):1300– 1308, July 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001495.

Li:2019:CSA

[LZZ19] Chenghao Li, Huyin Zhang, and Tianying Zhou. Coflow scheduling algorithm based density peaks clustering. Future Generation Computer Systems, 97(??):805– 813, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331844.

Mohiuddin:2019:SDA

[MAA^{+}19] Irfan Mohiuddin, Ahmad Almogren, Mohammed Al Qurishi, Mohammad Mehedi Hassan, Iehab Al Rassan, and Giancarlo Fortino. Secure distributed adaptive bin packing algorithm for cloud storage. Future Generation Computer Systems, 90(??):307–316, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18304035.

Mohamed:2015:AAM

[MAB^{+}15] Mohamed Mohamed, Mourad Amziani, Djamel Belaïd, Samir Tata, and Tarek Melliti. An autonomic approach to manage elasticity of business processes in the cloud. Future Generation Computer Systems, 50(??):49–61, September 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002106.

Ma:2014:MAM

[MAC14] Lin Ma, Kunal Agrawal, and Roger D. Chamberlain. A memory access model for highly-threaded many-core architectures. Future Generation Computer Systems, 30(??):202– 215, January 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001349.

Mohammed:2017:BCI

[MAÇ17] Adnan Saher Mohammed, Sahin Emrah Amrahov, and Fatih V. Çelebi. Bidirectional conditional insertion sort algorithm; an efficient progress on the classical insertion sort. Future Generation Computer Systems, 71(??):102–112, June 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301711.

Merlino:2016:MCS

[MAD^{+}16] Giovanni Merlino, Stamatis Arkoulis, Salvatore Distefano, Chrysa Papagianni, Antonio Puliafito, and Symeon Papavassiliou. Mobile crowdsensing as a service: a platform for applications on top of sensing clouds. Future Generation Computer Systems, 56(??): 623–639, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002976.

Mariani:2018:PCP

[MAJD18] Giovanni Mariani, Andreea Anghel, Rik Jongerius, and Gero Dittmann. Predicting cloud performance for HPC applications before deployment. Future Generation Computer Systems, 87(??):618–628, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312542.

Mann:2015:RRE

[Man15] Zoltán Ádám Mann. Rigorous results on the effectiveness of some heuristics for the consolidation of virtual machines in a cloud data center. Future Generation Computer Systems, 51(??): 1–6, October 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000965.

Mompean:2019:GBP

[MAPA19] Juan Mompeán, Juan L. Aragón, Pedro M. Prieto, and Pablo Artal. GPU-based processing of Hartmann–Shack images for accurate and highspeed ocular wavefront sensing. Future Generation Computer Systems, 91(??):177–190, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18310045.

Matsui:2018:IPS

[Mat18] Kanae Matsui. An information provision system to promote energy conservation and maintain indoor comfort in smart homes using sensed data by IoT sensors. Future Generation Computer Systems, 82(??):388–394, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17313559.

Mirza:2018:CIM

[MAY18] Qublai K. Ali Mirza, Irfan Awan, and Muhammad Younas. CloudIntell: An intelligent malware detection system. Future Generation Computer Systems, 86(??):1042–1053, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17314929.

Mousa:2019:CAC

[MBA19] Afaf Mousa, Jamal Bentahar, and Omar Alam. Context-aware composite SaaS using feature model. Future Generation Computer Systems, 99(??):376– 390, October 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1832507X.

Malawski:2010:IOS

[MBB10] Maciej Malawski, Tomasz Bartyński, and Marian Bubak. Invocation of operations from script-based Grid applications. Future Generation Computer Systems, 26(1):138–146, January 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Margolis:2011:TCP

[MBC^{+}11] Todd Margolis, Sheldon Brown, Tracy Cornish, Hector Bracho, Michael Stanton, Tereza Cristina Melo de Brito Carvalho, Fernando Frota Redigolo, Fábio Carneiro de Castro, Kunitake Kaneko, Jane de Almeida, Cicero Inacio da Silva, and Eunézio Antônio de Souza. Tri-continental premiere of 4K feature movie via network streaming at FILE 2009. Future Generation Computer Systems, 27(7): 924–934, July 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Milovanovic:2019:AIU

[MBL^{+}19] Stevan Milovanović, Zorica Bogdanović, Aleksandra Labus, Dusan Barać, and Marijana Despotović-Zrakić. An approach to identify user preferences based on social network analysis. Future Generation Computer Systems, 93(??):121–129, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18311208.

Maqsood:2018:CAC

[MBM18] Tahir Maqsood, Kashif Bilal, and Sajjad A. Madani. Congestion-aware core mapping for networkon-chip based systems using betweenness centrality. Future Generation Computer Systems, 82(??):459– 471, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X16308445.

Masip-Bruin:2018:MRC

[MBMTJR18] Xavi Masip-Bruin, Eva Marin-Tordera, Admela Jukan, and Guang-Jie Ren. Managing resources continuity from the edge to the cloud: Architecture and performance. Future Generation Computer Systems, 79 (part 3)(??):777– 785, February 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302686.

Maurer:2013:ARC

[MBS13] Michael Maurer, Ivona Brandic, and Rizos Sakellariou. Adaptive resource configuration for Cloud infrastructure management. Future Generation Computer Systems, 29(2):472– 487, February 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001525.

Mustafa:2015:RGC

[MBV^{+}15] H. D. Mustafa, B. M. Baveja, S. Vijayan, S. N. Merchant, and U. B. Desai. Replicating the geographical cloud: Provisioning omnipresence, omniscience and omnipotence. Future Generation Computer Systems, 47(??):1– 15, June 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1400257X.

Ma:2018:ACA

[MCA^{+}18] Lin Ma, Roger D. Chamberlain, Kunal Agrawal, Chen Tian, and Ziang Hu. Analysis of classic algorithms on highly-threaded many-core architectures.

Future Generation Computer Systems, 82(??):528– 543, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17302029.

Martins:2019:ISR

[MCAS19] Rolando Martins, Manuel E.

Correia, Luís Antunes, and Fernando Silva. Iris: Secure reliable live-streaming with opportunistic mobile edge cloud offloading. Future Generation Computer Systems, 101(??):272–292, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301037.

Molto:2016:AMB

[MCdA16] Germán Moltó, Miguel Caballer, and Carlos de Alfonso. Automatic memorybased vertical elasticity and oversubscription on cloud platforms. Future Generation Computer Systems, 56(??):1–10, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003155.

Moreau:2011:OPM

[MCF^{+}11] Luc Moreau, Ben Clifford, Juliana Freire, Joe Futrelle, Yolanda Gil, Paul Groth, Natalia Kwasnikowska, Simon Miles, Paolo Missier, Jim Myers, Beth Plale, Yogesh Simmhan, Eric Stephan, and Jan Van den Bussche. The Open Provenance Model core specification (v1.1). Future Generation Computer Systems, 27(6):743–756, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Munoz:2015:SCB

[MCG^{+}15] Diego Muñoz, Raymundo Cornejo, Francisco J. Gutierrez, Jesús Favela, Sergio F. Ochoa, and Mónica Tentori. A social cloud-based tool to deal with time and media mismatch of intergenerational family communication. Future Generation Computer Systems, 53(??):140–151, December 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001356.

Ma:2019:ASF

[MCJ19] Yi-Wei Ma, Jiann-Liang Chen, and Jia-Yi Jhou. Adaptive service function selection for Network Function Virtualization networking. Future Generation Computer Systems, 91(??):108–123, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17328406.

Ma:2016:QAQ

[MCL^{+}16] Shang-Pin Ma, Kuan Y.

Chang, Jing-Hong Lin, Chih-Chun Ma, and Jui-Hsaing Lin. QoS-aware query relaxation for service discovery with business rules. Future Generation Computer Systems, 60 (??):1–12, July 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16000133.

Mahmood:2018:ECC

[MCN^{+}18] Khalid Mahmood, Shehzad Ashraf Chaudhry, Husnain Naqvi, Saru Kumari, Xiong Li, and Arun Kumar Sangaiah. An elliptic curve cryptography based lightweight authentication scheme for smart grid communication. Future Generation Computer Systems, 81(??):557–565, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17309263.

Merino:2016:DQU

[MCR^{+}16] Jorge Merino, Ismael Caballero, Bibiano Rivas, Manuel Serrano, and Mario Piattini. A data quality in use model for Big Data. Future Generation Computer Systems, 63(??):123– 130, October 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003817.

Mhenni:2019:ADZ

[MCRB19] Abir Mhenni, Estelle Cherrier, Christophe Rosenberger, and Najoua Essoukri Ben Amara. Analysis of Doddington zoo classification for user dependent template update: Application to keystroke dynamics recognition. Future Generation Computer Systems, 97(??):210–218, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18331236.

Moura:2018:EAR

[MCSA18] Douglas L. L. Moura, Raquel S. Cabral, Thiago Sales, and Andre L. L. Aquino. An evolutionary algorithm for roadside unit deployment with betweenness centrality preprocessing. Future Generation Computer Systems, 88(??):776–784, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17323592.

Mao:2015:SBQ

[MCT^{+}15] Chengying Mao, Jifu Chen, Dave Towey, Jinfu Chen, and Xiaoyuan Xie. Searchbased QoS ranking prediction for web services in cloud environments. Future Generation Computer Systems, 50(??):111–126, September 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000254.

Mwilu:2016:DSR

[MCWP16] Odette Sangupamba Mwilu, Isabelle Comyn-Wattiau, and Nicolas Prat. Design science research contribution to business intelligence in the cloud — a systematic literature review. Future Generation Computer Systems, 63(??):108–122, October 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003623.

Mazzucco:2012:BEB

[MD12] Michele Mazzucco and Dmytro Dyachuk. Balancing electricity bill and performance in server farms with setup costs. Future Generation Computer Systems, 28(2):415–426, February 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000665.

Mariano:2019:FAR

[MDA^{+}19] Angelo Mariano, Giulio D’Amato, Fiorenzo Ambrosino, Giuseppe Aprea, Francesco Buonocore, Massimo Celino, Antonio Colavincenzo, Marco Fina, Agostino Funel, Simone Giusepponi, Guido Guarnieri, Filippo Palombi, Samuele Pierattini, Giovanni Ponti, Giuseppe Santomauro, Giovanni Bracco, and Silvio Migliori. Fast Access to Remote Objects 2.0: a renewed gateway to ENEAGRID distributed computing resources. Future Generation Computer Systems, 94(??):920–928, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17308907.

Mason:2018:PHC

[MDB^{+}18a] Karl Mason, Martin Duggan, Enda Barrett, Jim Duggan, and Enda Howley. Predicting host CPU utilization in the cloud using evolutionary neural networks. Future Generation Computer Systems, 86(??):162–173, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17322793.

Meddeb:2018:AAF

[MDB^{+}18b] Maroua Meddeb, Amine Dhraief, Abdelfettah Belghith, Thierry Monteil, Khalil Drira, and Sofien Gannouni. AFIRM: Adaptive forwarding based link recovery for mobility support in NDN/IoT networks. Future Generation Computer Systems, 87(??):351–363, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17314528.

Moens:2015:ARC

[MDD15] Hendrik Moens, Bart Dhoedt, and Filip De Turck. Allocating resources for customizable multi-tenant applications in clouds using dynamic feature placement. Future Generation Computer Systems, 53(??):63–76, December 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002010.

Mirzaei:2019:AAA

[MdFTGM19] O. Mirzaei, J. M. de Fuentes, J. Tapiador, and L. Gonzalez-Manzano. AndrODet: an adaptive Android obfuscation detector. Future Generation Computer Systems, 90(??):240–261, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18309312.

Montella:2019:WBA

[MDM^{+}19] Raffaele Montella, Diana Di Luccio, Livia Marcellino, Ardelio Galletti, Sokol Kosta, Giulio Giunta, and Ian Foster. Workflow-based automatic processing for Internet of Floating Things crowdsourced data. Future Generation Computer Systems, 94(??):103–119, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18307672.

Marinho:2019:NEF

[MdMMNS^{+}19] Leandro B. Marinho, Navar de M. M. Nascimento, João Wellington M. Souza, Mateus Valentim Gurgel, Pedro P. Rebouças Filho, and Victor Hugo C. de Albuquerque. A novel electrocardiogram feature extraction approach for cardiac arrhythmia classification. Future Generation Computer Systems, 97(??): 564–577, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18323446.

Mattoso:2015:DSH

[MDO^{+}15] Marta Mattoso, Jonas Dias, Kary A. C. S. Ocaña, Eduardo Ogasawara, Flavio Costa, Felipe Horta, Vítor Silva, and Daniel de Oliveira. Dynamic steering of HPC scientific workflows: a survey. Future Generation Computer Systems, 46(??): 100–113, May 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002519.

Marinho:2017:DSW

[MdOO^{+}17] Anderson Marinho, Daniel de Oliveira, Eduardo Ogasawara, Vitor Silva, Kary Ocaña, Leonardo Murta, Vanessa Braganholo, and Marta Mattoso. Deriving scientific workflows from algebraic experiment lines: a practical approach. Future Generation Computer Systems, 68(??):111–127, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630276X.

Misale:2018:PHP

[MDT^{+}18] Claudia Misale, Maurizio Drocco, Guy Tremblay, Alberto R. Martinelli, and Marco Aldinucci. PiCo: High-performance data analytics pipelines in modern C++. Future Generation Computer Systems, 87(??):392–403, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1732681X.

Maurer:2012:CBA

[MEBA12] Michael Maurer, Vincent C. Emeakaroha, Ivona Brandic, and Jörn Altmann. Cost-benefit analysis of an SLA mapping approach for defining standardized Cloud computing goods. Future Generation Computer Systems, 28(1): 39–47, January 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001051.

Mezghani:2016:CMT

[MED16] Emna Mezghani, Ernesto Exposito, and Khalil Drira. A collaborative methodology for tacit knowledge management: Application to scientific research. Future Generation Computer Systems, 54(??):450–455, January 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500182X.

Merlo:2013:SCA

[Mer13] A. Merlo. Secure cooperative access control on Grid. Future Generation Computer Systems, 29(2):497– 508, February 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001574.

Mery:2017:PSB

[Mér17] Dominique Méry. Playing with state-based models for designing better algorithms. Future Generation Computer Systems, 68(??): 445–455, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301066.

Mohamed:2019:EPS

[MEW^{+}19] Mohamed Mohamed, Robert Engel, Amit Warke, Shay Berman, and Heiko Ludwig. Extensible persistence as a service for containers. Future Generation Computer Systems, 97(??): 10–20, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18303546.

Ma:2019:GBS

[MFC^{+}19] Shang-Pin Ma, Chen-Yuan Fan, Yen Chuang, I-Hsiu Liu, and Ci-Wei Lan. Graph-based and scenario-driven microservice analysis, retrieval, and testing. Future Generation Computer Systems, 100(??):724–735, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19302614.

McGough:2014:CCE

[MFG^{+}14] A. Stephen McGough, Matthew Forshaw, Clive Gerrard, Stuart Wheater, Ben Allen, and Paul Robinson. Comparison of a cost-effective virtual cloud cluster with an existing campus cluster. Future Generation Computer Systems, 41(??):65–78, December 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001344.

Mills:2018:MPS

[MFL18] Nicholas Mills, F. Alex Feltus, and Walter B. Ligon III. Maximizing the performance of scientific data transfer by optimizing the interface between parallel file systems and advanced research networks. Future Generation Computer Systems, 79 (part 1)(??):190– 198, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17306854.

Malawski:2013:CMC

[MFN13] Maciej Malawski, Kamil Figiela, and Jarek Nabrzyski. Cost minimization for computational applications on hybrid cloud infrastructures. Future Generation Computer Systems, 29(7):1786–1794, September 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000186.

Mukhutdinov:2019:MAD

[MFSV19] Dmitry Mukhutdinov, Andrey Filchenkov, Anatoly Shalyto, and Valeriy Vyatkin. Multi-agent deep learning for simultaneous optimization for time and energy in distributed routing system. Future Generation Computer Systems, 94(??):587– 600, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18309087.

Monster:2017:CIN

[MFT^{+}17] Dan Mønster, Riccardo Fusaroli, Kristian Tylén, Andreas Roepstorff, and Jacob F. Sherson. Causal inference from noisy timeseries data — testing the convergent cross-mapping algorithm in the presence of noise and external influence. Future Generation Computer Systems, 73(??): 52–62, August 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16307427.

Murphy:2010:VOC

[MG10] Michael A. Murphy and Sebastien Goasguen. Virtual organization clusters: Selfprovisioned clouds on the Grid. Future Generation Computer Systems, 26(8): 1271–1281, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Missier:2011:WOP

[MG11] Paolo Missier and Carole Goble. Workflows to open provenance graphs, roundtrip. Future Generation Computer Systems, 27(6): 812–819, June 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Macias:2014:SNE

[MG14] Mario Macías and Jordi Guitart. SLA negotiation and enforcement policies for revenue maximization and client classification in cloud providers. Future Generation Computer Systems, 41(??):19–31, December 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000491.

Macias:2016:ATM

[MG16] Mario Macías and Jordi Guitart. Analysis of a trust model for SLA negotiation and enforcement in cloud markets. Future Generation Computer Systems, 55(??):460–472, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000667.

Meilander:2018:MSR

[MG18] Dominik Meiländer and Sergei Gorlatch. Modeling the scalability of realtime online interactive applications on clouds. Future Generation Computer Systems, 86(??):1019–1031, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17315625.

Makowski:2019:EVT

[MG19] Lukasz Makowski and Paola Grosso. Evaluation of virtualization and traffic filtering methods for container networks. Future Generation Computer Systems, 93(??):345–357, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18302371.

Mohammed:2018:RTC

[MGA^{+}18] Mazin Abed Mohammed, Mohd Khanapi Abd Ghani, N. Arunkumar, Raed Ibraheem Hamed, Mohamad Khir Abdullah, and M. A. Burhanuddin. A real time computer aided object detection of nasopharyngeal carcinoma using genetic algorithm and artificial neural network based on Haar feature fear. Future Generation Computer Systems, 89(??):539–547, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18313670.

Mutlag:2019:ETF

[MGA^{+}19] Ammar Awad Mutlag, Mohd Khanapi Abd Ghani, N. Arunkumar, Mazin Abed Mohammed, and Othman Mohd. Enabling technologies for fog computing in healthcare IoT systems. Future Generation Computer Systems, 90(??):62– 78, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314006.

Michon:2017:SBI

[MGG^{+}17] Étienne Michon, Julien Gossa, Stéphane Genaud, Léo Unbekandt, and Vincent Kherbache. Schlouder: a broker for IaaS clouds. Future Generation Computer Systems, 69(??):11– 23, April 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16303077.

Martis:2018:RAB

[MGL^{+}18] Roshan Joy Martis, Varadraj Prabhu Gurupur, Hong Lin, Aminul Islam, and Steven Lawrence Fernandes. Recent advances in big data analytics, Internet of Things and machine learning. Future Generation Computer Systems, 88(??):696–698, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18317709.

Mondejar:2013:CTI

[MGLPPJ13] Rubén Mondéjar, Pedro García-López, Carles Pairot, and Lluis Pamies-Juarez. CloudSNAP: a transparent infrastructure for decentralized Web deployment using distributed interception. Future Generation Computer Systems, 29(1):370–380, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001671.

Melab:2018:MCV

[MGMT18] N. Melab, J. Gmys, M. Mezmaz, and D. Tuyttens. Multi-core versus many-core computing for many-task branch-and-bound applied to big optimization problems. Future Generation Computer Systems, 82(??):472–481, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X16308706.

Moosavi:2016:EES

[MGN^{+}16] Sanaz Rahimi Moosavi, Tuan Nguyen Gia, Ethiopia Nigussie, Amir M. Rahmani, Seppo Virtanen, Hannu Tenhunen, and Jouni Isoaho. End-to-end security scheme for mobility enabled healthcare Internet of Things. Future Generation Computer Systems, 64(??):108–124, November 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300334.

Mateescu:2011:HCW

[MGR11] Gabriel Mateescu, Wolfgang Gentzsch, and Calvin J. Ribbens. Hybrid computing — where HPC meets grid and Cloud Computing. Future Generation Computer Systems, 27(5): 440–453, May 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Makaratzis:2018:EMC

[MGT18] Antonios T. Makaratzis, Konstantinos M. Giannoutakis, and Dimitrios Tzovaras. Energy modeling in cloud simulation frameworks. Future Generation Computer Systems, 79 (part 2)(??):715– 725, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17303229.

Mariotti:2018:SST

[MGV^{+}18] Mirko Mariotti, Osvaldo Gervasi, Flavio Vella, Alfredo Cuzzocrea, and Alessandro Costantini. Strategies and systems towards grids and clouds integration:a DBMS-based solution. Future Generation Computer Systems, 88(??):718–729, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302996.

Martin:2019:CFD

[MH19] Ignacio Martín and José Alberto Hernández. CloneSpot: Fast detection of Android repackages. Future Generation Computer Systems, 94(??):740–748, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317497.

Malik:2014:LBG

[MHC14] Sheheryar Malik, Fabrice Huet, and Denis Caromel. Latency based group discovery algorithm for network aware cloud scheduling. Future Generation Computer Systems, 31(??): 28–39, February 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001830.

Martin:2019:MLB

[MHdlS19] Ignacio Martín, José Alberto Hernández, and Sergio de los Santos. Machinelearning based analysis and classification of Android malware signatures. Future Generation Computer Systems, 97(??):295–305, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18325159.

Mallig:2016:MWE

[MHW^{+}16] Nicolai Mallig, Michael Heilig, Christine Weiss, Bastian Chlond, and Peter Vortisch. Modelling the weekly electricity demand caused by electric cars. Future Generation Computer Systems, 64(??):140–150, November 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16000273.

Mohd:2018:HDM

[MHY^{+}18] Bassam Jamil Mohd, Thaier Hayajneh, Khalil M. Ahmad Yousef, Zaid Abu Khalaf, and Md Zakirul Alam Bhuiyan. Hardware design and modeling of lightweight block ciphers for secure communications. Future Generation Computer Systems, 83(??):510– 521, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17304661.

Meng:2018:PAC

[MHZK18] Yunfei Meng, Zhiqiu Huang, Yu Zhou, and Changbo Ke. Privacy-aware cloud service selection approach based on P-Spec policy models and privacy sensitivities. Future Generation Computer Systems, 86(??):1–11, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17321271.

Mahmud:2016:CED

[MID16] Shahid Mahmud, Rahat Iqbal, and Faiyaz Doctor. Cloud enabled data analytics and visualization framework for healthshocks prediction. Future Generation Computer Systems, 65(??):169–181, December 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003271.

Miles:2011:MAM

[Mil11] Simon Miles. Mapping attribution metadata to the Open Provenance Model. Future Generation Computer Systems, 27(6):806– 811, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Malawski:2015:ACD

[MJDN15] Maciej Malawski, Gideon Juve, Ewa Deelman, and Jarek Nabrzyski. Algorithms for cost- and deadline-constrained provisioning for scientific workflow ensembles in IaaS clouds. Future Generation Computer Systems, 48 (??):ibc, July 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000059.

Ma:2018:BDR

[MJGW18] Zhaofeng Ma, Ming Jiang, Hongmin Gao, and Zhen Wang. Blockchain for digital rights management. Future Generation Computer Systems, 89(??):746–764, December 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18301614.

Maheshwari:2016:WPI

[MJM^{+}16] Ketan Maheshwari, Eun-Sung Jung, Jiayuan Meng, Vitali Morozov, Venkatram Vishwanath, and Rajkumar Kettimuthu. Workflow performance improvement using model-based scheduling over multiple clusters and clouds. Future Generation Computer Systems, 54(??):206–218, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000795.

Mofrad:2016:SLA

[MJRM16] Mohammad Hasanzadeh Mofrad, Omid Jalilian, Alireza Rezvanian, and Mohammad Reza Meybodi. Service level agreement based adaptive grid superscheduling. Future Generation Computer Systems, 55(??):62–73, February 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002654.

Medeiros:2016:ETTa

[MK16a] Claudia Bauzer Medeiros and Daniel S. Katz. eScience today and tomorrow. Future Generation Computer Systems, 56(??): 523–525, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003295.

Medeiros:2016:ETTb

[MK16b] Claudia Bauzer Medeiros and Daniel S. Katz. eScience today and tomorrow — Part 2. Future Generation Computer Systems, 59(??):93–94, June 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003945.

Mai:2017:DIS

[MK17] Vu Mai and Ibrahim Khalil. Design and implementation of a secure cloud-based billing model for smart meters as an Internet of Things using homomorphic cryptography. Future Generation Computer Systems, 72(??):327– 338, July 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301832.

Makkar:2019:CSF

[MK19a] Aaisha Makkar and Neeraj Kumar. Cognitive spammer: a framework for PageRank analysis with split by over-sampling and train by under-fitting. Future Generation Computer Systems, 90(??):381–404, January 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18305703.

Mavridis:2019:CCV

[MK19b] Ilias Mavridis and Helen Karatza. Combining containers and virtual machines to enhance isolation and extend functionality on cloud computing. Future Generation Computer Systems, 94(??):674– 696, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18305764.

Mauch:2013:HPC

[MKH13] Viktor Mauch, Marcel Kunze, and Marius Hillenbrand. High performance cloud computing. Future Generation Computer Systems, 29(6):1408–1416, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000647.

Marosi:2013:TVC

[MKK13] Attila Marosi, József Kovács, and Peter Kacsuk. Towards a volunteer cloud system. Future Generation Computer Systems, 29(6):1442– 1451, August 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000660.

Mishra:2011:HRL

[MKM11] Shakti Mishra, D. S. Kushwaha, and A. K. Misra. Hybrid reliable load balancing with MOSIX as middleware and its formal verification using process algebra. Future Generation Computer Systems, 27(5): 506–526, May 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Murugan:2019:SDD

[MKRD19] Muthukumar Murugan, Krishna Kant, Ajaykrishna Raghavan, and David H. C. Du. Software defined deduplicated replica management in scale-out storage systems. Future Generation Computer Systems, 97(??):340–354, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18323884.

Mokatren:2018:EPM

[MKS18] Moayad Mokatren, Tsvi Kuflik, and Ilan Shimshoni. Exploring the potential of a mobile eye tracker as an intuitive indoor pointing device: a case study in cultural heritage. Future Generation Computer Systems, 81(??):528–541, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1731470X.

Moin:2019:SID

[MKS^{+}19] Sana Moin, Ahmad Karim, Zanab Safdar, Kalsoom Safdar, Ejaz Ahmed, and Muhammad Imran. Securing IoTs in distributed blockchain: Analysis, requirements and open issues. Future Generation Computer Systems, 100(??):325–343, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18330851.

Manuali:2011:GNC

[ML11] C. Manuali and A. Laganà. GriF: a new collaborative framework for a Web service approach to grid empowered calculations. Future Generation Computer Systems, 27 (3):315–318, March 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Munoz:2017:DEB

[ML17] Cristina Muñoz and Pierre Leone. A distributed event-based system based on compressed fragmentediterated Bloom filters. Future Generation Computer Systems, 75(??):108–127, October 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302200.

Manashty:2019:LMN

[ML19] Alireza Manashty and Janet Light. Life Model: A novel representation of lifelong temporal sequences in health predictive analytics. Future Generation Computer Systems, 92(??):141– 156, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17326523.

Moscicki:2011:PMT

[MLBS11] J. T. Mościcki, M. Lamanna, M. Bubak, and P. M. A. Sloot. Processing moldable tasks on the grid: Late job binding with lightweight user-level overlay. Future Generation Computer Systems, 27(6):725–736, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mambretti:2011:HPD

[MLC^{+}11] Joe Mambretti, Mathieu Lemay, Scott Campbell, Hervé Guy, Thomas Tam, Eric Bernier, Bobby Ho, Michel Savoie, Cees de Laat, Ronald van der Pol, Jim Chen, Fei Yeh, Sergi Figuerola, Pau Minoves, Dimitra Simeonidou, Eduard Escalona, Norberto Amaya Gonzalez, Admela Jukan, Wolfgang Bziuk, Dongkyun Kim, et al. High Performance Digital Media Network (HPDMnet): An advanced international research initiative and global experimental testbed. Future Generation Computer Systems, 27(7):893–905, July 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mahmood:2018:PBA

[MLC^{+}18a] Khalid Mahmood, Xiong Li, Shehzad Ashraf Chaudhry, Husnain Naqvi, Saru Kumari, Arun Kumar Sangaiah, and Joel J. P. C. Rodrigues. Pairing based anonymous and secure key agreement protocol for smart grid edge computing infrastructure. Future Generation Computer Systems, 88(??):491–500, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17327188.

Manogaran:2018:MCB

[MLC18b] Gunasekaran Manogaran, Daphne Lopez, and Naveen Chilamkurti. In-mapper combiner based MapReduce algorithm for processing of big climate data. Future Generation Computer Systems, 86(??):433–445, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324639.

Ma:2013:CFP

[MLG13] Jianwei Ma, Wanyu Liu, and Tristan Glatard. A classification of file placement and replication methods on grids. Future Generation Computer Systems, 29(6):1395–1406, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000344.

Meana-Llorian:2017:IFL

[MLGGB^{+}17] Daniel Meana-Llorián, Cristian González García, B. Cristina Pelayo G-Bustelo, Juan Manuel Cueva Lovelle, and Nestor Garcia-Fernandez. IoFClime: the fuzzy logic and the Internet of Things to control indoor temperature regarding the outdoor ambient conditions. Future Generation Computer Systems, 76(??):275–284, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306598.

Ma:2015:CSD

[MLL15] Shang-Pin Ma, Ci-Wei Lan, and Chia-Hsueh Li. Contextual service discovery using term expansion and binding coverage analysis. Future Generation Computer Systems, 48(??): ibc, July 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001836.

Mendez:2016:PES

[MLM16] Alejandro Pérez Méndez, Rafael Marín López, and Gabriel López Millán. Providing efficient SSO to cloud service access in AAA-based identity federations. Future Generation Computer Systems, 58(??): 13–28, May 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003854.

Moca:2016:MCS

[MLSF16] Mircea Moca, Cristian Litan, Gheorghe Cosmin Silaghi, and Gilles Fedak. Multi-criteria and satisfaction oriented scheduling for hybrid distributed computing infrastructures. Future Generation Computer Systems, 55(??):428–443, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000849.

Ma:2018:SFG

[MLW^{+}18a] Xingpo Ma, Junbin Liang, Jianxin Wang, Sheng Wen, Tian Wang, Yin Li, Wenpeng Ma, and Chuanda Qi. Secure fine-grained spatiotemporal top-k queries in TMWSNs. Future Generation Computer Systems, 86(??):174–184, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17319441.

Mo:2018:ERS

[MLW^{+}18b] Yijun Mo, Bixi Li, Bang Wang, Laurence T. Yang, and Minghua Xu. Event recommendation in social networks based on reverse random walk and participant scale control. Future Generation Computer Systems, 79 (part 1)(??): 383–395, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17303059.

Martinelli:2010:UCG

[MM10] Fabio Martinelli and Paolo Mori. On usage control for GRID systems. Future Generation Computer Systems, 26(7):1032–1042, July 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Moscicki:2018:CSS

[MM18] Jakub T. Mościcki and Luca Mascetti. Cloud storage services for file synchronization and sharing in science, education and research. Future Generation Computer Systems, 78 (part 3)(??):1052–1054, January 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17319842.

Munoz-Martinez:2019:ITC

[MMAA19] Francisco Muñoz-Martínez, José L. Abellán, and Manuel E. Acacio. InsideNet: a tool for characterizing convolutional neural networks. Future Generation Computer Systems, 100(??):298–315, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302651.

Merelli:2018:LPP

[MMC^{+}18] Ivan Merelli, Lucia Morganti, Elena Corni, Carmelo Pellegrino, Daniele Cesini, Luca Roverelli, Gabriele Zereik, and Daniele D’Agostino. Low-power portable devices for metagenomics analysis: Fog computing makes bioinformatics ready for the Internet of Things. Future Generation Computer Systems, 88(??):467– 478, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17324123.

Matsunaga:2016:WEC

[MMF16] Andréa Matsunaga, Austin Mast, and José A. B. Fortes. Workforce-efficient consensus in crowdsourced transcription of biocollections information. Future Generation Computer Systems, 56(??):526–536, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002277.

Ma:2018:ATB

[MML^{+}18] Xindi Ma, Jianfeng Ma, Hui Li, Qi Jiang, and Sheng Gao. ARMOR: A trust-based privacypreserving framework for decentralized friend recommendation in online social networks. Future Generation Computer Systems, 79 (part 1)(??):82– 94, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17300547.

Mrowczynski:2018:BMF

[MMLO18] Piotr Mrówczyński, Jakub T. Mościcki, Massimo Lamanna, and Frederik Orellana. Benchmarking and monitoring framework for interconnected file synchronization and sharing services. Future Generation Computer Systems, 78 (part 3)(??):1083–1090, January 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17303436.

Munoz:2019:EAE

[MMPF19] Daniel-Jesus Munoz, José A.

Montenegro, Mónica Pinto, and Lidia Fuentes. Energyaware environments for the development of green applications for cyberphysical systems. Future Generation Computer Systems, 91(??):536–554, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18307295.

Muresano:2017:AEE

[MMRL17] Ronal Muresano, Hugo Meyer, Dolores Rexachs, and Emilio Luque. An approach for an efficient execution of SPMD applications on multi-core environments. Future Generation Computer Systems, 66(??):11–26, January 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302011.

Mian:2013:PDA

[MMVP13] Rizwan Mian, Patrick Martin, and Jose Luis Vazquez-Poletti. Provisioning data analytic workloads in a cloud. Future Generation Computer Systems, 29(6):1452–1458, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000209.

Michail:2019:PLS

[MMVS19] Dimitrios Michail, Antonios Makris, Iraklis Varlamis, and Mark Sawyer. Parallelization of largescale drug-protein binding experiments. Future Generation Computer Systems, 97(??):492–502, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1830058X.

Mao:2018:MMD

[MNC^{+}18] Kaili Mao, Jianwei Niu, Huan Chen, Lei Wang, and Mohammed Atiquzzaman. Mining of marital distress from microblogging social networks: a case study on Sina Weibo. Future Generation Computer Systems, 86(??):1481–1490, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17310543.

Mualla:2019:ABS

[MND^{+}19] Yazan Mualla, Amro Najjar, Alaa Daoud, Stéphane Galland, Christophe Nicolle, Ansar-Ul-Haque Yasar, and Elhadi Shakshuki. Agent-based simulation of unmanned aerial vehicles in civilian applications: a systematic literature review and research directions. Future Generation Computer Systems, 100(??):344–364, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18328462.

Maheshwari:2012:DEE

[MNV12] Nitesh Maheshwari, Radheshyam Nanduri, and Vasudeva Varma. Dynamic energy efficient data placement and cluster reconfiguration algorithm for MapReduce framework. Future Generation Computer Systems, 28(1):119– 127, January 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1100135X.

Megahed:2019:OCS

[MNY^{+}19] Aly Megahed, Ahmed Nazeem, Peifeng Yin, Samir Tata, Hamid Reza Motahari Nezhad, and Taiga Nakamura. Optimizing cloud solutioning design. Future Generation Computer Systems, 91(??):86–95, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18306150.

Martinez:2018:LND

[MOBD18] Aritz Martinez, Eneko Osaba, Miren Nekane Bilbao, and Javier Del Ser. Let nature decide its nature: On the design of collaborative hyperheuristics for decentralized ephemeral environments. Future Generation Computer Systems, 88(??):792–805, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17323555.

Morillo:2018:CWI

[MOFGP18] Pedro Morillo, Juan M. Orduña, Marcos Fernández, and Inmaculada García-Pereira. Comparison of WSN and IoT approaches for a real-time monitoring system of meal distribution trolleys: A case study. Future Generation Computer Systems, 87(??):242–250, October 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17315741.

Mazumdar:2017:PES

[MP17] Somnath Mazumdar and Marco Pranzo. Power efficient server consolidation for cloud data center. Future Generation Computer Systems, 70(??):4– 16, May 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308093.

Matri:2018:KSD

[MPC^{+}18] Pierre Matri, María S.

Pérez, Alexandru Costan, Luc Bougé, and Gabriel Antoniu. Keeping up with storage: Decentralized, write-enabled dynamic geo-replication. Future Generation Computer Systems, 86(??): 1093–1105, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

https://www.sciencedirect.

                                        com/science/article/pii/
                                        S0167739X17312402.

Martinez-Prieto:2015:SAR

[MPCAF15] Miguel A. Martínez-Prieto, Carlos E. Cuesta, Mario Arias, and Javier D. Fernández. The Solid architecture for real-time management of big semantic data. Future Generation Computer Systems, 47 (??):62–79, June 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1400209X.

Mate:2016:HIA

[MPF^{+}16] Alejandro Maté, Jesús Peral, Antonio Ferrández, David Gil, and Juan Trujillo. A hybrid integrated architecture for energy consumption prediction. Future Generation Computer Systems, 63(??):131–147, October 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300644.

Memos:2018:EAM

[MPI^{+}18] Vasileios A. Memos, Kostas E. Psannis, Yutaka Ishibashi, Byung-Gyu Kim, and B. B. Gupta. An Efficient Algorithm for Mediabased Surveillance System (EAMSuS) in IoT smart city framework. Future Generation Computer Systems, 83(??):619–628, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17307707.

Morabito:2018:LLE

[MPLM18] Roberto Morabito, Riccardo Petrolo, Valeria Loscrì, and Nathalie Mitton. LEGIoT: a lightweight edge gateway for the Internet of Things. Future Generation Computer Systems, 81(??):1–15, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17306593.

Malensek:2013:EGC

[MPP13] Matthew Malensek, Sangmi Lee Pallickara, and Shrideep Pallickara. Exploiting geospatial and chronological characteristics in data streams to enable efficient storage and retrievals. Future Generation Computer Systems, 29(4):1049–1061, June 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1200132X.

Messina:2016:TAS

[MPR^{+}16] F. Messina, G. Pappalardo, D. Rosaci, C. Santoro, and G. M. L. Sarné. A trustaware, self-organizing system for large-scale federations of utility computing infrastructures. Future Generation Computer Systems, 56(??):77–94, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002368.

Meng:2019:PPS

[MQL^{+}19] Shunmei Meng, Lianyong Qi, Qianmu Li, Wenmin Lin, Xiaolong Xu, and Shaohua Wan. Privacypreserving and sparsityaware location-based prediction method for collaborative recommender systems. Future Generation Computer Systems, 96(??): 324–335, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18318053.

Muzammal:2019:RBD

[MQN19] Muhammad Muzammal, Qiang Qu, and Bulat Nasrulin. Renovating blockchain with distributed databases: an open source system. Future Generation Computer Systems, 90(??):105–117, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18308732.

Migdal:2019:SMK

[MR19] Denis Migdal and Christophe Rosenberger. Statistical modeling of keystroke dynamics samples for the generation of synthetic datasets. Future Generation Computer Systems, 100(??):907–920, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331212.

Mesbahi:2017:HRA

[MRH17] Mohammad Reza Mesbahi, Amir Masoud Rahmani, and Mehdi Hosseinzadeh. Highly reliable architecture using the 80/20 rule in cloud computing datacenters. Future Generation Computer Systems, 77(??):77–86, December 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17312396.

Moonsamy:2014:MPP

[MRL14] Veelasha Moonsamy, Jia Rong, and Shaowu Liu. Mining permission patterns for contrasting clean and malicious Android applications. Future Generation Computer Systems, 36(??):122–132, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001933.

Muhuri:2019:ASR

[MRN19] Pranab K. Muhuri, Amit Rauniyar, and Rahul Nath. On arrival scheduling of real-time precedence constrained tasks on multiprocessor systems using genetic algorithm. Future Generation Computer Systems, 93(??):702–726, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18312810.

Morillo:2010:EPS

[MROD10] P. Morillo, S. Rueda, J. M. Orduña, and J. Duato. Ensuring the performance and scalability of peer-to-peer distributed virtual environments. Future Generation Computer Systems, 26(7):905–915, July 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Moreira:2018:SIP

[MRS^{+}18a] Mário W. L. Moreira, Joel J. P. C. Rodrigues, Arun K. Sangaiah, Jalal Al-Muhtadi, and Valery Korotaev. Semantic interoperability and pattern classification for a serviceoriented architecture in pregnancy care. Future Generation Computer Systems, 89(??):137–147, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17325438.

Muralidharan:2018:MIM

[MRS18b] Shapna Muralidharan, Abhishek Roy, and Navrati Saxena. MDP-IoT: MDP based interest forwarding for heterogeneous traffic in IoT-NDN environment. Future Generation Computer Systems, 79 (part 3)(??):892–908, February 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1731035X.

Martin:2019:MHR

[MRT^{+}19] Paul Martin, Laurent Remy, Maria Theodoridou, Keith Jeffery, and Zhiming Zhao. Mapping heterogeneous research infrastructure metadata into a unified catalogue for use in a generic virtual research environment. Future Generation Computer Systems, 101(??):1–13, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19302699.

Martins:2019:MFB

[MS19] Paulo Martins and Leonel Sousa. A methodical FHEbased cloud computing model. Future Generation Computer Systems, 95(??): 639–648, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18316819.

Maimaiti:2019:SAK

[MSA^{+}19] Palida Maimaiti, Li Feng Sen, Guhaer Aisilahong, Rukeya Maimaiti, and Wu Yun Yun. Statistical analysis with Kruskal Wallis test for patients with joint contracture. Future Generation Computer Systems, 92(??):419–423, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18322623.

Mehrotra:2016:TAP

[MSBA16] Rajat Mehrotra, Srishti Srivastava, Ioana Banicescu, and Sherif Abdelwahed. Towards an autonomic performance management approach for a cloud broker environment using a decompositioncoordination based methodology. Future Generation Computer Systems, 54(??):195–205, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000825.

Momenzadeh:2019:WSA

[MSE19] Zahra Momenzadeh and Faramarz Safi-Esfahani. Workflow scheduling applying adaptable and dynamic fragmentation (WSADF) based on runtime conditions in cloud computing. Future Generation Computer Systems, 90(??):327– 346, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18305909.

Matsuhashi:2012:TVF

[MSI^{+}12] Yohei Matsuhashi, Takahiro Shinagawa, Yoshiaki Ishii, Nobuyuki Hirooka, and Kazuhiko Kato. Transparent VPN failure recovery with virtualization. Future Generation Computer Systems, 28(1):78–84, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001026.

Montes:2013:GCA

[MSM^{+}13] Jesús Montes, Alberto Sánchez, Bunjamin Memishi, María S. Pérez, and Gabriel Antoniu. GMonE: a complete approach to cloud monitoring. Future Generation Computer Systems, 29(8):2026–2040, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000496.

Meyer:2018:OPS

[MSM^{+}18a] Hugo Meyer, Jose Carlos Sancho, Milica Mrdakovic, Wang Miao, and Nicola Calabretta. Optical packet switching in HPC. An analysis of applications performance. Future Generation Computer Systems, 82(??):606–616, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1730300X.

Muhammad:2018:ISU

[MSM^{+}18b] Khan Muhammad, Muhammad Sajjad, Irfan Mehmood, Seungmin Rho, and Sung Wook Baik. Image steganography using uncorrelated color space and its application for security of visual contents in online social networks. Future Generation Computer Systems, 86(??):951–960, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X16306768.

Monajemi:2018:IRC

[MSO18] Sadaf Monajemi, Saeid Sanei, and Sim-Heng Ong. Information reliability in complex multitask networks. Future Generation Computer Systems, 83(??):485–495, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17315194.

Madougou:2013:CWB

[MSS^{+}13] Souley Madougou, Shayan Shahand, Mark Santcroos, Barbera van Schaik, Ammar Benabdelkader, Antoine van Kampen, and Sílvia Olabarriaga. Characterizing workflow-based activity on a production einfrastructure using provenance data. Future Generation Computer Systems, 29(8):1931–1942, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000794.

Mizeranschi:2016:MMD

[MSS^{+}16] Alexandru E. Mizeranschi, Martin T. Swain, Raluca Scona, Quentin Fazilleau, Bartosz Bosak, Tomasz Piontek, Piotr Kopta, Paul Thompson, and Werner Dubitzky. MultiGrain/ MAPPER: a distributed multiscale computing approach to modeling and simulating gene regulation networks. Future Generation Computer Systems, 63(??):1–14, October 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300711.

Macker:2017:OAD

[MT17] Joseph P. Macker and Ian Taylor. Orchestration and analysis of decentralized workflows within heterogeneous networking infrastructures. Future Generation Computer Systems, 75(??):388–401, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300262.

Mencagli:2018:EPT

[MTD18] Gabriele Mencagli, Massimo Torquati, and Marco Danelutto. Elastic-PPQ: a two-level autonomic system for spatial preference query processing over dynamic data streams. Future Generation Computer Systems, 79 (part 3)(??):862– 877, February 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1730938X.

Muthuvelu:2013:TGP

[MVC^{+}13] Nithiapidary Muthuvelu, Christian Vecchiola, Ian Chai, Eswaran Chikkannan, and Rajkumar Buyya. Task granularity policies for deploying bagof-task applications on global grids. Future Generation Computer Systems, 29(1):170–181, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000878.

Munoz:2010:EAR

[MVCC10] Víctor Méndez Muñoz, Gabriel Amorós Vicente, Félix García Carballeira, and José Salt Cairols. Emergent algorithms for replica location and selection in data grid. Future Generation Computer Systems, 26(7):934–946, July 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Menychtas:2014:MAB

[MVG^{+}14] Andreas Menychtas, Jürgen Vogel, Andrea Giessmann, Anna Gatzioura, Sergio Garcia Gomez, Vrettos Moulos, Frederic Junker, Mathias Müller, Dimosthenis Kyriazis, Katarina Stanoevska-Slabeva, and Theodora Varvarigou. 4CaaSt marketplace: an advanced business environment for trading cloud services. Future Generation Computer Systems, 41(??):104–120, December 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000454.

Margaris:2018:QPU

[MVG18] Dionisis Margaris, Costas Vassilakis, and Panagiotis Georgiadis. Query personalization using social network information and collaborative filtering techniques. Future Generation Computer Systems, 78 (part 1)(??):440–450, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17303692.

Manogaran:2018:NAI

[MVL^{+}18a] Gunasekaran Manogaran, R. Varatharajan, Daphne Lopez, Priyan Malarvizhi Kumar, Revathi Sundarasekar, and Chandu Thota. A new architecture of Internet of Things and big data ecosystem for secured smart healthcare monitoring and alerting system. Future Generation Computer Systems, 82(??):375–387, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17305149.

Mencagli:2018:HFM

[MVL18b] Gabriele Mencagli, Marco Vanneschi, and Silvia Lametti. The homeforwarding mechanism to reduce the cache coherence overhead in nextgeneration CMPs. Future Generation Computer Systems, 82(??):493–509, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17300286.

Mukherjee:2012:CDD

[MW12] A. Mukherjee and P. Watson. Case for dynamic deployment in a gridbased distributed query processor. Future Generation Computer Systems, 28(1):171–183, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001002.

Mazurczyk:2019:CAN

[MWCK19] Wojciech Mazurczyk, Steffen Wendzel, Mehdi Chourib, and Jörg Keller. Countering adaptive network covert communication with dynamic wardens. Future Generation Computer Systems, 94(??):712–725, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18316133.

Ma:2018:LHS

[MWL18a] Liwen Ma, Jiaji Wu, and Chunyuan Li. Localization of a high-speed train using a speed model based on the gradient descent algorithm.

Future Generation Computer Systems, 85(??):201– 209, August 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1732455X.

Mukherjee:2018:FIS

[MWL^{+}18b] Bidyut Mukherjee, Songjie Wang, Wenyi Lu, Roshan Lal Neupane, Daniel Dunn, Yijie Ren, Qi Su, and Prasad Calyam. Flexible IoT security middleware for end-to-end cloudfog communication. Future Generation Computer Systems, 87(??):688–703, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17311470.

Munawar:2010:DUP

[MWMA10] Asim Munawar, Mohamed Wahib, Masaharu Munetomo, and Kiyoshi Akama. The design, usage, and performance of GridUFO: a Grid based Unified Framework for Optimization. Future Generation Computer Systems, 26(4):633–644, April 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Munir:2012:CEA

[MWPVB12] Kashif Munir, Michael Welzl, Marcelo Pasin, and Pascale Primet Vicat-Blanc. Combining explicit admission control and congestion control for predictable data transfers in grids. Future Generation Computer Systems, 28(7): 1121–1132, July 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11002111.

Ma:2014:SEE

[MWQ^{+}14] Xingkong Ma, Yijie Wang, Qing Qiu, Weidong Sun, and Xiaoqiang Pei. Scalable and elastic event matching for attributebased publish/subscribe systems. Future Generation Computer Systems, 36(??):102–119, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001982.

Ma:2019:FBC

[MWQ^{+}19] Ruixin Ma, Kai Wang, Tie Qiu, Arun Kumar Sangaiah, Dan Lin, and Hannan Bin Liaqat. Featurebased compositing memory networks for aspect-based sentiment classification in Social Internet of Things.

Future Generation Computer Systems, 92(??):879– 888, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17313286.

Ma:2015:RSB

[MWW^{+}15] Yan Ma, Haiping Wu, Lizhe Wang, Bormin Huang, Rajiv Ranjan, Albert Zomaya, and Wei Jie. Remote sensing big data computing: Challenges and opportunities. Future Generation Computer Systems, 51(??):47–60, October 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002234.

Ma:2012:GRA

[MWYC12] Jianliang Ma, Chunhao Wang, Baozhong Yu, and Tianzhou Chen. Global register alias table: Boosting sequential program on multi-core. Future Generation Computer Systems, 28(6):957–964, June 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001890.

Mammeri:2018:PSE

[MYBMM18] Souhila Mammeri, Mohand Yazid, Louiza Bouallouche-Medjkoune, and Asma Mazouz. Performance study and enhancement of multichannel access methods in the future generation VHT WLAN. Future Generation Computer Systems, 79 (part 2)(??):543– 557, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630807X.

Ma:2018:RAC

[MYHZ18] Li Ma, Wenyin Yang, Yingyu Huo, and Yong Zhong. Research on access control model of social network based on distributed logic. Future Generation Computer Systems, 83(??):173–182, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17316515.

Moon:2016:AIS

[MYK16] Seonghoon Moon, Juwan Yoo, and Songkuk Kim. Adaptive interface selection over cloud-based splitlayer video streaming via multi-wireless networks. Future Generation Computer Systems, 56(??):664– 674, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003027.

Mao:2019:IIP

[MYW^{+}19] Bo Mao, Yaodong Yang, Suzhen Wu, Hong Jiang, and Kuan-Ching Li. IO-Follow: Improving the performance of VM live storage migration with IO following in the cloud. Future Generation Computer Systems, 91(??):167–176, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18314584.

Mateos:2010:EGE

[MZC10] Cristian Mateos, Alejandro Zunino, and Marcelo Campo. On the evaluation of gridification effort and runtime aspects of JGRIM applications. Future Generation Computer Systems, 26(6):797–819, June 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mekki:2016:UUD

[MZD^{+}16] Kais Mekki, Ahmed Zouinkhi, William Derigent, Eric Rondeau, André Thomas, and Mohamed Naceur Abdelkrim. USEE: a uniform data dissemination and energy efficient protocol for communicating materials. Future Generation Computer Systems, 56(??):651– 663, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002952.

Ma:2017:MVC

[MZH^{+}17] Hua Ma, Haibin Zhu, Zhigang Hu, Wensheng Tang, and Pingping Dong. Multivalued collaborative QoS prediction for cloud service via time series analysis. Future Generation Computer Systems, 68(??): 275–288, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303971.

Meng:2019:ESF

[MZL^{+}19] Weizhi Meng, Liqiu Zhu, Wenjuan Li, Jinguang Han, and Yan Li. Enhancing the security of FinTech applications with mapbased graphical password authentication. Future Generation Computer Systems, 101(??):1018–1027, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302882.

Ma:2019:MMA

[MZP^{+}19] Xiao Ma, Jiangfeng Zeng, Limei Peng, Giancarlo Fortino, and Yin Zhang. Modeling multi-aspects within one opinionated sentence simultaneously for aspect-level sentiment analysis. Future Generation Computer Systems, 93 (??):304–311, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

https://www.sciencedirect.
com/science/article/pii/
S0167739X18319125.

MingRu:2019:MIC

[MZYA19] Kong MingRu, Qin Zheng, Song Kui Yan, and N. Arunkumar. Medical image classification algorithm based on principal component feature dimensionality reduction. Future Generation Computer Systems, 98(??):627–634, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18327183.

Noury:2019:AIC

[NA19] Amir Noury and Morteza Amini. An access and inference control model for time series databases.

Future Generation Computer Systems, 92(??):93– 108, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18306757.

Noor:2019:MLB

[NAAC19] Umara Noor, Zahid Anwar, Tehmina Amjad, and Kim-Kwang Raymond Choo. A machine learning-based FinTech cyber threat attribution framework using high-level indicators of compromise. Future Generation Computer Systems, 96(??):227–242, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326141.

Nawaratne:2018:SEI

[NAD^{+}18] Rashmika Nawaratne, Damminda Alahakoon, Daswin De Silva, Prem Chhetri, and Naveen Chilamkurti. Self-evolving intelligent algorithms for facilitating data interoperability in IoT environments. Future Generation Computer Systems, 86(??):421–432, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324706.

Naghibzadeh:2016:MSH

[Nag16] Mahmoud Naghibzadeh.

Modeling and scheduling hybrid workflows of tasks and task interaction graphs on the cloud. Future Generation Computer Systems, 65(??):33–45, December 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301406.

Nafi:2018:SDN

[NAGD18] Nazmus S. Nafi, Khandakar Ahmed, Mark A. Gregory, and Manoj Datta. Software defined neighborhood area network for smart grid applications. Future Generation Computer Systems, 79 (part 2)(??):500–513, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17311007.

Noor:2019:MLF

[NAM^{+}19] Umara Noor, Zahid Anwar, Asad Waqar Malik, Sharifullah Khan, and Shahzad Saleem. A machine learning framework for investigating data breaches based on semantic analysis of adversary’s attack patterns in threat intelligence repositories. Future Generation Computer Systems, 95(??): 467–487, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18306708.

N:2018:IML

[NBB18] Sumith N., Annappa B., and Swapan Bhattacharya. Influence maximization in large social networks: Heuristics, models and parameters. Future Generation Computer Systems, 89(??):777–790, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18301092.

Nathani:2012:PBR

[NCS12] Amit Nathani, Sanjay Chaudhary, and Gaurav Somani. Policy based resource allocation in IaaS cloud. Future Generation Computer Systems, 28(1): 94–103, January 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11000987.

NguyenGia:2019:EEF

[NDA^{+}19] Tuan Nguyen Gia, Imed Ben Dhaou, Mai Ali, Amir M. Rahmani, Tomi Westerlund, Pasi Liljeberg, and Hannu Tenhunen. Energy efficient fog-assisted IoT system for monitoring diabetic patients with cardiovascular disease. Future Generation Computer Systems, 93(??):198–211, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18314365.

Navarro:2018:CFA

[NDZ^{+}18a] Javier Navarro, Faiyaz Doctor, Víctor Zamudio, Rahat Iqbal, Arun Kumar Sangaiah, and Carlos Lino. Corrigendum to “Fuzzy adaptive cognitive stimulation therapy generation for Alzheimer’s sufferers: Towards a pervasive dementia care monitoring platform” [future gener. comput. syst. 88 (2018) 479–490]. Future Generation Computer Systems, 88(??):730–731, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317746.

See [NDZ^{+}18b].

Navarro:2018:FAC

[NDZ^{+}18b] Javier Navarro, Faiyaz Doctor, Víctor Zamudio, Rahat Iqbal, Arun Kumar Sangaiah, and Carlos Lino. Fuzzy adaptive cognitive stimulation therapy generation for Alzheimer’s sufferers: Towards a pervasive dementia care monitoring platform. Future Generation Computer Systems, 88(??):479–490, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1830089X.

See corrigenda [NDZ^{+}18a, NDZ^{+}19].

Navarro:2019:CFA

[NDZ^{+}19] Javier Navarro, Faiyaz Doctor, Víctor Zamudio, Rahat Iqbal, Arun Kumar Sangaiah, and Carlos Lino. Corrigendum to “Fuzzy adaptive cognitive stimulation therapy generation for Alzheimer’s sufferers: Towards a pervasive dementia care monitoring platform” [Future Gener. Comput. Syst. 88 (2018) 479–490]. Future Generation Computer Systems, 93(??):1074–1075, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18320727.

See [NDZ^{+}18b].

Nadeem:2013:OET

[NF13] Farrukh Nadeem and Thomas Fahringer. Optimizing execution time predictions of scientific workflow applications in the Grid through evolutionary programming. Future Generation Computer Systems, 29(4):926–935, June 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001975.

Nemeth:2010:SSD

[NFK10] Zsolt Németh, Thomas Fahringer, and Péter Kacsuk. Special section: DAP-SYS, workshop on distributed and parallel systems. Future Generation Computer Systems, 26(3): 471–472, March 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nesa:2018:NPS

[NGB18] Nashreen Nesa, Tania Ghosh, and Indrajit Banerjee. Non-parametric sequence-based learning approach for outlier detection in IoT. Future Generation Computer Systems, 82(??):412–421, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17314474.

Nawaz:2019:PMS

[NHH^{+}19] Falak Nawaz, Omar Hussain, Farookh Khadeer Hussain, Naeem Khalid Janjua, Morteza Saberi, and Elizabeth Chang. Proactive management of SLA violations by capturing relevant external events in a Cloud of Things environment. Future Generation Computer Systems, 95 (??):26–44, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18318065.

Nogoorani:2016:TTD

[NJ16] Sadegh Dorri Nogoorani and Rasool Jalili. TIRIAC: a trust-driven risk-aware access control framework for grid environments. Future Generation Computer Systems, 55(??):238–254, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000588.

Nguyen:2017:RTE

[NJ17] Duc T. Nguyen and Jai E. Jung. Real-time event detection for online behavioral analysis of big social data. Future Generation Computer Systems, 66(??):137–145, January 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300899.

Nguyen:2018:SFU

[NJ18] Hoang Long Nguyen and Jai E. Jung. Socio-Scope: A framework for understanding Internet of Social Knowledge. Future Generation Computer Systems, 83(??):358–365, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17325402.

Nguyen:2019:SED

[NJ19] Hoang Long Nguyen and Jason J. Jung. Social event decomposition for constructing knowledge graph. Future Generation Computer Systems, 100(??):10– 18, November 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302493.

Najafi:2019:VRS

[NJB19] Aniseh Najafi, Hamid Haj Seyyed Javadi, and Majid Bayat. Verifiable ranked search over encrypted data with forward and backward privacy. Future Generation Computer Systems, 101(??):410–419, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18323318.

Nawaz:2018:EDA

[NJH^{+}18] Falak Nawaz, Naeem Khalid Janjua, Omar Khadeer Hussain, Farookh Khadeer Hussain, Elizabeth Chang, and Morteza Saberi. Eventdriven approach for predictive and proactive management of SLA violations in the Cloud of Things. Future Generation Computer Systems, 84(??):78– 97, July 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1732280X.

Nou:2011:PAS

[NJHT11] Ramon Nou, Ferran Julià, Kevin Hogan, and Jordi Torres. A path to achieving a self-managed Grid middleware. Future Generation Computer Systems, 27(1):10–19, January 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nickolay:2018:BGB

[NJKF18] Sam Nickolay, Eun-Sung Jung, Rajkumar Kettimuthu, and Ian Foster. Bridging the gap between peak and average loads on science networks. Future Generation Computer Systems, 79 (part 1)(??):169– 179, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17309834.

Njogu:2013:CVB

[NJKH13] Humphrey Waita Njogu, Luo Jiawei, Jane Nduta Kiere, and Damien Hanyurwimfura. A comprehensive vulnerability based alert management approach for large networks. Future Generation Computer Systems, 29(1):27–45, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1200091X.

Newman:2015:RAF

[NK15] Peter Newman and Gerald Kotonya. A resource-aware framework for resourceconstrained service-oriented systems. Future Generation Computer Systems, 47(??):161–175, June 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001800.

Nayahi:2017:PUP

[NK17] J. Jesu Vedha Nayahi and V. Kavitha. Privacy and utility preserving data clustering for data anonymization and distribution on Hadoop. Future Generation Computer Systems, 74(??):393–408, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16304459.

Nengroo:2018:MLB

[NK18] Ab Shaqoor Nengroo and K. S. Kuppusamy. Machine learning based heterogeneous web advertisements detection using a diverse feature set. Future Generation Computer Systems, 89(??):68–77, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17328777.

Nalepa:2019:MPA

[NKB19] Grzegorz J. Nalepa, Krzysztof Kutt, and Szymon Bobek. Mobile platform for affective context-aware systems. Future Generation Computer Systems, 92(??):490– 503, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312207.

Nguyen:2016:AMA

[NKP16] Tuan Anh Nguyen, Dong Seong Kim, and Jong Sou Park. Availability modeling and analysis of a data center for disaster tolerance. Future Generation Computer Systems, 56(??):27– 50, March 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002824.

Na:2019:CCV

[NLLC19] Woongsoo Na, Demeke Shumeye Lakew, Jihoon Lee, and Sungrae Cho. Congestion control vs. link failure: TCP behavior in mmWave connected vehicular networks. Future Generation Computer Systems, 101(??):1213–1222, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19307800.

Ning:2016:CCP

[NLM^{+}16] Huansheng Ning, Hong Liu, Jianhua Ma, Laurence T. Yang, and Runhe Huang. Cybermatics: Cyber-physical-social-thinking hyperspace based science and technology. Future Generation Computer Systems, 56(??):504–522, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002356.

Nkenyereye:2019:TSP

[NLS19] Lewis Nkenyereye, Chi Harold Liu, and JaeSeung Song. Towards secure and privacy preserving collision avoidance system in 5G fog based Internet of Vehicles. Future Generation Computer Systems, 95(??): 488–499, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317588.

Nouri:2019:ADE

[NLV^{+}19] Seyed Mohammad Reza Nouri, Han Li, Srikumar Venugopal, Wenxia Guo, MingYun He, and Wenhong Tian. Autonomic decentralized elasticity based on a reinforcement learning controller for cloud applications. Future Generation Computer Systems, 94(??): 765–780, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18302826.

Neupane:2019:IDU

[NNC^{+}19] Roshan Lal Neupane, Travis Neely, Prasad Calyam, Nishant Chettri, Mark Vassell, and Ramakrishnan Durairajan. Intelligent defense using pretense against targeted attacks in cloud platforms. Future Generation Computer Systems, 93(??):609– 626, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18307805.

Niemi:2018:TGB

[NNLH18] Tapio Niemi, Jukka K.

Nurminen, Juha-Matti Liukkonen, and Ari-Pekka Hameri. Towards green big data at CERN. Future Generation Computer Systems, 81(??):103–113, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17313651.

Naik:2019:DLB

[NNRA19] Nenavath Srinivas Naik, Atul Negi, Tapas Bapu B. R., and R. Anitha. A data locality based scheduler to enhance MapReduce performance in heterogeneous environments. Future Generation Computer Systems, 90(??):423– 434, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18308379.

Nolasco:2019:SDT

[NO19] Diogo Nolasco and Jonice Oliveira. Subevents detection through topic modeling in social media posts. Future Generation Computer Systems, 93(??):290– 303, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18307611.

Nonaka:2018:FPI

[NOF18] Jorji Nonaka, Kenji Ono, and Masahiro Fujita. 234Compositor: A flexible parallel image compositing framework for massively parallel visualization environments. Future Generation Computer Systems, 82(??):647–655, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17302030.

Nalepa:2019:ACA

[NPH19] Grzegorz J. Nalepa, José Palma, and María Trinidad Herrero. Affective computing in ambient intelligence systems. Future Generation Computer Systems, 92(??):454–457, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18328528.

Nayak:2012:NAD

[NPP12] Sasmita Kumari Nayak, Sasmita Kumari Padhy, and Siba Prasada Panigrahi. A novel algorithm for dynamic task scheduling. Future Generation Computer Systems, 28(5): 709–717, May 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002354.

Nijim:2013:AEC

[NQQL13] Mais Nijim, Xiao Qin, Meikang Qiu, and Kenli Li. An adaptive energyconserving strategy for parallel disk systems. Future Generation Computer Systems, 29(1):196–207, January 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000994.

Nam:2015:MCR

[NRR^{+}15] Sungwon Nam, Khairi Reda, Luc Renambot, Andrew Johnson, and Jason Leigh. Multiusercentered resource scheduling for collaborative display wall environments. Future Generation Computer Systems, 45(??):162– 175, April 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001605.

Namasudra:2017:TES

[NRV^{+}17] Suyel Namasudra, Pinki Roy, Pandi Vijayakumar, Sivaraman Audithan, and Balamurugan Balusamy. Time efficient secure DNA based access control model for cloud computing environment. Future Generation Computer Systems, 73(??):90–105, August 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17300572.

Nekovee:2010:SLS

[NS10] Maziar Nekovee and Radhika S. Saksena. Simulations of large-scale WiFibased wireless networks:

Interdisciplinary challenges and applications. Future Generation Computer Systems, 26(3):514–520, March 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nabti:2017:QMG

[NS17a] Chemseddine Nabti and Hamida Seba. Querying massive graph data: a compress and search approach. Future Generation Computer Systems, 74(??):63– 75, September 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17305435.

Neves:2017:PST

[NS17b] Diogo Telmo Neves and João Luís Sobral. Parallel SuperFine — a tool for fast and accurate supertree estimation: Features and limitations. Future Generation Computer Systems, 67(??):441–454, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300814.

Nie:2019:KGE

[NS19] Binling Nie and Shouqian Sun. Knowledge graph embedding via reasoning over entities, relations, and text. Future Generation Computer Systems, 91(??):426–433, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17321593.

N:2019:MBA

[NSR^{+}19] Muthukumar N., Seshadhri Srinivasan, K. Ramkumar, Deepak Pal, Juri Vain, and Srini Ramaswamy. A model-based approach for design and verification of Industrial Internet of Things. Future Generation Computer Systems, 95(??): 354–363, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18321137.

Niewiadomska-Szynkiewicz:2014:DPM

[NSSA^{+}14] Ewa Niewiadomska-Szynkiewicz, Andrzej Sikora, Piotr Arabas, Mariusz Kamola, Marcin Mincer, and Joanna Kolodziej. Dynamic power management in energyaware computer networks and data intensive computing systems. Future Generation Computer Systems, 37(??):284–296, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002124.

Nishad:2019:ACH

[NUPA19] Anurag Nishad, Abhay Upadhyay, Ram Bilas Pachori, and U. Rajendra Acharya. Automated classification of hand movements using tunable-Q wavelet transform based filter-bank with surface electromyogram signals. Future Generation Computer Systems, 93(??):96– 110, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18316492.

Nagarajan:2011:DTE

[NV11] Aarthi Nagarajan and Vijay Varadharajan. Dynamic trust enhanced security model for trusted platform based services. Future Generation Computer Systems, 27(5):564–573, May 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Naserian:2018:PLP

[NWD^{+}18] Elahe Naserian, Xinheng Wang, Keshav Dahal, Zhi Wang, and Zaijian Wang. Personalized location prediction for group travellers from spatialtemporal trajectories. Future Generation Computer Systems, 83(??):278–292, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17318812.

Ni:2017:MEI

[NWL17] Wei Ni, Weigang Wu, and Keqin Li. A message efficient intersection control algorithm for intelligent transportation in smart cities. Future Generation Computer Systems, 76(??):339–349, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16304794.

Nedjah:2017:EYR

[NWMG17] N. Nedjah, R. S. Wyant, L. M. Mourelle, and B. B. Gupta. Efficient yet robust biometric iris matching on smart cards for data high security and privacy. Future Generation Computer Systems, 76(??):18– 32, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301085.

Negash:2019:TII

[NWT19] Behailu Negash, Tomi Westerlund, and Hannu Tenhunen. Towards an interoperable Internet of Things through a web of virtual things at the fog layer. Future Generation Computer Systems, 91(??):96–107, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18308859.

Niu:2015:NAS

[NZL^{+}15] Ben Niu, Xiaoyan Zhu, Qinghua Li, Jie Chen, and Hui Li. A novel attack to spatial cloaking schemes in location-based services. Future Generation Computer Systems, 49(??):125– 132, August 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002209.

Naranjo-Zolotov:2019:ESC

[NZOCJ^{+}19] Mijail Naranjo-Zolotov, Tiago Oliveira, Frederico Cruz-Jesus, José Martins, Ramiro Gonçalves, Frederico Branco, and Nuno Xavier. Examining social capital and individual motivators to explain the adoption of online citizen participation. Future Generation Computer Systems, 92(??):302–311, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18313396.

Orojloo:2017:MEC

[OA17] Hamed Orojloo and Mohammad Abdollahi Azgomi. A method for evaluating the consequence propagation of security attacks in cyber-physical systems. Future Generation Computer Systems, 67(??): 57–71, February 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302679.

Oguego:2018:UAM

[OAMS18] C. L. Oguego, J. C. Augusto, A. Muñoz, and M. Springett. Using argumentation to manage users’ preferences. Future Generation Computer Systems, 81(??):235–243, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630379X.

Ogiela:2017:NPI

[OB17] Marek R. Ogiela and Leonard Barolli. New paradigms for information and services management in grid and pervasive computing. Future Generation Computer Systems, 67(??):227–229, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16304095.

Ozsoydan:2019:SIB

[OB19] Fehmi Burcin Ozsoydan and Adil Baykasoglu. A swarm intelligence-based algorithm for the set-union knapsack problem. Future Generation Computer Systems, 93(??):560–569, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18300979.

Oyarzun:2018:ECC

[OBG^{+}18] G. Oyarzun, R. Borrell, A. Gorobets, F. Mantovani, and A. Oliva. Efficient CFD code implementation for the ARMbased Mont-Blanc architecture. Future Generation Computer Systems, 79 (part 3)(??):786–796, February 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17305733.

Ozkasap:2014:FHE

[OCCK14] Oznur Ozkasap, Emrah Cem, Sena Efsun Cebeci, and Tugba Koc. Flat and hierarchical epidemics in P2P systems: Energy cost models and analysis. Future Generation Computer Systems, 36(??):257–266, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1300188X.

Okeyo:2014:COT

[OCW14] George Okeyo, Liming Chen, and Hui Wang. Combining ontological and temporal formalisms for composite activity modelling and recognition in smart homes. Future Generation Computer Systems, 39(??):29–43, October 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000399.

Onorati:2019:SNE

[ODC19] Teresa Onorati, Paloma Díaz, and Belen Carrion. From social networks to emergency operation centers: a semantic visualization approach. Future Generation Computer Systems, 95(??):829–840, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17309354.

Ochoa:2014:DSU

[OdI14] Sergio F. Ochoa and Diego López de Ipiña. Distributed solutions for ubiquitous computing and ambient intelligence. Future Generation Computer Systems, 34(??):94–96, May 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000090.

Odelu:2017:PSA

[ODK^{+}17] Vanga Odelu, Ashok Kumar Das, Saru Kumari, Xinyi Huang, and Mohammad Wazid. Provably secure authenticated key agreement scheme for distributed mobile cloud computing services. Future Generation Computer Systems, 68(??):74–88, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303065.

Ocana:2013:DPC

[OdOD^{+}13] Kary A. C. S. Ocaña, Daniel de Oliveira, Jonas Dias, Eduardo Ogasawara, and Marta Mattoso. Designing a parallel cloud based comparative genomics workflow to improve phylogenetic analyses. Future Generation Computer Systems, 29(8):2205–2219, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000654.

Ostberg:2013:GCS

[ÖE13] Per-Olov Östberg and Erik Elmroth. GJMF

— a composable serviceoriented grid job management framework. Future Generation Computer Systems, 29(1):144–157, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000945.

Ostberg:2013:DSF

[ÖEE13] Per-Olov Östberg, Daniel Espling, and Erik Elmroth. Decentralized scalable fairshare scheduling. Future Generation Computer Systems, 29(1):130–143, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001355.

Ochoa:2017:CPS

[OFD17] Sergio F. Ochoa, Giancarlo Fortino, and Giuseppe Di Fatta. Cyber-physical systems, Internet of Things and big data. Future Generation Computer Systems, 75(??):82–84, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17311196.

Olaniyan:2018:OEC

[OFMZ18] Richard Olaniyan, Olamilekan Fadahunsi, Muthucumaru Maheswaran, and Mohamed Faten Zhani. Opportunistic edge computing: Concepts, opportunities and research challenges. Future Generation Computer Systems, 89(??):633–645, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18303388.

OLoughlin:2018:PBH

[OG18] John O’Loughlin and Lee Gillam. A performance brokerage for heterogeneous clouds. Future Generation Computer Systems, 87(??):831–845, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17309251.

Oh:2010:RTP

[OKF10] Sangyoon Oh, Jai-Hoon Kim, and Geoffrey Fox. Real-time performance analysis for publish/subscribe systems. Future Generation Computer Systems, 26 (3):318–323, March 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Olabarriaga:2010:SSM

[OM10] Sílvia D. Olabarriaga and Johan Montagnat. Special section: Medical imaging on grids. Future Generation Computer Systems, 26(4):678–680, April 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Orujov:2018:SBI

[OMD^{+}18] F. Orujov, R. Maskeliunas, R. Damasevicius, Wei Wei, and Ye Li. Smartphone based intelligent indoor positioning using fuzzy logic. Future Generation Computer Systems, 89(??):335– 348, December 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18305004.

Olaverri-Monreal:2019:CAS

[OMKM^{+}19] Cristina Olaverri-Monreal, Gerd Ch. Krizek, Florian Michaeler, Rene Lorenz, and Matthias Pichler. Collaborative approach for a safe driving distance using stereoscopic image processing. Future Generation Computer Systems, 95(??): 880–889, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17307124.

Ortiz-Martin:2019:FAI

[OMPSPL^{+}19] Lara Ortiz-Martin, Pablo Picazo-Sanchez, Pedro Peris-Lopez, Juan Tapiador, and Gerardo Schneider. Feasibility analysis of inter-pulse intervals based solutions for cryptographic token generation by two electrocardiogram sensors. Future Generation Computer Systems, 96(??):283– 296, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18330784.

Olanda:2013:HPS

[OPO13] Ricardo Olanda, Mariano Pérez, and Juan Manuel Orduña. Hybrid P2P schemes for remote terrain interactive visualization systems. Future Generation Computer Systems, 29(6):1522–1532, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12002087.

Oyekan:2017:RRT

[OPT^{+}17] John Oyekan, Vinayak Prabhu, Ashutosh Tiwari, Vinubha Baskaran, Mark Burgess, and Rob Mcnally. Remote real-time collaboration through synchronous exchange of digitised human-workpiece interactions. Future Generation Computer Systems, 67(??):83–93, February 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302692.

Ostad-Sharif:2019:TPS

[OSANAM19] Arezou Ostad-Sharif, Hamed Arshad, Morteza Nikooghadam, and Dariush Abbasinezhad-Mood. Three party secure data transmission in IoT networks through design of a lightweight authenticated key agreement scheme. Future Generation Computer Systems, 100(??):882–892, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18309452.

Ozsoy:2014:OLC

[OSC14] Adnan Ozsoy, Martin Swany, and Arun Chauhan. Optimizing LZSS compression on GPGPUs. Future Generation Computer Systems, 30(??):170–178, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001362.

Osman:2019:NBD

[Osm19] Ahmed M. Shahat Osman.

A novel big data analytics framework for smart cities. Future Generation Computer Systems, 91(??):620–633, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17307446.

Ouyang:2019:MSA

[OWX19] Xue Ouyang, Changjian Wang, and Jie Xu. Mitigating stragglers to avoid QoS violation for time-critical applications through dynamic server blacklisting. Future Generation Computer Systems, 101(??):831–842, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18311397.

Peternier:2014:IEU

[PAB^{+}14] Achille Peternier, Danilo Ansaloni, Daniele Bonetta, Cesare Pautasso, and Walter Binder. Improving execution unit occupancy on SMT-based processors through hardware-aware thread scheduling. Future Generation Computer Systems, 30(??):229–241, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001295.

Pradal:2017:ISW

[PAC^{+}17] Christophe Pradal, Simon Artzet, Jérôme Chopard, Dimitri Dupuis, Christian Fournier, Michael Mielewczik, Vincent Nègre, Pascal Neveu, Didier Parigot, Patrick Valduriez, and Sarah Cohen-Boulakia. InfraPhenoGrid: a scientific workflow infrastructure for plant phenomics on the grid. Future Generation Computer Systems, 67(??):341–353, February 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301820.

Palmieri:2013:SSD

[Pal13] Francesco Palmieri. Scalable service discovery in ubiquitous and pervasive computing architectures: a percolation-driven approach. Future Generation Computer Systems, 29(3): 693–703, March 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001604.

Pallickara:2016:SRA

[Pal16] Shrideep Pallickara. Some recent advances in utility and cloud computing. Future Generation Computer Systems, 56(??):315–316, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003714.

Pearce:2019:EDL

[PAL^{+}19] Olga Pearce, Hadia Ahmed, Rasmus W. Larsen, Peter Pirkelbauer, and David F. Richards. Exploring dynamic load imbalance solutions with the CoMD proxy application. Future Generation Computer Systems, 92(??):920–932, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17300560.

Parra-Arnau:2014:MPU

[PARMF14] Javier Parra-Arnau, David Rebollo-Monedero, and Jordi Forné. Measuring the privacy of user profiles in personalized information systems. Future Generation Computer Systems, 33(??):53–63, April 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1300006X.

Pedersen:2017:LSB

[PB17] Edvard Pedersen and Lars Ailo Bongo. Largescale biological metadatabase management. Future Generation Computer Systems, 67(??):481–489, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300231.

Ponciano:2018:ABC

[PB18] Lesandro Ponciano and Francisco Brasileiro. Agreementbased credibility assessment and task replication in human computation systems. Future Generation Computer Systems, 87(??):159–170, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17303618.

Piccialli:2018:ISC

[PBA18] Francesco Piccialli, Paolo Benedusi, and Flora Amato. S-InTime: a social cloud analytical service oriented system. Future Generation Computer Systems, 80(??):229–241, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16307397.

Perez:2011:SBA

[PBC^{+}11] Juan M. Marín Pérez, Jorge Bernal Bernabé, Jose M. Alcaraz Calero, Felix J. Garcia Clemente, Gregorio Martínez Pérez, and Antonio F. Gómez Skarmeta. Semantic-based authorization architecture for Grid. Future Generation Computer Systems, 27(1):40–55, January 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Podzimek:2016:RPL

[PBC^{+}16] Andrej Podzimek, Lubomír Bulej, Lydia Y. Chen, Walter Binder, and Petr Tuma. Robust partial-load experiments with Showstopper. Future Generation Computer Systems, 64(??):15– 38, November 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301078.

Podzimek:2017:RRP

[PBC^{+}17] Andrej Podzimek, Lubomír Bulej, Lydia Y. Chen, Walter Binder, and Petr Tuma. Reprint of “Robust partial-load experiments with Showstopper”. Future Generation Computer Systems, 72(??):81– 104, July 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306331.

Peng:2018:ITR

[PBL^{+}18] Hao Peng, Mengjiao Bao, Jianxin Li, Md Zakirul Alam Bhuiyan, Yaopeng Liu, Yu He, and Erica Yang. Incremental term representation learning for social network analysis. Future Generation Computer Systems, 86(??):1503–1512, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17310117.

Palmieri:2013:DSF

[PBV^{+}13] Francesco Palmieri, Luigi Buonanno, Salvatore Venticinque, Rocco Aversa, and Beniamino Di Martino. A distributed scheduling framework based on selfish autonomous agents for federated cloud environments. Future Generation Computer Systems, 29(6):1461–1472, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000289.

Peixoto:2017:WVS

[PC17] João Paulo Just Peixoto and Daniel G. Costa. Wireless visual sensor networks for smart city applications:a relevance-based approach for multiple sinks mobility. Future Generation Computer Systems, 76(??):51–62, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17309652.

Pan:2018:SAM

[PC18a] Weifeng Pan and Chunlai Chai. Structure-aware mashup service clustering for cloud-based Internet of Things using genetic algorithm based clustering algorithm. Future Generation Computer Systems, 87(??):267–277, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18302097.

Piccialli:2018:EFS

[PC18b] Francesco Piccialli and Angelo Chianese. Editorial for FGCS special issue: the Internet of Cultural Things: Towards a smart cultural heritage. Future Generation Computer Systems, 81(??):514–515, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1732873X.

Peng:2018:UGM

[PCC18] Weimin Peng, Aihong Chen, and Jing Chen. Using general master equation for feature fusion. Future Generation Computer Systems, 82(??):119–126, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17303412.

Pumma:2017:REF

[PcFP^{+}17] Sarunya Pumma, Wu chun Feng, Phond Phunchongharn, Sylvain Chapeland, and Tiranee Achalakul. A runtime estimation framework for ALICE. Future Generation Computer Systems, 72(??):65–77, July 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302972.

Perkovic:2019:LVL

[PCK19] Toni Perković, Mario Cagalj, and Tonko Kovacević. LISA: Visible light based initialization and SMS based authentication of constrained IoT devices. Future Generation Computer Systems, 97(??):105– 118, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18321083.

Pham:2011:HTF

[PD11] Van-Hau Pham and Marc Dacier. Honeypot trace forensics: The observation viewpoint matters. Future Generation Computer Systems, 27(5):539–546, May 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Paton:2012:UDA

[PdAF12] Norman W. Paton, Marcelo A. T. de Aragão, and Alvaro A. A. A. Fernandes. Utility-driven adaptive query workload execution. Future Generation Computer Systems, 28(7): 1070–1079, July 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11002123.

Panadero:2018:MCB

[PdASM18] Javier Panadero, Jesica de Armas, Xavier Serra, and Joan Manuel Marquès. Multi criteria biased randomized method for resource allocation in distributed systems: Application in a volunteer computing system. Future Generation Computer Systems, 82(??):29–40, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17315236.

Pucciani:2010:PSS

[PDDS10] Gianni Pucciani, Andrea Domenici, Flavia Donno, and Heinz Stockinger. A performance study on the synchronisation of heterogeneous Grid databases using CONStanza. Future Generation Computer Systems, 26(6):820–834, June 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Peng:2018:TIC

[PDH18] Limei Peng, Ahmad R.

Dhaini, and Pin-Han Ho. Toward integrated cloudfog networks for efficient IoT provisioning: Key challenges and solutions. Future Generation Computer Systems, 88(??):606– 613, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1830596X.

Ponto:2010:GSM

[PDK10] Kevin Ponto, Kai Doerr, and Falko Kuester. Gigastack: a method for visualizing giga-pixel layered imagery on massively tiled displays. Future Generation Computer Systems, 26(5):693–700, May 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ponto:2011:CMU

[PDW^{+}11] Kevin Ponto, Kai Doerr, Tom Wypych, John Kooker, and Falko Kuester. CGLXTouch: a multi-user multi-touch approach for ultra-high-resolution collaborative workspaces. Future Generation Computer Systems, 27(6):649–656, June 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pereira:2019:PCG

[PECA19] Adrián Arenal Pereira, Jordán Pascual Espada, Rubén González Crespo, and Sergio Ríos Aguilar. Platform for controlling and getting data from network connected drones in indoor environments. Future Generation Computer Systems, 92(??):656–662, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17317363.

Poonpakdee:2017:REM

[PF17] Pasu Poonpakdee and Giuseppe Di Fatta. Robust and efficient membership management in largescale dynamic networks. Future Generation Computer Systems, 75(??):85– 93, October 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302649.

Paulraj:2018:RAV

[PFPJ18] Getzi Jeba Leelipushpam Paulraj, Sharmila Anand John Francis, J. Dinesh Peter, and Immanuel Johnraja Jebadurai. Resource-aware virtual machine migration in IoT cloud. Future Generation Computer Systems, 85(??):173–183, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17322471.

Palmieri:2016:GBR

[PFRC16] Francesco Palmieri, Ugo Fiore, Sergio Ricciardi, and Aniello Castiglione. GRASP-based resource reoptimization for effective big data access in federated clouds. Future Generation Computer Systems, 54(??):168–179, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000345.

Price:2013:SWS

[PFS^{+}13] Simon Price, Peter A.

Flach, Sebastian Spiegler, Christopher Bailey, and Nikki Rogers. SubSift Web services and workflows for profiling and comparing scientists and their published works. Future Generation Computer Systems, 29(2):569–581, February 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11002238.

Perez:2010:BRS

[PGCC^{+}10] José M. Pérez, Félix García-Carballeira, Jesús Carretero, Alejandro Calderón, and Javier Fernández. Branch replication scheme: a new model for data replication in large scale data grids. Future Generation Computer Systems, 26(1): 12–20, January 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Perez:2019:ASA

[PGCML^{+}19] Salvador Pérez, Dan Garcia-Carrillo, Rafael Marín-López, José L. Hernández-Ramos, Rafael Marín-Pérez, and Antonio F. Skarmeta. Architecture of security association establishment based on bootstrapping technologies for enabling secure IoT infrastructures. Future Generation Computer Systems, 95(??):570–585, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18325573.

Perez:2018:RDN

[PGTBC18] Juan Luis Pérez, Alberto Gutierrez-Torre, Josep Ll. Berral, and David Carrera. A resilient and distributed near real-time traffic forecasting application for fog computing environments. Future Generation Computer Systems, 87(??):198–212, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1732678X.

Pipan:2010:UTO

[Pip10] Gregor Pipan. Use of the TRIPOD overlay network for resource discovery. Future Generation Computer Systems, 26(8): 1257–1270, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Poenaru:2018:AAF

[PIP18a] Andrei Poenaru, Roxana Istrate, and Florin Pop. AFT: Adaptive and fault tolerant peer-to-peer overlay — a user-centric solution for data sharing. Future Generation Computer Systems, 80(??):583–595, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301339.

Pop:2018:HHH

[PIP18b] Florin Pop, Alexandru Iosup, and Radu Prodan. HPS–HDS: High performance scheduling for heterogeneous distributed systems. Future Generation Computer Systems, 78 (part 1)(??):242–244, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17319659.

Pamies-Juarez:2013:NRG

[PJDO13] Lluis Pamies-Juarez, Anwitaman Datta, and Frédérique Oggier. In-network redundancy generation for opportunistic speedup of data backup. Future Generation Computer Systems, 29(6):1353–1362, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000460.

Papazachos:2011:GSM

[PK11] Zafeirios C. Papazachos and Helen D. Karatza. Gang scheduling in multicore clusters implementing migrations. Future Generation Computer Systems, 27(8):1153–1165, October 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Przybylo:2019:EBA

[PKA19] Jaromir Przybylo, Eliasz Kańtoch, and Piotr Augustyniak. Eyetrackingbased assessment of affectrelated decay of human performance in visual tasks. Future Generation Computer Systems, 92(??):504–515, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312001.

Pourmasoumi:2019:ECD

[PKB19] Asef Pourmasoumi, Mohsen Kahani, and Ebrahim Bagheri. The evolutionary composition of desirable execution traces from event logs. Future Generation Computer Systems, 98(??):78–103, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17328583.

Park:2014:ERT

[PKF14] Sanghyuk Park, Jai-Hoon Kim, and Geoffrey Fox. Effective real-time scheduling algorithm for cyber physical systems society. Future Generation Computer Systems, 32(??):253–259, March 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002136.

Prasad:2018:RPA

[PKI^{+}18] Abhinandan S. Prasad, David Koll, Jesus Omana Iglesias, Jordi Arjona Aroca, Volker Hilt, and Xiaoming Fu. RConf(PD): Automated resource configuration of complex services in the cloud. Future Generation Computer Systems, 87(??):639–650, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17314231.

Pramanik:2019:EDT

[PKP19] Himadri Sikhar Pramanik, Manish Kirtania, and Ashis K. Pani. Essence of digital transformation — manifestations at large financial institutions from North America. Future Generation Computer Systems, 95(??):323–343, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18308951.

Peddi:2017:ICB

[PKY^{+}17] Sri Vijay Bharat Peddi, Pallavi Kuhad, Abdulsalam Yassine, Parisa Pouladzadeh, Shervin Shirmohammadi, and Ali Asghar Nazari Shirehjini. An intelligent cloud-based data processing broker for mobile e-health multimedia applications. Future Generation Computer Systems, 66(??):71–86, January 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300632.

Panneerselvam:2018:IRF

[PLA18] John Panneerselvam, Lu Liu, and Nick Antonopoulos. InOt–RePCoN: Forecasting user behavioural trend in large-scale cloud environments. Future Generation Computer Systems, 80(??):322–341, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310130.

Perez:2011:FDS

[PLCGS11] Alejandro Pérez, Gabriel López, Óscar Cánovas, and Antonio F. Gómez-Skarmeta. Formal description of the SWIFT identity management framework. Future Generation Computer Systems, 27(8): 1113–1123, October 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Peris-Lopez:2018:EAC

[PLGMCdF18] Pedro Peris-Lopez, Lorena González-Manzano, Carmen Camara, and José María de Fuentes. Effect of attacker characterization in ECG-based continuous authentication mechanisms for Internet of Things. Future Generation Computer Systems, 81(??):67– 77, April 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17300407.

Pan:2018:ASJ

[PLL^{+}18] Weifeng Pan, Bing Li, Jing Liu, Yutao Ma, and Bo Hu. Analyzing the structure of Java software systems by weighted Kcore decomposition. Future Generation Computer Systems, 83(??):431–444, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17320940.

Panneerselvam:2018:IIT

[PLLA18] John Panneerselvam, Lu Liu, Yao Lu, and Nick Antonopoulos. An investigation into the impacts of task-level behavioural heterogeneity upon energy efficiency in cloud datacentres. Future Generation Computer Systems, 83(??):239–249, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1731960X.

Palau:2019:CPS

[PLLP19] Adrià Salvador Palau, Zhenglin Liang, Daniel Lütgehetmann, and Ajith Kumar Parlikad. Collaborative prognostics in social asset networks. Future Generation Computer Systems, 92(??):987–995, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17308105.

Peng:2019:EEC

[PLP^{+}19] Yuyang Peng, Jun Li, Sangdon Park, Konglin Zhu, Mohammad Mehedi Hassan, and Ahmed Alsanad. Energy-efficient cooperative transmission for intelligent transportation systems. Future Generation Computer Systems, 94(??):634–640, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18321319.

Peng:2019:MPP

[PLW^{+}19] Tao Peng, Qin Liu, Guojun Wang, Yang Xiang, and Shuhong Chen. Multidimensional privacy preservation in locationbased services. Future Generation Computer Systems, 93(??):312–326, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1732160X.

Pan:2019:LAF

[PLZX19] Yubiao Pan, Yongkun Li, Huizhen Zhang, and Yinlong Xu. Lifetimeaware FTL to improve the lifetime and performance of solid-state drives. Future Generation Computer Systems, 93(??):58– 67, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18300566.

Punithan:2014:PTT

[PM14] Dharani Punithan and R. I. (Bob) McKay. Phase transitions in two-dimensional daisyworld with smallworld effects — a study of local and long-range couplings. Future Generation Computer Systems, 33(??): 64–80, April 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000472.

Perl:2014:PPT

[PMBS14] H. Perl, Y. Mohammed, M. Brenner, and M. Smith. Privacy/performance tradeoff in private search on bio-medical data. Future Generation Computer Systems, 36(??):441–452, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002707.

Perez:2018:SCC

[PMCC18] Alfonso Pérez, Germán Moltó, Miguel Caballer, and Amanda Calatrava. Serverless computing for container-based architectures. Future Generation Computer Systems, 83(??): 50–59, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17316485.

Pham:2018:DHH

[PMDS18] Minh Pham, Yehenew Mengistu, Ha Do, and Weihua Sheng. Delivering home healthcare through a Cloud-based Smart Home Environment (CoSHE). Future Generation Computer Systems, 81(??):129– 140, April 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302194.

Pimenta:2011:CMN

[PMFH11] Marcelo Pimenta, Evandro Miletto, Luciano Flores, and Aurelio Hoppe. Cooperative mechanisms for networked music. Future Generation Computer Systems, 27(1):100–108, January 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Papamartzivanos:2018:DGT

[PMK18] Dimitrios Papamartzivanos, Félix Gómez Mármol, and Georgios Kambourakis. Dendron: Genetic trees driven rule induction for network intrusion detection systems. Future Generation Computer Systems, 79 (part 2)(??):558– 574, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X16305465.

Povedano-Molina:2013:DHA

[PMLVLS^{+}13] Javier Povedano-Molina, Jose M. Lopez-Vega, Juan M. Lopez-Soler, Antonio Corradi, and Luca Foschini. DARGOS: a highly adaptable and scalable monitoring architecture for multitenant clouds. Future Generation Computer Systems, 29(8):2041–2056, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000824.

Perez-Miguel:2013:HTC

[PMMAM13] Carlos Pérez-Miguel, Jose Miguel-Alonso, and Alexander Mendiburu. High throughput computing over peer-to-peer networks. Future Generation Computer Systems, 29(1):352–360, January 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001506.

Petcu:2013:PCA

[PMPC13] Dana Petcu, Georgiana Macariu, Silviu Panica, and Ciprian Craciun. Portable Cloud applications — from theory to practice. Future Generation Computer Systems, 29(6):1417–1430, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000210.

Pirro:2010:FDK

[PMT10] Giuseppe Pirró, Carlo Mastroianni, and Domenico Talia. A framework for distributed knowledge management: Design and implementation. Future Generation Computer Systems, 26(1):38–49, January 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pandey:2013:CCS

[PN13] Suraj Pandey and Surya Nepal. Cloud computing and scientific applications

— big data, scalable analytics, and beyond. Future Generation Computer Systems, 29(7):1774–1776, September 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000861.

Palanca:2013:DPS

[PNGFJ13] Javier Palanca, Marti Navarro, Ana García-Fornes, and Vicente Julian. Deadline prediction scheduling based on benefits. Future Generation Computer Systems, 29(1): 61–73, January 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001033.

Pervaiz:2014:UAQ

[PNZ14] Haris Pervaiz, Qiang Ni, and Charilaos C. Zarakovitis. User adaptive QoS aware selection method for cooperative heterogeneous wireless systems: a dynamic contextual approach. Future Generation Computer Systems, 39(??): 75–87, October 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000375.

Piechotta:2016:SDC

[POJ^{+}16] Chris Piechotta, Martin Grooss Olsen, Adam Enø Jensen, Joey W. Coleman, and Peter Gorm Larsen. A secure dynamic collaboration environment in a cloud context. Future Generation Computer Systems, 55(??):165–175, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002502.

Ponraj:2019:OVM

[Pon19] Anitha Ponraj. Optimistic virtual machine placement in cloud data centers using queuing approach. Future Generation Computer Systems, 93(??):338–344, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18304692.

Palmieri:2010:TFM

[PP10] Francesco Palmieri and Silvio Pardi. Towards a federated Metropolitan Area Grid environment: The SCoPE network-aware infrastructure. Future Generation Computer Systems, 26(8):1241–1256, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pop:2018:RBR

[PPA18] Florin Pop, Radu Prodan, and Gabriel Antoniu. RM-BDP: Resource management for big data platforms. Future Generation Computer Systems, 86(??):961–963, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18311245.

Pop:2016:AAT

[PPB16] Florin Pop and Maria Potop-Butucaru. ARMCO: Advanced topics in resource management for ubiquitous cloud computing: an adaptive approach. Future Generation Computer Systems, 54(??):79– 81, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002484.

Pennino:2019:PIS

[PPG19] Diego Pennino, Maurizio Pizzonia, and Federico Griscioli. Pipelineintegrity: Scaling the use of authenticated data structures up to the cloud. Future Generation Computer Systems, 100(??):618–647, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18328048.

Pittaras:2015:RDA

[PPL^{+}15] C. Pittaras, C. Papagianni, A. Leivadeas, P. Grosso, J. van der Ham, and S. Papavassiliou. Resource discovery and allocation for federated virtualized infrastructures. Future Generation Computer Systems, 42(??):55–63, January 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000089.

Pouya:2017:CHD

[PPLL17] Iman Pouya, Sander Pronk, Magnus Lundborg, and Erik Lindahl. Copernicus, a hybrid dataflow and peerto-peer scientific computing platform for efficient large-scale ensemble sampling. Future Generation Computer Systems, 71(??): 18–31, June 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16305337.

Perles:2018:EEI

[PPMM^{+}18] Angel Perles, Eva Pérez-Marín, Ricardo Mercado, J. Damian Segrelles, Ignacio Blanquer, Manuel Zarzo, and Fernando J. Garcia-Diego. An energyefficient Internet of Things (IoT) architecture for preventive conservation of cultural heritage. Future Generation Computer Systems, 81(??):566–581, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17313663.

Persico:2018:BBD

[PPPS18] Valerio Persico, Antonio Pescapé, Antonio Picariello, and Giancarlo Sperlí. Benchmarking big data architectures for social networks data processing using public cloud platforms. Future Generation Computer Systems, 89(??):98–109, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17328303.

Plageras:2018:EIB

[PPS^{+}18] Andreas P. Plageras, Kostas E.

Psannis, Christos Stergiou, Haoxiang Wang, and B. B. Gupta. Efficient IoT-based sensor BIG data collectionprocessing and analysis in smart buildings. Future Generation Computer Systems, 82(??):349–357, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17314127.

Papangelis:2019:ATS

[PPS^{+}19] Konstantinos Papangelis, Domenico Potena, Waleed W. Smari, Emanuele Storti, and Keqin Wu. Advanced technologies and systems for collaboration and computer supported cooperative work. Future Generation Computer Systems, 95(??):764–774, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19305011.

Pallickara:2012:TED

[PPZ12] Sangmi Lee Pallickara, Shrideep Pallickara, and Milija Zupanski. Towards efficient data search and subsetting of large-scale atmospheric datasets. Future Generation Computer Systems, 28(1):112–118, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11000926.

Petroni:2017:EUF

[PQBP17] Fabio Petroni, Leonardo Querzoni, Roberto Beraldi, and Mario Paolucci. Exploiting user feedback for online filtering in eventbased systems. Future Generation Computer Systems, 71(??):202–211, June 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16304125.

Pilla:2014:TAL

[PRC^{+}14] Laércio L. Pilla, Christiane P. Ribeiro, Pierre Coucheney, François Broquedis, Bruno Gaujal, Philippe O. A. Navaux, and Jean-François Méhaut. A topology-aware load balancing algorithm for clustered hierarchical multicore machines. Future Generation Computer Systems, 30(??):191–201, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001374.

Pereira:2019:ARF

[PRL^{+}19] Jorge Pereira, Leandro Ricardo, Miguel Luís, Carlos Senna, and Susana Sargento. Assessing the reliability of fog computing for smart mobility applications in VANETs. Future Generation Computer Systems, 94(??):317–332, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18307076.

Pungila:2014:EPA

[PRN14] Ciprian-Petrisor Pungilǎ, Mario Reja, and Viorel Negru. Efficient parallel automata construction for hybrid resourceimpelled data-matching. Future Generation Computer Systems, 36(??):31– 41, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001878.

Petri:2012:SLA

[PRS12] Ioan Petri, Omer F. Rana, and Gheorghe Cosmin Silaghi. Service level agreement as a complementary currency in peerto-peer markets. Future Generation Computer Systems, 28(8):1316–1327, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001737.

Paul:2013:ESC

[PRS^{+}13] P. Victer Paul, D. Rajaguru, N. Saravanan, R. Baskaran, and P. Dhavachelvan. Efficient service cache management in mobile P2P networks. Future Generation Computer Systems, 29(6):1505–1521, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1200221X.

Petri:2014:RAS

[PRSR14] Ioan Petri, Omer F. Rana, Gheorghe Cosmin Silaghi, and Yacine Rezgui. Risk assessment in service provider communities. Future Generation Computer Systems, 41(??):32–43, December 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001617.

Przymus:2014:ZMB

[PRW14] Piotr Przymus, Krzysztof Rykaczewski, and Ryszard Wiśniewski. Zebra mussels’ behaviour detection, extraction and classification using wavelets and kernel methods. Future Generation Computer Systems, 33(??):81–89, April 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000757.

Papaioannou:2010:RBE

[PS10] Thanasis G. Papaioannou and George D. Stamoulis. Reputation-based estimation of individual performance in collaborative and competitive grids. Future Generation Computer Systems, 26(8):1327–1335, October 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Prodan:2013:SCG

[PS13] Radu Prodan and Michael Sperk. Scientific computing with Google App Engine. Future Generation Computer Systems, 29(7):1851–1859, September 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000022.

Pietri:2019:PBA

[PS19] Ilia Pietri and Rizos Sakellariou. A Pareto-based approach for CPU provisioning of scientific workflows on clouds. Future Generation Computer Systems, 94(??):479–487, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18310264.

Pramsohler:2015:LIA

[PSBB15] Thomas Pramsohler, Simon Schenk, Andreas Barthels, and Uwe Baumgarten. A layered interfaceadaptation architecture for distributed componentbased systems. Future Generation Computer Systems, 47(??):113–126, June 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001812.

P:2019:RSR

[PSI19] Shaniba Asmi P., Kamalraj Subramaniam, and Nisheena V. Iqbal. A review of significant researches on prediction of preterm birth using uterine electromyogram signal. Future Generation Computer Systems, 98(??):135– 143, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18322349.

Paul:2012:QEG

[PSJ^{+}12] P. Victer Paul, N. Saravanan, S. K. V. Jayakumar, P. Dhavachelvan, and R. Baskaran. QoS enhancements for global replication management in peer to peer networks. Future Generation Computer Systems, 28(3):573–582, March 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000288.

Preissl:2010:TMS

[PSK^{+}10] Robert Preissl, Martin Schulz, Dieter Kranzlmüller, Bronis R. de Supinski, and Daniel J. Quinlan. Transforming MPI source code based on communication patterns. Future Generation Computer Systems, 26 (1):147–154, January 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pennycook:2019:IMP

[PSL19] S. J. Pennycook, J. D. Sewall, and V. W. Lee. Implications of a metric for performance portability. Future Generation Computer Systems, 92(??):947–958, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17300559.

Pan:2018:IKC

[PSLZ18] Weifeng Pan, Beibei Song, Kangshun Li, and Kejun Zhang. Identifying key classes in object-oriented software using generalized k-core decomposition. Future Generation Computer Systems, 81(??):188–202, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302492.

Pal:2016:DCA

[PSPP16] Rajesh Kumar Pal, Ierum Shanaya, Kolin Paul, and Sanjiva Prasad. Dynamic core allocation for energy efficient video decoding in homogeneous and heterogeneous multicore architectures. Future Generation Computer Systems, 56(??): 247–261, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002988.

Pagliarecci:2013:MCG

[PSS13] F. Pagliarecci, L. Spalazzi, and F. Spegni. Model checking grid security. Future Generation Computer Systems, 29(3):811–827, March 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002330.

Pillai:2018:LDE

[PSS^{+}18] Arvind Pillai, Rajkumar Soundrapandiyan, Swapnil Satapathy, Suresh Chandra Satapathy, Ki-Hyun Jung, and Rajakumar Krishnan. Local diagonal extrema number pattern: a new feature descriptor for face recognition. Future Generation Computer Systems, 81(??):297–306, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17314711.

Patton:2014:SSP

[PSW^{+}14] Evan W. Patton, Patrice Seyed, Ping Wang, Linyun Fu, F. Joshua Dein, R. Sky Bristol, and Deborah L. McGuinness. SemantEco: a semantically powered modular architecture for integrating distributed environmental and ecological data. Future Generation Computer Systems, 36(??): 430–440, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001969.

Pirbhulal:2019:JRA

[PSW^{+}19] Sandeep Pirbhulal, Oluwarotimi Williams Samuel, Wanqing Wu, Arun Kumar Sangaiah, and Guanglin Li. A joint resourceaware and medical data security framework for wearable healthcare systems. Future Generation Computer Systems, 95(??):382– 391, June 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18315474.

Piao:2019:PPG

[PSY^{+}19] Chunhui Piao, Yajuan Shi, Jiaqi Yan, Changyou Zhang, and Liping Liu. Privacy-preserving governmental data publishing: a fog-computing-based differential privacy approach.

Future Generation Computer Systems, 90(??):158– 174, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18300773.

Pavani:2016:DMS

[PT16] Gustavo Sousa Pavani and Rodrigo Izidoro Tinini. Distributed meta-scheduling in lambda grids by means of Ant Colony Optimization. Future Generation Computer Systems, 63(??): 15–24, October 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300826.

Pease:2018:IRT

[PTD^{+}18] Sarogini Grace Pease, Russell Trueman, Callum Davies, Jude Grosberg, Kai Hin Yau, Navjot Kaur, Paul Conway, and Andrew West. An intelligent realtime cyber-physical toolset for energy and process prediction and optimisation in the future industrial Internet of Things. Future Generation Computer Systems, 79 (part 3)(??):815– 829, February 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630382X.

Piparo:2018:SSI

[PTM^{+}18] Danilo Piparo, Enric Tejedor, Pere Mato, Luca Mascetti, Jakub Moscicki, and Massimo Lamanna. SWAN: a service for interactive analysis in the cloud. Future Generation Computer Systems, 78 (part 3)(??):1071–1078, January 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16307105.

Pirro:2012:DBS

[PTT12] Giuseppe Pirrò, Domenico Talia, and Paolo Trunfio. A DHT-based semantic overlay network for service discovery. Future Generation Computer Systems, 28(4): 689–707, April 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11002305.

Papadakis-Vlachopapadopoulos:2019:CSR

[PVGD^{+}19] Konstantinos Papadakis-Vlachopapadopoulos, Román Sosa González, Ioannis Dimolitsas, Dimitrios Dechouniotis, Ana Juan Ferrer, and Symeon Papavassiliou. Collaborative SLA and reputation-based trust management in cloud federations. Future Generation Computer Systems, 100(??):498–512, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18329248.

Pico-Valencia:2019:SMB

[PVHTP19] Pablo Pico-Valencia, Juan A.

Holgado-Terriza, and Patricia Paderewski. A systematic method for building Internet of Agents applications based on the Linked Open Data approach. Future Generation Computer Systems, 94(??): 250–271, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17321234.

Pandey:2012:ACE

[PVN^{+}12] Suraj Pandey, William Voorsluys, Sheng Niu, Ahsan Khandoker, and Rajkumar Buyya. An autonomic cloud environment for hosting ECG data analysis services. Future Generation Computer Systems, 28(1):147–154, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11000732.

Panah:2017:TAA

[PvSS17] Arezou Soltani Panah, Ron van Schyndel, and Timos Sellis. Towards an asynchronous aggregationcapable watermark for end-to-end protection of big data streams. Future Generation Computer Systems, 72(??):288–304, July 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302965.

Plaga:2019:SFD

[PWA^{+}19] Sven Plaga, Norbert Wiedermann, Simon Duque Anton, Stefan Tatschner, Hans Schotten, and Thomas Newe. Securing future decentralised industrial IoT infrastructures: Challenges and free open source solutions. Future Generation Computer Systems, 93(??):596–608, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18314043.

Pettit:2013:BEP

[PWB^{+}13] Christopher Pettit, Steve Williams, Ian Bishop, Jean-Philippe Aurambout, A. B. M. Russel, Anthony Michael, Subhash Sharma, David Hunter, Pang Choung Chan, Colin Enticott, Ann Borda, and David Abramson. Building an ecoinformatics platform to support climate change adaptation in Victoria. Future Generation Computer Systems, 29(2):624–640, February 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001385.

Pei:2017:EPU

[PWMX17] Xiaoqiang Pei, Yijie Wang, Xingkong Ma, and Fangliang Xu. Efficient inplace update with grouped and pipelined data transmission in erasure-coded storage systems. Future Generation Computer Systems, 69(??):24–40, April 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16304046.

Peng:2018:IID

[PWP^{+}18] Yong Peng, Xinghua Wang, Shuangling Peng, Helai Huang, Guangdong Tian, and Hongfei Jia. Investigation on the injuries of drivers and copilots in rearend crashes between trucks based on real world accident data in China. Future Generation Computer Systems, 86(??):1251–1258, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1731734X.

Polap:2018:MTL

[PWWD18] Dawid Polap, Marcin Woźniak, Wei Wei, and Robertas Damasevicius. Multi-threaded learning control mechanism for neural networks. Future Generation Computer Systems, 87(??):16–34, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18300931.

Pournaras:2017:SRS

[PYH17] Evangelos Pournaras, Mark Yao, and Dirk Helbing. Self-regulating supply-demand systems. Future Generation Computer Systems, 76(??):73–91, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16303946.

Pagnin:2018:HDB

[PYH^{+}18] Elena Pagnin, Anjia Yang, Qiao Hu, Gerhard Hancke, and Aikaterini Mitrokotsa. HB^{+}DB: Distance bounding meets human based authentication. Future Generation Computer Systems, 80(??):627–639, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301492.

Peng:2018:SNB

[PYM18] Sancheng Peng, Shui Yu, and Peter Mueller. Social networking big data: Opportunities, solutions, and challenges. Future Generation Computer Systems, 86(??):1456–1458, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18312184.

Pagan:2018:PTW

[PZA18] Josué Pagán, Marina Zapater, and José L. Ayala. Power transmission and workload balancing policies in eHealth mobile cloud computing scenarios. Future Generation Computer Systems, 78 (part 2)(??):587–601, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302133.

P:2019:MLB

[PZC19] Vinod P., Akka Zemmari, and Mauro Conti. A machine learning based approach to detect malicious Android apps using discriminant system calls. Future Generation Computer Systems, 94(??):333–350, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18306216.

Pop:2016:MAT

[PZY16] Florin Pop, Xiaomin Zhu, and Laurence T. Yang. MidHDC: Advanced topics on middleware services for heterogeneous distributed computing. Part 1. Future Generation Computer Systems, 56(??):734–735, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003738.

Pop:2017:MAT

[PZY17] Florin Pop, Xiaomin Zhu, and Laurence T. Yang. MidHDC: Advanced topics on middleware services for heterogeneous distributed computing. Part

2. Future Generation Computer Systems, 74(??):86– 89, September 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1730972X.

Quick:2013:DDM

[QC13] Darren Quick and Kim-Kwang Raymond Choo. Digital droplets: Microsoft SkyDrive forensic data remnants. Future Generation Computer Systems, 29(6):1378–1394, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000265.

Quick:2018:DFI

[QC18] Darren Quick and Kim-Kwang Raymond Choo. Digital forensic intelligence: Data subsets and open source intelligence (DFINT + OSINT): a timely and cohesive mix. Future Generation Computer Systems, 78 (part 2)(??):558–567, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308639.

Quarati:2016:DCS

[QCD16] Alfonso Quarati, Andrea Clematis, and Daniele D’Agostino. Delivering cloud services with QoS requirements: Business opportunities, architectural solutions and energy-saving aspects. Future Generation Computer Systems, 55(??):403–427, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000539.

Qiu:2018:QDS

[QCX18] Lirong Qiu, Feng Cai, and Guixian Xu. Quantum digital signature for the access control of sensitive data in the big data era. Future Generation Computer Systems, 86(??):372–379, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18301250.

Qiu:2019:NQE

[QCY^{+}19] Chao Qiu, Shaohua Cui, Haipeng Yao, Fangmin Xu, F. Richard Yu, and Chenglin Zhao. A novel QoS-enabled load scheduling algorithm based on reinforcement learning in software-defined energy internet. Future Generation Computer Systems, 92(??):43–51, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1830308X.

Qiang:2017:MUC

[QCYJ17] Weizhong Qiang, Feng Chen, Laurence T. Yang, and Hai Jin. MUC: Updating cloud applications dynamically via multi-version execution. Future Generation Computer Systems, 74(??):254–264, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003866.

Qiu:2019:UMV

[QCZH19] Xiru Qiu, Zhikui Chen, Liang Zhao, and Chengsheng Hu. Unsupervised multi-view non-negative for law data feature learning with dual graphregularization in smart Internet of Things. Future Generation Computer Systems, 100(??):523–530, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1931091X.

Qiu:2018:PUC

[QGT^{+}18] Meikang Qiu, Keke Gai, Bhavani Thuraisingham, Lixin Tao, and Hui Zhao. Proactive user-centric secure data scheme using attribute-based semantic access controls for mobile clouds in financial industry. Future Generation Computer Systems, 80(??): 421–429, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1600008X.

Qiu:2018:PPW

[QGX18] Meikang Qiu, Keke Gai, and Zenggang Xiong. Privacy-preserving wireless communications using bipartite matching in social big data. Future Generation Computer Systems, 87(??):772–781, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17301449.

Qamar:2019:MMA

[QKC19] Attia Qamar, Ahmad Karim, and Victor Chang. Mobile malware attacks: Review, taxonomy and future directions. Future Generation Computer Systems, 97(??):887–909, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331601.

Qian:2018:AAE

[QLM^{+}18] Yongfeng Qian, Jiayi Lu, Yiming Miao, Wen Ji, Renchao Jin, and Enmin Song. AIEM: AI-enabled affective experience management. Future Generation Computer Systems, 89(??):438–445, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18307143.

Qiu:2019:DCR

[QMCX19] Han Qiu, Gerard Memmi, Xuan Chen, and Jian Xiong. DC coefficient recovery for JPEG images in ubiquitous communication systems. Future Generation Computer Systems, 96 (??):23–31, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1832942X.

Quan:2012:APE

[QMSG12] Dang Minh Quan, Federico Mezza, Domenico Sannenli, and Raffaele Giafreda. T-Alloc: a practical energy efficient resource allocation algorithm for traditional data centers. Future Generation Computer Systems, 28(5): 791–800, May 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000719.

Qu:2019:STB

[QNM^{+}19] Qiang Qu, Ildar Nurgaliev, Muhammad Muzammal, Christian S. Jensen, and Jianping Fan. On spatiotemporal blockchain query processing. Future Generation Computer Systems, 98(??):208–218, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314213.

Quezada-Pina:2012:APJ

[QPTGG^{+}12] Ariel Quezada-Pina, Andrei Tchernykh, José Luis González-García, Adán Hirales-Carbajal, Juan Manuel Ramírez-Alcaraz, Uwe Schwiegelshohn, Ramin Yahyapour, and Vanessa Miranda-López. Adaptive parallel job scheduling with resource admissible allocation on twolevel hierarchical grids. Future Generation Computer Systems, 28(7):965–976, July 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000374.

Qiao:2018:CTC

[QRW^{+}18] Huidong Qiao, Jiangchun Ren, Zhiying Wang, Haihe Ba, and Huaizhe Zhou. Compulsory traceable ciphertextpolicy attribute-based encryption against privilege abuse in fog computing. Future Generation Computer Systems, 88(??):107– 116, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17328820.

Qureshi:2019:PBP

[Qur19] Basit Qureshi. Profilebased power-aware workflow scheduling framework for energy-efficient data centers. Future Generation Computer Systems, 94(??): 453–467, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18318491.

Qasha:2019:SPO

[QWCW19] Rawaa Qasha, Zhenyu Wen, Jacek Cała, and Paul Watson. Sharing and performance optimization of reproducible workflows in the cloud. Future Generation Computer Systems, 98(??):487–502, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314377.

Qian:2019:EYN

[QXZ^{+}19] Wang Qian, Wang Xiaoyi, Ye Zi, Yan Yuqing, and Xu Dieyi. Effects of Yin-nourishing and bloodcooling decoction on proteinuria and renal tubular damage in IgA nephropathy. Future Generation Computer Systems, 98(??):682–687, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326414.

Qi:2018:TSL

[QZD^{+}18] Lianyong Qi, Xuyun Zhang, Wanchun Dou, Chunhua Hu, Chi Yang, and Jinjun Chen. A two-stage localitysensitive hashing based approach for privacypreserving mobile service recommendation in crossplatform edge environment. Future Generation Computer Systems, 88(??):636–643, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18301420.

Qiu:2018:TES

[QZM^{+}18] Tie Qiu, Aoyang Zhao, Ruixin Ma, Victor Chang, Fangbing Liu, and Zhangjie Fu. A task-efficient sink node based on embedded multi-core SoC for Internet of Things. Future Generation Computer Systems, 82(??):656–666, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X16308032.

Rohitratana:2012:IPS

[RA12] Juthasit Rohitratana and Jörn Altmann. Impact of pricing schemes on a market for Software-as-a-Service and perpetual software. Future Generation Computer Systems, 28(8):1328–1339, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000842.

Riahi:2018:IEU

[RAA^{+}18] Hassen Riahi, Alberto Aimar, Alejandro Álvarez Ayllón, Justas Balcas, Diego Ciangottini, José M. Hernández, Oliver Keeble, Nicolò Magini, Andrea Manzi, Luca Mascetti, Marco Mascheroni, Andres Jorge Tanasijczuk, and Eric Wayne Vaandering. Integration of enduser cloud storage for CMS analysis. Future Generation Computer Systems, 78 (part 3)(??):1079–1082, January 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17306271.

Riaz:2019:PBP

[RAA^{+}19] Farhan Riaz, Muhammad Ajmal Azad, Junaid Arshad, Muhammad Imran, Ali Hassan, and Saad Rehman. Pervasive blood pressure monitoring using Photoplethysmogram (PPG) sensor. Future Generation Computer Systems, 98(??):120–130, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18327729.

Ray:2018:USO

[RACA18] Biplob R. Ray, Jemal Abawajy, Morshed Chowdhury, and Abdulhameed Alelaiwi. Universal and secure object ownership transfer protocol for the Internet of Things. Future Generation Computer Systems, 78 (part 2)(??):838– 849, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302182.

Ribeiro:2019:TFB

[RAdARP19] Igor C. G. Ribeiro, Célio Albuquerque, Antônio A. de A. Rocha, and Diego Passos. THOR: a framework to build an advanced metering infrastructure resilient to DAP failures in smart grids. Future Generation Computer Systems, 99(??):11–26, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326268.

Radhakrishna:2018:NFS

[RAKJ18] Vangipuram Radhakrishna, Shadi A. Aljawarneh, P. V. Kumar, and V. Janaki. A novel fuzzy similarity measure and prevalence estimation approach for similarity profiled temporal association pattern mining. Future Generation Computer Systems, 83(??):582– 595, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17303795.

Rao:2017:SEC

[Rao17] Y. Sreenivasa Rao. A secure and efficient ciphertextpolicy attribute-based signcryption for personal health records sharing in cloud computing. Future Generation Computer Systems, 67(??):133–151, February 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302746.

Rasmussen:2012:DSP

[RB12] Rune Rasmussen and Ross Brown. A deductive system for proving workflow models from operational procedures. Future Generation Computer Systems, 28(5): 732–742, May 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000131.

Ricci:2013:E

[RB13] Laura Ricci and Ranieri Baraglia. Editorial. Future Generation Computer Systems, 29(6):1459–1460, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000514.

Rodriguez:2018:SDW

[RB18] Maria A. Rodriguez and Rajkumar Buyya. Scheduling dynamic workloads in multi-tenant scientific workflow as a service platforms. Future Generation Computer Systems, 79 (part 2)(??):739–750, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17301681.

Rahman:2017:PPF

[RBA17] Farzana Rahman, Md Zakirul Alam Bhuiyan, and Sheikh Iqbal Ahamed. A privacy preserving framework for RFID based healthcare systems. Future Generation Computer Systems, 72(??):339–352, July 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301819.

Rycerz:2015:CES

[RBC^{+}15] Katarzyna Rycerz, Marian Bubak, Eryk Ciepiela, Daniel Harezlak, Tomasz Gubala, Jan Meizner, Maciej Pawlik, and Bartosz Wilk. Composing, execution and sharing of multiscale applications. Future Generation Computer Systems, 53(??):77–87, December 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002034.

Raghavendra:2018:ASD

[RBGA18] U. Raghavendra, N. Shyamasunder Bhat, Anjan Gudigar, and U. Rajendra Acharya. Automated system for the detection of thoracolumbar fractures using a CNN architecture. Future Generation Computer Systems, 85(??):184– 189, August 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17321544.

Ravi:2013:SCA

[RBJ^{+}13] Vignesh T. Ravi, Michela Becchi, Wei Jiang, Gagan Agrawal, and Srimat Chakradhar. Scheduling concurrent applications on a cluster of CPU–GPU nodes. Future Generation Computer Systems, 29(8):2262– 2271, October 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001167.

Rieffel:2014:PAP

[RBLvM14] Eleanor G. Rieffel, Jacob T. Biehl, Adam J. Lee, and William van Melle. Private aggregation for presence streams. Future Generation Computer Systems, 31(??):169– 181, February 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001088.

Ranjan:2013:MDP

[RBN13] Rajiv Ranjan, Rajkumar Buyya, and Surya Nepal. Model-driven provisioning of application services in hybrid computing environments. Future Generation Computer Systems, 29(5): 1211–1215, July 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000216.

Rajni:2013:BFB

[RC13] Rajni and Inderveer Chana.

Bacterial foraging based hyper-heuristic for resource scheduling in grid computing. Future Generation Computer Systems, 29(3): 751–762, March 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001781.

Rho:2018:SIT

[RC18] Seungmin Rho and Yu Chen. Social Internet of Things: Applications, architectures and protocols. Future Generation Computer Systems, 82(??):667–668, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18301158.

Rho:2019:SIT

[RC19] Seungmin Rho and Yu Chen. Social Internet of Things: Applications, architectures and protocols. Future Generation Computer Systems, 92(??):959–960, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18328589.

Rafailidis:2017:LSS

[RCM17] D. Rafailidis, E. Constantinou, and Y. Manolopoulos. Landmark selection for spectral clustering based on Weighted PageRank. Future Generation Computer Systems, 68(??):465– 472, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300504.

Reina:2018:EDL

[RCMT18] D. G. Reina, T. Camp, A. Munjal, and S. L. Toral. Evolutionary deployment and local search-based movements of 0th responders in disaster scenarios. Future Generation Computer Systems, 88(??):61– 78, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17325372.

Rodriguez-Covili:2011:TRA

[RCOP^{+}11]  Juan  Rodríguez-Covili,

Sergio F. Ochoa, José A. Pino, Valeria Herskovic, Jesus Favela, David Mejía, and Alberto L. Morán. Towards a reference architecture for the design of mobile shared workspaces. Future Generation Computer Systems, 27(1):109– 118, January 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ruan:2019:VMA

[RCTY19] Xiaojun Ruan, Haiquan Chen, Yun Tian, and Shu Yin. Virtual machine allocation and migration based on performance-to-power ratio in energy-efficient clouds. Future Generation Computer Systems, 100(??):380–394, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18321629.

Ren:2019:MSD

[RCW^{+}19] Lei Ren, Xuejun Cheng, Xiaokang Wang, Jin Cui, and Lin Zhang. Multiscale dense gate recurrent unit networks for bearing remaining useful life prediction. Future Generation Computer Systems, 94(??): 601–609, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1832096X.

Rasooli:2014:CCO

[RD14] Aysan Rasooli and Douglas G. Down. COSHH: a classification and optimization based scheduler for heterogeneous Hadoop systems. Future Generation Computer Systems, 36 (??):1–15, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000077.

Rajinikanth:2018:AEM

[RDSA18] V. Rajinikanth, Nilanjan Dey, Suresh Chandra Satapathy, and Amira S. Ashour. An approach to examine Magnetic Resonance Angiography based on Tsallis entropy and deformable snake model. Future Generation Computer Systems, 85(??):160–172, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17322537.

Rahmanian:2018:LAB

[RGAT18] Ali Asghar Rahmanian, Mostafa Ghobaei-Arani, and Sajjad Tofighy. A learning automata-based ensemble resource usage prediction algorithm for cloud computing environment. Future Generation Computer Systems, 79 (part 1)(??):54– 71, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17309378.

Rodero:2010:GBS

[RGC^{+}10] Ivan Rodero, Francesc Guim, Julita Corbalan, Liana Fong, and S. Masoud Sadjadi. Grid broker selection strategies using aggregated resource information. Future Generation Computer Systems, 26(1):72–86, January 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ros-Giralt:2018:ADS

[RGCCL18] Jordi Ros-Giralt, Alan Commike, Peter Cullen, and Richard Lethin. Algorithms and data structures to accelerate network analysis. Future Generation Computer Systems, 86(??):535–545, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1830222X.

Rico-Gallego:2016:ELM

[RGDML16] Juan-Antonio Rico-Gallego, Juan-Carlos Díaz-Martín, and Alexey L. Lastovetsky. Extending τ-lop to model concurrent MPI communications in multicore clusters. Future Generation Computer Systems, 61(??): 66–82, August 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300346.

Ramirez-Gallego:2018:OEB

[RGGH18] S. Ramírez-Gallego, S. García, and F. Herrera. Online entropy-based discretization for data streaming classification. Future Generation Computer Systems, 86(??):59–70, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17325815.

ReboucasFilho:2019:EPP

[RGM^{+}19] Pedro P. Rebouças Filho, Samuel L. Gomes, Navar M. Mendonça e Nascimento, Cláudio M. S. Medeiros, Fatma Outay, and Victor Hugo C. de Albuquerque. Energy production predication via Internet of Things based machine learning system. Future Generation Computer Systems, 97(??):180–193, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18328486.

Rahmani:2018:ESH

[RGN^{+}18] Amir M. Rahmani, Tuan Nguyen Gia, Behailu Negash, Arman Anzanpour, Iman Azimi, Mingzhe Jiang, and Pasi Liljeberg. Exploiting smart e-health gateways at the edge of healthcare Internet-of-Things: a fog computing approach. Future Generation Computer Systems, 78 (part 2)(??):641–658, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302121.

Rego:2018:SDN

[RGSL18] Albert Rego, Laura Garcia, Sandra Sendra, and Jaime Lloret. Software Defined Network-based control system for an efficient traffic management for emergency situations in smart cities. Future Generation Computer Systems, 88(??):243– 253, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17330364.

Rodriguez-Garcia:2014:CSE

[RGVGGSSZ14] Miguel Ángel Rodríguez-García, Rafael Valencia-García, Francisco García-Sánchez, and J. Javier Samper-Zapater. Creating a semantically-enhanced cloud services environment through ontology evolution. Future Generation Computer Systems, 32(??): 295–306, March 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001684.

Rahman:2016:SPV

[RHH^{+}16] Sk. Md. Mizanur Rahman, M. Anwar Hossain, Mohammad Mehedi Hassan, Atif Alamri, Abdullah Alghamdi, and Mukaddim Pathan. Secure privacy vault design for distributed multimedia surveillance system. Future Generation Computer Systems, 55(??):344–352, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1400212X.

Raza:2019:CAM

[RHH^{+}19] Muhammad Raza, Farookh Khadeer Hussain, Omar Khadeer Hussain, Ming Zhao, and Zia ur Rehman. A comparative analysis of machine learning models for quality pillar assessment of SaaS services by multi-class text classification of users’ reviews. Future Generation Computer Systems, 101(??):341–371, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19300196.

Ronkko:2015:APE

[RHKC15] Mauno Rönkkö, Jani Heikkinen, Ville Kotovirta, and Venkatachalam Chandrasekar. Automated preprocessing of environmental data. Future Generation Computer Systems, 45(??):13–24, April 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002040.

Rosas:2014:WSR

[RHMGC14] Erika Rosas, Nicolas Hidalgo, Mauricio Marin, and Veronica Gil-Costa. Web search results caching service for structured P2P networks. Future Generation Computer Systems, 30(??):254–264, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001325.

Raza:2017:SEE

[RHPV17] Shahid Raza, Tómas Helgason, Panos Papadimitratos, and Thiemo Voigt. SecureSense: End-to-end secure communication architecture for the cloud-connected Internet of Things. Future Generation Computer Systems, 77(??):40–51, December 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17312360.

Rajkumar:2019:SSI

[RJN^{+}19] G. Rajkumar, R. Jayabharathy, K. Narasimhan, N. Raju, M. Easwaran, V. Elamaran, Gustavo Ramirez-gonzalez, and Marlon Burbano-fernandez. Spectral and SNR improvement analysis of normal and abnormal heart sound signals using different windows. Future Generation Computer Systems, 92(??):438–443, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18318132.

Rathod:2019:CDG

[RJS^{+}19] Vishal Rathod, Natasha Jeppu, Samanvita Sastry, Shruti Singala, and Mohit P. Tahiliani. Co-CoA++: Delay gradient based congestion control for Internet of Things. Future Generation Computer Systems, 100(??):1053– 1072, November 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18308677.

Rodriguez:2018:DPA

[RKB18] Maria A. Rodriguez, Ramamohanarao Kotagiri, and Rajkumar Buyya. Detecting performance anomalies in scientific workflows using hierarchical temporal memory. Future Generation Computer Systems, 88(??):624–635, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17327292.

Ren:2017:MCT

[RLL^{+}17] Yongmao Ren, Jun Li, Lingling Li, Shanshan Shi, Jiang Zhi, and Haibo Wu. Modeling content transfer performance in information-centric networking. Future Generation Computer Systems, 74(??):12–19, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17305861.

Roman:2018:MEC

[RLM18] Rodrigo Roman, Javier Lopez, and Masahiro Mambo. Mobile edge computing, Fog et al.: a survey and analysis of security threats and challenges. Future Generation Computer Systems, 78 (part 2)(??):680–698, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16305635.

Ramachandran:2012:DAR

[RLP12] Karthick Ramachandran, Hanan Lutfiyya, and Mark Perry. Decentralized approach to resource availability prediction using group availability in a P2P desktop grid. Future Generation Computer Systems, 28(6):854–860, June 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X10001998.

Reuillon:2013:OWE

[RLRC13] Romain Reuillon, Mathieu Leclaire, and Sebastien Rey-Coyrehourcq. Open-MOLE, a workflow engine specifically tailored for the distributed exploration of simulation models. Future Generation Computer Systems, 29(8):1981–1990, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001027.

Roy:2011:ERM

[RM11] Sarbani Roy and Nandini Mukherjee. Efficient resource management for running multiple concurrent jobs in a computational grid environment. Future Generation Computer Systems, 27(8):1070– 1082, October 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Roshanbin:2016:AIU

[RM16] Narges Roshanbin and James Miller. ADAMAS: Interweaving Unicode and color to enhance CAPTCHA security. Future Generation Computer Systems, 55(??):289–310, February 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002386.

Riboni:2019:SBA

[RM19] Daniele Riboni and Marta Murtas. Sensor-based activity recognition: One picture is worth a thousand words. Future Generation Computer Systems, 101(??):709–722, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19303863.

Renambot:2016:SCP

[RMA^{+}16] Luc Renambot, Thomas Marrinan, Jillian Aurisano, Arthur Nishimoto, Victor Mateevitsi, Krishna Bharadwaj, Lance Long, Andy Johnson, Maxine Brown, and Jason Leigh. SAGE2: a collaboration portal for scalable resolution displays. Future Generation Computer Systems, 54(??):296–305, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001892.

Rezaeimehr:2018:TTC

[RMA^{+}18] Fatemeh Rezaeimehr, Parham Moradi, Sajad Ahmadian, Nooruldeen Nasih Qader, and Mahdi Jalili. TCARS: Time- and communityaware recommendation system. Future Generation Computer Systems, 78 (part 1)(??):419–429, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17305411.

Reyna:2018:BII

[RMC^{+}18] Ana Reyna, Cristian Martín, Jaime Chen, Enrique Soler, and Manuel Díaz. On blockchain and its integration with IoT. Challenges and opportunities. Future Generation Computer Systems, 88(??):173– 190, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17329205.

Rodero-Merino:2012:UCS

[RMCMD12] Luis Rodero-Merino, Eddy Caron, Adrian Muresan, and Frédéric Desprez. Using clouds to scale grid resources: An economic model. Future Generation Computer Systems, 28(4): 633–646, April 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001865.

Reyes:2010:MSG

[RMCN^{+}10] S. Reyes, C. Muñoz-Caro, A. Niño, R. Sirvent, and R. M. Badia. Monitoring and steering Grid applications with GRID superscalar. Future Generation Computer Systems, 26(4):645–653, April 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Roy:2018:AAE

[RMDB18] Deepsubhra Guha Roy, Bipasha Mahato, Debashis De, and Rajkumar Buyya. Application-aware end-to-end delay and message loss estimation in Internet of Things (IoT)–MQTT–SN protocols. Future Generation Computer Systems, 89(??):300–316, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17329990.

Rubio-Montero:2015:GEM

[RMHCMG15] A. J. Rubio-Montero, E. Huedo, F. Castejón, and R. Mayo-García. GWpilot: Enabling multi-level scheduling in distributed infrastructures with Grid-Way and pilot jobs. Future Generation Computer Systems, 45(??):25–52, April 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001873.

Rubio-Montero:2017:SMV

[RMHMG17] A. J. Rubio-Montero, E. Huedo, and R. Mayo-García. Scheduling multiple virtual environments in cloud federations for distributed calculations. Future Generation Computer Systems, 74(??):90– 103, September 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300656.

Re:2018:ESE

[RMJ^{+}18] Reginaldo Ré, Rômulo Manciola Meloca, Douglas Nassif Roma Junior, Marcelo Alexandre da Cruz Ismael, and Gabriel Costa Silva. An empirical study for evaluating the performance of multi-cloud APIs. Future Generation Computer Systems, 79 (part 2)(??):726– 738, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17301802.

Ralha:2019:MSD

[RML^{+}19] Célia Ghedini Ralha, Aldo H. D. Mendes, Luiz A. Laranjeira, Aletéia P. F. Araújo, and Alba C. M. A. Melo. Multiagent system for dynamic resource provisioning in cloud computing platforms. Future Generation Computer Systems, 94 (??):80–96, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18301742.

Rezaei-Mayahi:2019:TAP

[RMRSA19] Mehdi Rezaei-Mayahi, Mostafa Rezazad, and Hamid Sarbazi-Azad. Temperature-aware power consumption modeling in Hyperscale cloud data centers. Future Generation Computer Systems, 94(??):130–139, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18316182.

Rashid:2017:UAR

[RMSPP17] Zulqarnain Rashid, Joan Melià-Seguí, Rafael Pous, and Enric Peig. Using Augmented Reality and Internet of Things to improve accessibility of people with motor disabilities in the context of smart cities. Future Generation Computer Systems, 76(??):248–261, November 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306860.

Rodero-Merino:2010:IDS

[RMVG^{+}10] Luis Rodero-Merino, Luis M.

Vaquero, Victor Gil, Fermín Galán, Javier Fontán, Rubén S. Montero, and Ignacio M. Llorente. From infrastructure delivery to service management in clouds. Future Generation Computer Systems, 26(8):1226– 1240, October 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Riesinger:2018:NSP

[RNR18] Christoph Riesinger, Tobias Neckel, and Florian Rupp. Non-standard pseudo random number generators revisited for GPUs. Future Generation Computer Systems, 82(??):482–492, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X16307993.

Rehman:2019:HRN

[ROK19] Osama Rehman and Mohamed Ould-Khaoua. A hybrid relay node selection scheme for message dissemination in VANETs. Future Generation Computer Systems, 93(??):1– 17, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18320417.

Ravandi:2018:SOR

[RP18] Babak Ravandi and Ioannis Papapanagiotou. A self-organized resource provisioning for cloud block storage. Future Generation Computer Systems, 89(??):765–776, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17330121.

Rathore:2018:RTS

[RPA^{+}18] M. Mazhar Rathore, Anand Paul, Awais Ahmad, Naveen Chilamkurti, Won-Hwa Hong, and HyunCheol Seo. Real-time secure communication for smart city in high-speed big data environment. Future Generation Computer Systems, 83(??):638–652, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17317557.

Rabehaja:2019:DIS

[RPH19] Tahiry Rabehaja, Shantanu Pal, and Michael Hitchens. Design and implementation of a secure and flexible access-right delegation for resource constrained environments. Future Generation Computer Systems, 99(??):593–608, October 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331182.

Rodriguez:2010:PEA

[RPMG10] Gabriel Rodríguez, Xoán C.

Pardo, María J. Martín, and Patricia González. Performance evaluation of an application-level checkpointing solution on Grids.

Future Generation Computer Systems, 26(7):1012– 1023, July 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rahim:2019:SAB

[RQN^{+}19] Azizur Rahim, Tie Qiu, Zhaolong Ning, Jinzhong Wang, Noor Ullah, Amr Tolba, and Feng Xia. Social acquaintance based routing in vehicular social networks. Future Generation Computer Systems, 93(??): 751–760, April 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167739X17304326.

Rapuzzi:2018:BSA

[RR18] R. Rapuzzi and M. Repetto. Building situational awareness for network threats in fog/edge computing: Emerging paradigms beyond the security perimeter model. Future Generation Computer Systems, 85(??):235–249, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17327978.

Rahman:2010:CDW

[RRB10] Mustafizur Rahman, Rajiv Ranjan, and Rajkumar Buyya. Cooperative and decentralized workflow scheduling in global grids. Future Generation Computer Systems, 26(5):753– 768, May 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rusnak:2016:TNM

[RRH16] Vít Rusnák, Lukás Rucka, and Petr Holub. Toward natural multi-user interaction in advanced collaborative display environments. Future Generation Computer Systems, 54(??):313– 325, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000813.

Rajesh:2019:SCI

[RRKA19] K. Rajesh, V. Ramaswamy, K. Kannan, and N. Arunkumar. Satellite cloud image classification for cyclone prediction using dichotomous logistic regression based fuzzy hypergraph model. Future Generation Computer Systems, 98(??):688–696, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326669.

Rajovic:2014:TMC

[RRP^{+}14] Nikola Rajovic, Alejandro Rico, Nikola Puzovic, Chris Adeniyi-Jones, and Alex Ramirez. Tibidabo: Making the case for an ARM-based HPC system. Future Generation Computer Systems, 36(??):322– 334, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001581.

Rajasekaran:2010:SGA

[RRS10] M. Pallikonda Rajasekaran, S. Radhakrishnan, and P. Subbaraj. Sensor grid applications in patient monitoring. Future Generation Computer Systems, 26(4):569–575, April 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Riaz:2018:FNF

[RRU^{+}18] Shehroz Riaz, Maaz Rehan, Tariq Umer, Muhammad Khalil Afzal, Waqas Rehan, Ehsan Ullah Munir, and Tassawar Iqbal. FRP: a novel fast rerouting protocol with multilink-failure recovery for mission-critical WSN. Future Generation Computer Systems, 89(??):148–165, December 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18303935.

Reddy:2016:NSF

[RS16] P. Vijaya Vardhan Reddy and K. Shyamala. New scoring formula to rank hypervisors’ performance complementing with statistical analysis using DOE. Future Generation Computer Systems, 61(??):54– 65, August 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300255.

Rashidi:2017:HHQ

[RS17a] Shima Rashidi and Saeed Sharifian. A hybrid heuristic queue based algorithm for task assignment in mobile cloud. Future Generation Computer Systems, 68(??):331–345, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16304022.

Roy:2017:LOS

[RS17b] Dipanjan Roy and Anirban Sengupta. Low overhead symmetrical protection of reusable IP core using robust fingerprinting and watermarking during high level synthesis. Future Generation Computer Systems, 71(??):89– 101, June 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16305556.

Rosas:2014:DTW

[RSJ^{+}14] Claudia Rosas, Anna Sikora, Josep Jorba, Andreu Moreno, Antonio Espinosa, and Eduardo César. Dynamic tuning of the workload partition factor and the resource utilization in dataintensive applications. Future Generation Computer Systems, 37(??):162–177, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002665.

Rashti:2016:LHS

[RSK16] Mohammad Javad Rashti, Gerald Sabin, and Rajkumar Kettimuthu. Longhaul secure data transfer using hardware-assisted GridFTP. Future Generation Computer Systems, 56(??):265–276, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002940.

Rehmat:2018:UFA

[RSRA18] Muhammad Asim Rehmat, Muhammad Shahbaz, Asad Raza, and Haider Abbas. A unified framework for automated inspection of handheld safety critical devices in production assemblies. Future Generation Computer Systems, 88(??):342– 356, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18309051.

Raza:2018:ABP

[RSY^{+}18] Mudassar Raza, Muhammad Sharif, Mussarat Yasmin, Muhammad Attique Khan, Tanzila Saba, and Steven Lawrence Fernandes. Appearance based pedestrians’ gender recognition by employing stacked auto encoders in deep learning. Future Generation Computer Systems, 88(??):28– 39, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17322288.

Rao:2015:HBS

[RT15] K. Sunil Rao and P. Santhi Thilagam. Heuristics based server consolidation with residual resource defragmentation in cloud data centers. Future Generation Computer Systems, 50(??):87– 98, September 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001794.

Rajavel:2016:APB

[RT16] Rajkumar Rajavel and Mala Thangarathanam. Adaptive probabilistic behavioural learning system for the effective behavioural decision in cloud trading negotiation market. Future Generation Computer Systems, 58(??): 29–41, May 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500391X.

Ranjan:2017:NEI

[RTHB17] Rajiv Ranjan, Dhavalkumar Thakker, Armin Haller, and Rajkumar Buyya. A note on exploration of IoT generated big data using semantics. Future Generation Computer Systems, 76(??):495–498, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17313912.

Ren:2016:SSC

[RTS^{+}16] Shu Qin Ren, Benjamin Hong Meng Tan, Sivaraman Sundaram, Taining Wang, Yibin Ng, Victor Chang, and Khin Mi Mi Aung. Secure searching on cloud storage enhanced by homomorphic indexing. Future Generation Computer Systems, 65(??):102– 110, December 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300577.

Rho:2016:CPSb

[RVC16a] Seungmin Rho, Athanasios V. Vasilakos, and Weifeng Chen. Cyberphysical systems technologies and application — Part II. Future Generation Computer Systems, 61(??): 83–84, August 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300516.

Rho:2016:CPSa

[RVC16b] Seungmin Rho, Athanasios V. Vasilakos, and Weifeng Chen. Cyber physical systems technologies and applications. Future Generation Computer Systems, 56(??):436–437, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003325.

Regueiro:2017:SMO

[RVST17] Manuel A. Regueiro, José R. R. Viqueira, Christoph Stasch, and José A. Taboada. Semantic mediation of observation datasets through Sensor Observation Services. Future Generation Computer Systems, 67(??):47–56, February 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302722.

Rajasekaran:2013:SSI

[RW13] Sanguthevar Rajasekaran and Michal Wozniak. Special section on invited papers from NetCoM-2009. Future Generation Computer Systems, 29(1):241, January 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001288.

Ren:2018:SBC

[RW18] Lifang Ren and Wenjian Wang. An SVMbased collaborative filtering approach for Top-N web services recommendation. Future Generation Computer Systems, 78 (part 2)(??):531–543, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300389.

Rezaeifar:2019:RAF

[RWO^{+}19] Zeinab Rezaeifar, Jian Wang, Heekuck Oh, Suk-Bok Lee, and Junbeom Hur. A reliable adaptive forwarding approach in named data networking. Future Generation Computer Systems, 96(??): 538–551, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18316431.

Remenska:2013:UMC

[RWV^{+}13] Daniela Remenska, Tim A. C. Willemse, Kees Verstoep, Jeff Templon, and Henri Bal. Using model checking to analyze the system behavior of the LHC production grid. Future Generation Computer Systems, 29(8):2239–2251, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001180.

Ren:2018:RTE

[RWY^{+}18] Yizhi Ren, Gang Wang, Lanping Yu, Benyun Shi, Weitong Hu, and Zhen Wang. Rigorous or tolerant: The effect of different reputation attitudes in complex networks. Future Generation Computer Systems, 83(??):476–484, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17319490.

Rui:2019:SAF

[RWZ^{+}19] Lanlan Rui, Xiaotong Wang, Yao Zhang, Xiaomei Wang, and Xuesong Qiu. A self-adaptive and fault-tolerant routing algorithm for wireless sensor networks in microgrids. Future Generation Computer Systems, 100(??):35– 45, November 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18302504.

Rajasekaran:2019:AMH

[RYH^{+}19] Manikandan Rajasekaran, Abdulsalam Yassine, M. Shamim Hossain, Mohammed F. Alhamid, and Mohsen Guizani. Autonomous monitoring in healthcare environment: Rewardbased energy charging mechanism for IoMT wireless sensing nodes. Future Generation Computer Systems, 98(??):565–576, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326347.

Ranaldo:2016:CDU

[RZ16] Nadia Ranaldo and Eugenio Zimeo. Capacitydriven utility model for service level agreement negotiation of cloud services. Future Generation Computer Systems, 55(??):186– 199, February 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500062X.

Sengupta:2014:MSD

[SA14] Shubhashis Sengupta and K. M. Annervaz. Multi-site data distribution for disaster recovery — a planning framework. Future Generation Computer Systems, 41(??):53–64, December 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1400140X.

Silva:2019:CSP

[SA19] Fábio Silva and Cesar Analide. Computational sustainability and the PHESS platform: Using affective computing as social indicators. Future Generation Computer Systems, 92(??):329–341, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17328856.

Sunercan:2011:SAI

[SAC11] H. Kevser Sunercan, M. Nedim Alpdemir, and Nihan Kesim Cicekli. A systematic approach to the integration of overlapping partitions in service-oriented data grids. Future Generation Computer Systems, 27(6):667– 680, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Souza:2019:MDD

[SAG19] Thiago I. A. Souza, Andre L. L. Aquino, and Danielo G. Gomes. A method to detect data outliers from smart urban spaces via tensor analysis. Future Generation Computer Systems, 92(??):290– 301, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18307556.

Sanchez-Arias:2017:MSC

[SAGGB17] Gonzalo Sánchez-Arias, Cristian González García, and B. Cristina Pelayo G-Bustelo. Midgar: Study of communications security among smart objects using a platform of heterogeneous devices for the Internet of Things. Future Generation Computer Systems, 74(??):444–466, September 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301632.

Sanchez-Artigas:2010:EPB

[SAGL10] Marc Sànchez-Artigas and Pedro García-López. eSciGrid: a P2P-based escience Grid for scalable and efficient data sharing. Future Generation Computer Systems, 26(5):704– 719, May 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Senouci:2019:MVM

[SAH19] Oussama Senouci, Zibouda Aliouat, and Saad Harous. MCA-V2I: a multi-hop clustering approach over vehicle-to-Internet communication for improving VANETs performances.

Future Generation Computer Systems, 96(??):309– 323, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18329315.

Somasundaram:2010:CRB

[SAK^{+}10] Thamarai Selvi Somasundaram, Balachandar R. Amarnath, R. Kumar, P. Balakrishnan, K. Rajendar, R. Rajiv, G. Kannan, G. Rajesh Britto, E. Mahendran, and B. Madusudhanan. CARE Resource Broker: a framework for scheduling and supporting virtual resource management. Future Generation Computer Systems, 26(3): 337–347, March 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shafeeq:2019:PAD

[SAK19] Sehrish Shafeeq, Masoom Alam, and Abid Khan. Privacy aware decentralized access control system. Future Generation Computer Systems, 101(??):420–433, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18332308.

Saracevic:2019:NAS

[SAM^{+}19] Muzafer Saracević, Sasa Adamović, Vladislav Miskovic, Nemanja Macek, and Marko Sarac. A novel approach to steganography based on the properties of Catalan numbers and Dyck words. Future Generation Computer Systems, 100(??):186–197, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19300184.

Smara:2017:ATF

[SAPA17] Mounya Smara, Makhlouf Aliouat, Al-Sakib Khan Pathan, and Zibouda Aliouat. Acceptance test for fault detection in component-based cloud computing and systems. Future Generation Computer Systems, 70(??):74– 93, May 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302151.

Sarier:2018:MBI

[Sar18a] Neyire Deniz Sarier. Multimodal biometric Identity Based Encryption. Future Generation Computer Systems, 80(??):112–125, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302613.

Singh:2018:SDD

[SAR18b] Priyanka Singh, Nishant Agarwal, and Balasubramanian Raman. Secure data deduplication using secret sharing schemes over cloud. Future Generation Computer Systems, 88(??):156–167, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17327474.

Sureshkumar:2019:RSC

[SAVS19] Venkatasamy Sureshkumar, Ruhul Amin, V. R. Vijaykumar, and S. Raja Sekar. Robust secure communication protocol for smart healthcare system with FPGA implementation. Future Generation Computer Systems, 100(??):938–951, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18332448.

Simmhan:2011:AAS

[SB11] Yogesh Simmhan and Roger Barga. Analysis of approaches for supporting the Open Provenance Model: a case study of the Trident workflow workbench. Future Generation Computer Systems, 27(6): 790–796, June 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sampaio:2014:THA

[SB14] Altino M. Sampaio and Jorge G. Barbosa. Towards high-available and energyefficient virtual computing environments in the cloud. Future Generation Computer Systems, 40(??):30– 43, November 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001307.

Sotiriadis:2016:ICB

[SB16] Stelios Sotiriadis and Nik Bessis. An inter-cloud bridge system for heterogeneous cloud platforms. Future Generation Computer Systems, 54(??):180– 194, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000400.

Singh:2017:MBS

[SB17a] Hari Singh and Seema Bawa. A MapReducebased scalable discovery and indexing of structured big data. Future Generation Computer Systems, 73(??):32–43, August 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17304880.

Singh:2017:APM

[SB17b] Kalpana Singh and Lynn Batten. Aggregating privatized medical data for secure querying applications. Future Generation Computer Systems, 72(??): 250–263, July 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16306732.

Singh:2018:EBS

[SB18] Amritpal Singh and Shalini Batra. Ensemble based spam detection in social IoT using probabilistic data structures. Future Generation Computer Systems, 81(??):359–371, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17313134.

Salkenov:2019:CBA

[SB19a] Aldiyar Salkenov and Susmit Bagchi. Cloud based autonomous monitoring and administration of heterogeneous distributed systems using mobile agents. Future Generation Computer Systems, 99(??):527– 557, October 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331194.

Savelyev:2019:GET

[SB19b] Alexey Savelyev and Emre Brookes. GenApp: Extensible tool for rapid generation of web and native GUI applications. Future Generation Computer Systems, 94(??):929–936, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310063.

Serrano:2019:DNN

[SB19c] Emilio Serrano and Javier Bajo. Deep neural network architectures for social services diagnosis in smart cities. Future Generation Computer Systems, 100(??):122–131, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301918.

Smowton:2017:CEA

[SBA^{+}17] Christopher Smowton, Andoena Balla, Demetris Antoniades, Crispin Miller, George Pallis, Marios D. Dikaiakos, and Wei Xing. A cost-effective approach to improving performance of big genomic data analyses in clouds. Future Generation Computer Systems, 67(??):368–381, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003593.

Senturk:2018:RPF

[SBAD^{+}18] Izzet F. Senturk, P. Balakrishnan, Anas Abu-Doleh, Kamer Kaya, Qutaibah Malluhi, and Ümit V. Çatalyürek. A resource provisioning framework for bioinformatics applications in multi-cloud environments. Future Generation Computer Systems, 78 (part 1)(??):379–391, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301911.

Schwiegelshohn:2010:PGC

[SBB^{+}10] Uwe Schwiegelshohn, Rosa M. Badia, Marian Bubak, Marco Danelutto, Schahram Dustdar, Fabrizio Gagliardi, Alfred Geiger, Ladislav Hluchy, Dieter Kranzlmüller, Erwin Laure, Thierry Priol, Alexander Reinefeld, Michael Resch, Andreas Reuter, Otto Rienhoff, Thomas Rüter, Peter Sloot, Domenico Talia, Klaus Ullmann, Ramin Yahyapour, and Gabriele von Voigt. Perspectives on Grid computing. Future Generation Computer Systems, 26(8): 1104–1115, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Saia:2016:BST

[SBCF16] Roberto Saia, Ludovico Boratto, Salvatore Carta, and Gianni Fenu. Binary sieves: Toward a semantic approach to user segmentation for behavioral targeting. Future Generation Computer Systems, 64(??):186–197, November 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300838.

Soille:2018:VDI

[SBD^{+}18] P. Soille, A. Burger, D. De Marchi, P. Kempeneers, D. Rodriguez, V. Syrris, and V. Vasilev. A versatile data-intensive computing platform for information retrieval from big geospatial data. Future Generation Computer Systems, 81(??):30–40, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1730078X.

Serrano:2016:SGC

[SBK^{+}16] Damián Serrano, Sara Bouchenak, Yousri Kouki, Frederico Alvares de Oliveira, Jr., Thomas Ledoux, Jonathan Lejeune, Julien Sopena, Luciana Arantes, and Pierre Sens. SLA guarantees for cloud services. Future Generation Computer Systems, 54(??):233– 246, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000801.

Subba:2018:GTB

[SBK18] Basant Subba, Santosh Biswas, and Sushanta Karmakar. A game theory based multi layered intrusion detection framework for VANET. Future Generation Computer Systems, 82(??):12–28, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17314486.

Sultan:2018:ISS

[SBL18] Nazatul Haque Sultan, Ferdous Ahmed Barbhuiya, and Maryline Laurent. ICAuth: a secure and scalable owner delegated intercloud authorization. Future Generation Computer Systems, 88(??):319–332, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17324251.

Shen:2014:AOB

[SBLW14] Jun Shen, Ghassan Beydoun, Graham Low, and Lijuan Wang. Aligning ontology-based development with service oriented systems. Future Generation Computer Systems, 32(??):263–273, March 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001702.

Sotiriadis:2017:VMC

[SBP^{+}17] Stelios Sotiriadis, Nik Bessis, Euripides G. M. Petrakis, Cristiana Amza, Catalin Negru, and Mariana Mocanu. Virtual machine cluster mobility in inter-cloud platforms. Future Generation Computer Systems, 74(??):179–189, September 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300206;
 https://www.math.utah.
edu/pub/tex/bib/virtual-
machines.bib.

Sanati:2016:LEO

[SC16] Behnaz Sanati and Albert M. K. Cheng. LBBA: an efficient online benefit-aware multiprocessor scheduling for QoS via online choice of approximation algorithms. Future Generation Computer Systems, 59(??):125– 135, June 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003374.

Saia:2019:EBU

[SC19] Roberto Saia and Salvatore Carta. Evaluating the benefits of using proactive transformeddomain-based techniques in fraud detection tasks. Future Generation Computer Systems, 93(??):18– 32, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18306423.

Sitton-Candanedo:2019:REC

[SCAC^{+}19] Inés Sittón-Candanedo, Ricardo S. Alonso, Juan M. Corchado, Sara Rodríguez-González, and Roberto Casado-Vara. A review of edge computing reference architectures and a new global edge proposal. Future Generation Computer Systems, 99(??):278– 294, October 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1930264X.

Starlinger:2016:EES

[SCBK^{+}16] Johannes Starlinger, Sarah Cohen-Boulakia, Sanjeev Khanna, Susan B. Davidson, and Ulf Leser. Effective and efficient similarity search in scientific workflow repositories. Future Generation Computer Systems, 56(??):584–594, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002228.

Su:2011:VSM

[SCCS11] Yen-Liang Su, Po-Cheng Chen, Jyh-Biau Chang, and Ce-Kuen Shieh. Variablesized map and localityaware reduce on publicresource grids. Future Generation Computer Systems, 27(6):843–849, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Su:2018:PDN

[SCEC18] Xin Su, Aniello Castiglione, Christian Esposito, and Chang Choi. Power domain NOMA to support group communication in public safety networks. Future Generation Computer Systems, 84(??):228–238, July 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1730571X.

Sun:2018:BDI

[SCG^{+}18] Gang Sun, Victor Chang, Steven Guan, Muthu Ramachandran, Jin Li, and Dan Liao. Big data and Internet of Things — fusion for different services and its impacts. Future Generation Computer Systems, 86(??):1368–1370, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18311282.

Sun:2017:AAS

[SCH^{+}17] Yanming Sun, Min Chen, Long Hu, Yongfeng Qian, and Mohammad Mehedi Hassan. ASA: Against statistical attacks for privacyaware users in Location Based Service. Future Generation Computer Systems, 70(??):48–58, May 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302023.

Sun:2019:CBM

[SCH^{+}19] Hui Sun, Guodong Chen, Jianzhong Huang, Xiao Qin, and Weisong Shi. CalmWPC: a buffer management to calm down write performance cliff for NAND flash-based storage systems. Future Generation Computer Systems, 90(??):461–476, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1830829X.

Stefanic:2019:SFL

[SCJ^{+}19a] Polona Stefanic, Matej Cigale, Andrew C. Jones, Louise Knight, and Ian Taylor. Support for full life cycle cloud-native application management: Dynamic TOSCA and SWITCH IDE. Future Generation Computer Systems, 101(??):975–982, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19302766.

Stefanic:2019:SWN

[ŠCJ^{+}19b] Polona Štefanič, Matej Cigale, Andrew C. Jones, Louise Knight, Ian Taylor, Cristiana Istrate, George Suciu, Alexandre Ulisses, Vlado Stankovski, Salman Taherizadeh, Guadalupe Flores Salado, Spiros Koulouzis, Paul Martin, and Zhiming Zhao. SWITCH workbench: a novel approach for the development and deployment of time-critical microservice-based cloudnative applications. Future Generation Computer Systems, 99(??):197–212, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1831094X.

Smari:2014:EAB

[SCL14] Waleed W. Smari, Patrice Clemente, and Jean-François Lalande. An extended attribute based access control model with trust and privacy: Application to a collaborative crisis management system. Future Generation Computer Systems, 31(??):147–168, February 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300109X.

Sahni:2018:MMW

[SCL18] Yuvraj Sahni, Jiannong Cao, and Xuefeng Liu. MidSHM: a middleware for WSN-based SHM application using serviceoriented architecture. Future Generation Computer Systems, 80(??):263–274, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301280.

Shin:2019:UMS

[SCLC19] Jinmyeong Shin, Seok-Hwan Choi, Peng Liu, and Yoon-Ho Choi. Unsupervised multi-stage attack detection framework without details on single-stage attacks. Future Generation Computer Systems, 100(??):811–825, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18329212.

Smith:2012:ONR

[SCMS12] Jay Smith, Edwin K. P.

Chong, Anthony A. Maciejewski, and Howard Jay Siegel. Overlay network resource allocation using a decentralized marketbased approach. Future Generation Computer Systems, 28(1):24–35, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001361.

Skillen:2014:OUM

[SCN^{+}14] Kerry-Louise Skillen, Liming Chen, Chris D. Nugent, Mark P. Donnelly, William Burns, and Ivar Solheim. Ontological user modelling and semantic rule-based reasoning for personalisation of Help-On-Demand services in pervasive environments. Future Generation Computer Systems, 34(??):97– 109, May 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300246X.

Shen:2018:LML

[SCS^{+}18] Jian Shen, Shaohua Chang, Jun Shen, Qi Liu, and Xingming Sun. A lightweight multi-layer authentication protocol for wireless body area networks. Future Generation Computer Systems, 78 (part 3)(??):956–963, January 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306963.

Shah:2018:EIA

[SCY^{+}18] Syed Bilal Shah, Zhe Chen, Fuliang Yin, Inam Ullah Khan, and Niqash Ahmad. Energy and interoperable aware routing for throughput optimization in clustered IoTwireless sensor networks. Future Generation Computer Systems, 81(??):372– 381, April 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310488.

Su:2014:EEA

[SCZ^{+}14] Jinshu Su, Dan Cao, Baokang Zhao, Xiaofeng Wang, and Ilsun You. ePASS: an expressive attributebased signature scheme with privacy and an unforgeability guarantee for the Internet of Things. Future Generation Computer Systems, 33(??):11– 18, April 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002331.

Si:2019:MRA

[SCZ^{+}19] Huayou Si, Zhihui Chen, Wei Zhang, Jian Wan, Jilin Zhang, and Neal N. Xiong. A member recognition approach for specific organizations based on relationships among users in social networking Twitter. Future Generation Computer Systems, 92(??):1009–1020, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17304764.

Sannino:2018:DLA

[SD18] G. Sannino and G. De Pietro. A deep learning approach for ECG-based heartbeat classification for arrhythmia detection. Future Generation Computer Systems, 86(??):446–455, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17324548.

Song:2011:WFI

[SDC11] Xudong Song, Wanchun Dou, and Jinjun Chen. A workflow framework for intelligent service composition. Future Generation Computer Systems, 27(5): 627–636, May 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Solano:2017:SVM

[SDDG17] A. Solano, N. Duro, R. Dormido, and P. González. Smart vending machines in the era of Internet of Things. Future Generation Computer Systems, 76(??):215–220, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16304757.

Sperhac:2019:MCG

[SDF^{+}19] Jeanette M. Sperhac, Robert L. DeLeon, Thomas R. Furlani, Steven M. Gallo, Martins Innus, Matthew D. Jones, Jeffrey T. Palmer, Abani Patra, Benjamin D. Plessinger, Ryan Rathsam, Nikolay Simakov, Joseph P. White, Rudra Chakraborty, and Gregary Dean. Managing computational gateway resources with XDMoD. Future Generation Computer Systems, 98(??):154–166, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18310732.

Sun:2019:FCS

[SDH^{+}19] Le Sun, Hai Dong, Omar Khadeer Hussain, Farookh Khadeer Hussain, and Alex X. Liu. A framework of cloud service selection with criteria interactions. Future Generation Computer Systems, 94(??):749–764, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18311579.

Srinivas:2019:GRC

[SDK19] Jangirala Srinivas, Ashok Kumar Das, and Neeraj Kumar. Government regulations in cyber security: Framework, standards and recommendations. Future Generation Computer Systems, 92(??):178–188, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18316753.

Severino:2018:ITC

[SDST18] Gerardo Severino, Guido D’Urso, Maddalena Scarfato, and Gerardo Toraldo. The IoT as a tool to combine the scheduling of the irrigation with the geostatistics of the soils. Future Generation Computer Systems, 82(??):268–273, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324408.

Selvitopi:2019:LAL

[SDTA19] Oguz Selvitopi, Gunduz Vehbi Demirci, Ata Turk, and Cevdet Aykanat. Locality-aware and loadbalanced static task scheduling for MapReduce. Future Generation Computer Systems, 90(??):49–61, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17329266.

Sarwar:2013:TVR

[SDWS13] Muhammad S. Sarwar, T. Doherty, J. Watt, and Richard O. Sinnott. Towards a virtual research environment for language and literature researchers. Future Generation Computer Systems, 29(2):549– 559, February 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000805.

Sivagami:2019:IDF

[SE19] V. M. Sivagami and K. S.

Easwarakumar. An improved dynamic fault tolerant management algorithm during VM migration in cloud data center.

Future Generation Computer Systems, 98(??):35– 43, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18324208.

Shahmoradi:2019:MOC

[SEHS19] M. R. Shahmoradi, M. Ebrahimi, Z. Heshmati, and M. Salehi. Multilayer overlapping community detection using multi-objective optimization. Future Generation Computer Systems, 101(??):221–235, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19301578.

Sehgal:2011:UAL

[SEMJ11] Saurabh Sehgal, Miklos Erdelyi, Andre Merzky, and Shantenu Jha. Understanding application-level interoperability: Scalingout MapReduce over highperformance grids and clouds. Future Generation Computer Systems, 27(5): 590–599, May 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Simao:2019:GWS

[SEPV19] J. Simão, S. Esteves, André Pires, and L. Veiga.

GC-Wise: a self-adaptive approach for memoryperformance efficiency in Java VMs. Future Generation Computer Systems, 100(??):674–688, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18304898.

Salza:2019:SGA

[SF19] Pasquale Salza and Filomena Ferrucci. Speed up genetic algorithms in the cloud using software containers. Future Generation Computer Systems, 92(??):276–289, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324147.

Simonet:2015:ADP

[SFR15] Anthony Simonet, Gilles Fedak, and Matei Ripeanu. Active data: a programming model to manage data life cycle across heterogeneous systems and infrastructures. Future Generation Computer Systems, 53(??):25–42, December 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15001995.

Smaoui:2013:IVC

[SG13] Malek Smaoui and Marc Garbey. Improving volunteer computing scheduling for evolutionary algorithms. Future Generation Computer Systems, 29(1): 1–14, January 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000921.

Somasundaram:2014:CFS

[SG14] Thamarai Selvi Somasundaram and Kannan Govindarajan. CLOUDRB: a framework for scheduling and managing highperformance computing (HPC) applications in science cloud. Future Generation Computer Systems, 34 (??):47–65, May 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002884.

Subirats:2015:AFE

[SG15] Josep Subirats and Jordi Guitart. Assessing and forecasting energy efficiency on Cloud computing platforms. Future Generation Computer Systems, 45(??):70–94, April 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002428.

Sethi:2017:SWD

[SG17] Ricky J. Sethi and Yolanda Gil. Scientific workflows in data analysis: Bridging expertise across multiple domains. Future Generation Computer Systems, 75(??):256–270, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300195.

Sperhac:2019:VHG

[SG19] Jeanette M. Sperhac and Steven M. Gallo. VIDIA: a HUBzero gateway for data analytics education. Future Generation Computer Systems, 94(??):833–840, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17308531.

Singh:2018:BFB

[SGB^{+}18] Amritpal Singh, Sahil Garg, Shalini Batra, Neeraj Kumar, and Joel J. P. C. Rodrigues. Bloom filter based optimization scheme for massive data handling in IoT environment. Future Generation Computer Systems, 82(??):440–449, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17314516.

Singh:2019:PDS

[SGBK19] Amritpal Singh, Sahil Garg, Shalini Batra, and Neeraj Kumar. Probabilistic data structurebased community detection and storage scheme in online social networks. Future Generation Computer Systems, 94(??):173–184, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310300.

Guan:2019:MIF

[sGbKS19] Jian sheng Guan, Shao bo Kang, and Yuan Sun. Medical image fusion algorithm based on multi-resolution analysis coupling approximate spare representation. Future Generation Computer Systems, 98(??):201–207, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326396.

Sanchez-Garcia:2016:SSA

[SGGCR^{+}16] J. Sánchez-García, J. M.

García-Campos, D. G. Reina, S. L. Toral, and F. Barrero. On-siteDriverID: a secure authentication scheme based on Spanish eID cards for vehicular ad hoc networks. Future Generation Computer Systems, 64(??):50– 60, November 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301121.

Singh:2018:NCE

[SGJ18] Vishakha Singh, Indrajeet Gupta, and Prasanta K. Jana. A novel cost-efficient approach for deadline-constrained workflow scheduling by dynamic provisioning of resources. Future Generation Computer Systems, 79 (part 1) (??):95–110, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17307264.

Secretan:2010:AAP

[SGKC10] Jimmy Secretan, Michael Georgiopoulos, Anna Koufakou, and Kel Cardona. APHID: An architecture for private, highperformance integrated data mining. Future Generation Computer Systems, 26(7):891–904, July 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sun:2019:SRA

[SGL^{+}19] Dawei Sun, Shang Gao, Xunyun Liu, Fengyun Li, Xinqi Zheng, and Rajkumar Buyya. State and runtime-aware scheduling in elastic stream computing systems. Future Generation Computer Systems, 97(??):194–209, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18321897.

Simmhan:2011:SST

[SGM11] Yogesh Simmhan, Paul Groth, and Luc Moreau. Special section: The third provenance challenge on using the open provenance model for interoperability. Future Generation Computer Systems, 27(6):737– 742, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shuja:2017:CAE

[SGN^{+}17] Junaid Shuja, Abdullah Gani, Anjum Naveed, Ejaz Ahmed, and Ching-Hsien Hsu. Case of ARM emulation optimization for offloading mechanisms in Mobile Cloud Computing. Future Generation Computer Systems, 76(??):407– 417, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301558.

Sanchez-Garcia:2019:DPB

[SGRT19] J. Sánchez-García, D. G.

Reina, and S. L. Toral. A distributed PSO-based exploration algorithm for a UAV network assisting a disaster scenario. Future Generation Computer Systems, 90(??):129–148, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18303649.

Song:2018:QFS

[SGS^{+}18] Jun Song, Kun Gao, Xinyang Shen, Xiaotian Qi, Rui Liu, and Kim-Kwang Raymond Choo. QRFence: a flexible and scalable QR link security detection framework for Android devices. Future Generation Computer Systems, 88(??):663–674, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17324160.

Silva:2019:MOI

[SH19] Bhagya Nathali Silva and Kijun Han. Mutation operator integrated ant colony optimization based domestic appliance scheduling for lucrative demand side management. Future Generation Computer Systems, 100(??):557–568, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19301141.

Sharma:2016:ECP

[Sha16]  Sugam Sharma.  Ex-

panded cloud plumes hiding Big Data ecosystem. Future Generation Computer Systems, 59(??):63– 92, June 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16000054.

Salvadores:2010:SCA

[SHBP10] Manuel Salvadores, Pilar Herrero, José Luis Bosque, and María S. Peréz. A semantic collaborative awareness model to deal with resource sharing in grids. Future Generation Computer Systems, 26(2): 276–280, February 2010.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Somrak:2019:EVS

[SHH^{+}19] Andrej Somrak, Iztok Humar, M. Shamim Hossain, Mohammed F. Alhamid, M. Anwar Hossain, and Jože Guna. Estimating VR sickness and user experience using different HMD technologies: an evaluation study. Future Generation Computer Systems, 94(??): 302–316, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18325044.

Spencer:2010:RED

[SHJS^{+}10] R. E. Spencer, R. Hughes-Jones, M. Strong, S. Casey, A. Rushton, P. Burgess, S. Kershaw, and C. Greenwood. The role of ES-LEA in the development of eVLBI. Future Generation Computer Systems, 26 (1):111–119, January 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shi:2019:AMW

[SHL^{+}19a] Guolong Shi, Yigang He, Bing Li, Lei Zuo, Baiqiang Yin, Wenbo Zeng, and Farhan Ali. Analysis and modeling of wireless channel characteristics for Internet of Things scene based on geometric features. Future Generation Computer Systems, 101(??):492–501, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19300470.

Sun:2019:TTR

[SHL^{+}19b] Penghao Sun, Yuxiang Hu, Julong Lan, Le Tian, and Min Chen. TIDE: Timerelevant deep reinforcement learning for routing optimization. Future Generation Computer Systems, 99(??):401–409, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19305424.

Shao:2013:VOS

[SHLJ13] Zhiyuan Shao, Ligang He, Zhiqiang Lu, and Hai Jin. VSA: an offline scheduling analyzer for Xen virtual machine monitor. Future Generation Computer Systems, 29(8):2067–2076, October 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12002245.

Schnorr:2010:TIV

[SHN10] Lucas Mello Schnorr, Guillaume Huard, and Philippe O. A. Navaux. Triva: Interactive 3D visualization for performance analysis of parallel applications. Future Generation Computer Systems, 26(3):348– 358, March 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sahuquillo:2016:DET

[SHP^{+}16] Julio Sahuquillo, Houcine Hassan, Salvador Petit, José Luis March, and José Duato. A dynamic execution time estimation model to save energy in heterogeneous multicores running periodic tasks. Future Generation Computer Systems, 56(??):211–219, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002216.

Sedaghat:2016:DCD

[SHRE16] Mina Sedaghat, Francisco Hernández-Rodriguez, and Erik Elmroth. Decentralized cloud datacenter reconsolidation through emergent and topologyaware behavior. Future Generation Computer Systems, 56(??):51–63, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003039.

Sanin:2019:EBK

[SHS^{+}19] Cesar Sanin, Zhang Haoxi, Imran Shafiq, Md Maqbool Waris, Caterine Silva de Oliveira, and Edward Szczerbicki. Experience based knowledge representation for Internet of Things and cyber physical systems with case studies. Future Generation Computer Systems, 92(??):604– 616, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17316965.

Shah:2018:PCI

[SI18] Syed Ali Raza Shah and Biju Issac. Performance comparison of intrusion detection systems and application of machine learning to Snort system. Future Generation Computer Systems, 80(??):157–170, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17323178.

Saito:2019:HMD

[SI19] Kenji Saito and Mitsuru Iwamura. How to make a digital currency on a blockchain stable. Future Generation Computer Systems, 100(??):58–69, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18300475.

Smanchat:2013:SPS

[SIL^{+}13] Sucha Smanchat, Maria Indrawan, Sea Ling, Colin Enticott, and David Abramson. Scheduling parameter sweep workflow in the Grid based on resource competition. Future Generation Computer Systems, 29(5): 1164–1183, July 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000198.

Sipos:2012:PCW

[Sip12] Gergely Sipos. Protecting the consistency of workflow applications in collaborative development environments. Future Generation Computer Systems, 28(3): 500–512, March 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001695.

Sanchez-Iborra:2018:EIN

[SISGS18] Ramon Sanchez-Iborra, Jesus Sanchez-Gomez, and Antonio Skarmeta. Evolving IoT networks by the confluence of MEC and LP–WAN paradigms. Future Generation Computer Systems, 88(??):199–208, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17324159.

Saginbekov:2014:ECD

[SJ14] Sain Saginbekov and Arshad Jhumka. Efficient code dissemination in wireless sensor networks. Future Generation Computer Systems, 39(??):111–119, October 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002720.

Shmeis:2018:FCG

[SJ18] Zeinab Shmeis and Mohamad Jaber. Fine and coarse grained composition and adaptation of spark applications. Future Generation Computer Systems, 86(??):629–640, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17320873.

Shmeis:2019:RBO

[SJ19] Zeinab Shmeis and Mohamad Jaber. A rewritebased optimizer for spark. Future Generation Computer Systems, 98(??):586– 599, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18311944.

Son:2017:TLT

[SJL^{+}17] Yunsik Son, MyoungHwan Joung, Yong-Wook Lee, Oh-Heum Kwon, and Ha-Joo Song. Tag localization in a two-dimensional RFID tag matrix. Future Generation Computer Systems, 76(??):384–390, November 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300619.

Sun:2018:PSO

[SJL^{+}18] Hanlin Sun, Wei Jie, Jonathan Loo, Lizhe Wang, Sugang Ma, Gang Han, Zhongmin Wang, and Wei Xing. A parallel self-organizing overlapping community detection algorithm based on swarm intelligence for large scale complex networks. Future Generation Computer Systems, 89(??):265–285, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18301328.

Sartakhti:2013:NLB

[SJR13] Javad Salimi Sartakhti, Saeed Jalili, and Ali Gholami Rudi. A new lightbased solution to the Hamiltonian path problem. Future Generation Computer Systems, 29(2):520– 527, February 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001562.

Syed:2019:SHF

[SJSA19] Liyakathunisa Syed, Saima Jabeen, Manimala S., and Abdullah Alsaeedi. Smart healthcare framework for ambient assisted living using IoMT and big data analytics techniques. Future Generation Computer Systems, 101(??):136–151, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18321071.

Shao:2018:ELB

[SJTN18] Xun Shao, Masahiro Jibiki, Yuuichi Teranishi, and Nozomu Nishinaga. An efficient load-balancing mechanism for heterogeneous range-queriable cloud storage. Future Generation Computer Systems, 78 (part 3)(??):920–930, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16305581.

Srirama:2012:ASC

[SJV12] Satish Narayana Srirama, Pelle Jakovits, and Eero Vainikko. Adapting scientific computing problems to clouds using MapReduce. Future Generation Computer Systems, 28(1):184–192, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001075.

Schulte:2015:EBP

[SJV^{+}15] Stefan Schulte, Christian Janiesch, Srikumar Venugopal, Ingo Weber, and Philipp Hoenisch. Elastic Business Process Management: State of the art and open challenges for BPM in the cloud. Future Generation Computer Systems, 46 (??):36–50, May 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1400168X.

Stavrinides:2012:SRT

[SK12] Georgios L. Stavrinides and Helen D. Karatza. Scheduling real-time DAGs in heterogeneous clusters by combining imprecise computations and bin packing techniques for the exploitation of schedule holes. Future Generation Computer Systems, 28(7): 977–988, July 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12000556.

Sengupta:2018:FER

[SK18] Anirban Sengupta and Deepak Kachave. Forensic engineering for resolving ownership problem of reusable IP core generated during high level synthesis. Future Generation Computer Systems, 80(??):29– 46, March 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302108.

Stavrinides:2019:EEQ

[SK19] Georgios L. Stavrinides and Helen D. Karatza. An energy-efficient, QoS-aware and cost-effective scheduling approach for real-time workflow applications in cloud computing systems utilizing DVFS and approximate computations. Future Generation Computer Systems, 96(??):216– 226, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18327353.

Shirai:2011:MPL

[SKF^{+}11] Daisuke Shirai, Masahiko Kitamura, Tatsuya Fujii, Atsushi Takahara, Kunitake Kaneko, and Naohisa Ohta. Multi-point 4K/2K layered video streaming for remote collaboration. Future Generation Computer Systems, 27(7):986– 990, July 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Skowronski:2019:OBA

[Sko19] Rafał Skowroński. The open blockchain-aided multiagent symbiotic cyberphysical systems. Future Generation Computer Systems, 94(??):430–443, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18307520.

Somu:2017:CMR

[SKS17] Nivethitha Somu, Kannan Kirthivasan, and Shankar Sriram V. S. A computational model for ranking cloud service providers using hypergraph based techniques. Future Generation Computer Systems, 68(??): 14–30, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302734.

Singh:2018:MWT

[SKS^{+}18] Amit Kumar Singh, Basant Kumar, Sanjay Kumar Singh, S. P. Ghrera, and Anand Mohan. Multiple watermarking technique for securing online social network contents using back propagation neural network. Future Generation Computer Systems, 86(??):926–939, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16306628.

Seneviratne:2011:TPM

[SL11] Sena Seneviratne and David C. Levy. Task profiling model for load profile prediction. Future Generation Computer Systems, 27 (3):245–255, March 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Son:2019:TIN

[SL19] Heesuk Son and Dongman Lee. Towards interactive networking: Runtime message inference approach for incompatible protocol updates in IoT environments. Future Generation Computer Systems, 96(??):563– 578, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18319460.

Sun:2016:NTE

[SLA^{+}16] Gang Sun, Dan Liao, Vishal Anand, Dongcheng Zhao, and Hongfang Yu. A new technique for efficient live migration of multiple virtual machines. Future Generation Computer Systems, 55(??):74–86, February 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002848.

Sun:2017:EFA

[SLB^{+}17] Gang Sun, Dan Liao, Sitong Bu, Hongfang Yu, Zhili Sun, and Victor Chang. The efficient framework and algorithm for provisioning evolving VDC in federated data centers.

Future Generation Computer Systems, 73(??):79– 89, August 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308019.

Silva:2017:RDQ

[SLC^{+}17] Vítor Silva, José Leite, José J. Camata, Daniel de Oliveira, Alvaro L. G. A. Coutinho, Patrick Valduriez, and Marta Mattoso. Raw data queries during data-intensive parallel workflow execution. Future Generation Computer Systems, 75(??):402– 422, October 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300237.

Shen:2015:SMD

[SLD^{+}15] Dian Shen, Junzhou Luo, Fang Dong, Xiang Fei, Wei Wang, Guoqing Jin, and Weidong Li. Stochastic modeling of dynamic rightsizing for energy-efficiency in cloud data centers. Future Generation Computer Systems, 48(??):ibc, July 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001824.

Sun:2018:FSI

[SLD^{+}18] Guanglu Sun, Jiabin Li, Jian Dai, Zhichao Song, and Fei Lang. Feature selection for IoT based on maximal information coefficient. Future Generation Computer Systems, 89(??):606–616, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18304837.

Slezak:2014:SSA

[Śle14]  Dominik Ślezak.  Spe-

cial section on applications of intelligent data and knowledge processing technologies — preface. Future Generation Computer Systems, 33(??): 19–20, April 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002409.

Shi:2017:PDG

[SLG^{+}17] Zhan Shi, Junhao Li, Pengfei Guo, Shuangshuang Li, Dan Feng, and Yi Su. Partitioning dynamic graph asynchronously with distributed FENNEL. Future Generation Computer Systems, 71(??):32–42, June 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1730033X.

Su:2019:TUT

[SLH^{+}19] Yuhan Su, Xiaozhen Lu, Lianfen Huang, Xiaojiang Du, and Mohsen Guizani. Tac-U: a traffic balancing scheme over licensed and unlicensed bands for Tactile Internet. Future Generation Computer Systems, 97(??):41–49, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18317291.

Seol:2017:IMO

[SLK17] Soonuk Seol, Eun-Kyu Lee, and Wooseong Kim. Indoor mobile object tracking using RFID. Future Generation Computer Systems, 76(??):443–451, November 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302552.

Sun:2017:LLL

[SLL^{+}17] Gang Sun, Dan Liao, Hui Li, Hongfang Yu, and Victor Chang. L2P2: a location-label based approach for privacy preserving in LBS. Future Generation Computer Systems, 74(??):375–384, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302953.

Sun:2018:LLO

[SLL^{+}18] Gang Sun, Yayu Li, Yao Li, Dan Liao, and Victor Chang. Low-latency orchestration for workfloworiented service function chain in edge computing. Future Generation Computer Systems, 85(??):116– 128, August 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18300153.

Shi:2010:SHS

[SLS10] Fei Shi, Keqiu Li, and Yanming Shen. Seamless handoff scheme in Wi-Fi and WiMAX/heterogeneous networks. Future Generation Computer Systems, 26(8): 1403–1408, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sodhro:2019:AIB

[SLS^{+}19] Ali Hassan Sodhro, Zongwei Luo, Gul Hassan Sodhro, Muhammad Muzamal, Joel J. P. C. Rodrigues, and Victor Hugo C. de Albuquerque. Artificial intelligence based QoS optimization for multimedia communication in IoV systems. Future Generation Computer Systems, 95(??): 667–680, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18320314.

Sipos:2019:EAD

[SLSS19] Gergely Sipos, Giuseppe La Rocca, Diego Scardaci, and Peter Solagna. The EGI applications on demand service. Future Generation Computer Systems, 98(??):171–179, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314481.

Sakai:2019:RSE

[SLTK19] Yuki Sakai, Huimin Lu, Joo-Kooi Tan, and Hyoungseop Kim. Recognition of surrounding environment from electric wheelchair videos based on modified YOLOv2. Future Generation Computer Systems, 92(??):157–161, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18315887.

Sun:2011:FED

[SLW11] Xiaoxun Sun, Min Li, and Hua Wang. A family of enhanced (L, α)-diversity models for privacy preserving data publishing. Future Generation Computer Systems, 27(3):348– 356, March 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sun:2019:EET

[SLY^{+}19] Gang Sun, Yayu Li, Hongfang Yu, Athanasios V. Vasilakos, Xiaojiang Du, and Mohsen Guizani. Energy-efficient and traffic-aware service function chaining orchestration in multi-domain networks. Future Generation Computer Systems, 91(??):347–360, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1831848X.

Sun:2018:TPH

[SLZ^{+}18] Gang Sun, Dan Liao, Dongcheng Zhao, Zhili Sun, and Victor Chang. Towards provisioning hybrid virtual networks in federated cloud data centers. Future Generation Computer Systems, 87(??):457–469, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17313602.

Saito:2010:BSR

[SM10] Kenji Saito and Eiichi Morino. The brighter side of risks in peer-to-peer barter relationships. Future Generation Computer Systems, 26(8):1300–1316, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sood:2018:FCB

[SM18] Sandeep K. Sood and Isha Mahajan. Fog-cloud based cyber-physical system for distinguishing, detecting and preventing mosquito borne diseases. Future Generation Computer Systems, 88(??):764– 775, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17315868.

Souza:2018:TPS

[SMBMT^{+}18] V. B. Souza, X. Masip-Bruin, E. Marín-Tordera, S. Sànchez-López, J. Garcia, G. J. Ren, A. Jukan, and A. Juan Ferrer. Towards a proper service placement in combined Fog-to-Cloud (F2C) architectures. Future Generation Computer Systems, 87(??):1–15, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17323051.

Sanjuan:2018:WFP

[SMC18] Gemma Sanjuan, Tomàs Margalef, and Ana Cortés. Wind field parallelization based on Schwarz alternating domain decomposition method. Future Generation Computer Systems, 82(??):565–574, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1630872X.

Sahay:2019:CID

[SME^{+}19] Rishikesh Sahay, Weizhi Meng, D. A. Sepulveda Estay, Christian D. Jensen, and Michael Bruhn Barfod. CyberShip-IoT: a dynamic and adaptive SDN-based security policy enforcement framework for ships. Future Generation Computer Systems, 100(??):736–750, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1930367X.

See corrigendum [SME^{+}21].

Sahay:2021:CCI

[SME^{+}21] Rishikesh Sahay, Weizhi Meng, D. A. Sepulveda Estay, Christian D. Jensen, and Michael Bruhn Barfod. Corrigendum to “CyberShip-IoT: a Dynamic and Adaptive SDN-Based Security Policy Enforcement Framework for Ships” [Future Gener. Comput. Syst. 100 (2019) 736–750]. Future Generation Computer Systems, 118(??):492–494, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20326674.

See [SME^{+}19].

Safa:2019:DPB

[SMF^{+}19] Nader Sohrabi Safa, Carsten Maple, Steve Furnell, Muhammad Ajmal Azad, Charith Perera, Mohammad Dabbagh, and Mehdi Sookhak. Deterrence and prevention-based model to mitigate information security insider threats in organisations. Future Generation Computer Systems, 97(??):587–597, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331285.

Shang:2018:SAA

[SMG18] Fengjun Shang, Lin Mao, and Wenjuan Gong. Serviceaware adaptive link load balancing mechanism for software-defined networking. Future Generation Computer Systems, 81(??): 452–464, April 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1730403X.

Soares:2014:GBA

[SMM^{+}14] Christophe Soares, Rui S.

Moreira, Ricardo Morla, José Torres, and Pedro Sobral. A graph-based approach for interference free integration of commercial off-the-shelf elements in pervasive computing systems. Future Generation Computer Systems, 39(??): 3–15, October 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002999.

Sanchez:2012:AFE

[SMPC12] Alberto Sánchez, Jesús Montes, María S. Pérez, and Toni Cortes. An autonomic framework for enhancing the quality of data grid services. Future Generation Computer Systems, 28(7):1005–1016, July 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11002160.

Sarkar:2013:RRP

[SMRM13] Madhulina Sarkar, Triparna Mondal, Sarbani Roy, and Nandini Mukherjee. Resource requirement prediction using clone detection technique. Future Generation Computer Systems, 29(4):936–952, June 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001835.

Spillner:2013:COC

[SMS13] Josef Spillner, Johannes Müller, and Alexander Schill. Creating optimal cloud storage systems. Future Generation Computer Systems, 29(4):1062–1072, June 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001409.

Shakshuki:2014:WCP

[SMS14a] Elhadi M. Shakshuki, Haroon Malik, and Tarek Sheltami. WSN in cyber physical systems: Enhanced energy management routing approach using software agents. Future Generation Computer Systems, 31(??):93–104, February 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000502.

Smith:2014:MSR

[SMS14b] Jay Smith, Anthony A.

Maciejewski, and Howard Jay Siegel. Maximizing stochastic robustness of static resource allocations in a periodic sensor driven cluster. Future Generation Computer Systems, 33(??): 1–10, April 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002112.

Sheltami:2016:DCT

[SMS16] Tarek Sheltami, Muhammad Musaddiq, and Elhadi Shakshuki. Data compression techniques in wireless sensor networks. Future Generation Computer Systems, 64(??):151–162, November 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16000285.

Santana:2019:NBA

[SMS^{+}19] Clodomir J. Santana, Mariana Macedo, Hugo Siqueira, Anu Gokhale, and Carmelo J. A. Bastos-Filho. A novel binary artificial bee colony algorithm. Future Generation Computer Systems, 98(??):180– 196, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18325652.

Silva:2018:AGF

[SMSF18] Edelberto Franco Silva, Débora Christina Muchaluat-Saade, and Natalia Castro Fernandes. ACROSS: a generic framework for attribute-based access control with distributed policies for virtual organizations. Future Generation Computer Systems, 78 (part 1)(??):1–17, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17316060.

Sun:2016:CFF

[SMZ^{+}16] Le Sun, Jiangang Ma, Yanchun Zhang, Hai Dong, and Farookh Khadeer Hussain. Cloud-FuSeR: Fuzzy ontology and MCDM based cloud service selection. Future Generation Computer Systems, 57(??):42– 55, April 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003829.

Silva:2018:JML

[SNC18] Bruno Silva, Marco A. S.

Netto, and Renato L. F. Cunha. JobPruner: A machine learning assistant for exploring parameter spaces in HPC applications. Future Generation Computer Systems, 83(??):144–157, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17323294.

Stan:2019:CCG

[SNP19] Roxana Gabriela Stan, Catalin Negru, and Florin Pop. CloudWave: Content gathering network with flying clouds. Future Generation Computer Systems, 98(??):474–486, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18327705.

Snasel:2017:GTA

[SNXB17] Václav Snásel, Jana Nowaková, Fatos Xhafa, and Leonard Barolli. Geometrical and topological approaches to big data. Future Generation Computer Systems, 67(??):286–296, February 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301856.

Scheider:2017:WGD

[SOA17] Simon Scheider, Frank O. Ostermann, and Benjamin Adams. Why good data analysts need to be critical synthesists. determining the role of semantics in data analysis. Future Generation Computer Systems, 72(??):11– 22, July 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17303047.

Sanchez-Oro:2018:IGA

[SOD18] Jesús Sánchez-Oro and Abraham Duarte. Iterated greedy algorithm for performing community detection in social networks. Future Generation Computer Systems, 88(??):785– 791, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17323932.

Shokripour:2012:NMS

[SOIS12] Amin Shokripour, Mohamed Othman, Hamidah Ibrahim, and Shamala Subramaniam. New method for scheduling heterogeneous multi-installment systems. Future Generation Computer Systems, 28(8):1205–1216, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000611.

Suliman:2019:GPI

[SOM^{+}19] Ahmed Suliman, Hadi Otrok, Rabeb Mizouni, Shakti Singh, and Anis Ouali. A greedy-proof incentive-compatible mechanism for group recruitment in mobile crowd sensing. Future Generation Computer Systems, 101 (??):1158–1167, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19303619.

Sapountzi:2018:SND

[SP18a] Androniki Sapountzi and Kostas E. Psannis. Social networking data analysis tools & challenges. Future Generation Computer Systems, 86(??):893–913, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1630423X.

Sharma:2018:BBH

[SP18b] Pradip Kumar Sharma and Jong Hyuk Park. Blockchain based hybrid network architecture for the smart city. Future Generation Computer Systems, 86(??):650–655, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1830431X.

Santos:2019:ZRA

[SPD^{+}19] Igor L. Santos, Luci Pirmez, Flavia C. Delicato, Gabriel M. Oliveira, Claudio M. Farias, Samee U. Khan, and Albert Y. Zomaya. Zeus: A resource allocation algorithm for the cloud of sensors. Future Generation Computer Systems, 92(??):564– 581, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312761.

Santana-Perez:2017:REE

[SPdSR^{+}17] Idafen Santana-Perez, Rafael Ferreira da Silva, Mats Rynge, Ewa Deelman, María S. Pérez-Hernández, and Oscar Corcho. Reproducibility of execution environments in computational science using semantics and clouds. Future Generation Computer Systems, 67(??):354–367, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16000029.

Satria:2017:ROM

[SPJ17] Dimas Satria, Daihee Park, and Minho Jo. Recovery for overloaded mobile edge computing. Future Generation Computer Systems, 70(??):138–147, May 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302096.

Stergiou:2018:SII

[SPKG18] Christos Stergiou, Kostas E.

Psannis, Byung-Gyu Kim, and Brij Gupta. Secure integration of IoT and cloud computing. Future Generation Computer Systems, 78 (part 3)(??):964–975, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630694X.

Sanchez:2010:HPS

[SPMC10] Alberto Sánchez, María S. Pérez, Jesús Montes, and Toni Cortes. A high performance suite of data services for grids. Future Generation Computer Systems, 26(4):622–632, April 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shi:2010:AGA

[SPR^{+}10] Xuanhua Shi, Jean-Louis Pazat, Eric Rodriguez, Hai Jin, and Hongbo Jiang. Adapting grid applications to safety using faulttolerant methods: Design, implementation and evaluations. Future Generation Computer Systems, 26(2): 236–244, February 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sodhro:2018:CIP

[SPS18] Ali Hassan Sodhro, Sandeep Pirbhulal, and Arun Kumar Sangaiah. Convergence of IoT and product lifecycle management in medical health care. Future Generation Computer Systems, 86(??):380–391, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17328509.

Sirbu:2017:PPB

[SPSP17] Alexandru Sîrbu, Cristian Pop, Cristina Serbanescu, and Florin Pop. Predicting provisioning and booting times in a metalas-a-service system. Future Generation Computer Systems, 72(??):180–192, July 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630231X.

Sahoo:2018:EDL

[SPT^{+}18] Kshira Sagar Sahoo, Deepak Puthal, Mayank Tiwary, Joel J. P. C. Rodrigues, Bibhudatta Sahoo, and Ratnakar Dash. An early detection of low rate DDoS attack to SDN based data center networks using information distance metrics. Future Generation Computer Systems, 89(??):685– 697, December 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18309543.

Saadat:2012:PNP

[SR12] Nazanin Saadat and Amir Masoud Rahmani. PDDRA: a new pre-fetching based dynamic data replication algorithm in data grids. Future Generation Computer Systems, 28(4):666– 681, April 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002081.

Sousa:2019:DLD

[SR19] Maria José Sousa and Álvaro Rocha. Digital learning: Developing skills for digital transformation of organizations. Future Generation Computer Systems, 91(??):327–334, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18311191.

Sanchez:2019:AIE

[SRdlPG19] Antonio J. Sánchez, Sara Rodríguez, Fernando de la Prieta, and Alfonso González. Adaptive interface ecosystems in smart cities control systems. Future Generation Computer Systems, 101(??):605–620, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302808.

Somu:2018:TCO

[SRKS18] Nivethitha Somu, Gauthama Raman M. R., Kannan Kirthivasan, and Shankar Sriram V. S. A trust centric optimal service ranking approach for cloud service selection. Future Generation Computer Systems, 86(??):234–252, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18301262.

Spivak:2018:STA

[SRN^{+}18] Anton Spivak, Andrew Razumovskiy, Denis Nasonov, Alexander Boukhanovsky, and Anton Redice. Storage tier-aware replicative data reorganization with prioritization for efficient workload processing. Future Generation Computer Systems, 79 (part 2)(??): 618–629, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17305502.

Sharma:2019:MLB

[SRP19] Pradip Kumar Sharma, Shailendra Rathore, and Jong Hyuk Park. Multilevel learning based modeling for link prediction and users’ consumption preference in online social networks. Future Generation Computer Systems, 93(??):952–961, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17308014.

Sajjad:2015:SDA

[SRZD15] Ali Sajjad, Muttukrishnan Rajarajan, Andrea Zisman, and Theo Dimitrakos. A scalable and dynamic application-level secure communication framework for inter-cloud services. Future Generation Computer Systems, 48(??): ibc, July 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000357.

Salimi:2013:BSC

[SS13] Hadi Salimi and Mohsen Sharifi. Batch scheduling of consolidated virtual machines based on their workload interference model. Future Generation Computer Systems, 29(8):2057– 2066, October 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000332.

Singh:2017:CBM

[SS17] Sarbjeet Singh and Jagpreet Sidhu. Compliancebased multi-dimensional trust evaluation system for determining trustworthiness of cloud service providers. Future Generation Computer Systems, 67(??):109–132, February 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630262X.

Shah:2019:CUC

[SSA^{+}19] Saleh Shah, Babar Shah, Adnan Amin, Feras Al-Obeidat, Francis Chow, Fernando Joaquim Lopes Moreira, and Sajid Anwar. Compromised user credentials detection in a digital enterprise using behavioral analytics. Future Generation Computer Systems, 93(??):407–417, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18312524.

Syed:2013:PGC

[SSB13] Raheel Hassan Syed, Maxime Syrame, and Julien Bourgeois. Protecting grids from cross-domain attacks using security alert sharing mechanisms. Future Generation Computer Systems, 29(2):536–547, February 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001501.

Souza:2019:KTU

[SSC^{+}19] Renan Souza, Vítor Silva, Jose J. Camata, Alvaro L. G. A. Coutinho, Patrick Valduriez, and Marta Mattoso. Keeping track of user steering actions in dynamic workflows. Future Generation Computer Systems, 99(??):624–643, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18308240.

Sanchez-Sepulveda:2019:VII

[SSFFR19] Mónica Sanchez-Sepulveda, David Fonseca, Jordi Franquesa, and Ernesto Redondo. Virtual interactive innovations applied for digital urban transformations. mixed approach. Future Generation Computer Systems, 91(??):371–381, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18313372.

Sassi:2017:DBN

[SSG17] Najla Sassi, Kais Ben Salah, and Khaled Ghédira. DOC–BRelax: a new multi-agent system to solve distributed constraint optimization problems. Future Generation Computer Systems, 73(??):44–51, August 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630108X.

Said:2019:TAN

[SSG19] Mostafa Said, Ahmed Shalaby, and Fayez Gebali. Thermal-aware networkon-chips: Single- and crosslayered approaches. Future Generation Computer Systems, 91(??):61–85, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18312160.

Saberi:2019:SMB

[SSHC19] Zahra Saberi, Morteza Saberi, Omar Hussain, and Elizabeth Chang. Stackelberg model based game theory approach for assortment and selling price planning for small scale online retailers. Future Generation Computer Systems, 100(??):1088–1102, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1930336X.

Simic:2019:OPO

[SSI19] Visnja Simic, Boban Stojanovic, and Milos Ivanovic. Optimizing the performance of optimization in the cloud environment — an intelligent auto-scaling approach. Future Generation Computer Systems, 101(??):909–920, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18321496.

Shen:2019:POS

[SSJ19] Bo Shen, Yulong Shen, and Wen Ji. Profit optimization in service-oriented data market: a Stackelberg game approach. Future Generation Computer Systems, 95(??):17– 25, June 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18318193.

Siddiqua:2019:IIC

[SSK^{+}19] Ayesha Siddiqua, Munam Ali Shah, Hasan Ali Khattak, Ikram Ud Din, and Mohsen Guizani. iCAFE: Intelligent Congestion Avoidance and Fast Emergency services. Future Generation Computer Systems, 99(??):365–375, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302377.

Skalkowski:2013:QBS

[SSKK13] Kornel Skalkowski, Renata Slota, Dariusz Król, and Jacek Kitowski. QoS-based storage resources provisioning for grid applications. Future Generation Computer Systems, 29(3): 713–727, March 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001677.

Silaghi:2012:TCS

[SSL12] Gheorghe Cosmin Silaghi, Liviu Dan Serban, and Cristian Marius Litan. A time-constrained SLA negotiation strategy in competitive computational grids. Future Generation Computer Systems, 28(8):1303–1315, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002251.

Smit:2013:DAL

[SSL13] Michael Smit, Bradley Simmons, and Marin Litoiu. Distributed, applicationlevel monitoring for heterogeneous clouds using stream processing. Future Generation Computer Systems, 29(8):2103–2114, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1300023X.

Si:2019:IIS

[SSL^{+}19] Haiping Si, Changxia Sun, Yanling Li, Hongbo Qiao, and Lei Shi. IoT information sharing security mechanism based on blockchain technology. Future Generation Computer Systems, 101(??):1028–1040, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19312725.

Swain:2010:PFG

[SSLF^{+}10] Martin Swain, Cândida G. Silva, Nuno Loureiro-Ferreira, Vitaliy Ostropytskyy, João Brito, Olivier Riche, Frederick Stahl, Werner Dubitzky, and Rui M. M. Brito. P-found: Grid-enabling distributed repositories of protein folding and unfolding simulations for data mining. Future Generation Computer Systems, 26(3):424– 433, March 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sun:2017:STT

[SSP17] Hongyang Sun, Patricia Stolf, and Jean-Marc Pierson. Spatio-temporal thermal-aware scheduling for homogeneous highperformance computing datacenters. Future Generation Computer Systems, 71(??):157–170, June 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301966.

Silva:2019:CTP

[SSRQ19] Ademir Silva, Kátia Silva, Antonio Rocha, and Flavio Queiroz. Calculating the trust of providers through the construction weighted Sec-SLA. Future Generation Computer Systems, 97(??):873–886, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831570X.

Sharma:2019:FAE

[SSSJ19a] Yogesh Sharma, Weisheng Si, Daniel Sun, and Bahman Javadi. Failure-aware energy-efficient VM consolidation in cloud computing systems. Future Generation Computer Systems, 94(??):620–633, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831700X.

Sturm:2019:BBR

[SSSJ19b] Christian Sturm, Jonas Scalanczi, Stefan Schönig, and Stefan Jablonski. A blockchain-based and resource-aware process execution engine. Future Generation Computer Systems, 100(??):19–34, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18327158.

Shafiq:2017:TEB

[SSST17] Syed Imran Shafiq, Cesar Sanin, Edward Szczerbicki, and Carlos Toro. Towards an experience based collective computational intelligence for manufacturing. Future Generation Computer Systems, 66(??):89– 99, January 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301091.

Satpathy:2018:SAS

[SST18] Suchismita Satpathy, Bibhudatta Sahoo, and Ashok Kumar Turuk. Sensing and actuation as a service delivery model in cloud edge centric Internet of Things. Future Generation Computer Systems, 86(??):281–296, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17320642.

Song:2019:ERM

[SSW^{+}19] Zhiting Song, Yanming Sun, Jiafu Wan, Lingli Huang, Yan Xu, and Ching-Hsien Hsu. Exploring robustness management of Social Internet of Things for customization manufacturing. Future Generation Computer Systems, 92(??):846–856, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1731018X.

Smari:2013:RDH

[SSZ13] Waleed W. Smari, Luca Spalazzi, and Yacine Zemali. Recent developments in high performance computing and security: an editorial. Future Generation Computer Systems, 29(3): 782–787, March 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001628.

Sun:2017:ARS

[SSZ^{+}17] Shengtao Sun, Weijing Song, Albert Y. Zomaya, Yang Xiang, Kim-Kwang Raymond Choo, Tejal Shah, and Lizhe Wang. Associative retrieval in spatial big data based on spreading activation with semantic ontology. Future Generation Computer Systems, 76(??):499–509, November 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16304137.

Sashi:2011:DRD

[ST11] K. Sashi and Antony Selvadoss Thanamani. Dynamic replication in a data grid using a Modified BHR Region Based Algorithm. Future Generation Computer Systems, 27(2):202– 210, February 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shao:2017:VBR

[STA17a] Zhou Shao, David Taniar, and Kiki Maulana Adhinugraha. Voronoi-based range-kNN search with Map Grid in a mobile environment. Future Generation Computer Systems, 67(??):305–314, February 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300498.

Sta:2017:QED

[Sta17b] Hatem Ben Sta. Quality and the efficiency of data in “Smart-Cities”. Future Generation Computer Systems, 74(??):409–416, September 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308081.

Simpkin:2019:CDT

[STB^{+}19] Chris Simpkin, Ian Taylor, Graham A. Bent, Geeth de Mel, Swati Rallapalli, Liang Ma, and Mudhakar Srivatsa. Constructing distributed timecritical applications using cognitive enabled services. Future Generation Computer Systems, 100(??):70– 85, November 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18311762.

Singaraju:2015:ASN

[STC15] Janardhan Singaraju, Ajithkumar Thamarakuzhi, and John A. Chandy. Active storage networks: Using embedded computation in the network switch for cluster data processing. Future Generation Computer Systems, 45(??):149–160, April 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002143.

Saber:2018:VRH

[STMV18] Takfarinas Saber, James Thorburn, Liam Murphy, and Anthony Ventresque. VM reassignment in hybrid clouds for large decentralised companies: A multi-objective challenge. Future Generation Computer Systems, 79 (part 2)(??):751– 764, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301164.

Sun:2010:OOO

[Sun10]  Xiaoping Sun.  OSLN:

An Object-Oriented Semantic Link Network language for complex object description and operation. Future Generation Computer Systems, 26(3):389– 399, March 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Smanchat:2015:TWS

[SV15] Sucha Smanchat and Kanchana Viriyapant. Taxonomies of workflow scheduling problem and techniques in the cloud. Future Generation Computer Systems, 52(??):1–12, November 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001776.

Sahni:2016:WPA

[SV16] Jyoti Sahni and Deo Prakash Vidyarthi. Workflow-and-platform aware task clustering for scientific workflow execution in cloud environment. Future Generation Computer Systems, 64(??):61–74, November 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301182.

Sciacca:2019:VSG

[SVB^{+}19] Eva Sciacca, Fabio Vitello, Ugo Becciani, Alessandro Costa, Akos Hajnal, Peter Kacsuk, Zoltan Farkas, Istvan Marton, Sergio Molinari, Anna Maria Di Giorgio, Eugenio Schisano, Scige John Liu, Davide Elia, Stefano Cavuoti, Giuseppe Riccio, and Massimo Brescia. VIALACTEA science gateway for Milky Way analysis. Future Generation Computer Systems, 94(??):947–956, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17309561.

Sharma:2019:CAC

[SVK19] Varun Kumar Sharma, Lal Pratap Verma, and Mahesh Kumar. CL-ADSP: Cross-layer adaptive data scheduling policy in mobile ad-hoc networks. Future Generation Computer Systems, 97(??): 530–563, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17324044.

Scott:2010:SLV

[SVN^{+}10a] Stephen L. Scott, Geoffroy Vallée, Thomas Naughton, Anand Tikotekar, Christian Engelmann, and Hong Ong. System-level virtualization research at Oak Ridge National Laboratory. Future Generation Computer Systems, 26(3): 304–307, March 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sundari:2010:GMB

[SVN10b] M. Sivagama Sundari, Sathish S. Vadhiyar, and Ravi S. Nanjundiah. Grids with multiple batch systems for performance enhancement of multicomponent and parameter sweep parallel applications. Future Generation Computer Systems, 26(2): 217–227, February 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schweinsberg:2017:ACS

[SW17] Kai Schweinsberg and Lutz Wegner. Advantages of complex SQL types in storing XML documents. Future Generation Computer Systems, 68(??):500–507, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300267.

Sheikhalishahi:2016:MDJ

[SWG^{+}16] Mehdi Sheikhalishahi, Richard M.

Wallace, Lucio Grandinetti, José Luis Vazquez-Poletti, and Francesca Guerriero.

A multi-dimensional job scheduling. Future Generation Computer Systems, 54(??):123–131, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500076X.

Sher:2013:LSA

[SWW^{+}13] Anna A. Sher, Ken Wang, Andrew Wathen, Philip John Maybank, Gary R. Mirams, David Abramson, Denis Noble, and David J. Gavaghan. A local sensitivity analysis method for developing biological models with identifiable parameters: Application to cardiac ionic channel modelling. Future Generation Computer Systems, 29(2):591–598, February 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001725.

Song:2018:FRD

[SWW^{+}18] Tao Song, Jiajun Wang, Jiewei Wu, Ruhui Ma, Alei Liang, Tao Gu, and Zhengwei Qi. FastDesk: a remote desktop virtualization system for multi-tenant. Future Generation Computer Systems, 81(??):478–491, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17304776.

Sha:2018:SCO

[SWY^{+}18] Kewei Sha, Wei Wei, T. Andrew Yang, Zhiwei Wang, and Weisong Shi. On security challenges and open issues in Internet of Things. Future Generation Computer Systems, 83(??):326–337, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324883.

Sun:2013:CEF

[SYAL13] Gang Sun, Hongfang Yu, Vishal Anand, and Lemin Li. A cost efficient framework and algorithm for embedding dynamic virtual network requests. Future Generation Computer Systems, 29(5):1265–1277, July 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001586.

Sun:2018:DFB

[SYCH18] Bangyong Sun, Nianzeng Yuan, Congjun Cao, and Jon Y. Hardeberg. Design of four-band multispectral imaging system with one single-sensor. Future Generation Computer Systems, 86(??):670–679, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18300840.

Sani:2019:CSF

[SYJ^{+}19a] Abubakar Sadiq Sani, Dong Yuan, Jiong Jin, Longxiang Gao, Shui Yu, and Zhao Yang Dong. Cyber security framework for Internet of Thingsbased energy Internet. Future Generation Computer Systems, 93(??):849–859, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1730660X.

Sani:2019:NFL

[SYJ^{+}19b] Abubakar Sadiq Sani, Dong Yuan, Jiong Jin, Longxiang Gao, Shui Yu, and Zhao Yang Dong. A new fuzzy logic based node localization mechanism for wireless sensor networks. Future Generation Computer Systems, 93(??):799– 813, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17303886.

Sharma:2019:CTR

[SYJA19] Vishal Sharma, Ilsun You, Dushantha Nalin K. Jayakody, and Mohammed Atiquzzaman. Cooperative trust relaying and privacy preservation via edgecrowdsourcing in Social Internet of Things. Future Generation Computer Systems, 92(??):758–776, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312748.

Sookhak:2017:ABD

[SYK^{+}17] Mehdi Sookhak, F. Richard Yu, Muhammad Khurram Khan, Yang Xiang, and Rajkumar Buyya. Attribute-based data access control in mobile cloud computing: Taxonomy and open issues. Future Generation Computer Systems, 72(??):273–287, July 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302795.

Sun:2018:DDA

[SYL18] ShengYao Sun, WenBin Yao, and XiaoYong Li. DARS: a dynamic adaptive replica strategy under high load Cloud-P2P. Future Generation Computer Systems, 78 (part 1)(??): 31–40, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17315972.

Sun:2019:RFR

[SYQ^{+}19] ShengYao Sun, WenBin Yao, BaoJun Qiao, Ming Zong, Xin He, and XiaoYong Li. RRSD: a file replication method for ensuring data reliability and reducing storage consumption in a dynamic Cloud–P2P environment. Future Generation Computer Systems, 100(??):844–858, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308568.

Shen:2019:BMD

[SYT^{+}19] Rongbo Shen, Kezhou Yan, Kuan Tian, Cheng Jiang, and Ke Zhou. Breast mass detection from the digitized X-ray mammograms based on the combination of deep active learning and self-paced learning. Future Generation Computer Systems, 101(??):668–679, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18323938.

Song:2017:SAM

[SYW17] Jun Song, Fan Yang, and Lizhe Wang. Secure authentication in motion: a novel online payment framework for drive-thru Internet. Future Generation Computer Systems, 76(??):146–158, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301960.

Shen:2017:RDP

[SYY^{+}17] Wenting Shen, Guangyang Yang, Jia Yu, Hanlin Zhang, Fanyu Kong, and Rong Hao. Remote data possession checking with privacy-preserving authenticators for cloud storage. Future Generation Computer Systems, 76(??):136– 145, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16304939.

Shahriar:2012:TTP

[SZ12] Hossain Shahriar and Mohammad Zulkernine. Trustworthiness testing of phishing websites: a behavior model-based approach. Future Generation Computer Systems, 28(8):1258–1271, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000045.

Song:2017:TTO

[SZD^{+}17] Jeungeun Song, Yin Zhang, Kui Duan, M. Shamim Hossain, and Sk Md Mizanur Rahman. TOLA: Topic-oriented learning assistance based on cyberphysical system and big data. Future Generation Computer Systems, 75(??):200–205, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301650.

Stirm:2019:RCE

[SZG^{+}19] Claire Stirm, Rich Zink, Sandra Gesing, Michael Zentner, and Damion Junk. REMEDI central

— expanding and sustaining a medical device community. Future Generation Computer Systems, 101(??):372–380, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19303127.

Shen:2016:MSH

[SZK16] Bo Shen, Xingshe Zhou, and Mucheol Kim. Mixed scheduling with heterogeneous delay constraints in cyber-physical systems. Future Generation Computer Systems, 61(??):108– 117, August 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003349.

Slawik:2018:EUC

[SZK18] Mathias Slawik, Begüm Ilke Zilci, and Axel Küpper. Establishing user-centric cloud service registries. Future Generation Computer Systems, 87(??):846–867, October 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18304813.

Sun:2018:NBD

[SZR18] Xin Sun, Xishun Zhao, and Livio Robaldo. Normbased deontic logic for access control, some computational results. Future Generation Computer Systems, 79 (part 1)(??):295– 302, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1730136X.

Sakr:2019:ESI

[SZV19] Sherif Sakr, Albert Zomaya, and Athanasios V. Vasilakos. Editorial for special issue of FGCS special issue on “Benchmarking big data systems”. Future Generation Computer Systems, 96 (??):32–34, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19303048.

Sun:2019:EEI

[SZW^{+}19] Mengyu Sun, Zhangbing Zhou, Junping Wang, Chu Du, and Walid Gaaloul. Energy-efficient IoT service composition for concurrent timed applications. Future Generation Computer Systems, 100(??):1017–1030, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19302274.

Thilak:2018:CAB

[TA18] K. Deepa Thilak and A. Amuthan. Cellular automata-based improved ant colony-based optimization algorithm for mitigating DDoS attacks in VANETs. Future Generation Computer Systems, 82(??):304–314, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1732215X.

Tapia:2019:SBE

[TA19] Juan E. Tapia and Claudia Arellano. Soft-biometrics encoding conditional GAN for synthesis of NIR periocular images. Future Generation Computer Systems, 97(??):503–511, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18330966.

Trani:2018:ECC

[TAB^{+}18] Luca Trani, Malcolm Atkinson, Daniele Bailo, Rossana Paciello, and Rosa Filgueira. Establishing core concepts for information-powered collaborations. Future Generation Computer Systems, 89(??):421–437, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17327619.

Thirunarayan:2014:CTM

[TAHS14] Krishnaprasad Thirunarayan, Pramod Anantharam, Cory Henson, and Amit Sheth. Comparative trust management with applications: Bayesian approaches emphasis. Future Generation Computer Systems, 31(??):182–199, February 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001052.

Tserpes:2012:RMS

[TAKV12] Konstantinos Tserpes, Fotis Aisopos, Dimosthenis Kyriazis, and Theodora Varvarigou. A recommender mechanism for service selection in serviceoriented environments. Future Generation Computer Systems, 28(8):1285–1294, October 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002263.

Tao:2010:CCP

[Tao10] Jie Tao. Comprehensive cache performance tuning with a toolset. Future Generation Computer Systems, 26(1):167–174, January 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tonyali:2018:PPP

[TAS^{+}18] Samet Tonyali, Kemal Akkaya, Nico Saputro, A. Selcuk Uluagac, and Mehrdad Nojoumian. Privacypreserving protocols for secure and reliable data aggregation in IoT-enabled Smart Metering systems. Future Generation Computer Systems, 78 (part 2)(??):547–557, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17306945.

Toffetti:2017:SMC

[TBB^{+}17] Giovanni Toffetti, Sandro Brunner, Martin Blöchlinger, Josef Spillner, and Thomas Michael Bohnert. Self-managing cloud-native applications: Design, implementation, and experience. Future Generation Computer Systems, 72(??):165–179, July 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302977.

Tierney:2016:SSH

[TBdL16] Brian Tierney, Mehmet Balman, and Cees de Laat. Special section on highperformance networking for distributed data-intensive science. Future Generation Computer Systems, 56(??): 262–264, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003192.

Tao:2010:SST

[TBK^{+}10] Jie Tao, Arndt Bode, Andreas Knüpfer, Dieter Kranzlmüller, Jens Volkert, and Roland Wismüller. Special section: Tools for program development and analysis in computational science. Future Generation Computer Systems, 26 (1):135–137, January 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tao:2019:EPA

[TBR^{+}19] Hai Tao, Md Zakirul Alam Bhuiyan, Md Arafatur Rahman, Guojun Wang, Tian Wang, Md. Manjur Ahmed, and Jing Li. Economic perspective analysis of protecting big data security and privacy. Future Generation Computer Systems, 98(??):660–671, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18319915.

Tan:2018:ARM

[TBS^{+}18] Jen Hong Tan, Sulatha V.

Bhandary, Sobha Sivaprasad, Yuki Hagiwara, Akanksha Bagchi, U. Raghavendra, A. Krishna Rao, Biju Raju, Nitin Shridhara Shetty, Arkadiusz Gertych, Kuang Chua Chua, and U. Rajendra Acharya. Age-related macular degeneration detection using deep convolutional neural network. Future Generation Computer Systems, 87(??):127–135, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17319167.

Trilles:2017:DOS

[TCB^{+}17] Sergio Trilles, Andrea Calia, Óscar Belmonte, Joaquín Torres-Sospedra, Raúl Montoliu, and Joaquín Huerta. Deployment of an open sensorized platform in a smart city context. Future Generation Computer Systems, 76(??):221– 233, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16305519.

Tolosana-Calasanz:2018:MDD

[TCBC18] Rafael Tolosana-Calasanz, José Ángel Bañares, and José-Manuel Colom. Modeldriven development of data intensive applications over cloud resources. Future Generation Computer Systems, 87(??):888–909, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17329473.

Tolosana-Calasanz:2016:RMB

[TCBPR16] Rafael Tolosana-Calasanz, José Ángel Bañares, Congduc Pham, and Omer F. Rana. Resource management for bursty streams on multi-tenancy cloud environments. Future Generation Computer Systems, 55(??):444–459, February 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000679.

Tai:2018:HKF

[TCC18] Chih-Hua Tai, Ching-Tang Chang, and Yue-Shan Chang. Hybrid knowledge fusion and inference on cloud environment. Future Generation Computer Systems, 87(??):568–579, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17314176.

Tomas:2011:NAM

[TCCC11] Luis Tomás, Agustín C.

Caminero, Carmen Carrión, and Blanca Caminero. Network-aware meta-scheduling in advance with autonomous self-tuning system. Future Generation Computer Systems, 27(5): 486–497, May 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Triboan:2019:SBA

[TCCW19] Darpan Triboan, Liming Chen, Feng Chen, and Zumin Wang. A semanticsbased approach to sensor data segmentation in realtime activity recognition. Future Generation Computer Systems, 93(??):224– 236, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18303947.

Tang:2014:SIR

[TCG14] Xueyan Tang, Wentong Cai, and Rick Siow Mong Goh. Special issue: Recent Advances in Parallel and Distributed Systems, ICPADS 2012 selected papers. Future Generation Computer Systems, 30(??): 169, January 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002215.

Tsai:2019:EUP

[TCH19] Ming-Fong Tsai, Ping Chen, and Yap Jia Hong. Enhancing the utilization of public bike sharing systems using return anxiety information. Future Generation Computer Systems, 92(??):961–971, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17312682.

Thilakanathan:2014:PSM

[TCN^{+}14] Danan Thilakanathan, Shiping Chen, Surya Nepal, Rafael Calvo, and Leila Alem. A platform for secure monitoring and sharing of generic health data in the Cloud. Future Generation Computer Systems, 35(??):102–113, June 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001908.

Tudoran:2016:JEH

[TCN^{+}16] Radu Tudoran, Alexandru Costan, Olivier Nano, Ivo Santos, Hakan Soncu, and Gabriel Antoniu. Jet-Stream: Enabling high throughput live event streaming on multi-site clouds. Future Generation Computer Systems, 54(??):274–291, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000333.

Tomas:2012:GBA

[TCR^{+}12] Luis Tomás, Agustín C.

Caminero, Omer Rana, Carmen Carrión, and Blanca Caminero. A GridWay-based autonomic network-aware metascheduler. Future Generation Computer Systems, 28(7): 1058–1069, July 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002196.

Truica:2018:BTK

[TDBR18] Ciprian-Octavian Truica, Jérôme Darmont, Alexandru Boicea, and Florin Radulescu. Benchmarking top-k keyword and top-k document processing with T^{2}K^{2} and T^{2}K^{2}D^{2}. Future Generation Computer Systems, 85(??):60–75, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17323580.

Tang:2014:LCM

[TDC^{+}14] C.-J. Tang, M.-R. Dai, C.- C. Chuang, Y.-S. Chiu, and W. S. Lin. A load control method for small data centers participating in demand response programs. Future Generation Computer Systems, 32(??): 232–245, March 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001659.

Tan:2018:ESA

[TDFZ18] Zhipeng Tan, Li Du, Dan Feng, and Wei Zhou. EML: an I/O scheduling algorithm in largescale-application environments. Future Generation Computer Systems, 78 (part 3)(??):1091–1100, January 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17306258.

Tang:2017:TTE

[TDLC17] Mingdong Tang, Xiaoling Dai, Jianxun Liu, and Jinjun Chen. Towards a trust evaluation middleware for cloud service selection. Future Generation Computer Systems, 74(??):302–312, September 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1600011X.

Teylo:2017:HEA

[TdPF^{+}17] Luan Teylo, Ubiratam de Paula, Yuri Frota, Daniel de Oliveira, and Lúcia M. A. Drummond. A hybrid evolutionary algorithm for task scheduling and data assignment of data-intensive scientific workflows on clouds. Future Generation Computer Systems, 76(??):1– 17, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17309883.

Tchana:2016:SCE

[TDSH16] Alain Tchana, Noel De Palma, Ibrahim Safieddine, and Daniel Hagimont. Software consolidation as an efficient energy and cost saving solution. Future Generation Computer Systems, 58 (??):1–12, May 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003891.

Teimourikia:2017:ODR

[TF17] Mahsa Teimourikia and Mariagrazia Fugini. Ontology development for runtime safety management methodology in smart work environments using ambient knowledge. Future Generation Computer Systems, 68(??):428–441, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302333.

Tang:2018:CBI

[TF18] Rui Tang and Simon Fong. Clustering big IoT data by metaheuristic optimized mini-batch and parallel partition-based DGC in Hadoop. Future Generation Computer Systems, 86(??):1395–1412, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17322677.

Tsamoura:2011:DEL

[TGM11] Efthymia Tsamoura, Anastasios Gounaris, and Yannis Manolopoulos. Decentralized execution of linear workflows over Web services. Future Generation Computer Systems, 27 (3):341–347, March 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tang:2019:SCM

[TGM^{+}19a] Yayuan Tang, Kehua Guo, Jianhua Ma, Yutong Shen, and Tao Chi. A smart caching mechanism for mobile multimedia in information centric networking with edge computing. Future Generation Computer Systems, 91(??):590–600, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18311841.

Tortonesi:2019:TID

[TGM^{+}19b] Mauro Tortonesi, Marco Govoni, Alessandro Morelli, Giulio Riberto, Cesare Stefanelli, and Niranjan Suri. Taming the IoT data deluge: An innovative information-centric service model for fog computing applications. Future Generation Computer Systems, 93(??):888–902, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17306702.

Tanbeer:2017:SRP

[THA^{+}17] Syed Khairuzzaman Tanbeer, Mohammad Mehedi Hassan, Ahmad Almogren, Mansour Zuair, and Byeong-Soo Jeong. Scalable regular pattern mining in evolving body sensor data. Future Generation Computer Systems, 75(??):172–186, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300851.

Trinh:2012:EEL

[THT12] Tuan Anh Trinh, Helmut Hlavacs, and Domenico Talia. Energy efficiency in large-scale distributed systems. Future Generation Computer Systems, 28(5): 743–744, May 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002342.

Tsumoto:2014:SBB

[TIHT14] Shusaku Tsumoto, Haruko Iwata, Shoji Hirano, and Yuko Tsumoto. Similaritybased behavior and process mining of medical practices. Future Generation Computer Systems, 33(??): 21–31, April 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002318.

Tung:2018:SOF

[TJ18] Cheng-Huang Tung and En-Yih Jean. Stroke-order-free on-line Chinese character recognition by stroke adjustment of two-layer bipartite weighted matching. Future Generation Computer Systems, 81(??):219– 234, April 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16307075.

Tao:2010:SDO

[TJWS10] Yongcai Tao, Hai Jin, Song Wu, and Xuanhua Shi. Scalable DHT- and ontology-based information service for large-scale grids. Future Generation Computer Systems, 26(5): 729–739, May 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tang:2015:SAS

[TJZ^{+}15] Zhuo Tang, Lingang Jiang, Junqing Zhou, Kenli Li, and Keqin Li. A selfadaptive scheduling algorithm for reduce start time. Future Generation Computer Systems, 43–44(??): 51–60, February 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001599.

Taylor:2018:CSP

[TKA^{+}18a] Simon J. E. Taylor, Tamas Kiss, Anastasia Anagnostou, Gabor Terstyanszky, Peter Kacsuk, Joris Costes, and Nicola Fantini. The CloudSME simulation platform and its applications: a generic multi-cloud platform for developing and executing commercial cloud-based simulations. Future Generation Computer Systems, 88(??):524– 539, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17329102.

Thirukrishna:2018:REE

[TKA18b] J. T. Thirukrishna, S. Karthik, and V. P. Arunachalam. Revamp energy efficiency in homogeneous wireless sensor networks using optimized radio energy algorithm (OREA) and power-aware distance source routing protocol. Future Generation Computer Systems, 81(??):331– 339, April 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17322082.

Terstyanszky:2014:ESW

[TKK^{+}14] Gabor Terstyanszky, Tamas Kukla, Tamas Kiss, Peter Kacsuk, Akos Balasko, and Zoltan Farkas. Enabling scientific workflow sharing through coarsegrained interoperability. Future Generation Computer Systems, 37(??):46– 59, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000417.

Tao:2015:NNT

[TKR^{+}15] Jie Tao, Joanna Kolodziej, Rajiv Ranjan, Prem Prakash Jayaraman, and Rajkumar Buyya. A note on new trends in dataaware scheduling and resource provisioning in modern HPC systems. Future Generation Computer Systems, 51(??):45–46, October 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001089.

Theodoropoulos:2014:ESI

[TKRA14] Georgios Theodoropoulos, Kostas Katrinis, Rolf Riesen, and Shoukat Ali. Editorial: Special issue on Extreme Scale Parallel Architectures and Systems. Future Generation Computer Systems, 30(??):44– 45, January 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002203.

Temelkovski:2019:EMD

[TKTG19] Damjan Temelkovski, Tamas Kiss, Gabor Terstyanszky, and Pamela Greenwell. Extending molecular docking desktop applications with cloud computing support and analysis of results. Future Generation Computer Systems, 97(??):814–824, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18314092.

Tsai:2019:SSE

[TL19] Chun-Wei Tsai and Shih-Jui Liu. SEIM: Search economics for influence maximization in online social networks. Future Generation Computer Systems, 93(??):1055–1064, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1730849X.

Tang:2015:EHI

[TLC^{+}15] Shaohua Tang, Bo Lv, Guomin Chen, Zhiniang Peng, Adama Diene, and Xiaofeng Chen. Efficient hardware implementation of PMI+ for low-resource devices in mobile cloud computing. Future Generation Computer Systems, 52(??):116–124, November 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002489.

Tang:2011:SSA

[TLL^{+}11] Xiaoyong Tang, Kenli Li, Guiping Liao, Kui Fang, and Fan Wu. A stochastic scheduling algorithm for precedence constrained tasks on Grid. Future Generation Computer Systems, 27(8):1083–1091, October 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Teng:2019:NCD

[TLL^{+}19] Haojun Teng, Yuxin Liu, Anfeng Liu, Neal N. Xiong, Zhiping Cai, Tian Wang, and Xuxun Liu. A novel code data dissemination scheme for Internet of Things through mobile vehicle of smart cities. Future Generation Computer Systems, 94(??):351–367, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18319289.

Tao:2017:GTM

[TLSC17] Xiuting Tao, Guoqiang Li, Daniel Sun, and Hongming Cai. A game-theoretic model and analysis of data exchange protocols for Internet of Things in clouds. Future Generation Computer Systems, 76(??):582– 589, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16308378.

Tarkhaneh:2019:IDE

[TM19] Omid Tarkhaneh and Irene Moser. An improved differential evolution algorithm using Archimedean spiral and neighborhood search based mutation approach for cluster analysis. Future Generation Computer Systems, 101(??):921–939, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19310337.

Tariq:2019:ADS

[TMB^{+}19] Muhammad Tariq, Hammad Majeed, Mirza Omer Beg, Farrukh Aslam Khan, and Abdelouahid Derhab. Accurate detection of sitting posture activities in a secure IoT based assisted living environment. Future Generation Computer Systems, 92(??):745–757, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1731333X.

Tian:2015:PBA

[TMDZ15] Feng Tian, Tian Ma, Bo Dong, and Qinghua Zheng. PWLM_{3}-based automatic performance model estimation method for HDFS write and read operations. Future Generation Computer Systems, 50(??):127–139, September 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500028X.

Tian:2019:KNC

[TMJH19] Donghai Tian, Rui Ma, Xiaoqi Jia, and Changzhen Hu. KEcruiser: a novel control flow protection for kernel extensions. Future Generation Computer Systems, 100(??):1–9, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331273.

Taura:2013:DIG

[TMM^{+}13] Kenjiro Taura, Takuya Matsuzaki, Makoto Miwa, Yoshikazu Kamoshida, Daisaku Yokoyama, Nan Dun, Takeshi Shibata, Choi Sung Jun, and Jun’ichi Tsujii. Design and implementation of GXP make — a workflow system based on make. Future Generation Computer Systems, 29(2):662–672, February 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001087.

Tordsson:2012:CBM

[TMMVL12] Johan Tordsson, Rubén S. Montero, Rafael Moreno-Vozmediano, and Ignacio M. Llorente. Cloud brokering mechanisms for optimized placement of virtual machines across multiple providers. Future Generation Computer Systems, 28(2):358–367, February 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001373.

Tormo:2015:DFS

[TMP15] Ginés Dólera Tormo, Félix Gómez Mármol, and Gregorio Martínez Pérez. Dynamic and flexible selection of a reputation mechanism for heterogeneous environments. Future Generation Computer Systems, 49(??):113– 124, August 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001289.

Trapero:2017:NAM

[TMS^{+}17] Ruben Trapero, Jolanda Modic, Miha Stopar, Ahmed Taha, and Neeraj Suri. A novel approach to manage cloud security SLA incidents. Future Generation Computer Systems, 72(??): 193–205, July 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301844.

Tarneberg:2017:DAP

[TMW^{+}17] William Tärneberg, Amardeep Mehta, Eddie Wadbro, Johan Tordsson, Johan Eker, Maria Kihl, and Erik Elmroth. Dynamic application placement in the mobile cloud network. Future Generation Computer Systems, 70(??):163–177, May 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302060.

Tarus:2017:HKB

[TNY17] John K. Tarus, Zhendong Niu, and Abdallah Yousif.

A hybrid knowledge-based recommender system for elearning based on ontology and sequential pattern mining. Future Generation Computer Systems, 72 (??):37–48, July 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17303254.

Tao:2017:OBD

[TOD17] Ming Tao, Kaoru Ota, and Mianxiong Dong. Ontology-based data semantic management and application in IoT- and cloud-enabled smart homes. Future Generation Computer Systems, 76(??):528– 539, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630615X.

Torkestani:2013:HRP

[Tor13] Javad Akbari Torkestani. A highly reliable and parallelizable data distribution scheme for data grids. Future Generation Computer Systems, 29(2):509– 519, February 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001549.

Torroglosa:2018:MFI

[TOS18] Elena Torroglosa, Jordi Ortiz, and Antonio Skarmeta. Matching federation identities, the eduGAIN and STORK approach. Future Generation Computer Systems, 80(??):126–138, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300523.

Tchernykh:2014:AEE

[TPBS14] Andrei Tchernykh, Johnatan E.

Pecero, Aritz Barrondo, and Elisa Schaeffer. Adaptive energy efficient scheduling in Peer-to-Peer desktop grids. Future Generation Computer Systems, 36(??): 209–220, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001568.

Tong:2019:TCF

[TQL^{+}19] Chao Tong, Ji Qi, Yu Lian, Jianwei Niu, and Joel J. P. C. Rodrigues. TimeTrustSVD: A collaborative filtering model integrating time, trust and rating information. Future Generation Computer Systems, 93(??):933–941, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17307732.

Turabieh:2018:DRR

[TSAER18] Hamza Turabieh, Amer Abu Salem, and Noor Abu-El-Rub. Dynamic L-RNN recovery of missing data in IoMT applications. Future Generation Computer Systems, 89(??):575–583, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18307490.

Toosi:2018:RPD

[TSB18] Adel Nadjaran Toosi, Richard O. Sinnott, and Rajkumar Buyya. Resource provisioning for data-intensive applications with deadline constraints on hybrid clouds using Aneka. Future Generation Computer Systems, 79 (part 2)(??):765–775, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301863.

Torres:2011:SMD

[TSBH11] E. Torres, D. Segrelles, I. Blanquer, and V. Hernández. Service monitoring and differentiation techniques for resource allocation in the grid, on the basis of the level of service. Future Generation Computer Systems, 27(8):1142–1152, October 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tang:2018:TAC

[TSD18] Kun Tang, Ronghua Shi, and Jian Dong. Throughput analysis of cognitive wireless acoustic sensor networks with energy harvesting. Future Generation Computer Systems, 86(??):1218–1227, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17315327.

Terroso-Saenz:2019:OIP

[TSGVRGS19] Fernando Terroso-Saenz, Aurora González-Vidal, Alfonso P. Ramallo-González, and Antonio F. Skarmeta. An open IoT platform for the management and analysis of energy data. Future Generation Computer Systems, 92(??):1066–1079, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17304181.

Tso:2019:TOO

[Tso19] Raylin Tso. Two-in-one oblivious signatures. Future Generation Computer Systems, 101(??):467–475, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1930202X.

That:2015:PAP

[TSOB15] M. T. T. That, S. Sadou, F. Oquendo, and I. Borne. Preserving architectural pattern composition information through explicit merging operators. Future Generation Computer Systems, 47(??):97–112, June 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001654.

Tennant:2017:SRT

[TSRG17] Mark Tennant, Frederic Stahl, Omer Rana, and João Bártolo Gomes. Scalable real-time classification of data streams with concept drift. Future Generation Computer Systems, 75(??):187–199, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17304685.

Tian:2019:DDM

[TSS^{+}19] Zhihong Tian, Shen Su, Wei Shi, Xiaojiang Du, Mohsen Guizani, and Xiang Yu. A data-driven method for future Internet route decision modeling. Future Generation Computer Systems, 95(??): 212–220, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18311518.

Tibermacine:2016:SAC

[TSTD16] Chouki Tibermacine, Salah Sadou, Minh Tu Ton That, and Christophe Dony. Software architecture constraint reuseby-composition. Future Generation Computer Systems, 61(??):37–53, August 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630019X.

Tokmakoff:2016:ARF

[TSTL16] Andrew Tokmakoff, Ben Sparrow, David Turner, and Andrew Lowe. Aus-Plots Rangelands field data collection and publication: Infrastructure for ecological monitoring. Future Generation Computer Systems, 56(??):537–549, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002782.

Tang:2017:ACC

[TSWL17] Yu Tang, Hailong Sun, Xu Wang, and Xudong Liu. Achieving convergent causal consistency and high availability for cloud storage. Future Generation Computer Systems, 74(??):20–31, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17305940.

Tchana:2013:TLA

[TTB^{+}13] Alain Tchana, Giang Son Tran, Laurent Broto, Noel DePalma, and Daniel Hagimont. Two levels autonomic resource management in virtualized IaaS. Future Generation Computer Systems, 29(6):1319– 1332, August 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000277.

Tang:2014:SLB

[TTC^{+}14] Cheng-Hsien Tang, Meng-Feng Tsai, Shan-Hao Chuang, Jen-Jung Cheng, and Wei-Jen Wang. Shortestlinkage-based parallel hierarchical clustering on main-belt moving objects of the solar system. Future Generation Computer Systems, 34(??):26–46, May 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002938.

Teabe:2015:ECA

[TTH15] Boris Teabe, Alain Tchana, and Daniel Hagimont. Enforcing CPU allocation in a heterogeneous IaaS. Future Generation Computer Systems, 53(??):1–12, December 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001880.

Toor:2019:EPA

[TuIS^{+}19] Asfa Toor, Saif ul Islam, Nimra Sohail, Adnan Akhunzada, Jalil Boudjadar, Hasan Ali Khattak, Ikram Ud Din, and Joel J. P. C. Rodrigues. Energy and performance aware fog computing: a case of DVFS and green renewable energy. Future Generation Computer Systems, 101(??):1112–1121, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19310234.

Turek:2018:EBD

[Tur18] Wojciech Turek. Erlangbased desynchronized urban traffic simulation for high-performance computing systems. Future Generation Computer Systems, 79 (part 2)(??):645– 652, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17311810.

Tiwari:2016:IAC

[TV16] Pawan Kumar Tiwari and Deo Prakash Vidyarthi. Improved auto control ant colony optimization using lazy ant approach for grid scheduling problem. Future Generation Computer Systems, 60(??):78– 89, July 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300012.

Thai:2018:STR

[TVB18] Long Thai, Blesson Varghese, and Adam Barker. A survey and taxonomy of resource optimisation for executing bag-of-task applications on public clouds.

Future Generation Computer Systems, 82(??):1– 11, May 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17305071.

Tiplea:2013:RAM

[TVV13] Ferucio Laurentiu Tiplea, Loredana Vamanu, and Cosmin Vârlan. Reasoning about minimal anonymity in security protocols. Future Generation Computer Systems, 29(3):828–842, March 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000349.

Taal:2019:PSD

[TWdLZ19] Arie Taal, Junchao Wang, Cees de Laat, and Zhiming Zhao. Profiling the scheduling decisions for handling critical paths in deadline-constrained cloud workflows. Future Generation Computer Systems, 100(??):237–249, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18311634.

Tan:2019:AAI

[TWG^{+}19] Zhenhua Tan, DanKe Wu, Tianhan Gao, Ilsun You, and Vishal Sharma. AIM: Activation increment minimization strategy for preventing bad information diffusion in OSNs. Future Generation Computer Systems, 94(??):293–301, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18319058.

Tay:2018:TSA

[TWW^{+}18] Jia Jun Tay, M. L. Dennis Wong, Ming Ming Wong, Cishen Zhang, and Ismat Hijazin. A tree search algorithm for low multiplicative complexity logic design. Future Generation Computer Systems, 83(??):132– 143, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17320010.

Tian:2018:VBD

[TWZP18] Ling Tian, Hongyu Wang, Yimin Zhou, and Chengzong Peng. Video big data in smart city: Background construction and optimization for surveillance video processing. Future Generation Computer Systems, 86(??):1371–1382, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17321805.

Tang:2014:TPS

[TX14] Shaohua Tang and Lingling Xu. Towards provably secure proxy signature scheme based on Isomorphisms of Polynomials. Future Generation Computer Systems, 30(??):91– 97, January 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001179.

Tian:2011:TRR

[TY11] Chunqi Tian and Baijian Yang. R^{2} Trust, a reputation and risk based trust management framework for large-scale, fully decentralized overlay networks. Future Generation Computer Systems, 27(8):1135–1141, October 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tong:2018:NDL

[TYWZ18] Chao Tong, Xiang Yin, Shili Wang, and Zhigao Zheng. A novel deep learning method for aircraft landing speed prediction based on cloudbased sensor data. Future Generation Computer Systems, 88(??):552–558, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18310641.

Taheri:2013:HNN

[TZBK13] Javid Taheri, Albert Y. Zomaya, Pascal Bouvry, and Samee U. Khan. Hopfield neural network for simultaneous job scheduling and data replication in grids. Future Generation Computer Systems, 29(8):1885–1900, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000800.

Tian:2019:MHR

[TZD^{+}19] Daxin Tian, Chuang Zhang, Xuting Duan, Yunpeng Wang, Jianshan Zhou, and Zhengguo Sheng. A multi-hop routing protocol for video transmission in IoVs based on cellular attractor selection. Future Generation Computer Systems, 95(??):713–726, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18318077.

Tao:2018:MLC

[TZL^{+}18] Ming Tao, Jinglong Zuo, Zhusong Liu, Aniello Castiglione, and Francesco Palmieri. Multi-layer cloud architectural model and ontology-based security service framework for IoTbased smart homes. Future Generation Computer Systems, 78 (part 3)(??):1040– 1051, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16305775.

Tang:2018:IDP

[TZLL18] Zhuo Tang, Xiangshen Zhang, Kenli Li, and Keqin Li. An intermediate data placement algorithm for load balancing in Spark computing environment. Future Generation Computer Systems, 78 (part 1)(??):287–301, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302126.

Taneja:2018:TBM

[TZQ18] Shubbhi Taneja, Yi Zhou, and Xiao Qin. Thermal benchmarking and modeling for HPC using big data applications. Future Generation Computer Systems, 87(??):372–381, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17323622.

Taheri:2014:PFJ

[TZST14] Javid Taheri, Albert Y. Zomaya, Howard Jay Siegel, and Zahir Tari. Pareto frontier for job execution and data transfer time in hybrid clouds. Future Generation Computer Systems, 37(??):321–334, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002847.

Uriarte:2019:DGD

[UDST19] Rafael Brundo Uriarte, Rocco De Nicola, Vincenzo Scoca, and Francesco Tiezzi. Defining and guaranteeing dynamic service levels in clouds. Future Generation Computer Systems, 99(??):27–40, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18320144.

Uta:2018:MNA

[UDvdW^{+}18] Alexandru Uta, Ove Danner, Cas van der Weegen, Ana-Maria Oprescu, Andreea Sandu, Stefania Costache, and Thilo Kielmann. MemEFS: A network-aware elastic in-memory runtime distributed file system. Future Generation Computer Systems, 82(??):631–646, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17303783.

Urbieta:2017:ACA

[UGBM^{+}17] A. Urbieta, A. González-Beltrán, S. Ben Mokhtar, M. Anwar Hossain, and L. Capra. Adaptive and context-aware service composition for IoT-based smart cities. Future Generation Computer Systems, 76(??):262–274, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308688.

Ullah:2019:UEH

[UKK^{+}19] Sana Ullah, Ki-Il Kim, Kyong Hoon Kim, Muhammad Imran, Pervez Khan, Eduardo Tovar, and Farman Ali. UAV-enabled healthcare architecture: Issues and challenges. Future Generation Computer Systems, 97(??):425–432, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18318247.

Ullah:2019:ARU

[UMUB19] Amin Ullah, Khan Muhammad, Ijaz Ul Haq, and Sung Wook Baik. Action recognition using optimized deep autoencoder and CNN for surveillance data streams of nonstationary environments. Future Generation Computer Systems, 96(??):386– 397, July 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18318533.

Ubik:2016:CPT

[UNM^{+}16] Sven Ubik, Jiri Navratil, Jiri Melnikov, Boncheol Goo, Faridah Noor Mohd Noor, Alain Baumann, Jaroslav Hrb, Claudio Allocchio, and Gerard Castillo. Cyber performances, technical and artistic collaboration across continents. Future Generation Computer Systems, 54(??):306–312, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002198.

Uribe-Perez:2017:NCS

[UPP17] N. Uribe-Pérez and C. Pous.

A novel communication system approach for a Smart City based on the human nervous system. Future Generation Computer Systems, 76(??):314–328, November 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308640.

UlHassan:2019:PPB

[URC19] Muneeb Ul Hassan, Mubashir Husain Rehmani, and Jinjun Chen. Privacy preservation in blockchain based IoT systems: Integration issues, prospects, challenges, and future research directions. Future Generation Computer Systems, 97(??): 512–529, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326542.

Umer:2019:IRM

[URKM19] Tariq Umer, Mubashir Husain Rehmani, Ahmed E. Kamal, and Lyudmila Mihaylova. Information and resource management systems for Internet of Things: Energy management, communication protocols and future applications. Future Generation Computer Systems, 92(??):1021–1027, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18329194.

urRehman:2019:RBD

[uRYS^{+}19] Muhammad Habib ur Rehman, Ibrar Yaqoob, Khaled Salah, Muhammad Imran, Prem Prakash Jayaraman, and Charith Perera. The role of big data analytics in industrial Internet of Things. Future Generation Computer Systems, 99(??):247–259, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18313645.

Uta:2016:ODL

[USK16] Alexandru Uta, Andreea Sandu, and Thilo Kielmann. Overcoming data locality: an in-memory runtime file system with symmetrical data distribution. Future Generation Computer Systems, 54(??):144– 158, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000308.

Uflacker:2011:SNA

[UZ11] Matthias Uflacker and Alexander Zeier. A semantic network approach to analyzing virtual team interactions in the early stages of conceptual design. Future Generation Computer Systems, 27(1):88–99, January 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Valls:2012:DBP

[VAdlP12] Marisol García Valls, Alejandro Alonso, and Juan Antonio de la Puente. A dualband priority assignment algorithm for dynamic QoS resource management. Future Generation Computer Systems, 28(6):902–912, June 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001907.

Vanmechelen:2014:ECS

[VAR14] Kurt Vanmechelen, Jörn Altmann, and Omer Rana. Economics of computing services. Future Generation Computer Systems, 41(??):17–18, December 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001691.

Varghese:2018:NGC

[VB18] Blesson Varghese and Rajkumar Buyya. Next generation cloud computing: New trends and research directions. Future Generation Computer Systems, 79 (part 3)(??):849–861, February 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302224.

Vijayakumar:2018:CEP

[VCD^{+}18] P. Vijayakumar, Victor Chang, L. Jegatha Deborah, Balamurugan Balusamy, and P. G. Shynu. Computationally efficient privacy preserving anonymous mutual and batch authentication schemes for vehicular ad hoc networks. Future Generation Computer Systems, 78 (part 3)(??):943–955, January 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630663X.

Vijayakumar:2018:KMK

[VCDK18] P. Vijayakumar, Victor Chang, L. Jegatha Deborah, and Bharat S. Rawal Kshatriya. Key management and key distribution for secure group communication in mobile and cloud network. Future Generation Computer Systems, 84(??):123–125, July 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18305363.

Vaquero:2019:RCN

[VCE^{+}19] Luis M. Vaquero, Felix Cuadrado, Yehia Elkhatib, Jorge Bernal-Bernabe, Satish N. Srirama, and Mohamed Faten Zhani. Research challenges in nextgen service orchestration. Future Generation Computer Systems, 90(??): 20–38, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18303157.

Vecchiola:2012:DDP

[VCKB12] Christian Vecchiola, Rodrigo N. Calheiros, Dileban Karunamoorthy, and Rajkumar Buyya. Deadlinedriven provisioning of resources for scientific applications in hybrid clouds with Aneka. Future Generation Computer Systems, 28(1):58–65, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000896.

Verba:2019:MIB

[VCL^{+}19] Nandor Verba, Kuo-Ming Chao, Jacek Lewandowski, Nazaraf Shah, Anne James, and Feng Tian. Modeling industry 4.0 based fog computing environments for application analysis and deployment. Future Generation Computer Systems, 91(??):48–60, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18303297.

Vu:2016:PBB

[VD16] Trong-Tuan Vu and Bilel Derbel. Parallel Branchand-Bound in multi-core multi-CPU multi-GPU heterogeneous environments. Future Generation Computer Systems, 56(??): 95–109, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003222.

vanderHam:2015:NIM

[vdHSL^{+}15] Jeroen van der Ham, József Stéger, Sándor Laki, Yiannos Kryftis, Vasilis Maglaris, and Cees de Laat. The NOVI information models. Future Generation Computer Systems, 42(??):64–73, January 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002811.

Vescoukis:2012:SOA

[VDK12] Vassilios Vescoukis, Nikolaos Doulamis, and Sofia Karagiorgou. A service oriented architecture for decision support systems in environmental crisis management. Future Generation Computer Systems, 28(3): 593–604, March 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000380.

vanderLee:2019:IGM

[vdLLE19] Tim van der Lee, Antonio Liotta, and Georgios Exarchakos. Interference graphs to monitor and control schedules in low-power WPAN. Future Generation Computer Systems, 93(??):111–120, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17329862.

vanderPol:2016:AST

[vdPGZ^{+}16] Ronald van der Pol, Bart Gijsen, Piotr Zuraniewski, Daniel Filipe Cabaça Romão, and Marijke Kaat. Assessment of SDN technology for an easy-to-use VPN service. Future Generation Computer Systems, 56(??):295– 302, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002903.

Varvarigou:2012:INA

[VDTK12] Theodora Varvarigou, Anastasios Doulamis, Konstantinos Tserpes, and Dimosthenis Kyriazis. Infrastructure and networkaware Grids and service oriented architectures. Future Generation Computer Systems, 28(3):525–526, March 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001877.

Veiga:2018:BEE

[VEET18] Jorge Veiga, Jonatan Enes, Roberto R. Expósito, and Juan Touriño. BDEv 3.0: Energy efficiency and microarchitectural characterization of big data processing frameworks. Future Generation Computer Systems, 86(??):565–581, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17325360.

Veiga:2016:FME

[VETT16] Jorge Veiga, Roberto R.

Expósito, Guillermo L. Taboada, and Juan Touriño. Flame-MR: an eventdriven architecture for MapReduce applications. Future Generation Computer Systems, 65(??):46– 56, December 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301868.

Vaisband:2018:HDI

[VF18] Boris Vaisband and Eby G.

Friedman. Heterogeneous 3-D ICs as a platform for hybrid energy harvesting in IoT systems. Future Generation Computer Systems, 87(??):152–158, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17321635.

Villarreal:2014:MUA

[VFHB14] Vladimir Villarreal, Jesus Fontecha, Ramon Hervas, and Jose Bravo. Mobile and ubiquitous architecture for the medical control of chronic diseases through the use of intelligent devices: Using the architecture for patients with diabetes. Future Generation Computer Systems, 34(??): 161–175, May 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300277X.

Vrbsky:2013:DPC

[VGC^{+}13] Susan V. Vrbsky, Michael Galloway, Robert Carr, Rahul Nori, and David Grubic. Decreasing power consumption with energy efficient data aware strategies. Future Generation Computer Systems, 29(5): 1152–1163, July 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12002361.

Vijayakumar:2019:MNE

[VGD^{+}19] Pandi Vijayakumar, S. Milton Ganesh, Lazarus Jegatha Deborah, SK Hafizul Islam, Mohammad Mehedi Hassan, Abdulahmeed Alelaiwi, and Giancarlo Fortino. MGPV: a novel and efficient scheme for secure data sharing among mobile users in the public cloud. Future Generation Computer Systems, 95(??): 560–569, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326578.

Vazquez:2010:FTE

[VHML10] Constantino Vázquez, Eduardo Huedo, Rubén S. Montero, and Ignacio M. Llorente. Federation of TeraGrid, EGEE and OSG infrastructures through a metascheduler. Future Generation Computer Systems, 26(7):979–985, July 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vazquez:2011:UCG

[VHML11] Constantino Vázquez, Eduardo Huedo, Rubén S. Montero, and Ignacio M. Llorente. On the use of clouds for grid resource provisioning. Future Generation Computer Systems, 27(5):600–605, May 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vinh:2016:CSA

[Vin16] Phan Cong Vinh. Concurrency of self-* in autonomic systems. Future Generation Computer Systems, 56(??):140–152, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001090.

Vidhyalakshmi:2017:CFE

[VK17] R. Vidhyalakshmi and Vikas Kumar. CORE framework for evaluating the reliability of SaaS products. Future Generation Computer Systems, 72(??): 23–36, July 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302960.

Visegradi:2014:EEG

[VKK14] Ádám Visegrádi, József Kovács, and Peter Kacsuk. Efficient extension of gLite VOs with BOINC based desktop grids. Future Generation Computer Systems, 32(??):13–23, March 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300229X.

vonKistowski:2019:MRE

[vKLA^{+}19] Jóakim von Kistowski, Klaus-Dieter Lange, Jeremy A. Arnold, John Beckett, Hansfried Block, Mike Tricker, Sanjay Sharma, Johann Pais, and Samuel Kounev. Measuring and rating the energy-efficiency of servers. Future Generation Computer Systems, 100(??):579–589, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18312123.

Vora:2019:TTI

[VKT^{+}19] Jayneel Vora, Shriya Kaneriya, Sudeep Tanwar, Sudhanshu Tyagi, Neeraj Kumar, and M. S. Obaidat. TILAA: Tactile Internetbased Ambient Assistant Living in fog environment. Future Generation Computer Systems, 98(??):635– 649, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326839.

vanKessel:2013:UTD

[vKvWD^{+}13] Timo van Kessel, Ben van Werkhoven, Niels Drost, Jason Maassen, Henri E. Bal, and Frank J. Seinstra. User transparent data and task parallel multimedia computing with Pyxis-DT. Future Generation Computer Systems, 29(8):2252– 2261, October 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001192.

Venkataramanan:2019:HCS

[VL19] V. Venkataramanan and S. Lakshmi. Hardware cosimulation of LTE physical layer for mobile network applications. Future Generation Computer Systems, 99(??):124–133, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1832702X.

Vaquero:2013:MDF

[VLAC^{+}13] L. M. Vaquero, S. Sae Lor, J. M. Alcaraz-Calero, D. Niyato, S. Clayman, D. Audsin, and R. Nadarajan. On measuring disturbances in the force: Advanced cloud monitoring systems. Future Generation Computer Systems, 29(8):2007–2008, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000848.

Visheratin:2018:HSA

[VMN^{+}18] Alexander A. Visheratin, Mikhail Melnik, Denis Nasonov, Nikolay Butakov, and Alexander V. Boukhanovsky. Hybrid scheduling algorithm in early warning systems. Future Generation Computer Systems, 79 (part 2)(??): 630–642, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1730540X.

Vigne:2012:SMA

[VMSRM12] Ralph Vigne, Juergen Mangler, Erich Schikuta, and Stefanie Rinderle-Ma. A structured marketplace for arbitrary services. Future Generation Computer Systems, 28(1):48–57, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001063.

Vahdat-Nejad:2019:CAC

[VNAMM19] Hamed Vahdat-Nejad, Elham Asani, Zohreh Mahmoodian, and Mohammad Hossein Mohseni. Context-aware computing for mobile crowd sensing: a survey. Future Generation Computer Systems, 99(??):321–332, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18329583.

Villalba:2017:PBC

[VOCHC17] Luis Javier García Villalba, Ana Lucila Sandoval Orozco, Jocelin Rosales Corripio, and Julio Hernandez-Castro. A PRNU-based counter-forensic method to manipulate smartphone image source identification techniques. Future Generation Computer Systems, 76(??):418– 427, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16305532.

Voith:2012:QSP

[VOS12] Thomas Voith, Karsten Oberle, and Manuel Stein. Quality of service provisioning for distributed data center inter-connectivity enabled by network virtualization. Future Generation Computer Systems, 28 (3):554–562, March 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X11000392.

Villalba:2017:APA

[VOV17] Luis Javier García Villalba, Ana Lucila Sandoval Orozco, and Jorge Maestre Vidal. Advanced Payload Analyzer Preprocessor. Future Generation Computer Systems, 76(??):474–485, November 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16304782.

Vassilakis:2018:ESE

[VPA^{+}18] Costas Vassilakis, Vassilis Poulopoulos, Angeliki Antoniou, Manolis Wallace, George Lepouras, and Martin Lopez Nores. exhiSTORY: Smart exhibits that tell their own stories. Future Generation Computer Systems, 81(??): 542–556, April 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17324093.

Veloudis:2019:ASD

[VPP^{+}19] Simeon Veloudis, Iraklis Paraskakis, Christos Petsos, Yannis Verginadis, Ioannis Patiniotakis, Panagiotis Gouvas, and Gregoris Mentzas. Achieving security-by-design through ontology-driven attributebased access control in cloud environments. Future Generation Computer Systems, 93(??):373–391, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17320782.

Vouros:2010:SIS

[VPT^{+}10] George A. Vouros, Andreas Papasalouros, Konstantinos Tzonas, Alexandros Valarakos, Konstantinos Kotis, Jorge-Arnulfo Quiané-Ruiz, Philippe Lamarre, and Patrick Valduriez. A semantic information system for services and traded resources in Grid e-markets. Future Generation Computer Systems, 26(7):916–933, July 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vasile:2015:RAH

[VPT^{+}15] Mihaela-Andreea Vasile, Florin Pop, Radu-Ioan Tutueanu, Valentin Cristea, and Joanna Kolodziej. Resource-aware hybrid scheduling algorithm in heterogeneous distributed computing. Future Generation Computer Systems, 51(??): 61–71, October 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002532.

Vivekanandan:2012:BFO

[VR12] K. Vivekanandan and D. Ramyachitra. Bacteria foraging optimization for protein sequence analysis on the grid. Future Generation Computer Systems, 28(4):647–656, April 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001944.

Verstaevel:2016:PES

[VRGR16] Nicolas Verstaevel, Christine Régis, Marie-Pierre Gleizes, and Fabrice Robert. Principles and experimentations of self-organizing embedded agents allowing learning from demonstration in ambient robotics. Future Generation Computer Systems, 64(??):78– 87, November 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630067X.

Vilela:2019:PEF

[VRS^{+}19] Pedro H. Vilela, Joel J.

P. C. Rodrigues, Petar Solic, Kashif Saleem, and Vasco Furtado. Performance evaluation of a fogassisted IoT solution for ehealth applications. Future Generation Computer Systems, 97(??):379–386, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18323458.

Varalakshmi:2013:TDA

[VS13] P. Varalakshmi and S. Thamarai Selvi. Thwarting DDoS attacks in grid using information divergence. Future Generation Computer Systems, 29(1):429–441, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11002093.

Venkatasubramanian:2019:ECM

[VS19] D. Venkatasubramanian and B. Shanthi. An efficient control method for PMSM drive using an optimized indirect matrix converter. Future Generation Computer Systems, 98(??):12–17, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18326852.

Viegas:2019:BRT

[VSBN19] Eduardo Viegas, Altair Santin, Alysson Bessani, and Nuno Neves. BigFlow: Real-time and reliable anomaly-based intrusion detection for high-speed networks. Future Generation Computer Systems, 93 (??):473–485, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        https://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X18307635.

Verbelen:2013:GPA

[VSDD13] Tim Verbelen, Tim Stevens, Filip De Turck, and Bart Dhoedt. Graph partitioning algorithms for optimizing software deployment in mobile cloud computing. Future Generation Computer Systems, 29(2):451– 459, February 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001513.

Venkatesh:2019:QID

[VSKS19] K. Venkatesh, L. N. B.

Srinivas, M. B. Mukesh Krishnan, and A. Shanthini. QoS improvisation of delay sensitive communication using SDN based multipath routing for medical applications. Future Generation Computer Systems, 93(??):256–265, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18322325.

Volety:2019:CBW

[VSM^{+}19] Tejaswi Volety, Shalabh Saini, Thomas McGhin, Charles Zhechao Liu, and Kim-Kwang Raymond Choo. Cracking Bitcoin wallets: I want what you have in the wallets. Future Generation Computer Systems, 91(??):136–143, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18302929.

Vissers:2014:DDS

[VSP^{+}14] Thomas Vissers, Thamarai Selvi Somasundaram, Luc Pieters, Kannan Govindarajan, and Peter Hellinckx. DDoS defense system for Web services in a cloud environment. Future Generation Computer Systems, 37 (??):37–45, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1400048X.

Vasudevan:2017:PBA

[VTTK17] Meera Vasudevan, Yu-Chu Tian, Maolin Tang, and Erhan Kozan. Profilebased application assignment for greener and more energy-efficient data centers. Future Generation Computer Systems, 67(??):94–108, February 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302606.

Vinh:2016:NIC

[VV16] Phan Cong Vinh and Emil Vassev. Nature-inspired computation and communication: a formal approach. Future Generation Computer Systems, 56(??): 121–123, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003246.

VandenBossche:2011:EBF

[VVB11] Ruben Van den Bossche, Kurt Vanmechelen, and Jan Broeckhove. An evaluation of the benefits of fine-grained value-based scheduling on general purpose clusters. Future Generation Computer Systems, 27(1):1–9, January 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

VandenBossche:2013:OCE

[VVB13a] Ruben Van den Bossche, Kurt Vanmechelen, and Jan Broeckhove. Online cost-efficient scheduling of deadline-constrained workloads on hybrid clouds. Future Generation Computer Systems, 29(4):973– 985, June 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12002324.

Verboven:2013:BBS

[VVB13b] Sam Verboven, Kurt Vanmechelen, and Jan Broeckhove. Black box scheduling for resource intensive virtual machine workloads with interference models. Future Generation Computer Systems, 29(8):1871– 1884, October 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300099X.

VandenBossche:2015:IRC

[VVB15] Ruben Van den Bossche, Kurt Vanmechelen, and Jan Broeckhove. IaaS reserved contract procurement optimisation with load prediction. Future Generation Computer Systems, 53(??):13–24, December 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002009.

VanHeddeghem:2012:DCC

[VVC^{+}12] Ward Van Heddeghem, Willem Vereecken, Didier Colle, Mario Pickavet, and Piet Demeester. Distributed computing for carbon footprint reduction by exploiting low-footprint energy availability. Future Generation Computer Systems, 28(2):405–414, February 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000859.

vanWerkhoven:2019:KTS

[vW19] Ben van Werkhoven. Kernel Tuner: a searchoptimizing GPU code autotuner. Future Generation Computer Systems, 90(??):347–358, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18313359.

vanWerkhoven:2014:OCO

[vWMBS14] Ben van Werkhoven, Jason Maassen, Henri E. Bal, and Frank J. Seinstra. Optimizing convolution operations on GPUs using adaptive tiling. Future Generation Computer Systems, 30(??): 14–26, January 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001829.

Wang:2018:LRI

[Wan18a] Zhiwei Wang. Leakage resilient ID-based proxy reencryption scheme for access control in fog computing. Future Generation Computer Systems, 87(??):679–685, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17310075.

Wang:2018:PPA

[Wan18b] Zhiwei Wang. A privacypreserving and accountable authentication protocol for IoT end-devices with weaker identity. Future Generation Computer Systems, 82(??):342–348, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17307495.

Wang:2019:PSK

[Wan19]  Zhiwei Wang.  Prov-

ably secure key-aggregate cryptosystems with auxiliary inputs for data sharing on the cloud. Future Generation Computer Systems, 93(??):770–776, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17306970.

Wohrer:2014:MOL

[WBJM14] Alexander Wöhrer, Peter Brezany, Ivan Janciak, and Eduard Mehofer. Modeling and optimizing largescale data flows. Future Generation Computer Systems, 31(??):12–27, February 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002148.

Wettinger:2016:SDA

[WBKL16] Johannes Wettinger, Uwe Breitenbücher, Oliver Kopp, and Frank Leymann. Streamlining DevOps automation for cloud applications using TOSCA as standardized metamodel. Future Generation Computer Systems, 56(??):317–332, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002496.

Wu:2019:VBC

[WBR19] Caesar Wu, Rajkumar Buyya, and Kotagiri Ramamohanarao. Valuebased cloud price modeling for segmented business to business market. Future Generation Computer Systems, 101(??):502–523, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301797.

Weng:2014:EFG

[WC14] Tsung-Hsi Weng and Chung-Ping Chung. Exploiting fine-grain parallelism in the H.264 deblocking filter by operation reordering. Future Generation Computer Systems, 37(??): 76–87, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002367.

Wang:2018:GBH

[WCB^{+}18] Yingxue Wang, Yanan Chen, Md Zakirul Alam Bhuiyan, Yu Han, Shenghui Zhao, and Jianxin Li. Gaitbased human identification using acoustic sensor and deep neural network. Future Generation Computer Systems, 86(??):1228–1237, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17314760.

Wu:2014:GEE

[WCC14] Chia-Ming Wu, Ruay-Shiung Chang, and Hsin-Yu Chan. A green energyefficient scheduling algorithm using the DVFS technique for cloud datacenters. Future Generation Computer Systems, 37(??): 141–147, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001234.

Wang:2016:ECA

[WCC^{+}16] Bei Wang, Yuxia Cheng, Wenzhi Chen, Qinming He, Yang Xiang, Mohammad Mehedi Hassan, and Abdulhameed Alelaiwi. Efficient consolidation-aware VCPU scheduling on multicore virtualization platform. Future Generation Computer Systems, 56(??): 229–237, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002605.

Wang:2015:WBA

[WCF^{+}15] Zhenhua Wang, Haopeng Chen, Ying Fu, Delin Liu, and Yunmeng Ban. Workload balancing and adaptive resource management for the swift storage system on cloud. Future Generation Computer Systems, 51(??):120–131, October 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002404.

Wang:2018:MDC

[WCH^{+}18] Bin Wang, Chao Chen, Ligang He, Bo Gao, Jiadong Ren, Zhangjie Fu, Songling Fu, Yongjian Hu, and Chang-Tsun Li. Modelling and developing conflict-aware scheduling on large-scale data centres. Future Generation Computer Systems, 86(??):995–1007, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17315649.

Wang:2010:DRS

[WCHL10] Chien-Min Wang, Hsi-Min Chen, Chun-Chen Hsu, and Jonathan Lee. Dynamic resource selection heuristics for a nonreserved bidding-based Grid environment. Future Generation Computer Systems, 26(2):183–197, February 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wright:2010:SVE

[WCKW10] H. Wright, R. H. Crompton, S. Kharche, and P. Wenisch. Steering and visualization: Enabling technologies for computational science. Future Generation Computer Systems, 26(3):506–513, March 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wang:2017:TAF

[WCL^{+}17a] Jianfeng Wang, Xiaofeng Chen, Jin Li, Jiaolian Zhao, and Jian Shen. Towards achieving flexible and verifiable search for outsourced database in cloud computing. Future Generation Computer Systems, 67(??):266–275, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301108.

Wang:2017:PSO

[WCL^{+}17b] Jin Wang, Yiquan Cao, Bin Li, Hye jin Kim, and Sungyoung Lee. Particle swarm optimization based clustering algorithm with mobile sink for WSNs. Future Generation Computer Systems, 76(??):452–457, November 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302540.

Wang:2019:EMI

[WCM^{+}19] Shuangyan Wang, Salvatore Cuomo, Gang Mei, Wuyi Cheng, and Nengxiong Xu. Efficient method for identifying influential vertices in dynamic networks using the strategy of local detection and updating. Future Generation Computer Systems, 91(??):10–24, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18307738.

West:2012:TCW

[WCVL12] Andrew G. West, Jian Chang, Krishna K. Venkatasubramanian, and Insup Lee. Trust in collaborative Web applications. Future Generation Computer Systems, 28(8):1238–1251, October 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11000100.

Wang:2018:RAW

[WCW18] Yang Wang, Hongjie Cen, and Sunan Wang. Resource allocation of wireless backhaul in heterogeneous network based on the large-scale MIMO. Future Generation Computer Systems, 88(??):117–126, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18302851.

Wen:2019:ROC

[WCWC19] Ping Wen, Shengduo Chen, Jiarui Wang, and Wei Che. Receiver Operating Characteristics (ROC) analysis for decreased disease risk and elevated treatment response to pegylatedinterferon in chronic hepatitis B patients. Future Generation Computer Systems, 98(??):372–376, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1832692X.

See retraction notice [WCWC20].

Wen:2020:RNR

[WCWC20] Ping Wen, Shengduo Chen, Jiarui Wang, and Wei Che. Retraction notice to “Receiver Operating Characteristics (ROC) Analysis for decreased disease risk and elevated treatment response to pegylatedinterferon in chronic hepatitis B patient” [Future Gener. Comput. Syst. 98 (2019) 372–376]. Future Generation Computer Systems, 107(??):1148, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306518.

See [WCWC19].

Wang:2018:NIA

[WDD18] Yufeng Wang, Wenyong Dong, and Xueshi Dong. A novel ITö algorithm for influence maximization in the large-scale social networks. Future Generation Computer Systems, 88(??):755–763, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17315522.

Wang:2018:PAS

[WDJC18] YuPeng Wang, Xin Dai, Jason J. Jung, and Chang Choi. Performance analysis of smart cultural heritage protection oriented wireless networks. Future Generation Computer Systems, 81(??):593–600, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17305587.

Wazid:2019:DSK

[WDKV19] Mohammad Wazid, Ashok Kumar Das, Neeraj Kumar, and Athanasios V. Vasilakos. Design of secure key management and user authentication scheme for fog computing services. Future Generation Computer Systems, 91(??):475–492, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18303959.

Weekley:2016:CHQ

[WdL16] Jeffrey D. Weekley and Cees de Laat. CineGrid, high quality media streaming and processing on advanced photonic networks. Future Generation Computer Systems, 54(??):292– 295, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002344.

Wozniak:2019:MJW

[WDR^{+}19] Justin M. Wozniak, Matthieu Dorier, Robert Ross, Tong Shu, Tahsin Kurc, Li Tang, Norbert Podhorszki, and Matthew Wolf. MPI jobs within MPI jobs: a practical way of enabling task-level fault-tolerance in HPC workflows. Future Generation Computer Systems, 101(??):576–589, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1830757X.

Wu:2019:SAS

[WDW^{+}19] Dapeng Wu, Lingli Deng, Honggang Wang, Keyu Liu, and Ruyan Wang. Similarity aware safety multimedia data transmission mechanism for Internet of Vehicles. Future Generation Computer Systems, 99(??):609–623, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18318211.

Wang:2019:SFE

[WDZ19] Xiaofen Wang, Hong-Ning Dai, and Ke Zhang. Secure and flexible economic data sharing protocol based on ID-based dynamic exclusive broadcast encryption in economic system. Future Generation Computer Systems, 99(??):177–185, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18316145.

Weinberg:2011:PSH

[Wei11] Richard Weinberg. Producing and streaming high resolution digital movies of microscopic subjects. Future Generation Computer Systems, 27(7):906– 913, July 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wenyin:2010:DPT

[WFQ^{+}10] Liu Wenyin, Ning Fang, Xiaojun Quan, Bite Qiu, and Gang Liu. Discovering phishing target based on Semantic Link Network. Future Generation Computer Systems, 26(3):381– 388, March 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wong:2013:UFD

[WG13] Adam K. L. Wong and Andrzej M. Goscinski. A unified framework for the deployment, exposure and access of HPC applications as services in clouds. Future Generation Computer Systems, 29(6):1333–1344, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000307.

Wang:2019:LNF

[WGC19] Hao Wang, Chaonian Guo, and Shuhan Cheng. LoC

— a new financial loan management system based on smart contracts. Future Generation Computer Systems, 100(??):648–655, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18324233.

Weinreich:2015:ESK

[WGM15] Rainer Weinreich, Iris Groher, and Cornelia Miesbauer. An expert survey on kinds, influence factors and documentation of design decisions in practice. Future Generation Computer Systems, 47(??): 145–160, June 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002556.

Wu:2019:BFD

[WGX^{+}19] Shaoen Wu, Hanqing Guo, Junhong Xu, Shangyue Zhu, and Honggang Wang. In-band full duplex wireless communications and networking for IoT devices: Progress, challenges and opportunities. Future Generation Computer Systems, 92(??):705–714, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17307902.

Wei:2019:ITP

[WH19] Zequn Wei and Jin-Kao Hao. Iterated two-phase local search for the set-union knapsack problem. Future Generation Computer Systems, 101(??):1005–1017, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19306569.

Wang:2019:EBT

[WHBC19] Tong Wang, Azhar Hussain, Muhammad Nasir Mumtaz Bhutta, and Yue Cao. Enabling bidirectional traffic mobility for ITS simulation in smart city environments. Future Generation Computer Systems, 92(??):342–356, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18303807.

Wu:2019:IAM

[WHCW19] Binghong Wu, Kuangrong Hao, Xin Cai, and Tong Wang. An integrated algorithm for multi-agent faulttolerant scheduling based on MOEA. Future Generation Computer Systems, 94 (??):51–61, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18309749.

Wang:2018:MMI

[WHCZ18] Jingjing Wang, Xiaoguang Han, Zengqiang Chen, and Qing Zhang. Model matching of input/output asynchronous sequential machines based on the semitensor product of matrices. Future Generation Computer Systems, 83(??):468– 475, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17304557.

Wang:2013:HVE

[WHMO13] Miao Wang, Viliam Holub, John Murphy, and Patrick O’Sullivan. High volumes of event stream indexing and efficient multi-keyword searching for cloud monitoring. Future Generation Computer Systems, 29(8):1943–1962, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001131.

Wang:2017:ERS

[WHS^{+}17] Shangguang Wang, Lin Huang, Lei Sun, Ching-Hsien Hsu, and Fangchun Yang. Efficient and reliable service selection for heterogeneous distributed software systems. Future Generation Computer Systems, 74(??):158–167, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003970.

Wang:2018:PPR

[WHS^{+}18] Huaqun Wang, Debiao He, Yanfei Sun, Neeraj Kumar, and Kim-Kwang Raymond Choo. PAT: A precise reward scheme achieving anonymity and traceability for crowdcomputing in public clouds. Future Generation Computer Systems, 79 (part 1)(??): 262–270, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1630735X.

Wang:2016:SOM

[WHW16] Yi-Rong Wang, Kuo-Chan Huang, and Feng-Jian Wang. Scheduling online mixed-parallel workflows of rigid tasks in heterogeneous multi-cluster environments. Future Generation Computer Systems, 60 (??):35–47, July 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16000261.

Woodman:2017:APP

[WHW17] Simon Woodman, Hugo Hiden, and Paul Watson. Applications of provenance in performance prediction and data storage optimisation. Future Generation Computer Systems, 75(??):299–309, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300213.

Wu:2017:SDP

[WHYZ17] Renke Wu, Linpeng Huang, Peng Yu, and Haojie Zhou. SunwayMR: a distributed parallel computing framework with convenient dataintensive applications programming. Future Generation Computer Systems, 71 (??):43–56, June 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300584.

Wu:2018:EDF

[WHYZ18] Renke Wu, Linpeng Huang, Peng Yu, and Haojie Zhou. EDAWS: a distributed framework with efficient data analytics workspace towards discriminative services for critical infrastructures. Future Generation Computer Systems, 81(??): 78–93, April 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17311184.

Wu:2019:RRH

[WHZ19] Renke Wu, Linpeng Huang, and Haojie Zhou. RHKV: An RDMA and HTM friendly key-value store for data-intensive computing. Future Generation Computer Systems, 92(??):162– 177, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18301080.

Wang:2010:MSA

[WHZL10] Jue Wang, Changjun Hu, Jilin Zhang, and Jianjiang Li. Message scheduling for array re-decomposition on distributed memory systems. Future Generation Computer Systems, 26(2): 281–290, February 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wang:2018:MPI

[WJLW18] Feng Wang, Wenjun Jiang, Xiaolin Li, and Guojun Wang. Maximizing positive influence spread in online social networks via fluid dynamics. Future Generation Computer Systems, 86(??):1491–1502, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17310233.

Wang:2018:MLP

[WJS^{+}18] Meisong Wang, Prem Prakash Jayaraman, Ellis Solaiman, Lydia Y. Chen, Zheng Li, Song Jun, Dimitrios Georgakopoulos, and Rajiv Ranjan. A multi-layered performance analysis for cloud-based topic detection and tracking in big data applications. Future Generation Computer Systems, 87(??):580–590, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17315935.

Wang:2017:SPA

[WJZ^{+}17] Yuzhu Wang, Jinrong Jiang, He Zhang, Xiao Dong, Lizhe Wang, Rajiv Ranjan, and Albert Y. Zomaya. A scalable parallel algorithm for atmospheric general circulation models on a multi-core cluster. Future Generation Computer Systems, 72(??):1– 10, July 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17302017.

Wang:2013:EAP

[WKC^{+}13] Lizhe Wang, Samee U. Khan, Dan Chen, Joanna Kolodziej, Rajiv Ranjan, Cheng zhong Xu, and Albert Zomaya. Energyaware parallel task scheduling in a cluster. Future Generation Computer Systems, 29(7):1661–1670, September 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000484.

Widodo:2017:NCD

[WLA17a] Ryan N. S. Widodo, Hyotaek Lim, and Mohammed Atiquzzaman. A new content-defined chunking algorithm for data deduplication in cloud storage. Future Generation Computer Systems, 71(??): 145–156, June 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16305829.

Widodo:2017:SSD

[WLA17b] Ryan N. S. Widodo, Hyotaek Lim, and Mohammed Atiquzzaman. SDM: Smart deduplication for mobile cloud storage. Future Generation Computer Systems, 70(??):64– 73, May 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302084.

Witanto:2018:ASD

[WLA18a] Joseph Nathanael Witanto, Hyotaek Lim, and Mohammed Atiquzzaman. Adaptive selection of dynamic VM consolidation algorithm using neural network for cloud resource management. Future Generation Computer Systems, 87(??):35–42, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17323336.

Witanto:2018:SGF

[WLA18b] Joseph Nathanael Witanto, Hyotaek Lim, and Mohammed Atiquzzaman. Smart government framework with geo-crowdsourcing and social media analysis. Future Generation Computer Systems, 89(??):1– 9, December 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18304461.

Weyers:2011:ICR

[WLB11] Benjamin Weyers, Wolfram Luther, and Nelson Baloian. Interface creation and redesign techniques in collaborative learning scenarios. Future Generation Computer Systems, 27 (1):127–138, January 2011.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wang:2019:MKS

[WLGL19] Haiyu Wang, Xuelian Li, Juntao Gao, and Wei Li. MOBT: a kleptographically-secure hierarchicaldeterministic wallet for multiple offline Bitcoin transactions. Future Generation Computer Systems, 101(??):315–326, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1733011X.

Wang:2016:ATN

[WLH16] Jue Wang, Chun Liu, and Yuehui Huang. Auto tuning for new energy dispatch problem: a case study. Future Generation Computer Systems, 54(??):501– 506, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000576.

Wu:2019:QQD

[WLH^{+}19] Ziyan Wu, Zhihui Lu, Patrick C. K. Hung, Shih-Chia Huang, Yu Tong, and Zhenfang Wang. QaMeC: A QoS-driven IoVs application optimizing deployment scheme in multimedia edge clouds. Future Generation Computer Systems, 92(??):17–28, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18317540.

Wu:2018:EEH

[WLHH18] WenTai Wu, WeiWei Lin, Ching-Hsien Hsu, and LiGang He. Energy-efficient Hadoop for big data analytics and computing: a systematic review and research insights. Future Generation Computer Systems, 86(??):1351–1367, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17318174.

Wang:2019:IKF

[WLL^{+}19a] Jun Wang, Jing Luo, Xiaozhu Liu, Yongkai Li, Shubo Liu, Rongbo Zhu, and Ashiq Anjum. Improved Kalman filter based differentially private streaming data release in cognitive computing. Future Generation Computer Systems, 98(??):541–549, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18330838.

Wang:2019:TTA

[WLL^{+}19b] Ximeng Wang, Yun Liu, Jie Lu, Fei Xiong, and Guangquan Zhang. Tru-GRC: Trust-aware group recommendation with virtual coordinators. Future Generation Computer Systems, 94(??):224–236, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18320764.

Wang:2016:NFS

[WLLF16] Dingxian Wang, Xiao Liu, Hangzai Luo, and Jianping Fan. A novel framework for semantic entity identification and relationship integration in large scale text data. Future Generation Computer Systems, 64(??):198–210, November 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002551.

Wu:2017:DIS

[WLML17] Suzhen Wu, Kuan-Ching Li, Bo Mao, and Minghong Liao. DAC: Improving storage availability with Deduplication-Assisted Cloudof-Clouds. Future Generation Computer Systems, 74(??):190–198, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300073.

Wang:2010:FFD

[WLP10] Kai Wang, Jianhua Li, and Li Pan. Fast file dissemination in peer-to-peer networks with upstream bandwidth constraint. Future Generation Computer Systems, 26(7):986–1002, July 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Woo:2018:RIS

[WLP18] Min Woo Woo, JongWhi Lee, and KeeHyun Park. A reliable IoT system for personal healthcare devices. Future Generation Computer Systems, 78 (part 2)(??):626–640, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17305423.

Wu:2010:OPS

[WLQ10] Tiejun Wu, Maozhen Li, and Man Qi. Optimizing peer selection in BitTorrent networks with genetic algorithms. Future Generation Computer Systems, 26(8):

1151–1156, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wu:2018:NMG

[WLRL18] Xiaoban Wu, Peilong Li, Yongyi Ran, and Yan Luo. Network measurement for 100 GbE network links using multicore processors. Future Generation Computer Systems, 79 (part 1)(??):180– 189, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17307690.

Wu:2018:LRT

[WLS^{+}18] Fan Wu, Xiong Li, Arun Kumar Sangaiah, Lili Xu, Saru Kumari, Liuxi Wu, and Jian Shen. A lightweight and robust two-factor authentication scheme for personalized healthcare systems using wireless medical sensor networks. Future Generation Computer Systems, 82(??):727–737, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1730523X.

Wang:2018:DAM

[WLW^{+}18] Wei Wang, Yuanyuan Li, Xing Wang, Jiqiang Liu, and Xiangliang Zhang. Detecting Android malicious apps and categorizing benign apps with ensemble of classifiers. Future Generation Computer Systems, 78 (part 3)(??):987–994, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300742.

Wei:2019:SQA

[WLW^{+}19] Xiaohui Wei, Yuanyuan Liu, Xingwang Wang, Bingyi Sun, Shang Gao, and Jon Rokne. A survey on quality-assurance approximate stream processing and applications. Future Generation Computer Systems, 101(??):1062– 1080, December 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19314414.

Wu:2018:ABM

[WLXZ18] Ying Wu, MuHua Liu, Rui Xue, and Rui Zhang. Attribute-based multi-function verifiable computation. Future Generation Computer Systems, 78 (part 3)(??):995–1004, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302546.

Wang:2011:PHB

[WLYL11] Xiaogang Wang, Junzhou Luo, Ming Yang, and Zhen Ling. A potential HTTPbased application-level attack against Tor. Future Generation Computer Systems, 27(1):67–77, January 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wang:2014:INA

[WLZ^{+}14] Yan Wang, Wenzao Li, Jiliu Zhou, Xiaohua Li, and Yifei Pu. Identification of the normal and abnormal heart sounds using wavelet-time entropy features based on OMS-WPD. Future Generation Computer Systems, 37(??): 488–495, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1400034X.

Wang:2016:OMD

[WLZ^{+}16] Shangguang Wang, Tao Lei, Lingyan Zhang, Ching-Hsien Hsu, and Fangchun Yang. Offloading mobile data traffic for QoS-aware service provision in vehicular cyber-physical systems. Future Generation Computer Systems, 61(??):118– 127, August 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003179.

Wu:2019:BLA

[WLZ^{+}19] Pin Wu, Zhihui Lu, Quan Zhou, Zhidan Lei, Xiaoqiang Li, Meikang Qiu, and Patrick C. K. Hung. Bigdata logs analysis based on seq2seq networks for cognitive Internet of Things. Future Generation Computer Systems, 90(??):477– 488, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18309233.

Watanobe:2014:HIA

[WM14] Yutaka Watanobe and Nikolay Mirenkov. Hybrid intelligence aspects of programming in *AIDA algorithmic pictures. Future Generation Computer Systems, 37(??):417–428, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002951.

Wei:2018:PAU

[WMA18] Jianliang Wei, Fei Meng, and N. Arunkumar. A personalized authoritative user-based recommendation for social tagging. Future Generation Computer Systems, 86(??):355–361, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18303078.

Waqas:2019:CTA

[WMA^{+}19] Muhammad Waqas, Saif Ur Rehman Malik, Saeed Akbar, Adeel Anjum, and Naveed Ahmad. Convergence time analysis of OSPF routing protocol using social network metrics. Future Generation Computer Systems, 94(??):62– 71, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1732753X.

Wannous:2017:TOI

[WMBV17] Rouaa Wannous, Jamal Malki, Alain Bouju, and Cécile Vincent. Trajectory ontology inference considering domain and temporal dimensions — application to marine mammals. Future Generation Computer Systems, 68(??):491– 499, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1600025X.

Wang:2019:SGP

[WMC19] Shuangyan Wang, Gang Mei, and Salvatore Cuomo. A simple and generic paradigm for creating complex networks using the strategy of vertex selectingand-pairing. Future Generation Computer Systems, 100(??):994–1004, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19304960.

Wu:2018:AUG

[WMJW18] Song Wu, Chao Mei, Hai Jin, and Duoqiang Wang. Android unikernel: Gearing mobile code offloading towards edge computing. Future Generation Computer Systems, 86(??):694–703, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17329734.

Wang:2016:GCC

[WMK16] Yongge Wang, Qutaibah M. Malluhi, and Khaled MD Khan. Garbled computation in cloud. Future Generation Computer Systems, 62(??):54–65, September 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003416.

Wen:2014:DMM

[WMLS14] Jing Wen, Yan Ma, Peng Liu, and Shengtao Sun. Distributed multipliers in MWM for analyzing job arrival processes in massive HPC workload datasets. Future Generation Computer Systems, 37(??):335– 344, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002732.

Wang:2016:UMC

[WMQ^{+}16] Bin Wang, Ruhui Ma, Zhengwei Qi, Jianguo Yao, and Haibing Guan. A user mode CPU–GPU scheduling framework for hybrid workloads. Future Generation Computer Systems, 63(??):25–36, October 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300553.

Wang:2017:CES

[WMX^{+}17] Xu An Wang, Jianfeng Ma, Fatos Xhafa, Mingwu Zhang, and Xiaoshuang Luo. Cost-effective secure e-health cloud system using identity based cryptographic techniques. Future Generation Computer Systems, 67(??):242–254, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302588.

Wang:2018:PHP

[WMY^{+}18] Lizhe Wang, Yan Ma, Jining Yan, Victor Chang, and Albert Y. Zomaya. pipsCloud: High performance cloud computing for remote sensing big data management and processing. Future Generation Computer Systems, 78 (part 1)(??):353–368, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301923.

Wang:2010:PIA

[WN10] Yufeng Wang and Akihiro Nakao. Poisonedwater: An improved approach for accurate reputation ranking in P2P networks. Future Generation Computer Systems, 26(8):1317–1326, October 2010. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wang:2019:GAA

[WNR19] Lei Wang, Jianwei Niu, and Joel J. P. C. Rodrigues. GMA: An adult account identification algorithm on Sina Weibo using behavioral footprints. Future Generation Computer Systems, 93(??):942– 951, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17308804.

Witkowski:2013:PPC

[WOPW13] M. Witkowski, A. Oleksiak, T. Piontek, and J. Weglarz. Practical power consumption estimation for real life HPC applications. Future Generation Computer Systems, 29(1):208– 217, January 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001392.

Walulya:2018:VMC

[WPGN^{+}18] Ivan Walulya, Dimitris Palyvos-Giannas, Yiannis Nikolakopoulos, Vincenzo Gulisano, Marina Papatriantafilou, and Philippas Tsigas. Viper: a module for communicationlayer determinism and scaling in low-latency stream processing. Future Generation Computer Systems, 88(??):297–308, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17326791.

Wu:2016:TDB

[WPJ16] Song Wu, Yaqiong Peng, and Hai Jin. Time Donating Barrier for efficient task scheduling in competitive multicore systems. Future Generation Computer Systems, 54(??):469– 477, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000977.

Wu:2018:OSQ

[WPS^{+}18] Wanqing Wu, Sandeep Pirbhulal, Arun Kumar Sangaiah, Subhas Chandra Mukhopadhyay, and Guanglin Li. Optimization of signal quality over comfortability of textile electrodes for ECG monitoring in fog computing based medical applications. Future Generation Computer Systems, 86(??):515–526, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18300621.

Wang:2019:PRP

[WPY19] Hongbing Wang, Shunshun Peng, and Qi Yu. A parallel refined probabilistic approach for QoS-aware service composition. Future Generation Computer Systems, 98(??):609–626, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314262.

Wang:2014:AUC

[WQ14] Lan Wang and Wenyu Qu. Advances in ubiquitous computing and communications. Future Generation Computer Systems, 38(??):11–12, September 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000806.

Wang:2015:HRR

[WQG15] Jihe Wang, Meikang Qiu, and Bing Guo. High reliable real-time bandwidth scheduling for virtual machines with hidden Markov predicting in telehealth platform. Future Generation Computer Systems, 49(??):68–76, August 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1400154X.

Wang:2019:ICR

[WQZ19] Jian Wang, Kuoyuan Qiao, and Zhiyong Zhang. An improvement for combination rule in evidence theory. Future Generation Computer Systems, 91(??):1–9, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18300839.

Wang:2017:MLS

[WRCC17] Ru Wang, Seungmin Rho, Bo-Wei Chen, and Wandong Cai. Modeling of large-scale social network services based on mechanisms of information diffusion: Sina Weibo as a case study. Future Generation Computer Systems, 74(??):291–301, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300620.

Wright:2019:PMB

[Wri19] Steven A. Wright. Performance modeling, benchmarking and simulation of high performance computing systems. Future Generation Computer Systems, 92(??):900–902, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18328590.

Wang:2015:STT

[WRK^{+}15] Lizhe Wang, Rajiv Ranjan, Joanna Kolodziej, Albert Zomaya, and Leila Alem. Software tools and techniques for big data computing in healthcare clouds. Future Generation Computer Systems, 43–44(??): 38–39, February 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002283.

Wu:2010:ISA

[WS10] Chao-Chin Wu and Ren-Yi Sun. An integrated security-aware job scheduling strategy for large-scale computational Grids. Future Generation Computer Systems, 26(2):198–206, February 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wesner:2015:SST

[WSB^{+}15] Stefan Wesner, Lutz Schubert, Rosa M. Badia, Antonio Rubio, Pier Paolucci, and Roberto Giorgi. Special section on terascale computing. Future Generation Computer Systems, 53(??):88–89, December 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002381.

Wang:2019:ITS

[WSC^{+}19] Bo Wang, Ying Song, Jie Cao, Xiao Cui, and Ling Zhang. Improving task scheduling with parallelism awareness in heterogeneous computational environments. Future Generation Computer Systems, 94(??):419–429, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18312056.

Wei:2016:RRR

[WSH^{+}16] Hongxing Wei, Zhenzhou Shao, Zhen Huang, Renhai Chen, Yong Guan, Jindong Tan, and Zili Shao. RT-ROS: a realtime ROS architecture on multi-core processors. Future Generation Computer Systems, 56(??):171–178, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001831.

Wang:2019:MEM

[WSL^{+}19] Jia Wang, Houbing Song, Jianqiang Li, Qiuzhen Lin, and Lee-Ming Cheng. Modular exponential multivariate sequence and its application to lightweight security design. Future Generation Computer Systems, 98(??):435–443, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18315401.

Wang:2018:PPS

[WSN18] Shuo Wang, Richard Sinnott, and Surya Nepal. Privacy-protected statistics publication over social media user trajectory streams. Future Generation Computer Systems, 87(??):792–802, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17300596.

Wu:2016:BPK

[WSQ^{+}16] Qianhong Wu, Yang Sun, Bo Qin, Jiankun Hu, Weiran Liu, Jianwei Liu, and Yong Ding. Batch public key cryptosystem with batch multi-exponentiation. Future Generation Computer Systems, 62(??):196– 204, September 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003933.

Wahid:2018:BDA

[WSQ^{+}18] Abdul Wahid, Munam Ali Shah, Faisal Fayyaz Qureshi, Hafsa Maryam, Rahat Iqbal, and Victor Chang. Big data analytics for mitigating broadcast storm in vehicular content centric networks. Future Generation Computer Systems, 86(??):1301–1320, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17322616.

Wang:2010:RFR

[WSU^{+}10] Jian Wang, Ruimin Shen, Carsten Ullrich, Heng Luo, and Changyong Niu. Resisting free-riding behavior in BitTorrent. Future Generation Computer Systems, 26(8):1285–1299, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wu:2019:SFB

[WSY^{+}19] Qiang Wu, Jun Shen, Binbin Yong, Jianqing Wu, Fucun Li, Jinqiang Wang, and Qingguo Zhou. Smart fog based workflow for traffic control networks. Future Generation Computer Systems, 97(??):825–835, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18325809.

Wang:2018:TAO

[WSZC18] Haitao Wang, Lihua Song, Guomin Zhang, and Hui Chen. Timetable-aware opportunistic DTN routing for vehicular communications in battlefield environments. Future Generation Computer Systems, 83(??):95–103, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17303205.

Wandelt:2018:QQR

[WSZH18] Sebastian Wandelt, Xiaoqian Sun, Massimiliano Zanin, and Shlomo Havlin. QRE: Quick Robustness Estimation for large complex networks. Future Generation Computer Systems, 83(??):413– 424, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17302169.

Wang:2014:AUB

[WTG^{+}14] Zhenhua Wang, Lai Tu, Zhe Guo, Laurence T. Yang, and Benxiong Huang. Analysis of user behaviors by mining large network data sets. Future Generation Computer Systems, 37(??):429–437, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000405.

Woodley:2019:PKT

[WTG^{+}19] Alan Woodley, Ling-Xiang Tang, Shlomo Geva, Richi Nayak, and Timothy Chappell. Parallel K-Tree: a multicore, multinode solution to extreme clustering. Future Generation Computer Systems, 99(??):333– 345, October 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17319271.

Wang:2017:PVI

[WTM^{+}17] Junchao Wang, Arie Taal, Paul Martin, Yang Hu, Huan Zhou, Jianmin Pang, Cees de Laat, and Zhiming Zhao. Planning virtual infrastructures for time critical applications with multiple deadline constraints. Future Generation Computer Systems, 75(??):365– 375, October 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301905.

Wimmer:2013:AEA

[WTP^{+}13] Jason Wimmer, Michael Towsey, Birgit Planitz, Ian Williamson, and Paul Roe. Analysing environmental acoustic data through collaboration and automation. Future Generation Computer Systems, 29(2):560–568, February 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1200057X.

Wang:2013:GHM

[WTR^{+}13] Lizhe Wang, Jie Tao, Rajiv Ranjan, Holger Marten, Achim Streit, Jingying Chen, and Dan Chen. G-Hadoop: MapReduce across distributed data centers for dataintensive computing. Future Generation Computer Systems, 29(3):739–750, March 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001744.

Wu:2014:IFF

[WTS14] Mu-En Wu, Raylin Tso, and Hung-Min Sun. On the improvement of Fermat factorization using a continued fraction technique. Future Generation Computer Systems, 30(??):162– 168, January 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001222.

Wapet:2019:PPN

[WTTH19] Lavoisier Wapet, Alain Tchana, Giang Son Tran, and Daniel Hagimont. Preventing the propagation of a new kind of illegitimate apps. Future Generation Computer Systems, 94(??): 368–380, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18313505.

Wu:2016:EES

[Wu16] Jun Wu. Energy-efficient scheduling of real-time tasks with shared resources. Future Generation Computer Systems, 56(??): 179–191, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15001879.

Wang:2011:MDE

[WW11] Guojun Wang and Jie Wu. Multi-dimensional evidence-based trust management with multi-trusted paths. Future Generation Computer Systems, 27(5): 529–538, May 2011. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wu:2013:SBS

[WW13] Kaigui Wu and Changze Wu. State-based search strategy in unstructured P2P. Future Generation Computer Systems, 29(1):381–386, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001415.

Wu:2019:BAM

[WWA19] Zhichao Wu, Han Wang, and N. Arunkumar. Bayesian analysis model for the use of anesthetic analgesic drugs in cancer patients based on geometry reconstruction. Future Generation Computer Systems, 93(??):170–175, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18321939.

Wang:2014:NMO

[WWC14] Xiaoli Wang, Yuping Wang, and Yue Cui. A new multi-objective bilevel programming model for energy and locality aware multi-job scheduling in cloud computing. Future Generation Computer Systems, 36(??):91– 101, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002689.

Wang:2013:RPR

[WWCN13] Pengyao Wang, Jianqin Wang, Ying Chen, and Guangyuan Ni. Rapid processing of remote sensing images based on cloud computing. Future Generation Computer Systems, 29(8):1963–1968, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001015.

Wang:2014:ADE

[WWD^{+}14] Guodong Wang, Yulei Wu, Ke Dou, Yongmao Ren, and Jun Li. AppTCP: the design and evaluation of application-based TCP for e-VLBI in fast long distance networks. Future Generation Computer Systems, 39(??):67–74, October 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300280X.

Wang:2018:ASA

[WWDF18] Huaqun Wang, Zhiwei Wang, and Josep Domingo-Ferrer. Anonymous and secure aggregation scheme in fog-based public cloud computing. Future Generation Computer Systems, 78 (part 2)(??):712–719, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302558.

Wang:2019:LPS

[WWG19a] Jihe Wang, Danghui Wang, and Bing Guo. A lowpower sensor polling for aggregated-task context on mobile devices. Future Generation Computer Systems, 98(??):362–371, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18312706.

Wang:2019:TED

[WWG^{+}19b] Ning Wang, Zhigang Wang, Yu Gu, Yubin Bao, and Ge Yu. TSH: Easy-to-be distributed partitioning for large-scale graphs. Future Generation Computer Systems, 101(??):804–818, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17327723.

Wang:2017:CAA

[WWH^{+}17] Xing Wang, Wei Wang, Yongzhong He, Jiqiang Liu, Zhen Han, and Xiangliang Zhang. Characterizing Android apps’ behavior for effective detection of malapps at large scale. Future Generation Computer Systems, 75(??): 30–45, October 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17307720.

Wang:2019:DFG

[WWH^{+}19] Haifang Wang, Zhongjie Wang, Sihang Hu, Xiaofei Xu, Shiping Chen, and Zhiying Tu. DUSKG: a fine-grained knowledge graph for effective personalized service recommendation. Future Generation Computer Systems, 100(??):600–617, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1832911X.

Wei:2019:CEP

[WWP19] Li Wei, Chongling Wu, and Xie Peng. Clinical effect of percutaneous vertebroplasty (PVP) in spinal surgery on senile osteoporotic compressible fractures of the spine. Future Generation Computer Systems, 98(??):197–200, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326359.

See retraction notice [WWP20].

Wei:2020:RNC

[WWP20] Li Wei, Chongling Wu, and Xie Peng. Retraction notice to “Clinical effect of percutaneous vertebroplasty (pvp) in spinal surgery on senile osteoporotic compressible fractures of the spine” [Future Gener. Comput. Syst. 98 (2019) 197–200]. Future Generation Computer Systems, 107(??):1149, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2030652X.

See [WWP19].

Wang:2018:LAS

[WWQ^{+}18] Jihe Wang, Danghui Wang, Meikang Qiu, Yao Chen, and Bing Guo. A localityaware shuffle optimization on fat-tree data centers. Future Generation Computer Systems, 89(??):31– 43, December 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17327310.

Wang:2016:TIA

[WWRS16] Yongzhi Wang, Jinpeng Wei, Shaolei Ren, and Yulong Shen. Toward integrity assurance of outsourced computing — a game theoretic perspective. Future Generation Computer Systems, 55(??):87– 100, February 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002630.

Wang:2019:IBF

[WWSL19] TianTian Wang, KeChao Wang, XiaoHong Su, and Zhang Lei. Invariant based fault localization by analyzing error propagation. Future Generation Computer Systems, 94(??):549– 563, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18319186.

Wu:2016:PBP

[WWT^{+}16] Fuhui Wu, Qingbo Wu, Yusong Tan, Rongzhen Li, and Wei Wang. PCP-B^{2}: Partial critical path budget balanced scheduling algorithms for scientific workflow applications. Future Generation Computer Systems, 60(??):22– 34, July 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16000066.

Wei:2018:CMJ

[WWTF18] Xianglin Wei, Tongxiang Wang, Chaogang Tang, and Jianhua Fan. Collaborative mobile jammer tracking in multi-hop wireless network. Future Generation Computer Systems, 78 (part 3)(??):1027–1039, January 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306951.

Wang:2017:DDB

[WWVJ17] Yufeng Wang, Li Wei, Athanasios V. Vasilakos, and Qun Jin. Device-to-device based mobile social networking in proximity (MSNP) on smartphones: Framework, challenges and prototype. Future Generation Computer Systems, 74(??):241–253, September 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003337.

Wang:2016:PSR

[WWW^{+}16] Yujue Wang, Qianhong Wu, Duncan S. Wong, Bo Qin, Jian Mao, and Yong Ding. Provably secure robust optimistic fair exchange of distributed signatures. Future Generation Computer Systems, 62(??):29–39, September 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300565.

Wan:2017:POT

[WWX^{+}17] Long Wan, Lijun Wei, Naixue Xiong, Jinjiang Yuan, and Jiacai Xiong. Pareto optimization for the two-agent scheduling problems with linear non-increasing deterioration based on Internet of Things. Future Generation Computer Systems, 76(??):293–300, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302990.

Wu:2018:RII

[WWZ18] Shaofei Wu, Mingqing Wang, and Yuntao Zou. Research on Internet information mining based on agent algorithm. Future Generation Computer Systems, 86(??):598–602, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18302905.

Wen:2019:ECA

[WWZ^{+}19] Yiping Wen, Zhibin Wang, Yu Zhang, Jianxun Liu, Buqing Cao, and Jinjun Chen. Energy and cost aware scheduling with batch processing for instance-intensive IoT workflows in clouds. Future Generation Computer Systems, 101(??):39–50, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331066.

Wang:2019:EOP

[WWZC19] Zhuowei Wang, Hao Wang, Wuqing Zhao, and Lianglun Cheng. Energy optimization of parallel programs in a heterogeneous system by combining processor core-shutdown and dynamic voltage scaling. Future Generation Computer Systems, 92(??):198– 209, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18305569.

Wang:2018:RAD

[WWZZ18] Yan Wang, Zhensen Wu, Yuanjian Zhu, and Pei Zhang. Research on anomaly detection algorithm based on generalization latency of telecommunication network. Future Generation Computer Systems, 85(??):9–18, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324627.

Wu:2018:HPC

[WXGM18] Yulei Wu, Yang Xiang, Jingguo Ge, and Peter Muller. High-performance computing for big data processing. Future Generation Computer Systems, 88(??):693–695, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317679.

Wang:2016:SHI

[WXLY16] Wei Wang, Peng Xu, Hui Li, and Laurence Tianruo Yang. Secure hybridindexed search for high efficiency over keyword searchable ciphertexts. Future Generation Computer Systems, 55(??):353–361, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001484.

Wang:2017:HIT

[WXPL17] Mei Wang, Meng Xiao, Sancheng Peng, and Guohua Liu. A hybrid index for temporal big data. Future Generation Computer Systems, 72(??):264–272, July 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302527.

Wang:2015:DCA

[WXYL15] Wei Wang, Peng Xu, Laurence Tianruo Yang, and Hui Li. A design for cloudassisted Fair-Play Management System of online contests with provable security. Future Generation Computer Systems, 52(??):137–146, November 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002738.

Wang:2018:RAM

[WXZ^{+}18a] Qian Wang, Wei Xiong, Yin Zhang, Ning Pan, Zhuo Yu, Enmin Song, and Chih-Cheng Hung. Remote analysis of myocardial fiber information in vivo assisted by cloud computing. Future Generation Computer Systems, 85(??):146–159, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17326602.

Wang:2018:SAC

[WXZ^{+}18b] Tao Wang, Jiwei Xu, Wenbo Zhang, Zeyu Gu, and Hua Zhong. Selfadaptive cloud monitoring with online anomaly detection. Future Generation Computer Systems, 80(??): 89–101, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1730376X.

Wu:2011:NMA

[WXZL11] Jun Wu, Xin Xu, Pengcheng Zhang, and Chunming Liu. A novel multi-agent reinforcement learning approach for job scheduling in Grid computing. Future Generation Computer Systems, 27(5):430–439, May 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wei:2017:RAD

[WY17] Shang-Chia Wei and Wei-Chang Yeh. Resource allocation decision model for dependable and costeffective grid applications based on grid bank. Future Generation Computer Systems, 77(??):12–28, December 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17303941.

Wang:2019:IHC

[WY19] Xiuhui Wang and Ke Yan.

Immersive human-computer interactive virtual environment using large-scale display system. Future Generation Computer Systems, 96(??):649–659, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17316229.

Wang:2011:OMR

[WYBS11] Xiaofeng Wang, Chee Shin Yeo, Rajkumar Buyya, and Jinshu Su. Optimizing the makespan and reliability for workflow applications with reputation and a look-ahead genetic algorithm. Future Generation Computer Systems, 27(8): 1124–1134, October 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wu:2017:DBN

[WYH^{+}17] Hao Wu, Kun Yue, Ching-Hsien Hsu, Yiji Zhao, Binbin Zhang, and Guoying Zhang. Deviation-based neighborhood model for context-aware QoS prediction of cloud and IoT services. Future Generation Computer Systems, 76(??):550–560, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16304034.

Wang:2019:PAO

[WYJ^{+}19] Xiangming Wang, Dong Yang, Xin Jian, Min Chen, and Joze Guna. Performance analysis and optimization for coverage enhancement strategy of narrow-band Internet of Things. Future Generation Computer Systems, 101(??):434–443, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19307411.

Wu:2018:CQP

[WYL^{+}18] Hao Wu, Kun Yue, Bo Li, Binbin Zhang, and Ching-Hsien Hsu. Collaborative QoS prediction with context-sensitive matrix factorization. Future Generation Computer Systems, 82(??):669–678, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17304570.

Wu:2013:TQA

[WZ13] Quanwang Wu and Qingsheng Zhu. Transactional and QoS-aware dynamic service composition based on ant colony optimization. Future Generation Computer Systems, 29(5): 1112–1119, July 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12002300.

Wu:2016:FAR

[WZ16] Xi Wu and Huibiao Zhu.

Formalization and analysis of the REST architecture from the process algebra perspective. Future Generation Computer Systems, 56(??):153–168, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002861.

Wei:2018:RSI

[WZ18] Pengcheng Wei and Zhen Zhou. Research on security of information sharing in Internet of Things based on key algorithm. Future Generation Computer Systems, 88(??):599– 605, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18303054.

Wu:2017:ESI

[WZCH17] Libing Wu, Yubo Zhang, Kim-Kwang Raymond Choo, and Debiao He. Efficient and secure identity-based encryption scheme with equality test in cloud computing. Future Generation Computer Systems, 73(??): 22–31, August 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16305313.

Zhong:2019:SRN

[wZcZN^{+}19] Xiao wu Zhong, Ming cai Zhao, Guo Ning, Liu Meirong, and Guo Zixuan. Synthesis and research of a novel near infrared nanoprobe targeting liver cancer. Future Generation Computer Systems, 98(??):536–540, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326438.

See retraction notice [wZcZN^{+}20].

Zhong:2020:RNS

[wZcZN^{+}20] Xiao wu Zhong, Ming cai Zhao, Guo Ning, Liu Meirong, and Guo Zixuan. Retraction notice to “Synthesis and research of a novel near infrared nanoprobe targeting liver cancer” [Future Gener. Comput. Syst. 98 (2019) 536–540]. Future Generation Computer Systems, 107(??):1147, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306506.

See [wZcZN^{+}19].

Wang:2019:EMR

[WZE19] Tao Wang, Zhigao Zheng, and Mohamed Elhoseny. Equivalent mechanism: Releasing location data with errors through differential privacy. Future Generation Computer Systems, 98(??):600–608, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18323513.

Wang:2019:UTP

[WZF^{+}19] Linbo Wang, Hui Zhen, Xianyong Fang, Shaohua Wan, Weiping Ding, and Yanwen Guo. A unified two-parallel-branch deep neural network for joint gland contour and segmentation learning. Future Generation Computer Systems, 100(??):316–324, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19306533.

Wang:2018:FBS

[WZH^{+}18] Tian Wang, Jiyuan Zhou, Minzhe Huang, MD Zakirul Alam Bhuiyan, Anfeng Liu, Wenzheng Xu, and Mande Xie. Fog-based storage technology to fight with cyber threat. Future Generation Computer Systems, 83(??):208–218, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17323166.

Wang:2019:PBB

[WZH^{+}19] Yurong Wang, Guijuan Zhang, Xiaoqian Hao, Yi Ma, Min Ma, Xianxin Yan, Xuefeng Jiang, Fengjie Bie, and Naijun Yuan. Potential biomarker for breast cancer screening: A systematic review and meta-analysis. Future Generation Computer Systems, 91(??):518–526, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18318351.

Wei:2018:CEA

[WZL18] Xiao Wei, Daniel Dajun Zeng, and Xiangfeng Luo. Concept evolution analysis based on the dissipative structure of concept semantic space. Future Generation Computer Systems, 81(??):384–394, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17308476.

Wang:2018:UMA

[WZM^{+}18] Zi Wang, Zhiwei Zhao, Geyong Min, Xinyuan Huang, Qiang Ni, and Rong Wang. User mobility aware task assignment for mobile edge computing. Future Generation Computer Systems, 85(??):1– 8, August 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17318587.

Wu:2018:PPB

[WZML18] Suzhen Wu, Weidong Zhu, Bo Mao, and Kuan-Ching Li. PP: Popularitybased proactive data recovery for HDFS RAID systems. Future Generation Computer Systems, 86(??):1146–1153, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17305204.

Wu:2018:DWA

[WZS^{+}18] Lin Wu, Qingfeng Zhuge, Edwin Hsing-Mean Sha, Xianzhang Chen, and Linfeng Cheng. DWARM: a wear-aware memory management scheme for inmemory file systems. Future Generation Computer Systems, 88(??):1– 15, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17322410.

Wan:2019:MDD

[WZW^{+}19a] Shaohua Wan, Yu Zhao, Tian Wang, Zonghua Gu, Qammer H. Abbasi, Kim-Kwang, and Kim-Kwang Raymond Choo. Multi-dimensional data indexing and range query processing via Voronoi diagram for Internet of Things. Future Generation Computer Systems, 91(??):382–391, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18311956.

Wang:2019:IGP

[WZW19b] Shaohua Wang, Yang Zhong, and Erqi Wang. An integrated GIS platform architecture for spatiotemporal big data. Future Generation Computer Systems, 94(??):160–172, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17319283.

Wang:2018:TLP

[WZWC18] Zhuowei Wang, Wuqing Zhao, Hao Wang, and Lianglun Cheng. Threelevel performance optimization for heterogeneous systems based on software prefetching under power constraints. Future Generation Computer Systems, 86(??):51–58, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17325943.

Wu:2018:SOO

[WZWW18] Dapeng Wu, Feng Zhang, Honggang Wang, and Ruyan Wang. Securityoriented opportunistic data forwarding in mobile social networks. Future Generation Computer Systems, 87(??):803–815, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301206.

Wang:2019:RSP

[WZY^{+}19] Weijie Wang, Gaopeng Zhang, Luming Yang, V. S. Balaji, V. Elamaran, and N. Arunkumar. Revisiting signal processing with spectrogram analysis on EEG, ECG and speech signals. Future Generation Computer Systems, 98(??):227–232, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18326220.

Wang:2016:RAI

[WZZ16] Mingzhong Wang, Liehuang Zhu, and Zijian Zhang.

Risk-aware intermediate dataset backup strategy in cloud-based data intensive workflows. Future Generation Computer Systems, 55(??):524–533, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001575.

Xhafa:2010:CMH

[XA10] Fatos Xhafa and Ajith Abraham. Computational models and heuristic methods for Grid scheduling problems. Future Generation Computer Systems, 26(4):608–621, April 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Xue:2010:WTM

[XAW^{+}10] Yong Xue, Jianwen Ai, Wei Wan, Yingjie Li, Ying Wang, Jie Guang, Linlu Mei, Hui Xu, Qiang Li, and Linyan Bai. Workload and task management of Grid-enabled quantitative aerosol retrieval from remotely sensed data. Future Generation Computer Systems, 26(4):590–598, April 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Xhafa:2014:SIP

[XB14] Fatos Xhafa and Leonard Barolli. Semantics, intelligent processing and services for big data. Future Generation Computer Systems, 37(??):201–202, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000296.

Xing:2010:ABS

[XCGD10] Wei Xing, Oscar Corcho, Carole Goble, and Marios D. Dikaiakos. An ActOn-based semantic information service for Grids. Future Generation Computer Systems, 26(3):324– 336, March 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Xiao:2018:CCP

[XCS^{+}18] Xi Xiao, Chunhui Chen, Arun Kumar Sangaiah, Guangwu Hu, Runguo Ye, and Yong Jiang. CenLocShare: A centralized privacy-preserving location-sharing system for mobile online social networks. Future Generation Computer Systems, 86(??):863–872, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17301723.

Xu:2019:WAL

[XCZ^{+}19] Jiajie Xu, Jing Chen, Rui Zhou, Junhua Fang, and Chengfei Liu. On workflow aware location-based service composition for personal trip planning. Future Generation Computer Systems, 98(??):274–285, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326694.

Xu:2017:BAE

[XDH^{+}17] Guan Xu, Bin Dai, Benxiong Huang, Jun Yang, and Sheng Wen. Bandwidthaware energy efficient flow scheduling with SDN in data center networks. Future Generation Computer Systems, 68(??):163–174, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303028.

Xue:2012:PEO

[XDHL12] Zhenghua Xue, Xiaoshe Dong, Leijun Hu, and Jianhui Li. A performance and energy optimization mechanism for cooperationoriented multiple server clusters. Future Generation Computer Systems, 28(5): 801–810, May 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11000720.

Xiang:2015:CCS

[XDWL15] Yang Xiang, Beniamino Di Martino, Guilin Wang, and Jin Li. Cloud computing: Security, privacy and practice. Future Generation Computer Systems, 52(??):59–60, November 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15002137.

Xia:2019:ACD

[XFJ^{+}19] Wen Xia, Dan Feng, Hong Jiang, Yucheng Zhang, Victor Chang, and Xiangyu Zou. Accelerating content-defined-chunking based data deduplication by exploiting parallelism. Future Generation Computer Systems, 98(??):406– 418, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18320053.

Xie:2016:OAS

[XFLL16] Yulai Xie, Dan Feng, Yan Li, and Darrell D. E. Long. Oasis: an active storage framework for object storage platform. Future Generation Computer Systems, 56(??):746–758, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002642.

Xu:2016:AEM

[XFM16] Chao Xu, Zhiyong Feng, and Zhaopeng Meng. Affective experience modeling based on interactive synergetic dependence in big data. Future Generation Computer Systems, 54(??):507–517, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000527.

Xie:2016:UID

[XFTZ16] Yulai Xie, Dan Feng, Zhipeng Tan, and Junzhe Zhou. Unifying intrusion detection and forensic analysis via provenance awareness. Future Generation Computer Systems, 61(??): 26–36, August 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300188.

Xu:2019:DTL

[XHL^{+}19] Xiaolong Xu, Nan Hu, Tao Li, Marcello Trovati, Francesco Palmieri, Georgios Kontonatsios, and Aniello Castiglione. Distributed temporal link prediction algorithm based on label propagation. Future Generation Computer Systems, 93(??):627–636, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17322653.

Xiang:2019:MLB

[XHW19] Wei Xiang, Tao Huang, and Wanggen Wan. Machine learning based optimization for vehicle-to-infrastructure communications. Future Generation Computer Systems, 94(??): 488–495, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18318326.

Xu:2015:RSV

[XJWW15] Xiaolin Xu, Hai Jin, Song Wu, and Yihong Wang. Rethink the storage of virtual machine images in clouds. Future Generation Computer Systems, 50(??):75– 86, September 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001885.

Xu:2018:VDP

[XJY^{+}18] Haiyong Xu, Gangyi Jiang, Mei Yu, Ting Luo, Zongju Peng, Feng Shao, and Hao Jiang. 3D visual discomfort predictor based on subjective perceivedconstraint sparse representation in 3D display system. Future Generation Computer Systems, 83(??): 85–94, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17303655.

Xu:2019:SDD

[XJZ^{+}19] Xianghua Xu, Jiancheng Jin, Shanqing Zhang, Lingjun Zhang, Shiliang Pu, and Zongmao Chen. Smart data driven traffic sign detection method based on adaptive color threshold and shape symmetry. Future Generation Computer Systems, 94(??): 381–391, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18311804.

Xie:2018:FPP

[XKBA18] Rongjun Xie, Ibrahim Khalil, Shahriar Badsha, and Mohammed Atiquzzaman. Fast and peer-to-peer vital signal learning system for cloud-based healthcare.

Future Generation Computer Systems, 88(??):220– 233, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17328169.

Xie:2013:P

[XKC13] Bin Xie, Anup Kumar, and Chi-Ming Chen. Preface. Future Generation Computer Systems, 29(7):1680– 1681, September 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1300085X.

Xin:2018:DII

[XKJ^{+}18] Jin Xin, Zhao Kaixuan, Ji Jiangtao, Du Xinwu, Ma Hao, and Qiu Zhaomei. Design and implementation of intelligent transplanting system based on photoelectric sensor and PLC. Future Generation Computer Systems, 88(??):127–139, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18306344.

Xiong:2019:FMA

[XL19] Juan Xiong and Zhixiang Liu. Fuzzy meta association rules based on hierarchy theory based analysis of epidemic incidence of hand, foot and mouth disease in children. Future Generation Computer Systems, 91(??):574–578, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18318302.

Xu:2014:APA

[XLL^{+}14] Yuemei Xu, Yang Li, Tao Lin, Zihou Wang, Guoqiang Zhang, Hui Tang, and Song Ci. An adaptive per-application storage management scheme based on manifold learning in information centric networks. Future Generation Computer Systems, 36(??): 170–179, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001957.

Xiong:2018:ASR

[XLL^{+}18a] Wei Xiong, Zhihui Lu, Bing Li, Bo Hang, and Zhao Wu. Automating smart recommendation from natural language API descriptions via representation learning. Future Generation Computer Systems, 87(??):382–391, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18306678.

Xu:2018:PPE

[XLL18b] Chen Xu, Peilong Li, and Yan Luo. A programmable policy engine to facilitate time-efficient science DMZ management. Future Generation Computer Systems, 89(??):515–524, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18302279.

Xu:2019:COM

[XLL^{+}19a] Xiaolong Xu, Qingxiang Liu, Yun Luo, Kai Peng, Xuyun Zhang, Shunmei Meng, and Lianyong Qi. A computation offloading method over big data for IoT-enabled cloudedge computing. Future Generation Computer Systems, 95(??):522–533, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18319770.

Xu:2019:DBB

[XLL^{+}19b] Xiwei Xu, Qinghua Lu, Yue Liu, Liming Zhu, Haonan Yao, and Athanasios V. Vasilakos. Designing blockchain-based applications: a case study for imported product traceability. Future Generation Computer Systems, 92(??):399–406, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18314298.

Xu:2019:PLB

[XLL^{+}19c] Zheng Xu, Xiangfeng Luo, Yunhuai Liu, Lin Mei, and Chuanping Hu. Power law based foundation for the measurement of discrimination information for human knowledge representation. Future Generation Computer Systems, 96(??): 617–627, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16304447.

Xu:2017:APS

[XLW^{+}17] Xiaofei Xu, Zhizhong Liu, Zhongjie Wang, Quan Z. Sheng, Jian Yu, and Xianzhi Wang. S-ABC: a paradigm of service domain-oriented artificial bee colony algorithms for service selection and composition. Future Generation Computer Systems, 68(??):304–319, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16303053.

Xu:2014:MTE

[XLZ^{+}14] Zheng Xu, Xiangfeng Luo, Shunxiang Zhang, Xiao Wei, Lin Mei, and Chuanping Hu. Mining temporal explicit and implicit semantic relations between entities using Web search engines. Future Generation Computer Systems, 37(??): 468–477, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002069.

Xu:2018:SSI

[XLZ18] Zheng Xu, Yunhuai Liu, and Hui Zhang. Special section on intelligent sensing and applications for cyber-physical systems. Future Generation Computer Systems, 81(??):382– 383, April 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17328625.

Xu:2019:MMR

[XPL19] Yi Xu, Lei Peng, and Guang-Yao Li. Multi modal registration of structural features and mutual information of medical image. Future Generation Computer Systems, 93(??): 499–505, April 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0167739X18318144.

Xu:2018:PSB

[XRPT18] Xiaolong Xu, Hanzhong Rong, Ella Pereira, and Marcello Trovati. Predatory Search-based Chaos Turbo Particle Swarm Optimisation (PS-CTPSO): a new particle swarm optimisation algorithm for Web service combination problems. Future Generation Computer Systems, 89(??):375–386, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17330340.

Xi:2015:SSC

[XSMS15] Ning Xi, Cong Sun, Jianfeng Ma, and Yulong Shen. Secure service composition with information flow control in service clouds. Future Generation Computer Systems, 49(??):142–148, August 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002696.

Xu:2019:FEI

[XTF^{+}19] Gaoxiang Xu, Zhipeng Tan, Dan Feng, Laurence T. Yang, Wei Zhou, Xinyan Zhang, Yang Zhang, and Jie Xu. FvRS: Efficiently identifying performancecritical data for improving performance of big data processing. Future Generation Computer Systems, 91(??):157–166, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18306381.

Xi:2019:PWD

[XTL^{+}19] Teng Xi, Ye Tian, Xiong Li, Hui Gao, and Wendong Wang. Pixel-wise depth based intelligent station for inferring finegrained PM 2.5. Future Generation Computer Systems, 92(??):84–92, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18307209.

Xu:2018:QGR

[XTT18] Xiaoyong Xu, Maolin Tang, and Yu-Chu Tian. QoS-guaranteed resource provisioning for cloudbased MapReduce in dynamical environments. Future Generation Computer Systems, 78 (part 1)(??): 18–30, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17303163.

Xu:2019:DAB

[XTZ^{+}19] Qian Xu, Chengxiang Tan, Wenye Zhu, Ya Xiao, Zhijie Fan, and Fujia Cheng. Decentralized attribute-based conjunctive keyword search scheme with online/offline encryption and outsource decryption for cloud computing. Future Generation Computer Systems, 97(??): 306–326, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18329613.

Xie:2016:DDT

[XWJ^{+}16] Xia Xie, Hongguang Wang, Hai Jin, Feng Zhao, Xijiang Ke, and Laurence Tianruo Yang. DTA: Dynamic topology algorithms in content-based Publish/ Subscribe. Future Generation Computer Systems, 54(??):159–167, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000278.

Xue:2019:EDS

[XWjZyF19] Xiao Xue, Shufang Wang, Le jun Zhang, and Zhi yong Feng. Evaluating of dynamic service matching strategy for social manufacturing in cloud environment. Future Generation Computer Systems, 91(??):311–326, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17328881.

Xu:2015:KSL

[XWL^{+}15] Zheng Xu, Xiao Wei, Xiangfeng Luo, Yunhuai Liu, Lin Mei, Chuanping Hu, and Lan Chen. Knowle: a semantic link network based system for organizing large scale online news events. Future Generation Computer Systems, 43–44(??):40–50, February 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000636.

Xu:2018:SNT

[XWL^{+}18] Rongbin Xu, Yeguo Wang, Haoyu Luo, Futian Wang, Ying Xie, Xiao Liu, and Yun Yang. A sufficient and necessary temporal violation handling point selection strategy in cloud workflow. Future Generation Computer Systems, 86(??):464–479, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17330637.

Xu:2018:IEE

[XWM18] Fangliang Xu, Yijie Wang, and Xingkong Ma. Incremental encoding for erasure-coded cross-datacenters cloud storage. Future Generation Computer Systems, 87(??):527–537, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17314450.

Xu:2019:ACP

[XWRZ19] Yang Xu, Guojun Wang, Ju Ren, and Yaoxue Zhang. An adaptive and configurable protection framework against Android privilege escalation threats. Future Generation Computer Systems, 92(??):210– 224, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18307775.

Xiao:2019:FSM

[XWW19] Ruowei Xiao, Zhanwei Wu, and Dongyu Wang. A finite-state-machine model driven service composition architecture for Internet of Things rapid prototyping. Future Generation Computer Systems, 99(??):473– 488, October 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18322052.

Xu:2017:EIR

[XWX^{+}17] Jiwei Xu, Tao Wang, Xiaozhao Xing, Wenbo Zhang, and Hua Zhong. Efficient image restoration of virtual machines with reference count based rewriting and caching. Future Generation Computer Systems, 77(??):87–96, December 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301218.

Xiang:2019:UUR

[XWZ^{+}19] Qiao Xiang, X. Tony Wang, J. Jensen Zhang, Harvey Newman, Y. Richard Yang, and Y. Jace Liu. Unicorn: Unified resource orchestration for multidomain, geo-distributed data analytics. Future Generation Computer Systems, 93(??):188–197, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18302413.

Xiao:2014:AAM

[XX14] Zhifeng Xiao and Yang Xiao. Achieving Accountable MapReduce in cloud computing. Future Generation Computer Systems, 30(??):1–13, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001465.

Xue:2019:DDS

[XXB19] Jingting Xue, Chunxiang Xu, and Lanhua Bai. DStore: a distributed system for outsourced data storage and retrieval. Future Generation Computer Systems, 99(??):106–114, October 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18324610.

Xu:2019:ECE

[XXQ^{+}19] Xiaolong Xu, Yuan Xue, Lianyong Qi, Yuan Yuan, Xuyun Zhang, Tariq Umer, and Shaohua Wan. An edge computing-enabled computation offloading method with privacy preservation for Internet of Connected Vehicles. Future Generation Computer Systems, 96(??):89–100, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18318703.

Xu:2019:IDR

[XXX^{+}19] Fang Xu, Qiong Xu, Zenggang Xiong, Nan Xiao, Yong Xie, Min Deng, and Huibing Hao. Intelligent distributed routing scheme based on social similarity for mobile social networks. Future Generation Computer Systems, 96(??): 472–480, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301256.

Xu:2015:IBH

[XY15] Heyang Xu and Bo Yang. An incentive-based heuristic job scheduling algorithm for utility grids. Future Generation Computer Systems, 49(??):1–7, August 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000370.

Xie:2018:RAO

[XYLZ18] Junping Xie, Minhua Yang, Jinhai Li, and Zhong Zheng. Rule acquisition and optimal scale selection in multi-scale formal decision contexts and their applications to smart city. Future Generation Computer Systems, 83(??):564– 581, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17303722.

Xu:2019:SIC

[XYML19] Shengmin Xu, Guomin Yang, Yi Mu, and Ximeng Liu. A secure IoT cloud storage system with fine-grained access control and decryption key exposure resistance. Future Generation Computer Systems, 97(??):284–294, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18320922.

Xiang:2011:SST

[XZ11] Yang Xiang and Wanlei Zhou. Special section: Trusted computing. Future Generation Computer Systems, 27(5):527–528, May 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Xu:2014:AFN

[XZ14a] Bei Xu and Hai Zhuge. Automatic faceted navigation. Future Generation Computer Systems, 32(??): 187–197, March 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12002233.

Xu:2014:MBP

[XZ14b] Li Xu and Yuexin Zhang. Matrix-based pairwise key establishment for wireless mesh networks. Future Generation Computer Systems, 30(??):140–145, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001453.

Xu:2016:ABI

[XZ16] Bei Xu and Hai Zhuge. An angle-based interest model for text recommendation. Future Generation Computer Systems, 64(??):211– 226, November 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300887.

Xiao:2019:EEC

[XZL^{+}19] Chunhua Xiao, Lei Zhang, Weichen Liu, Neil Bergmann, and Yuhua Xie. Energyefficient crypto acceleration with HW/SW codesign for HTTPS. Future Generation Computer Systems, 96(??):336–347, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831197X.

Xiong:2019:PPH

[XZP^{+}19] Hu Xiong, Yanan Zhao, Li Peng, Hao Zhang, and Kuo-Hui Yeh. Partially policy-hidden attributebased broadcast encryption with secure delegation in edge computing. Future Generation Computer Systems, 97(??):453–461, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19300172.

Xie:2019:NDN

[XZW^{+}19] Ying Xie, Yuanwei Zhu, Yeguo Wang, Yongliang Cheng, Rongbin Xu, Abubakar Sadiq Sani, Dong Yuan, and Yun Yang. A novel directional and non-local-convergent particle swarm optimization based workflow scheduling in cloudedge environment. Future Generation Computer Systems, 97(??):361–378, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18321332.

Xia:2014:BBI

[XZZ^{+}14] Feng Xia, Xuhai Zhao, Jianhui Zhang, Jianhua Ma, and Xiangjie Kong. BeeCup: a bio-inspired energy-efficient clustering protocol for mobile learning. Future Generation Computer Systems, 37(??): 449–460, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1300294X.

Xu:2018:SBT

[XZZ^{+}18] Jian Xu, Hang Zhang, Wubai Zhou, Rouying He, and Tao Li. Signature based trouble ticket classification. Future Generation Computer Systems, 78 (part 1)(??):41–58, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16308056.

Xiong:2019:NSC

[XZZ^{+}19] Huiyuan Xiong, Minghui Zhang, Ronghui Zhang, Xionglai Zhu, Lu Yang, Xuemei Guo, and Bowen Cai. A new synchronous control method for dual motor electric vehicle based on cognitive-inspired and intelligent interaction. Future Generation Computer Systems, 94(??):536– 548, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18315279.

Yaqoob:2019:FCR

[YAA^{+}19] Tahreem Yaqoob, Azka Arshad, Haider Abbas, Muhammad Faisal Amjad, and Narmeen Shafqat. Framework for calculating Return on Security Investment (ROSI) for securityoriented organizations. Future Generation Computer Systems, 95(??):754–763, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18312081.

Younas:2018:END

[YAGG18] Muhammad Younas, Irfan Awan, George Ghinea, and Tor-Morten Grønli. Editorial: New developments in cloud and IoT. Future Generation Computer Systems, 86(??):723–725, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1831121X.

Yazdanpanah:2015:PHR

[YÁJG^{+}15] Fahimeh Yazdanpanah, Carlos Álvarez, Daniel Jiménez-González, Rosa M. Badia, and Mateo Valero. Picos: a hardware runtime architecture support for OmpSs. Future Generation Computer Systems, 53(??):130–139, December 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002702.

Yahyaoui:2014:NNF

[YAO14] Hamdi Yahyaoui, Mohamed Almulla, and Hala S. Own. A novel nonfunctional matchmaking approach between fuzzy user queries and real world Web services based on rough sets. Future Generation Computer Systems, 35 (??):27–38, June 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002975.

Younas:2016:ITC

[YAP16] Muhammad Younas, Irfan Awan, and Antonio Pescape. Internet of Things and cloud services. Future Generation Computer Systems, 56(??):605–606, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003726.

Yaseen:2018:CBS

[YARH18] Muhammad Usman Yaseen, Ashiq Anjum, Omer Rana, and Richard Hill. Cloudbased scalable object detection and classification in video streams. Future Generation Computer Systems, 80(??):286–298, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301929.

Yu:2018:TLR

[YAX^{+}18] Zuoxia Yu, Man Ho Au, Qiuliang Xu, Rupeng Yang, and Jinguang Han. Towards leakage-resilient fine-grained access control in fog computing. Future Generation Computer Systems, 78 (part 2)(??):763– 777, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301310.

Ye:2013:NCG

[YC13] Deshi Ye and Jianhai Chen. Non-cooperative games on multidimensional resource allocation. Future Generation Computer Systems, 29(6):1345–1352, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000320.

Yang:2019:III

[YCD^{+}19] Chao-Tung Yang, Shuo-Tsung Chen, Walter Den, Yun-Ting Wang, and Endah Kristiani. Implementation of an intelligent indoor environmental monitoring and management system in cloud. Future Generation Computer Systems, 96(??): 731–749, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18304187.

Yin:2019:ADM

[YCH19] Zhida Yin, Haopeng Chen, and Fei Hu. An advanced decision model enabling two-way initiative offloading in edge computing. Future Generation Computer Systems, 90(??):39– 48, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17329527.

Yang:2019:IRT

[YCL^{+}19] Chao-Tung Yang, Shuo-Tsung Chen, Jung-Chun Liu, Yao-Yu Yang, Karan Mitra, and Rajiv Ranjan. Implementation of a realtime network traffic monitoring service with network functions virtualization. Future Generation Computer Systems, 93(??):687– 701, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1830311X.

Yao:2015:LAB

[YCT15] Xuanxia Yao, Zhi Chen, and Ye Tian. A lightweight attribute-based encryption scheme for the Internet of Things. Future Generation Computer Systems, 49(??): 104–112, August 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002039.

Yu:2018:DDM

[YCX18] Fangxiaoqi Yu, Haopeng Chen, and Jinqing Xu. DMPO: Dynamic mobilityaware partial offloading in mobile edge computing. Future Generation Computer Systems, 89(??):722– 735, December 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17329813.

Ye:2018:FPP

[YCXW18] Ayong Ye, Qiuling Chen, Li Xu, and Wei Wu. The flexible and privacypreserving proximity detection in mobile social network. Future Generation Computer Systems, 79 (part 1)(??):271– 283, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16307518.

Yang:2010:IMI

[YCY10] Chao-Tung Yang, Chiu-Hsiung Chen, and Ming-Feng Yang. Implementation of a medical image file accessing system in coallocation data grids. Future Generation Computer Systems, 26(8):1127–1140, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yu:2018:DSG

[YCZJ18] Han Yu, Hongming Cai, Jun Zhou, and Lihong Jiang. Data service generation framework from heterogeneous printed forms using semantic link discovery. Future Generation Computer Systems, 79 (part 2)(??):514–527, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17305034.

Ye:2018:CKS

[YD18] Jun Ye and Yong Ding.

Controllable keyword search scheme supporting multiple users. Future Generation Computer Systems, 81 (??):433–442, April 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X17308063.

Yi:2018:INI

[YDD^{+}18] Jiao-Hong Yi, Suash Deb, Junyu Dong, Amir H. Alavi, and Gai-Ge Wang. An improved NSGA–III algorithm with adaptive mutation operator for big data optimization problems. Future Generation Computer Systems, 88(??):571–585, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18306137.

Yamaoka:2011:VHR

[YDK11] So Yamaoka, Kai-Uwe Doerr, and Falko Kuester. Visualization of highresolution image collections on large tiled display walls. Future Generation Computer Systems, 27(5):498–505, May 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yan:2016:TSP

[YDNV16] Zheng Yan, Wenxiu Ding, Valtteri Niemi, and Athanasios V. Vasilakos. Two schemes of privacy-preserving trust evaluation. Future Generation Computer Systems, 62(??):175–189, September 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1500343X.

Yue:2019:NSB

[YDQC19] Wenying Yue, Peng Du, Xiongwen Quan, and Ying Chen. A new strategy based on approximate dynamic programming to maximize the net revenue of IaaS cloud providers with limited resources. Future Generation Computer Systems, 101(??):1168– 1186, December 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831906X.

Yount:2019:MLS

[YDT19] Charles Yount, Alejandro Duran, and Josh Tobin. Multi-level spatial and temporal tiling for efficient HPC stencil computation on many-core processors with large shared caches. Future Generation Computer Systems, 92(??):903– 919, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17304648.

Yang:2013:AHA

[YFY^{+}13] Donghua Yang, Yuqiang Feng, Ye Yuan, Xixian Han, Jinbao Wang, and Jianzhong Li. Ad-hoc aggregate query processing algorithms based on bitstore for query intensive applications in cloud computing. Future Generation Computer Systems, 29(7):1725–1735, September 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000623.

Yadav:2018:TWR

[YG18] Neeraj Yadav and Major Singh Goraya. Twoway ranking based service mapping in cloud environment. Future Generation Computer Systems, 81(??):

53–66, April 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17310804.

Yasar:2016:SIE

[YGS16] Ansar-Ul-Haque Yasar, Stephane Galland, and Elhadi Shakshuki. Special issue: Emerging, ambient and ubiquitous systems 2015. Future Generation Computer Systems, 64(??):75–77, November 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302102.

Yu:2019:LEW

[YGY^{+}19] Ruiguo Yu, Jie Gao, Mei Yu, Wenhuan Lu, Tianyi Xu, Mankun Zhao, Jie Zhang, Ruixuan Zhang, and Zhuo Zhang. LSTM-EFG for wind power forecasting based on sequential correlation features. Future Generation Computer Systems, 93(??):33– 42, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18314420.

Yu:2016:RIS

[YGYW16] Xiaoshan Yu, Huaxi Gu, Yintang Yang, and Kun Wang. RingCube — an incrementally scale-out optical interconnect for cloud computing data center. Future Generation Computer Systems, 54(??):41– 51, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002186.

Yang:2018:CTF

[YH18] Shao-Jun Yang and Xinyi Huang. Certain types of M-fuzzifying matroids: A fundamental look at the security protocols in RFID and IoT. Future Generation Computer Systems, 86(??):582–590, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1732294X.

Yeh:2019:RIP

[YH19] Tsozen Yeh and Hsinyi Huang. Realizing integrated prioritized service in the Hadoop cloud system. Future Generation Computer Systems, 100(??):176–185, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18325779.

Yaqoob:2019:ITF

[YHA^{+}19] Ibrar Yaqoob, Ibrahim Abaker Targio Hashem, Arif Ahmed, S. M. Ahsan Kazmi, and Choong Seon Hong. Internet of Things forensics: Recent advances, taxonomy, requirements, and open challenges. Future Generation Computer Systems, 92(??):265–275, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18315644.

Yuan:2019:UBP

[YHH^{+}19] Weiwei Yuan, Kangya He, Guangjie Han, Donghai Guan, and Asad Masood Khattak. User behavior prediction via heterogeneous information preserving network embedding. Future Generation Computer Systems, 92(??):52– 58, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1831567X.

Yang:2016:EHA

[YHL16] Xu Yang, Xinyi Huang, and Joseph K. Liu. Efficient handover authentication with user anonymity and untraceability for mobile cloud computing. Future Generation Computer Systems, 62(??):190–195, September 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003088.

Yang:2019:IIG

[YHL^{+}19] Xuan Yang, Guangjie Han, Li Liu, Aihua Qian, and Wenbo Zhang. IGRC: An improved grid-based joint routing and charging algorithm for wireless rechargeable sensor networks. Future Generation Computer Systems, 92(??):837–845, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312232.

Yildiz:2017:EFF

[YIA17] Orcun Yildiz, Shadi Ibrahim, and Gabriel Antoniu. Enabling fast failure recovery in shared Hadoop clusters: Towards failure-aware scheduling. Future Generation Computer Systems, 74(??):208–219, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300280.

Yen:2014:SIH

[YJHZ14] Neil Y. Yen, Qin Jin, Ching-Hsien Hsu, and Qiangfu Zhao. Special issue on “Hybrid intelligence for growing Internet and its applications”. Future Generation Computer Systems, 37(??):401– 403, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14000624.

Yu:2019:MDD

[YJL^{+}19] Min Yu, Jianguo Jiang, Gang Li, Chenzhe Lou, Yunzheng Liu, Chao Liu, and Weiqing Huang. Malicious documents detection for business process management based on multilayer abstract model. Future Generation Computer Systems, 99(??):517–526, October 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18326700.

Yang:2018:DIR

[YJS18] Jiachen Yang, Bin Jiang, and Houbing Song. A distributed image-retrieval method in multi-camera system of smart city based on cloud computing. Future Generation Computer Systems, 81(??):244–251, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17321362.

Yang:2018:SRB

[YJY^{+}18] Xiaomin Yang, Lihua Jian, Binyu Yan, Kai Liu, Lei Zhang, and Yiguang Liu. A sparse representation based pansharpening method. Future Generation Computer Systems, 88(??):385–399, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1732424X.

Yun:2017:MHA

[YK17] Unil Yun and Donggyu Kim. Mining of high average-utility itemsets using novel list structure and pruning strategy. Future Generation Computer Systems, 68(??):346–360, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16304733.

Yildirim:2013:MTS

[YKK13] Esma Yildirim, Jangyoung Kim, and Tevfik Kosar. Modeling throughput sampling size for a cloudhosted data scheduling and optimization service. Future Generation Computer Systems, 29(7):1795–1807, September 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13000174.

Yuksel:2017:RIP

[YKÖ17] Buket Yüksel, Alptekin Küpçü, and Öznur Özkasap. Research issues for privacy and security of electronic health services. Future Generation Computer Systems, 68(??):1–13, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302667.

Yun:2016:SWB

[YL16] Unil Yun and Gangin Lee. Sliding window based weighted erasable stream pattern mining for stream data applications. Future Generation Computer Systems, 59(??):1– 20, June 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003969.

Yu:2018:RSB

[YL18] Wei Yu and Shijun Li. Recommender systems based on multiple social networks correlation. Future Generation Computer Systems, 87(??):312–327, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1830164X.

Yuan:2018:ESD

[YLA18] Bo Yuan, Lu Liu, and Nick Antonopoulos. Efficient service discovery in decentralized online social networks. Future Generation Computer Systems, 86(??):775–791, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17306465.

Ye:2016:RMU

[YLG^{+}16] Wei Ye, Ruixuan Li, Xiwu Gu, Yuhua Li, and Kunmei Wen. Role mining using answer set programming. Future Generation Computer Systems, 55(??):336– 343, February 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002118.

Yuan:2019:NSP

[YLH^{+}19] Weiwei Yuan, Chenliang Li, Guangjie Han, Donghai Guan, Li Zhou, and Kangya He. Negative sign prediction for signed social networks. Future Generation Computer Systems, 93(??):962–970, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17307768.

Yang:2014:MMG

[YLHJ14] Chao-Tung Yang, Jung-Chun Liu, Kuan-Lung Huang, and Fuu-Cheng Jiang. A method for managing green power of a virtual machine cluster in cloud. Future Generation Computer Systems, 37(??): 26–36, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000466.

Yang:2017:HSC

[YLJ^{+}17] Lei Yang, Weichen Liu, Weiwen Jiang, Chao Chen, Mengquan Li, Peng Chen, and Edwin H. M. Sha. Hardware–software collaboration for dark silicon heterogeneous many-core systems. Future Generation Computer Systems, 68(??): 234–247, March 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303090.

Yang:2018:EFA

[YLJL18] Bo Yang, Zhiyong Li, Shilong Jiang, and Keqin Li. Envy-free auction mechanism for VM pricing and allocation in clouds. Future Generation Computer Systems, 86(??):680–693, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17330418.

Yang:2019:RLS

[YLL^{+}19] Aimin Yang, Yifan Li, Chenshuai Liu, Jie Li, Yuzhu Zhang, and Jiahao Wang. Research on logistics supply chain of iron and steel enterprises based on block chain technology. Future Generation Computer Systems, 101(??):635–645, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19309124.

Yang:2015:HSP

[YLN15] Ji-Jiang Yang, Jian-Qiang Li, and Yu Niu. A hybrid solution for privacy preserving medical data sharing in the cloud environment. Future Generation Computer Systems, 43–44(??):74–86, February 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001253.

Yan:2015:TDF

[YLVY15] Zheng Yan, Jun Liu, Athanasios V. Vasilakos, and Laurence T. Yang. Trustworthy data fusion and mining in Internet of Things. Future Generation Computer Systems, 49(??): 45–46, August 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15000862.

Yang:2018:BCS

[YLWW18] Yongjian Yang, Wenbin Liu, En Wang, and Hengzhi Wang. Beaconing control strategy based on game theory in mobile crowdsensing. Future Generation Computer Systems, 86(??):222–233, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17320009.

Yuan:2013:ARM

[YMD^{+}13] Xiaoqun Yuan, Geyong Min, Yi Ding, Qiong Liu, Jinhong Liu, Hao Yin, and Qing Fang. Adaptive resource management for P2P live streaming systems. Future Generation Computer Systems, 29(6):1573–1582, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12001756.

Yahyaoui:2013:TCB

[YMLT13] Hamdi Yahyaoui, Zakaria Maamar, Erbin Lim, and Philippe Thiran. Towards a community-based, social network-driven framework for Web services management. Future Generation Computer Systems, 29(6):1363–1377, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000319.

Yang:2013:ECS

[YMW13] Erica Yang, Brian Matthews, and Michael Wilson. Enhancing the core scientific metadata model to incorporate derived data. Future Generation Computer Systems, 29(2):612– 623, February 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001427.

Yan:2018:CBR

[YMW^{+}18] Jining Yan, Yan Ma, Lizhe Wang, Kim-Kwang Raymond Choo, and Wei Jie. A cloud-based remote sensing data production system. Future Generation Computer Systems, 86(??):1154–1166, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17303035.

Yuan:2017:GTB

[YMY^{+}17] Xiaoqun Yuan, Geyong Min, Laurence T. Yang, Yi Ding, and Qing Fang. A game theory-based dynamic resource allocation strategy in Geo-distributed Datacenter Clouds. Future Generation Computer Systems, 76(??):63– 72, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16307750.

Yi:2018:SCS

[YN18] Haibo Yi and Zhe Nie. Side-channel security analysis of UOV signature for cloud-based Internet of Things. Future Generation Computer Systems, 86(??):704–708, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18304151.

Yun:2019:EAI

[YNLY19] Unil Yun, Hyoju Nam, Gangin Lee, and Eunchul Yoon. Efficient approach for incremental high utility pattern mining with indexed list structure. Future Generation Computer Systems, 95(??):221–239, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18315814.

Yang:2012:BOC

[YNSM12] Xiaoyu Yang, Bassem Nasser, Mike Surridge, and Stuart Middleton. A business-oriented Cloud federation model for realtime applications. Future Generation Computer Systems, 28(8):1158–1167, October 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000386.

Yu:2014:SAM

[YNY^{+}14] Yong Yu, Lei Niu, Guomin Yang, Yi Mu, and Willy Susilo. On the security of auditing mechanisms for secure cloud storage. Future Generation Computer Systems, 30(??):127–132, January 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001040.

Yeom:2012:MAA

[YP12] Kiwon Yeom and Ji-Hyung Park. Morphological approach for autonomous and adaptive systems based on self-reconfigurable modular agents. Future Generation Computer Systems, 28(3): 533–543, March 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11000306.

Yan:2012:AWS

[YPCK12] Lijun Yan, Jeng-Shyang Pan, Shu-Chuan Chu, and Muhammad Khurram Khan. Adaptively weighted sub-directional two-dimensional linear discriminant analysis for face recognition. Future Generation Computer Systems, 28(1):232–235, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1000230X.

Yu:2014:FAC

[YPHZ14] Xu Yu, Li Peng, Zhixing Huang, and Hai Zhuge. A framework for automated construction of resource space based on background knowledge. Future Generation Computer Systems, 32(??):222–231, March 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001623.

Ye:2019:ISA

[YPJ19] Chaochao Ye, Julong Pan, and Qun Jin. An improved SSO algorithm for cyberenabled tumor risk analysis based on gene selection. Future Generation Computer Systems, 92(??):407– 418, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18307453.

Yang:2017:EEV

[YPLZ17] Ting Yang, Haibo Pen, Wei Li, and Albert Y. Zomaya. An energy-efficient virtual machine placement and route scheduling scheme in data center networks. Future Generation Computer Systems, 77(??):1–11, December 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310579.

Yin:2019:SCM

[YQZ^{+}19] Hui Yin, Zheng Qin, Jixin Zhang, Lu Ou, Fangmin Li, and Keqin Li. Secure conjunctive multi-keyword ranked search over encrypted cloud data for multiple data owners. Future Generation Computer Systems, 100(??):689–700, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17321192.

Yoo:2016:CRD

[YS16] Hyunguk Yoo and Taeshik Shon. Challenges and research directions for heterogeneous cyber-physical system based on IEC 61850: Vulnerabilities, security requirements, and security architecture. Future Generation Computer Systems, 61(??):128–136, August 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003064.

Yang:2015:AMI

[YSC^{+}15] Chao-Tung Yang, Wen-Chung Shih, Lung-Teng Chen, Cheng-Ta Kuo, Fuu-Cheng Jiang, and Fang-Yie Leu. Accessing medical image file with coallocation HDFS in cloud. Future Generation Computer Systems, 43–44(??): 61–73, February 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001563.

Yan:2019:DPP

[YSC^{+}19] Zhiqiang Yan, Bo Sun, Yu Chen, Lijie Wen, Lei Hu, Jianmin Wang, Mingji Yang, and Lu Wang. Decomposed and parallel process discovery: a framework and application. Future Generation Computer Systems, 98(??):392–405, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18327110.

Yassine:2019:IBD

[YSHM19] Abdulsalam Yassine, Shailendra Singh, M. Shamim Hossain, and Ghulam Muhammad. IoT big data analytics for smart homes with fog and cloud computing. Future Generation Computer Systems, 91(??):563–573, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18311099.

Yousef:2019:ITL

[YSL19] Khalil M. Ahmad Yousef, Ali Shatnawi, and Mohammad Latayfeh. Intelligent traffic light scheduling technique using calendarbased history information. Future Generation Computer Systems, 91(??):124–135, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18306241.

Yao:2018:SEM

[YSZW18] Shihong Yao, Arun Kumar Sangaiah, Zhigao Zheng, and Tao Wang. Sparsity estimation matching pursuit algorithm based on restricted isometry property for signal reconstruction. Future Generation Computer Systems, 88(??):747– 754, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17303813.

Yao:2019:HRU

[YTQ19] Qingrong Yao, Wu Tian, and Li Qiu. Highresolution ultrasound images in gouty arthritis to evaluate relationship between tophi and bone erosion. Future Generation Computer Systems, 98(??):131–134, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831851X.

See retraction notices [YTQ20, YWG^{+}19].

Yao:2020:RNH

[YTQ20] Qingrong Yao, Wu Tian, and Li Qiu. Retraction notice to “Highresolution ultrasound images in gouty arthritis to evaluate relationship between tophi and bone erosion” [future gener. comput. syst. 98 (2018) 131– 134]. Future Generation Computer Systems, 106 (??):689, May 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20305689.

See [YTQ19].

Yeo:2010:AMP

[YVCB10] Chee Shin Yeo, Srikumar Venugopal, Xingchen Chu, and Rajkumar Buyya. Autonomic metered pricing for a utility computing service. Future Generation Computer Systems, 26(8): 1368–1380, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yeh:2012:EBR

[YW12] Wei-Chang Yeh and Shang-Chia Wei. Economic-based resource allocation for reliable Grid-computing service based on Grid Bank. Future Generation Computer Systems, 28(7):989– 1002, July 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000581.

Yang:2018:HBH

[YWCC18] Wenyin Yang, Guojun Wang, Kim-Kwang Raymond Choo, and Shuhong Chen. HEPart: A balanced hypergraph partitioning algorithm for big data applications. Future Generation Computer Systems, 83(??):250–268, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17316473.

Yuan:2010:VMM

[YWF^{+}10] Yulai Yuan, Yongwei Wu, Xiao Feng, Jing Li, Guangwen Yang, and Weimin Zheng. VDB– MR: MapReduce-based distributed data integration using virtual database. Future Generation Computer Systems, 26(8):1418– 1425, October 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yanli:2019:IKG

[YWG^{+}19] Ji Yanli, Lu Wei, Che Guowei, Yang Mei, and Liu Lunxu. Inhibition of KRAS gene mutation on non-small cell lung cancer and its effect on circulating tumor cells. Future Generation Computer Systems, 98(??):104–108, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326426.

See retraction notices [YWG^{+}20, JLC^{+}20].

Yanli:2020:RNI

[YWG^{+}20] Ji Yanli, Lu Wei, Che Guowei, Yang Mei, and Liu Lunxu. Retraction notice to “Inhibition of KRAS gene mutation on non-small cell lung cancer and its effect on circulating tumor cells” [Future Gener. Comput. Syst. 98 (2019) 104–108]. Future Generation Computer Systems, 107(??): 1143, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306361.

See [YWG^{+}19].

Yang:2018:SAD

[YWJ^{+}18] Kang Yang, Rui Wang, Yu Jiang, Houbing Song, Chenxia Luo, Yong Guan, Xiaojuan Li, and Zhiping Shi. Sensor attack detection using history based pairwise inconsistency. Future Generation Computer Systems, 86(??):392–402, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17322306.

Yan:2019:IFF

[YWJ^{+}19] Hongyang Yan, Yu Wang, Chunfu Jia, Jin Li, Yang Xiang, and Witold Pedrycz. IoT-FBAC: Function-based access control scheme using identity-based encryption in IoT. Future Generation Computer Systems, 95(??): 344–353, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1830997X.

Yang:2017:MRT

[YWL^{+}17] Jiachen Yang, Huanling Wang, Zhihan Lv, Wei Wei, Houbing Song, Melike Erol-Kantarci, Burak Kantarci, and Shudong He. Multimedia recommendation and transmission system based on cloud platform. Future Generation Computer Systems, 70(??): 94–103, May 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630200X.

Yao:2019:DLM

[YWLL19] Chenhui Yao, Shuodong Wu, Zhuo Liu, and Peng Li. A deep learning model for predicting chemical composition of gallstones with big data in medical Internet of Things. Future Generation Computer Systems, 94(??):140–147, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18324877.

Yu:2017:SSS

[YWY^{+}17] Chen Yu, Namin Wang, Laurence T. Yang, Dezhong Yao, Ching-Hsien Hsu, and Hai Jin. A semi-supervised social relationships inferred model based on mobile phone data. Future Generation Computer Systems, 76(??):458–467, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306665.

Yang:2018:MSM

[YWZ^{+}18] Jiachen Yang, Chang Wang, Qiming Zhao, Bin Jiang, Zhihan Lv, and Arun Kumar Sangaiah. Marine surveying and mapping system based on cloud computing and Internet of Things. Future Generation Computer Systems, 85(??):39–50, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17327048.

Yu:2016:CDI

[YXA^{+}16] Yong Yu, Liang Xue, Man Ho Au, Willy Susilo, Jianbing Ni, Yafang Zhang, Athanasios V. Vasilakos, and Jian Shen. Cloud data integrity checking with an identity-based auditing mechanism from RSA. Future Generation Computer Systems, 62(??):85– 91, September 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300164.

Yang:2018:PBC

[YXA^{+}18] Rupeng Yang, Qiuliang Xu, Man Ho Au, Zuoxia Yu, Hao Wang, and Lu Zhou. Position based cryptography with location privacy: a step for Fog Computing. Future Generation Computer Systems, 78 (part 2)(??):799– 806, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17311032.

Ye:2018:ISS

[YXD18] Jun Ye, Zheng Xu, and Yong Ding. Image search scheme over encrypted database. Future Generation Computer Systems, 87(??):251–258, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324354.

Ying:2018:CHA

[YXY18] Sun Ying and Zhang Xue-Ying. Characteristics of human auditory model based on compensation of glottal features in speech emotion recognition. Future Generation Computer Systems, 81(??):291–296, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17312979.

Yan:2018:FDA

[YXZG18a] Zheng Yan, Haomeng Xie, Peng Zhang, and Brij B. Gupta. Flexible data access control in D2D communications. Future Generation Computer Systems, 82(??):738–751, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1730835X.

Yao:2018:MMS

[YXZG18b] Hong Yao, Muzhou Xiong, Deze Zeng, and Junfang Gong. Mining multiple spatial-temporal paths from social media data. Future Generation Computer Systems, 87(??):782–791, October 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301176.

Yoon:2011:SBC

[YY11] Eun-Jun Yoon and Kee-Young Yoo. A secure broadcasting cryptosystem and its application to grid computing. Future Generation Computer Systems, 27(5):620–626, May 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yao:2014:EEI

[YYD^{+}14] Dezhong Yao, Chen Yu, Anind K. Dey, Christian Koehler, Geyong Min, Laurence T. Yang, and Hai Jin. Energy efficient indoor tracking on smartphones. Future Generation Computer Systems, 39(??): 44–54, October 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002963.

Yuan:2010:DPS

[YYLC10] Dong Yuan, Yun Yang, Xiao Liu, and Jinjun Chen. A data placement strategy in scientific cloud workflows. Future Generation Computer Systems, 26(8): 1200–1214, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yang:2019:SSA

[YYS^{+}19] Shiqi Yang, Di Yin, Xia Song, Xiaoming Dong, Gunasekaran Manogaran, George Mastorakis, Constandinos X. Mavromoustakis, and Jordi Mongay Batalla. Security situation assessment for massive MIMO systems for 5G communications. Future Generation Computer Systems, 98(??):25– 34, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19300731.

Yang:2019:ARL

[YYW^{+}19] Jun Yang, Xinghui You, Gaoxiang Wu, Mohammad Mehedi Hassan, Ahmad Almogren, and Joze Guna. Application of reinforcement learning in UAV cluster task scheduling. Future Generation Computer Systems, 95(??):140–148, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18325299.

You:2012:WFQ

[YZ12] Guohua You and Ying Zhao. A weighted-fair-queuing (WFQ)-based dynamic request scheduling approach in a multi-core system. Future Generation Computer Systems, 28(7): 1110–1120, July 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11002135.

Yan:2019:DDE

[YZC^{+}19] Hui Yan, Xiongtao Zhang, Huangke Chen, Yun Zhou, Weidong Bao, and Laurence T. Yang. DEED: Dynamic energy-efficient data offloading for IoT applications under unstable channel conditions. Future Generation Computer Systems, 96(??):425–437, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18322143.

Yang:2018:PPF

[YZG^{+}18] Yang Yang, Xianghan Zheng, Wenzhong Guo, Ximeng Liu, and Victor Chang. Privacy-preserving fusion of IoT and big data for e-health. Future Generation Computer Systems, 86(??):1437–1455, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324391.

Yildiz:2018:IEB

[YZI18] Orcun Yildiz, Amelie Chi Zhou, and Shadi Ibrahim. Improving the effectiveness of burst buffers for big data processing in HPC systems with Eley. Future Generation Computer Systems, 86(??):308–318, September 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17317302.

Yang:2018:CDD

[YZL^{+}18] Yang Yang, Xianghan Zheng, Ximeng Liu, Shangping Zhong, and Victor Chang. Cross-domain dynamic anonymous authenticated group key management with symptommatching for e-health social system. Future Generation Computer Systems, 84(??):160–176, July 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1730554X.

Yang:2019:BBL

[YZL^{+}19] Mengmeng Yang, Tianqing Zhu, Kaitai Liang, Wanlei Zhou, and Robert H. Deng. A blockchain-based location privacy-preserving crowdsensing system. Future Generation Computer Systems, 94(??):408–418, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18320909.

Yang:2014:IIV

[YZLQ14] Hailong Yang, Qi Zhao, Zhongzhi Luan, and Depei Qian. iMeter: an integrated VM power model based on performance profiling. Future Generation Computer Systems, 36(??): 267–286, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001532.

Yu:2015:RDP

[YZN^{+}15] Yong Yu, Yafang Zhang, Jianbing Ni, Man Ho Au, Lanxiang Chen, and Hongyu Liu. Remote data possession checking with enhanced security for cloud storage. Future Generation Computer Systems, 52(??):77–85, November 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14001903.

Yu:2014:GSO

[YZW14] Hongliang Yu, Fan Zhang, and Yongwei Wu. Granary: a sharing oriented distributed storage system. Future Generation Computer Systems, 38(??):47– 60, September 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001660.

Yang:2018:ERB

[YZW^{+}18] Anjia Yang, Yunhui Zhuang, Jian Weng, Gerhard Hancke, Duncan S. Wong, and Guomin Yang. Exploring relationship between indistinguishability-based and unpredictability-based RFID privacy models. Future Generation Computer Systems, 82(??):315–326, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17318939.

Yao:2018:EJC

[YZWG18] Shihong Yao, Zhigao Zheng, Tao Wang, and Qingfeng Guan. An efficient joint compression and sparsity estimation matching pursuit algorithm for artificial intelligence application. Future Generation Computer Systems, 86(??):603–613, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18302139.

Yuan:2019:GCS

[YZZC19] Zhe Yuan, Xuguang Zhang, Wenqin Zhuang, and Jianxin Chen. Green content sharing mode: D2D coordination multiple points transmission. Future Generation Computer Systems, 92(??):252–264, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18313566.

Zhang:2013:NBE

[ZA13] Kan Zhang and Nick Antonopoulos. A novel bartering exchange ring based incentive mechanism for peer-to-peer systems. Future Generation Computer Systems, 29(1):361– 369, January 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001166.

Zounmevo:2014:FRC

[ZA14] Judicael A. Zounmevo and Ahmad Afsahi. A fast and resource-conscious MPI message queue mechanism for large-scale jobs. Future Generation Computer Systems, 30(??):265– 290, January 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001489.

Zapater:2014:NED

[ZAA^{+}14] Marina Zapater, Patricia Arroba, José L. Ayala, José M. Moya, and Katzalin Olcoz. A novel energy-driven computing paradigm for e-health scenarios. Future Generation Computer Systems, 34(??): 138–154, May 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13002768.

Zarrin:2015:DSF

[ZAB15] Javad Zarrin, Rui L. Aguiar, and João Paulo Barraca. Dynamic, scalable and flexible resource discovery for largedimension many-core systems. Future Generation Computer Systems, 53(??):119–129, December 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002714.

Zhou:2018:MSV

[ZAC^{+}18] Zhou Zhou, Jemal Abawajy, Morshed Chowdhury, Zhigang Hu, Keqin Li, Hongbing Cheng, Abdulhameed A. Alelaiwi, and Fangmin Li.

Minimizing SLA violation and power consumption in cloud data centers using adaptive energy-aware algorithms. Future Generation Computer Systems, 86(??):836–850, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17316059.

Zikria:2018:SRP

[ZAI^{+}18] Yousaf Bin Zikria, Muhammad Khalil Afzal, Farruh Ishmanov, Sung Won Kim, and Heejung Yu. A survey on routing protocols supported by the Contiki Internet of Things operating system. Future Generation Computer Systems, 82(??):200–219, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17324299.

Zhu:2017:ROL

[ZaTZ^{+}17] Hongfei Zhu, Yu an Tan, Xiaosong Zhang, Liehuang Zhu, Changyou Zhang, and Jun Zheng. A roundoptimal lattice-based blind signature scheme for cloud services. Future Generation Computer Systems, 73(??):106–114, August 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17301620.

See cryptoanalysis [CJS19].

Ziafat:2019:HSO

[ZB19] Hasan Ziafat and Seyed Morteza Babamir. A hierarchical structure for optimal resource allocation in geographically distributed clouds. Future Generation Computer Systems, 90(??):539–568, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17328844.

Zhang:2017:LEA

[ZBCT17] Zhe Zhang, Brian Bockelman, Dale W. Carder, and Todd Tannenbaum. Lark: an effective approach for software-defined networking in high throughput computing clusters. Future Generation Computer Systems, 72(??):105–117, July 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300541.

Zurita:2014:UCD

[ZBF14] Gustavo Zurita, Nelson Baloian, and Jonathan Frez. Using the cloud to develop applications supporting geo-collaborative Situated Learning. Future Generation Computer Systems, 34(??):124–137, May 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002173.

Zheng:2014:ABA

[ZBL^{+}14] Jin Zheng, Md Zakirul Alam Bhuiyan, Shaohua Liang, Xiaofei Xing, and Guojun Wang. Auction-based adaptive sensor activation algorithm for target tracking in wireless sensor networks. Future Generation Computer Systems, 39(??): 88–99, October 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002781.

Zhou:2019:GPG

[ZCDV19] Jun Zhou, Zhenfu Cao, Xiaolei Dong, and Thanos Vasilakos. GTSIM-POP: Game theory based secure incentive mechanism and patient-optimized privacypreserving packet forwarding scheme in m-healthcare social networks. Future Generation Computer Systems, 101(??):70–82, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18315681.

Zhang:2017:IPP

[ZCH^{+}17] Yin Zhang, Min Chen, Dijiang Huang, Di Wu, and Yong Li. iDoctor: Personalized and professionalized medical recommendations based on hybrid matrix factorization. Future Generation Computer Systems, 66(??):30–35, January 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X15003842.

Zhang:2015:TLA

[ZCK^{+}15] Fan Zhang, Junwei Cao, Samee U. Khan, Keqin Li, and Kai Hwang. A tasklevel adaptive MapReduce framework for real-time streaming data in healthcare applications. Future Generation Computer Systems, 43–44(??):149–160, February 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14001319.

Zhang:2014:MOS

[ZCL^{+}14] Fan Zhang, Junwei Cao, Keqin Li, Samee U. Khan, and Kai Hwang. Multiobjective scheduling of many tasks in cloud platforms. Future Generation Computer Systems, 37(??): 309–320, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001854.

Zhang:2018:EAC

[ZCL^{+}18] Peng Zhang, Zehong Chen, Joseph K. Liu, Kaitai Liang, and Hongwei Liu. An efficient access control scheme with outsourcing capability and attribute update for fog computing. Future Generation Computer Systems, 78 (part 2)(??):753–762, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16307567.

Zheng:2019:PPC

[ZCL^{+}19] Xu Zheng, Zhipeng Cai, Guangchun Luo, Ling Tian, and Xiao Bai. Privacy-preserved community discovery in online social networks. Future Generation Computer Systems, 93(??):1002–1009, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17307896.

Zhang:2018:EPT

[ZCLW18] Shaobo Zhang, Kim-Kwang Raymond Choo, Qin Liu, and Guojun Wang. Enhancing privacy through uniform grid and caching in location-based services. Future Generation Computer Systems, 86(??):881–892, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17312384.

Zalila:2019:MDC

[ZCM19] Faiez Zalila, Stéphanie Challita, and Philippe Merle. Model-driven cloud resource management with OCCIware. Future Generation Computer Systems, 99(??):260–277, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18306071.

Zhao:2016:NPC

[ZCQ^{+}16] Hui Zhao, Min Chen, Meikang Qiu, Keke Gai, and Meiqin Liu. A novel pre-cache schema for high performance Android system. Future Generation Computer Systems, 56(??):

766–772, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15001806.

Zhang:2016:OVR

[ZCS^{+}16] Qizhi Zhang, Haopeng Chen, Yuxi Shen, Sixiang Ma, and Heng Lu. Optimization of virtual resource management for cloud applications to cope with traffic burst. Future Generation Computer Systems, 58(??):42–55, May 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003957.

Zhang:2011:WSW

[ZCW11] Donglai Zhang, Paul Coddington, and Andrew Wendelborn. Web services workflow with result data forwarding as resources. Future Generation Computer Systems, 27(6):694– 702, June 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zimba:2019:BNB

[ZCW19] Aaron Zimba, Hongsong Chen, and Zhaoshun Wang. Bayesian network based weighted APT attack paths modeling in cloud computing. Future Generation Computer Systems, 96(??): 525–537, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18308768.

Zhang:2018:DDB

[ZCX^{+}18] Yifei Zhang, Hongming Cai, Boyi Xu, Athanasios T. Vasilakos, and Chengxi Huang. Data driven business rule generation based on fog computing. Future Generation Computer Systems, 89(??):494–505, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18308781.

Zhang:2018:RAS

[ZCYZ18] Jun Zhang, Aniello Castiglione, Laurence Tianruno Yang, and Yan Zhang. Recent advances in security and privacy in social big data. Future Generation Computer Systems, 87(??):686–687, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18313050.

Zheng:2018:CAM

[ZCZ^{+}18] Ruijuan Zheng, Jing Chen, Mingchuan Zhang, Qingtao Wu, Junlong Zhu, and Huiqiang Wang. A collaborative analysis method of user abnormal behavior based on reputation voting in cloud environment. Future Generation Computer Systems, 83(??):60– 74, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17321258.

Zaki:2013:ADU

[ZDL^{+}13] Zulkifly Mohd Zaki, Peter M. Dew, Lydia M. S. Lau, Andrew R. Rickard, Jenny C. Young, Tahir Farooq, Michael J. Pilling, and Chris J. Martin. Architecture design of a userorientated electronic laboratory notebook: a case study within an atmospheric chemistry community. Future Generation Computer Systems, 29(8):2182–2196, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1300071X.

Zhao:2019:IOS

[ZDL^{+}19] Yibing Zhao, Feng Ding, Jining Li, Lie Guo, and Wanfeng Qi. The intelligent obstacle sensing and recognizing method based on D-S evidence theory for UGV. Future Generation Computer Systems, 97(??): 21–29, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18329911.

Zhao:2019:LCL

[ZDM^{+}19] Zitian Zhao, Hancong Duan, Geyong Min, Yue Wu, Zilei Huang, Xian Zhuang, Hao Xi, and Meirong Fu. A lighten CNN-LSTM model for speaker verification on embedded devices. Future Generation Computer Systems, 100(??):751–758, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18330620.

Zhou:2016:IBP

[ZDW^{+}16] Yunya Zhou, Hua Deng, Qianhong Wu, Bo Qin, Jianwei Liu, and Yong Ding. Identity-based proxy re-encryption version 2: Making mobile access easy in cloud. Future Generation Computer Systems, 62(??):128–139, September 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003076.

Zou:2018:MUA

[ZDW^{+}18] Caifeng Zou, Huifang Deng, Jiafu Wan, Zhongren Wang, and Pan Deng. Mining and updating association rules based on fuzzy concept lattice. Future Generation Computer Systems, 82(??):698–706, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17312451.

Zadin:2016:NBI

[ZF16] Abedalmotaleb Zadin and Thomas Fevens. Neighborhoodbased interference minimization for stable positionbased routing in mobile ad hoc networks. Future Generation Computer Systems, 64(??):88–97, November 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300668.

Zhong:2017:LES

[ZFC17] Hong Zhong, Yaming Fang, and Jie Cui. LBBSRT: an efficient SDN load balancing scheme based on server response time. Future Generation Computer Systems, 68(??):183–190, March 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303727.

Zhong:2018:RLE

[ZFC18] Hong Zhong, Yaming Fang, and Jie Cui. Reprint of “LBBSRT: an efficient SDN load balancing scheme based on server response time”. Future Generation Computer Systems, 80(??):409–416, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17325475.

Zhou:2018:SAE

[ZFH^{+}18] Yukun Zhou, Dan Feng, Yu Hua, Wen Xia, Min Fu, Fangting Huang, and Yucheng Zhang. A similarity-aware encrypted deduplication scheme with flexible access control in the cloud. Future Generation Computer Systems, 84(??):177–189, July 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17309238.

Zhang:2018:EFG

[ZFS^{+}18] Liang Zhang, Yuanyuan Feng, Peiyi Shen, Guangming Zhu, Wei Wei, Juan Song, Syed Afaq Ali Shah, and Mohammed Bennamoun. Efficient finer-grained incremental processing with MapReduce for big data. Future Generation Computer Systems, 80(??):102–111, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17303138.

Zhang:2014:CMD

[ZFW14] Quan Zhang, Dan Feng, and Fang Wang. Courier: Multi-dimensional QoS guarantees for the consolidated storage system. Future Generation Computer Systems, 37(??):97– 107, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001271.

Zhang:2018:LFV

[ZFY18] Fei Zhang, Xiaoming Fu, and Ramin Yahyapour. LayerMover: Fast virtual machine migration over WAN with three-layer image structure. Future Generation Computer Systems, 83(??):37–49, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630797X.

Zhang:2018:SMS

[ZG18] Zhiyong Zhang and Brij B. Gupta. Social media security and trustworthiness: Overview and new direction. Future Generation Computer Systems, 86(??):914–925, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16303879.

Zakarya:2019:MEP

[ZG19] Muhammad Zakarya and Lee Gillam. Managing energy, performance and cost in large scale heterogeneous datacenters using migrations. Future Generation Computer Systems, 93(??):529–547, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17326249.

Zahid:2017:ENI

[ZGB^{+}17] Feroz Zahid, Ernst Gunnar Gran, Bartosz Bogdański, Bjørn Dag Johnsen, and Tor Skeie. Efficient network isolation and load balancing in multi-tenant HPC clusters. Future Generation Computer Systems, 72(??):145–162, July 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16300735.

Zhang:2018:LAR

[ZGL^{+}18] Feifei Zhang, Jidong Ge, Zhongjin Li, Chuanyi Li, Chifong Wong, Li Kong, Bin Luo, and Victor Chang. A load-aware resource allocation and task scheduling for the emerging cloudlet system. Future Generation Computer Systems, 87(??):438–456, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17316771.

Zeng:2019:CLS

[ZGL19] Yue Zeng, Songtao Guo, and Guiyan Liu. Comprehensive link sharing avoidance and switch aggregation for software-defined data center networks. Future Generation Computer Systems, 91(??):25–36, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1830832X.

Zhou:2013:ARS

[ZGS^{+}13] Zhangbing Zhou, Walid Gaaloul, Lei Shu, Samir Tata, and Sami Bhiri. Assessing the replaceability of service protocols in mediated service interactions. Future Generation Computer Systems, 29(1):287– 299, January 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X11001464.

Zhang:2019:ADW

[ZGV19] James Zhang, Robert Gardner, and Ilija Vukotic. Anomaly detection in wide area network meshes using two machine learning algorithms. Future Generation Computer Systems, 93(??):418–426, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18302267.

Zhang:2010:CRU

[ZGZ^{+}10] Daqiang Zhang, Minyi Guo, Jingyu Zhou, Dazhou Kang, and Jiannong Cao. Context reasoning using extended evidence theory in pervasive computing environments. Future Generation Computer Systems, 26(2):207–216, February 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhuge:2017:AMC

[ZH17] Hai Zhuge and Lei He.

Automatic maintenance of category hierarchy. Future Generation Computer Systems, 67(??):1–12, February 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302710.

Zhang:2017:NAV

[ZHHC17] Weizhe Zhang, Shuo Han, Hui He, and Huixiang Chen. Network-aware virtual machine migration in an overcommitted cloud. Future Generation Computer Systems, 76(??):428– 442, November 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1630053X.

Zhou:2018:PAC

[ZHHQ18] Yi Zhou, Fazhi He, Neng Hou, and Yimin Qiu. Parallel ant colony optimization on multi-core SIMD CPUs. Future Generation Computer Systems, 79 (part 2)(??):473– 487, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16304289.

Zhang:2018:IBR

[ZHL^{+}18] Junbao Zhang, Haojun Huang, Yan Luo, Yinting Fan, and Guan Yang. Immunization-based redundancy elimination in Mobile Opportunistic Networks-Generated big data. Future Generation Computer Systems, 79 (part 3)(??):920– 927, February 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17311147.

Zhuge:2010:SSS

[Zhu10] Hai Zhuge. Special section: Semantic Link Network. Future Generation Computer Systems, 26(3): 359–360, March 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhuge:2014:CPS

[Zhu14] Hai Zhuge. Cyber-Physical Society — the science and engineering for future society. Future Generation Computer Systems, 32(??): 180–186, March 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002185.

Zhu:2018:IBD

[Zhu18] Dingju Zhu. IOT and big data based cooperative logistical delivery scheduling method and cloud robot system. Future Generation Computer Systems, 86(??):709–715, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18301924.

Zhang:2019:MOS

[ZHW19] Hong Zhang, Hai Huang, and Liqiang Wang. Meteor: Optimizing Spark-on-Yarn for short applications. Future Generation Computer Systems, 101(??):262–271, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18311373.

Zineddine:2018:OSQ

[Zin18] Mhamed Zineddine. Optimizing security and quality of service in a realtime operating system using multi-objective bat algorithm. Future Generation Computer Systems, 87(??):102–114, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17309597.

Zhou:2014:DDA

[ZJW^{+}14] Wei Zhou, Weijia Jia, Sheng Wen, Yang Xiang, and Wanlei Zhou. Detection and defense of application-layer DDoS attacks in backbone web traffic. Future Generation Computer Systems, 38(??):36–46, September 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X13001672.

Zissis:2012:ACC

[ZL12] Dimitrios Zissis and Dimitrios Lekkas. Addressing cloud computing security issues. Future Generation Computer Systems, 28(3): 583–592, March 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X10002554.

Zhang:2013:DAB

[ZL13] Yiming Zhang and Ling Liu. Distance-aware bloom filters: Enabling collaborative search for efficient resource discovery. Future Generation Computer Systems, 29(6):1621–1630, August 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1200163X.

Zhu:2018:ARR

[ZL18] Qian Zhu and Yazhuo Li. Agricultural research recommendation algorithm based on consumer preference model of e-commerce. Future Generation Computer Systems, 88(??):151– 155, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18307039.

Zheng:2014:ABD

[ZLG^{+}14] Long Zheng, Yanchao Lu, Minyi Guo, Song Guo, and Cheng-Zhong Xu. Architecture-based design and optimization of genetic algorithms on multiand many-core systems. Future Generation Computer Systems, 38(??):75– 91, September 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002082.

Zheng:2016:VMC

[ZLL^{+}16] Qinghua Zheng, Rui Li, Xiuqi Li, Nazaraf Shah, Jianke Zhang, Feng Tian, Kuo-Ming Chao, and Jia Li. Virtual machine consolidated placement based on multi-objective biogeography-based optimization. Future Generation Computer Systems, 54(??):95–122, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000564.

Zhang:2017:APO

[ZLL17a] Shiwen Zhang, Qin Liu, and Yaping Lin. Anonymizing popularity in online social networks with full utility. Future Generation Computer Systems, 72(??): 227–238, July 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301170.

Zhu:2017:RSM

[ZLL^{+}17b] Xiao Zhu, Duo Liu, Liang Liang, Kan Zhong, Linbo Long, Meikang Qiu, Zili Shao, and Edwin H.-M. Sha. Revisiting swapping in mobile systems with SwapBench. Future Generation Computer Systems, 74(??):265–275, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16301376.

Zhang:2019:ELG

[ZLL^{+}19] Lingyue Zhang, Huilin Li, Yannan Li, Yong Yu, Man Ho Au, and Baocang Wang. An efficient linkable group signature for payer tracing in anonymous cryptocurrencies. Future Generation Computer Systems, 101(??):29–38, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19303115.

Zhao:2015:ESS

[ZLR^{+}15] Yong Zhao, Youfu Li, Ioan Raicu, Shiyong Lu, Wenhong Tian, and Heng Liu. Enabling scalable scientific workflow management in the Cloud. Future Generation Computer Systems, 46 (??):3–16, May 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002179.

Zhang:2019:CSA

[ZLT^{+}19] Shaobo Zhang, Xiong Li, Zhiyuan Tan, Tao Peng, and Guojun Wang. A caching and spatial Kanonymity driven privacy enhancement scheme in continuous location-based services. Future Generation Computer Systems, 94 (??):40–50, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17321398.

Zhang:2010:MPO

[ZLTY10] Junwei Zhang, Bu-Sung Lee, Xueyan Tang, and Chai-Kiat Yeo. A model to predict the optimal performance of the Hierarchical Data Grid. Future Generation Computer Systems, 26(1):1–11, January 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhu:2018:ADP

[ZLXZ18] Tianqing Zhu, Gang Li, Ping Xiong, and Wanlei Zhou. Answering differentially private queries for continual datasets release. Future Generation Computer Systems, 87(??):816–827, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17300183.

Zhou:2019:LIB

[ZLY^{+}19] Lu Zhou, Xiong Li, Kuo-Hui Yeh, Chunhua Su, and Wayne Chiu. Lightweight IoT-based authentication scheme in cloud computing circumstance. Future Generation Computer Systems, 91(??):244–251, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18307878.

Zhang:2013:ASD

[ZLZ13] Youhui Zhang, Yanhua Li, and Weimin Zheng. Automatic software deployment using user-level virtualization for cloud-computing. Future Generation Computer Systems, 29(1):323– 329, January 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001518.

Zhang:2015:CDA

[ZME^{+}15] Fan Zhang, Qutaibah M.

Malluhi, Tamer Elsayed, Samee U. Khan, Keqin Li, and Albert Y. Zomaya. CloudFlow: a data-aware programming model for cloud workflow applications on modern HPC systems. Future Generation Computer Systems, 51(??): 98–110, October 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002222.

Ziegeldorf:2018:SAD

[ZMH^{+}18] Jan Henrik Ziegeldorf, Roman Matzutt, Martin Henze, Fred Grossmann, and Klaus Wehrle. Secure and anonymous decentralized Bitcoin mixing. Future Generation Computer Systems, 80(??):448– 466, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16301297.

Ziemke:2019:BTI

[ZMN19] Dominik Ziemke, Simon Metzler, and Kai Nagel. Bicycle traffic and its interaction with motorized traffic in an agent-based transport simulation framework. Future Generation Computer Systems, 97(??):30– 40, August 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17320447.

Zamboulis:2010:QPE

[ZMP10] Lucas Zamboulis, Nigel Martin, and Alexandra Poulovassilis. Query performance evaluation of an architecture for finegrained integration of heterogeneous Grid data sources. Future Generation Computer Systems, 26(8): 1073–1091, October 2010.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zeng:2016:EAT

[ZMTT16] Gang Zeng, Yutaka Matsubara, Hiroyuki Tomiyama, and Hiroaki Takada. Energyaware task migration for multiprocessor real-time systems. Future Generation Computer Systems, 56(??):220–228, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15002319.

Zhang:2019:SAP

[ZMZ^{+}19] Xingwei Zhang, Rui Mo, Huijun Zhao, Xi Luo, and Yunsheng Yang. Statistical analysis of photodynamic therapy and stent drainage for unresectable cholangiocarcinoma. Future Generation Computer Systems, 91(??):511–517, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18318296.

See retraction notice [ZMZ^{+}20].

Zhang:2020:RNS

[ZMZ^{+}20] Xingwei Zhang, Rui Mo, Huijun Zhao, Xi Luo, and Yunsheng Yang. Retraction notice to “Statistical analysis of photodynamic therapy and stent drainage for unresectable cholangiocarcinoma” [Future Gener. Comput. Syst. 91 (2019) 511–517]. Future Generation Computer Systems, 107(??):1145, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306397.

See [ZMZ^{+}19].

Zarakovitis:2012:PCS

[ZN12] Charilaos C. Zarakovitis and Qiang Ni. A performance comparative study on the implementation methods for OFDMA cross-layer optimization. Future Generation Computer Systems, 28(6):923– 929, June 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001932.

Zhao:2018:IOG

[ZNC^{+}18] Qingjuan Zhao, Jianwei Niu, Huan Chen, Lei Wang, and Mohammed Atiquzzaman. An indicative opinion generation model for short texts on social networks. Future Generation Computer Systems, 86(??):1471–1480, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17310245.

Zambrano:2017:TIT

[ZPPE17] A. M. Zambrano, I. Perez, C. Palau, and M. Esteve. Technologies of Internet of Things applied to an Earthquake Early Warning System. Future Generation Computer Systems, 75(??):206–215, October 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16303909.

Zheng:2018:COD

[ZQB^{+}18] Wei Zheng, Yingsheng Qin, Emmanuel Bugingo, Dongzhan Zhang, and Jinjun Chen. Cost optimization for deadline-aware scheduling of big-data processing jobs on clouds. Future Generation Computer Systems, 82(??):244–255, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17317296.

Zhu:2014:EPP

[ZRZ^{+}14] Tianqing Zhu, Yongli Ren, Wanlei Zhou, Jia Rong, and Ping Xiong. An effective privacy preserving algorithm for neighborhoodbased collaborative filtering. Future Generation Computer Systems, 36(??): 142–155, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001647.

Zang:2018:CMB

[ZRZL18] Wenke Zang, Liyan Ren, Wenqian Zhang, and Xiyu Liu. A cloud model based DNA genetic algorithm for numerical optimization problems. Future Generation Computer Systems, 81(??):465–477, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17304697.

Zhang:2019:SSA

[ZRZR19] Yaocheng Zhang, Wei Ren, Tianqing Zhu, and Yi Ren. SaaS: a situational awareness and analysis system for massive Android malware detection. Future Generation Computer Systems, 95(??):548–559, June 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18309427.

Zhuge:2010:STS

[ZS10] Hai Zhuge and Yunchuan Sun. The schema theory for semantic link network. Future Generation Computer Systems, 26(3): 408–420, March 2010. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhuge:2016:SKG

[ZS16] Hai Zhuge and Xiaoping Sun. Semantics, knowledge and grids on Big Data. Future Generation Computer Systems, 64(??):163– 164, November 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302114.

Zvara:2019:ODD

[ZSBB19] Zoltán Zvara, Péter G. N.

Szabó, Barnabás Balázs, and András Benczúr. Optimizing distributed data stream processing by tracing. Future Generation Computer Systems, 90(??):578–591, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17325141.

Zhang:2019:DBR

[ZSFZ19] Yunyi Zhang, Zhan Shi, Dan Feng, and Xiu-Xiu Zhan. Degree-biased random walk for large-scale network embedding. Future Generation Computer Systems, 100(??):198–209, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19300378.

Zhao:2019:SSD

[ZSGJ19] Feng Zhao, Yu Shen, Xiangyu Gui, and Hai Jin. SDBPR: Social distanceaware Bayesian personalized ranking for recommendation. Future Generation Computer Systems, 95(??): 372–381, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18320478.

Zhu:2012:SSG

[ZSH12] Rongbo Zhu, Zhili Sun, and Jiankun Hu. Special section: Green computing. Future Generation Computer Systems, 28(2):368– 370, February 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X11001221.

Zareie:2019:INR

[ZSJ19] Ahmad Zareie, Amir Sheikhahmadi, and Mahdi Jalili.

Influential node ranking in social networks based on neighborhood diversity. Future Generation Computer Systems, 94(??):120– 129, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18319009.

Zhao:2019:MAS

[ZSL^{+}19a] Dongcheng Zhao, Gang Sun, Dan Liao, Shizhong Xu, and Victor Chang. Mobile-aware service function chain migration in cloud-fog computing. Future Generation Computer Systems, 96(??):591–604, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18325329.

Zhou:2019:AFG

[ZSL^{+}19b] Lu Zhou, Chunhua Su, Zhen Li, Zhe Liu, and Gerhard P. Hancke. Automatic fine-grained access control in SCADA by machine learning. Future Generation Computer Systems, 93(??):548–559, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17325761.

Zheng:2018:GDP

[ZSMS18] Zhigao Zheng, Nitin Saxena, K. K. Mishra, and Arun Kumar Sangaiah. Guided dynamic particle swarm optimization for optimizing digital image watermarking in industry applications. Future Generation Computer Systems, 88(??):92–106, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18307945.

Zhou:2017:KIC

[ZSP17] Qunzhi Zhou, Yogesh Simmhan, and Viktor Prasanna. Knowledgeinfused and consistent Complex Event Processing over real-time and persistent streams. Future Generation Computer Systems, 76(??):391–406, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16304769.

Zhao:2019:CSM

[ZSQ^{+}19] Linchang Zhao, Zhaowei Shang, Anyong Qin, Taiping Zhang, Ling Zhao, Yu Wei, and Yuan Yan Tang. A cost-sensitive meta-learning classifier: SPFCNN-Miner. Future Generation Computer Systems, 100(??):1031–1043, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301876.

Zhou:2018:SHT

[ZSS^{+}18] Lu Zhou, Chunhua Su, Xin Sun, Xishun Zhao, and Kim-Kwang Raymond Choo. Stag hunt and trust emergence in social networks. Future Generation Computer Systems, 88(??):168–172, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17327206.

Zhou:2018:TPW

[ZSW^{+}18a] Lu Zhou, Chunhua Su, Yamin Wen, Weijie Li, and Zheng Gong. Towards practical whitebox lightweight block cipher implementations for IoTs. Future Generation Computer Systems, 86(??):507–514, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17326705.

Zuo:2018:CSA

[ZSW^{+}18b] Cong Zuo, Jun Shao, Guiyi Wei, Mande Xie, and Min Ji. CCA-secure ABE with outsourced decryption for fog computing. Future Generation Computer Systems, 78 (part 2)(??):730– 738, January 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16304745.

Zhang:2015:PHT

[ZSX^{+}15] Tao Zhang, XiangZheng Sun, Wei Xue, Nan Qiao, Huang Huang, JiWu Shu, and Weimin Zheng. ParSA: High-throughput scientific data analysis framework with distributed file system. Future Generation Computer Systems, 51(??):111–119, October 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X14002088.

Zhang:2014:DBS

[ZSZ14] Li Zhang, Xiaoping Sun, and Hai Zhuge. Densitybased spatial keyword querying. Future Generation Computer Systems, 32(??):211–221, March 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000356.

Zhu:2019:DCW

[ZT19] Zhaomeng Zhu and Xueyan Tang. Deadline-constrained workflow scheduling in IaaS clouds with multiresource packing. Future Generation Computer Systems, 101(??):880–893, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1832363X.

Zhan:2019:HRS

[ZTC^{+}19] Yan Zhan, Pandu R.

Tadikamalla, James A. Craft, Jiansha Lu, Jijun Yuan, Zhi Pei, and Shiyun Li. Human reliability study on the door operation from the view of deep machine learning. Future Generation Computer Systems, 99(??):143–153, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326098.

Zhou:2018:TTA

[ZTD^{+}18] Yi Zhou, Shubbhi Taneja, Gautam Dudeja, Xiao Qin, Jifu Zhang, Minghua Jiang, and Mohammed I.

Alghamdi. Towards thermalaware Hadoop clusters. Future Generation Computer Systems, 88(??):40– 54, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17301188.

Zhang:2017:SED

[ZTKF17] Yi-Fan Zhang, Yu-Chu Tian, Wayne Kelly, and Colin Fidge. Scalable and efficient data distribution for distributed computing of all-to-all comparison problems. Future Generation Computer Systems, 67(??):152–162, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302825.

Zhang:2019:SAM

[ZTKX19] Fuquan Zhang, Jeyarajan Thiyagalingam, Thia Kirubarajan, and Shuwen Xu. Speed-adaptive multicopy routing for vehicular delay tolerant networks. Future Generation Computer Systems, 94(??):392– 407, May 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314122.

Zhang:2019:QCE

[ZTL^{+}19] Fan Zhang, Xuxin Tang, Xiu Li, Samee U. Khan, and Zhijiang Li. Quantifying cloud elasticity with container-based autoscaling. Future Generation Computer Systems, 98(??):672–681, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18307842.

Zhu:2010:CCC

[ZW10] Yuqing Zhu and Jianmin Wang. Client-centric consistency formalization and verification for system with large-scale distributed data storage. Future Generation Computer Systems, 26(8): 1180–1188, October 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhang:2018:MEE

[ZWDP18] Liang Zhang, Wenji Wu, Phil DeMar, and Eric Pouyoul. mdtmFTP and its evaluation on ESNET SDN testbed. Future Generation Computer Systems, 79 (part 1)(??):199– 204, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17306489.

Zhu:2019:CTB

[ZWGC19] Liehuang Zhu, Yulu Wu, Keke Gai, and Kim-Kwang Raymond Choo. Controllable and trustworthy blockchain-based cloud data management. Future Generation Computer Systems, 91(??):527–535, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18311993.

Zhang:2017:CBV

[ZWHC17] Jiangtao Zhang, Xuan Wang, Hejiao Huang, and Shi Chen. Clustering based virtual machines placement in distributed cloud computing. Future Generation Computer Systems, 66(??): 1–10, January 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16302035.

Zhu:2018:EHM

[ZWJ^{+}18] Longting Zhu, Jigang Wu, Guiyuan Jiang, Long Chen, and Siew-Kei Lam. Efficient hybrid multicast approach in wireless data center network. Future Generation Computer Systems, 83(??):27– 36, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17311160.

Zeng:2019:TSS

[ZWJ^{+}19a] Peng Zeng, Zhaowei Wang, Zhengyi Jia, Linghe Kong, Dong Li, and Xi Jin. Timeslotted software-defined industrial Ethernet for realtime quality of service in Industry 4.0. Future Generation Computer Systems, 99(??):1–10, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18311427.

Zhou:2019:URI

[ZWJ19b] Xiaokang Zhou, Bo Wu, and Qun Jin. User role identification based on social behavior and networking analysis for information dissemination. Future Generation Computer Systems, 96(??):639–648, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1730818X.

Zhang:2013:OPD

[ZWL13] Chaoyi Zhang, Muqing Wu, and Linlin Luan. An optimal PSO distributed precoding algorithm in QRD-based multirelay system. Future Generation Computer Systems, 29(1):107–113, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X12000933.

Zhang:2016:ESM

[ZWL^{+}16] Jilin Zhang, Jian Wan, Fangfang Li, Jie Mao, Li Zhuang, Junfeng Yuan, Enyi Liu, and Zhuoer Yu. Efficient sparse matrixvector multiplication using cache oblivious extension quadtree storage format. Future Generation Computer Systems, 54(??):490– 500, January 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15000606.

Zuo:2019:LBN

[ZWMC19] Yuan Zuo, Yulei Wu, Geyong Min, and Laizhong Cui. Learning-based network path planning for traffic engineering. Future Generation Computer Systems, 92(??):59–67, March 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18313244.

Zhang:2019:IIP

[ZWQ^{+}19] Yin Zhang, Haoyu Wen, Feier Qiu, Zie Wang, and Haider Abbas. iBike: Intelligent public bicycle services assisted by data analytics. Future Generation Computer Systems, 95(??): 187–197, June 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18322787.

Zhao:2012:PAP

[ZWS^{+}12] Baokang Zhao, Dan Wang, Zili Shao, Jiannong Cao, and Jinshu Su. Privacy aware publishing of successive location information in sensor networks. Future Generation Computer Systems, 28(6):913– 922, June 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1100166X.

Zhang:2013:TTD

[ZWW^{+}13] Shengzhi Zhang, Wenjie Wang, Haishan Wu, Athanasios V. Vasilakos, and Peng Liu. Towards transparent and distributed workload management for large scale Web servers. Future Generation Computer Systems, 29(4): 913–925, June 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001872.

Zhang:2018:SAC

[ZWWL18] Shunxiang Zhang, Zhongliang Wei, Yin Wang, and Tao Liao. Sentiment analysis of Chinese micro-blog text based on extended sentiment dictionary. Future Generation Computer Systems, 81(??):395–403, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17307835.

Zhong:2019:SVL

[ZWX^{+}19] Lin Zhong, Qianhong Wu, Jan Xie, Jin Li, and Bo Qin. A secure versatile light payment system based on blockchain. Future Generation Computer Systems, 93(??):327– 337, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18302656.

Zhou:2019:STC

[ZWYH19] Qimin Zhou, Hao Wu, Kun Yue, and Ching-Hsien Hsu. Spatio-temporal context-aware collaborative QoS prediction. Future Generation Computer Systems, 100(??):46–57, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18332473.

Zhang:2018:ORQ

[ZWZ18] Bin Zhang, Xiaoyang Wang, and Zhigao Zheng. The optimization for recurring queries in big data analysis system with MapReduce. Future Generation Computer Systems, 87(??):549–556, October 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17310208.

Zhao:2019:NAS

[ZWZ19] Weiguo Zhao, Liying Wang, and Zhenxing Zhang. A novel atom search optimization for dispersion coefficient estimation in groundwater. Future Generation Computer Systems, 91(??):601–610, February 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18306575.

Zhao:2019:SCT

[ZXD^{+}19] Xiaogang Zhao, Zenggang Xiong, Ling Ding, Xuemin Zhang, and Fang Xu. A smart coordinated temperature feedback controller for energyefficient data centers. Future Generation Computer Systems, 93(??):506–514, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18321599.

Zhang:2014:EFH

[ZXJ^{+}14] Xiaojun Zhang, Chunxiang Xu, Chunhua Jin, Run Xie, and Jining Zhao. Efficient fully homomorphic encryption from RLWE with an extension to a threshold encryption scheme. Future Generation Computer Systems, 36(??):180–186, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13002422.

Zou:2014:CSL

[ZXL14] Yinlong Zou, Wei Xue, and Shenshen Liu. A case study of large-scale parallel I/O analysis and optimization for numerical weather prediction system. Future Generation Computer Systems, 37(??):378– 389, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14000053.

Zhang:2018:LBW

[ZXL^{+}18] Yang Zhang, Zuocheng Xing, Cang Liu, Chuan Tang, and Qinglin Wang. Locality based warp scheduling in GPGPUs. Future Generation Computer Systems, 82(??):520–527, May 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17302728.

Zhang:2019:CRF

[ZXM^{+}19] Hanqi Zhang, Xi Xiao, Francesco Mercaldo, Shiguang Ni, Fabio Martinelli, and Arun Kumar Sangaiah. Classification of ransomware families with machine learning based on N-gram of opcodes. Future Generation Computer Systems, 90(??):211–221, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18307325.

Zhang:2018:AKE

[ZXW^{+}18] Yuexin Zhang, Yang Xiang, Tao Wang, Wei Wu, and Jian Shen. An over-the-air key establishment protocol using keyless cryptography. Future Generation Computer Systems, 79 (part 1)(??):284– 294, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1630752X.

Zeng:2019:SMP

[ZXW19] Guosun Zeng, Huanliang Xiong, and Canghai Wu. A scalable method of parallel tasks after the extension of machine systems based on equal change rate. Future Generation Computer Systems, 101(??):680–693, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19300056.

Zhang:2018:VPA

[ZXWA18] Yuexin Zhang, Yang Xiang, Wei Wu, and Abdulhameed Alelaiwi. A variant of password authenticated key exchange protocol. Future Generation Computer Systems, 78 (part 2)(??):699–711, January 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302145.

Zhang:2018:OAM

[ZXZL18] Jixian Zhang, Ning Xie, Xuejie Zhang, and Weidong Li. An online auction mechanism for cloud computing resource allocation and pricing based on user evaluation and cost. Future Generation Computer Systems, 89(??):286– 299, December 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17317983.

Zikria:2018:ITI

[ZYA^{+}18] Yousaf Bin Zikria, Heejung Yu, Muhammad Khalil Afzal, Mubashir Husain Rehmani, and Oliver Hahm. Internet of Things (IoT): Operating system, applications and protocols design, and validation techniques. Future Generation Computer Systems, 88(??):699–706, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317710.

Zhang:2018:BAT

[ZYB^{+}18] Dongzhan Zhang, Wenjing Yan, Emmanuel Bugingo, Wei Zheng, and Jinjun Chen. A benchmark approach and its toolkit for online scheduling of multiple deadline-constrained workflows in big-data processing systems. Future Generation Computer Systems, 85(??):222–234, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17323531.

Zhao:2019:DLB

[ZYC^{+}19] Xianlong Zhao, Kexin Yang, Qimei Chen, Duo Peng, Hao Jiang, Xianze Xu, and Xinzhuo Shuang. Deep learning based mobile data offloading in mobile edge computing systems. Future Generation Computer Systems, 99(??):346– 355, October 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304406.

Zhang:2019:HEA

[ZYCZ19] Junhua Zhang, Dong Yuan, Lizhen Cui, and Bing Bing Zhou. A highly efficient algorithm towards optimal data storage and regeneration cost in multiple clouds. Future Generation Computer Systems, 99(??):459–472, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18322295.

Zhang:2017:PPT

[ZYK17] Lifang Zhang, Zheng Yan, and Raimo Kantola. Privacypreserving trust management for unwanted traffic control. Future Generation Computer Systems, 72(??): 305–318, July 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16302308.

Zhu:2015:HPF

[ZYTC15] Feng Zhu, Yiping Yao, Wenjie Tang, and Dan Chen. A high performance framework for modeling and simulation of largescale complex systems. Future Generation Computer Systems, 51(??):132–141, October 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002520.

Zhang:2018:SPA

[ZYW^{+}18] Mingchuan Zhang, Meiyi Yang, Qingtao Wu, Ruijuan Zheng, and Junlong Zhu. Smart perception and autonomic optimization: a novel bioinspired hybrid routing protocol for MANETs. Future Generation Computer Systems, 81(??):505–513, April 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17304545.

Zhan:2018:HPV

[ZYZ^{+}18] Dongyang Zhan, Lin Ye, Hongli Zhang, Binxing Fang, Huhua Li, Yang Liu, Xiaojiang Du, and Mohsen Guizani. A highperformance virtual machine filesystem monitor in cloud-assisted cognitive IoT. Future Generation Computer Systems, 88(??):209–219, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18304394.

Zhang:2015:ITS

[ZZ15] Jie Zhang and Futai Zhang. Information-theoretical secure verifiable secret sharing with vector space access structures over bilinear groups and its applications. Future Generation Computer Systems, 52(??):109–115, November 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X14002477.

Zhang:2019:TMM

[ZZ19] Cheng Zhang and Zixuan Zheng. Task migration for mobile edge computing using deep reinforcement learning. Future Generation Computer Systems, 96(??):111–118, July 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18329674.

Zhang:2019:CPR

[ZZBP19] Huibing Zhang, Hao Zhong, Weihua Bai, and Fang Pan. Cross-platform rating prediction method based on review topic. Future Generation Computer Systems, 101(??):236–245, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308660.

Zheng:2019:TGP

[ZZBZ19] Wei Zheng, Chengxin Zhang, Eric W. Bell, and Yang Zhang. I-TASSER gateway: a protein structure and function prediction server powered by XSEDE. Future Generation Computer Systems, 99(??):73–85, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18314705.

Zhou:2014:PBM

[ZZC14] Jingyu Zhou, Yunlong Zhang, and Jia Cheng. Preference-based mining of top-K influential nodes in social networks. Future Generation Computer Systems, 31(??):40–47, February 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X12001471.

Zhou:2018:ESH

[ZZC18] Lu Zhou, Youwen Zhu, and Kim-Kwang Raymond Choo. Efficiently and securely harnessing cloud to solve linear regression and other matrix operations. Future Generation Computer Systems, 81(??):404– 413, April 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17308075.

Zamani:2018:CMS

[ZZDM^{+}18] Ali Reza Zamani, Mengsong Zou, Javier Diaz-Montes, Ioan Petri, Omer Rana, and Manish Parashar.

A computational model to support in-network data analysis in federated ecosystems. Future Generation Computer Systems, 80(??):342–354, March 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310580.

Zhang:2018:STI

[ZZF18] PeiYun Zhang, MengChu Zhou, and Giancarlo Fortino. Security and trust issues in fog computing: a survey. Future Generation Computer Systems, 88(??):16– 27, November 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17329722.

Zhai:2019:EBS

[ZZF^{+}19] Xuemeng Zhai, Wanlei Zhou, Gaolei Fei, Cai Lu, Sheng Wen, and Guangmin Hu. Edge-based stochastic network model reveals structural complexity of edges. Future Generation Computer Systems, 100 (??):1073–1087, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301530.

Zhang:2016:KBD

[ZZH^{+}16] Jiangtao Zhang, Lingmin Zhang, Hejiao Huang, Zeo L. Jiang, and Xuan Wang. Key based data analytics across data centers considering bi-level resource provision in cloud computing. Future Generation Computer Systems, 62(??):40–50, September 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16300528.

Zhang:2018:MCD

[ZZH^{+}18] Bo Zhang, Qian Zhang, Zhenhua Huang, Meizi Li, and Luqun Li. A multicriteria detection scheme of collusive fraud organization for reputation aggregation in social networks. Future Generation Computer Systems, 79 (part 3)(??):797–814, February 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16308366.

Zhang:2017:CMV

[ZZJ17] Dalu Zhang, Dejiang Zhou, and Xiang Jin. Chunk mode VM migration in XIA and triple-way pipeline for performance optimization. Future Generation Computer Systems, 74(??):32–40, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17306660.

Zhao:2016:PHR

[ZZJY16] Feng Zhao, Yajun Zhu, Hai Jin, and Laurence T. Yang. A personalized hashtag recommendation approach using LDA-based topic model in microblog environment. Future Generation Computer Systems, 65(??):196– 206, December 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X15003258.

Zou:2010:CTV

[ZZL^{+}10] Deqing Zou, Weide Zheng, Jinjiu Long, Hai Jin, and Xueguang Chen. Constructing trusted virtual execution environment in P2P grids. Future Generation Computer Systems, 26(5):769–775, May 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhou:2018:EAC

[ZZLH18] Zhangbing Zhou, Deng Zhao, Lu Liu, and Patrick C. K. Hung. Energy-aware composition for wireless sensor networks as a service. Future Generation Computer Systems, 80(??): 299–310, March 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17303266.

Zhou:2018:CAS

[ZZLR18] Xiaokang Zhou, Albert Y. Zomaya, Weimin Li, and Ivan Ruchkin. Cybermatics: Advanced strategy and technology for cyberenabled systems and applications. Future Generation Computer Systems, 79 (part 1)(??):350– 353, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17321428.

Zhou:2018:DSH

[ZZLZ18] Ke Zhou, Jiangfeng Zeng, Yu Liu, and Fuhao Zou. Deep sentiment hashing for text retrieval in social CIoT. Future Generation Computer Systems, 86(??):362–371, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18302383.

Zou:2013:DIT

[ZZQ^{+}13] Deqing Zou, Wenrong Zhang, Weizhong Qiang, Guofu Xiang, Laurence Tianruo Yang, Hai Jin, and Kan Hu. Design and implementation of a trusted monitoring framework for cloud platforms. Future Generation Computer Systems, 29(8):2092–2102, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000046.

Zhou:2019:MCM

[ZZS^{+}19] Xiumin Zhou, Gongxuan Zhang, Jin Sun, Junlong Zhou, Tongquan Wei, and Shiyan Hu. Minimizing cost and makespan for workflow scheduling in cloud using fuzzy dominance sort based HEFT. Future Generation Computer Systems, 93(??):278– 289, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18314080.

Zhang:2018:ELM

[ZZSZ18] Wei Zhang, Xinchang Zhang, Huiling Shi, and Longquan Zhou. An efficient latency monitoring scheme in software defined networks. Future Generation Computer Systems, 83(??):303–309, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1732397X.

Zhang:2019:LAL

[ZZX^{+}19] Chuan Zhang, Liehuang Zhu, Chang Xu, Kashif Sharif, Xiaojiang Du, and Mohsen Guizani. LPTD: Achieving lightweight and privacy-preserving truth discovery in CIoT. Future Generation Computer Systems, 90(??):175–184, January 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18304497.

Zhang:2018:PEP

[ZZXL18] Chuan Zhang, Liehuang Zhu, Chang Xu, and Rongxing Lu. PPDP: An efficient and privacypreserving disease prediction scheme in cloud-based e-healthcare system. Future Generation Computer Systems, 79 (part 1)(??): 16–25, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X16304277.

Zhang:2019:LAS

[ZZY^{+}19] Liping Zhang, Lanchao Zhao, Shuijun Yin, Chi-Hung Chi, Ran Liu, and Yixin Zhang. A lightweight authentication scheme with privacy protection for smart grid communications. Future Generation Computer Systems, 100(??):770–778, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310398.

Zhu:2017:TDV

[ZZZ17] Wei Zhu, Yi Zhuang, and Long Zhang. A threedimensional virtual resource scheduling method for energy saving in cloud computing. Future Generation Computer Systems, 69(??):66–74, April 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306586.

Zhang:2019:EPB

[ZZZC19] Xinbo Zhang, Dongzhan Zhang, Wei Zheng, and Jinjun Chen. An enhanced priority-based scheduling heuristic for DAG applications with temporal unpredictability in task execution and data transmission. Future Generation Computer Systems, 100(??):428–439, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18311919.